A Complete Bibliography of Publications in the British Journal for the History of Science: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.65

Title word cross-reference

$105.00 [Web18b]. $119.00 [Gul18]. $125.00 [San15a]. $126.00 [Sør17]. $135.00 [Car18]. $139.00 [Ako19]. $146.00 [Pos14]. $147.00 [Dun14]. $16.00 [Fix16]. $17.99 [Smo10]. $175.00 [Car18]. $18.95 [Mey11]. $21.00 [Ren10]. $21.95 [Ren19]. $22.95 [Ale11, Pis17]. $22.95/£ [Bad19]. $24.95 [LD18, Pha10, Smi12, Sta11d, SM19]. $25.00 [Ayl18, Leh12a, Wil12]. $26.00 [Goo10b]. $26.99/£ [Hol16c]. $27.50 [Lan15b]. $27.95 [Hol16b, Mar17a, Mar18b]. $28.00 [Far12]. $28.95 [Ree16a]. $29.95 [Arn16, Moo17, Wai18]. $30.00 [Boa18, Bod19, Can16, Goo11a, Web15]. $30.00/£ [Ksi16]. $32.95 [Hay19b]. $34.00 [Sac19]. $34.95 [Bow16, Coz16, Gif16, Lou19, May17, Smi19]. $35.00 [Dra17, Far16, Fer11a, Fre10, Gra11, Mac11a, Ric10, Rot16, Sho18, TS14, Van14, Web15, vN19]. $35.00/£ [Smi16a]. $36.00 [Eve18]. $39.00 [Gou12c]. $39.50 [Sle12]. $39.95 [Can19b, Fra17, Mar18a, Rud13]. $39.95/£ [Joh16, Pea19]. $40.00 [Hen17, Prk18, San16, Wad19, Web18a]. $40.00/£ [Jam18, SJ19]. $45.00 [Ayl19, Dun16b, Fin10a, Ins19, Jar17, Kel17, Lee15b, Lou16b, Nal12, Ree19a, Sac16b, Shi16, Sim17a, Tun19, Wal16a]. $45.00/£ [Alb15, Con15, Egm13, Roo15]. $49.00 [Wyl11]. $49.95 [Hen19, Wil11a]. $50.00 [Gil15a, Ins19, Leh12a, Man17, O’C17, Rad14, Sta16a, Tam16b]. $54.95 [Aga17b]. $54.99 [Fix19]. $55.00 [Art11, Bel18b, Cur18, Ebe18, Hem19, Mai12a, Rie16, Sch19, Yal17].

$55.00/EUR40.95 [Bec18c]. $56.70 [Esp10]. $60.00 [Far19, Ser12, Smi11c]. $61.95 [Ayl17]. $65.00 [Ade10, Ber12, Fal18, Fin10a, Kao12, Smi11d, Ste11b]. $65.00/£ [Das18]. $69.95 [Ree15]. $72.00 [Tim12]. $75.00 [Fin10a, War19]. $85.00 [Met11]. _{2} [Boa15].

0-226-66737-9 [Web14]. 0-230-55314-1 [Goo10a]. 0-8264-1868-6 [Ade10].

1-62273-738-5 [Dan19]. 1-84467-442-8 [Kid11a]. 15.00/ [Goo11a]. 150/$215.00 [Cuo12]. 1596 [Eve13]. 16.99/$27.00 [Cal12]. 1610 [Coo11]. 1630 [Tar15]. 17.50/$25.00 [Aga17a]. 17.99 [Osl10]. 1700 [Sma13]. 1713 [Sch15b]. 1730 [Far13, McC19]. 1750 [Wal11b]. 1760 [Ste14b]. 1789 [Wit11b]. 17th [AM11, Fix19, Mar12b]. 18.99 [Tay14]. 1800 [Bui19, Edd08, Rob10a, Sim17a, Smi11d, Tar12, Tee11]. 1820 [Bar17, Gil11]. 1830 [Gil15a, Ksi15]. 1837 [Mca16]. 1840 [Mar18b]. 1840s [McM18]. 1841 [Pis16]. 1848 [Vos11]. 1850 [Bod18, Fin17, Joh15, Köb16, Mor10a, Sta12b, Web14]. 1851 [Can11, Ebe18, Mac13]. 1857 [Kne10]. 1859 [Hoc16a, Les15, Lou16b]. 1860 [Daw12, Far11]. 1860s [Eff17]. 1870s [Toa11, Toa12]. 1871 [Daw15c]. 1875 [Daw16b]. 1880 [Bak15b]. 1880s [Eff17, OS13]. 1881 [Mil11c]. 1883 [Rad13]. 1896 [CGC13a]. 18th [Bak15a, HSSR13]. 19.95 [Rob14]. 1900 [Ric10, Tur16a]. 1901 [Kee12]. 1904 [Jon10b]. 1914 [Ada10, Ber18, Cow10, For10, Han16, Lou15, Sal15, SS15]. 1916 [Bow18]. 1918 [Hoc14, Kaa18]. 1920s [Nav15b, Waz14]. 1926 [Dar11]. 1929-1959 [Szu18]. 1930 [Les13b, Pro11, Ree16b, San17, Sta19c]. 1930s [Pet14]. 1935 [Gli12]. 1939 [Big10, Hug10, Som11]. 1940 [MY18, Pis17]. 1940s [McM18]. 1945 [Fot18, Ham15, Lee15b, vL16]. 1950 [Fin11, GR18, Joh16, Smi18a]. 1950s [Ada16, AG17, Hog12, Jon14a]. 1951 [Gli12]. 1958 [Wol14]. 1960 [Hom16, Lou19, Tho16a]. 1960s [Bin12, Chi08, Kri12, Lan14, Til10, Wis12]. 1970 [CRB09, Ric10]. 1975 [Sto14]. 1980s [Cor12, Mye17]. 19th [GH13, Sia18, Sim14, Sim13]. 19th-20th [Sim13].

2/1 [Chr10]. 2/2 [Chr10]. 2000

[Far10b, Mor18, Ros16, Win15]. 2006 [Ree17]. 2010 [Adl16]. 20th [BFGG11, GH13, Nav16, Sia18, Sim14, Sim13, BS15]. 21.00/$30.00 [Bow17, Ree18]. 22.50 [Val14]. 223-634 [Chr10]. 25.95 [Nav14a]. 28.00/$40.00 [Gil16b]. 29.00 [Cle14]. 2nd [Ren19].

30.00/$45.00 [Nal15]. 30.95 [Smi14]. 32.50 [Gre14b]. 32.95 [Jam18, Tat17b]. 35.00/$50.00 [Whi15]. 39.99 [Hen15b]. 3rd [Kir15, Mol14].

41.50/$60 [Wal19a].

55.00 [Tam14a].

60.00/$90.95 [Jim19]. 60.00/$99.00 [Boa12]. 62 [Ree10]. 64 [Smi11b]. 64.99/$99.99 [Ree17]. 65.00 [Row14]. 65.00/$99.00 [Nav15b].

81-7824-376-8 [Ban14a]. 88 [Ken19].

90.00 [Ric14]. 978-0-19-046322-9 [Mar17a]. 978-0-19-065820-5 [Gou17]. 978-0-19-518750-2 [Pic10b]. 978-0-19-536840-6 [Boo12]. 978-0-19-536841-3 [Boo12]. 978-0-19-537314-1 [Bjö11]. 978-0-19-537389-9 [Opi14a]. 978-0-19-569705-6 [Bha11]. 978-0-19-818313-6 [Eve13].

978-0-19-870305-1 [Kot16]. 978-0-19-870459-1 [Hen15a]. 978-0-19-870480-5 [Bla17]. 978-0-19-870513-0 [Kni16]. 978-0-19-871288-6 [Dih16]. 978-0-19-871470-5 [Mca16]. 978-0-19-871581-8 [Mor15b]. 978-0-19-871648-8 [Kip16]. 978-0-19-871781-4 [Nas16]. 978-0-19-872528-2 [Ste16]. 978-0-19-873677-6 [Reg17]. 978-0-19-874685-0 [Bad16]. 978-0-19-874795-6 [Tys17]. 978-0-19-875763-4 [Pis16]. 978-0-19-879-498-1 [Bix19]. 978-0-19-881617-1 [Tat18]. 978-0-19-921312-2 [Pet10a]. 978-0-19-923484-4 [Mil11b]. 978-0-19-923795-1 [Biv10]. 978-0-19-927820-6 [Tee11]. 978-0-19-954136-2 [Biv10]. 978-0-19-954161-4 [Pic10a]. 978-0-19-954470-7 [Vos11]. 978-0-19-954624-4 [Biv10]. 978-0-19-956096-7 [Biv10]. 978-0-19-956272-5 [Hug10]. 978-0-19-956664-8 [Pyl10]. 978-0-19-956773-7 [Bur11]. 978-0-19-956787-4 [Ful10a]. 978-0-19-956889-5 [Mil11c]. 978-0-19-957431-5 [Mor13]. 978-0-19-958256-3 [Sle11a]. 978-0-19-958352-2 [Val11]. 978-0-19-958458-1 [Wad11]. 978-0-19-958658-5 [Und12]. 978-0-19-958856-5 [Mil11d]. 978-0-19-9588626 [Mil13]. 978-0-19-958977-7 [Lan12]. 978-0-19-959493-1 [Lan11a]. 978-0-19-959667-6 [Mac13]. 978-0-19-959903-5 [Lou17]. 978-0-19-964303-5 [Daw15b]. 978-0-19-964437-7 [Dij12]. 978-0-19-965472-7 [Eva13].

978-0-19-965573-1 [Win15]. 978-0-19-966058-2 [Mar15]. 978-0-19-966396-5 [Hen13a]. 978-0-19-966437-5 [Hen15b]. 978-0-19-966454-2 [Bjö14]. 978-0-19-966926-4 [KK14]. 978-0-19-967526-5 [Kni15]. 978-0-19-967851-8 [Can15]. 978-0-19-968718-3 [Lau15]. 978-0-19-968942-2 [Mil15]. 978-0-19-969491-4 [Mor13]. 978-0-19-969625-3 [Sch15a]. 978-0-19-976649-9 [Gou13b]. 978-0-19-979496-6 [Tay14]. 978-0-19-985711-1 [Nav15a]. 978-0-19-992264-2 [Hen14]. 978-0-19-997815-1 [Nav16]. 978-0-19-998731-3 [Cle16]. 978-0-19-998933-1 [Pos15a]. 978-0-1987-9725-8 [Can19a]. 978-0-1988-3337-6 [HA19b]. 978-0-1992-9979-9 [Man16]. 978-0-1996-5787-2 [SS16]. 978-0-224-06262-6 [Und13]. 978-0-226-00110-4 [Car12]. 978-0-226-01729-7 [Kus15]. 978-0-226-01780-8 [Cap15]. 978-0-226-04601-3 [Ree18]. 978-0-226-04663-1 [Rei15]. 978-0-226-05853-5 [SS13a]. 978-0-226-05893-1 [Lan15b]. 978-0-226-06652-3 [Ste10]. 978-0-226-06863-3 [Boo11]. 978-0-226-06867-1 [San15b]. 978-0-226-07472-6 [Wyl11]. 978-0-226-07840-3 [Ber15]. 978-0-226-07966-0 [Alb15]. 978-0-226-08020-8 [Web15]. 978-0-226-08602-6 [Coz15]. 978-0-226-10656-4 [Roo15]. 978-0-226-10849-0 [Con15]. 978-0-226-11181-0 [Val14]. 978-0-226-12648-7 [Bou15]. 978-0-226-15596-8 [Smi16a].

978-0-226-18420-3 [Tam16b]. 978-0-226-18448-7 [Can16]. 978-0-226-20393-5 [Ksi16]. 978-0-226-24377-1 [Web17]. 978-0-226-24573-7 [Jam17]. 978-0-226-25917-8 [Wil16b]. 978-0-226-26145-4 [Wal16a]. 978-0-226-27554-3 [Bar17]. 978-0-226-27651-9 [Daw12]. 978-0-226-28820-8 [Prk17]. 978-0-226-28879-6 [Shi16]. 978-0-226-28953-3 [Smi10c]. 978-0-226-29322-6 [Rie16]. 978-0-226-30292-8 [Hen17]. 978-0-226-30442-7 [Tur14]. 978-0-226-31720-5 [Rad17]. 978-0-226-31783-0 [Bec12]. 978-0-226-31944-5 [Ebe17]. 978-0-226-33015-0 [Mar16]. 978-0-226-33273-4 [O’C17]. 978-0-226-33998-6 [Gil16b]. 978-0-226-3473205 [Man17]. 978-0-226-35136-0 [Boa18]. 978-0-226-35918-2 [Wil12]. 978-0-226-37291-4 [Aga17a]. 978-0-226-38442-9 [Bow17]. 978-0-226-38859-5 [Kel17]. 978-0-226-40118-8 [Mac11a]. 978-0-226-42953-3 [Lou16b]. 978-0-226-44990-6 [Pro11]. 978-0-226-45486-3 [Web18a]. 978-0-226-46220-2 [Dra17]. 978-0-226-46529-6 [Egm13]. 978-0-226-46568-5 [Mar12c]. 978-0-226-47078-8 [Nal12]. 978-0-226-47114-3 [Rih12]. 978-0-226-48114-2 [Hem19]. 978-0-226-48726-7 [And12a]. 978-0-226-49697-9 [Sch19]. 978-0-226-49806-5 [Stu10]. 978-0-226-50628-9 [Rap12]. 978-0-226-52081-0 [Mae12]. 978-0-226-52228-9 [Bar13b]. 978-0-226-52857-1 [Eve18].

978-0-226-53503-6 [Hol14a]. 978-0-226-538446-4 [vN19]. 978-0-226-54570-7 [Rad14]. 978-0-226-56069-4 [Mai12a]. 978-0-226-67534-3 [Tes11]. 978-0-226-67760-6 [Roo12b]. 978-0-226-68295-2 [San14]. 978-0-226-70096-0 [SS12]. 978-0-226-70711-2 [Mon10]. 978-0-226-71214-7 [Web10]. 978-0-226-72005-0 [Ren10]. 978-0-226-73860-4 [Goo10b]. 978-0-226-74855-9 [She13]. 978-0-226-76783-3 [San12]. 978-0-226-76886-1 [Ste14b]. 978-0-226-81664-7 [Cal13]. 978-0-226-85586-8 [Pin12]. 978-0-226-87182-0 [Coe10]. 978-0-226-89377-8 [Buc11]. 978-0-226-90258-6 [Bal14b]. 978-0-226-90709-3 [Goo12a]. 978-0-226-92199-0 [Boo14]. 978-0-226-92284-3 [Van14]. 978-0-226-92371-0 [Cle14]. 978-0-226-92374-1 [Whi15]. 978-0-226-92414-4 [Gre14b]. 978-0-2260-3898-8 [Wal15]. 978-0-2260-8860-0 [ES16]. 978-0-2260-9216-4 [Whi16]. 978-0-2261-1029-5 [Bel18b]. 978-0-2261-4383-5 [San16]. 978-0-2262-3584-4 [Far16]. 978-0-2262-4766-3 [Sho18]. 978-0-2262-5884-3 [Hai16]. 978-0-2264-0207-9 [Ebe18]. 978-0-2264-7612-4 [Sac19]. 978-0-2265-0185-7 [Bod19]. 978-0-2265-1000-2 [Ayl18]. 978-0-2265-5161-6 [Wal19a]. 978-0-22650-123-9 [Web18b]. 978-0-2266-1025-2 [Ins19]. 978-0-230-00892-2 [Ree10]. 978-0-230-11147-9 [Nor12]. 978-0-230-20281-8 [Qui12].

978-0-230-21817-8 [Cal12]. 978-0-230-23009-5 [Dun11]. 978-0-230-23067-5 [Smi11b]. 978-0-230-23172-6 [Aga13a]. 978-0-230-23313-3 [Fin11]. 978-0-230-25032-1 [San13]. 978-0-230-27618-5 [Kno14]. 978-0-230-27963-6 [Ase17]. 978-0-230-28427-2 [Mye12]. 978-0-230-28987-1 [Hoc14]. 978-0-230-30196-2 [Sim14]. 978-0-230-30801-5 [McG15]. 978-0-230-34050-3 [Mat13]. 978-0-230-34364-1 [Far13]. 978-0-230-35616-0 [Joh13a]. 978-0-230-55521-1 [Ada10]. 978-0-230-57484-7 [Mor10a]. 978-0-231-14538-1 [Sea10]. 978-0-231-14766-8 [Ren15]. 978-0-231-17102-1 [Tys17]. 978-0-231-17684-2 [Pea18]. 978-0-231-17954-6 [Das18]. 978-0-231-51952-6 [Ren15]. 978-0-2311-6484-9 [Mil16]. 978-0-252-08205-4 [Mil17]. 978-0-253-22032-5 [Pha10]. 978-0-262-01317-8 [Art11]. 978-0-262-01338-3 [Wol11]. 978-0-262-01373-4 [Dar11]. 978-0-262-01398-7 [Bud11]. 978-0-262-01458-8 [Mai11]. 978-0-262-01478-6 [Gou12b]. 978-0-262-01649-0 [And12c]. 978-0-262-01675-9 [Mil12]. 978-0-262-01726-1 [Sar14]. 978-0-262-01746-6 [Pin13]. 978-0-262-01826-5 [Lan14]. 978-0-262-02657-4 [Jon10b]. 978-0-262-02819-6 [Lea16]. 978-0-262-02876-9 [Boo17]. 978-0-262-03394-7 [Ric16]. 978-0-262-03418-0 [Tat17a]. 978-0-262-03443-2 [San17]. 978-0-262-03454-8 [Nic17a].

978-0-262-03477-7 [Mod17]. 978-0-262-03509-5 [Prk18]. 978-0-262-03526-2 [Jam18]. 978-0-262-03554-5 [Tat17b]. 978-0-262-03618-4 [Nic17b]. 978-0-262-07293-9 [Web10]. 978-0-262-08378-2 [Tur10]. 978-0-262-12308-2 [Fin10a]. 978-0-262-13503-0 [Fin10a]. 978-0-262-19590-4 [Fin10a]. 978-0-262-20168-1 [Söd10]. 978-0-262-20172-8 [Söd10]. 978-0-262-20176-6 [Söd10]. 978-0-262-51578-8 [Sar13]. 978-0-262-52538-1 [Lo15]. 978-0-262-52653-1 [Aga15]. 978-0-2620-3503-2 [Wad19]. 978-0-2620-3823-2 [SJ19]. 978-0-2620-3861-4 [Bad19]. 978-0-266-19499-8 [Sac16a]. 978-0-295-99320-1 [Sta16a]. 978-0-300-12381-4 [Sno10]. 978-0-300-12439-2 [Fin10c]. 978-0-300-13685-2 [Hol10]. 978-0-300-13911-2 [Cun10b]. 978-0-300-15091-9 [Sle10]. 978-0-300-15093-3 [Wal11b]. 978-0-300-15197-8 [Far11]. 978-0-300-15911-0 [SS11]. 978-0-300-16042-0 [Far10a]. 978-0-300-16991-1 [Ser14a]. 978-0-300-17934-7 [Die15]. 978-0-3002-2268-5 [Bod19]. 978-0-313-33190-9 [Ade10]. 978-0-374-53499-8 [Fix16]. 978-0-393-06563-3 [Smi10b]. 978-0-415-56142-6 [Pri13c]. 978-0-415-58059-5 [Roe17]. 978-0-415-80871-2 [Wil16a]. 978-0-415-85642-3 [BS16]. 978-0-4157-0296-6 [D’A16]. 978-0-465-01723-2 [Far12]. 978-0-4650-5599-9 [Hol16c]. 978-0-470-67154-2 [Ham14b].

978-0-472-05356-8 [Ren19]. 978-0-500-25218-5 [Mar17b]. 978-0-500-51936-3 [Ric18]. 978-0-520-26278-2 [Chi12]. 978-0-520-26903-3 [Mul13]. 978-0-520-28802-7 [Boo17]. 978-0-520-29339-7 [Lou19]. 978-0-521-11301-4 [Pea12]. 978-0-521-17052-9 [BZ11a]. 978-0-521-19531-7 [San15b]. 978-0-521-51757-7 [Rad13]. 978-0-521-67387-7 [Cha10]. 978-0-521-70255-3 [Osl10]. 978-0-521-71184-5 [Hoq10]. 978-0-521-74509-3 [Rad13]. 978-0-521-76751-4 [Tot10]. 978-0-521-82170-4 [Sta11c]. 978-0-521-82834-5 [Hus15]. 978-0-521-84324-9 [Tau11]. 978-0-521-85181-7 [Rob10b]. 978-0-521-87079-5 [Hoq10]. 978-0-521-87906-4 [SJ11]. 978-0-521-88484-6 [Sta11c]. 978-0-521-88809-7 [Rad13]. 978-0-521-89204-9 [Jam10a]. 978-0-52119-236-1 [Met11]. 978-0-571-23160-7 [Jam11a]. 978-0-674-02424-3 [Pic10b]. 978-0-674-02663-6 [Cha10]. 978-0-674-03166-1 [Har10a]. 978-0-674-04661-0 [Par11]. 978-0-674-04936-9 [Wes12b]. 978-0-674-05233-8 [Hal14b]. 978-0-674-05568-1 [Pri12]. 978-0-674-06587-1 [Nav14a]. 978-0-674-06663-2 [LN14]. 978-0-674-08866-5 [Tho16b]. 978-0-674-08905-1 [HA16]. 978-0-674-28614-6 [Joh16]. 978-0-674-50464-6 [Dih17]. 978-0-674-66041-0 [Alb16]. 978-0-674-72488-4 [Les15]. 978-0-674-72871-4 [Ren15]. 978-0-674-73682-5 [Tur16b].

978-0-6745-0423-3 [Sac16b]. 978-0-6747-2526-3 [Vol16]. 978-0-691-01186-8 [Car11]. 978-0-691-09182-2 [Jon10a]. 978-0-691-11927-4 [Bec18c]. 978-0-691-13983-8 [Rob14]. 978-0-691-14014-8 [Ale11]. 978-0-691-14301-5 [Coo14]. 978-0-691-14507-5 [Sau13]. 978-0-691-14552-5 [Smi12]. 978-0-691-14576-1 [Mur12]. 978-0-691-1478-7 [Eic12b]. 978-0-691-14782-6 [SS17]. 978-0-691-14905-9 [Hol15]. 978-0-691-15271-4 [Hen13b]. 978-0-691-15478-7 [Ksi13]. 978-0-691-15574-6 [vL13b]. 978-0-691-16528-8 [Hah16]. 978-0-691-16935-4 [LD18]. 978-0-7123-5880-4 [Doh14]. 978-0-7188-9276-0 [Gar14]. 978-0-7190-3357-5 [SS15]. 978-0-7190-7797-5 [Sta12b]. 978-0-7190-7976-4 [Mis11]. 978-0-7190-8114-9 [Mac10]. 978-0-7190-8261-0 [And12b]. 978-0-7190-9032-4 [Jar16]. 978-0-7190-9144-5 [Kir15]. 978-0-7294-0948-3 [Sha10]. 978-0-7294-0992-6 [Dun12a]. 978-0-7294-1055-7 [dC14]. 978-0-7294-1077-9 [Hei15]. 978-0-7456-3676-4 [For10]. 978-0-7456-4147-8 [Ren10]. 978-0-7456-5190-3 [Kid11b]. 978-0-7456-8204-4 [Gri15]. 978-0-7486-3940-3 [Daw10]. 978-0-7486-3965-6 [Les13b]. 978-0-7524-4867-1 [Van11b]. 978-0-7524-5283-8 [Van11b]. 978-0-7546-0131-0 [Jam13]. 978-0-7546-0612-3 [Gar11]. 978-0-7546-3250-4 [Big10]. 978-0-7546-553908 [Mai12c].

978-0-7546-5705-7 [Til10]. 978-0-7546-5761-3 [Hen10]. 978-0-7546-5987-7 [Bow10]. 978-0-7546-6150-4 [Jar10]. 978-0-7546-6269-3 [Far10b]. 978-0-7546-6332-4 [Rob10a]. 978-0-7546-6338-6 [Ber11a]. 978-0-7546-6632-5 [Wil11b]. 978-0-7546-6699-8 [Wit11b]. 978-0-7546-6854-1 [Tar12]. 978-0-7546-6960.9 [Row14]. 978-0-75462-2914-6 [Bur13]. 978-0-75462-2916-0 [Bur13]. 978-0-8014-4803-4 [Rot16]. 978-0-8014-4864-5 [Dun12a]. 978-0-8014-7480-4 [Row10]. 978-0-8018-8693-5 [Dry10]. 978-0-8018-9040-6 [Ham10]. 978-0-8018-9313-1 [Str11]. 978-0-8018-9359-9 [Mil11a]. 978-0-8018-9466-4 [Leh12a]. 978-0-8018-9696-5 [Sta11d]. 978-0-8018-9739-9 [Wil14]. 978-0-8018-9904-1 [Eic12a]. 978-0-8020-9284-7 [Jam11b]. 978-0-8047-4504-8 [Ser12]. 978-0-8047-5358-6 [Smi11d]. 978-0-8047-6289-2 [Mey11]. 978-0-8047-6865-8 [TS14]. 978-0-8047-8297-5 [Sin15]. 978-0-8047-8411-5 [Lee15b]. 978-0-8122-4191-4 [Stu11]. 978-0-8122-4305-5 [Par12]. 978-0-8122-4781-7 [Sch17b]. 978-0-8122-4919-4 [Cur18]. 978-0-8122-4921-7 [Bec18a]. 978-0-8122-4959-0 [McC19]. 978-0-8130-5443-8 [Smi19]. 978-0-8135-5465-5 [Whi16]. 978-0-8139-3060-2 [Sle12]. 978-0-8165-3187-5 [May17]. 978-0-8223-4640-1 [Hig10]. 978-0-8223-4731-6 [Rus11a]. 978-0-8223-5744-5 [vL16].

978-0-8229-4378-6 [Ste11b]. 978-0-8229-4443-0 [Dun16b]. 978-0-8229-4445-4 [Ayl17]. 978-0-8229-4450-8 [Fal18]. 978-0-8229-4452-2 [Sim17a]. 978-0-8229-4466-9 [Jar17]. 978-0-8229-4475-1 [Ree19a]. 978-0-8229-4529-1 [Ayl19]. 978-0-8229-4538-3 [Hen19]. 978-0-8229-4558-1 [Ins19]. 978-0-8229-4565-9 [Tun19]. 978-0-8229-6290-8 [Ree16a]. 978-0-8229-6363-9 [Hol16b]. 978-0-8264-3875-1 [YM10]. 978-0-86193-321-1 [Hoc16a]. 978-0-86341-965-2 [Goo11a]. 978-0-86698-415-7 [Tim12]. 978-0-86698-447-8 [Edw14]. 978-0-87071-867-0 [Pis17]. 978-0-87169-267-2 [Gil15a]. 978-0-87328-249-9 [Smi14]. 978-0-88135-387-7 [Wil11a]. 978-0-89327-970-7 [Smo10]. 978-0-906450-18-5 [Keh12]. 978-0-916101-67-1 [Kao12]. 978-0-9556983-9-2 [Sal14]. 978-0-9565295-1-0 [Sim17b]. 978-0-9567795-8-8 [Les13a]. 978-0-9574530-2-9 [Les14]. 978-0-9935510-0-0 [Bil17]. 978-0226-05845-0 [Arb17]. 978-0226-30570-7 [Gou12c]. 978-07-123-5868-2 [Dun15a]. 978-0822944546 [Fin17]. 978-0903874-43-4 [Roo12c]. 978-1-107-00397-2 [Jam12]. 978-1-107-01221-9 [Cuo13]. 978-1-107-01457-2 [Cuo12]. 978-1-107-01744-3 [Nav15b]. 978-1-107-02480-9 [Ree14]. 978-1-107-02595-0 [Tam14a]. 978-1-107-02842-5 [Daw15a]. 978-1-107-03312-2 [Ree17]. 978-1-107-03361-0 [Nal15].

978-1-107-03617-8 [Daw14]. 978-1-107-07052-3 [Dun16a]. 978-1-107-12697-8 [Kum16]. 978-1-107-13436-2 [Daw16b]. 978-1-107-54560-1 [Huf16]. 978-1-107-61983-8 [Joh15]. 978-1-107-65815-8 [Bar15b]. 978-1-107-68802-5 [Som15]. 978-1-1070-3360-3 [Tat16]. 978-1-1071-3490-4 [Lan19]. 978-1-118-62077-9 [Daw16a]. 978-1-137-03186-0 [Wol14]. 978-1-137-27207-2 [Ham14a]. 978-1-137-31131-3 [Ksi15]. 978-1-137-31173-3 [Kaa18]. 978-1-137-32156-5 [Wes15]. 978-1-137-34092-4 [Aga14]. 978-1-137-37242-0 [Mil15]. 978-1-137-37771-5 [Ber17]. 978-1-137-38443-0 [Hoc16a]. 978-1-137-40651-4 [Car16b]. 978-1-137-43872-0 [Har16]. 978-1-137-46744-7 [Ros16]. 978-1-137-46747-8 [Doh16]. 978-1-137-49272-2 [Kee17]. 978-1-137-49768-0 [Tra16]. 978-1-137-51139-3 [GP18]. 978-1-137-52063-0 [Köb16]. 978-1-137-54638-8 [Lin17]. 978-1-137-56760-5 [Ham16]. 978-1-137-56761-1 [Pay16]. 978-1-137-59687-1 [Rob17]. 978-1-137-59769-4 [Duf18]. 978-1-1374-0687-3 [Sta19c]. 978-1-1375-9797-7 [Ber18]. 978-1-138-02359-8 [Kum18]. 978-1-138-67938-2 [D’A18]. 978-1-138-69696-9 [Wal17]. 978-1-138-82817-9 [Bod18]. 978-1-4020-4559-2 [Siv10]. 978-1-4039-8724-2 [Rod15a, Rod15b]. 978-1-4039-8802-7 [Lan15a]. 978-1-4039-8803-4 [Daw17]. 978-1-4051-5625-7 [Sch17a].

978-1-409-41886-3 [Wag14]. 978-1-4094-0038-7 [Dys14]. 978-1-4094-0075-2 [Bak11]. 978-1-4094-0613-6 [Win13]. 978-1-4094-2021-7 [Tur12b]. 978-1-4094-2642-4 [Gou13a]. 978-1-4094-2868-8 [Gou12a]. 978-1-4094-3301-9 [BT14]. 978-1-4094-3580-8 [Sto14]. 978-1-4094-4921-8 [Sch15b]. 978-1-4094-5237-9 [Hig14a]. 978-1-4094-6335-1 [Wag15]. 978-1-4094-6643-7 [Dun15b]. 978-1-4094-6815-8 [Wal16b]. 978-1-4094-6922-3 [OP16]. 978-1-421-41608-3 [Bow16]. 978-1-4214-0457-8 [Joh14]. 978-1-4214-0468-4 [Ban15]. 978-1-4214-0522-3 [War13]. 978-1-4214-0530-8 [Joh13b]. 978-1-4214-0749-4 [Gib14]. 978-1-4214-0771-5 [Hol13a]. 978-1-4214-0790-6 [Kir13]. 978-1-4214-0981-8 [Sta15]. 978-1-4214-1022-7 [Tur15]. 978-1-4214-1131-6 [Ser15]. 978-1-4214-1316-7 [Mos16b]. 978-1-4214-1340-2 [Web15]. 978-1-4214-1562-8 [Bro16]. 978-1-4214-1591-8 [Aga17b]. 978-1-4214-1593-2 [Coz16]. 978-1-4214-1604-5 [Gif16]. 978-1-4214-1783-7 [Arn16]. 978-1-4214-1903-9 [Fra17]. 978-1-4214-1954-1 [Moo17]. 978-1-4214-2186-5 [Mar18a]. 978-1-4214-2618-1 [Can19b]. 978-1-4214-2658-7 [Pea19]. 978-1-4331-2319-1 [Ken19]. 978-1-4411-1697-0 [Fan12]. 978-1-4411-6976-1 [Zai16]. 978-1-4426-2652-2 [Adl16]. 978-1-4426-4395-6 [Ber12]. 978-1-4438-0357-1 [Gas10].

978-1-4438-2680-8 [Kee12]. 978-1-4438-2918-2 [Mar12b]. 978-1-4696-1290-4 [Ree16b]. 978-1-4696-2947-6 [Wai18]. 978-1-4696-3173-8 [Hay19b]. 978-1-4696-3287-2 [Mar18b]. 978-1-4704-1076-6 [Hol16a]. 978-1-4704-1469-6 [Sør17]. 978-1-4704-1493-1 [Bar16]. 978-1-472-46410-1 [Big18]. 978-1-4724-1276-8 [Røs15b]. 978-1-4724-1954-5 [Bil15a]. 978-1-4724-2232-3 [Coz17]. 978-1-4724-2257-6 [Stu17]. 978-1-4724-2810-3 [Mox16a]. 978-1-4724-3425-8 [Sim16]. 978-1-4724-4652-7 [Feo18]. 978-1-4724-6622-8 [Bec16]. 978-1-4724-7872-6 [Dun17b]. 978-1-4724-8461-1 [Fot18]. 978-1-4724-8829-9 [Ful18]. 978-1-47244436-3 [Goo17]. 978-1-4742-8308-3 [Smi18a]. 978-1-4773-0359-7 [Yal17]. 978-1-4875-0367-3 [War19]. 978-1-5036-0344-8 [SM19]. 978-1-5036-0544-2 [Far19]. 978-1-5095-1892-0 [Tho19]. 978-1-5261-0144-0 [Hal17]. 978-1-5261-1088-6 [Law17]. 978-1-55238-705-4 [Rob15]. 978-1-5800-46595-3 [Smi18b]. 978-1-58046-408-6 [Sta14]. 978-1-58046-476-5 [San15a]. 978-1-58046-527-4 [Han16]. 978-1-60497-751-6 [Bud13a]. 978-1-60618-920-7 [Smi11c]. 978-1-60618-991-7 [Ric10]. 978-1-60618-994-8 [Fre10]. 978-1-60618-995-5 [Gra11]. 978-1-61168-044-7 [Rud13]. 978-1-7802-3640-7 [WM18a]. 978-1-78023-033-7 [Cha14]. 978-1-78023-375-8 [Far15].

978-1-78023-824-1 [Bec18b]. 978-1-78074-532-9 [Bar15a]. 978-1-78326-973-0 [Mor18]. 978-1-7849-9413-6 [Wal18]. 978-1-7849-9512-6 [MY18]. 978-1-7849-9516-4 [Bui19]. 978-1-78683-004-3 [Hun18]. 978-1-78707-138-4 [Jim19]. 978-1-7873-5005-2 [Hay18]. 978-1-7891-4085-9 [Fla19]. 978-1-84384-421-1 [Lou16a]. 978-1-84465-204-4 [Tos11]. 978-1-84465-587-8 [Daw16c]. 978-1-84511-705-4 [Mac11b]. 978-1-84519-381-2 [Stu11]. 978-1-84519-481-9 [Man13b]. 978-1-84614-035-8 [McO10]. 978-1-84614-6572 [Ros17]. 978-1-8463-1496-4 [Ban12]. 978-1-84631-959-4 [Bar15c]. 978-1-84882-554-3 [Ens12]. 978-1-84885-366-9 [Wig11]. 978-1-84885-518-2 [Moo13]. 978-1-8489-3420-7 [Ham15]. 978-1-8489-3518-1 [Sch16]. 978-1-8489-3527-3 [Tho16a]. 978-1-84893-008-7 [Coo11]. 978-1-84893-016-2 [Gil11]. 978-1-84893-020-9 [Wit11a]. 978-1-84893-030-8 [Boa12]. 978-1-84893-100-8 [Goo12b]. 978-1-84893-126-8 [Ric12]. 978-1-84893-308-8 [Ken13]. 978-1-84893-362-0 [Sal15]. 978-1-84893-364-4 [Raj15]. 978-1-84893-394-1 [Daw15c]. 978-1-84893-423-8 [Bak15b]. 978-1-84893-426-9 [Rad14]. 978-1-84893-428-3 [Hom16]. 978-1-84893-430-6 [Ham15]. 978-1-84893-443-6 [Lou15]. 978-1-84893-463-1 [Jac15]. 978-1-84971-161-6 [Wal12c]. 978-1-84973-326-7 [And12d].

978-1-85196-636-3 [Alb11a]. 978-1-85196-653-0 [Cul11]. 978-1-85196-658-5 [Alb11a]. 978-1-85196-907-4 [Hul10]. 978-1-85196-970-8 [Smo10]. 978-1-85196-974-6 [Mar12a]. 978-1-85196-975-3 [Cow10]. 978-1-86189-375-8 [Han17]. 978-1-86189-804-3 [Bel12]. 978-1-893616-00-4 [Fer11a]. 978-1-900289-96-2 [Joh11]. 978-1-905267-47-7 [Elt12]. 978-1-906165-08-6 [WM11]. 978-1-9094-0060-3 [Kus18]. 978-1-909662-25-4 [Tur16a]. 978-1-920898-80-9 [Bro11a]. 978-142140936-8 [Cle15]. 978-1472469168 [Cle17]. 978-2-02-040298-9 [Man13a]. 978-2-13-056626-7 [Esp10]. 978-2-251-34512-3 [Chr10]. 978-2-251-34513-0 [Chr10]. 978-2-271-09184-0 [Bus18]. 978-2-4060-8072-5 [Gis19]. 978-2-503-52584-6 [Boa10]. 978-2-7116-2330-3 [Hed13]. 978-3-0300-1637-1 [Fix19]. 978-3-0343-0708-6 [Bar13a]. 978-3-0348-0285-7 [Buk14]. 978-3-05-004528-3 [Sta11a]. 978-3-05-006478-9 [vL15]. 978-3-11-025306-1 [GG12]. 978-3-319-14552-5 [Tam16a]. 978-3-319-26246-8 [Lyk16]. 978-3-319-32454-8 [Bow18]. 978-3-319-39647-7 [Sia18]. 978-3-319-54653-7 [Nic19]. 978-3-3192-5925-3 [Gul18]. 978-3-3194-9162-2 [Ako19]. 978-3-428-13849-4 [Cap14]. 978-3-515-09601-0 [Mic11]. 978-3-515-09602-7 [Gli11]. 978-3-515-09606-5 [Wal11a]. 978-3-515-09842-7 [Dun12b].

978-3-515-10046-5 [Ser14b]. 978-3-525-31705-1 [Web13]. 978-3-540-69221-8 [Hen11b]. 978-3-631-71992-3 [Pri19b]. 978-3-643-90095-1 [Egm12]. 978-3-7913-4142-2 [Coe10]. 978-3-8353-0594-6 [Web11]. 978-3-8353-0596-0 [Hen11a]. 978-3-8353-1071-1 [Web14]. 978-3-8442-5871-4 [Sau15]. 978-3-8471-0263-2 [Hig16]. 978-3-938944-63-9 [Mar14a]. 978-3938016-14-5 [vL13a]. 978-8-7730-4417-9 [Hun19]. 978-8-8400-1869-0 [Cle18]. 978-8-8858-0311-4 [Pri19a]. 978-80-87378-67-0 [Gli12]. 978-80-87378-73-1 [Gli12]. 978-83-7656-022-9 [Ree15]. 978-84-370-9446-5 [Car16a]. 978-84-7584-657-6 [Ric11b]. 978-87-7112-594-8 [Goo18]. 978-87-7124-014-6 [Gou14]. 978-87-7934-311-5 [Hig11]. 978-87-7934-317-7 [Hal10]. 978-88-222-6198-5 [Col13]. 978-88-222-6254-7 [Tar15]. 978-88-8492-564-0 [Kne10]. 978-9-004-2639-4 [Smi16b]. 978-9-004-26421-2 [Wri16]. 978-9-004-28679-5 [Loi16]. 978-9-0042-1696-9 [Ver16]. 978-9-0042-7437-2 [Gou16]. 978-9-0043-1009-4 [Car18]. 978-9-0043-1065-0 [Car18]. 978-9-0043-2089-5 [Ben19a]. 978-90-04-16645-5 [Pha10]. 978-90-04-17743-7 [Dub10]. 978-90-04-18107-6 [Van11c]. 978-90-04-20703-5 [Fon12]. 978-90-04-20712-7 [Sto12]. 978-90-04-20935-0 [Sma13]. 978-90-04-21150-6 [Bak14]. 978-90-04-21870-3 [Dob14].

978-90-04-21871-0 [Roo12a]. 978-90-04-22193-2 [Hig13]. 978-90-04-22918-1 [Sho14]. 978-90-04-23058-3 [Fal13a]. 978-90-04-24390-3 [Bak15a]. 978-90-04-24441-2 [Pos14]. 978-90-04-25178-6 [Mos16a]. 978-90-04-25296-7 [Dun14]. 978-90-04-26023-4 [Car16a]. 978-90-04-26330-7 [AN18]. 978-90-04-26439-7 [Hig15]. 978-90-04-27665-2 [Sha11]. 978-90-04-32492-3 [Dun17a]. 978-90-420-2658-2 [Nes10]. 978-90-420-2728-2 [Som11]. 978-90-420-3747-2 [Eic15]. 978-90-481-2361-2 [Sum11]. 978-90-481-3541-7 [Pum12]. 978-90-481-8644-0 [Rap11]. 978-90-5201-656-6 [Sim13]. 978-91-87351-10-5 [Røs15a]. 978-91-981948-0-7 [Gil16a]. 978-94-007-0036-9 [Oos12]. 978-94-007-3931-4 [Boa15]. 978-94-007-5350-1 [Ric14]. 978-981-4299-37-4 [Gil12]. 979-185918-497-4 [Cul13]. 987-2-8124-1419-0 [Ver15a].

A. [Elt12, Eve18, Stu11]. A.D [Dun15b]. Aalborg [Goo18]. Aarhus [Gou14, Hal10, Hig11]. Aaron [Arb17, Leh12a]. Abena [Coz15]. aberration [Fis10]. Abetted [Stu11, Ros10]. Abingdon [BS16, Big18, Bod18, D’A18, Daw16c, Dun17b, Feo18, Ful18, Kum18, Roe17, Wal17]. abortion [Kir17]. Abraham [Pha10, Prk18]. absolute [BG11]. Abt [Car12]. Abundance [Cha10, OR08]. Academic [Man13b, Røs15a, Smi18a, Stu11]. Academicae [Cla14]. Academy [Gli12]. Accademia [Wil11a]. Access [Sau15]. Accidents [Ros16]. Accompanying [Car16a]. account [Van11a]. accounting [Neu18]. Accumulation [Egm12, Rob11]. Ackerknecht [Moo17]. Ackermann [Hig15]. Acolytes [Phi12, Web14]. across [Ako19, Bar16, BM11, Gil16a, Hol14b, Par12, Røs15a, THB12]. Act [ES16, Ter14]. Action [Bak15a, HSSR13]. actions [A^{+}12, Ser14b]. Activists [Mar12b, AM11]. activities [Sim19a]. Acumen [Tos11]. Adam [Ayl18, BS16, Car16a, Met11]. Adaptation [San15a, CR14]. address [Rad16]. Adelman [Cul11]. Administrative [San10b]. Adrian [Mac11a, McO10]. Adriana [War19]. Advanced [Gil16a]. Advancement [Mis11, Toa11, Toa12, Wit10]. adventure [Koh08, Smo10]. Adventures [Tun19]. Advertising [Bak11, Wig10]. Adviser [Smi10a]. Aepyornis [And13b]. Aerospace [Smi14, Wes12a]. Aesthetics [Dra17, GP18, Sta16a, Bra15]. Affair [Cal13, Man14, Tur12a]. Africa [Coz15, Dri10, Dub10, Hui09, Mac11b, OA14]. Africans [Sar14]. After [Ric13b, BG14, Bou15, Eic15]. Against [Fey10, Gil15b, Tur16a, Kid11a]. Agassi [Ric14]. Age [Bak11, Daw10, Dun12a, Elt12, Far10a, Far16, Fin17, Goo17, Jac15, Jam18, Mar12a, Mil13, Sta11c, Wit11a, Bro10b, Bur12, Car10, Des12, Hem15, Hol09, Jac13, JJ14, Mil08, ML10b, Sch11, Sec14, Sta10, Wig10, Fox09, Kni15, Røs15b, Smi19]. Agents [Qui12, GS10]. Ages [Rie16, Bar08, Bec18a, CG12, DeV09, Kle13, Osl10, Sea10, Fal13a]. Agnes [Feo19]. Agricultural [Cha10, OR08]. agriculture [Uek10, Web18a]. Aguaplano [Pri19a]. Agustí [Far10b]. Ägypter [BW13, vL13a]. Ahmad [SS11]. Air [Dan19, Ree17]. Air-Related [Dan19]. Airs [Ree16a, FJB14]. Aitor [Hug10]. Akademie [Sta11a, vL15]. Akademie-Verlag [Sta11a]. al [Win15]. Alain [Chr10]. Alain-Philippe [Chr10]. Alan [Cha10, Eve13, Fin10a, Fin10b, App12, Sum11, vL13b]. Albert [Mil11a, Hun10]. Alberti [Lou17, Mac10, Wad11]. Albritton [Gil16b]. Alchemical [Cle17]. Alchemy [Sac16b, Sea10, DeV09, Pri13b, Ray15, San14, Tim12, LG11]. Aldershot [Big10, Bow10, Far10b, Hen10, Jar10, Rob10a, Til10]. Alembics [McC19]. Aleppo [Mil11d]. Alex [Fal18]. Alexander [Bar15a, Coe10, D’A18, Dry10, Fix16, Jim19, Mil11d, Par11, Aga13a, Coe10, Joh11, KOW12, Lac09, Liv15, Mar12c, Mar14a, Rap12, Wal09]. Alexander-Skipnes [D’A18]. Alexandra [dC14]. Alferov [Mai11]. Alff [McC19]. Alfred [Les15, Ben13b, Cow13, Fie10, Fla19, Les13a, Les14, Smi12, SDS13, WC13]. Algebra [Hol15, KP14]. Algebraic [Bar16, Hol14b]. Algeria [Eff17]. Alice [Hig16, Ren19]. Alisha [Tar12]. Alison [Bak14, Bal14b, BS16, Bjö11, Doh14, Dun17a, Ren15]. Aliveness [Smo10, Ayr08]. Allais [Heu15]. Allen [Ade10, Gar14, McO10, Ros17]. Allende [And12c, Med11]. allergy [Smi15, Mil16]. Allies [Sør17]. ally [Yal11]. Almost [Rei15, EKD^{+}13]. Alois [Gli12, Š^{+}11]. Alper [Tam16b]. altägyptische [BW13, vL13a]. alten [BW13, vL13a]. Alternative [Dun12a, Lov10]. Aluminum [Arn16, Mar18a]. Alun [Doh16]. am [Pri19b]. Amanda [Mon10]. amateur [Cur14]. Amazonian [Opi14a, Hol13b]. America [Bow16, Pet13, Pro11, Rie16, Wal17, Wis12, Wyl11, Bri10, Cha10, Coe10, Fra17, Hol13a, Koh10, Lan14, O’C17, Pau07, Pet10b, Pod13, Ree19a, Sol13, Vic11, Wal09, Whi16, Wol13, Whi15]. American [Aga17a, Bar16, Bel17, Boo14, Cha10, Cur18, Daw10, Fre10, Gil15a, Gra11, Hen19, Hol16a, Kir13, Les13b, Les13a, Ree16b, Ric10, Sch17a, Ser14a, Smi11c, Sør17, Tun19, Web18a, Wes12b, Wil16a, Wol14, Abt11, Ada16, Car12, Cas11, Com12, Fer12, Gab16, Hol09, LaF13, Lüt19, Moh13, OR08, Ste14a, Ter13, SDS13]. Americas [Coe10, KOW12, Lac09, Mar14a]. Amherst [Bud13a]. Amir [Bar15a, Fix16, Par11]. among [Fix19, Yal17]. Amparo [Coz17, Ric11b]. Ampère [Fal18]. Amsterdam [Eic15, Nes10, Som11]. Ana [Tam16a]. analogical [Lam11]. analogy [Lam11]. analysts [Pic14, Can15]. Anatomical [Mar17b, Mit12, Wal16b, Ber13, KZ15, Wag14]. Anatomis’d [Ber11a, Cun10a]. Anatomist [Ber11a, Mae12, Cun10a, Mes10]. Anatomists [Sho18, JA15]. Anatomy [Egm13, Stu11, Wal17, Bat10, Ber11b, Kus11, Pod13, Ros10, Eic12a, Han17]. Ancestors [Jam11b, Hes09]. Ancient [Cuo13, D’A16, Dys14, Eve18, Jon10a, Mur12, RVG12, Rob10b, Tau11, BW13, BJ10, Che12, Cre10, Hud14, Leh07, Llo12, Rob08, BW13]. Andean [Sim17a]. Anderson [Goo10a, Ham10, Hen14, Jam13, Rud13]. André [Bak15a, Van11c]. Andreas [Goo18]. Andréolle [Mar12b]. Andrew [Ayl17, Ber18, Biv10, Boo17, Fin11, Hen15b, Tys17, Wit11b, Ber11a, Gar11, Gou17, Jam11a, Und12, Van11b, vL13b]. Anduaga [Hug10]. anew [Rus14b, Far15]. Angela [Ksi15, Cap15]. Angeles [Chi12]. Angelina [Aga14]. Angelique [Eic15]. Anglo [Les13b, Fer12]. Anglo-American [Les13b]. Angola [San10b]. Angus [Mai12a]. Animal [Gib15, Gou13b, SFFC12, Cow13, Mye17, Kni16]. Animals [Adl16, Eic15, Hoc16a, Kus18, Mca16, Sho18, Sle12, WM18a, Cow14, Der15, Ric13b, Rit10]. Animating [Wel17]. Animi [Gat17]. Anita [Dun15b, Sho18]. Anja [AN18]. Anja-Silvia [AN18]. Ann [Fin10b, Gil15a, Ree16a, Ren19, AN18]. Anna [Sim13, Fon12, Ham16, Mes10, Pay16, Web14]. Anne [MY18, Bow18, Daw14, Dys14, Smi10b, Ver15a]. Anne-Lise [Ver15a]. Annette [Buk14]. Annie [Ber18].

Annotated [Les15, Sto12, FF11, WC13]. Anstey [Lan12]. answer [Mur14]. answers [C^{+}12]. Anthony [AN18, Hay19b, Mai12c, Wal11b]. Anthrax [Sta11d, Jon10c]. Anthropologists [SS16, JA15, Lar14]. Anthropology [Bar13b, Bel12, Daw15c, MR13, SS12, SS13b, Lar13, Liv15, Mia12, Pin11, Qur11]. anthropometric [MR13, Lun13]. anti [GR18]. anti-racist [GR18]. Antibiotic [Coz16, Pod15]. Antilles [Dew10]. antipodean [JA15]. Antiquarianism [Ksi15, Byr13]. Antiquity [Dij12, Dun16a, Hol15, Sho14, Dar12, HKZ12, KP14, Sto13, Gre14b]. Antoniadi [RC11]. Antonio [Coz17, Ric11b, Wil11a]. Anxiety [Bad19]. Apes [Jam11b, Pet18, Hes09]. Apocalyptic [Cur18]. Apothecaries [Sim19a]. Appel [vL13b]. Applebaum [Hig13]. appliances [Jon16b]. Approach [Som15, Ekh10, Lam13]. Appropriation [Smi16b, Yal11]. Aquarium [vN19]. Arab [Mil11c, Kla09]. Arabatzis [Tam16a]. Arabic [Mim15]. arachnologist [Roo11, Fon12]. Aratus [Leh12a]. Arber [Feo19]. Arbor [Ren19]. archaeology [Bul16, Eff17]. Archibald [Bro11a, Pea18, Mac09]. Archimedes [Cuo12, N^{+}11]. Architects [Smi10a]. Architecture [Wal11b, GJ09a, GJ09b]. Archive [Fin10b, Mul13, Mue11]. Arctic [Rob15, Cap13]. Argentina [Coz17]. Argument [Egm13, Hen17, Kus11, Lam11]. Aristotelian [Gul18, Van11c, Fri11]. Aristotelianism [Car18]. Aristotle [Mos16b, Big18, Gre14a, Leh19, Mar14b]. Arizona [Edw14, May17, Tim12]. Ark [vN19]. Armageddon [Smi10a]. Armin [Gli12, vTSŠ^{+}11]. Arms [Sto14, Wal12a]. Armstrong [Fla19]. army [Pye12]. Arnaud [Bar17]. Arnold [Dar11, Kum16, Set10]. Arrizabalaga [Wit11b]. Arsdall [Dys14]. arsenide [Spi11, Spi12b]. Art [Bel17, Bel18b, Cuo12, Der12, Dun15a, Hah16, Hay18, Kus15, Mon10, Røs15b, Tur15, Van14, WM18a, Hun13, JJ14, Kel12, Kem19, McG13, Mer12, Mer13, QM13, Ree09, Sch13, Sum14a, Ber12]. Arthur [Doh14, Ham14b, Kus18, Ayr12]. Arthurian [Lou16a]. Artifact [Mur12, BJ10]. artificial [BH19]. Artisan [Ben19a, Rab17]. Arts [Buc11, Elt12, Far10a, Fox09, Lov10, Pow12, Wer10, Big18, Roo12b]. Arturo [Kus18]. Arwen [Kir13]. Asare [Coz15]. Ash [Hoc14]. Ashgate [Bak11, Bec16, BT14, Ber11a, Big10, Bil15a, Bow10, Bur13, Coz17, Dun15b, Dys14, Far10b, Gar11, Goo17, Gou12a, Gou13a, Hen10, Hig14a, Jam13, Jar10, Mai12c, Mox16a, OP16, Rob10a, Røs15b, Row14, Sch15b, Sim16, Sto14, Stu17, Tar12, Til10, Tur12b, Wag14, Wag15, Wal16b, Wil11b, Win13, Wit11b]. Ashok [Aga14]. Asia [Ako19]. Asian [Pha10, Abr09]. aside [AG17]. Aspects [Mox16a]. assayers [KT19]. Assembling [Law15a, Bux18]. assessment [CC11b, Sys18]. Association [Mis11, Toa11, Toa12, Wit10, Kus18, Lau12, Mac15, Szu18]. Astbury [Hen15a, Hal14c]. Asthma [Jac09, Biv10]. Astroculture [Aga13a, Gep12]. Astrology [Duf18, Dun16b, LN14, Van14, Azz12, Gre14a, McG13, QM13]. Astronomer [Reg17, JS08, Spi15]. astronomers [Pum11]. Astronomes [Chr10, JS08]. Astronomía [Car16a, Nav14b]. Astronomical [CG12, Fal13a, Nas10]. Astronomy [Duf18, Hig10, Kir15, Leh07, Nal15, ABS10, Dic13, Don14, Mol14, Nav14b, RC11, Rob10b]. Atapuerca [Hoc16b]. atelier [BFGG11, Sim13]. Athanasius [Gre14b, Sto12, FF11, Sto13]. Athena [Pye12]. Atlantic [Mar18b, Par12, Sim17a, BM11, Dew10, Mur14]. Atmosphere [Fin10c, Jay09]. Atmospheric [Ric16]. Atom [Cha09, Sum11, Joh12, Lau12].

Atome [Hen11b, Rec10]. Atomic [Cap15, Lau15, Pha10, Pha13, Sta11c, Abr09, Car10, Cre13, Des12, Gro17, McR14, See12, Ban14a, Lau12]. Atoms [Cha09, Lau12, Rec10, Sum11, Tat18]. attachment [Vic11]. attitudes [Bur12]. AU$59.95 [Bro11a]. Aubin [Hig10, Sør17]. audiometric [VM18]. Audra [Hol13a]. Ausgabe [BW13]. ausgehenden [Pri19b]. Austin [Yal17]. Australia [Dev16, Hal17]. Authority [Daw14, Mac16, Wal19a, DeW13, RC11]. authors [Ano10n, Ano11n, Ano12n, Ano13n, Ano14p, Ano15j, Wat14]. authorship [Csi17, Mim15]. Auto [Hig16]. Autobiography [Bjö14, FF11, Man13c, Sto12]. autodidacticism [BZ11b, Wil14]. Autopsy [Wag14, Mit12]. avant [Gis19]. Averroe [Bur11]. Averroës [Gla09]. Aviva [Sch19]. Avner [Wil14]. Ayesha [Shi16]. Ayres [Ham14b, Smo10]. Azzolini [LN14].

B [Ban12, GP18, Lar13, Lou17, Mac11b, Moo13, Ser15, Wig11]. B. [Goo12b, Mar17a]. Babbage [Fis14]. babies [McL12, Mai12a]. Back [Ano10a, Ano10c, Ano10e, Ano10g, Ano11a, Ano11c, Ano11e, Ano11g, Ano12a, Ano12c, Ano12e, Ano12g, Ano13a, Ano13c, Ano13e, Ano13g, Ano14a, Ano14b, Ano14e, Ano14g, Ano14i, Ano15a, Ano15c, Ano15e, Ano15g, Ano16a, Ano16c, Ano16e, Ano16g, Ano17a, Ano17c, Ano17e, Ano18a, Ano19a, Ano19c, Ano19e]. backwards [Tam14b]. Bacon [Eve13, Ric14, Aga13b, Mer12, Mer13, Mor17, SK12, Wee19]. Bacteriology [Cha12, GF09, Sta14, Str11]. Bad [Ken12, Gou13a]. Badino [Sau15]. Baeyer [Dih17]. Baldwin [Wal16a]. balloon [Doh17]. Ballooning [Gil11, Lyn10]. Ballstadt [Sha10]. Balmer [Coz17]. Baltimore [Aga17b, Arn16, Ban15, Bow16, Bro16, Can19b, Cle15, Coz16, Dry10, Eic12a, Fra17, Gib14, Gif16, Ham10, Hol13a, Joh13b, Joh14, Kir13, Leh12a, Mar18a, Mil11a, Moo17, Mos16b, Pea19, Ser15, Sta11d, Sta15, Str11, Tur15, War13, Web15, Wil14]. Bank [Ric18]. Banken [Ree17]. Baptist [Cle17]. Barbara [Ree15]. Barcelona [Ric11b]. Bards [Daw10, Hol09]. Bark [Sim17a]. Barrett [Sau13]. Barrier [Fer11a, Bus10, Jon16b]. Bartlett [Osl10]. Barton [Wal19a]. based [Tur38]. Basel [Fix19]. Bashford [Bjö11, Ren15]. Basic [Far12, Hol16c]. Basil [Goo11a]. Basingstoke [Ada10, Aga13a, Ber17, Cal12, Car16b, Daw17, Doh16, Dun11, Fin11, Goo10a, Ham14a, Ham16, Hoc14, Hoc16a, Joh13a, Kee17, Kno14, Köb16, Ksi15, Lan15a, Mat13, McG15, Mil15, Mor10a, Mye12, Nor12, Pay16, Qui12, Rod15a, Rod15b, Ros16, San13, Sim14, Smi11b, Wes15]. Bates [Stu11]. Bateson [Gli12, Hay19b, Rus14a, Š^{+}11]. Battista [Pri19a]. battles [Ech12]. BBC [Boo12, Nie10a, Jon14a, Jon16a]. be [KT19, Rav11, Swe15]. Beach [Wil11a]. Beagle [CV09, Keh12, Ste11a, Rad13]. Bears [Law15b]. Beatty [Ken13]. Beautiful [Hal14d, vL16]. Beauty [Tys17]. Became [Nav16, Ser14a, Wil16b, BS15, Com12, Csi17]. Beck [Sch16]. Becoming [Coo14]. Beddoes [Fin10c, Ful18, Jay09]. Bede [Ban12, KW10]. Beeckman [Cle15, vB13]. Been [Ayl17, THB12]. bees [Leh19]. before [Car16b, Löw11]. Beginnings [BS16, Wil12, Goo16, Hug11, Sac19]. Behavior [Arb17]. behaviourist [Pet10b]. Beholding [Bui19]. Being [Llo12, Stu11, Ros10, Eva13, Sar14]. Belfast [Bar08, FW15]. Beliefs [Cur18, Som15, Lam13, Kot16]. Belknap [Ren15]. Bell [Lou16b]. Belles [Chr10]. Bellorini [Bec16]. Ben [Wil14, Far15, Sal15, vN19]. Ben-Zaken [Wil14]. Bender [Ser12]. Benjamin [Das18, Rob15, Sin15, Ale11, Bar13a, Bax13, Cap13, Dra17, Tat17a]. Bennett [Daw15a, Dun14]. Bentley [Jam12]. Benton [Les14]. Berber [Eff17]. Beretta [Col13, Wil11a]. Berg [Wal11a]. Bergwik [Kee17]. Berkel [Cle15]. Berkeley [Chi12, Mul13, Smi14]. Berkowitz [Ree19a]. Berlin [Cap14, Egm12, GG12, Gul18, Mar14a, Sau15, Siv10, Sta11a, vL15, SV17]. Bernard [Daw15b, Fal13a, Hed13, Jac15, Ree19a, Bow10, Daw15a, Daw16a, Mur14, Pos14, Smi16b]. Bernd [Rad14]. Bernhardt [Nic17a]. Bertin [Mit18]. Bertoloni [Bel18b, Eic12a]. Bertrams [Ree14]. Best [Ken19]. Bestandsaufnahme [Hen11a, VW10]. Bethe [Nav14a, Sch12]. Better [Ber12, Buk14, Der12, LOvT12]. Betts [Lin17]. Between [Lee15a, Mar11, San17, Sno10, Tho16a, Bar17, BJ10, Can19b, Hun09, JA15, Kel10, Kem19, Kot16, LMS13, MHO13, Mur12, Pos14, Rus11a, Sal15, VPW15, Rap12]. Beyond [Die15, Wag14, vN19, BDR13, Mit12, Nas10, THB12, TR14, Bod19, Har16]. Bezoars [Doh14, HWM12]. Biagioli [Goo12a]. Bibliothèque [Boa10]. Biess [Bou15]. Big [Ban14a, Røs15a, Pha13]. Bigg [Hig10]. Bild [BDS15]. Bildphysiologie [Kle13, vL15]. Bill [Lou16b]. Bind [Hay19b]. Binzley [Can19b]. Bioethics [Fer13, Das18, Rod15a, Rod15b]. biogeochemical [OS13]. Biographical [Man13b, Cro12]. Biographie [Hen11b, Rec10]. Biography [Hig16, Hul10, Gra13, Ham09, Jac09, Jac08, Mox16b, Rad09, Rec10, Rin04, Scu09, Tat09, Biv10, Hen13b, Ren10]. bioinformatics [Ste13, Web15]. Biological [Cha10, Joh13a, Mye12, Sto14, Dev16, Fri11, OR08, Ren12, Wal12a, Wil11c]. biologie [Gis19]. biologies [End18a]. Biology [Ber17, Erl11, GS12, Ham15, Jar17, Rad13, San12, Cur14, Esp13, HD13, Pet14, Rus11b, SF11, Ber15, Erl11, Erl13, Jar16, San12, San13]. Biomedical [Nov12, Dav19, Ban15]. Biomedicine [Lee15b, CC11a, DiM13]. biomedizinischen [Fan10, Sta11a]. biosphere [OS13]. Biotech [Web15, Cur14, Hug11, Ras14, Wil12]. biotechnology [Mye17, Yi15, Ree18, San16]. Biotopia [End18b]. Birds [Wal18, And13b, CGC13a, CGC13b, Kus18]. Birgit [Hai16]. Birkhäuser [Buk14, Fix19, Sia18]. Birth [Nic17a, Tur14, Gor12]. Birthing [Wil11b]. bis [Hen11b, Rec10]. Bitter [OA14, Coz15]. Bivins [Ste16]. Bjerknes [Ric16]. BJH [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14c, Ano14a, Ano14d, Ano14b, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano18b, Ano18a, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e]. Black [Ban14a, AJ12, Jam13]. Blacked [Gal12]. Blacked-out [Gal12]. Blackness [Mar18b]. Blackwell [Daw16a, Ham14b, Sch17a]. Blair [AN18]. Blancke [Bro16]. Blay [Bus18]. Bleichmar [Par12, Smi11d, SS13a]. Blood [HKZ12, Neu18, Pea19, Ban14b, Sho14]. blood-group [Ban14b]. Bloom [Opi14a, Hol13b]. Bloomington [Pha10]. Bloomsbury [Smi18a, Zai16, Hay18]. Blue [Wes12a, OG17, Smi14]. Blume [Law17]. board [Ben19b]. Boas [Fie10]. Boddice [Mil15, Mil17]. Bodies [Kno14, Lee15b, Sin15, Wil11b, Ben13a, DiM13, Pac12, Doh16]. Body [Bui19, Han17, Hen13a, Lou16a, Lou17, Pyl10, Ree16a, Smi10b, Wag15, Elm13, FJB14, Gar09a, Gar09b, Har08c, Kin13, Kle13]. Boerhaave [Roo12b, Pow12]. Bohemia [Gli11, Sim10]. Bologna [Cap14, Man14]. Boltzmann [Bad19, Tat18]. Bomb [Pha10, Abr09, Mag12]. bombardier [Ekh10]. Bombay [Asp11]. bombing [Ver19]. Bone [O’C17, Alb16]. Boner [Oos12, Tar15]. Bones [Man17, Ree16d]. Boogert [Mil11d]. Book [AN18, Ada10, Ade10, Adl16, Aga13a, Aga14, Aga15, Aga17a, Aga17b, Ako19, Alb11a, Alb15, Alb16, Ale11, And12a, And12d, And12c, And12b, Arb17, Arn16, Art11, Ase17, Ayl17, Ayl18, Ayl19, Bad16, Bad19, Bak11, Bak14, Bak15b, Bak15a, Bal14b, Ban14a, Ban15, Ban12, Bar13a, Bar13b, Bar15a, Bar15b, Bar16, Bar17, Bar15c, Bec12, Bec16, Bec18a, Bec18b, Bec18c, BT14, Bel12, Bel17, Bel18b, BZ11a, Ben19a, Ber11a, Ber12, Ber15, Ber17, Ber18, BS16, Bha11, Big10, Big18, Bil15a, Bil17, Biv10, Bix19, Bjö11, Bjö14, Bla17, Boa10, Boa12, Boa15, Boa18, Bod18, Bod19, Boo11, Boo12, Boo14, Boo17, Bou15, Bow16, Bow18, Bow10, Bow17, Bro11a, Bro16, Buc11, Bud11, Bud13a, Bui19, Buk14, Bur11, Bur13, Bus18, Cal12, Cal13, Can15, Can16, Can19b, Can19a]. Book [Cap14, Cap15, Car16a, Car18, Car16b, Car11, Car12, Cha14, Cha10, Chi12, Chr10, Cle14, Cle15, Cle16, Cle17, Cle18, Coe10, Col13, Con15, Coo11, Coo14, Cow10, Coz15, Coz16, Coz17, Cul11, Cul13, Cun10b, Cuo12, Cuo13, Cur18, D’A16, D’A18, Dan19, Dar11, Das18, Daw17, Daw10, Daw12, Daw14, Daw15a, Daw15b, Daw15c, Daw16a, Daw16b, Daw16c, Die15, Dih16, Dih17, Dij12, Dob14, Doh14, Doh16, Dra17, Dry10, Dub10, Duf18, Dun11, Dun12a, Dun12b, Dun14, Dun15a, Dun15b, Dun16a, Dun16b, Dun17a, Dun17b, Dys14, ES16, Ebe17, Ebe18, Edw14, Egm12, Egm13, Eic12a, Eic12b, Eic15, Elt12, Ens12, Esp10, Eva13, Eve13, Eve18, Fal18, Fal13a, Fan12, Far10a, Far10b, Far11, Far12, Far13, Far15, Far16, Far19, Feo18, Fer11a, Fin10a, Fin10b, Fin10c]. Book [Fin11, Fin17, Fix16, Fix19, Fla19, Fon12, For10, Fot18, Fra17, Fre10, Ful18, Ful10a, GP18, Gar11, Gar14, Gas10, Gib14, Gif16, Gil15a, Gil16a, Gil12, Gil16b, Gil11, Gis19, Gli11, Gli12, Goo10a, Goo11a, Goo12a, Goo18, Goo17, Goo10b, Goo12b, Gou12a, Gou13a, Gou13b, Gou16, Gou17, Gou12b, Gou12c, Gou14, Gra11, GG12, Gre14b, Gri15, Gul18, Hah16, Hai16, Hal10, Hal14b, Hal17, Ham14a, Ham10, Ham14b, Ham15, Ham16, Han17, Han16, Har16, Har10a, Hay18, Hay19b, Hed13, Hei15, Hem19, Hen14, Hen15a, Hen15b, Hen10, Hen13a, Hen17, Hen11a, Hen11b, Hen13b, Hen19, Hig10, Hig11, Hig13, Hig14a, Hig15, Hig16, HA16, Hoc14, Hoc16a, Hol13a, Hol14a, Hol15, Hol16a, Hol16b, Hol10, Hol16c, Hom16, HA19b, Hoq10, Huf16, Hug10]. Book [Hul10, Hun18, Hun19, Hus15, Ins19, Jac15, Jam10a, Jam11a, Jam11b, Jam12, Jam13, Jam17, Jam18, Jar10, Jar16, Jar17, Jim19, Joh15, Joh13a, Joh11, Joh13b, Joh14, Joh16, Jon10a, Jon10b, Kaa18, Kao12, KK14, Kee12, Kee17, Keh12, Kel17, Ken13, Ken19, Kid11a, Kid11b, Kip16, Kir13, Kir15, Kne10, Kni15, Kni16, Kno14, Köb16, Kot16, Ksi13, Ksi15, Ksi16, Kum16, Kum18, Kus15, Kus18, Lan11a, Lan12, Lan15a, Lan19, Lan14, Lan15b, LN14, Lau15, Law17, Lea16, Lee15b, Leh12a, Les13b, Les13a, Les14, Les15, LD18, Lin17, Lo15, Loi16, Lou15, Lou16a, Lou16b, Lou17, Lou19, Lyk16, Mac11b, Mac13, Mac10, Mac11a, Mae12, Mai12a, Mai11, Mai12c, Man13a, Man13b, Man17, Man16, Mar16, Mar17a, Mar18a, Mar12a, Mar15, Mar12b]. Book [Mar12c, Mar14a, Mar17b, Mar18b, Mat13, May17, McC19, McG15, McO10, Mca16, Met11, Mey11, Mic11, Mil11a, Mil11b, Mil12, Mil13, Mil15, Mil16, Mil17, Mil11c, Mil11d, Mis11, Mod17, Mon10, Moo17, Moo13, Mor10a, Mor18, Mor15b, Mor13, Mos16b, Mos16a, Mox16a, Mul13, MY18, Mur12, Mye12, Nal12, Nal15, Nas16, Nav14a, Nav15a, Nav15b, Nav16, Nes10, Nic17a, Nic17b, Nic19, Nor12, O’C17, OP16, Oos12, Opi14a, Osl10, Par11, Par12, Pay16, Pea18, Pea12, Pea19, Pet10a, Pha10, Pic10a, Pic10b, Pin12, Pin13, Pis16, Pis17, Pos14, Pos15a, Pri12, Pri19a, Pri13c, Pri19b, Prk17, Prk18, Pro11, Pum12, Pyl10, Qui12, Rad13, Rad14, Rad17, Raj15, Rap11, Rap12, Ree14, Ree16a, Ree16b, Ree17, Ree18, Ree19a, Ree15, Ree10, Reg17, Rei15, Ren19, Ren10]. Book [Ren15, Ric12, Ric11b, Ric18, Ric10, Ric14, Ric16, Rie16, Rih12, Rob10a, Rob14, Rob15, Rob17, Rob10b, Rod15a, Rod15b, Roe17, Roo12a, Roo12b, Roo12c, Roo15, Ros17, Ros16, Røs15b, Røs15a, Rot16, Row10, Row14, Rud13, Rus11a, Sac16a, Sac16b, Sac19, Sal14, Sal15, San12, San13, San14, San15a, San16, San17, Sar13, Sar14, SJ19, Sau13, Sau15, SS11, Sch17a, Sch15a, Sch15b, Sch16, Sch17b, Sch19, Sea10, SS12, SS13a, SS15, SS16, SS17, Ser12, Ser14b, Ser14a, Ser15, Sha10, Sha11, She13, Shi16, Sho14, Sho18, Sia18, Sim14, Sim17a, Sim13, Sim16, Sim17b, Sin15, Siv10, Sle10, Sle11a, Sle12, Sma13, Smi10b, Smi10c, Smi11d, Smi11c, Smi11b, Smi12, Smi14, Smi16b, Smi16a, Smi18a, Smi18b, Smi19, Smo10, Sno10, Söd10, Som11]. Book [Som15, Sør17, Sta11a, Sta11c, Sta11d, Sta12b, Sta14, Sta15, Sta16a, Sta19c, Ste10, Ste11b, Ste14b, Ste16, SM19, Sto14, Sto12, Str11, Stu10, Stu11, Stu17, Sum11, SJ11, Tam14a, Tam16b, Tam16a, Tar12, Tar15, Tat17a, Tat17b, Tat16, Tat18, Tau11, Tay14, Tee11, Tes11, Tho16a, Tho19, TS14, Tho16b, Til10, Tim12, Tos11, Tot10, Tra16, Tun19, Tur16a, Tur10, Tur14, Tur15, Tur16b, Tur12b, Tys17, Und12, Und13, Val14, Val11, Van11b, Van11c, Van14, Ver15a, Ver16, Vol16, Vos11, Wad11, Wad19, Wag14, Wag15, Wai18, Wal16a, Wal18, Wal19a, Wal11a, Wal12c, Wal16b, Wal17, Wal15, Wal11b, War13, War19, Web10, Web11, Web13, Web14, Web15, Web17, Web18a, Web18b, Wes15, Wes12b, Whi15, Whi16, Wig11, Wil11a, Wil11b]. Book [Wil16a, Wil12, Wil14, Wil16b, Win15, Win11, Win13, Wit11a, Wit11b, Wol11, Wol14, WM11, WM18a, Wri16, Wyl11, Yal17, YM10, Zai16, dC14, vN19, vL13a, vL13b, vL15, vL16, Kus11, LMSR09, oRLM^{+}09]. Books [Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano15i, Bec18b, Bel12, Cha14, Doh14, Far12, Far15, Fla19, Gar11, Hol16c, Kni15, Tho19, Wal18, WM18a, HWM12, Nas10, RCK10, Sec14]. Boole [Val10]. Borchardt [vL13a]. Bord [Mor10a]. borders [THB12]. Bordogna [Ste10]. boredom [Don14]. Bostanci [Ayl18]. Boston [AN18, Ben19a, Car16a, Car18, Dob14, Dun17a, Fon12, GG12, Pha10, Sma13, Sto12]. botanic [Hic19]. Botanical [Coe10, Mul13, Ric18, SS13a, Smo10, Ble12, Lac09, Mue11]. Botanist [Gil16a]. Botany [Bil17, dC14, Bal19, Coo12, Feo19, Kus11, Bec16, Egm13]. Bouk [Wil16b]. Boulter [Hay18]. Boulton [Tom12a, Tom13a]. Bound [Bak14, Mar17b, ML^{+}12]. bound/PLC [Mar17b]. Boundaries [Ste10, Bor08, Røs15a]. boundary [Kem19]. Bourke [Mil15]. Bouvier [Sim13]. Bowlby [Vic11]. Bowler [Boo11, San15b]. Box [Bod19]. Boydell [Hoc16a]. Boyle [Mox16a, Aga13b, Hun09, Ric14, Sno10, WM18b, Mox16a]. Bradley [Bud13a, Fis10]. Brain [Smi18b, Sta16a]. Branch [Smi10a]. Brandon [San12]. Brandsma [Lou16a]. Brandt [San17]. Bras [Arn16]. Break [Fer11a, Bus10]. Breast [Kaa13, Raj15]. Breasted [Car12, Abt11]. Bredekamp [Hai16]. Breed [Pea19]. Breeding [Adl16, Ber12, Der12, Der15]. Breidbach [Gli11, Gli12]. Brenni [Dun17a]. Brepols [Boa10]. Brewer [Lou16a]. Brewing [Sum13, Bak15b]. Brian [Coz17, Dun15b, Dun16a, Nor12, Smi18a, Wal12c]. Bridger [Tur16b]. Briefwechsel [GG12, RRL12]. Briggs [Tys17]. Brighton [Man13b, Stu11]. Brigitte [Buk14, Kee17]. Brill [AN18, Bak14, Bak15a, Ben19a, Car16a, Car18, Dob14, Dub10, Dun14, Dun17a, Fal13a, Fon12, Gou16, Hig13, Hig15, Loi16, Mos16a, Pha10, Pos14, Roo12a, Sha11, Sho14, Sma13, Smi16b, Sto12, Van11c, Ver16, Wri16]. Brinkman [Wyl11]. Britain [Bod18, Bow10, Doh16, Far11, Hoc16a, Kaa18, Kee12, Kel17, Ken13, Lou15, Mis11, Mor10a, MY18, Mye12, O’C17, Pos14, Ree19a, Ric12, Ros16, SS15, Smi11b, Sta19c, Tat17b, Wil11c, AG17, Alb11b, Bea11, Ber13, Boo11, Bor09, Bow09, BW14, Cow14, Doh17, Elm13, Eps14, Gou12a, Har09, Hen13a, Hes11, Hig14a, Jam10a, Jon13, JA15, Lig09, LMS13, Lor13, Man15, McL12, Mis13, Moo10, Mos14, Nav10, Ort09, Pea19, Qur11, Sch17b, Sem12, SS12, Smi10a, Ste16, Sto14, Szu18, TV11, Tom13b, VM18, Vol16, Wad11, Wal19b, Wal12a, Wat14, Wes15, Wit10]. British [And12b, Ayl19, Ber18, BS16, Daw10, Daw15c, Doh14, Dun15a, Far19, Hug10, Jar16, Keh12, McA13, McC19, Mis11, Ree10, Rot16, Smi16b, Sta14, Tho16b, And09, Big10, Bin12, Cha12, Cor12, Erl11, Erl13, Heg11, Hig14b, Hog12, HL12, Hol09, Hug12, Hut08, JJ14, Jon16b, Jon16a, Lau12, Mac15, Mag12, Mur14, Mye17, OG17, Pod13, Ree16d, Ren12, Røs15b, Sch08a, SS13b, Toa11, Toa12, Wil15a, Wil15b, Wit10, Zai16, Joh13a]. Brno [Gli12]. Brock [Kip16]. Bronn [Web10, Gli08]. Bronowski [Des12]. Broomhall [Bod18]. Brosco [Ser15]. Brotóns [Car16a]. Brown [Goo12b, Jon10b, Sta12b, Und12]. Browne [LD18, Rad13]. Bruce [Mil11a]. Brunswick [Whi16]. Brush [Nav16]. Brusius [Die15]. Brussels [Sim13]. brutal [Mad16]. Bryan [Gro17]. BSHS [Keh12]. Buchwald [Ksi13, Mur12, Sch15a]. Bud [Pic10a]. Buffalo [Ber12]. Bugge [Hig11]. Bugs [Cla09, Sle10]. Building [Gil16a, Lee15b, DiM13, Hig14b, SV17, Nic19, Ree17, Web18b]. bulk [Coh18]. Bunn [Joh13b]. Bur [AGF13]. burden [Spi15]. Burke [Stu11, Ros10]. Burkhardt [Daw16b]. Burlington [Big10]. Burnham [Bow16]. Burns [Pos15a]. Burnstock [Fin10b]. Busch [Fer11a]. Business [Daw12, Web18b, Fyf12, Wal18]. Buss [Smi19]. Butler [Tur16a, Gil15b]. Butterfield [Jam12, Ben11]. butterfly [Zal14]. Button [SJ19]. Byrne [Sau13, Ksi15].

C [Aga13a, Bow16, Bro16, Hal10, Jam17, Joh15, Joh13b, Kus15, Par12, Rob17, Roo12b, Roo12c, Sch15b, Ste14b, Tur10, Vol16, Wes12b, Bor09, Bro11b, Kas14, Leg13, Mor10a, Nie14, OS13, Ros16, Sim19a, Sim17b, Som11, Sta12b, Wol09]. C. [Gou13b, Tur12b]. c.1704 [Sch15b]. c.1890 [MY18]. CA [Far19, Lee15b, Lou19, Sin15, Smi11d, SM19]. cabinet [Hei18]. Cabinets [BT13, Dun14]. Cadwallader [Rot16]. Caen [Ver19]. Caesar [Car18]. Cahan [Hem19]. Cain [Alb15, Ric10]. Calculation [TS14, Cam12]. calculator [Mor15a, CC11b]. Calendar [Tam14a, Per12]. Calendars [Rob10b, Leh07]. Calgary [Rob15]. California [Boo17, Chi12, Lou19, Mul13, Smi14, Smi14, Wes12a]. Calinger [Bec18c]. Callahan [Aga14]. Calvi [Tun19]. Calvin [Ban12]. Cambria [Bud13a]. Cambridge [Aga15, Alb16, And12d, And12c, Art11, Bad19, Bar15b, Bel17, BZ11a, Boo17, Bud11, Cha10, Cuo12, Cuo13, Dar11, Daw14, Daw15a, Daw16b, Dih17, Dun16a, Fin10a, For10, Gar14, Gas10, Gou12b, Gri15, Hal14b, Har10a, HA16, Hoq10, Huf16, Hus15, Jam10a, Jam12, Jam18, Joh15, Joh16, Jon10b, Kee12, Kum16, Lan19, Lan14, LN14, Lea16, Les15, Lo15, Lou16a, Mai11, Mar12b, Met11, Mil12, Mod17, Nal15, Nav14a, Nav15b, Nic17a, Nic17b, Osl10, Par11, Pea12, Pic10b, Pin13, Pri12, Prk18, Rad13, Ree14, Ree17, Ren10, Ren15, Ric16, Rob10b, Roo12c, Sac16b, San15b, San17, Sar13, Sar14, SJ19, Söd10, Som15, Sta11c, SJ11, Tam14a, Tat17a, Tat17b, Tat16, Tau11, Tho19, Tho16b, Tot10, Tur10, Tur16b, Vol16, Wad19, Web10, Wes12b, Wol11]. Cambridge [HJ13, HR09, JL14, OPR11, Pet14, RR09]. Camiller [Ren10]. Camilleri [Sta11c]. Camp [Fot18]. Campaign [Bur12]. Campe [TS14]. can [Mer19, Swe15]. Canaan [Fer11a]. Canadian [Esp10]. Canales [Coz17, Ric11b]. Cancer [Daw17, Ham14a, Lan15a, Loi16, Raj15, TlT12, AF16, C^{+}12, Kaa13, Tim13, Ham14a]. candle [FJ11, Mor13]. canine [BW14]. Cans [Arn16]. Cantor [Mac13, San15a]. Cape [Rob15, Und13, Cap13]. Capelotti [Rob15]. Capital [Joh15, Jac14, Qui16]. Captain [Fer11a, Bus10]. capture [Par18]. Cardano [Ech11]. Cardiff [Hun18]. Care [Bow16]. Career [Gil16a, Rei15, EKD^{+}13]. Career-Building [Gil16a]. Carin [Ree19a]. Carl [Gil16a, GG12, Mar18a, Mar19, RRL12]. Carlo [Kus18]. Carlsbad [Vog10]. Carola [SJ11]. Carolina [Hay19b, Mar18b, Ree16b, Wai18]. Caroline [Win11, Mca16]. Carolus [Coo11, Egm10]. Carolyne [Lou16a]. Carrara [Feo18]. Carson [Sta11c]. Carsten [Ham14a, Lan15a]. Carter [Fre10]. Cartesian [Dih16, Dar14]. cartography [Sva18]. Casalini [Big18]. Case [Bar13a, Car11, Pic10b, Tay14, Cow08, Gol10, O’B13, Olo14, War12]. Casey [Lea16]. Casper [Smi18b]. Cassidy [Wes12b]. Castle [Kir15, Mol14]. Catalogue [Lou15, Roo12c, Jon13, OPR11]. cataloguing [Csi17]. Catching [Ter14, ES16]. Category [Bar17]. Catelijne [Lo15]. Caterina [Kus18]. Catherine [Sør17, Daw15b, Kno14, Røs15b, Sha11]. Catholic [Wal15]. Catholicism [Ade10, Cul13, Ade10, HA08, O’L06, O’L12]. Catholics [FW15]. Cathryn [Sta11c]. Cathy [Bod19, Smi10c]. Cause [Gou13a, McO10, Des09, Ken12]. Cavendish [Law15b]. Cavert [Loi16]. Cayleff [Fra17]. CDN$29.95 [Jam11b]. Cécile [Ver19]. Celebrity [Far16, Hem15]. Celestial [TB13, Van13, Tar15]. Celina [Far10a]. Cell [Cap14, Sin15, Ben13a, MR12]. Celtic [Cul13, O’L12]. censorship [Gal12, Kir17, Tar13]. Center [Tim12]. Centre [Edw14, Und13, Mon12]. Centres [GCA10, Egm12, Rob11, Wit11b]. centrifuge [Kri12]. Centuries [Hen17, BFGG11, GH13, Hig19, Bar13a, Sia18, Sim14, Sim13, War12]. Centuries-Long [Hen17]. Century [And12a, Ayl19, Ban14a, Bla17, Boo11, Coz17, Cul11, Doh16, Dub10, Duf18, Dun14, Egm13, Eic12a, Fin10a, Fix19, Gar14, Goo10b, Gou12a, Hig10, Hoc16a, Hun19, Ken13, Kir15, Kni16, Man16, Mar12b, Mic11, Mor15b, Mye12, Nav16, O’C17, Ree19a, SS12, Smi14, Stu11, Tam16b, Wad11, War13, Wes12b, And13b, AM11, Ban14b, Bea11, Bee19, Bel18a, BT13, Bri10, Bux18, CR14, Car09, Cas11, Cla14, Col16, Cow14, Csi17, Dar12, Don14, End18b, Fox12a, Fra16, Gep12, Gib15, Jon14b, Kaa13, KP14, KT19, LMS13, Mol14, Mos14, Pac12, Pet14, Pha13, RC14, RC11, Sch15c, Tam14b, Tar13, Ter14, TW10, Wal19b, Wes12a, Wil15a, Wil11c, Yal15, Yam15, ZR15, Ade09, Aga13a, Alb11b, Alb15, All12, ABS10, Bat10, Ber11b, BS16, Bow09]. Century [BS15, Can19a, Cla09, CC11a, Dij12, ES16, Fre10, Gas10, GMWR13, Gil16a, HSSR13, Hol15, Hui09, JS08, Kno14, Kus11, LW11, LMSR09, Mar12a, McA13, oRLM^{+}09, Mül11, Mur10, Pos14, Pye12, Qur11, Rad14, Rec13, San15a, San10b, Sch08b, Sem12, Ter10, Vol08, War19, Wyl11, Zai16, dC09, Bak15a, Raj15]. Certainty [Art11, Gui09]. Ch [KK14]. Châtellier [Boa10]. Chabás [Fal13a]. Chakrabarti [Sta14]. Challenge [SS11, Dal10]. Challenges [Ser14b, A^{+}12]. Challenging [Ver16, Dho15]. Chalmers [Sum11]. Cham [Ako19, Nic19, Sia18]. Chambers [Daw12, Fyf12]. Chancellor [Rad13]. Chaney [Hay19b]. Chang [Ser14b, Boa15]. Change [Bon11, Lan14, Les15, Win15, FSS^{+}14, WC13, Wis12, Oos12]. Changed [Dun17a, Row10, Edg09]. Changing [Sho14, Bur12, HKZ12, Pos15b]. Chapel [Hay19b, Mar18b, Ree16b, Wai18]. character [Fri11]. Charles [And12a, CJ10, Daw16b, Lou16b, Sim16, Sim17b, Bra12, CV09, Cun10b, Van09, Dil11, Dil12, Dil13, Gis19, Jim19, Kir15, Les13a, Mis11, Mit18, Rad13, Sta19a, Wol11, Yan13, Bud13a]. Charlotte [Hig10, Cal12]. Charlottesville [Sle12]. charting [SM12]. charts [Sys18]. Chasing [Sch13, Tur15]. Chatto [Alb11a, Bak15b, Boa12, Coo11, Cow10, Cul11, Daw15c, Gil11, Goo12b, Ham15, Hom16, Hul10, Jac15, Ken13, Lou15, Mar12a, Rad14, Raj15, Ric12, Sal15, Sch16, Smo10, Tho16a, Wit11a]. Chemical [Boa12, Mor13, Roo12b, Sto14, FJ11, Löw11, McE10, Pow12, Wal12a]. Chemicals [Ree16b, Ste14a]. Chemist [Mar12a, Mil08, Row12, Win13]. Chemistry [And12d, Kip16, Mor18, Rob10a, Roo12b, Sac16a, Edd08, Jol11, Löw11, Pow12, RH12, Lyk16]. Chemla [Cuo13]. chez [Nav10]. Chhem [Ser14b]. Chicago [Aga17a, Alb15, And12a, Arb17, Ayl18, Bal14b, Bar13b, Bar17, Bec12, Bel18b, Ber15, Boa18, Bod19, Boo11, Boo14, Bou15, Bow17, Buc11, Cal13, Can16, Cap15, Car12, Cle14, Coe10, Con15, Coz15, Daw12, Dra17, ES16, Ebe17, Ebe18, Egm13, Eve18, Far16, Gil16b, Goo12a, Goo10b, Gou12c, Gre14b, Hai16, Hem19, Hen17, Hol14a, Ins19, Jam17, Kel17, Ksi16, Kus15, Lan15b, Lou16b, Mac11a, Mae12, Mai12a, Man17, Mar16, Mar12c, Mon10, Nal12, O’C17, Pin12, Prk17, Pro11, Rad14, Rad17, Rap12, Ree18, Rei15, Ren10, Rie16, Rih12, Roo12b, Roo15, Sac16a, Sac19, San12, San14, San15b, San16, Sch19, SS12, SS13a, She13, Shi16, Sho18, Smi10c, Smi16a, Ste10, Ste14b, Stu10, Tam16b, Tes11, Tur14, Val14, Van14, Wal16a, Wal19a]. Chicago [Wal15, Web10, Web14, Web15, Web17, Web18a, Web18b, Whi15, Whi16, Wil12, Wil16b, Wyl11, vN19]. Chichester [Ham14b, Pea18, Sch17a]. chickens [Der12, Ber12]. Chico [Far19]. Child [Sle11a, Shu10, Vic11]. childbirth [AG17, Kir17]. Childhood [Wai18, RSA^{+}17]. Children [Pro11, Koh10, Nie14]. Chile [And12c, Med11]. chimie [Jol11, Hed13]. Chimisso [Til10]. China [Mul13, Pos14, Chi12, Goo10b, Jin19, LMS13, Mar16, Mou17, Mue11, Mul10, Sch08b]. Chinese [D’A16, SM19, Hud14]. Chitra [Die15]. Cholera [Ham09, Biv10]. Chotkowski [Boo14]. Chris [Joh13a, Nic17a]. Christensen [KK14]. Christian [Chr13, Cle17, Dih17, KK14, SS16, Lar14]. Christianity [Jam11b, Hes09]. Christina [San17, Til10]. Christine [Gil16a, Law17, Les13b]. Christoph [Hus15, Car16a, Qui12, Str11]. Christopher [Gou14, Bar16, Bel12, Biv10, Fre10, Tot10]. chromosome [Rus14a]. chymical [Yam15]. Cieli [Pri19a]. Ciencia [GC09, Ric11b, Ric11b]. Cimento [Wil11a]. Cinchona [Sim17a]. cinema [Ell17, Kir11, OG16, OG17, Gou12b]. cinematic [Kir17]. ciphers [Bia12]. Circle [Fix19, GP18]. Circulating [Few10]. Circulation [Pos14, Wri16, LMS13, Mar19, Raj10, SM12, Vog10]. Citizenship [Wol14, Ter13]. Citti [Col13]. City [Hol16a, Dud14]. Civic [Alb11a, Fin09, FW15, Lun13]. civil [Smi10a]. civilization [Bar14, BF13, Ksi13, Mil17]. claims [RC11]. Claire [Ada10, Bla17, Lou15]. Clarendon [Eve13]. Clark [Sle10]. Clarke [Sch15b, Kas14]. Classic [Rad17]. Classical [Rob10b, Wag15, Kin13, Leh07]. Classification [Chi12, Nal15, Dic13, McM18, Mul10, Ros18]. Classiques [Gis19, Ver15a]. Claudine [Man13a]. Claus [Sho14]. cleanliness [Bux18]. Clerc [Rab17]. Clercq [Hig11]. Clericuzio [Wil11a]. Clerk [FMW14, Hen15b, Lam11, Sta12a]. Clinic [Fot18]. Clinical [Nes10, Kic09]. Clio [Sal14]. Clock [Hen17]. cloth [Bad19, Bec18a, Bec18b, Bec18c, Bel17, Bel18b, Bod19, Bow17, Cur18, Das18, Die15, Dun16b, Ebe17, Ebe18, Gil16b, Hay19b, Hem19, Hen19, Ins19, Kel17, Man17, McC19, Mod17, Nic17a, O’C17, Prk17, Ric16, Sac19, San16, San17, SJ19, Sch19, War19, Web18b]. Clower [Gou13b]. Club [Jar17, Wal19a]. Clubs [Wol14, Ter13]. Clunies [And18]. Clusius [Coo11, Egm10]. CNRS [Bus18]. co [Kri12]. co-production [Kri12]. Coat [Hen15a, Hal14c]. Coats [Gou12b, Kir11]. Cocco [Cle14]. Coelestis [Ver10]. Coen [Val14]. Coetlogon [Dun12a, Lov10]. Coffin [Fer11a, Bus10]. cognition [Edd17]. Cohen [Huf16, Man13a, Whi16, Fei19, Sta19b]. Cohen-Cole [Whi16]. Coil [Lou17]. Coimbra [Big18]. Cold [Aga15, Cal13, Har16, Hol13a, Lin17, Nic19, Rei15, Rob17, Sar13, SC12, Tur16b, Whi16, OK14, Tur12a, Bri15, CC14, EKD^{+}13, Hec11, Sol13, TR14, Wol13, Wri16, Mat13]. Colden [Rot16]. Cole [Stu10, Whi16]. Coleridge [GP18]. Colin [And12d, Mor15b].

Collaboration [Aga14, Ins19, Mod17, Røs15a, Ada16, Car18, KCM13]. Collaborative [Buk14, LOvT12]. Collapse [Lan11a]. Collected [Sau13, BB12]. Collecting [BM11, War19, Par12, Ros17]. Collection [Kus18, Sim10]. collections [HWM12, KZ15, Lip14, Bil17, Doh14, Wal16b]. Collectors [Ham10, And08]. College [Big18, Cul13, Mor18, Rud13]. Colleges [O’L12]. Collegium [Gil16a]. Collins [Gri15, Nic17b, vN19]. Colonial [Hal17, May17, Asp11, San10b]. Colonialism [Dub10, Hui09]. Colonies [Rob14, McC12]. colour [Die15]. Columbia [Das18, Mil16, Pea18, Ren15, Sea10, Tys17]. Comas [RC11]. Combe [Dyd15, Jen15]. comes [Bro10b]. Comet [Car16a]. Comfort [Ser14a]. Coming [Mul10, Chi12]. commentary [BB12, Sau13]. Commerce [Whi15, Pet13]. Commission [Stu11, Ros10]. Commoditization [Bak11, Wig10]. common [Dyd15]. Communication [Ber18, Smi16b]. Communism [Lin17]. Communist [Mai11]. Communities [Ade09, Sir13, Sør17, Ste16, Bel18a, BH19, Wal19b, Cul11, Gib14]. Community [Nic19, Web17, JS17, Win19a, Sch19]. Companion [Daw16a, Hoq10, Hus15, JL14, Sch17a, HR09, RR09]. Company [Smi10b, Noa14, Win19a]. Comparative [Ako19, Huf16, Nas16, Coh15, Leg13, Pod13]. Compass [GJ09a, GJ09b, Wal11b]. Competing [Lan14, Wol13, Wis12, Hol13a]. Complex [Vol16, Eps14]. Complexity [Bow11]. Complicated [Car11, Gol10]. Composing [Gou17]. composition [Fis10]. Computer [Boo17, Nic17a]. Computing [Ban15, San13, GS12, Mah11, Nov12, Pri12, Tat17b]. Concentration [Fot18]. conceptions [Del09]. Concepts [Sho14, HKZ12]. conference [A^{+}12]. Confession [Sch19]. confidentiality [Bia12]. conflict [Toa11, Toa12, Yal11]. conflicts [Wot12, Sal14]. Conformist [Hen13a, Elm13]. Congresses [Hig14a, Mis13]. Conimbricensis [Big18]. Conjectures [Fis10]. conjunction [Elt12]. Conjuring [Tra16]. connections [And13b]. Conquered [Hol16c, Han15]. conquest [Pet14]. Conscience [Und12, Bro12]. Consciousness [Hay19b]. consensus [Dic13, Nal15]. Conservation [vN19]. Constitution [Jen15]. Constructing [BT14, Mai12b]. Consumers [McG15]. Contagious [Ste16]. containing [See12]. contemplationes [Cla14]. Contemporaries [Jac15]. Contemporary [Bur13, Gli12]. Contentious [Stu10, LCMJ08]. contest [Jon14a]. contested [Wat14]. Context [Wil11a, Leg13, San10b]. Contexts [Sal15, MHO13]. Contingency [Ayl17]. Contingency/Inevitability [Ayl17]. Continuity [Oos12, Win15, Bon11, FSS^{+}14]. Continuum [Ade10, Fan12, YM10]. contraceptive [Szu18]. contraceptives [Löw11]. contraddizione [Pri19a]. contribution [Wal11b]. Contributions [Bud13a, Lyk16, Bra12]. Control [Loi16, Sto14, Ale08, Dev16, Hil12, Wal12a, Dry10]. Controversy [Arb17, Jam10a, Mon10, Mur12, Nal15, Nav15b, BJ10, Dic13, Jon16b, Ort09, Pum11, Ree09, Waz14, Dun15b, GMML14, Wal17]. Conway [Gif16]. Cook [Ric18, dC14]. cooperation [CGC13a, CGC13b]. Coopersmith [Aga17b]. Coopmans [Lo15]. Copenhagen [Hun19]. Copenhaver [Dun16a]. Copernican [Nay14, Van13]. Copernicus [Van11c, Mos16a]. Coral [EJ13, Bar15c]. cordis [Neu18]. Corinne [Lou16a]. Cork [Cul13]. Cornelis [Bui19]. Cornell [Dun12a, Rot16, Row10]. corporation [Mox19]. corps [Mit18]. correctly [Spi15]. Correspondence [AJ12, Daw16b, Gli12, Jam13, Kne10, Sch15b, Sto12, CMC08, FF11, Kas14, RRL12, San11, Š^{+}11, vTSŠ^{+}11]. Corrigendum [CGC13a, Rod15b]. Corsi [Kne10]. Corvallis [Pis17]. Cosmic [Gou16, Pri15]. Cosmography [Tes11, Por09]. Cosmological [Tat16, Dan14]. cosmology [Bon11, Oos12]. Cosmopolitanism [Pis17, Lip14]. Cosmos [Coe10, Dun16b, GS13, Mar15, Sch19, Gou17, Wal09]. Costa [Gas10, Les15]. Could [Ayl17]. Council [Jon16a, McG19]. Counterculture [Aga17a]. County [Mar17a]. Coupain [Ree14]. Couples [Buk14, LOvT12]. Courtier [Sho18]. Courtney [Web18a]. Cover [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14c, Ano14a, Ano14d, Ano14b, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano18b, Ano18a, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Bec18b, Sch19, Web18b]. coverage [Nie10b]. Cowan [Pic10b]. Cowie [Hoc16a]. Cows [Sle12, Rit10]. Craciun [War19]. Cracking [Tay14, O’B13]. Crackpots [Nor12]. Craft [Smi16a, Ver15b]. Crafting [Set10, Dar11]. Craig [Mos16b]. Crane [Wil16a]. craniology [Eff17]. craniometry [Goo16]. Cravens [Mat13]. Crawford [Sim17a]. Creager [Cap15]. created [Hol13b, Opi14a]. Creating [Ber12, OR08, SF11, Der12, Cha10, San12]. Creation [Boo12, Car12, Sal14, Yal17, Abt11, Erl11, Erl13, Nie10a, Wot12]. Creationism [Bla14, Bro16]. Creative [Goo12a, BJW11]. Creators [Lea16]. creatures [Str19]. credible [Mag12]. credit [Fra16]. Creese [Tau11]. Crimean [Nes10, Kic09]. crimes [Ros10, Stu11]. Crippa [Fix19]. Cristiano [Ben19a, Big18]. Cristina [Bec16]. Critical [Daw15b, Mar12c, Gat17, HMLE15]. criticism [Gai17]. critics [Lig09, Bow10]. critique [Chr10, Bus18]. Crook [Man13b, Ros16]. Cross [Bel17, Bro11a, Wil14, BZ11b, Llo09, Mac09]. Cross-cultural [Wil14, BZ11b, Llo09]. Cross-currents [Bel17]. crossing [Mor17]. crossroads [Mou17]. Crown [Dun16b]. Crowther [Met11, HA19a, Jon16a]. crusade [Ene14, Rat16]. Cryptozoology [Nor12]. CT [Ade10, Bod19, Fer11a]. Cuba [Mar12c, Von11]. Cultural [Bud11, Fin11, Goo12a, Hah16, Hol16b, Jam10a, Mos16a, OP16, Rad14, War13, Yal17, BZ11b, BJW11, CGC13a, CGC13b, Dew14, Fox12a, L^{+}10, Llo09, MWR12, Ort09, SS10, Sha12, Wil14]. Culture [Alb11a, Cle14, D’A18, Dun16a, Ell10, Far11, Far16, Kno14, Mac10, Man13b, Man17, Mar12a, McC19, MY18, Mye12, Rap12, Rot16, Rud13, Sac16b, SS13a, Ser12, Sta12b, Tam14a, Tur15, Wai18, Wig11, Zai16, Alb09, AD12, ABS10, Ban11, BM10, Bin12, Ble12, Bro11b, Coc13, Cro12, Fin09, FW15, Har09, Hem15, Hog12, HL12, Hug12, LZ14, Mar11, Mül11, Pac12, Per12, Ray15, Sch13, Wil11c, Daw15a, Hig10, Lou17]. Cultures [Jam10a, Nas16, Par12, Pha10, Ree19a, Abr09, BM11, Doh17, KT19, Mag12, Ort09, Rec13, Sel08, Sta16b, Siv10]. Cunningham [Ber11a, Gar11, Wit11b]. curating [Nie14]. Cure [Smi10b, Har08c]. Curie [Far16, Hem15]. Curiosities [Wad11, Alb11b, Kus18]. Curiosity [BZ11a, Ros17, Tys17, Huf11]. Curious [Ebe17, War19]. current [San11]. currents [Bel17]. Cursus [Big18]. Curtain [Bar16, AG17, Hol14b]. Cuts [Mye17]. cutting [Mye17]. Cybernetic [Med11, And12c]. cycle [Mar19]. Cycles [Egm12, Rob11]. cyclonic [Mah18]. Cypess [Ebe17].

D

[Bak14, Cle17, Dun17a, Elt12, Hen19, Lau15, Nal12, Rei15, Rob10a, Roe17, Roo12c, Sta11d, Swe15, Tur14, Wil11b, Wyl11]. D.S [Lou16a]. d’abaco [Ekh10]. Daguerreotype [Bel17]. daily [Nie14]. Dallal [SS11]. d’Almeida [Mit18]. Dalton [Gro17]. Dan [KK14, Wil16b]. Danger [Ros16]. dangereuses [Ada16]. Dangerous [Bar15a, Fix16, Ale14]. dangers [SM12]. Daniel [Bou15, Car18, Fin17, Mos16a, Bec18b, Can15, Gre14b, Pyl10, Ros18, Wit11a]. Daniela [SS13a, vL13a, BW13, Par12, Smi11d, vL13a]. Danielson [Tat16]. dans [Boa10, BFGG11, Sim13, Vol08]. Danske [Hun19]. Darin [Dun16b]. Dario [Tar15]. Dark [Pea12, San10a]. Darrigol [Dih16, Dij12, Tat18]. Dartmouth [Rud13]. Darwin [Bil15a, CRB09, Daw16b, Eic15, Goo12b, Jim19, Pea18, Rad13, Ric10, San15b, Smi11c, And18, Bow17, Bro10a, CV09, Van09, Daw10, Des09, Dil11, Dil12, Dil13, DR09, End16, Hay18, HR09, Hol09, Hoq10, Koh08, LD18, McO10, Pri13a, Ric13b, San15b, Sch10, Smo10, Sta19a, Van11b, Yan13]. Darwinian [Bow10, Lan15b, Feo19, Ric13a]. Darwinians [Lig09]. Darwinism [Web10, Bil15b, Del09, Gli08]. darwinisme [Del09, Esp10]. Darwins [Smo10, Van11b, Ayr08, Pat09]. Daryn [Rih12, Rob10b]. Dassow [Coe10, Mar14a]. Daston [Hen10, Rad17, Rei15]. Data [Web15, vL16, Csi17, Hal14d, Van11a, Ste13]. Data-Driven [Web15, Ste13]. date [Mim15]. Daughters [Ray15, Sac16b]. David [Fin10a, Gra11, Hen14, Mac11b, Ric18, Smo10, Aga17a, And12a, And13a, CJ10, Dri10, For10, Gou12b, Hem19, Hig10, Ins19, Kum16, Mar12a, McC19, Rie16, San15a, Sch16, She13, Sør17, Tau11, Tur16a, Wes12b]. Davide [Fix19]. Davies [Ebe18]. Davy [Boa18]. Dawn [Ale10, Ben19a, Kni15, Lau15, Par11, Smo10, Ayr08, McR14, Sec14]. Dawson [O’C17]. Day [Tee11, GT09]. Days [See12, Wil16b]. Dayton [Gou12c]. DC [Wal12c]. Deal [HA19a]. dealers [Pod13]. Dean [Die15]. Death [Han17, Jon10c, Sta11d, Del10]. Debate [Fix19, Mur12, BJ10, Hes09, San11, Jam11b]. Debates [Hir11, Mos16a, Pes12, Roo12a, Rab17, SM12, Vol08]. Debating [Tam16b, Yal15, Bow17]. debats [Vol08, Boa10]. Deborah [Val14]. Debt [Jim19]. Deceived [End16]. December [CJ10]. Deception [Whi15, Pet13]. decision [Wyn11, Wal12c]. decision-making [Wyn11, Wal12c]. d’Edition [Chr10]. Deep [Ayl19, Ksi16, Sta15, Bul16, Nie10b, Rud14, Tur13]. Defection [Cal13, Tur12a]. defence [Pes12, Smi10a, Spi11, Spi12b]. Defending [Gou10, Pum12]. Defining [Web14, Phi12]. definition [Tar13]. del [Pri19a, Wil11a]. Delbourgo [Ros17]. DeLee [Gai17]. Deleted [San15b]. Delft [ZR15]. Delhi [Bha11]. Delia [Smi18b]. Delisle [Esp10]. Della [Tar13]. Democracy [Das18, Mat13, SC12]. Denis [Roe17]. Denise [Web14, Hol10]. Denmark [Hal10, KKNN08]. Dennis [Dun12a, Lov10, Tat16]. Department [Gli12, Mor18]. depleted [San15b]. Derr [Les13a]. Derry [Adl16, Ber12]. Descartes [Bal19, Jol11, Man14, Hed13]. Descended [Ric10, CRB09]. Descent [Con15, Hal14a]. Description [Dun12a, Sta10]. Design [Mai12a, Mar18a, Bux18, McL12, Sta12a, Ber18]. designing [Few10]. Desmond [McO10]. Detail [WM18a]. Detection [Nic17b]. detector [Bun12, Joh13b]. Determined [Fer12, Les13b]. Deutsche [Hen11a, VW10]. deutschen [Uek10, Web13]. Developer [Lea16]. Development [Arb17, Roe17, Sle11a, BW14, Col16, GR18, Mak13, OS13, OR08, Shu10, Spi11, Spi12b, Cha10]. devices [Tur08b, Söd10]. DeVun [Sea10]. Dew [Mil11b]. Dewar [Win13, Row12]. Dewis [OP16]. Dewitt [Daw14]. Dhondt [Ver16]. Diabetes [Tat09, Biv10]. diadem [Cou17]. diagnosis [Wee19]. diagram [BM10, Ser12]. Diagrams [Gul18]. Dialogue [Tho19]. Diane [Mur12, Ser15]. Diarmid [Alb11a, Goo17]. Diary [Les13a, SDS13]. Dick [Nal15]. Dictatorships [Coz17]. Did [Rih12, Leh12b]. Diderot [Bal08, Sha10]. Dietrich [Ber15, Rud13, Tys17]. differ [Pum11]. Difference [Mar18b]. different [Mak13]. Differing [MR12, Cap14]. diffraction [Nav10]. Diffusion [Man13b, Cro12]. Digges [Pum11]. Digitizing [Ban15, Nov12]. Dilemma [Lea16]. DiMoia [Lee15b]. Dinge [Pri19b]. Dinosaur [Wyl11, Bri10]. Diogo [Ase17]. Dioskuri [Gli12, vTSŠ^{+}11]. diplomacy [Ada16, Pod13]. Dirk [Ase17]. Disability [Tur12b, Goo11b, McG19, GP18]. Disarmament [Bur12, Sto14, Wal12a]. Disaster [Ako19, Val14, Coe13]. Disasters [Ako19]. Discarded [Tat17b]. discharge [Mül11]. Disciple [Smi11c, Sch10]. disciplinarity [Bil18, Ree16d]. disciplinary [HJ13]. Disciplinas [Nav14b, Car16a]. Discipline [Ber11a, SV17, Cun10a, Mit18, Pos15b, Sep12, She13]. Disciplines [Llo09, Mac10, Alb09, Ful10a, Nav14b, Prk17]. Discontents [Prk17]. Discourse [Car16a]. Discovered [Ksi16, Pum11, Rud14]. Discoverers [Win11]. Discovery [Dic13, Ebe17, Tur13, SS17, Sta15, Nal15]. Discussing [Ayl17]. discussion [RSA^{+}17]. Disease [Bel18b, Eic12a, Han16, Ken13, SS17, SM19, Str11, Wit11a, Bea11, Ber11b, GF09, Ken12, PB13, Sch11, Tim13, Lan15a, Ser15, Gou13a]. Diseases [Bow18]. disorder [Bea11, Ken13].

Display [Alb15, Jon10b, Ree19a, AKM14, Ber13, Bro09, Mit12, RC14, Hig15, Wag14]. Displayed [Han17]. Disputed [Lan15b, Ric13a]. disputes [RC11]. disputing [Vog10]. Disrupted [Kni16, Gib15]. Dissection [Wag14, Mit12]. dissertation [AK16, Tur38]. distances [Ben19b]. distemper [BW14]. District [Gil16b]. divergent [Del09]. divergentes [Del09]. Divine [Smi11a, Leh19, Eic12b]. Dixon [Rot16, Van11b]. DKK [Hun19]. DKr [Goo18]. DNA [Stu10, Bud13b, GS12, LCMJ08, San13]. Doctors [Mil11d, Few10]. Doctrine [Gou13b, SFFC12, Wee19]. dodo [Law15a]. Doel [Rob17]. Does [Mon13, Leh19, Hol14a]. Dog [Pea19]. dogs [BW14]. Doig [Hei15]. Doing [Dun12b, HW11]. Dollond [Dun15b, GMML14]. Domain [San17]. Domenico [Bel18b, CC11a, Eic12a]. Domestic [Han16]. Domesticating [Goo08, Hei18, Cow10]. Domesticity [Kee17]. Don [Ade10, Cul13]. Donald [Buk14, Kee17]. Dongen [Wri16]. Donington [Joh11]. Donna [Mar12b]. don’t [McG19]. Doogab [San16]. Dordrecht [Rap11]. Double [Hay19b, Hen15a, Hal14c]. double-helix [Hal14c]. Douglas [Lou19, SS15]. Dove [Coz15]. Dr [Fin10c, Gar14]. Dr. [All12, Jay09]. drama [OG17]. drawing [Kem19, AG17]. drawn [Swe15]. Dresser [Wal18]. Dritsas [Mac11b]. Driven [Web15, Ste13]. Drug [Sim17a]. Drugs [Sim14, GH13]. drums [Pet18]. Duda [Hol16a]. Duel [Ale10, Par11]. Duke [Hig10, LN14, Rus11a, vL16, Azz12]. Duncan [Car18, Mye12]. Duncker [Cap14]. Dünkel [Hai16]. Dunn [Köb16]. Dupin [Bud13a, Bra12]. Dupré [Bak14]. Durham [Hig10, Rus11a, Tos11, vL16]. during [Ano17g, Csi17, Egm12, Har16, Mag12, Nic19, Rob11, TR14]. Dutch [Ban11].

DVDs [Fin10b].

e-book [Ren15]. E.

[Lar13, RC11, Sim17b, Ste14b]. E.H [Cle16]. E.M [Alb15]. each [Mer19]. Eagleton [Sha11]. Earl [Kir15, Mol14]. Early [Bel17, Boo11, Can19a, Cle14, Cle16, D’A18, Dob14, Egm12, Eve13, Ham16, Hen10, Hol15, Mos16b, Oos12, Par12, Pin12, Pos14, Roo15, Row14, RH12, Sac16b, San12, Sch16, Sch17b, Shi16, Sho14, SK12, Tur12b, Wag15, Web17, Wil11b, Wit11a, AGF13, Bec15, Bin12, BM11, Bon11, Bow09, Coc13, DS08, End18b, Fox12b, Goo11b, HKZ12, KP14, Kin13, Lam11, Law15a, LMS13, Mac15, Man12, Mar14b, Nav10, OG16, Opi14b, Par18, Raj10, Ray15, Rob11, RC11, Sch11, SF11, Swe15, VB11, Wer13, Wil15b, Yam15, Yeo14, ZR15, And12d]. Early-Modern [Pos14, LMS13]. Earth [Bak14, Far12, Hig10, Rus11b, Web18a, Bas14, Poo10, ABS10, Bur12, Fer11b, Ksi16, ML^{+}12, Rad13, Ren15, Rud14, WM11]. Earth-Bound [Bak14, ML^{+}12]. Earthquake [Val14, Coe13]. Earthscan [Wal12c]. East [Mak13, Win19a]. Eastern [Rob10b, Leh07, Noa14]. Eating [Spa12, Ste14b]. eb [Nav14b]. Ebenstein [Mar17b]. Echoes [Pri15, Gou16]. Echterhölter [Web14]. Eclecticism [Web17]. eclipse [Pet14]. Ecole [Boa10]. Ecological [Hay19b]. Ecology [Ham14b, And13a, Ayr12, Hen14]. Economic [BT14, GR18, Mai12b]. economies [Wil15a]. Economy [BT14, Joh15, Jac14, Mai12b, Ren11]. Ecstasy [Car11, Gol10]. ed [Aga13a, Ako19, Ayl19, Bar13a, Bar15c, Bha11, Bod18, Bur13, Can19a, Car18, Cuo13, D’A18, Daw16a, Dob14, Dun11, Dun15b, Egm12, Eic15, Fot18, Gil15a, Joh11, Kee12, Kid11b, Kir15, Les15, LD18, Mil15, Mor13, Oos12, Pha10, Pin13, Rad13, Ree15, Røs15b, Row14, San15b, Sar13, Sch16, Siv10, Smi14, Smi16b, Söd10, Ver15a, Ver16, Wag14, Wri16, vL13b]. Eddy [Rob10a]. Eden [And12c, Law15b]. Edge [Tat17b, Mye17]. Edgerton [Row10]. Edinburgh [Daw10, Edd08, Elt12, Les13b, Rob10a, Stu11, Bat10, Mye17, Ros10, Stu11]. Edison [Sch13, Tur15]. Editing [Wal19b]. Edition [BW13, Car16a, Hoq10, Les15, Sau15, Siv10, Tim12, Gat17, GG10, HMLE15, LG11, WC13, Daw15b, Mar12c, Chr10]. Editions [Bus18, Man13a]. editor [Ano17g, GG10, YM10]. Editore [Col13, Tar15]. Editorial [Ree19b]. editors [Daw16b]. Edizioni [Cle18, Kne10]. Edmund [Rad13, Rad14, San15a]. edn [Ren19]. eds [AN18, Aga15, Aga17a, And12a, Ayl17, Bak14, Bak15a, Bec12, Ber15, Bjö11, Bou15, Bro16, Bui19, Buk14, Can19b, Cap14, Cle16, Col13, Coz17, Cuo12, Daw15a, Daw15b, Daw16b, Daw16c, Die15, Doh14, Dun12b, Dun14, Dun17a, Dys14, Ebe18, Eve13, Far10b, Fin10a, Fin11, Fin17, Gar11, Gli11, Gli12, Goo12a, Goo18, Goo17, Gou12c, Gou14, Gra11, GG12, Hai16, Ham14a, Ham16, Har16, Hen15b, Hen10, Hen11a, Hig10, Hig11, Hig15, Hig16, Hoc14, Hoq10, Hus15, Jac15, Jam13, Kee17, Köb16, Kot16, Law17, Les13a, Lin17, Lo15, Lou16a, Lou19, Mar12b, Mar12c, Mar14a, Mat13, Mic11, Par12, Pay16, Pet10a, Pos14, Qui12, Rad13, Rad14, Rad17, Ree16a, Ree19a, Ric11b, Ric10, Rob17, Roo12c, Ros16, Røs15a, Rud13, Sal15, San12, San15a, Sau13, Sau15, Sch15a]. eds [Ser14b, Sho14, Sia18, Sim14, Sim13, Sim16, Smi11d, Sør17, Ste11b, SJ11, Tam16a, Tar12, Tar15, Tho16a, Tot10, Wal11a, Wal16b, War19, Wil11a, Wit11b, YM10, vN19, vL13a]. Educated [Rud13, AD12]. Education [Big18, Hoc16a, Lyk16, Van11c, Wit11b, Wol14, Cow14, Edd17, Gra15, GCA10, Hei18, Lau12, Ter13]. Educator [Bud13a, Bra12]. Edward [Bil17, Sto12, Jar16]. Efficiency [Dry10, Ale08]. Efram [Ayl19, Daw15c]. Eggheads [Nor12]. Egmond [Coo11, WM18a]. Egypt [Mur10]. Egyptian [Mur12, BW13, BJ10, Gre14b, Mos14, Sto13]. Egyptians [BW13]. Egyptologist [Car12, Abt11]. Egyptology [Man13b, Cro12]. Ehrenhaft [San11]. eight [Van11a]. Eighteenth [Bar13a, Doh16, Dub10, Dun14, ES16, Gar14, Gas10, Gil16a, Gou12a, Ken13, Kni16, Mor15b, Pac12, Raj15, War19, Bea11, BT13, Gib15, Hig19, Jon14b, Kaa13, Ter14, Wil15a, Yam15, All12, Hui09, Kno14, Mur10, San10b, Sem12, Ter10, War12, dC09]. Eighteenth-Century [Doh16, Dub10, Dun14, Gar14, Gou12a, Ken13, Kni16, Pac12, Bea11, BT13, Gib15, Yam15, All12, Hui09, Kno14, Mur10, Sem12, Ter10]. Eigner [Ste11b]. Einleitung [BW13, vL13a]. Einstein [Joh14, Mil11a, Hun10, Hus15, Ill12, JL14, Nav15b, San11, Waz14]. Elaine [Tar12]. Elborg [Str11]. Eleanor [Jon10a, Pet10a]. Electra [Ell17]. Electric [Sim13, Sta15, BFGG11, Ker14, Tur13]. Electrical [GS08, YM10, Bax13]. Electricity [Cow10, Goo08, Mah09, Goo11a]. électrique [BFGG11, Sim13]. Electroacoustics [Jam18]. Electroconvulsive [Wal17]. Electrodynamics [Fal18]. Electron [Nav10]. Electronic [Bec18b]. Elena [GG12]. Elephants [Gil15a, Pri11]. Elisabeth [Hen11a]. Elite [Jon14a, Rob14, McC12]. elites [Llo09]. Elixir [Bec18a]. Elizabeth [Fin10b, Ham14a, Loi16, Han17, Sch17b]. Ellen [Ebe18]. Elliot [Cro12, Man13b]. Elliott [Wig11]. Ellis [Kaa18, Mur14]. Elmer [Hen13a]. embedded [Pye12]. Embryo [Das18]. Embryologie [Web11]. embryos [Wel17]. Emergence [San16, Yi15]. Emiliano [Ayl17]. Emily [GP18, Keh12]. Emotion [Mil15, Bod14]. Emotional [FSS^{+}14, Win15]. Emotions [Bod18, Bou15, Eic15, BG14, Ric13b, Lou16a]. Empathy [Hoc16a, Cow14, Bod19]. Empire [Ben19a, Ber18, Hoc14, Hug10, Opi14a, Rot16, Sar13, SS12, SS13a, And09, Ble12, Hec11, Hol13b, Kel12, McA13, Qur11, Wil15b, Yam15, And09, AS12, Goo17, McA13, Tam16b, Tho16b, Vog10, Yal15]. Empires [Köb16, Smi11d, B^{+}09]. Empiricism [Fre10, Car09]. Encounter [Goo17]. Encounters [Hal17, Sal15, War19, MHO13, Pri11]. Encyclopaedia [Sel08, Siv10]. Encyclopedia [Dun12a, San15b, Lov10, Sta10, Dun12a]. Encyclopédie [Hei15, Doi13]. Encylopédie [Doi13]. End [Cun10b, Sea10, DeV09, JS08, McA13, Web08]. Endt [Bar15c]. Endt-Jones [Bar15c]. Engine [Ins19]. Engineer [Bud13a, Rap11, Bra12, Mai12c]. Engineering [Gif16, Goo11a, San16, Yi15]. Engineers [Wis12, Lan14]. Engines [Kel17]. England [Bow18, Daw15b, Hal14a, Han16, Hig11, Ken19, Mca16, Mit12, Rob15, Sta12b, Tee11, Wag14, Wal11b, Alb11a, All12, Bla17, Bro11b, Cap13, Cla09, Con15, Eag10, Ell10, Fox12b, Gar14, GJ09a, GJ09b, GT09, Mac16, Nay10, PdCB10, Row14, Sha11, Wig11]. English [Roo15, Mon13, Yeo14]. Enlightened [Bil15a, Pri13a]. Enlightenment [Ber11a, Dun12a, Ell10, Far12, Ful18, May17, Mil12, Mit12, Rot16, Sta10, Ste14b, Vos11, Wag14, Lev09, Tom12b, Bec18c, Ber13, Ble12, Cun10a, Dun16a, Far10a, Far19, Fer11b, Fox09, Jar10, Rat09, Sac16a, Spa12, Web17, Wig11, SS13a]. enrichment [Kri12]. Enrique [Far10b]. Entangled [Hec11, Sar13]. enterprise [Ras14, Web15]. Enterprises [Köb16]. Entertainment [Tra16, Cow14, Hoc16a]. enthusiasm [Joh12]. entities [WM18b]. entomology [Wal19b]. Entropy [Bad19]. Environment [Rad13, Wil16a, Col16, Pos15b, Roe17]. Environmental [Roe17]. Epigenetics [HA16, Ber17, Jar17]. Epilogue [GS18]. Epistemic [Can19a, Lan19, Col16, ML10a]. Epistemology [Mey11, Rhe10]. Epistolary [Gib14, Sir13]. Epitomised [Gar11, RCK10]. Época [Car16a, Nav14b]. Epstein [Vol16]. Equation [Bad19]. equilibration [Wil15a]. Era [Coz16, Lip14, Mag12, Pod15, Cur18]. Erasmus [Bil15a, Pri13a]. Eric [Kid11b, Bad19, Kem19, Tys17]. Erich [Gli12, Š^{+}11, vTSŠ^{+}11]. Erickson [Rei15]. Erik [Mul13, Gif16, Jar17, Tho16b]. Erling [Gil12]. Ernst [Ree14, Web10, Gli08, Ric08]. Erratum [Dil13, Gam18]. Erwin [Moo17]. Esbester [Ros16]. espaänola [Nav14b]. Española [Car16a, Ric11b]. Esposito [Ham15]. ESS [Røs15a]. Essary [San15b]. Essay [Mar12c, Von11, Fei19, Rhe10, Sta19b, Mey11]. Essays [AN18, Dys14, Sle12, Tam16a, RVG12, Rit10, Car16a]. Essential [Gou12c, PDG11]. establishing [Boo15b]. Estella [Mar17a]. età [Cle18]. Ether [Can19a]. Ethics [Tur16b, Bri15, Gab16]. Ethnic [Chi12, Sva18, Mul10]. ethnographic [Pet18]. Ette [Mar12c, Mar14a]. Etudes [Boa10]. Eufrasia [Kus18]. Eugenics [Ber17, Les13b, BL10, Fer12, Ren11, Bjö11, Wal15]. Eugenio [Gou13b]. Euler [Bec18c]. EUR [Buk14, Cap14, Col13, Cul13, Fal13a, Web14]. EUR104.00 [Car18]. EUR105.00 [Dob14]. EUR115.00 [Smi16b, Wri16]. EUR125.00/$150.00 [Dun17a]. EUR129.00 [Fon12, Van11c]. EUR135.00 [Car18]. EUR149.95 [GG12]. EUR150.00/$205.00 [Mos16a]. EUR170.00 [Kus18]. EUR184.00 [Sto12]. EUR19.50 [Ric11b]. EUR19.80 [Mar14a]. EUR20.00 [Pri19a]. EUR217.00 [Sho14]. EUR22.00 [Bus18]. EUR229.00 [Bak15a]. EUR23.00 [Man13a]. EUR24.00 [Hed13]. EUR248.00 [AN18]. EUR25.00 [Cle18, Kne10]. EUR28.00 [Mil17].

EUR32.00 [Tar15]. EUR34.95 [Egm12]. EUR36.80 [vL13a]. EUR37.84 [Sau15]. EUR38.00 [Ser14b]. EUR39.00 [Web11]. EUR39.90 [Hen11a]. EUR40.00 [Chr10]. EUR44.00 [Mic11]. EUR44.99 [Hig16]. EUR45.00 [Loi16]. EUR46.00 [Wal11a]. EUR48.00 [Gli11]. EUR49.00 [Dun12b, Gou16]. EUR49.95 [Web13]. EUR59.80 [Sta11a]. EUR65.00 [Boa10]. EUR74.95 [Pri19b]. EUR74.95/£ [Pri19b]. EUR79.00 [Pha10]. EUR85.00 [Chr10, Eic15]. EUR93.00 [Gis19]. EUR95.00/$110.00 [Ben19a]. EUR99.00 [Bak14, Dub10, Hig13, Roo12a, Sha11, Sma13]. EUR99.80 [vL15]. Europe [Gil11, Hen10, Köb16, Mod17, Wit11b, Ako19, Bec15, BT13, Ber11a, Bla14, Bra15, Bro16, Cun10a, DS08, Dun14, Goo11b, GCA10, HKZ12, Lin17, Lyn10, McG15, Sch16, Shi16, Sho14, Sta16a, Tur12b]. European [Aga13a, Buc11, Car18, CGC13a, Fal13a, Far10b, Joh15, Wil11a, CGC13b, CG12, Gep12, Jac14, Ker14, PNGP09, Wer10]. Europeans [Ase17]. Eva [Far16]. Evaluating [GS10, Qui12]. evangelical [Sta12a]. Eve [Met11, Tar15, TB13]. Evelyn [Rus11a]. ever [THB12]. Everest [Val10, Heg13]. Everett [Sau13, BB12]. Evert [Tho16a]. Everyday [Arn16]. Evidence [Boa15, Kin13, Wag15]. Evocative [Tur07, Söd10]. Evolution [Ayl19, Con15, Jam11b, LZ14, Mil17, Pea18, vN19, Bar14, Bil15b, Col16, Cow13, Del09, Hal14a, Hes09, Hol09, Van11a, Daw10, Daw15a, Del09]. Evolutionary [Hen14, Lan15b, Lig09, Ric10, Rus11b, San15b, She13, Smo10, Web10, And13a, CRB09, Edd17, Koh08, Ric08, Ric13a, Sep12, Bow10, Rad13]. Evolutionist [Ben13b, Les14, Smi16b]. exceedingly [Pum11]. Excellence [Wit11b, GCA10]. Excellent [Tys17]. exceptionalism [Dav19]. Exchange [Goo17, Yal17].

Exchanges [Par12, BM11]. Exchanging [Ase17]. Exclusion [Wal12c, Wyn11]. Exegesis [Tat18]. Exercise [Gar14, All12, Gra15, Neu18]. Exercitationes [Car18]. Exhibiting [Cow14, Hoc16a]. Exhibition [Can11, Goo17, Mac13, Lau12, Pri11]. Exhibitions [SS12, Cou17, Qur11]. existence [BG11]. Existential [Cur18]. Exotericae [Car18]. Exotic [Mca16]. expansion [Doi13, Hei15]. Expedition [Far12, Dew10, Fer11b, Few10, Nie10b]. Expeditionary [Mac11b, Dri10]. Expeditions [Mil11c, Rob15, SS13a, Ble12, Cap13, Kla09, NHR12, Gou14]. Experience [Gib14, Mul13, Sta19c, Hog12, Mue11, Sir13]. Experiences [Ako19]. Experiment [Eic12a, Kus15, Ber11b, Hun13, Ins19]. Experimental [Ber11a, Boa18, Dun14, Far19, Heg13, Nes10, San17, BT13, Cun10a, Cur14, Erl11, Erl13, Kic09, Mit18, Mox16b, Win19a]. experimentation [Mer12, Mer13]. Experimenting [Rad16]. Experiments [Dun12b, Fal18, Fin10c, Joh14, Sta15, Wai18, Gra15, HW11, Ill12, Jay09, Sch15c, Tur13, Das18, Man16]. Expertise [Tho16a, SM12, VPW15]. Experts [Gri15, Cla09, Col14]. Explained [Huf16, Coh15, Rav11]. explaining [Dav19]. Exploiting [Ase17]. Exploration [Coe10, Lou16b, Mul13, Row14, Fox12b, KWB15, Lac09, Lüt19, Mue11, Sum14b, Gif16, Sim16]. Exploratory [Fal18]. Explored [San17]. Explorer [Ben13b, Les14, Cap13, Rob15]. Exploring [Rob17]. Expositions [Jon10b, Bro09]. extinction [Cow13]. Extraordinary [HA19b, Lam13, Som15]. Eye [Rud13, AD12, WM18a]. Eyes [Mil11d].

F [Sle10, Tur12b]. F. [McR14]. F.E [Lau15]. Faber [Jam11a]. Facil [Ham16, Pay16]. facing [Tam14b]. Facsimile [Les15, WC13]. facts [Lüt19, Swe15]. Faidra [Far10b]. Failed [Cha09, Sum11]. Fairs [Wol14, Ter13]. Faith [Ban12, Lar14, SS16]. Fall [Aga17b]. Falling [Tur08a, Söd10]. Fallwell [Hom16]. Family [Ree14, BCH^{+}13, Hog12, Szu18]. Fangerau [Ser14b, Sta11a]. Far [San15b]. Fara [Bix19]. Faraday [Fal18, Lam11, Mor13, Jam10b]. Fari [Car16b]. Farmacopea [Cle18]. Farming [Mar16]. Farnham [Bak11, Bec16, BT14, Ber11a, Bil15a, Bur13, Coz17, Dun15b, Dys14, Gar11, Goo17, Gou12a, Gou13a, Hig14a, Jam13, Mai12c, Mox16a, OP16, Røs15b, Row14, Sch15b, Sim16, Sto14, Stu17, Tar12, Tur12b, Wag14, Wag15, Wal16b, Wil11b, Win13, Wit11b]. Farrar [Fix16]. Fascimile [Fin10a]. Fascism [Wad19]. fascismo [GC09, Ric11b]. Fascist [Wad19]. Fashionable [Ken13, Bea11]. Fashioned [Hay19a]. Fashioning [Doh16]. Fate [Wal16b, KZ15]. Father [Hen14, And13a]. favourite [Ano17g]. Fax [Aga17b]. Faxed [Aga17b]. Fay [Lou17]. Fco [Coz17, Ric11b]. fear [Wea12, Hal14b]. feared [Hog12]. feather [CGC13a, CGC13b]. February [Fie10]. Feder [BTW09, Wal11a]. Feeling [Bod18, Win15, FSS^{+}14]. Feingold [Ksi13]. Feld [Uek10, Web13]. Femininity [Jon09, Ada10]. Fennî [Bil15b]. Ferber [Rod15a, Rod15b]. Ferdinand [Gli12, Š^{+}11]. Ferguson [Les13b]. Fernando [Pin12]. Ferreiro [Far12]. Fever [Fre10, Car09]. Few [May17]. Feyerabend [Kid11b, Kid11a]. Ficino [Gou16, Pri15]. fiction [Cor12, End16]. Field [Gou14, Mon10, Boo15a, Bro10b, NHR12, Ree09, Uek10]. Fieldwork [Gou14, NHR12]. Fifteenth [Fin10a, LMSR09, oRLM^{+}09]. Fifteenth-Century [Fin10a, LMSR09, oRLM^{+}09]. Fifty [Aga14, Rad17, KCM13]. Fight [Gai17, Reg17]. Figural [Pin13, Spi12a]. Figure [Bak15a, HSSR13, Sha12]. Film [OG16, Ell17, OGE17]. filo [Cle18]. Filosofia [Car16a, Nav14b]. Fin [Sta16a, Chr10, JS08, Bra15]. Fin-de [Sta16a]. Fin-de-Siècle [Bra15]. Final [Sac19]. Finding [Liv15]. Findlen [Kus18]. Fine [Joh11, Mar09]. Finger [Gou13b]. fingerprinting [LCMJ08, Stu10]. Finnegan [Alb11a, Goo17]. Fire [Gra11, HHR09]. Fireworks [Wer10, Buc11]. firm [BCH^{+}13, Ree14]. First [Bix19, Fon12, Hig13, Joh15, Ric18, HA12, Jac14, Mad16, Nay14, Roo11, Ros18, Asp11]. Fish [Sta15, Tur13, Arn16]. fitness [Swe15]. FitzRoy [And18]. Flamsteed [Spi15]. Flavia [Pri19a]. Fleming [Ree16a, Ric16]. Fletcher [Sto12]. flight [Doh17]. Flood [Hen15b]. Flora [Rob15, Cap13]. Florence [Col13, Tar15]. Florian [Gli12]. Florida [Smi19, Mur14]. Florike [Coo11, WM18a]. Florilegium [Ric18]. Floris [Fei19, Huf16, Sta19b]. Flower [Opi14a, Hol13b]. flows [Hil12]. fly [Wil15a]. Fogel [San12]. Folkes [Roo17]. Following [Ter10]. font [RC13, Ver15a]. Fontes [Gas10]. Food [Mil16, Ste14b, Smi15, Spa12, Web18b]. force’ [Lam11]. Forces [Nav14a, Sch12]. Forensic [BS16]. forensics [O’B13, Tay14]. Foreword [Rad13]. Forgan [Bud11]. Forged [Rie16]. Forgotten [Hen15a, Rob15, Cap13, Hal14c, SM19]. Form [Dob14, Hol10, Web11, Gig09, Man12]. Formation [Wil16a]. Fors [Sac16a]. Forster [Str11]. Forum [Ano19g]. Fossil [Pod13, She13, Goo16, Sep12]. fossile [Coh11, Man13a]. Fossils [CMC08, Kne10]. Fostering [Tom12a, Tom13a]. Foundation [Dun12a, Hei15, Sha10, dC14]. Foundations [Ric16, Whi16, Sol13]. Fox [BDR13, Die15, Far10a, Pis17, Row14, Rus11a, Sch15a, Sim13, War13]. Frackowiak [Fin10b]. Fragments [Bal14b, Win12]. France [Bud13a, Esp10, Kee12, Sør17, Tam14a, Til10, Vos11, Wil11b, Bra12, Chi08, Dew09, Fox12a, Joh11, Lev09, Mar09, Mil11b, Per12, Sim19b, TV11, Tra16, War13]. Francesca [Ste10]. Franceschi [Hig16]. Francesco [Col13, Gou16, Cla14, Pri15]. Francis [Daw16c, Eve13, Goo10a, Ric14, SK12, Aga13b, Dun15b, GMML14, Lun13, Mar15, Mer12, Mer13, Mor17, Ren11, Wee19]. Franciscans [Bec18a]. Franco [Fin10a]. Frank [Mor13, Bou15, Lou16a, Mor13, Web13]. Frankfurt [Pri19b]. Franklin [Bax13, Jam11a, Lam09]. Franz [Dun12b, Gli11, Mic11, Ser14b, Wal11a]. Fraser [Sim16]. fray [Ech11]. Frederick [Daw16b]. Fredrik [Gil16b]. French [Gil15a, Ada16, BFGG11, Del09, Dew10, Eff17, JS08, Jol11, Lip14, Mit18, Pri11, RC13, Ver19, Vol08]. Freud [Gal12, Smi12, Tau10]. Frevertet [Win15]. Frey [Mar14a]. Friedrich [Car18, GG12, RRL12, Fal18]. fringe [Gor12, Tur14]. Frith [Fin10b]. Fritz [Ren10]. Frixione [Gou13b]. Fröhliche [Hen11b, Rec10]. Front [Ano10b, Ano10d, Ano10f, Ano10h, Ano11b, Ano11d, Ano11f, Ano11h, Ano12b, Ano12d, Ano12f, Ano12h, Ano13b, Ano13d, Ano13f, Ano13h, Ano14c, Ano14d, Ano14f, Ano14h, Ano14j, Ano15b, Ano15d, Ano15f, Ano15h, Ano16b, Ano16d, Ano16f, Ano16h, Ano17b, Ano17d, Ano17f, Ano18b, Ano19b, Ano19d, Ano19f]. frontier [Ben13a, Lan19, Sin15]. Frontiers [Pis17]. Frontispiece [Dun17b]. FRS [Bro11a, Mac09]. frühen [Pri19b]. Fruits [Pau07, Cha10]. Fuechtner [Lou19]. ful [Bui19]. fulfilment [Val10]. Fuller [Tos11]. Fullilove [Web18a]. funding [Mye17]. Future [Dih17, McC12, Mon13, Rob14]. Futures [Joh13a, Ren12]. Fyfe [Daw12].

G [Ada10, Boa12, Cap14, Cha14, Gib14, Gli08, Gli12, HA19a, Hoc14, Jam13, Jon16a, Nav16, Ree17, Web10, Wol11, Wol14]. G.

[Esp10, Eva13, Ful10a, Š^{+}11]. Gabrielle [Sar13, Sar14]. Gain [Cha09, Sum11]. Gainesville [Smi19]. Galan [Far10b]. Galathea [Nie10b]. Galen [GWW09, Tot10]. Galileo [Duf18, Ebe17, Rap11, Rap12, Val11, Gat17, Mar11, May10]. Galina [Nes10]. Gallery [Nie14]. gallium [Spi11, Spi12b]. Galton [Lun13, Ren11]. Game [TS14, Cam12]. Gamwell [Hah16]. garage [Cur14]. Garber [Pyl10]. García [San13]. García-Sancho [San13]. Garden [Smo10, Cur14, Koh08, Law15b]. gardeners [Hic19]. Gardening [OP16, Dew14, End18b]. Gardens [Del10]. Gardiner [Hic19]. Garnier [Gis19, Ver15a]. gas [Kri12, Mül11]. Gases [Ree17]. Gaslight [Mil12, Tom12a, Tom12b, Tom13a]. gasping [Heg13]. Gate [Olo14]. Gate-keeping [Olo14]. Gates [Joh09, Mac11a]. Gaudillière [Sim14]. Gaukroger [Lan11a, Pis16]. Gausemeier [Rad14]. Gauss [GG12, RRL12, RRL12]. Gavin [Ree10]. Gavroglu [Tam16a]. Gavrus [Smi18b]. Gay [Mor18]. Gayle [Ren19]. Gayon [Esp10]. Gee [Dun15b]. Gender [Cow10, Hig16, Wil11b, Goo08]. Gene [Ayl18, Ras14, Web15]. genealogical [Ech12]. Genealogische [Web14, Ech12]. Genealogy [Prk17, Eff17, Wal13, Nav15a]. Genentech [Hug11, Wil12]. General [Nic19]. Genes [Ser14a, Com12]. Genèse [Boa10, Vol08]. Genesis [Tur10, Har08a, Vol08]. Genetic [Pic10b, San16, Ser15, Cow08, PB13, Yi15]. Genetics [Ayl18, Gli12, Ban14b, End18b, Har08b, SV17, Zal14, Arb17, Pic10b]. Genetics/‘Mendelianum [Gli12]. genie [Mah18]. Genius [Bec18c, Prk18, Jay09, Fin10c]. Genoa [Ree15, Zip13]. genome [Hil12]. Genomics [Stu17]. genre [End18b, Hig16]. Gentleman [Hun18]. geo [JS08, Chr10, JS08]. geo-heliocentric [JS08]. géo-héliocentrique [Chr10, JS08]. geodesy [Ker14]. Geoffrey [Joh13b, Nav15b, Fin10b, Mac13]. Geographical [Roe17]. Geographies [Byr13, Ell10, Lan11b, LW11, Sar13, Wig11, Hec11, And12a, Nal12, Ksi15]. Geography [Sma13, Ste10, Wit10, Bor08, Hay19a, Rat16, Mis11, Sim16]. geological [CGC13a, CGC13b]. geomagnetism [Rat16]. Geopolitics [Hug10, Ren15, And09, Bas14]. George [Smi11c, Dyd15, Gra15, Jen15, Sch10]. Georgian [Wig11, Ell10, Hic19, Mac16]. Georgiana [Cle17]. Georgina [Sch17a]. Geosciences [Har16, TR14]. Geppert [Aga13a]. Gerbino [Wal11b]. Gere [Bod19, Smi10c]. German [Hom16, Rad14, Web10, BW13, Ech12, Fal13b, Gli08, Kle13, Rec10, RRL12, Uek10, VW10]. Germano [BT14]. Germanus [FF11, Sto12]. Germany [Cap14, Hig11, Kel17, Mic11, PdCB10, Som11, Web14, MR12, Pet18, Phi12, SV17, TW10, Wol09]. Gerovitch [Hol16b]. Gerrit [Ako19]. ghastly [Heg13]. Ghost [HA16]. Ghur [AGF13]. Giambattista [Tar13]. giant [And13b]. Gibson [HA19b, Kni16]. Gideon [Dob14]. Giffard [Kel17]. Gigante [Hol10]. Gill [Tot10]. Gillespie [Bel17]. Gillott [Tur16a]. Gilmore [Loi16]. Gingras [Tho19]. Giorgio [Bak14]. Giovanni [Pri19a]. Girl [Ver19]. Giroux [Fix16]. given [Bar08, Fin10b]. Giving [Val10]. Glasgow [Sim17b]. Glasner [Bur11]. Glennie [Tee11]. Gliboff [Web10]. Global [Aga15, Bas14, BZ11a, Fre10, Goo17, Hol14a, Joh16, Law17, Lou19, Pos15a, Sar13, Sar14, Shi16, Smi18a, Web18a, And13b, Bur16, Car09, GR18, Hec11, Huf11, Mon13, Ogl15, OK14, Fin17, Ren15, Smi16b]. Globalizing [Ase17]. Globes [Sum14b, Sum14a, Dun15a]. God [SS16, Sno10, Wal15, Ben11, Hun09, Jam12, Lar14]. Goddu [Van11c]. Godless [Cul13, O’L12]. Gods [Kot16]. Goeing [AN18]. Goethe [Jim19]. Goettingen [Hig16]. Goldstein [Car11, Sør17, Fal13a]. Golinski [Boa18]. Gómez [Coz17, Ric11b]. Good [Bod19]. Gooday [Cow10, Sta19c, YM10]. Goodey [Tur12b]. Gooding [CJ10, Ric18]. Gordin [Rei15, Tur14]. Gordon [Pos14, Wal11b, Rad13]. Gorée [Dew10]. Gosse [Sal14, Wot12]. Göttingen [Hen11a, Web11, Web13, Web14, vL13a]. Goulding [Pum12]. Governing [Ros16]. government [Mag12]. Govoni [Hig16]. Gowan [O’C17]. Gradmann [Qui12, Str11]. Graeme [Cow10, Sta19c, YM10]. Grafton [AN18, Cro12, Man13b]. Graham [Dys14, Gou12c, HA16, Mar15, Han16, Pri13c]. Graham-Smith [Mar15]. Grains [Hol16c, Han15]. Granada [Car16a]. Grant [Qui16]. Grapí [Dan19]. Grati [Gat17]. Gravitation [Nic19]. Gravitational [Nic17b]. Gravity [Nic17b]. Gray [Hen13b]. Great [Mac13, Rie16, Vol16, KT19, Can11, Eps14, Nas10]. Greater [Bod19]. Greatrakes [Hen13a, Elm13]. Greco [Mur12]. Greece [Tam14b]. Greek [Dij12, Tau11, Cre10, Dar12, Mak13]. Green [Mar16, Gil16b]. Greene [Ren19]. Greenland [Rob17]. Greenough [Law17]. Greenwich [Hig14b, Hig19, Ker14, Spi15]. Greenwood [Ade10]. Gregor [Gar11, RCK10]. Gregory [Fin10b, Fix19, Hay19b, Hoq10, Pro11, Van11b, Far13]. Grell [Wit11b]. Griffith [Mor18]. Grigson [Mca16]. Griset [Sim13]. Groovy [Aga17a]. Gross [Bou15, Wol11]. Group [Rob14, Ban14b, McC12]. Grove [Hun18]. Grown [Jar16]. Growth [She13, Mac16, Sep12]. Grund [Tim12]. Grundlagenforschung [Fan10, Sta11a]. Gruyter [GG12]. Guagnini [Sim13]. Guatemala [May17]. Guatemalan [Few10]. Guenther [Prk17]. Guerre [Chr10, JS08]. Guerrini [Sho18]. Guessoum [Moo13]. Guhr [AGF13]. Guicciardini [Art11]. guidance [Swe15]. guide [Pri11]. Guns [Sør17]. Gur [AGF13]. Gutenberg [Mac11a, Joh09]. Guttmann [Kem19]. Guy [Jam10a].

H [Bec12, Boa15, Bud11, Coe10, Coz16, Eic12b, Gou12c, Gou14, Hig10, Kip16, Les13a, Mil11d, Moo17, Pea12, Prk18, Roo12c, Mag12]. H-bomb [Mag12]. H. [Fei19, Gli08, Huf16, Sta19b, Web10]. Hård [McG15]. Habsburg [Hoc14, AS12, Vog10]. Hacking [Bar15b]. Haeckel [Web10, Gli08, Ric08]. Hale [Con15]. half [Vol08]. Halford [Hay19a]. Hall [Fie10, Hen15a]. Hallam [Han17, Web15, Ree18]. Halle [Web17]. Halpern [vL16]. Hamilton [Mat13, Swe15]. Hamlin [Biv10]. Handbook [Bjö11, BF14, Pet10a, Sch15a, BL10, RS09]. Hands [Pro11, Koh10, Dun17a]. Hands-On [Pro11, Koh10]. Hanley [Bow18]. Hannah [Mor18]. Hanover [Rud13]. Hans [Bro16, Doh14, HWM12, Nav14a, Rad14, Ros17, Ayl18, Cap14, Chr13, Dih17, KK14, Mey11, Sch12]. Hans-Jörg [Rad14, Cap14, Mey11]. Hansen [Hol16c]. Hansteen [Ene14]. Happy [Rec10]. Harbsmeier [Gou14]. hardback [Ada10, Ade10, Aga13a, Alb11a, Alb15, Alb16, Ale11, And12a, And12c, And12b, Art11, Ayl17, Bad16, Bak11, Bak14, Bak15b, Bak15a, Bal14b, Ban14a, Ban15, Bar16, Bar17, Bec16, BT14, BZ11a, Ber11a, Ber12, Ber17, BS16, Bha11, Big10, Bil15a, Biv10, Bjö11, Bjö14, Bla17, Boa12, Boa15, Boa18, Boo11, Boo12, Boo14, Boo17, Bow10, Bro16, Buc11, Bud11, Bud13a, Buk14, Bur11, Bur13, Cal13, Can16, Cap15, Car16a, Car18, Car16b, Car11, Car12, Cha10, Chi12, Cle14, Cle16, Cle17, Coe10, Con15, Coo11, Coo14, Cow10, Coz16, Coz17, Cul11, Cul13, Cun10b, Cuo12, Cuo13, D’A16, Dar11, Daw17, Daw10, Daw12, Daw14, Daw15a, Daw15b, Daw15c, Daw16a, Daw16b, Daw16c, Dih16, Dih17, Dij12, Dob14, Doh14, Doh16, Dry10, Dub10, Dun11, Dun12a]. hardback [Dun14, Dun15a, Dun15b, Dun16a, Dys14, ES16, Edw14, Egm13, Eic12b, Eic15, Ens12, Eva13, Eve13, Fal13a, Fan12, Far10a, Far10b, Far11, Far12, Far13, Far15, Far16, Fer11a, Fin10a, Fin10c, Fin11, Fin17, Fon12, Fra17, Ful10a, Gar11, Gas10, Gib14, Gif16, Gil16a, Gil11, Gis19, Goo10a, Goo17, Goo12b, Gou12a, Gou13a, Gou13b, Gou16, Gou17, Gou12b, Gou14, GG12, Gre14b, Hah16, Hai16, Hal10, Hal17, Ham14a, Ham10, Ham15, Ham16, Han17, Han16, Har16, Har10a, Hen14, Hen15a, Hen10, Hen13a, Hen17, Hen11b, Hen13b, Hig11, Hig13, Hig14a, Hig15, Hig16, HA16, Hoc14, Hoc16a, Hol14a, Hol15, Hol16a, Hol10, Hol16c, Hom16, HA19b, Hoq10, Hug10, Hul10, Hus15, Jac15, Jam10a, Jam11a, Jam11b, Jam12, Jam13, Jam17, Jar10, Jar16, Jar17, Jim19]. hardback [Joh13a, Joh11, Joh13b, Joh14, Joh16, Jon10a, Jon10b, Kao12, KK14, Kee12, Kee17, Ken13, Kir13, Kir15, Kni15, Kni16, Kno14, Köb16, Kot16, Ksi13, Ksi15, Ksi16, Kum16, Kus15, Lan11a, Lan12, Lan15a, Lan19, Lan14, LN14, Lau15, Lea16, Lee15b, Leh12a, Les13b, Les15, Lin17, Lou15, Lou16a, Lou16b, Lou17, Lyk16, Mac11b, Mac13, Mac10, Mac11a, Mae12, Mai12a, Mai11, Mai12c, Man16, Mar16, Mar17a, Mar12a, Mar15, Mar12b, Mar12c, Mat13, McG15, McO10, Mca16, Met11, Mil11b, Mil12, Mil13, Mil15, Mil16, Mil11c, Mil11d, Mis11, Mon10, Mor10a, Mor15b, Mor13, Mos16b, Mos16a, Mur12, Mye12, Nal12, Nal15, Nas16, Nav14a, Nav15b, Nav16, Nes10, Nor12, OP16, Oos12, Opi14a, Par11, Par12, Pay16, Pea12, Pet10a, Pic10b, Pin12]. hardback [Pin13, Pis16, Pos14, Pri12, Pro11, Pum12, Pyl10, Qui12, Rad13, Rad14, Raj15, Rap11, Rap12, Ree14, Ree17, Ree15, Ree10, Reg17, Rei15, Ren10, Ren15, Ric12, Ric14, Rie16, Rih12, Rob10a, Rob14, Rob17, Rob10b, Roe17, Roo12a, Roo12b, Roo12c, Roo15, Ros17, Ros16, Røs15b, Røs15a, Rot16, Row14, Sac16a, Sac16b, Sal15, San13, San14, San15a, San15b, Sar14, Sau13, Sau15, SS11, Sch17a, Sch15a, Sch15b, Sch16, Sch17b, Sea10, SS12, SS13a, SS15, SS16, SS17, Ser12, Ser14a, Sha11, She13, Shi16, Sho14, Sho18, Sim14, Sim17a, Sim16, Siv10, Sle10, Sle11a, Sle12, Sma13, Smi10b, Smi10c, Smi11d, Smi11b, Smi14, Smi16a, Smi18a, Smo10, Sno10, Söd10, Som11, Sør17, Sta11a, Sta11c, Sta11d, Sta12b, Sta14, Sta15, Sta16a, Ste10, Ste11b]. hardback [Ste14b, Ste16, Sto14, Sto12, Str11, Stu10, Stu11, Stu17, Sum11, SJ11, Tam14a, Tam16b, Tam16a, Tar12, Tat17a, Tat16, Tau11, Tay14, Tee11, Tes11, Tho16a, Tho16b, Til10, Tim12, Tot10, Tra16, Tun19, Tur16a, Tur10, Tur14, Tur15, Tur16b, Tur12b, Tys17, Und12, Und13, Val14, Val11, Van11c, Van14, Ver16, Vol16, Vos11, Wad11, Wag14, Wag15, Wal16a, Wal19a, Wal16b, Wal15, Wal11b, War13, Web10, Web13, Web14, Web15, Web17, Web18a, Wes15, Wes12b, Whi15, Whi16, Wig11, Wil11a, Wil11b, Wil16a, Wil12, Wil14, Wil16b, Win15, Win11, Win13, Wit11a, Wit11b, Wol11, Wol14, WM11, WM18a, Wyl11, Yal17, YM10, vL13a, vL13b, vL15]. hardcover [AN18, Aga17b, Ase17, Ayl19, Ben19a, Ber18, Big18, Bix19, Bod19, Bow18, Bui19, Can19a, D’A18, Duf18, Dun17b, Fal18, Far19, Feo18, Fot18, Ful18, GP18, Goo18, Gul18, Kaa18, Ken19, Kum18, Kus18, Law17, LD18, Mar18a, Mor18, Mox16a, MY18, Nic17b, Pea18, Pea19, Pri19b, Prk18, Ree19a, Ric18, Sia18, Smi18b, Smi19, Sta19c, Tat18, Wad19, Wal18, Wal17]. Hardesty [Hei15]. Hardin [Can19b]. Hare [Stu11, Ros10]. Harman [Ber15, Far11]. Harmonic [Tau11, Cre10]. Harmony [Gou17, Sch19]. Harper [Pic10b, Rob17, Tur10]. Harriet [Eve13, Sle12]. Harrington [Smi10b]. Harriot [Row14, Fox12b]. Harrison [Bec12, Can16]. Harry [Gri15, Nic17b].

Harvard [Alb16, Cha10, Dih17, Hal14b, Har10a, HA16, Joh16, LN14, Les15, Nav14a, Par11, Pic10b, Pri12, Sac16b, Tho16b, Tur16b, Vol16, Wes12b]. Harvey [Kus18, Gre14a, Neu18]. Hasok [Boa15]. Hautes [Boa10]. Haven [Bod19, Bui19, Cun10b, Die15, Far10a, Far11, Fin10c, Hol10, SS11, Ser14a, Sle10, Sno10, Wal11b]. Hawking [Bar13b, Mia12, Bar13b]. Haynes [Lou19]. Hayton [Dun16b]. Hayy [Wil14, BZ11b]. hazard [SM12]. Hazel [Sal15]. Head [Wol11, Gro09, Hul10, Jac08]. Healers [Dys14, RVG12]. Healing [Coz15, Fra17, Ham16, Hen13a, Pay16, Elm13, OA14]. Health [Bro09, Bux18, Gar14, Han16, Lee15b, All12, DiM13, Bow16, CC11a, Jon10b]. Hearing [Sta19c]. Heart [Ser14a, Com12]. Heather [Kaa18, Ken13, Som11]. Heaven [Fin10c, Jay09]. Heavens [Hig10, ABS10]. Heaviside [Goo11a, Mah09]. Hecht [Sar13, Sar14]. Hedesan [Cle17]. Heering [Dun12b, Gra11]. Heggie [And12b]. Heidelberg [Boa15, Buk14, Ens12, Hen11b, Lyk16, Oos12, Pum12, Ric14, Siv10, Sum11, Tam16a]. Heilbron [Bad16, Val11]. Heim [SJ11]. Heiner [Ser14b, Sta11a]. Heinous [Stu11, Ros10]. Heisenberg [Hen11b, Rec10, Cam09, Car10, Rec10, Sta11c]. Helaine [Siv10]. Helen [Sho14, Daw17, Hoc16a, Wag15]. Hélène [Bar13b]. Helge [Hal10, Hun19]. heliocentric [JS08]. héliocentrique [Chr10, JS08]. Heliocentrism [Van11c]. helix [Hal14c, Hen15a]. Helmholtz [Hem19]. Helmont [Cle17]. Helmut [Hen11b]. Helmuth [Mic11]. Helped [Jim19, Ver19]. Hemmungs [Far16]. Henk [Ste11b]. Henri [Gra13, Hen13b]. Henrietta [Kus18]. Henrik [Bro16, Sia18]. Henry [Abt11, BDR13, Car12, Die15, Hal10, Hul10, Jac08, Wal18]. Hentschel [Nas16]. her [Opi14b, Vic11]. Herbal [Feo18].

Herbert [Jam12, Ben11, Daw16c, FT15]. Herbs [RVG12, Dys14]. Heredity [Ber17, Col16, Cow08, Har10b, Les13b, Rad14, SV17, Fer12, GMWR13, Jen15, MWR12, Rus14a, Pic10b, San17]. heretical [Liv15]. Herman [Roo12b, Pow12]. Hermione [Kel17]. Hero [Jam11a, Lam09]. Heroes [Ale10, Par11]. Herschel [Win11]. Hesketh [Jam11b, Ric12]. Hess [Ade10, Can15, Sim14, Pic14]. Hessler [Bud11]. heterodoxy [Bax13]. Hevelius [Ver10]. Heymann [Rob17]. Heywood [Mai12c]. Hicks [Gou17, Tat17b]. Hidden [Har10a, Mic08]. Higgins [Gro17]. Higgitt [Köb16]. Higgott [Wal11b]. Higher [CC11a]. Hilary [SM19]. Hill [Hay19b, Mar18b, Ree16b, Wai18]. Hip [Goo10a, ANPR07]. Hirai [Roo12a]. Hiro [Roo12a]. Hispanic [SS13a, Ble12]. histoire [Bus18, Gis19, RC13, Ver15a]. Historia [Fer11a]. historian [Feo19]. Historians [Ban12, Mer19, OG16, RC13]. Historic [Bur13, Loi16]. Historical [Ale11, Ham14a, Mar18a, Pea12, Ree16a, Rod15a, Rod15b, Som15, Cur14, Fer13, FJB14, Lam13, San10a, TlT12, War10, Ako19, Ver12]. Historicizing [Mey11, Rhe10, Ayl19, Goo18]. historiens [RC13, Ver15a]. Histories [Gra11, Kum16, Mah11, HHR09, Roy12, Ham16, Pay16, Pri12]. historiographical [Ver12]. Historiography [Boa12, AF16, McE10, Sia18]. History [Ako19, Alb11a, Alb15, And12b, Bad16, Bar13a, Bar16, Bec18b, Bel18b, BS16, Bjö11, Boa12, BF14, Bur13, Cap15, Cle14, Coo11, Coo14, Cuo13, D’A16, Dan19, Daw17, Daw16a, Dij12, Dun11, Dun12a, Dun12b, Dun15a, Dun17b, ES16, Fan12, Fin10b, Fin11, Fra17, Gil15a, Gli12, GS08, Goo10a, Gou13a, Gou14, Hai16, Har10a, Hen17, Hol15, Ink08, Ins19, Jam12, Joh13a, Joh13b, Keh12, Kip16, Ksi16, Lan15a, Lan15b, Lou17, Lou19, Lyk16, Mca16, Mil16, Moo17, Mor18, Mor13, Mos16b, Pet10a, Pic10b, Rad13, Rad14, Ree14, Ren15, Ric12, Ric10, Sal14, San13, SJ19, Sau15, Sch17a, Sch15a, Sch17b, Ser15, Sim19b, Sim17b, Siv10, Smi10b, Smi18a, Smi18b, Söd10, Sta11d, Stu10, Stu11, Tam16a, Tat17a, Tee11, Til10, Tra16, Tur12b, Van11b]. History [Wad19, Wal16a, Web15, Wes12b, Wil11a, Wil16a, Wil14, Win19b, Wol11, YM10, vL16, ANPR07, BN13, Bal15, BL10, Bas14, Bel18a, BZ11b, Ben11, BCH^{+}13, Bod14, BDS15, Bun12, CRB09, Cas11, Che12, Chi08, Coc13, Coh15, Cre13, Dal10, Dar12, Dho15, Egm10, FJ11, Fin09, Gam17, Gam18, GS12, GT09, Goo11b, Gou10, Gro09, Hal14d, Han15, Har08b, Har08c, Hec17, HW11, Heg11, Hes11, Heu15, Hol14b, Hud14, Jon10c, KP14, Ken12, KKNN08, LaF13, Law15a, Lip14, Lov10, LCMJ08, Mar14b, McE10, McM18, McW11, Mic08, ML10a, Mor15a, Mor10b, Mou17, MWR12, NHR12, Nie14, O’L06, PDG11, PB13, Pri11, RC14, Ren12, RC13, Ric11a, Ric13a, RSA^{+}17, Rit10, Rob08, RS09, Ros10, Rud14, Rus11b, SS10, Sel08, SM12, Sha12, Smi15, Ste13]. history [Sum14a, Sun13, Ter10, Ter14, Tim13, THB12, Tur08b, Uek10, Wal19b, War12, Wer10, Win12, Wot12, Ade10, Bal14b, Bod19, Boo14, Bow16, Buc11, Gou12c, Hah16, Hal10, Hol16c, Huf16, Jon10a, Law17, Man17, Nas16, Pri13c, Pum12, SS11, Sle12, Mil15]. Hitler [Can15, Lan15b, Pic14, Ric13a]. Hittorf [Mül11]. Hjalmar [Sac16a]. Hjermitslev [Bro16]. HMS [Keh12, Ste11a]. Hobbins [Hal17]. Hochadel [Gra11]. Hodge [Hoq10]. Hodos [Fer11a]. Hogarth [Mar18b]. Hogben [HA19a]. Hogg [Zai16]. Hole [Wol11, Gro09]. Holenstein [Bak15a]. Holland [Hig11, PdCB10]. Hollings [Bar16]. Hollywood [Gou12b, Kir11]. Holmberg [Law17]. Holmes [Daw10, Tay14, O’B13]. Holway [Opi14a]. Homburg [Ree14]. Home [Smi10a]. homme [Del09]. Honor [AN18, Dys14, Tam16a, RVG12]. Honour [Fra16]. Hooke [Law15b, Mox16b]. Hope [Pic10b, Cow08]. hopeful [HA19a]. Hopkins [Aga17b, Arn16, Ban15, Bow16, Bro16, Can19b, Cle15, Coz16, Dry10, Eic12a, Fra17, Gib14, Gif16, Ham10, Hol13a, Joh13b, Joh14, Kir13, Leh12a, Mar18a, Mil11a, Moo17, Mos16b, Pea19, Ser15, Sta11d, Sta15, Str11, Tur15, War13, Web15, Wil14]. Horizon [Boo15b]. Horning [Tur15]. Horrocks [Hig13]. Horst [Hai16]. Horstmanshoff [Sho14]. Horticultural [Cha10, Pau07]. Hoskin [Win11]. Hospital [Kem19]. Hoßfeld [Gli11, Gli12]. Houches [Ver19]. House [Cou17]. Howorth [Joh12]. Hubert [Bak15a]. Hudecek [D’A16]. Hudson [And12d, Ric18, Mar17b]. Huff [BZ11a]. Huffine [Smi11d]. Hugh [Ful18]. Hughes [Wil12]. Hugo [Gli12, Š^{+}11]. Huigen [Dub10]. Human [Alb16, Ber17, Das18, Egm13, GMWR13, Goo10b, Goo12b, Hol16c, Joh15, Loi16, Lou17, Man16, Mat13, McO10, Pis16, Ser14a, Whi16, Bro10a, CC14, Com12, Des09, Goo16, Han15, Jac14, Kus11, Mad16, Man15, Sch08b, Sch15c, SC12, Rad14]. Humanism [Feo18, Hir11, Roo12a]. humanist [Des12]. Humanities [Gli12]. Humanity [Eva13, May17, Llo12]. Humans [Ayl19, Sho18]. Humblot [Cap14]. Humboldt [Coe10, Jim19, Mar12c, Mar14a, KOW12, Lac09, Wal09]. Humfrey [Tim12, LG11]. Humphry [Boa18]. Hunger [Hol15]. Hunt [Mil11a]. Hunter [Doh14, Dun17b, Kus15, Mox16a, Sno10]. Hurlbut [Das18]. Hurt [Bui19]. Hutchins [Big10]. Hutchison [Sal15]. Huxley [HA19a]. Huygens [Fix19]. Hvidfelt [Hal10]. Hybrid [Kri12, Sle12, Rit10]. Hyde [Jam17]. Hydrostatics [Ech11]. Hygiene [Wes15, Bux18, Tom13b]. Hypatia [Pum12, Gou10]. Hysteria [Bar17, Gol10, Scu09, Biv10, Car11].

Hysterical [Mic08, Har10a].

Ian [Kus18, YM10, Bar15b, Jam11b, Ric12]. Iberian [CGC13a, CGC13b]. Ibn [Wil14, BZ11b]. Ibn-Yaqzan [BZ11b]. IBSN [McG15]. Ice [Rie16, Rob17]. ichneumon [Wil15a]. Idea [Aga13b, Can19b, Fan12, Gou13a, Ric14, Ken12, McW11]. ideas [Del09]. identities [HJ13]. Identity [Goo10b, Hig14a, Sta12b, Bro11b, Mis13, Sch08b, Hol16b]. idols [Wee19]. If [Win19b]. II [Kel17, Egm12, Mai11, Nes10, Lau15, McR14]. III [Gra15]. Ilkley [DR09, Van11b]. Illinois [Mil17]. Illness [SM19, Mic08, Har10a]. illus [Die15, Mor18]. Illusion [Smi11b, Moo10]. illustrated [Bro11a]. Illustrating [Bel18a]. illustration [Pes12]. Illustrations [Bel18b]. Illy [Joh14]. Iltis [Gli12, Š^{+}11]. im [Pri19b, VW10]. Image [Dun17b, Egm13, Hai16, Wal15, BDS15, Kus11, Kle13]. imagery [BDS15, Hai16]. Images [Hol16b, Smi16a, WM18a, Bel18a, Nas10, Pet18, Ver10, Ver15b]. Imagination [Far19, Cla14]. Imaging [Ser14b, A^{+}12, Man13c]. Imagining [Gep12, San15b, Shi16, Lan19, Aga13a]. Immanuel [Tur14, Gor12]. Imperative [Har16, Liv15, TR14]. imperceptible [WM18b]. Imperial [Bro11a, Mor18, Nes10, Sim17a, Kic09, Lee15a, Mac09, Pet18, Roy12, Bil18]. Imperialism [Fre10, Pha10, Car09, Sin08]. impetus [Fri11]. Impossibility [Fix19]. Impossible [Tho19]. Improbable [Mur12, BJ10]. Improvement [Ell10, Ins19, Wig11]. Inc. [Mox19]. inclusive [Swe15]. Incorporated [Bar13b, Mia12]. Incredibilis [Sto12, FF11]. Index [Ano10n, Ano11n, Ano12n, Ano13n, Ano14p, Ano15j]. India [Bha11, Kum18, Pha10, Pos14, Sta14, Win19a, Abr09, Ban14a, Cha12, KB09, Kum16, LMS13, Pha13, Sin08]. Indian [Ham16, Pay16, Mah18]. Indiana [Pha10]. Indians [Few10]. individual [Hog12, Wil16b]. Indo [Bil18, KMB18]. Indo-Pacific [Bil18, KMB18]. industie [BFGG11]. Industrial [Bud11, L^{+}10, Noa14, Ree17, Vol16, Bul16, Eps14, SM12, Tom12a, Tom13a]. industrial-hazard [SM12]. industrie [Sim13]. Industry [Far10a, Hug10, Mil12, Ree16b, Ree17, Sim13, And09, BFGG11, Dew14, Fox09, Löw11, Ste14a, Tom12a, Tom12b, Tom13a, OP16]. Inequality [Tat17b]. Inevitability [Ayl17]. Infanticide [Mon10, Ree09]. Infectious [Han16]. Infinite [Tim12, LG11]. Infinitesimal [Ale14, Bar15a, Fix16]. Influence [Bud13a, Bra12, Gro17]. Influences [McG13, QM13, Van14]. Information [GS08, YM10, Hem15, Hil12, Far16]. Infrastructure [Goo18]. Ingrid [D’A18]. Inkster [YM10]. Inner [Söd10, Tur08b]. Innes [Lou16b]. Innocent [Wai18]. Innovation [Aga17a, Ber15, Boo17, Bud11, Cha10, BFGG11, Erl11, Erl13, HD13, L^{+}10, Llo09, OR08, Sim13]. Inquiry [HA19b, Rih12, Leh12b]. Inquisition [May10, Tar13]. INR [Ban14a]. insects [Cla09, Ter10]. Insights [Ric10, CRB09]. Inspiring [Dan19]. Institute [CC11a, Gli12, SV17, Abt11, Car12]. Institution [Goo11a]. Institutional [Jon16a]. Institutions [Erl11, Erl13, McM18]. Instrumental [Ebe17, Ver10]. Instruments [Dun12b, Dun17a, Hig15, Joh11, Sim16, AKM14, HW11, Mar09, Boo17]. Insurrection [Hen19]. integrated [Van11a]. integrity [KT19]. Intellectual [BZ11a, Ben13b, Far16, Goo12a, Huf11, Les14, Mac11a, Ren10, Rot16, Tur12b, BJW11, Des12, Goo11b, Hem15, Joh09, Pes12, Rin04]. intellectuals [Ban11]. Intelligence [Jam17, Kus15, Tur12b, Goo11b, Hun13, Lüt19, Pet10b]. Intentions [McC19]. inter [Nie10b]. interaction [Hil12]. Interactions [Kot16]. interdisciplinary [RSA^{+}17]. Interests [Pis17]. intermediaries [Hic19]. International [Aga14, Jon10b, Lyk16, Mic11, Ree17, Bro09, KCM13, Man15, Mou17, TW10, Wal12b]. internationalen [Fan10, Sta11a]. Internet [Tat17a]. Interpretation [Boa12, Sau13, Sta11c, Cam09, BB12, McE10]. intervention [Mer12, Mer13]. Interventions [Han16]. Interviews [Fin10b]. interwar [BW14, McL12]. introducing [Ben19b]. Introduction [Chr10, HL12, Raj10, Ree16c, Sta16b, THB12, BW13, Jam10b, OGE17, VM12, Kip16, Mor13, Ree18]. Intrusive [Han16]. Inventing [Eps14, Pow12, Vol16, Ric16, Roo12b]. Invention [Bar17, Ens12, Gil11, Pea19, Fra16, Lyn10, Pri10]. Inventions [Ebe17, Ill12, Joh14]. inventory [VW10]. Investigation [Bla17]. Invisible [AK16, Gou16, Jar10, Pri15, Rat09]. Ionosphere [Hug10, And09]. Iqbal [Bur13]. Iraq [Jon10a, Rob08]. Ireland [Ade09, Cul11, Kir15, Mol14]. Irish [O’L12, Cul13]. Irmgard [Ser14b]. Iron [Bar16, Hol14b]. Isaac [Art11, Cle15, Gui09, vB13]. ISBN [AN18, Ada10, Ade10, Adl16, Aga13a, Aga14, Aga15, Aga17a, Aga17b, Ako19, Alb11a, Alb15, Alb16, Ale11, And12a, And12d, And12c, And12b, Arb17, Arn16, Art11, Ase17, Ayl17, Ayl18, Ayl19, Bad16, Bad19, Bak11, Bak14, Bak15b, Bak15a, Bal14b, Ban14a, Ban15, Ban12, Bar13a, Bar13b, Bar15a, Bar15b, Bar16, Bar17, Bar15c, Bec12, Bec16, Bec18a, Bec18b, Bec18c, BT14, Bel12, Bel17, Bel18b, BZ11a, Ben19a, Ber11a, Ber12, Ber15, Ber17, Ber18, BS16, Bha11, Big10, Big18, Bil15a, Bil17, Biv10, Bix19, Bjö11, Bjö14, Bla17, Boa10, Boa12, Boa15, Boa18, Bod18, Bod19, Boo11, Boo12, Boo14, Boo17, Bou15, Bow16, Bow18, Bow10, Bow17, Bro11a, Bro16, Buc11, Bud11, Bud13a, Bui19, Buk14, Bur11, Bur13, Bus18, Cal12, Cal13, Can15, Can16, Can19b, Can19a]. ISBN [Cap14, Cap15, Car16a, Car18, Car16b, Car11, Car12, Cha14, Cha10, Chi12, Chr10, Cle14, Cle15, Cle16, Cle17, Cle18, Coe10, Col13, Con15, Coo11, Coo14, Cow10, Coz15, Coz16, Coz17, Cul11, Cul13, Cun10b, Cuo12, Cuo13, Cur18, D’A16, D’A18, Dan19, Dar11, Das18, Daw17, Daw10, Daw12, Daw14, Daw15a, Daw15b, Daw15c, Daw16a, Daw16b, Daw16c, Die15, Dih16, Dih17, Dij12, Dob14, Doh14, Doh16, Dra17, Dry10, Dub10, Duf18, Dun11, Dun12a, Dun12b, Dun14, Dun15a, Dun15b, Dun16a, Dun16b, Dun17a, Dun17b, Dys14, ES16, Ebe17, Ebe18, Edw14, Egm12, Egm13, Eic12a, Eic12b, Eic15, Elt12, Ens12, Esp10, Eva13, Eve13, Eve18, Fal18, Fal13a, Fan12, Far10a, Far10b, Far11, Far12, Far13, Far15, Far16, Far19, Feo18, Fer11a, Fin10a, Fin10c, Fin11]. ISBN [Fin17, Fix16, Fix19, Fla19, Fon12, For10, Fot18, Fra17, Fre10, Ful18, Ful10a, GP18, Gar11, Gar14, Gas10, Gib14, Gif16, Gil15a, Gil16a, Gil12, Gil16b, Gil11, Gis19, Gli11, Gli12, Goo10a, Goo11a, Goo12a, Goo18, Goo17, Goo10b, Goo12b, Gou12a, Gou13a, Gou13b, Gou16, Gou17, Gou12b, Gou12c, Gou14, Gra11, GG12, Gre14b, Gri15, Gul18, Hah16, Hai16, Hal10, Hal14b, Hal17, Ham14a, Ham10, Ham14b, Ham15, Ham16, Han17, Han16, Har16, Har10a, Hay18, Hay19b, Hed13, Hei15, Hem19, Hen14, Hen15a, Hen15b, Hen10, Hen13a, Hen17, Hen11a, Hen11b, Hen13b, Hen19, Hig10, Hig11, Hig13, Hig14a, Hig15, Hig16, HA16, Hoc14, Hoc16a, Hol13a, Hol14a, Hol15, Hol16a, Hol16b, Hol10, Hol16c, Hom16, HA19b, Hoq10, Huf16, Hug10, Hul10]. ISBN [Hun18, Hun19, Hus15, Ins19, Jac15, Jam10a, Jam11a, Jam11b, Jam12, Jam13, Jam17, Jam18, Jar10, Jar16, Jar17, Jim19, Joh15, Joh13a, Joh11, Joh13b, Joh14, Joh16, Jon10a, Jon10b, Kaa18, Kao12, KK14, Kee12, Kee17, Keh12, Kel17, Ken13, Ken19, Kid11a, Kid11b, Kip16, Kir13, Kir15, Kne10, Kni15, Kni16, Kno14, Köb16, Kot16, Ksi13, Ksi15, Ksi16, Kum16, Kum18, Kus15, Kus18, Lan11a, Lan12, Lan15a, Lan19, Lan14, Lan15b, LN14, Lau15, Law17, Lea16, Lee15b, Leh12a, Les13b, Les13a, Les14, Les15, LD18, Lin17, Lo15, Loi16, Lou15, Lou16a, Lou16b, Lou17, Lou19, Lyk16, Mac11b, Mac13, Mac10, Mac11a, Mae12, Mai12a, Mai11, Mai12c, Man13a, Man13b, Man17, Man16, Mar16, Mar17a, Mar18a, Mar12a, Mar15, Mar12b, Mar12c]. ISBN [Mar14a, Mar17b, Mar18b, Mat13, May17, McC19, McO10, Mca16, Met11, Mey11, Mic11, Mil11a, Mil11b, Mil12, Mil13, Mil15, Mil16, Mil17, Mil11c, Mil11d, Mis11, Mod17, Mon10, Moo17, Moo13, Mor10a, Mor18, Mor15b, Mor13, Mos16b, Mos16a, Mox16a, Mul13, MY18, Mur12, Mye12, Nal12, Nal15, Nas16, Nav14a, Nav15a, Nav15b, Nav16, Nes10, Nic17a, Nic17b, Nic19, Nor12, O’C17, OP16, Oos12, Opi14a, Osl10, Par11, Par12, Pay16, Pea18, Pea12, Pea19, Pet10a, Pha10, Pic10a, Pic10b, Pin12, Pin13, Pis16, Pis17, Pos14, Pos15a, Pri12, Pri19a, Pri13c, Pri19b, Prk17, Prk18, Pro11, Pum12, Pyl10, Qui12, Rad13, Rad14, Rad17, Raj15, Rap11, Rap12, Ree14, Ree16a, Ree16b, Ree17, Ree18, Ree19a, Ree15, Ree10, Reg17, Rei15, Ren19, Ren10, Ren15, Ric12]. ISBN [Ric11b, Ric18, Ric10, Ric14, Ric16, Rie16, Rih12, Rob10a, Rob14, Rob15, Rob17, Rob10b, Rod15a, Rod15b, Roe17, Roo12a, Roo12b, Roo12c, Roo15, Ros17, Ros16, Røs15b, Røs15a, Rot16, Row10, Row14, Rud13, Rus11a, Sac16a, Sac16b, Sac19, Sal14, Sal15, San12, San13, San14, San15a, San15b, San16, San17, Sar13, Sar14, SJ19, Sau13, Sau15, SS11, Sch17a, Sch15a, Sch15b, Sch16, Sch17b, Sch19, Sea10, SS12, SS13a, SS15, SS16, SS17, Ser12, Ser14b, Ser14a, Ser15, Sha10, Sha11, She13, Shi16, Sho14, Sho18, Sia18, Sim14, Sim17a, Sim13, Sim16, Sim17b, Sin15, Siv10, Sle10, Sle11a, Sle12, Smi10b, Smi10c, Smi11d, Smi11c, Smi11b, Smi12, Smi14, Smi16b, Smi16a, Smi18a, Smi18b, Smi19, Smo10, Sno10, Söd10, Som11, Som15, Sør17]. ISBN [Sta11a, Sta11c, Sta11d, Sta12b, Sta14, Sta15, Sta16a, Sta19c, Ste10, Ste11b, Ste14b, Ste16, SM19, Sto14, Sto12, Str11, Stu10, Stu11, Stu17, Sum11, SJ11, Tam14a, Tam16b, Tam16a, Tar12, Tar15, Tat17a, Tat17b, Tat16, Tat18, Tau11, Tay14, Tee11, Tes11, Tho16a, Tho19, TS14, Tho16b, Til10, Tim12, Tos11, Tot10, Tra16, Tun19, Tur16a, Tur10, Tur14, Tur15, Tur16b, Tur12b, Tys17, Und12, Und13, Val14, Val11, Van11b, Van11c, Van14, Ver15a, Ver16, Vol16, Vos11, Wad11, Wad19, Wag14, Wag15, Wai18, Wal16a, Wal18, Wal19a, Wal11a, Wal12c, Wal16b, Wal17, Wal15, Wal11b, War13, War19, Web10, Web11, Web13, Web14, Web15, Web17, Web18a, Web18b, Wes15, Wes12b, Whi15, Whi16, Wig11, Wil11a, Wil11b, Wil16a, Wil12]. ISBN [Wil14, Wil16b, Win15, Win13, Wit11a, Wit11b, Wol11, Wol14, WM11, WM18a, Wri16, Wyl11, Yal17, YM10, Zai16, dC14, vN19, vL13a, vL13b, vL15, vL16]. ISBN-13 [Col13, Cuo13]. Islam [Bur13, Bil15b, Dal10, Gue11, Moo13, SS11]. Islamic [Bur13]. Island [Mar12c, Von11, Par18]. islands [KMB18, Par18]. Ismail [Bil15b]. issue [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14c, Ano14a, Ano14d, Ano14b, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano18b, Ano18a, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e]. Italian [McG13, QM13, Van14, BCP12, CC11a]. Italy [Coz17, Feo18, Cap14, Cle14, Coc13, CC11a, Duf18, Ebe17, Mar11, MR12, Rap12, Ray15, Sac16b, Tar13]. Items [Gai17]. Ithaca [Dun12a, Rot16, Row10]. itinerant [JA15]. Ito [Hoc16a]. Itty [Pha10]. Iurii [Mai12c]. iv [Cal13, Mil13, Mil17]. Ivory [Sha12]. Iwan [Hun18]. ix [Ada10, Ako19, Arn16, Bad19, Bar13b, Ber15, Bod19, Cle15, Coo14, Cow10, D’A18, Dan19, Daw14, Die15, Ens12, Far13, Fix19, Gil11, Goo12b, Hal14b, Hen15a, Hen19, HA19b, Ins19, Joh15, Joh13b, Ksi16, Lan19, Lo15, Mey11, Mon10, Mor10a, Pha10, Pic10a, Pic10b, Ren10, Roo12c, San12, Sar13, SJ19, Sch15a, Söd10, Sta12b, Ste11b, Stu11, Tar12, Til10, Tun19, Wal15, War19, vL16]. ixv [Ksi15]. Izquierdo [Kus18].

J [Ade10, Alb16, And12a, Boo11, Boo17, Bow17, Cha10, Con15, Dun11, Dun17a, Gli12, Gou14, Gra11, Hed13, Hol13a, Hol15, Hoq10, Jar10, Ksi16, Lan15b, Lou16b, Mac10, Mil11a, Mis11, Mor13, Nal15, Oos12, Rad17, Ren15, Rie16, Rob15, Š^{+}11, Sim17a, Smi14, Smi18a, Sta11b, Sta15, Tar15, Tim12, Und13, Wad11, Web10, Web18b]. J. [Das18, HA19a, Jon16a, Mil11c, Mon12, Pea18, RC11, Sle10, Val11]. J.B [Roe17]. J.L. [Bad16]. Jackson [Biv10, Mil13]. Jacob [Joh15, Des12]. Jacomien [Gou16]. Jacqueline [Pet10a]. Jacques [Sta11a, Coo12, Fan10, dC14]. Jacyna [Hul10]. Jagdish [Pha10]. Jahnavi [Ban14a]. Jain [Kum18]. James [Bil17, Car12, Daw16b, Gou13a, Ins19, Loi16, McO10, Mil11a, Mor13, Row12, Ste10, Win13, YM10, vN19, Abt11, Bak15b, Bor08, Ebe18, Far15, Fis10, FMW14, Hen15b, Hun10, Kni15, Lam11, Les15, Mar12a, Mil08, Ree16a, Ric16, Ros17, Rus14b, Sta12a, Tay14]. Jamie [Sch15b, Whi16]. Jan [Cle17, Hoc14, Boa18, Car11]. Jane [vN19]. Janello [Ben19a]. Janet [CC11b, Lo15, Rad13, LD18, Smi16a]. Janina [Web11]. Janssen [Hus15]. Japan [Lee15a]. Japanese [Smi18a]. Jardine [Car16a, Chr10]. Jared [Smi19]. Jasper [Ako19]. Jaszi [Goo12a]. Jaume [Sau15, Can19a]. Jay [Fin10c]. Jean [Chr10, Esp10, Jam13, Coo12, Sim14, dC14]. Jean-Jacques [Coo12, dC14]. Jean-Paul [Sim14]. Jed [Ksi13, Mur12, Sch15a]. Jeff [Dun12a, Can19b, Wil16a]. Jeffrey [Goo17, Ser15, Bak11, Car12, Sau13]. Jennifer [Dry10]. Jens [Wal11a]. Jeremiah [Hig13, Hig13]. Jeremy [Hen13b]. Jeroen [Wri16]. Jessica [Hen17]. Jet [Gif16, Kel17]. Jim [Rad13, Dun14]. Jirí [Gli12, D’A16]. Joachim [Ren10]. Joan [Gou12c, Web17]. Joanna [Mil15, Dun12a, Mar17b]. Jockeys [Web15, Ras14]. Joe [Ric18, Mor10a, Ric10]. Joel [Smi11c]. Johann [Mül11]. Johannes [Reg17]. John [Aga15, And12d, Dys14, Edd08, Gil16b, Goo10a, Lou16b, Rad13, Rob10a, Sch10, Sea10, Smi11c, Tot10, Aga14, And18, Ans11, Boa12, Bow16, Daw10, DeV09, Fer11a, Gro17, KWB15, Lan12, Lee15b, Mod17, Mur14, OPR11, Roo12b, Roo12c, Rot16, Ser12, Spi15, Sto14, Sto12, Vic11, Win13, Sim17b]. Johns [Aga17b, Arn16, Ban15, Bow16, Bro16, Can19b, Cle15, Coz16, Dry10, Eic12a, Fra17, Gib14, Gif16, Ham10, Hol13a, Joh13b, Joh14, Kir13, Leh12a, Mac11a, Mar18a, Mil11a, Moo17, Mos16b, Pea19, Ser15, Sta11d, Sta15, Str11, Tur15, War13, Web15, Wil14]. Johnson [Bad19, Gar14, Ree16a, All12]. Johnston [Bak14, Wal11b]. Jolivette [Røs15b]. Joly [Hed13]. Jon [Wit11b]. Jonathan [Goo17, Qui12, Und13, Aga17b, Hoq10, Roe17, SS17, Wal17, Wes15, Zai16]. Jones [Ada10, Bar15c, Fin10b, Jam13, Lou15, Lou19, Mor15b, Sta11d]. Jonsson [Gil16b]. Jordan [Stu10]. Jörg [Ayl18, Rad14, Cap14, Mey11]. Joris [Tho16a]. José [Fal13a]. Josefowicz [Mur12]. Joseph [Ful18, Jam13, Kao12, Und12, AJ12, Ban15, Bro12, Gai17, Hen19, Ric18, Ric14]. Josephson [Arn16, Mai11]. Journal [Boa10, Erl11, Erl13, Hig11, Bal15, Wat14, PdCB10, Wal16a, Vol08]. Journalism [Tun19]. Joyce [Gul18]. József [Joh14]. Jubilees [Ver16, Dho15]. Judy [Rei15]. Jugend [Hen11b, Rec10]. Juler [Jar16]. Julia [Gar14]. Julian [HA19a, Kem19]. Juliana [Cul11]. Julie [Pea19, Goo10a, Jon10b]. Julie-Marie [Pea19]. Julius [Car18]. July [Fie10]. Jurassic [Wyl11, Bri10]. Jürgen [Tam16a]. Jurisprudence [Hen10, DS08]. Justin [Cle16, Eic12b].

K. [Nal12]. Kaartinen [Raj15]. Kaat [Tho16a]. Kai [Ste11b]. Kaiser [Aga17a, SJ11, HSW09]. Kaiserfeld [Røs15a]. Karel [Bui19]. Karen [Hol15, Sta19c, Alb15]. Kargon [Bud11]. Karin [Car18, GG12]. Karine [Cuo13]. Karns [Dry10]. Karsten [Pri19b]. Kassler [Sch15b]. Katherine [Vol16]. Kathie [Kus18]. Kathleen [Stu10, Hei15, Met11]. Kathryn [Ree16b]. Katja [Prk17]. Katrina [Die15]. Katz [Hol15]. Keeper [Bro12, Und12]. keeping [Olo14, Bal14a]. Keevak [Coo14]. Keighren [Lou16b]. Keller [Rus11a]. Kelly [Web17]. Kendall [Ban12]. Kennaway [Gou13a]. Kenneth [Lau15, Ree14]. Kennett [Bil17]. Kepler [Chr10, Reg17, Sch19]. Kernels [Hol16c, Han15]. Kerridge [VM18]. Kersten [Hen15a]. Kevin [Smi11d]. Kew [Mac15]. Keynes [Rad13]. Kichigina [Nes10]. kidney [C^{+}12]. killing [Yan13]. King [Sho14, Wag15]. Kingdom [Hol16c, Han15]. Kirby [Gou12b]. Kircher [Gre14b, Sto12, FF11, Sto13]. Kirk [Cha14, Wil11b]. Kiss [Nic17b]. kitchens [Heg13]. Kjaergaard [Hal10, Bro16]. Klaas [Cle15]. Klaus [vL13a, Nas16]. Klaver [Mil11c]. Klein [Rei15]. Kleist [Cam12, TS14]. Klemm [vL15]. Knew [Ren19]. Knight [Eve13, For10]. Knoeff [Wal16b]. Knossos [Ger09, Smi10c]. Know [Rih12, Leh12b]. Knowing [Bar13b, Bec15, Nal12, Lan11b, Mad16, Mia12, Sch16]. Knowledge [Aga17a, Bak11, Bak15a, BT14, Bow18, Bur13, Cha09, Cle17, Daw12, Ebe18, Far13, Far19, Gas10, Goo17, Hoc14, Hui09, Joh15, Mat13, Mod17, Mox19, Pos14, RSA^{+}17, Van11a, Sch17b, Smi16a, Sum11, Tar12, War19, AS12, Asp11, BH19, Ber13, BW13, Bor08, BS15, Few10, Fyf12, Hic19, Hil12, HSSR13, Jac14, KT19, Kri12, LR11, LMS13, Lyn12, Mai12b, Mor17, Nav14b, Pet18, San10b, Sim19a, SC12, Uek10, Ver15b, Wig10, dC09, Dub10, Nav16, Ste10, Tot10, Wri16, Yal17, GWW09]. Knowledges [Alb11a, Nay10]. Knox [Stu11, Bat10]. Koby [Nav15b]. Koch [Str11, GF09]. Kohlstedt [Pro11]. Kohn [Smo10]. Kongelige [Hun19]. Korea [Lee15b, DiM13]. Körper [Kle13, vL15]. Kr [Gil16a]. Kragh [Hal10, Hun19, Sia18]. Kremer [Hig15]. Krieg [BTW09, Wal11a]. Krige [Aga14, Aga15, Mod17]. Kristian [Hal10, Gou14, Sta11c]. Kristin [Smi11d, Tur10]. Kristine [Rob17]. Kuhn [Rad17]. Kulturgeschichte [Web11]. Kumar [Bha11]. Kuni [Car18]. Kurt [Hig11, Sha10]. Kuru [Ham10, And08]. Kusukawa [Egm13, Gar11]. Kutzinski [Mar14a, Mar12c]. Kyle [Feo18].

L [Ade10, Bec12, Buk14, Can19b, Cha10, Hen11a, Hig15, Hol14a, Jar17, Kee17, Lou15, Mor13, Rei15, Val11]. L. [Hul10]. Laak [Ase17]. Lab [Bix19, Kir11, Gou12b]. laboratoire [BFGG11, Sim13]. Laboratory [Gif16, GF09, Nes10, Sim13, Sta14, BFGG11, Cha12, Kic09, Van11a, Szu18, Lun13, Str11]. Lachapelle [Tra16]. Lack [Coe10, Hen14, And13a]. ladies [End14]. Lady [Mae12, Mes10]. Laertes [Ric11b]. LaFollette [Boo14]. Lake [Gil16b]. Lalli [Nic19]. Lamarckism [Tur16a, Gil15b]. Lamb [SS17]. Lambert [Jam11a]. Lamont [Som15]. Lancelot [HA19a]. Landwirtschaft [Uek10, Web13]. Lane [McO10, Nal12, Ros17]. Lang [Bar13a, Jim19, Ken19, Pri19b, Sim13, WM11]. Language [Hol14a, Rob10a, Edd08, Mon13, Rec10]. Lansley [Jim19]. Lanzoni [Bod19]. Largent [Sch17a]. Larrington [Lou16a]. Larry [Ful18, Pos14]. Larsen [SS16]. Late [Fal13a, Feo18, Ken13, Rap12, Sea10, Bea11, DeV09, Hic19, Hog12, KT19, Mar11, CG12]. Later [Bec18a, Bee19]. latest [Coh18]. Latin [Pod13, Tun19]. Laura [Mar14a, Coe10, Fin11, Web18b]. Laureate [Mai11]. Laurence [Fer11a, Kee12]. Law [Hen10, Les15, DS08, WC13]. Lawrence [Kot16, Wil11a, Mac11b, San14]. Laws [Hen10, DS08, DP17]. Leah [Sea10]. learn [Mer19]. Learned [Gib14, Yal15, Rab17, Sir13, Tam16b]. Learning [AN18, HW11, Pis17, Llo09, Dun12b]. Leben [Hen11b, Rec10]. LeContra [Chr10]. Lecture [Les13a, Bar08, SDS13]. Leech [KP13, Cha14]. Leeuwen [Gul18]. Legacies [May14, Ren15, FT15, Daw16c]. Legacy [Boo12, Mos16a, Gat17, Nie10a, RH12, Val10, And12d]. Legal [Goo12a, BJW11]. legend [Ins19]. Legenda [Tur16a]. Legitimizing [Røs15a]. Lehner [Hus15]. Lehoux [Rih12, Rob10b]. Leibniz [Eic12b, Fix19, Gar09a, Gar09b, Mor15a, Pyl10, Smi11a]. Leiden [AN18, Bak14, Bak15a, Ben19a, Car16a, Car18, Dob14, Dub10, Dun14, Dun17a, Fal13a, Fon12, Gou16, Hig13, Hig15, Loi16, Mos16a, Pha10, Pos14, Roo12a, Sha11, Sho14, Sma13, Smi16b, Sto12, Van11c, Ver16, Wri16]. Leigh [Cap13, Rob15]. Lemov [Rei15]. Léna [Ayl17]. Lenin [Mai11]. Leo [Col13, Tar15]. Leon [Wal15]. Leonelli [Ste11b]. Leong [Tar12]. Leonhard [Bec18c]. Leopold [Mar17a]. Leroux [Car11, Gol10]. Leslie [Mil12]. lessons [OG16]. Letters [Duf18, Gli12, Smi11c, Sch10, Š^{+}11]. Lettres [Chr10]. Levere [Ful18]. Levin [Bud11]. Levine [Bjö11, Fal18]. Levitt [Vos11]. Lewis [Edw14]. lexicon [Zip13, Ree15]. Lexicons [Win15, FSS^{+}14]. Ley [Smi19]. liaisons [Ada16]. Liber [Mim15]. Liberal [Mat13, SC12]. liberalism [Edd17]. Librairie [Hed13]. Library [Doh14, Dun15a]. libri [Cle18]. Liddel [Car18]. Lie [Joh13b, Bun12]. lies [McG19]. Life [Ban15, Bec18a, Bud13b, Car12, Cle16, Cre13, Gig09, Gil12, Gil16b, Ham14b, Hem19, Hen14, Hen15b, Hir11, Ins19, Jam12, Jam17, Jar17, Mae12, Mai11, Mar19, May14, Mox16a, Mox16b, Mye12, Prk18, RC14, Rad13, Rec10, Ren15, Roo12a, Ros17, Rud13, Sim17b, Smi11c, Ste13, Sto12, Und12, Und13, Web10, Web15, Abt11, AD12, And13a, Ayr12, Bal19, Bas14, Ben19b, Ben11, Bro12, Del10, FF11, FMW14, Mes10, Mon12, Nor10, Nov12, OGE17, Ras14, Ric08, Rus11b, Sch10, Smi11a, Swe15, Wel17, Wil11c, Alb15, Cap15, Eic12b, Hol10, Web15, Gai17]. Light [Eve18, Mil11a, MY18, Cou17, Fis10, Hay19a, Hun10, Pum11]. lighthouses [Cou17, ML10b, Elt12]. Lightman [Bow10, Daw15a, Daw15b, Daw16a, Jac15, Pos14, Ree19a, Smi16b]. Lightning [Gra11, HHR09]. Lights [Elt12, ML10b]. lii [Fin10a]. Like [CGC13b, Smi16a, Ver15b, CGC13a]. Lily [Opi14a, Hol13b]. Limitless [Rob14, McC12]. Limits [Sac16a, Law15b, Spi11, Spi12b, Tar13]. Linear [Row10, Edg09]. lines [Lam11, OG17]. Linnaean [Bil17, AK16].

Linnaeans [Hod11]. Linnean [Bil17].

Linstrum [Tho16b]. Lisa [Stu11, Cur18, Lan19]. Lisbon [Val14, Coe13]. Lise [Ver15a]. Lissa [Egm12]. Lister [Fon12, Roo11]. lists [Csi17]. LIT [Egm12]. Literary [Far19, Tun19]. Literature [Cal12, Jam10a, Lou16a, Sal15, Sle11a, Sle11b, Tam14a, TS14, Cam12, MHO13, Ort09, Per12, Shu10, Cal12, Dra17]. Live [Bec18b]. Liverpool [Ban12, Bar15c, Zal14]. Liversidge [Bro11a, Mac09]. Lives [Hig16, Loi16, BW14, Csi17, Fis14, Nie14]. Living [Hen17]. Livingstone [And12a, Mac11b, Dri10]. LLC [Fer11a]. Lloyd [Eva13, Ful10a]. Locality [Vog10, Raj10]. Localization [Prk17]. localizing [Olo14]. Locating [AF16, Mur10]. Lock [Tim12]. Locke [Lan12, Ans11]. Lockhart [Ebe18]. Loeb [Fan10, Sta11a]. Logic [Ens12, vL13b, App12, Pri10, Tur38, Val10]. Lomonosov [Mai12c]. London [Ade10, Alb11a, Alb15, And12a, And12c, Art11, Ase17, Ayl18, Bak15b, Bal14b, Bar13b, Bar15a, Bec12, Bec18b, Bel12, Bel18b, Ber12, Ber15, Ber18, Bil17, Boa12, Boa15, Bod19, Boo11, Boo12, Boo14, Bou15, Bow18, Buc11, Bud11, Buk14, Cal13, Can16, Car12, Cha14, Cha10, Chi12, Cle14, Cle17, Coe10, Con15, Coo11, Cow10, Coz15, Cul11, Cun10b, Dar11, Daw12, Daw15c, Doh14, Dra17, Duf18, Dun12a, Dun15a, ES16, Ebe18, Egm13, Ens12, Eve18, Fan12, Far10a, Far11, Far13, Far15, Far16, Fin10a, Fin10b, Fin10c, Fla19, Fot18, GP18, Gas10, Gil12, Gil11, Goo11a, Goo12a, Goo10b, Goo12b, Gou12b, Gou12c, Gre14b, Hal14b, Ham15, Han17, Har10a, Hay18, Hig10, Hol14a, Hol10, Hom16, Hul10, Ins19, Jac15, Jam11a, Jam18, Jon10b, Kaa18, Ken13, Kid11a]. London [Kid11b, Kne10, Ksi16, Kus15, Kus18, Lan14, Lan15b, LN14, Lou15, Lou16b, Mac11a, Mae12, Mai12a, Mai11, Mar12a, Mar12c, Mar17b, McO10, Mil12, Mil11c, Moo13, Mor18, Nal12, Nav14a, Nic17b, Oos12, Par11, Pic10b, Pin12, Pin13, Pri12, Pri13c, Prk18, Pro11, Pum12, Rad14, Raj15, Rap12, Ree15, Ree10, Rei15, Ren10, Ren15, Ric12, Ric18, Ric14, Rih12, Roo12b, Roo12c, Roo15, Ros17, Row10, Rus11a, Sac16a, Sac19, Sal15, San12, San14, San15b, Sar13, SS11, Sch16, Sch19, SS12, SS13a, Ser14a, She13, Sle10, Sle12, Smi10b, Smi10c, Smi18a, Smo10, Sno10, Söd10, Sta19c, Ste10, Ste14b, Stu10, Sum11, Tes11, Tho16a, Tur16a, Tur10, Tur14, Und13, Val14, Van14, Vol16, Wad19, Wal12c, Wal15, Wal11b, Web10, Web14]. London [Web17, Wes12b, Whi15, Wig11, Wil16a, Wil12, Wit11a, Wol11, WM18a, Wyl11, YM10, Zai16, vN19, Bee19, BH19, CMC08, Cou17, Hig19, Hoc16a, Hun13, Ito14, KT19, Kus15, Rad14, Sim19a, Win19a, Yam15, dC09]. Long [Aga14, Bjö14, Fin10a, Hen17, War19, Zai16, Man13c]. longitude [Ker14]. looking [Law15b]. Lord [Bil17, Nas10]. Loren [HA16]. Lorentz [Gai17]. Lorenz [Hun19, Gam17, Gam18]. Lorimer [SS15]. Lorraine [Rad17, Rei15, Hen10]. losing [BG11]. Loss [Keh12, Sta19c, Ste11a]. Lost [Ham10, Hol16a, Joh13a, Rei15, Tat17b, Tat16, And08, Dud14, EKD^{+}13, Ren12, Dan14]. Lothar [Pin13]. Loudons [OP16, Dew14]. Louis [Boa10, Mil11b, Sho18, Boo12, Dew09]. Louise [Hig14a]. Loveland [Dun12a]. Low [Bak14, Bud11, Dun15b, Dun17a, Elt12]. LP [Sch13, Tur15]. Luana [Big18]. Lucia [Col13]. Ludvig [Hun19]. Ludwig [Tat18, vL13a]. lunar [Ben19b]. Lund [Røs15a]. Lung [Lan15a, Heg13, Tim13]. Lutterworth [Gar14]. Lvov [Hol16a, Dud14]. lxi [Eve13]. lxxiv [Gar11]. lxxvii [AN18]. Lykknes [Buk14]. Lynall [Far13]. Lynch [Lo15, Stu10]. Lynda [Ken19, Nav15a]. Lynn [Gil11, Hah16]. Lynne [Hom16]. Lysenko [HA16].

M [Ade10, Bou15, Dys14, Far11, Fin10a, Gif16, Gou13b, Lan19, Lou16b, Lou19, Mac10, Mar14a, Met11, Mil11c, RVG12, Rot16, RC11, San14, Sch17a, Sle10, VM18, Wad11, Wal15, Wil11a]. M. [Tam16b]. MA [AN18, Aga15, Alb16, And12c, Art11, Bad19, Bel17, Boo17, Bud11, Cha10, Dar11, Dih17, Dob14, Fin10a, For10, Gou12b, Hal14b, Har10a, HA16, Jam18, Joh16, Jon10b, Lan14, LN14, Lea16, Les15, Lo15, Mai11, Mil12, Mod17, Nav14a, Nic17a, Nic17b, Par11, Pic10b, Pin13, Pri12, Prk18, Ren10, Ren15, Ric16, Sac16b, San17, Sar13, Sar14, SJ19, Söd10, Tat17a, Tat17b, Tho16b, Tur10, Tur16b, Vol16, Wad19, Web10, Wes12b, Wol11]. Mabberley [Ric18]. MacDonald [Sim16]. MacGregor [Doh14]. Machina [Ver10]. Machine [Joh13b, Stu10, Bun12, LCMJ08, OGE17, Aga17b]. Machines [Eic12b, Ens12, Mor15a, Pos15b, Pri10, Smi11a]. Mackensen [Pri19b]. Mackinder [Hay19a]. Maclay [Kem19]. MacLeod [Bro11a]. MacMillan [Ham16, Ada10, Aga13a, Aga14, Ase17, Ber17, Ber18, Bow18, Cal12, Car16b, Daw17, Doh16, Duf18, Dun11, Far13, Fin11, GP18, Goo10a, Ham14a, Har16, Hoc14, Hoc16a, Joh13a, Kaa18, Kee17, Kno14, Köb16, Ksi15, Lan15a, Lin17, Mat13, Mil15, Mor10a, Mye12, Nor12, Pay16, Qui12, Ree10, Rob17, Rod15a, Rod15b, Ros16, San13, Sim14, Smi11b, Sta19c, Tra16, Wes15, Wol14]. Macsotay [Bui19]. Mad [Stu11, Bat10]. Made [Hen19, Jar16]. madnes [Wee19]. maestro [Ekh10]. Magic [Cun10b, Tra16, Hoc16b, Web08, Dun16a]. Magical [McW11, Fan12]. Magnetic [Car09, Ene14, Man13c, Rat16, San11, Fre10]. magnetometer [Ene14]. Maharaj [Aga14]. Mahon [Goo11a]. Mahoney [Pri12]. Maienschein [vN19]. Maifreda [BT14]. Main [Pri19b]. Make [Opi14a, Hol13b, Mur14, Pet18]. Makers [WM11, Poo10]. Makes [Hen17]. Making [Bal15, BJW11, Boa18, BS15, Bur13, Coo11, CC11a, Daw15c, Far15, For10, Gas10, Hem15, Hol16b, Ins19, Kee17, Mac15, Mar18b, Nav14a, Ric12, SS13b, War19, Bel18a, Ber13, Egm10, Hes11, Kni09, Llo09, Lun13, Mye17, RC13, Rus14b, Sch12, Wyn11, dC09, vB13, Cle15, Far16, Ful10a, Goo12a, Kel17, Nav16, Wal16a, Wal12c]. Malden [For10, Ren10]. Male [Har10a, Mic08]. Mallemant [Rab17]. Malpighi [Eic12a, Ber11b]. Malthus [Con15, Hal14a, May14, Ren15]. Man [Boa18, Goo10b, Hen15a, San12, Stu11, Del09, Hal14c, Ros10, Sch08b, Jen15, SF11]. Managing [Sta19c]. Mancall [Par12]. Manchester [And12b, Bui19, Hal17, Jar16, Kir15, Law17, Les13a, Les14, Mac10, Mis11, MY18, SS15, Sta12b, Wal18, Alb09]. Manfred [Sho14]. Mankind [Gou12a, Sem12]. Manners [Mor10a, Bor09]. Manning [Dob14, Fin17]. Mansfield [Bjö14, Fin10b]. Mansoor [Lyk16]. Mantra [Dry10, Ale08]. Manufacturers [Goo10a, ANPR07]. Manuscript [Edw14, Gat17, Lew12, LMSR09, oRLM^{+}09, Fin10a]. Manuscripts [Eag10, Sha11]. Manzolini [Mae12, Mes10]. Map [CGC13a, CGC13b]. Mapping [Ase17]. maps [Par18]. Mara [Hig15]. Marc [Jar10]. Marcacci [Pri19a]. Marcel [Boo14]. Marcello [Eic12a, Ber11b]. Marco [Col13, Wil11a]. Margaret [Bud13a, Sta11b, Adl16, Ber12, Joh15, Law15b]. Margherita [Duf18]. Maria [Nal12, Ase17, Gou12a, Tes11]. Marianne [Man17]. Marie [Far16, Fie10, Fon12, Pea19, Gil16a, Hem15, Tat17b]. Marie-Christine [Gil16a]. ‘Mariner [CC11b]. Mario [Goo12a]. Marion [Bar15c]. Maritime [Fin10a, LMSR09, oRLM^{+}09]. Marjo [Raj15]. Mark [Biv10, Eve18, Hen15b, Kot16, Mat13, Mic11, Mil13, Sch17a, SJ11, Whi16, Daw16c, Ens12, Har10a]. Market [Web18b]. Marklund [Goo18]. Marotta [CC11a]. Marr [Joh11, Rap12]. marriage [Ver19]. Marrinan [Ser12]. Mars [Nal12, Gif16, Lan11b, Smi16a, Ver15b]. Marsden [Sal15]. Marshall [Daw15b]. Marsilio [Gou16, Pri15]. Martha [Goo12a, May17]. Martin [Bak15a, Fon12, Hen19, Mos16b, Bil15a, Ksi16, MR13, Roo11, Roo17]. Martina [Bud11, Sia18]. Martyrs [Ale10, Par11]. Mary [War19, ES16, Val10, Van14]. Masculinity [Kaa18]. Masha’allah [Mim15]. Massimiliano [Sau15]. Mastermind [Goo11a, Mah09]. Masterminding [Der15, Adl16]. Matemáticas [Car16a, Nav14b]. Material [Par12, BM11, Mül11]. Materialist [Dra17]. materiality [RSA^{+}17]. Materials [Ins19]. Mathematical [Art11, Bar15a, Bar16, Bec18c, Cuo13, Fix16, Hol16a, Mor17, Pum12, Rap12, Sør17, Wal11b, Ale14, Bee19, Che12, GJ09a, GJ09b, Gou10, GG10, Gui09, Mar11]. Mathematician [Bud13a, Bra12]. Mathematicians [Ben19b, Bee19]. Mathematics [Ada10, Ale10, Bar13a, Bar15b, D’A16, D’A18, Hol14b, Joh11, Kot16, Rob08, Row14, Sia18, Sør17, Cla14, Dud14, Fox12b, Hac14, Hud14, Jon09, Mar09, Mor17, Nav14b, RS09, War10, War12, Ale11, Bar16, Hah16, Hol16a, Jon10a, Kot16, Par11, Pet10a]. Matteo [Rap11]. Matter [Cle15, Hir11, Man12, Pea12, Roo12a, Sac16a, AGF13, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14c, Ano14a, Ano14d, Ano14b, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano18b, Ano18a, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, San10a, vB13, Dob14, Hen19]. matters [Rud14, Ksi16]. Matthew [Kus15, Lan14, Mil16, Rob10a, Sim17a]. Matthias [Rob17, Wal11a]. Matus [Bec18a]. Maudsley [Kem19]. Maurice [Bud11]. Maurits [Mil11d]. Maurizio [Ber17, Ham15]. Maverick [Goo11a, Mah09]. Max [Rad09, Ren10, Rin04, Spi12a, Pin13]. Maximilian [Dun16b]. Maxwell [Hen15b, FMW14, Lam11, Sta12a]. may [Swe15]. Maya [Few10]. Mayhew [Ren15]. Mazzolini [Cap14]. McBurney [Kus18]. McCartney [Hen15b, Kot16]. McConnell [Dun15b]. McCormick [Keh12, Ste11a]. McCray [Aga17a, Rob14]. McCulloch [Prk18]. McEvoy [Boa12]. McGee [Fin10a]. McGhie [Wal18]. McGrath [Van14]. McGregor [Kus18]. McKinnell [Loi16]. McLaren [Mai12a]. McNally [Stu10]. McOuat [Pos14]. McRae [Lau15]. McVeigh [Smi18a]. McWilliams [Fan12]. Me [O’C17]. mean [Nie14]. meaning [Del09]. Measure [Fer11b, Far12]. measurement [BW13, McG19, MR13]. measuring [Ker14, Mad16]. Mechanical [Cle15, Gra15, Tat18, Bal19, Spi15, vB13]. Mechanics [Gul18, Sau13, Sta11c, Tat18, Cam09, BB12]. Mechanism [Ber11b, Lan11a, Eic12a]. media [Nie10b]. Medical [Bar17, Edd08, Egm13, Fot18, Hir11, Lou15, Pic10b, Ree15, Rob10a, Sta12b, Str11, Stu11, Wad11, Wal17, Wit11b, Alb11b, Bat10, Bro11b, Gab16, GF09, GCA10, Har08b, Hic19, Jon13, Jon16b, Kus11, Sta16b, Vog10, Zal14, Zip13, A^{+}12, McG19, Roo12a, Ser14b]. Medicalizing [Mar18b]. Medicine [Bec16, Cun10b, For10, Ful18, Gib14, Hal17, Ham16, Jac08, Jon10b, May17, Nes10, Pay16, Rad14, Siv10, Sle11a, Sta12b, Sta14, Ste16, Tar12, Wit11a, Yam15, Bro09, Bro11b, Cha12, Com12, Cre13, Har08c, Heg11, Kic09, Kni09, LR11, Mac16, Roy12, Sch11, Sel08, Shu10, Sir13, Web08, And12b, Bow18, Cap15, Daw17, Feo18, Hul10, Moo17, Ser14a, Smi10b, SM19, Stu17]. Medieval [Bur11, Dys14, Edw14, Feo18, Gou17, Ham16, Lou16a, RVG12, Sha11, Tim12, Fri11, Gla09, Eag10]. Medina [And12c]. Mediterranean [Dys14, RVG12]. Meeting [Mis13, Hig14a]. Megan [Les13a]. Mel [Ric18]. Meli [Bel18b, Eic12a]. Melinda [Wal16a]. Meloni [Ber17]. Meltzer [Rie16]. Melvyn [Jam17]. members [Sim19a]. Memoirs [Olo14]. Memories [Hol16b]. Memory [Win12, Yan13, Bal14b]. Men [Daw17, Daw14, Har10a, DeW13, KT19, Mic08, Pos15b]. Menagerie [Mca16]. mend [Bax13]. Mendel [Gli12, Š^{+}11]. Mendelian [Gli12, vTSŠ^{+}11]. Mendelianum [Gli12]. Mendelism [Gli11, Sim10]. Mental [Tom13b, Ell17, Wes15]. Meredith [Duf18, Sac16b]. Mesoamerican [May17]. Messange [Rab17]. Messbarger [Mae12]. metalliferous [AGF13]. metals [Yam15]. Metaphors [Str19]. Metaphysical [Daw15b, HMLE15]. meteorological [Wil15b]. meteorology [Har08a, Ric16, Tur10]. method [Fey10, Gui09, Art11, Kid11a, RC13]. Méthode [RC13, Coh11, Man13a, Ver15a]. Méthodique [Hei15, Doi13]. methods [Hay19a, McM18]. Metropolis [Smi14, Wes12a]. metropolitan [Mox19]. Mialet [Bar13b]. Micale [Har10a]. Michael [Bec12, Ber15, Bow17, Daw16c, Doh14, Fin10a, Gou14, Hen10, Hoq10, Jac15, Lo15, Rei15, Ric10, Rud13, Ser12, Whi15, Coo14, Dun17b, Gil11, Gou12c, Hay18, Jam10b, Jam12, LMSR09, oRLM^{+}09, Mor13, Mox16a, Pea19, Pri12, San15b, Sno10, Sta12b, Stu10, Tur14, Win11]. Michal [Gli12, Gli11, Gli12]. Michel [Bus18, Hus15]. Michigan [Ren19]. microscopes [Qui16]. Microscopy [Jar10, Rat09]. mid [Ban14b, Pet14]. mid-twentieth [Ban14b]. mid-twentieth-century [Pet14]. Middle [Bec18a, CG12, Fal13a, Osl10, Sea10, Bar08, DeV09, Kle13]. Midwifery [Hom16, Fal13b]. Migration [Ste16]. Miguel [San13, Car16a]. Mikael [McG15]. Mike [Ros16, Van11b, Fin10c]. Miklós [Gli11]. Milan [Cle18, Azz12, LN14]. Milena [Nav15b]. Militär [BTW09, Wal11a]. Military [Vol16, Eps14]. Miller [Ins19, Kus18, Loi16, Mar12a, Web18b, Loi16]. Mind [Can15, Dra17, Lou16a, Rei15, Sle11a, Smi10b, Smi18b, Til10, Whi16, Chi08, Chr13, CC14, EKD^{+}13, Gal12, Har08c, Pic14, Ric13b, Shu10, Sta19a, Tur08a, Eic15, KK14, Söd10]. Mind-Body [Smi10b, Har08c]. Minds [Tho16b]. Mineral [Kni16, Gib15]. Mineralogy [Rob10a, Edd08]. Miniati [Hig15]. Mining [Kel12, Sac16a]. Minteer [vN19]. Minuzzi [Cle18]. miraculous [Elm13, Hen13a]. Mirage [Rus11a, Kel10]. Miri [Yal17]. Miriam [Bud11]. Mirjam [Die15]. Mirror [Row10, Edg09]. Misbehaving [Arb17]. Miskell [Hig14a]. Mission [Cun10b, Web08]. Misticall [LG11, Tim12]. Mitchell [Wag14, Hoc14]. Mittelalter [Kle13, Pri19b, vL15]. mixed [Ekh10]. moa [And13b]. model [Bud13b]. Modern [Ade10, Aga13b, Ale10, Bal14b, Bar15a, Chi12, Cle14, Cle16, D’A18, Dob14, Ebe17, Egm12, Eve18, Fal13b, Fix16, For10, Ham16, Hen10, HA19b, Huf16, Kee17, Kum16, Mil15, Moo13, Mos16b, Mur12, Oos12, Par11, Par12, Pay16, Pea19, Pin12, Pos14, Ric14, Ric16, Roo15, Ros16, Row14, Sac16b, Sch16, Sch17b, Shi16, Sho14, Tra16, Tur10, Tur14, Tur12b, Wag15, Wes15, Wil11b, Wit11a, Ale14, AGF13, Bec15, BM11, Bod14, Bon11, BJ10, Coc13, Coh15, DS08, Fox12b, Goo11b, Gor12, Gue11, Har08a, HKZ12, Kin13, Kni09, Law15a, LMS13, Llo12, Man12, Mar14b, Mul10, Nav14b, O’L06, Par18, Raj10, Rav11, Ray15, Rob11, Sch11, Tom13b, VB11, Wer13, Win12, Yeo14, Hom16]. moderna [Nav14b, Car16a, Cle18]. Modernism [Hul10, Sta16a, Bra15, Ger09, Jac08, Boo17, Smi10c]. Modernist [Jar16]. Modernity [Bud11, Can19a, Ell10, Fin11, For10, Lan11a, Pis16, Wig11, Kni09, L^{+}10, SS10]. Mogens [Goo18]. Mohun [Kir13]. moitié [Boa10, Vol08]. Molecular [San12, San13, GS12, SF11]. Molecules [Man17]. Molinari [Mar12b]. Mollan [Kir15]. Møller [Hig11]. monad [Gar09a, Gar09b, Pyl10]. Moncada [Fin10b]. mondo [Pri19a]. money [Neu18, Van11a]. Monica [LN14]. Monk [Und13]. Monkeynut [Hen15a, Hal14c]. Monks [Eag10, Sha11]. Monochord [Tau11, Cre10]. Monograph [Keh12]. monster [Bin12]. Monsters [O’C17]. Montgomery [Hol14a, Sch17a]. Monumenta [Gat17]. Mooney [Han16]. Moore [McO10, Smi11b]. Moral [DeW13, Hen10, DS08, Gra15, Daw14]. Morales [Kus18]. Morales-Muñiz [Kus18]. Morality [Mil17]. Morals [Swe15]. Morandi [Mae12, Mes10]. Moravia [Gli11, Sim10]. Moravian [Gli12]. Morbid [Alb11b, Wad11]. Mordechai [Ksi13]. Morette [Ver19]. Morgan [Dra17]. Morocco [Ada16]. Morris [Dun11, Jim19]. Morrison [Bak14, Dun15b, Dun17a, Elt12]. Morrison-Low [Bak14, Dun15b, Dun17a, Elt12]. Mortal [Lou17]. Morus [Hun18]. Moses [Bus10, Fer11a]. Mosley [Car16a]. Mossensohn [Yal17]. Most [Ins19, Stu11, Mad16, Ros10, ZR15]. mother [Vic11, Reg17]. Motion [Bur13, Cle15, Pin13, Spi12a, vB13]. motu [Neu18]. mountain [Hoc16b]. Mountains [Kum18]. move [Hod11]. moves [JS17]. movie [Kir17]. Moving [Kum18, Ell17]. MRI [Man13c, Bjö14]. Mrs [Gar14, CC11b]. Mrs. [All12]. Much [Ren19]. Mueggler [Mul13]. Mulino [Cap14]. Mulino/Duncker [Cap14]. Mullaney [Chi12]. Müller [Ayl18, Gli11, Rad14, Rad14, San17, Ser14b]. Multinational [Ree14, BCH^{+}13]. mummies [Mos14]. Muñiz [Kus18]. Murder [Stu11, Bat10, Yan13]. Murders [Stu11, Ros10]. Muriel [Joh12]. Murray [Lou16b, KWB15]. Museum [Cuo12, Gli12, Han17, Hen11a, VW10, Bux18, Nie14, Ros18, Alb09, Dun11, Mac10, Mor10b, Nie14, Sun13, VW10]. Museums [Alb15, Elt12, Ree19a, Wad11, Wyl11, Alb11b, Bri10, Qui16, RC14, Alb16]. Music [Boo17, Ebe17, Ebe18, Gou13a, Gou16, Sem12, Ken12, Pri15, Bec18b, Gou12a]. Musik [Pri19b]. Muslim [Moo13, Gue11, Yal11]. Musschenbroek [DP17]. Mutio [Rap12, Mar11]. Mutualism [Con15, Hal14a]. Muzaffar [Bur13, Bur13]. Myth [Sac19]. Mythologies [Hol16b].

N [And12a, Pha10]. N.H [Cap15]. Nachrufen [Ech12, Web14]. Nachtomy [Cle16]. name [Mur14]. Names [Tim12, LG11]. Nancy [Gib14, Rud13]. Nanette [Car11, Gol10]. nanoscale [ML10a]. Nanotechnologies [Rob14, McC12]. Naomi [Aga15]. Napoleone [Kus18]. narrating [Hoc16b]. narrative [Ree16d]. Narratives [Zai16, Nie10b]. NASA [Aga14, KCM13]. Nasser [Sac19]. Natalie [Cuo12]. Nathaniel [Ser14a]. Nation [Chi12, Dun11, Lee15b, Sch17b, Tat17a, DiM13, Mor10b, Mul10]. Nation-Building [Lee15b, DiM13]. National [Elt12, Pis17, SJ11, Bal14a, Hig14b, Nie10b, Wal12b, HSW09, VW10]. Nationalization [Hoc14, AS12, Car16b]. Nationalsozialismus [Hen11a, VW10]. Natural [Alb15, Bur11, Car16a, Coo11, DS08, ES16, Far13, Fin09, Gil15a, Gou12c, Hen10, Hir11, Ins19, Kni16, Kus18, Lan12, Osl10, Pis16, RCK10, Roe17, Roo12a, Row14, Sal14, Sch17b, Sha10, Sim17b, Web14, Web18b, AG17, Ans11, Bal08, Bar08, Bel18a, BH19, BCP12, Egm10, Fox12b, Gam17, Gam18, Gib15, Gla09, Law15a, Law15b, Leh19, Lip14, Lyn12, Nav14b, PDG11, Phi12, Pri11, RC14, Sun13, Ter10, Ter14, Wal19b, Wil15a, Wot12, Alb11a, Gar11, Hen10, Mox19]. natural-history [Lip14, Wal19b]. Naturali [BCP12, Col13]. Naturalism [Bow10, Jac15, Lig09, Rec13]. Naturalist [Gou12c, Sim17b, PDG11]. Naturalness [Tys17]. Nature [Adl16, Alb09, Ban12, Bec12, ES16, Edw14, Egm13, Far11, Fon12, Hen10, KK14, Pro11, Rus11a, Sch16, Wal16a, Web14, Web18b, Wil16a, Bec15, Chr13, CC14, DS08, Der15, DP17, Har09, HNS11, Hod11, Kel10, Kel12, KW10, Koh10, Kus11, Lew12, Phi12, Roo11, Ros18, SC12, Ter14, Vic11, Fra17, Jim19, Mac10, Tar13, Whi16, Mat13, Bal14a, Bal15]. Naturopathic [Fra17]. Naturphilosophie [Gam17, Gam18]. Naturwissenschaftler [Ech12, Web14]. naval [Leg13, Pos15b]. Navarro [Can19a, Car16a, Sau15]. Navicula [Eag10, Sha11]. navigation [Lam09, Sum14b, Jam11a]. Navigational [Köb16]. Naylor [Alb11a]. Nazi [Can15, Fot18, Pic14]. NC [Hay19b, Hig10, vL16]. Neanderthal [Mad16]. Near [Rob10b, Hig19, Leh07]. Neary [Goo10a]. Necessity [Dih16, Dar14]. Need [Hol14a, Mon13]. Needham [Mou17]. needs [Rav11]. Neelam [Bha11]. Negotiating [Kir13, Moh13]. Neil [Cha14, Pea19]. Nelson [Boo17, Sim17b]. néo [Del09, Del09, Esp10]. Neo-Darwinism [Del09]. néo-darwinisme [Del09, Esp10]. Neoplatonic [Cla14]. Nervous [Bea11, Fin11, Har10a, Mic08, SS10, Ken13]. Netherlands [Egm12, Rob11]. Network [Tat17a, Pet14]. Networks [Car18, Hig19, ML^{+}12, Sim19a, Bak14].

Netz [Cuo12]. Neuro [Prk17]. Neuro-disciplines [Prk17]. Neurology [SS10, Fin11]. neurophysiology [SFFC12, Gou13b]. Neuroscience [Fin10b, Gro09, Wol11]. Neuzeit [Pri19b]. Never [Mag12]. Newcastle [Gas10, Kee12, Mar12b]. Newcastle-upon-Tyne [Gas10]. newly [Pum11]. news [Lüt19]. Newton [Art11, Bak11, Sch15b, BF13, Coh18, Cow13, GG10, Gui09, Kas14, Ksi13, Lev16, Roo17, Wig10]. Newtonian [Fei19, Sta19b]. Nexus [Whi16, Sol13]. NHS [Ste16]. Niaz [Lyk16]. Niccolò [Art11]. Nicholas [Car16a, Chr10, Ree14, Chr10, Mil11b]. Nicholson [Wat14]. Nicolas [Web15]. Nidhal [Moo13]. Niebur [Boo12]. Nielsen [Gou14, Hal10]. Nieto [Far10b]. Nieto-Galan [Far10b]. Nigel [Cuo12, Tee11]. Nineteenth [And12a, Ayl19, Bar13a, Cul11, Dij12, Fre10, Hig10, Hoc16a, Hun19, Kir15, Mar12a, O’C17, Ree19a, SS12, Stu11, Tam16b, Wad11, War13, And13b, Bel18a, Bux18, Car09, Col16, Cow14, Csi17, Dar12, Don14, Fox12a, Fra16, Mol14, Mos14, Tam14b, Wal19b, Yal15, Ade09, Alb11b, ABS10, Bat10, Cla09, LW11, McA13, Mül11, Pye12, Qur11, Rec13, War12]. Nineteenth-Century [And12a, Ayl19, Cul11, Hig10, Hoc16a, Hun19, Kir15, O’C17, Ree19a, SS12, Stu11, Tam16b, Wad11, War13, And13b, Bel18a, Bux18, Cow14, Fox12a, Mol14, Mos14, Tam14b, Wal19b, Yal15, Ade09, Alb11b, ABS10, Bat10, Cla09, LW11, McA13, Mül11, Pye12, Qur11, Rec13]. NJ [Bec18c, Hah16, LD18, Whi16]. No [Fin10b, Keh12, Ste11a, Wat14]. Nobel [Rec10, Gil12, Nor10]. Nobelpreis [Hen11b, Rec10]. nobility [Opi14b]. Noble [Rit10, Sle12]. Noel [Cuo12]. nomenclature [Opi14b]. nomenclatures [Ban14b]. Non [Siv10, Sel08]. Non-Western [Siv10, Sel08]. Nonproliferation [Mod17]. Noon [See12].

Nordic [Røs15a]. Norrby [Gil12]. North [Fin10b, Hay19b, Ksi15, Les13a, Mar18b, Pro11, Ree16b, Sch15b, Wai18, Byr13, Kas14, Koh10, SDS13]. Northern [Car18, ML10b, Elt12]. Norton [Smi10b]. Norwich [Keh12]. Notebook [Les15, WC13]. Notebooks [Rad13, Yeo14, CV09, Roo15]. Notes [Sch15b, Chr10, Kas14]. Notorious [Stu11, Ros10]. novel [DeW13, Daw14]. Novelties [Tar15, TB13]. November [Ban15, CJ10, Sta11b]. Nuclear [Bur12, Cal13, Hal14b, McR14, Moo10, Røs15b, Sar14, Sch12, Smi11b, Sto14, Und12, Wal12c, Zai16, Bin12, Bro12, Cor12, Hec17, Hog12, HL12, Hug12, Joh12, JJ14, Lau12, Mag12, Tur12a, Wal12a, Wea12, Wyn11, Lau15, Nav14a, Smi11b]. Numbered [Wil16b]. Numbers [Bec12, Can19b, Tur10, Bar14, Har08a]. Nursery [Kee12, TV11]. Nurture [Rus11a, Kel10]. Nuts [Hol16c, Han15]. NY [Bud13a, Han16, Rot16, San15a, Smi18b, Sta14]. nymphs [Opi14b].

O [Boa15, Fin10a]. Oakland

[Boo17, Lou19]. Oberheim [Kid11b]. Obituaries [Fie10, Ech12]. Obituary [CJ10]. Object [Can19a, ML10a]. Objects [Mac10, Söd10, Alb09, ML10a, Tur07, Tur08a]. O’Brien [Tay14]. Observation [Hig13, Nas10, Web17, Mah18]. Observations [Pic12, HA12]. Observatories [Big10, Hig10, Hig11, PdCB10, ABS10, Hut08, McA13]. Observatory [Spi15, Hig14b, Hig14b, Hig19, Mac15]. Observed [Mil11d]. Observer [Hig11, PdCB10]. Observers [Val14, Coe13]. Occult [Fan12, McW11]. Ocean [Ham16, Pay16, Mah18]. Oceans [Gil16a]. O’Connor [Sal15]. October [Fie10]. odd [Jar15]. Oddi [Rap12, Mar11]. O’Dell [Røs15a]. O’Donnell [Lea16]. Oedipus [Gre14b, Sto13]. off [Spi11, Spi12b]. Office [Smi10a, Ano17g]. Official [Zai16]. Ogle [Joh16]. Ohad [Cle16]. Oikonomia [BT14, Mai12b]. Oken [Gam17, Gam18]. Olaf [Gli11, Gli12]. Old [Sch11, Wit11a]. Oldenziel [McG15]. Oldfield [Roe17]. Ole [Wit11b]. O’Leary [Ade10, Cul13]. Oliver [Gra11, Dij12, Goo11a, Mah09]. Olivier [Dih16, Tat18]. Olmstead [Cha10]. Olschki [Col13, Tar15]. Omens [Bec12, HNS11]. Omodeo [Car18, Mos16a]. One [Bix19, HA19b, Kus18, Wag15, Kin13, Mit18]. One-Sex [Wag15, Kin13]. Oneworld [Bar15a]. Onion [Wai18]. Onno [Kus18]. Onwards [Tam14b]. Open [Sau15, Whi16, CC14]. Openness [Ver12, Bia12]. Operations [Ens12, Pri10]. Opitz [Buk14, Kee17]. Oppenheimer [Mon12, Und13]. Opponents [Nav15b, Waz14]. Optical [Vos11, Lev09, ZR15]. Optics [Dij12, Van14, Dar12, McG13, QM13, Eve18]. oral [Mer19]. orbe [Mim15]. orchids [End16]. Order [Edw14, Gou16, Van13, Gib15, Leh19, Lew12, Pri15, Kni16]. Ordidge [Fin10b]. ordinals [Tur38]. ordinary [Hei18]. Ordnung [Pri19b]. Oregon [Pis17]. Oren [Ber15]. Oreskes [Aga15]. Organic [Hol10, Jar17, Les15, Ree16b, Gig09, Ste14a, WC13]. organicism [Pet14]. organism [HA19a]. Organisms [Man17, Wad19]. organization [Spi12a, Pin13]. organizer [Pet14]. Oriental [Car12, Abt11]. Orientalism [Dew09, Mil11b]. Origin [Goo12b, Hoq10, Ksi13, BF13, Ene14, Gro17, JS08, Dil11, RR09, Dil12, Dil13, Jin19]. original [Mim15]. Originally [Rad14]. origine [Chr10, JS08]. Origins [Fal18, Gou13b, Mar12a, Mil12, Pin12, Web10, WM11, Bro10a, CGC13a, CGC13b, Des09, Edd17, Gli08, Man15, Mil08, Poo10, Ren11, SFFC12, Tom12b, VB11, McO10, Pea18]. Orit [vL16]. Ornithology [Wal18]. Oronce [Joh11, Mar09]. Orphanage [Web17]. Ørsted [KK14, Chr13]. Orthodox [Mak13]. Ortolano [Jam10a]. Osbeck [Lan19]. Osler [Sta11b]. Osseo [Coz15]. Osseo-Asare [Coz15]. Oswald [Roo12c]. Other [Kus18, Mer19, ML10a, Str19]. Ottmar [Mar14a]. Otto [Hig10]. Ottoman [Mil11d, Tam16b, Bil15b, Yal15]. Ottomans [Yal17]. Our [Ben13b, Les14, Row10, Wil16b, BW14, Edg09, Fer11b, Nie14]. ours [Pum11]. outdoors [Hod11]. Outer [Aga13a, Gep12]. Outlining [Cal12]. Outsider [HD13, Ber15]. Outward [Dra17]. Own [Ben13b, Bix19, Les14]. ownership [Jon14a, Mac16]. Oxford [Ale11, Bad16, Bar13a, Bec18c, Bha11, Biv10, Bix19, Bjö11, Bjö14, Bla17, Boo12, Bur11, Can15, Can19a, Car11, Cle16, Daw15b, Daw16a, Dih16, Dij12, Dun12a, Eic12b, Eva13, Eve13, Ful10a, Gou13b, Gou17, Hei15, Hen14, Hen15a, Hen15b, Hen13a, Hen13b, HA19b, Hug10, Jam11b, Jim19, Jon10a, KK14, Kip16, Kni15, Kni16, Kot16, Ksi13, Lan11a, Lan12, Lau15, LD18, Lou17, Mac13, Man16, Mar17a, Mar15, Mca16, Mil11b, Mil13, Mil15, Mil11c, Mil11d, Mor15b, Mor13, Mur12, Nas16, Nav15a, Nav16, Opi14a, Pet10a, Pic10a, Pic10b, Pis16, Pos15a, Pyl10, Reg17, Rob14, Sau13, Sch15a, SS16, SS17, Sha10, Sle11a, Smi12, Ste16, Tat18, Tay14, Tee11, Tys17, Und12, Val11, Vos11, Wad11, Win15, Win11, WM11, dC14, vL13b]. Oxford [BL10, BF14, Hes09, RS09, SK12]. oxygen [Heg13].

P [Dun16a, Kir13, Lee15b, Mor18, Ser15, vN19, vL13a]. P. [Far11, Rob15, Roo12c]. PA [Ayl19, Dun16b, Fal18, Hen19, Ree19a, Sim17a]. Pablo [Tun19]. Pacific [Bil18, KMB18, Lüt19, Par18]. Package [Sta11d, Jon10c]. Packham [Kno14]. Paderborn [Ben19a]. Padua [Feo18]. Pain [Bod14, Bui19, Mil15, Bod19, Mil15]. Painkillers [Mil15]. Pakistan [Abr09, Pha10]. palaeontology [Goo16]. Paleobiology [She13, Sep12]. Paleolithic [Rie16]. Paleontology [Wyl11, Bri10]. Palgrave [Ada10, Aga13a, Aga14, Ase17, Ber17, Ber18, Bow18, Cal12, Car16b, Daw17, Doh16, Duf18, Dun11, Far13, Fin11, GP18, Goo10a, Ham14a, Ham16, Har16, Hoc14, Hoc16a, Joh13a, Kaa18, Kee17, Kno14, Köb16, Ksi15, Lan15a, Lin17, Mat13, McG15, Mil15, Mor10a, Mye12, Nor12, Pay16, Qui12, Ree10, Rob17, Rod15a, Rod15b, Ros16, San13, Sim14, Smi11b, Sta19c, Tra16, Wes15, Wol14]. palimpsest [N^{+}11, Cuo12]. Palmira [Gas10]. Pamela [Fin10a, Fin10a]. Panic [SJ19]. Pankaj [Kum18]. Panofsky [Arb17]. Panopticon [Bod19]. Paola [Hig16]. Paolo [Dun17a, Nay14]. Papanelopoulu [Far10b]. Paper [Mul13, San12, Fis10, Mue11, SF11]. paperback [Adl16, Aga14, Aga15, Aga17a, Ako19, And12d, Arb17, Arn16, Ayl18, Ban12, Bar13a, Bar13b, Bar15a, Bar15b, Bar15c, Bec12, Bel12, Ber15, Bil17, Boa10, Bod18, Boo12, Boo17, Bou15, Bow16, Bro11a, Bus18, Cal12, Can15, Can19b, Cap14, Car16a, Car18, Cha14, Cha10, Chr10, Cle15, Cle18, Col13, Coz15, Dan19, Dra17, Dun12a, Dun12b, Dun17a, Egm12, Eic12a, Elt12, Esp10, Eve18, Fix16, Fix19, Fla19, For10, Fre10, Gar14, Gil15a, Gil12, Gli11, Gli12, Goo11a, Goo12a, Goo10b, Gou12c, Gra11, Gri15, Hal14b, Ham14b, Hay18, Hed13, Hei15, Hig10, Hol13a, Hol16b, Hoq10, Huf16, Hun18, Hun19, Joh15, Keh12, Kid11a, Kid11b, Kip16, Kne10, Lan15b, Leh12a, Les13a, Les14, Lo15, Loi16, Lou19, Man13a, Man13b, Mar14a, Mar18b, May17, Mey11, Mic11, Mil11a, Mil17]. paperback [Moo17, Moo13, Mor13, Mul13, Nav15a, Pha10, Pic10a, Pis17, Pos15a, Pri19a, Pri13c, Rad13, Rad17, Ree16a, Ree16b, Ree18, Ren19, Ren10, Ric11b, Ric10, Rob15, Rod15a, Rod15b, Row10, Rud13, Rus11a, Sal14, San12, Sar13, Ser14b, Ser15, Sha10, Sim13, Sim17b, Sin15, Smi11c, Smi12, Smi16b, Som15, SM19, Tar15, Tho19, TS14, Tos11, Van11b, Ver15a, Wai18, Wal11a, Wal12c, Web11, Web14, Web15, Wri16, Zai16, dC14, vN19, vL16]. Papers [Daw15b, Ano17g, Csi17, GG10, HMLE15, Sim10]. Paracelsus [Cun10b, Web08]. Parade [SS12, Qur11]. Paradise [Tat16, Dan14]. Paradox [Ser15, Heu15, PB13]. Parapegmata [Rob10b, Leh07]. Parapsychology [Som11, Wol09]. Parasites [Dev16]. Pare [Gai17]. Paris [Bus18, Chr10, Esp10, Gis19, Hed13, Kne10, Man13a, Mur12, Ste14b, Ver15a, BJ10, CMC08, Jon14b, Ker14, Mor15b, Sho18, Spa12]. Parker [Gou12c]. Parliamentary [Cou17]. Parshall [Hol15]. Parsons [Kir15, Mol14]. part [Mit18]. Participation [Wal12c, Lun13, Wyn11]. partnerships [Nie10b, Zal14]. party [BG11]. Pascal [Sim13, TS14, Cam12]. Pasetti [Col13]. Passage [Coe10, Eve18, Wal09]. Past [Ano17g, Dih16, Ric12, Dar14, Far18, Hes11, Rad16, Ren12, Rie16, Joh13a]. Patent [Dun15b, GMML14, Jon16b]. patenting [Gab16, Mac16, Sta16b]. Patents [Joh14, Ill12]. Path [Fra17, Van11c]. Pathological [Bel18b]. Pathology [Wag14, AF16, Mit12]. paths [Gue19]. Pathways [Ham14a, TlT12]. Patients [Goo10a, Ham14a, ANPR07, TlT12]. Patricia [Bix19]. Patrick [Aga17a, Mil11d, Ren10, Rob14, Tar15, Fin17, Fla19, Oos12]. Patriots [Tam16b, Yal15]. Patrizi [Gou16, Pri15]. Patronage [Whi16, Sol13]. Patterns [Boa12, McE10]. Patteson [Boo17]. Pattison [Van11b]. Paul [Ade10, Cha10, Gou12c, Kid11a, Law17, Ser15, Arn16, Fot18, Kid11b, Lin17, Mai11, Man13b, Rei15, Sim14, Tee11, Wig11, Wyl11]. Paula [Ase17, Kus18, Smi11d]. Pauly [Cha10]. Payne [Ken19]. Pearl [Gar11, RCK10]. Pearls [Dud14, Hol16a]. Pedagogy [Rud13, Sau15, AD12, BN13].

Peder [Har16]. Pedersen [Hig11]. Peeters [Tho16a]. Peking [Goo10b, Sch08b]. Pemberton [Cha14, Pea19]. Penicillin [Bud08, Pic10a]. Penn [Sch17b]. Pennsylvania [Bec18a, Cur18, McC19, Par12, Stu11]. Penultimate [Tys17]. People [Ben13a, Goo10b, McG15, Lau12, Sch08b, Sin15]. Peoples [Qur11, SS12]. Pepys [Par18]. Perceived [Pin13, Spi12a]. perception [Kle13]. Perceptions [Tam14a, Per12]. Percivall [Ken19]. Perdiguero [Far10b]. Pere [Dan19]. Perfection [Ser14a, Com12]. Performance [Tho16a, VPW15]. Performing [Bro11b, Bui19, Sta12b]. Period [D’A18, Egm12, Ham16, Pay16, Rob11]. periodicals [Bel18a, Wal19b]. Periphery [Far10b, PNGP09]. Permanent [Ban14a]. Perovic [Tam14a]. Person [Mil16, Smi15]. Personalized [Stu17]. Perspective [BZ11a, Row10, BJW11, BG14, Bur16, Cro12, Cur14, Edg09, Fer13, FJB14, Huf11, San10a, TW10, Bou15, Goo12a, Man13b, Mar18a, Mic11, Pea12, Pos15a, Ree16a, Rod15a, Rod15b]. Perspectives [Bur13, Dun11, Hen15b, Roe17, Smi18a, Ste11b, dRLE09, FMW14, Llo09, Mor10b, TlT12, Bur13, Ham14a]. Perugia [Pri19a]. Peter [Bar13a, Bro16, Fin10b, Gil16a, Goo12a, Hal10, Hig11, Jim19, Ken19, Par12, Pic10b, Pri19b, San15b, Sau13, Sim13, Wal11a, Wit11b, WM11, Ade10, Bec12, Bjö14, Boo11, Can16, Dun11, Dun12b, Gra11, Hal17, Ham14b, Hen13a, Lan12, Smi14, Smo10, Som15, Tim12]. Peters [Tat17a]. Peterson [Jar17]. Pettit [Whi15]. Petty [Edw14, Lew12]. Phaenomena [APoC10, Leh12a]. Phalkey [Ban14a]. Pharmaceutical [Gab16]. pharmakon [Win19b]. Philadelphia [Bec18a, Cur18, Fre10, Gil15a, Gra11, Kao12, McC19, Par12, Ric10, Sch17b, Smi11c, Stu11]. Philip [Ins19, Mar12a, Cha10]. Philippa [Bjö11]. Philippe [Chr10]. Phillip [San12]. Phillips [Web14]. philology [Bil18]. Philosopher [Cha09, Sum11, Bal08, Cam09, Tau10, Sha10, Smi12, Sta11c]. Philosophical [Boa10, Fre10, Gar11, Gil15a, Gra11, Ric10, Smi11c, Ste11b, dRLE09, RCK10, Vol08]. philosophie [Gis19]. Philosophies [Esp10, Del09]. Philosophique [Hed13]. Philosophy [Bar15b, Cha09, Cle15, Cle17, Dun14, Gar11, Hir11, Ins19, Lyk16, Mos16b, Roo12a, Row14, Ser14b, Ste10, Sum11, Til10, Van11c, A^{+}12, Ans11, BT13, Bor08, Chi08, DS08, Fox12b, Gla09, Hac14, Law15b, Lev16, Man12, Mar14b, Nav14b, RCK10, vB13, Bur11, Cle16, Dob14, Hen10, Lan12]. phlogiston [BG11]. Photography [Bel12, Pin11, BDR13, MR13, Die15]. phrenological [Sys18]. Phrenology [Jen15]. Phyllis [VM18]. Physical [Hig14b, San12, SF11]. Physician [Gil16a]. Physicist [Nav14a, Sta11c, Cam09, Sch12, Hun19]. Physics [BF14, Bur11, Dar14, Hen19, JS17, Mil11a, Sau15, TW10, Wes12b, BN13, Bal14a, Cas11, Gla09, Hay19a, Hun10, Mit18, Nav10, Tur12a, Ver19, Bad16, Cal13, Dih16, Dih17, Mic11, Sch15a]. Physiological [Sta16a, Bra15]. Physiology [Kle13, Nes10, Sho14, Heg13, HKZ12, Kic09]. phytonymy [Bil18]. Piccolomini [Cla14]. Pick [Can15]. Pickering [Alb11a, Ayl17, Bak15b, Boa12, Coo11, Cow10, Cul11, Daw15c, Gil11, Goo12b, Ham15, Hom16, Hul10, Jac15, Ken13, Lou15, Mar12a, Rad14, Raj15, Ric12, Sal15, Sch16, Smo10, Tho16a, Wit11a]. Pickstone [Goo10a]. pictures [Ell17]. Picturing [Kus11, Rem11, Egm13, Kao12]. Pierre [Mit18]. Piers [Con15, Wag14]. Pieter [DP17, Ver16]. Pietism [Web17]. Pietro [Car18, Kne10, Mos16a]. Pigs [Wad19]. Pill [Löw11]. Pinney [Bel12]. Pioneers [Mar12b, AM11]. Pips [Hol16c, Han15].

Piracy [Joh09, Mac11a]. Pisa [Kne10, CMC08]. Pittsburgh [Ayl17, Ayl19, Dun16b, Fal18, Fin17, Hen19, Hol16b, Ins19, Jar17, Ree16a, Ree19a, Sim17a, Ste11b, Tun19]. PKU [Ser15, PB13]. Place [Ree16a, FJB14, KT19, Pea18]. placement [OG17]. Places [Hig14a, Mis13]. Placing [Alb11a, Leg13, Nay10]. Plains [Cha10, Pau07]. Planet [Ree16a, FJB14, Lan11b, Nal12]. Planning [Szu18]. Plant [Hol16c, Smo10, Ayr08, Han15]. Plants [Bec16, Coz15, Smo10, WM18a, Ayr08, Bal19, OA14]. Platonic [Gou17]. Playing [HHR09, Gra11]. PLC [Mar17b]. Pleasure [Bod19, SJ19]. Plotnick [SJ19]. Pluralism [Boa15]. plurality [Sta16b]. Plus [Kne10]. Podolsky [Coz16]. Poetics [Bla17, Duf18]. Poetry [Bil15a, Daw10, Hol09, Pri13a]. Poincaré [Hen13b, Gra13]. Point [Bur11, BG11, Gla09]. Poison [Kum16, Man16, Mil16, Sch15c, Smi15]. Polar [Jam11a, Lam09]. Policies [Coz17]. Politeness [Doh16]. Politic [Hen13a, Elm13]. Political [BT14, Hal14a, Mar12c, Mor10a, Von11, Vos11, Bor09, Lev09, Mai12b, Ren11, Ber17, Con15]. Politics [And12c, Arn16, Con15, D’A16, Das18, Dun16b, Far13, Hen13a, Jam10a, LN14, Law17, Man17, Mic11, Ree16b, SJ19, War13, Web18b, Whi16, Azz12, CC14, Dev16, Edd17, Eff17, Elm13, Fox12a, Hal14a, Hud14, Lun13, Lyn12, Mac15, McM18, Med11, Mou17, Ort09, Pac12, Per12, Sol13, Ste14a, Sva18, TW10, Kno14, Tam14a]. Polity [For10, Gri15, Kid11b, Ren10, Tho19]. Pollution [Kum16]. Polymathy [Car18]. Pontecorvo [Cal13, Tur12a]. Poochigian [Leh12a]. Poole [WM11]. Popular [Goo10b, Wai18, Bin12, Ree16d, Sch08b]. Popularisation [Kee12, TV11]. Popularization [Boo11, Bow09]. Popularizing [PNGP09, Far10b].

Population [Ren15, Bas14]. Porta [Tar13]. Portuguese [Smi11d, B^{+}09]. Portuondo [Tes11]. posguerra [Ric11b]. positional [Don14]. Possibility [Fan12, McW11]. Post [Nes10, Ste16, Kic09, Vic11]. Post-Crimean [Nes10, Kic09]. Post-war [Ste16, Vic11]. Postcolonial [Nie10b]. Postgenomics [Ayl18]. Postwar [Jam10a, Ort09]. Pott [Ken19]. £7.99 [Mor13]. Power [Mil11a, Wal19a, Asp11, Hun10, JS17, Sum14b, Zal14, SJ19]. Powered [Sta15, Fyf12, Tur13, Daw12]. Powers [Roo12b]. pow’r [Opi14b]. Pp [AN18, Ada10, Ade10, Adl16, Aga13a, Aga14, Aga15, Aga17a, Aga17b, Ako19, Alb11a, Alb15, Alb16, Ale11, And12a, And12d, And12c, And12b, Arb17, Arn16, Art11, Ase17, Ayl17, Ayl18, Ayl19, Bad16, Bad19, Bak11, Bak14, Bak15b, Bak15a, Bal14b, Ban14a, Ban15, Ban12, Bar13a, Bar13b, Bar15a, Bar15b, Bar16, Bar17, Bar15c, Bec12, Bec16, Bec18a, Bec18b, Bec18c, BT14, Bel12, Bel17, Bel18b, BZ11a, Ben19a, Ber11a, Ber12, Ber15, Ber17, Ber18, BS16, Bha11, Big10, Big18, Bil15a, Bil17, Biv10, Bix19, Bjö11, Bjö14, Bla17, Boa10, Boa12, Boa15, Boa18, Bod18, Bod19, Boo11, Boo12, Boo14, Boo17, Bou15, Bow16, Bow18, Bow10, Bow17, Bro11a, Bro16, Buc11, Bud11, Bud13a, Bui19, Buk14, Bur11, Bur13, Bus18, Cal12, Cal13, Can15, Can16, Can19b, Can19a]. Pp [Cap14, Cap15, Car16a, Car18, Car16b, Car11, Car12, Cha14, Cha10, Chi12, Chr10, Cle14, Cle15, Cle16, Cle17, Cle18, Coe10, Col13, Con15, Coo11, Coo14, Cow10, Coz15, Coz16, Coz17, Cul11, Cul13, Cun10b, Cuo12, Cuo13, Cur18, D’A16, D’A18, Dan19, Dar11, Das18, Daw17, Daw10, Daw12, Daw14, Daw15a, Daw15b, Daw15c, Daw16a, Daw16b, Daw16c, Die15, Dih16, Dih17, Dij12, Dob14, Doh14, Doh16, Dra17, Dry10, Dub10, Duf18, Dun11, Dun12a, Dun12b, Dun14, Dun15a, Dun15b, Dun16a, Dun16b, Dun17a, Dun17b, Dys14, ES16, Ebe17, Ebe18, Edw14, Egm12, Egm13, Eic12a, Eic12b, Eic15, Elt12, Ens12, Esp10, Eva13, Eve13, Eve18, Fal18, Fal13a, Fan12, Far10a, Far10b, Far11, Far12, Far13, Far15, Far16, Far19, Feo18, Fer11a, Fin10a, Fin10c, Fin11]. Pp [Fin17, Fix16, Fix19, Fla19, Fon12, For10, Fot18, Fra17, Fre10, Ful18, Ful10a, GP18, Gar11, Gar14, Gas10, Gib14, Gif16, Gil15a, Gil16a, Gil12, Gil16b, Gil11, Gis19, Gli11, Gli12, Goo10a, Goo11a, Goo12a, Goo18, Goo17, Goo10b, Goo12b, Gou12a, Gou13a, Gou13b, Gou16, Gou17, Gou12b, Gou12c, Gou14, Gra11, GG12, Gre14b, Gri15, Gul18, Hah16, Hai16, Hal10, Hal14b, Hal17, Ham14a, Ham10, Ham14b, Ham15, Ham16, Han17, Han16, Har16, Har10a, Hay18, Hay19b, Hed13, Hei15, Hem19, Hen14, Hen15a, Hen15b, Hen10, Hen13a, Hen17, Hen11a, Hen11b, Hen13b, Hen19, Hig10, Hig11, Hig13, Hig14a, Hig15, Hig16, HA16, Hoc14, Hoc16a, Hol13a, Hol14a, Hol15, Hol16a, Hol16b, Hol10, Hol16c, Hom16, HA19b, Hoq10, Huf16, Hug10, Hul10]. Pp [Hun18, Hun19, Hus15, Ins19, Jac15, Jam10a, Jam11a, Jam11b, Jam12, Jam13, Jam17, Jam18, Jar10, Jar16, Jar17, Jim19, Joh15, Joh13a, Joh11, Joh13b, Joh14, Joh16, Jon10a, Jon10b, Kaa18, Kao12, KK14, Kee12, Kee17, Keh12, Kel17, Ken13, Ken19, Kid11a, Kid11b, Kip16, Kir13, Kir15, Kne10, Kni15, Kni16, Kno14, Köb16, Kot16, Ksi13, Ksi15, Ksi16, Kum16, Kum18, Kus18, Lan11a, Lan12, Lan15a, Lan19, Lan14, Lan15b, LN14, Lau15, Law17, Lea16, Lee15b, Leh12a, Les13b, Les13a, Les14, Les15, LD18, Lin17, Lo15, Loi16, Lou15, Lou16a, Lou16b, Lou17, Lou19, Lyk16, Mac11b, Mac13, Mac10, Mac11a, Mae12, Mai12a, Mai11, Mai12c, Man13a, Man13b, Man17, Man16, Mar16, Mar17a, Mar18a, Mar12a, Mar15, Mar12b, Mar12c, Mar14a]. Pp [Mar17b, Mar18b, Mat13, May17, McC19, McG15, McO10, Mca16, Met11, Mey11, Mic11, Mil11a, Mil11b, Mil12, Mil13, Mil15, Mil16, Mil17, Mil11c, Mil11d, Mis11, Mod17, Mon10, Moo13, Mor10a, Mor18, Mor15b, Mor13, Mos16b, Mos16a, Mox16a, Mul13, MY18, Mur12, Mye12, Nal12, Nal15, Nas16, Nav14a, Nav15a, Nav16, Nes10, Nic17a, Nic17b, Nic19, Nor12, O’C17, OP16, Oos12, Opi14a, Osl10, Par11, Par12, Pay16, Pea18, Pea12, Pea19, Pet10a, Pha10, Pic10a, Pic10b, Pin12, Pin13, Pis16, Pis17, Pos15a, Pri12, Pri19a, Pri13c, Pri19b, Prk17, Prk18, Pro11, Pum12, Pyl10, Qui12, Rad13, Rad14, Rad17, Raj15, Rap11, Rap12, Ree14, Ree16a, Ree16b, Ree17, Ree18, Ree19a, Ree15, Ree10, Reg17, Rei15, Ren19, Ren10, Ren15, Ric12, Ric11b, Ric18]. Pp [Ric10, Ric14, Ric16, Rie16, Rih12, Rob10a, Rob14, Rob15, Rob17, Rob10b, Rod15a, Rod15b, Roe17, Roo12a, Roo12b, Roo12c, Roo15, Ros17, Ros16, Røs15b, Røs15a, Rot16, Row10, Row14, Rud13, Rus11a, Sac16a, Sac16b, Sac19, Sal14, Sal15, San12, San13, San14, San15a, San15b, San16, San17, Sar13, Sar14, SJ19, Sau13, Sau15, SS11, Sch17a, Sch15a, Sch15b, Sch16, Sch17b, Sch19, Sea10, SS12, SS13a, SS15, SS16, SS17, Ser12, Ser14b, Ser14a, Ser15, Sha10, Sha11, She13, Shi16, Sho14, Sho18, Sia18, Sim14, Sim17a, Sim13, Sim16, Sim17b, Sin15, Siv10, Sle10, Sle11a, Sle12, Sma13, Smi10b, Smi10c, Smi11d, Smi11c, Smi11b, Smi12, Smi14, Smi16b, Smi16a, Smi18a, Smi18b, Smi19, Smo10, Sno10, Söd10, Som11, Som15, Sør17, Sta11a]. Pp [Sta11c, Sta11d, Sta12b, Sta14, Sta15, Sta16a, Sta19c, Ste10, Ste11b, Ste14b, Ste16, SM19, Sto14, Sto12, Str11, Stu10, Stu11, Stu17, Sum11, SJ11, Tam14a, Tam16b, Tam16a, Tar12, Tar15, Tat17a, Tat17b, Tat16, Tat18, Tau11, Tay14, Tee11, Tes11, Tho16a, Tho19, TS14, Tho16b, Til10, Tim12, Tos11, Tot10, Tra16, Tun19, Tur16a, Tur10, Tur14, Tur15, Tur16b, Tur12b, Tys17, Und12, Und13, Val14, Val11, Van11b, Van11c, Van14, Ver15a, Ver16, Vol16, Vos11, Wad11, Wad19, Wag14, Wag15, Wai18, Wal16a, Wal18, Wal19a, Wal11a, Wal12c, Wal16b, Wal17, Wal15, Wal11b, War13, War19, Web10, Web11, Web13, Web14, Web15, Web17, Web18a, Web18b, Wes15, Wes12b, Whi15, Whi16, Wig11, Wil11a, Wil11b, Wil16a, Wil12, Wil14]. Pp [Wil16b, Win15, Win11, Win13, Wit11a, Wit11b, Wol11, Wol14, WM11, WM18a, Wri16, Wyl11, Yal17, YM10, Zai16, dC14, vN19, vL13a, vL13b, vL15, vL16, Moo17, Pos14]. Practical [Bee19, Joh14, Ill12]. Prácticas [Car16a, Nav14b]. Practice [Bak15a, Dar11, ES16, Fin17, Lo15, Sør17, Wal11b, Bee19, BH19, Bow11, CVLW14, GJ09a, GJ09b, HSSR13, Ree16d, Set10, Ter14]. practices [Ech12, Nav14b, Pod13]. practitioners [Jon16b, ZR15]. Prague [Gli12]. Praktiken [Ech12, Web14]. pratiche [Cle18]. Pratik [Sta14]. Prayer [Mil15]. Preaching [Toa11, Toa12]. precedents [Fri11]. precept [May10]. Preface [Boa10, Esp10]. pregnancy [Kir17, OG17]. Prehistoric [O’C17, Eff17]. Prehistory [Alb16, Bar14, Goo16, Hoc16b, Ree16d]. préparatoires [Chr10]. Present [Mar12b, AM11, Dar14, Dih16]. presentation [Ver10]. Presidential [Rad16]. Press [Ade10, Adl16, Aga15, Aga17a, Aga17b, Alb15, Alb16, Ale11, And12a, And12c, And12b, Arb17, Arn16, Art11, Ayl17, Ayl18, Ayl19, Bad16, Bad19, Bal14b, Ban15, Ban12, Bar13b, Bar15b, Bar17, Bar15c, Bec12, Bec18a, Bec18c, Bel17, Bel18b, BZ11a, Ber12, Ber15, Bha11, Biv10, Bix19, Bjö11, Bjö14, Bla17, Boa18, Bod19, Boo11, Boo12, Boo14, Boo17, Bou15, Bow16, Bow17, Bro11a, Bro16, Buc11, Bud11, Bud13a, Bui19, Bur11, Cal13, Can15, Can16, Can19b, Can19a, Cap15, Car11, Car12, Cha10, Chi12, Cle14, Cle15, Cle16, Coe10, Con15, Coo14, Coz15, Coz16, Cul13, Cun10b, Cuo12, Cuo13, Cur18, Dan19, Dar11, Das18, Daw10, Daw12, Daw14, Daw15a, Daw15b, Daw16b, Die15, Dih16, Dih17, Dij12, Dra17, Dry10, Dun12a, Dun16a, Dun16b, ES16, Ebe17]. Press [Ebe18, Egm13, Eic12a, Eic12b, Eva13, Eve13, Eve18, Fal18, Far10a, Far11, Far16, Far19, Fin10a, Fin10c, Fin17, Fra17, Ful10a, Gar14, Gib14, Gif16, Gil12, Gil16b, Goo11a, Goo12a, Goo18, Goo10b, Gou13b, Gou17, Gou12b, Gou12c, Gou14, Gre14b, Gri15, Hah16, Hai16, Hal10, Hal14b, Hal17, Ham10, Han16, Har10a, Hay18, Hay19b, Hem19, Hen14, Hen15a, Hen15b, Hen13a, Hen17, Hen13b, Hen19, Hig10, HA16, Hol13a, Hol14a, Hol15, Hol16b, Hol10, HA19b, Hoq10, Huf16, Hug10, Hun18, Hus15, Ins19, Jam10a, Jam11b, Jam12, Jam17, Jam18, Jar16, Jar17, Joh15, Joh13b, Joh14, Joh16, Jon10a, Jon10b, Kao12, KK14, Kel17, Kip16, Kir13, Kir15, Kne10, Kni15, Kni16, Kot16, Ksi13, Ksi16, Kum16, Kus15, Lan11a, Lan12, Lan19, Lan14, Lan15b]. Press [LN14, Lau15, Law17, Lea16, Lee15b, Leh12a, Les13b, Les13a, Les14, Les15, LD18, Lo15, Lou16b, Lou17, Lou19, Mac13, Mac10, Mac11a, Mae12, Mai12a, Mai11, Man13b, Man17, Man16, Mar16, Mar17a, Mar18a, Mar15, Mar12c, Mar18b, May17, McC19, Mca16, Met11, Mey11, Mil11a, Mil11b, Mil12, Mil13, Mil15, Mil16, Mil17, Mil11c, Mil11d, Mis11, Mod17, Mon10, Moo17, Mor15b, Mor13, Mos16b, Mul13, MY18, Mur12, Nal12, Nal15, Nas16, Nav14a, Nav15a, Nav15b, Nav16, Nic17a, Nic17b, O’C17, OP16, Opi14a, Osl10, Par11, Par12, Pea18, Pea12, Pea19, Pet10a, Pha10, Pic10a, Pic10b, Pin12, Pin13, Pis16, Pis17, Pos15a, Pri12, Prk17, Prk18, Pro11, Pyl10, Rad13, Rad14, Rad17, Rap12, Ree14, Ree16a, Ree16b, Ree17, Ree18, Ree19a, Reg17]. Press [Rei15, Ren19, Ren10, Ren15, Ric16, Rie16, Rih12, Rob14, Rob15, Rob10b, Roo12b, Roo15, Røs15a, Rot16, Row10, Rud13, Rus11a, Sac16a, Sac16b, Sac19, San12, San14, San15a, San15b, San16, San17, Sar13, Sar14, SJ19, Sau13, SS11, Sch15a, Sch17b, Sch19, Sea10, SS12, SS13a, SS15, SS16, SS17, Ser12, Ser14a, Ser15, She13, Shi16, Sho18, Sim17a, Sin15, Sle10, Sle11a, Sle12, Smi10c, Smi11d, Smi12, Smi14, Smi16a, Smi18b, Smi19, Sno10, Söd10, Som15, Sta11c, Sta11d, Sta12b, Sta14, Sta15, Sta16a, Ste10, Ste11b, Ste14b, Ste16, SM19, Str11, Stu10, Stu11, SJ11, Tam14a, Tam16b, Tat17a, Tat17b, Tat16, Tat18, Tau11, Tay14, Tee11, Tes11, TS14, Tho16b, Tot10, Tun19, Tur10, Tur14, Tur15, Tur16b, Tys17]. Press [Und12, Val14, Val11, Van11b, Van14, Vol16, Vos11, Wad11, Wad19, Wai18, Wal16a, Wal18, Wal19a, Wal15, Wal11b, War13, War19, Web10, Web14, Web15, Web17, Web18a, Web18b, Wes12b, Whi15, Whi16, Wil12, Wil14, Wil16b, Win15, Win11, Wol11, Wyl11, Yal17, vN19, vL13b, vL16, Bin12, Dew14]. Presses [Esp10]. Prestel [Coe10]. prestige [RC11]. Preston [Bla17, Roo12c]. preuve [Coh11, Man13a]. Prevention [Loi16]. Price [Gli12, Fin10b]. Priestley [Ens12]. Priestman [Bil15a]. Primatology [Mon10, Ree09]. prime [AGF13]. Prince [Gil15a]. princely [Rat16]. Princeton [Ale11, Bec18c, Car11, Coo14, Eic12b, Hah16, Hen13b, Hol15, Jon10a, Ksi13, LD18, Mur12, Rob14, Sau13, SS17, Smi12, Win11, vL13b, App12]. Principe [San14, Wil11a]. Prins [Gou16]. Print [Bak15b, Joh11, Lou16b, KWB15, Mar09, Sum13]. printing [Ban14b]. Priorities [Lan19]. Priority [RC11, Bia12]. Private [Hol16b]. Prize [Rec10]. Prizes [Gil12, Nor10]. Probability [Tat18, TS14, Cam12]. Problem [Pea12, Ekh10, Lam13, Pet10b, San10a, Ayl17, Som15]. procedure [Nas10]. proceedings [A^{+}12]. process [Boo15b]. product [OG17]. Production [Goo12a, Mat13, BJW11, Kri12, Rat16, SC12]. Professionals [Tur16a, Gil15b]. Professor [Sta11b]. Profit [Web18a]. Prognostication [Van13]. Programm [Fan10, Sta11a]. Programmed [Tat17b].

Programmes [Sto14, Wal12a].

programming [Pri10, Ens12]. progress [AF16, Far18, Jen15]. Progressive [Mil12, Tom12b]. Project [Daw16b, Fin10b, Pet14]. projector [Yam15]. Projects [McC19]. promotion [Dev16]. Proof [Cuo13, Che12]. propagation [Val10]. proper [Hic19]. Property [Far16, Goo12a, Mac11a, BJW11, Hem15, Joh09]. Prophecy [DeV09, Sea10]. Prophet [May14, Smi19, Ren15]. Prophets [Nav15a, Smi10c, Ger09, Wal13]. Proportion [Leh19]. proposal [Ver19]. Propulsion [Gif16]. Prostate [Daw17]. protagonists [AM11, Mar12b]. Protect [Cor12]. Protect/Protest [Cor12]. protection [Cow13, Sta16b]. Proteins [San13, GS12]. Protest [Cor12]. Protestant [Met11]. Providence [Bar16, Hol16a, Sør17, Bax13]. provinces [JS17]. Provincial [Sta12b, Bro11b]. Provoked [Mur12, BJ10]. ps [Mim15]. Pseudo [Tur14]. Pseudo-science [Tur14]. pseudoscience [Gor12]. Psychiatry [Wes15, Kem19, Tom13b]. Psychical [Som11, Wol09]. Psychoanalysis [Prk17]. Psychological [Lan19, Som15, Lam13]. Psychology [Pri13c, Smi18a, Tho16b, Tur12b, Whi15, Goo11b, Pet13, Ric11a, VB11, Pin12]. Public [Ben13b, Han16, Hol16b, Les14, Mye12, Nav15b, San17, Sta11c, Car10, Dev16, Ros18, RC11, Waz14, Wil11c]. publication [Bal14a]. Publications [Rad13, Wil11a, Van09]. Publics [Pha10, Abr09]. published [Ano17g, Rad14]. Publishers [Boa10, Kus18]. Publishing [Daw12, Gas10, Kee12, Ken19, Lou16b, Lyk16, Mor18, Sal14, Bal14a, Fyf12, KWB15, Olo14]. Pulse [Sta16a, Bra15]. Pulses [Hol16c, Han15]. Pure [Jam17]. Pursued [Rob14, McC12]. Pursuing [Hun10, Mil11a]. Pursuit [Can15, Coz16, Sch19, Pic14, Pod15].

Pursuits [Dih16, Dar14]. Pushing [SJ19]. Puzzle [Pea18]. Pyrotechnic [Buc11, Wer10].

Q [Ebe18]. QBism [Dih17]. Quantum [Dar11, Dih16, Dih17, Moo13, Sau13, Sau15, Sta11c, BN13, Cam09, Dar14, BB12, Gue11, Nav10, Set10]. Quarks [Bad16]. quarrel [JS08]. quarter [Mar17b]. Queen [Bar08]. Quelle [Ver15a, RC13]. Querelle [Chr10, JS08]. Quest [Cle17, Gif16, Jar10, McO10, Opi14a, AGF13, Des09, Hol13b, Rat09]. Question [Moo13, Van13, Gue11]. Questions [Lan15b, Tys17, C^{+}12, Ric13a]. Quinlan [Van14]. Quinlan-McGrath [Van14]. Quintessence [Bad16]. Quotable [LD18]. Qureshi [SS12].

R [Arn16, Bak11, Ber15, Bow18, Bud11, Eva13, Fal13a, Feo18, Ful10a, Gil11, Hal14b, Hen14, Hig16, Jam13, Ken13, Lan12, Mai11, Ric16, Rud13, San12, Sia18, Sto14, Val14]. R&D [Spi11, Spi12b]. raccoon [Pet10b]. Race [Ayl19, GR18, Lau15, McO10, Ric11a, SS15, Bal14a, Bro10a, Des09, Goo16, JA15, Lor13, McM18, McR14, Goo12b, Pri13c]. Rachel [SJ19]. Racial [Coo14, Fot18, Les13b, Mar18b, Sch08a, Fer12, Ree10]. Racism [Alb16, Pri13c, Ric11a]. racist [GR18]. Rader [Alb15]. Radiation [Ren19]. Radick [Hoq10, Van11b]. Radio [GS13, Hug10, And09, Mar15]. Radioisotopes [Cap15, Cre13]. Radiophonic [Boo12, Nie10a]. Radkau [Ren10]. ragioni [Ekh10]. Railway [And12d, RH12]. Railways [Mai12c]. Ralf [Ree17]. Ralph [Sal15]. Ramachandran [Shi16]. Ramalingam [Die15]. Ramsden [Rad14, San15a]. Ramus [Pum12, Gou10]. Rana [Mar18b]. Ranikhet [Ban14a]. Rankin [Tar12]. Raphael [Mar11]. rare [ZR15]. Rasmussen [Web15]. Ratcliff [Jar10]. Rational [Coz16, Pod15].

Rationalist [Dih16, Dar14]. Rationality [Wyn11, EKD^{+}13, Rei15, Wal12c]. Ray [Duf18, Roo12c, Sac16b, OPR11, Und13]. Raymond [Hen15b, Ree17]. Rays [MY18, McG19]. Re [Lan19, Gal12, Leg13]. Re-imagining [Lan19]. re-placing [Leg13]. re-territorialization [Gal12]. Read [Ale11, Wil11b, War10]. reader [KB09, Bha11]. Readers [Kni15, Sec14]. Reading [And18, BZ11b, KK14, Van11c, Chr13, Wil14]. Readings [Gou12c, PDG11]. Reaktion [Bec18b, Bel12, Cha14, Far15, Fla19, Han17, WM18a]. Realism [Boa15]. Reality [Ken13, Smi11b, Bea11, Moo10]. reappraisal [Rus14a]. Reason [Rei15, vL16, EKD^{+}13, Hal14d, Kel12]. reasoning [Mah18]. Reassessing [Fot18, Gro17]. reassessment [CC11b]. Réaumur [ES16, Ter14]. Rebecca [Rei15, Ebe17, Mae12, Wai18]. Rebekah [Köb16]. Rebel [Prk18]. Rebels [McG15]. rebuild [Ver19]. Recalcitrance [Can19a]. Recalcitrant [Lan15a, Tim13]. received [Ano10i, Ano10j, Ano10k, Ano10l, Ano11i, Ano11j, Ano11k, Ano11l, Ano12i, Ano12j, Ano12k, Ano12l, Ano13i, Ano13j, Ano13k, Ano13l, Ano14k, Ano14l, Ano14m, Ano14n]. Reception [Mos16a, Fis10]. Rechenberg [Hen11b]. Recherche [Sim13, BFGG11]. Recombinant [San16, Yi15]. reconceived [Har10b]. Reconciling [Moo13, Gue11]. Reconstructing [DiM13, O’C17, Lee15b]. Record [She13, Nas10, Sep12]. Recording [Tur15, Sch13]. recovering [Ech11]. recreating [OGE17]. Recycling [Wer13]. Red [Nal12, Lan11b, Mar16]. redefinition [Lip14]. Rediscovery [Pum12, Gou10]. Redman [Alb16]. Rees [Mon10]. reference [Fis10]. Refined [Doh16]. Reflections [Rad17, Ser14b, A^{+}12]. Reflexive [Pri13c, Ric11a]. Reform [Coz16, Roo12b, Gam17, Gam18, Pod15, Pow12, Ful18]. Reformation [Met11]. Reformer [Car18]. Regal [Nor12]. Regeneration [Les13b, Fer12]. regionalism [Lee15a]. Regionalizing [Nay10, Alb11a]. Regt [Ste11b]. Regulating [Kir17, Sim14, GH13]. Regulation [Stu11, Bat10]. Reich [GG12]. Reidy [Jac15]. Reimagining [Stu17]. Reinventing [Mor15a]. Reisch [Gar11]. Related [Dan19, Rob10b, Leh07]. Relations [SS15, Lor13]. relationship [Dho15, Ver16]. Relativity [Nav15b, Nic19, Waz14]. Religieuses [Boa10]. Religion [Bec18a, Can11, Can19b, Lin17, Mos16b, Mur12, Sma13, Tho19, Tys17, BJ10, Lar13, Mar14b, Yal11, Can16, Mac13]. Religious [Kot16, Sal14, Kum18, Wot12]. Relocating [GR18, Tam16a]. Reluctant [Smi12, Tau10]. remarks [GG10, Ver12]. Remmert [Kao12, Sia18]. Renaissance [Bec16, BT14, Car18, Edw14, Joh11, LN14, McG13, Mos16a, Pum12, QM13, Rap12, Roo12a, Row10, Sir13, Tim12, Azz12, Car18, Edg09, Gib14, Gou10, Hir11, Kle13, Mai12b, Mar09, Mar11, Van14, vL15]. Renato [Cap14]. Renn [Tam16a]. Renwick [Joh13a]. replacement [ANPR07, Goo10a]. Replication [Leg13]. Representation [CVLW14, Wel17, WM18b, Lo15]. Reprint [BW13, vL13a]. Reproduction [McL12, Wil11b, Kir17, OGE17, Mai12a]. reputation [Mar19]. Reputations [Kne10, CMC08, Fis10]. Rereading [Sep12, She13]. Research [BN13, Cap14, Das18, Fot18, HJ13, McG19, Mic11, Sim13, Som11, BFGG11, Bri15, Dav19, Hil12, MR12, Mon13, Mül11, Nie10b, Noa14, TW10, Wil15b, Wol09, Sau15, SV17, Tur16b]. Reshaped [Far12, Fer11b]. Resistance [Coz16, SS15, Lor13, Pod15]. resonance [Man13c]. resource [Ada16]. respiratory [McG19]. Response [Fei19, Sta19b, Bil15b]. responses [Nav10]. Restless [Hen17]. Restoration [Dun17b, Hen13a, Kus15, WM11, Elm13, Hun13, Poo10]. Rethinking [Gil15b, Tur16a]. Rethy [Ser14b]. retinue [Pye12]. Retrospectives [Gue19, Sim19b]. revelation [Liv15]. reversed [Dav19]. Reviel [Cuo12]. Review [AN18, Ada10, Aga17a, Aga17b, Ako19, Alb16, Ale11, And12a, And12b, Arb17, Arn16, Art11, Ase17, Ayl17, Ayl18, Ayl19, Bad16, Bad19, Bak11, Bak14, Ban12, Bar17, Bec12, Bec16, Bec18a, Bec18b, Bec18c, BT14, Bel12, Bel17, Bel18b, BZ11a, Ben19a, Ber11a, Ber17, Ber18, BS16, Bha11, Big10, Big18, Bil17, Bix19, Bjö11, Bla17, Boa10, Boa12, Boa18, Bod18, Boo11, Boo12, Boo14, Bow16, Bow18, Bow10, Bow17, Bro11a, Buc11, Bud11, Bui19, Bur11, Bus18, Cal12, Can16, Can19b, Can19a, Car16b, Car11, Cha14, Cle16, Cle17, Cle18, Coo11, Cow10, Coz17, Cul11, Cun10b, Cuo12, Cur18, D’A18, Dan19, Dar11, Das18, Daw17, Daw10, Daw16a, Daw16b, Dih17, Doh16, Dra17, Dry10, Dub10, Duf18, Dun11, Dun12b, Dun16a, Dun16b, Dun17a, Dun17b, Ebe17]. Review [Ebe18, Egm12, Eic12a, Eic12b, Elt12, Ens12, Esp10, Eve18, Fal18, Far10a, Far10b, Far11, Far19, Feo18, Fer11a, Fin10b, Fin10c, Fin11, Fin17, Fix16, Fix19, Fla19, Fon12, For10, Fot18, Fra17, Fre10, Ful18, Ful10a, GP18, Gar11, Gas10, Gib14, Gif16, Gil16a, Gil12, Gil16b, Gil11, Gis19, Gli11, Goo10a, Goo11a, Goo12a, Goo18, Goo17, Goo10b, Goo12b, Gou12a, Gou17, Gou12b, Gou12c, Gou14, Gra11, GG12, Gul18, Hah16, Hal10, Hal14b, Hal17, Ham10, Ham16, Han17, Han16, Har10a, Hay18, Hay19b, Hem19, Hen10, Hen17, Hen11a, Hen11b, Hen19, Hig10, Hig11, Hig16, HA16, Hol16b, Hol10, HA19b, Hug10, Hul10, Hun18, Hun19, Jam10a, Jam11a, Jam11b, Jam17, Jam18, Jar10, Jar17, Jim19, Joh11, Jon10a, Jon10b, Kaa18, Kao12, Kee12, Kee17]. Review [Keh12, Kel17, Ken19, Kid11a, Kid11b, Kip16, Kne10, Kno14, Köb16, Kot16, Kum16, Kum18, Kus18, Lan11a, Lan12, Lan19, LN14, Law17, Lea16, Leh12a, LD18, Lin17, Loi16, Lou16a, Lou17, Lou19, Lyk16, Mac11b, Mac10, Mac11a, Mae12, Mai11, Mai12c, Man17, Mar16, Mar17a, Mar18a, Mar12a, Mar12b, Mar14a, Mar17b, Mar18b, May17, McC19, McG15, McO10, Mca16, Met11, Mey11, Mic11, Mil11a, Mil11b, Mil12, Mil17, Mil11c, Mil11d, Mis11, Mod17, Mon10, Moo17, Mor10a, Mor18, Mos16a, Mox16a, MY18, Mur12, Mye12, Nes10, Nic17a, Nic17b, Nic19, O’C17, Oos12, Osl10, Par11, Par12, Pay16, Pea18, Pea12, Pea19, Pet10a, Pic10a, Pin12, Pis16, Pis17, Pri19a, Pri19b, Prk17, Prk18, Pro11, Pum12, Pyl10, Qui12, Rad17, Rap11, Rap12, Ree17]. Review [Ree18, Ree19a, Ree10, Reg17, Ren19, Ric12, Ric11b, Ric18, Ric10, Ric16, Rie16, Rih12, Rob10a, Rob17, Rob10b, Roe17, Roo12a, Roo12b, Roo12c, Ros17, Ros16, Rot16, Row10, Rus11a, Sac16a, Sac19, Sal14, San15b, San17, SJ19, SS11, Sch17a, Sch17b, Sch19, Sea10, SS17, Ser12, Ser14a, Sha10, Sha11, Shi16, Sho14, Sho18, Sia18, Sim17a, Sim16, Sim17b, Siv10, Sle10, Sle11a, Sle12, Sma13, Smi10b, Smi10c, Smi11d, Smi11c, Smi11b, Smi12, Smi16b, Smi18a, Smi18b, Smi19, Sno10, Som11, Sør17, Sta11a, Sta11d, Sta12b, Sta19c, Ste10, Ste11b, Ste16, SM19, Sto12, Str11, Stu10, Stu17, Sum11, SJ11, Tam16a, Tar12, Tat17a, Tat17b, Tat18, Tau11, Tee11, Tes11, Tho19, Tho16b, Til10, Tim12, Tos11, Tot10, Tra16, Tun19, Tur10, Tur12b]. Review [Und12, Val11, Van11c, Vos11, Wad11, Wad19, Wag14, Wai18, Wal18, Wal19a, Wal11a, Wal17, Wal11b, War19, Web11, Web17, Web18a, Web18b, Wes12b, Wig11, Wil11a, Wil11b, Wil16a, Wil16b, Win11, Wit11a, Wit11b, Wol11, WM11, WM18a, Wri16, Wyl11, Yal17, YM10, Zai16, vN19, Fei19, Sta19b]. reviewed [Ano10m, Ano11m, Ano12m, Ano13m, Ano14o, Ano15i]. Reviews [Ade10, Adl16, Aga13a, Aga14, Aga15, Alb11a, Alb15, And12d, And12c, Bak15b, Bak15a, Bal14b, Ban14a, Ban15, Bar13a, Bar13b, Bar15a, Bar15b, Bar16, Bar15c, Ber12, Ber15, Bil15a, Biv10, Bjö14, Boa15, Bod19, Boo17, Bou15, Bro16, Bud13a, Buk14, Bur13, Cal13, Can15, Cap14, Cap15, Car16a, Car18, Car12, Cha10, Chi12, Chr10, Cle14, Cle15, Coe10, Col13, Con15, Coo14, Coz15, Coz16, Cul13, Cuo13, D’A16, Daw12, Daw14, Daw15a, Daw15b, Daw15c, Daw16c, Die15, Dih16, Dij12, Dob14, Doh14, Dun12a, Dun14, Dun15a, Dun15b, Dys14, ES16, Edw14, Egm13, Eic15, Eva13, Eve13, Fal13a, Fan12, Far12, Far13, Far15, Far16, Fin10a, Gar14, Gil15a, Gli12, Gou13a, Gou13b, Gou16, Gre14b, Gri15, Hai16, Ham14a, Ham14b, Ham15, Har16, Hed13, Hei15, Hen14]. Reviews [Hen15a, Hen15b, Hen13a, Hen13b, Hig13, Hig14a, Hig15, Hoc14, Hoc16a, Hol13a, Hol14a, Hol15, Hol16a, Hol16c, Hom16, Hoq10, Huf16, Hus15, Ins19, Jac15, Jam12, Jam13, Jar16, Joh15, Joh13a, Joh13b, Joh14, Joh16, KK14, Ken13, Kir13, Kir15, Kni15, Kni16, Ksi13, Ksi15, Ksi16, Kus15, Lan15a, Lan14, Lan15b, Lau15, Lee15b, Les13b, Les13a, Les14, Les15, Lo15, Lou15, Lou16b, Mac13, Mai12a, Man13a, Man13b, Man16, Mar15, Mar12c, Mat13, Mil13, Mil15, Mil16, Moo13, Mor15b, Mor13, Mos16b, Mul13, Nal12, Nal15, Nas16, Nav14a, Nav15a, Nav15b, Nav16, Nor12, OP16, Opi14a, Pha10, Pic10b, Pin13, Pos14, Pos15a, Pri12, Pri13c, Rad13, Rad14, Raj15, Ree14, Ree16a, Ree16b, Ree15, Rei15, Ren10, Ren15, Ric14, Rob14, Rob15]. Reviews [Rod15a, Rod15b, Roo15, Røs15b, Røs15a, Row14, Rud13, Sac16b, Sal15, San12, San13, San14, San15a, San16, Sar13, Sar14, Sau13, Sau15, Sch15a, Sch15b, Sch16, SS12, SS13a, SS15, SS16, Ser14b, Ser15, She13, Sim14, Sim13, Sin15, Smi14, Smi16a, Smo10, Söd10, Som15, Sta11c, Sta14, Sta15, Sta16a, Ste14b, Sto14, Stu11, Tam14a, Tam16b, Tar15, Tat16, Tay14, Tho16a, TS14, Tur16a, Tur14, Tur15, Tur16b, Tys17, Und13, Val14, Van11b, Van14, Ver15a, Ver16, Vol16, Wag15, Wal16a, Wal12c, Wal16b, Wal15, War13, Web10, Web13, Web14, Web15, Wes15, Whi15, Whi16, Wil12, Wil14, Win15, Win13, Wol14, dC14, vL13a, vL13b, vL15, vL16]. Revision [Hei15, Doi13]. revisited [CVLW14, Lo15, Mar17a]. Reviving [Hud14, D’A16]. Revolution [BZ11a, Boa12, Kum18, Lip14, Mar16, Mor15b, Pos15a, Tar15, Bur16, Dan14, Huf11, Jon14b, Mai12b, McE10, Rem11, TB13, BT14, Ben19a, Boo17, Bud11, Kao12, L^{+}10, Tat16, Tom12a, Tom13a]. Revolutionaries [And12c, Med11]. Revolutionary [Mai12c, Tam14a, Per12]. Revolutionizing [Alb15, RC14]. Revolutions [Fin17, Rad17]. Rey [Ver15a]. Rhees [Gra11]. Rheinberger [Ayl18, Cap14, Mey11, Rad14]. Rhesus [Ban14b]. rhetoric [Tam14b, Toa11, Toa12, Ver10]. Rhetorical [Nav15a, Wal13]. Rhode [Cha10]. Rhodes [Fin10a, LMSR09, oRLM^{+}09]. Rhodri [Edw14]. Rhys [Hun18]. RI [Bar16, Hol16a, Sør17]. Ricardo [Goo12b]. Riccioli [Pri19a]. Rich [Bar15c, EJ13]. Richard [Daw15b, Esp10, Fin10b, Hig15, Rad13, Köb16, Mox16b, Olo14, Smi11b, Stu17]. Richards [Bow17, Hoq10, Lan15b, Pri13c, Rad17, Web10]. Richardson [Eic15]. Richter [Coe13, Val14]. Riddle [Dys14]. Ries [Gou14]. Rights [Sin15, Ben13a]. Rina [Wal16b]. Ringer [Ren10]. Rise [Aga17b, Ale10, Hal14b, Hay19b, Huf16, Lan11a, Par11, Web15, Wil16b, Coh15, Ras14, Wea12]. Risk [Moh13, Wil16b, Kir13]. Riskin [Hen17]. Risks [Ros16]. Ritual [Wal12c, Wyn11]. Ritvo [Sle12]. Road [Bjö14, Hen15a, Mul13, Hal14c, Man13c, Mue11]. Rob [Mil15, Mil17]. Robert [Biv10, Hoq10, Hun18, Keh12, Mon12, Mox16a, Ren15, Ric14, Sch15a, Sim13, Str11, Stu11, Und13, Wal16b, Web10, Aga13b, Bat10, Bow17, Cha14, GF09, Lan15b, Law15b, Loi16, Mox16b, Osl10, Pea12, Pic10a, Pis17, Pum12, Qui16, Rad17, Row14, Sta16a, Ste11a, War13, WM18b]. Roberta [Ste16]. Roberto [Nic19]. Roberts [Egm12, Har16]. Robertson [Liv15]. Robots [Mai12a, Smi16a, McL12, Ver15b]. Robson [Jon10a, Pet10a]. Rochester [Han16, San15a, Smi18b, Sta14]. rod [HHR09, Gra11]. Rodger [Ree16a]. Rodopi [Eic15, Nes10]. Rodríguez [Ric11b]. Roger [Fin10b, Sch15b, Tys17, Big10, Kas14, Sal14]. Rogers [Fer11a, Bus10]. Roland [Dun12b, Jam18]. role [GG10, Hod11, Wat14]. Rolfe [Kus18]. Roman [Ade10, Hol16a, May10, O’L06, Tar13]. Romanes [Smi11c, Sch10]. Romans [Rih12, Leh12b, Mur14]. Romantic [Bil15a, Esp13, Ksi15, Byr13, Pri13a, Ham15]. Romanticism [Gig09, Hol10]. Romantics [Jim19]. Ronald [Bec12, Can19b, Bec18c, Rob17]. Rood [Fin17]. Rooms [Alb16]. Roos [Fon12]. Roosevelt [HA19a]. Roots [Coz15, Jar17, Joh13a, OA14, Ren12]. Ropodi [Som11]. Roselló [Kus18]. Roselló-Izquierdo [Kus18]. Roses [Gil15a, Pri11]. Rosner [Stu11]. Ross [And18]. Rossby [Ric16]. Rosse [Kir15, Mol14, Nas10]. Rossi [Fin10a]. Rotblat [Und12, Bro12]. Rothman [Sch19]. Rothmann [Car16a]. Rothschild [Dav19]. round [RSA^{+}17]. Roussanova [GG12]. Rousseau [dC14, Coo12]. Routes [Ber15, Cap14, HD13, MR12]. Routledge [BS16, Big18, Bod18, Cle17, D’A16, D’A18, Daw16c, Dun17b, Feo18, Fot18, Ful18, Kum18, Pri13c, Roe17, Wal17, Wil16a]. Rover [Smi16a, Ver15b]. Rowlinson [Win13]. Roy [Bro11a]. Royal [And12d, Bro11a, Dun17b, Elt12, Gas10, Kus18, Hig19, Mac15, Mox19, Spi15, dC09].

Rublack [Reg17]. Rüdiger [Goo18, TS14]. Rudolf [MR13]. Rudwick [Ksi16]. Ruha [Sin15]. Rule [Wal11b, GJ09a, GJ09b]. Ruling [Tho16b]. Runaway [Hay19b]. Rupecissa [Sea10]. Rupert [Fie10]. Rupescissa [DeV09]. Ruprecht [Web13]. Ruse [Bow17, Hoq10, Ric10, San15b]. Rush [Wyl11, Bri10]. Ruskin [Gil16b]. Russel [Ben13b, Fla19, Les13a, Les14, Les15, SDS13]. Russell [And12d, Far15, Mil11d, Rad13]. Russia [Mai12c, HA16, Kic09, Nes10, RRL12]. Russian [Roe17]. Russland [GG12, RRL12]. Rutgers [Whi16]. Ruth [Stu10, Bur11, McG15, Pic10b, Wal19a]. Ruthless [Win13, Row12]. Rutter [Fin10b]. Ryan [Lou19].

S [Bec18c, Chi12, Col13, Hal10, Har10a, Hul10, Jac15, Ksi16, Nav14a, Nav15b, Pic10b, Prk18, Sal14, Smi11c, Smi19, Tar15]. Saberes [Car16a, Nav14b]. Sabina [Ste11b]. Sabine [Bar17]. Sabrina [Cle18]. Sachiko [Gar11, Egm13]. Sachse [SJ11]. Sacred [McO10, Des09, Sma13]. Sadiah [SS12]. Sadowsky [Wal17]. Safeguarding [Joh12]. Safety [Kir13, Ros16, Moh13]. Sagamore [Wil11a]. Sagan [Mar19]. Sakamoto [Car18]. Sake [AN18]. Salisbury [Fin11]. Sally [Pro11, Sle11a, Wil12]. salt [Vog10]. Salutary [dC14, Coo12]. Salvador [Fin10b]. Salvarani [Big18]. Samuel [Sch15b, Alb16, Gil15b, Kas14, Mac10, Row10, Tur16a, Wad11]. Sancho [San13]. Sand [Mar17a]. Sander [Web10]. Sanders [Pea12]. sanitation [Bux18]. Sanja [Tam14a]. Sara [Dun17a]. Sarah [Bel17, Feo18, OP16, Rod15a, Rod15b, Tur16b]. Saraiva [Wad19]. Sarpi [Nay14]. Sarrocchi [Duf18]. Sasquatch [Nor12]. Satellite [Bak14, ML^{+}12]. Satire [Far13, Lyn12]. Saunders [Lou16a]. Savage [Bar14]. Savannah [Fer11a, Bus10]. Savant [Bak15a, War13, Fox12a, HSSR13]. savants [Boa10, Vol08, Vol08]. Savile [Pum12, Gou10]. Saving [BW14]. Sayer [Sta19c]. Scaliger [Car18]. sceptic [Feo19]. sceptre [Opi14b]. Schäfer [Wit11a]. Schaffer [Ree10]. Schattengefechte [Ech12, Web14]. Schechner [Dun17a]. scheme [Bax13]. Schenk [Ako19]. Schindler [Gli12, Š^{+}11]. Schmalzer [Goo10b, Mar16]. Schmidt [Man16, Tur15]. Schneider [Hai16, Sia18]. Scholars [Gas10, Gou14, HSSR13, Kee12, Mar12b, NHR12, Bak15a]. Scholarship [Sma13, Coh18, Fei19, Sta19b]. scholastic [Lev16]. School [Edd08, Hol16a, Rob10a, Dud14, Ver19, Zal14]. Schwartz [Pic10b, Smi11c]. Schweber [Nav14a]. Schwert [BTW09, Wal11a]. Science [Ada10, Ade10, Aga17a, Aga13b, Alb11a, Alb15, Arb17, Bak11, Bak15b, Ban14a, Bec18a, Ber12, Ber17, BS16, Bha11, BG14, Bix19, B^{+}09, Boa12, Boo11, Bor09, Bow09, Bro11a, Bur13, Can16, Can19b, Cap15, Cha09, Cle14, Cul11, Cul13, Daw14, Dob14, Dra17, Dun11, Dun12b, Dun17b, Ebe18, Ell10, Ens12, Far10b, Far13, Far19, For10, Ful18, Ful10b, Goo10b, Goo12b, Gou12a, Gou12b, Hay18, Hen19, Hig10, Hig16, Hol13a, Hol14a, Huf16, Ins19, Jam10a, Jam12, Jam18, Kaa18, Kee12, Keh12, Kid11b, Kni15, Kni16, KKNN08, KMB18, Ksi15, Kus15, LaF13, Lan11a, Lan19, Lor13, Lou19, Mac11b, Man17, Man16, Mar12a, Mar12b, Mat13, Mil13, Mil17, Mis11, Mor10a, Mor10b, Mye12, Nas16, Nav16, O’L12, OK14, Pis16, Pro11, Rad17, Rec10, Ree10, Ric12, Ric14]. Science [Ric16, Rie16, Rih12, Rob17, Røs15a, Rot16, Roy12, Sac19, Sal15, San17, SS11, Ser14a, Sim17a, Sin15, Sin08, Siv10, Sle11a, Söd10, Som11, Sta11c, Ste10, Stu11, Sum11, Sys18, TV11, Tam16b, Tam16a, Tar12, Tay14, Ter13, Tes11, Til10, Tos11, Tra16, Tys17, Val14, Vol08, Wai18, War13, Web14, Web15, Whi15, Whi16, Wig11, Wil11a, Wri16, Yal11, Ade09, And13b, AM11, ABS10, Ayr08, Ban11, Bat10, Ben13a, Ben11, BCP12, Bia12, Boo15b, Bor08, Bro10a, BS15, BJ10, Byr13, Car10, Chi08, Cla09, Coc13, Coe13, Coh15, Com12, Coo12, Cre13, Cre10, Dal10, DeW13, Der12, Dev16, Don14, Dri10, End16, FO11, FW15, Fox12a, Gib15, GR18, Gue11, HNS11, HW11, Hes11, HA08, Hun09, Hun13, Jac13, Jon09, Jon14a, Kir11]. science [Kni09, Koh10, KB09, Leg13, Leh12b, LR11, LW11, Löw11, Lyn12, Mac15, Mac09, Man15, Man12, MHO13, Mar14b, McE10, Mer19, Mon13, Mou17, Mur14, Mye17, Nay10, O’B13, O’L06, OG16, Opi14b, Ort09, PNGP09, Pes12, Pet13, Pha13, Phi12, Por09, Pri10, RC14, Raj10, Ras14, Rav11, Ree09, RC13, RSA^{+}17, Ros10, Sch08a, Sch08b, Sch15c, Sec14, Sel08, Sem12, Shu10, Sim19b, Sle11b, SC12, Sol13, Sta12a, Sum13, Szu18, Tam14b, THB12, Tur14, Tur08a, VM18, Wal12b, Wer13, Wig10, Wil11c, Wit10, Wol13, Wol09, Yal15, Yeo14, Ade10, Aga15, And12a, Ayl17, Bec12, Boa10, Boa18, Boo11, Boo14, Bou15, Cal12, Coz17, Dan19, Daw16a, Dun11, Hal10, Hem19, HA19b, Hun18, Kee12, Kee17, Kum18, Lin17, Lyk16, Mai11, Mis11, Mon10]. Science [Moo13, Mor10a, Mor10b, Mos16b, Mur12, Nic17a, Nie14, Pha10, Pis17, Prk18, Ree19a, Sch17a, SS15, Smi11d, Smo10, Sno10, Tau11, Tho19, Toa11, Toa12, Wal19a, Wit10, Wol14, WM18a, Yal17, dC14, Roo15]. Sciences [Boa10, Buc11, Cle16, Eic12b, Gli12, Pin12, Ste14b, Whi16, AD12, CC14, JA15, Lov10, LOvT12, Mur10, Nor10, RC13, Smi11a, Spa12, Ver15a, VB11, Wer10, Ayl19, Buk14, Gil12, Rud13, Smi18b, Bus18]. Scientific [AKM14, Alb16, BT14, BZ11a, Ben19a, Bla17, Dew10, Dun17a, Ell10, Fin17, Gil12, Gri15, Hai16, Hal17, Hen13b, Hig14a, Hoc14, Jac15, Kao12, Kla09, Kne10, Les13a, Les14, Lo15, Mar16, Mor18, Pos15a, Rad17, Roe17, Sac16b, Smi10a, Sta11a, Sta15, Ste11b, Tar15, Tay14, Tra16, Wal16a, Web17, Wig11, dRLE09, And18, AS12, Bal15, Ban11, Bel18a, BDS15, Bur16, CMC08, Col14, CVLW14, Csi17, Fan10, Gra13, Hic19, Huf11, Lip14, Mai12b, Mis13, Mor17, O’B13, Olo14, Rad16, Rat16, Ray15, Rec10, Rec13, Rem11, Ren11, Roo17, Sta19a, TB13, Tur13, Vol08, Wal19b, Wat14, Hig15, Mak13, Mil11c]. Scientist [Cha09, Lau12, Sum11, VPW15, Tho16a]. Scientists [Ber15, Bri15, Gou12b, Ham10, Hay18, NHR12, Rob14, Wal13, WM11, And08, Ech12, HD13, Kir11, McC12, Poo10, Pye12, RRL12, Gou14, Nav15a, Tur16b]. Scienze [BCP12, Col13]. scope [Hay19a]. Scot [Qui16]. Scotland [Elt12, Alb11a, Fin09]. Scott [Coz16, Hol14a]. Scottish [Elt12, Mil11d, Sim17b, Hic19, ML10b]. Scouler [Sim17b]. Scouts [Ver19]. screen [AG17]. screening [Cow08, Pic10b]. screw [Spi15]. Scull [Biv10]. Sculpture [Jar16]. Scurvy [SS17]. sea [Ben19b, Nie10b]. Sean [Pri12, Cle14]. Search [Coz15, Mil13, WM11, Jac13, OA14, Poo10, RC11]. Searches [Wol14, Ter13]. Searching [Nor12]. Seattle [Sta16a]. Second [Bud11, Hoq10, Pha10, Siv10, Wyl11, Bri10, Vol08, L^{+}10, Sin08]. seconde [Boa10, Vol08]. Secord [Daw16b, Kni15, Rad13]. secrecy [Bia12, Ver12, VM12]. Secret [Lea16, Por09, Sch15c, Ada16, Man16, Tes11]. Secrets [Gre14b, LR11, Ren19, San14, Kel12, Pri13b, Sto13, Tar12]. Secularism [Rec13]. secularization [Toa11, Toa12]. Seeds [Hol16c, Web18a, Han15]. Seeing [Nal12, Ver15b, Lan11b, Spi15, Smi16a]. Seeking [Kas14, Sch15b]. Seen [Hig13, HA12]. Segonds [Chr10]. segreti [Cle18]. Sein [GG12, Fan10, RRL12, Sta11a]. Seized [Lip14]. Sekerák [Gli12]. selected [Sim10]. Selection [Sto12, FF11]. Selective [Hil12]. Self [Boa18, Doh16, Sys18, Ver10]. self-assessment [Sys18]. Self-Fashioning [Doh16]. self-presentation [Ver10]. Selin [Siv10]. Selling [Bak11, Wig10]. Selskab [Hun19]. selves [Lun13]. Semi [Gou12a]. semigroups [Hol14b, Bar16]. Seneca [BCP12, Col13]. sens [Del09]. Sense [Web10, Dyd15, Ric08]. Sensibility [Lan11a]. Sepkoski [She13]. September [Sta11b]. Sequence [Web15, Gai17, Ste13]. Sequencing [San13, GS12]. Sera [Ayl19, Daw15c]. Sera-Shriar [Ayl19, Daw15c]. Series [Cal12, YM10]. sermons [Toa11, Toa12]. Serrano [Ric11b]. set [Cuo12]. Seth [Dar11]. settlements [Wil15b]. Seuil [Man13a]. Sevan [Wol14]. Seven [See12]. Seventeenth [Bar13a, Bla17, Duf18, Eic12a, Bee19, Hig19, KT19, ZR15, Ber11b, War12]. Seventeenth-Century [Bla17, Duf18, Eic12a, Bee19, KT19, ZR15, Ber11b]. Sex [Mai12a, Szu18, Wag15, End16, Kin13, McL12]. Sexual [Lou19, Löw11]. Seysenegg [Gli12, Š^{+}11, vTSŠ^{+}11]. Shack [Mar17a]. Shadow [Vos11, Lev09, Ech12]. Shail [Fin11]. Shaky [Sol13, Whi16]. Shalev [Sma13]. Shank [Bec12]. Shape [Jim19]. Shaped [Bar15a, Fix16, Hol16c, HA19b, McG15, Ale14, Han15]. Shapin [Win19b]. Shaping [Ayr12, Coe10, GT09, Lan11a, Mod17, Pis16, Tur12b, Cou17, Goo11b, Wal19b, Wal09, Ham14b, Tee11]. Sharing [Mod17]. Sharon [Wal15]. Shaun [Joh11]. Shaw [Roe17]. Sheehan [Smi11d]. Shefer [Yal17]. Shefer-Mossensohn [Yal17]. Sherlock [Tay14, O’B13]. Sherry [Söd10, Söd10]. Shih [Ser14b]. Shih-Chang [Ser14b]. ships [Ech11]. Shipwreck [Cap13, Rob15]. Shock [San15a, CR14, Jar15]. Shocking [Sta15, Tur13]. Short [Bad16, Coo14, Kip16, Moo17, Mor13, Pic10b, Ser15, Sta11d, Wes12b, Cas11, Har08b, Jam10b, Jon10c, PB13]. Shorter [Rad13, Van09]. Show [O’C17]. Shrewsbury [Pat09, Van11b]. Shriar [Ayl19, Daw15c]. Shuttleworth [Sle11a]. Siberian [JS17]. Sibum [Hig10]. side [Ker14]. Siècle [Sta16a, Vol08, Boa10, JS08, Bra15, Chr10]. siècles [Sim13, BFGG11]. Siegfried [Dub10]. Sight [Eve18]. Sigrid [Goo10b, Mar16]. silence [Pos15b]. Silke [Hig15]. Silvan [Nav14a]. Silver [Fin10b]. Silvia [AN18]. Simões [Tam16a]. Simon [Alb11a, Ins19, Lau15, Qui12, Stu10, Buc11, McR14, Ree15, Zip13]. Simone [Cal13, Car16b, Har16]. Simple [Gil16b]. Simunek [Gli11, Gli12]. Sinanthropus [Man15]. since [DiM13, Hal14d, Lee15b, vL16]. Singapore [Ben19a, Gil12]. Singular [Gas10, dC09]. Sinha [Pha10]. Sir [Bil17, Doh14, Fin10b, Hul10, HWM12, Jac08, Row12, Win13]. Siraisi [Gib14]. SIRI [Les13a, Les14]. sistema [Pri19a]. Sixteenth [Egm13, Cla14, KT19, Tar13, JS08, Kus11]. sixteenth- [KT19]. Sixteenth-Century [Egm13, Cla14, Tar13, Kus11]. Skepticism [Van13]. skilful [Spi15]. skill [KT19]. Skills [Bak14, ML^{+}12]. Skinner [Bod19]. Skipnes [D’A18]. skulls [Mad16, Pet18]. Skuncke [Gil16a]. Sky [Smi14, Wes12a]. Slain [SS16, Lar14]. Slava [Hol16b]. Slavery [McO10, Des09]. Sleigh [Cal12]. slippery [Str19]. slips [Ros18]. Slipstream [Tys17]. Sloan [San12]. Sloane [Doh14, HWM12, Ros17]. Small [Sta11d, Jon10c]. smallpox [Few10]. Smile [Mor15b, Jon14b]. Smith [Bil17, Cle16, Eic12b, Eve18, Gou13b, Les13a, Lin17, Man13b, Mar15, Mil16, Rob15, SM19, Wil12, Cap13, Cro12, Liv15]. Snakes [Hal17]. Snezana [Kot16]. Soaking [MY18]. Sociabilities [Bod18]. Sociable [Sch17b]. Social [Ber17, Dry10, Joh13b, Jon10a, Mat13, Rad14, Whi16, Win19b, Ale08, Bun12, HA19a, Pet14, Ren11, Rob08, RC11, SC12, Sol13, Vic11]. social-scientific [Ren11]. Socialism [HSW09, SJ11, VW10]. Socialist [Mar16]. Sociedad [Car16a, Nav14b]. Société [Chr10].

Societies [Alb11a, Fin09, Leh07, Llo12, Rob10b]. Society [And12d, Bar16, Bil17, Bro11a, Daw15b, Dun17b, Elt12, Erl11, Fre10, Gas10, Gil15a, Gra11, Hol16a, HA19b, Keh12, Mye12, Rad14, Ree18, Ree10, Ric10, Roo12c, Sim17b, Smi11c, Sør17, SJ11, Tam16b, Tho16a, dC09, Ban11, Moh13, Nav14b, Sch08a, VPW15, Wil11c, Yal15, Erl13, HMLE15, HSW09, Mac15, Mox19, Sim19a, Kir13]. Socio [Kum18]. Socio-religious [Kum18]. Sociological [Asp11, Ham14a, TlT12]. Sociology [Joh13a, Ren11, Ren12]. Sofia [Dun14]. Sofie [Tra16]. softcover [Nic19]. Solà [RC11]. Solander [Ros18]. Soler [Ayl17]. Solid [Hen19]. Solovey [Mat13, Whi16]. Solvay [BCH^{+}13, Ree14]. Solving [Pea18]. some [Fis10, GG10]. Something [Bar15c, EJ13]. Sommer [Man17, vL13a]. Sommerfeld [Dar11, Set10]. Sons [Fin10c, Jay09]. Sophie [Bud11]. Sørensen [Sia18]. Soul [Hir11, Pin12, VB11, Roo12a]. Souls [Ham10, And08]. Sound [Boo12, Boo17, Tur15, Boo15a, Nie10a, Sch13, Ebe18, Jam18]. Sounding [Boo15a, Pos15b]. South [Bro11a, Dub10, Lee15b, Abr09, DiM13, GR18, Hui09, Pha10]. Southampton [Sal14]. Southern [Bro11a, Smi14, Mah18, Mac09, Wes12a]. Soviet [Tat17a, Hol16b, JS17]. Soviets [Hol13a, Wol13]. Space [Aga13a, Han16, Hol16b, Rob14, Rus11a, Smi19, Gep12, Kel10, KCM13, McC12, Aga14]. Spaces [Bil15a, Gal12, Mah18, Pri13a, Bod18, Goo17]. Spain [Coz17, Few10, Hoc16b]. Spalding [Mar12b]. Spanish [Ben19a, Sim17a, Smi11d, Tes11, B^{+}09, Nav14b, Por09]. Spark [Tur13, Sta15]. Sparks [Ell17]. Spary [Ste14b]. speak [Hes11, Ric12]. speaking [Ban14b]. Special [Boo12, Nie10a]. Species [Les15, Pea18, Hoq10, Dil11, RR09, Dil12, Dil13, Jin19, WC13]. Specimens [Ros18]. Spectacular [Stu11, Ros10]. speech [Sha12]. speed [Bal14a]. Spencer [Daw16c, FT15, Hal14b]. Spencerism [Smi16b]. sphere [Car10, Sta11c]. Spillmann [Pin13]. spin [Spi11, Spi12b]. spin-off [Spi11, Spi12b]. Spinning [Fan10, Sta11a]. Spiral [Nas10]. Spirit [Gou13b, HA19b, SFFC12]. Spirits [Les13b, Fer12]. Spiritualist [Les13b, Fer12]. Sport [Gar14, All12]. Sports [And12b, Arn16, Heg11]. Sprache [Hen11b, Rec10]. Sprat [Dun17b]. Springer [Ako19, Boa15, Ens12, Gul18, Hen11b, Lyk16, Nic19, Oos12, Pum12, Rap11, Ric14, Sia18, Siv10, Sum11, Tam16a]. Springfield [Mil17]. Squaring [Fix19]. St [Kao12]. Stability [GS08, YM10, Jac13, Mil13]. Staffan [Ayl18, Kee17, Rad14, Rad14, San17]. Stage [Tra16]. Stahl [Fin10a]. Stalnaker [Dun12a]. Stanback [GP18]. Standardization [YM10, MR13, VM18, GS08]. Standardizing [Qui12, GS10]. Standards [GS08, YM10]. Stanford [Boo17, Far19, Lee15b, Mey11, San16, Ser12, Sin15, Smi11d, SM19, TS14, Yi15]. Stanley [Gou13b]. star [Mar19, Pum11]. Stargazers [McA13]. Stars [LN14, Azz12]. State [Ban14a, Hen19, Hol13a, Pha10, Pis17, Tam16b, Tho16a, War13, Abr09, Fox12a, Kel12, Lau12, Pha13, Rat16, VPW15, Wol13, Yal15]. Statement [GR18]. States [Jon10b, Smi11b, VM12, Vol16, Ban15, Bro09, Eps14, Kel17, Moo10, Nov12, Wai18, Wil16a]. statesman [Bra12, Bud13a]. Statistical [Wil16b]. Statistico [Tat18]. Statistico-Mechanical [Tat18]. Steam [Fer11a, Fyf12, Ins19, Mar12a, Mil08, Bus10, Daw12]. Steam-powered [Fyf12, Daw12]. Steamship [Fer11a, Bus10]. Stedall [Pet10a]. Steel [Keh12]. Steen [Ree16b]. Stefaan [Bro16]. Stefan [Hen11a]. Steiner [Dun12b, Gli11, Mic11, Ser14b, Wal11a]. Steinke [Bak15a]. Steinle [Fal18]. Stem [Cap14, Sin15, Ben13a, MR12]. Stepchildren [Som11, Wol09]. Stephen [Bak14, Bar13b, Lin17, Wal11b, Lan11a, Mia12, Nav16, Pis16, Smi18b]. stepping [Par18]. Steve [Lo15, Tos11]. Steven [Nal15, Win19b]. Stevens [Ree18, Web15]. Stewart [Eve13, Ful18, Pos14, Ren19]. Sticks [Arn16]. Stock [Coh18]. Stockholm [Bjö14, Man13c]. Stokes [Ree17]. Stolleis [Hen10]. Stolzenberg [Gre14b]. Stone [Cha09, Sum11, Par18]. stones [Ree16d]. Stories [Ree16d, Wil11b, Kir17, Mar17a]. storm [Mah18]. Story [Bjö14, Mil15, Sac19, Man13c]. Straits [Wil15b]. Strange [Pea19, Rei15, EJ13, EKD^{+}13, Bar15c]. Strano [Bak14]. strategies [Bal14a]. Straus [Fix16]. Stress [CR14, Mil13, Jac13, San15a]. strife [Jon16a]. Strongly [Sta15, Tur13]. Stroud [Van11b]. Structure [Rad17]. Struggle [Wal15, Web10, Par18, Ric08]. Stuart [Law17, Fan12]. Stuber [Bak15a]. Studholme [Ric18]. Studies [Bar13a, Edw14, Fin10a, Gli12, Gou14, Mox16a, Rad13, Rad14, Ric10, Tim12, Boo15a, CRB09, Llo12, NHR12, Pes12, War12, Bur13]. Studio [Tur15, Sch13]. Study [Car18, Gil16a, Mis11, Pro11, Sto12, Tim12, Web10, FF11, Gli08, Koh10, LG11, Wit10]. studying [Hod11]. Sturm [Rei15]. Stuttgart [Dun12b, Gli11, Mic11, Ser14b, Wal11a]. Style [Mor10a, Bor09]. Styles [Hai16, BDS15]. subject [Mia12, Bar13b]. Sublime [Gil11, Lyn10]. subsistence [Fra16]. Substance [Hen19, Pyl10, Gar09a, Gar09b]. subtleties [Mor17]. Subverting [Mar14b, Mos16b]. Succeeded [Cha09, Sum11]. Sue [Gil15a]. Suffrage [Bix19]. sujet [Chr10, JS08]. Sul [Cle18]. Suman [Dar11]. summer [Ver19]. Sumner [Bak15b, YM10]. Sun [Hig13, HA12]. Sundials [Eag10, Sha11]. sunken [Ech11]. Supernatural [Osl10, Bar08]. supplied [KT19]. Support [Mic11, TW10]. Surgeon [Ken19]. Surgeons [ANPR07, Goo10a]. surgery [Fra16]. Surman [Hoc14]. Surveillance [Han16, Har16, Ada16, TR14]. Survey [CG12, Fal13a]. survival [Yam15]. Susan [Tur15, Bod18, Bod19, Fra17, Sta11d]. Susanna [HA19b]. Susannah [Kni16]. Susanne [SJ11]. Sussex [Man13b, Stu11]. Sustainable [Mar18a]. Sven [Bak14]. Sweat [Sho14, HKZ12]. Swedish [Gil16a]. Swift [Far13, Lyn12]. Swimming [All12, Gar14]. Swims [Tys17]. Sydney [Bro11a]. Sympathy [Mil17]. synthesis [Cla14]. Synthetic [Ban11, Ree16b, Ste14a]. Syria [Mil11d]. System [vL13b, JS08, Rab17]. systematics [Lee15a]. système [Chr10, JS08]. Systems [Ber13, Fin11, Tur38, App12, Ros18, SS10].

T [Aga13a, Dun11, Hen15a, Les15, Smi18b]. table [RSA^{+}17]. Tables [CG12, Fal13a]. Tacitus [Kel12]. Takashi [Hoc16a]. Taking [Roo17]. Talairach [Kee12]. Talairach-Vielmas [Kee12]. Talas [Dun14]. Talbot [Die15, BDR13]. Talent [Wol14, Ter13]. Tales [Wol11, Gro09]. Taming [KP14, Hol15]. Tania [MY18]. Tanja [vL15]. Tansley [Ayr12, Ham14b]. Tara [Prk18]. Tartaglia [Ech11, Ekh10]. Tatiana [Opi14a]. Tattersall [Biv10]. Tauber [Smi12]. Tauris [Mac11b, Moo13, Wig11]. Taylor [Daw16c, CC11b, Olo14]. Tchernetska [Cuo12]. Teaching [Koh10, Sun13, Bux18, HW11, Dun12b, Pro11]. Teams [Smi16a, Ver15b]. Tears [Sho14, HKZ12]. Technical [Hai16, BDS15]. Technique [Mye12, Wil11c]. techniques [RC13, Ter10, Ver15a]. Technische [BDS15]. technocracy [Ban11]. Technological [Cur18, Mod17]. Technologies [GS08, Arn16, YM10]. Technologists [Tat17b]. Technology [Aga15, And12c, Bak15b, Ber18, Boo17, Cow10, Far10b, For10, Goo11a, Hol13a, Lan14, Mil11a, Nas16, Rob17, Sim16, Siv10, Tur15, YM10, Goo08, Hun10, Ink08, Kni09, Med11, OK14, PNGP09, RC13, Sch13, Sel08, Spi11, Spi12b, Sta16b, Sum13, Wis12, Wol13, Bel17, Doh16]. Technopolitics [Sar13, Hec11]. Technoscientific [Wad19, Hil12]. Ted [Hen14, Les14]. Telegraph [Noa14, Ker14]. Telegraphy [Car16b]. Telescope [Dun15b, GMML14, Row10, Edg09]. Telescopes [Bak14, McA13, ML^{+}12]. televising [Boo15b]. Television [Boo14, Boo15b, LaF13, OG17]. tell [McG19]. Tempe [Edw14, Tim12]. terapeutiche [Cle18]. Termessos [vL13a]. Terms [Chi12, Mul10]. Terrall [ES16, War19]. territorialization [Gal12]. Territories [Can16]. Terry [Fin10b]. Terzian [Wol14]. terzo [Pri19a]. Tesfaye [Ham16, Pay16]. Tessicini [Tar15]. Test [Mai12a, ML10a, McL12]. Test-Tube [Mai12a, McL12]. testing [OG17]. Texas [Yal17]. Text [Egm13, Gul18, Kus11, Pum11, RSA^{+}17]. textbooks [BN13, Sau15]. textes [Chr10]. Texts [Ban12, Rob10b, Leh07]. Th [Wal11a, Vol08]. Thames [Mar17b, Ric18]. Their [Bow10, Kot16, Stu11, Ano17g, Col16, Lig09, Ros10]. Them [Stu11, Ros10]. Theodore [Tam16a]. theologie [Vol08, Boa10]. Theology [Far13, Hen10, DS08, Dil11, Dil12, Dil13, Lyn12, Vol08]. Theoretical [Hun19, Jar17, Ver19]. Theories [Nav16, BS15]. Theory [Bar15a, Bar16, Dar11, Fix16, Nav15b, Ale14, Cla14, Ell17, Fri11, Gro17, Hol14b, Nay14, Pri15, Ric13a, Rus14a, Set10, Swe15, Vic11, Gou16, Lan15b, Waz14].

Therapeutic [Qui12, GS10]. therapeutics [Pod15, Coz16]. Therapy [MY18, Wal17]. There [Bar15b, Hac14]. Theresa [Vos11]. thesi [Yal11]. thesis [App12, vL13b]. Thiel [Wal11a]. Thin [OG17]. Things [Ban12, Hen17, Söd10, KW10, ML10a, Nie14, Tur07]. Think [Söd10, Leh19, Tur07]. Thinker [Ben13b, Les14]. Thinking [Fan12, McW11, Coo14]. third [Fis10]. Thomas [Dun17b, Ful18, Hig11, Rei15, Røs15a, Boo17, Chi12, Fin10b, Fox12b, Pum11, Row14]. Thomson [Nav10]. Thor [Hol16c]. thorn [Ker14]. Thought [Jam12, Mox16a, Roe17, Ben11, Doh17, Ric08, Ell17, San15b, Web10]. Thousand [Hal10, KKNN08]. thousand-year [KKNN08]. Thrale [Gar14, All12]. threat [See12]. Threats [Cur18]. Three [San12, SF11]. Three-Man [San12, SF11]. Thrift [Tee11]. Thrifty [Ins19]. Thümmler [Gli12]. Thunberg [Gil16a]. Tiago [Wad19]. Tibet [Mue11, Mul13]. Tick [Hen17]. tidal [Nay14]. tie [Vic11]. Tiger [O’L12, Cul13]. Tiggelen [Buk14, Kee17]. Tim [Tot10]. Time [Ayl19, Joh16, Pea18, Sea10, Tam14a, Ano17g, BW13, Bul16, DeV09, Ogl15, Per12, Web08, Cun10b]. Timekeeping [Tee11, BW13, GT09]. Timeless [Gou12c, PDG11]. Times [Ben13b, Les14, AGF13, KW10, Pri13a, Ban12, Bil15a]. Timmermann [Ham14a, Lan15a]. Timothy [Dys14, SS16]. Tindley [Ber18]. Tinkerers [McG15]. Tissue [Mye12, Wil11c]. Tita [Far19]. Toby [BZ11a]. Today [Fin10b]. tolerance [Pes12]. Tom [Røs15a, Bil17, Ros16]. Tomas [Bui19]. Tomorrow [Fin10b, Hog12]. Tomory [Mil12]. Toms [Wes15]. Tony [Boa10]. Too [Ren19]. Tookey [VM18]. tools [Ter10]. Toon [Ham14a]. Toronto [Adl16, Ber12, Jam11b, War19]. Torpedo [Eps14, Vol16]. Torrens [Ful18]. Torriani [Ben19a]. Total [Goo10a, ANPR07]. toto} [Wel17]. Tour [Hig11, PdCB10, Roo17, SDS13, Les13a]. Tower [Sha12]. Toxic [FJB14, Kum16, Ree16a]. trace [Coh11, Man13a]. Trade [Lou15, Sim19a, Jon13, Sar14]. Tradition [Moo13, Van11c, Gue11]. traditions [Che12, Mor17, Cuo13]. traduction [Chr10]. tragedy [Bud08, Pic10a]. Tragic [Jam11a, Web10, Lam09, Ric08]. Train [Lau12]. Trans [Ren10, Ban12, Hig13, Leh12a]. Transactions [Boa10, Vol08]. Transatlantic [Bou15, Goo10a, Wri16, ANPR07, BG14]. Transcription [Fin10a, Les15, WC13]. Transcultural [Ako19]. Transdisciplinary [Prk18, McM18]. Transformation [Cha10, Joh16, Gab16, Gli08, Ogl15, Pau07, Mos16a, Web10]. transformative [Hod11]. Transformed [SM19]. Transforming [Jam18]. Transit [Hig13, HA12]. Transition [Ree19a]. Translated [Ayl18, Ban12, Big18, Fal18, Nav15b, Str11]. Translation [Car16a, Fin10a, Jin19, Sto12, Web10, AK16, FF11, Gat17, Gli08, Olo14, RCK10]. transmission [Mor15a]. transmutation [Jin19]. transnational [Kri12, Pes12, SM12, THB12, Wal12b]. Transparency [Vos11, Lev09]. Transporting [Doh17]. Travancore [Rat16]. Travel [Ksi15, Les13a, SDS13, Wit11b, Byr13, Dew10, GCA10, HJ13]. Travellers [Dub10, Hui09]. Travels [KWB15, Roo17, Lou16b]. Treasurer [Bil17]. Treasures [Ric18]. Treatise [Tim12, Lew12, Edw14, LG11]. treatment [Jon16b]. Trees [Mai12a, McL12]. Trevelyan [Kem19]. Trevor [Ful18]. Trial [Wag15, Kin13]. Trinity [Cou17]. triplicity [Leh19]. Trischler [Mic11]. Triumph [Hol16c, Pic10a, Bud08, Han15]. Trizio [Ayl17]. tropics [Cha12, Sta14]. troubled [San11]. True [Stu11, Ros10]. Trust [Fin10b, Kus18, KT19]. Truth [Joh13b, LCMJ08, Sch15b, Bun12, Kas14, Uek10, Stu10, Win19b]. Tschermak [Gli12, Š^{+}11, vTSŠ^{+}11]. Tschermak-Seysenegg [Gli12, Š^{+}11, vTSŠ^{+}11]. Tsun [D’A16, Hud14]. Tube [Mai12a, McL12]. Tucson [May17]. Turchetti [Cal13, Har16]. Turing [vL13b, App12, Nic17a]. Turkel [Sta15]. Turkle [Söd10]. Turn [Wyl11, Bri10, Col16, Spi15]. Turnhout [Boa10]. Turning [Bur11, Ham10, And08, Gla09]. Tuscany [Bec16]. Tutton [Stu17]. Twentieth [Aga13a, Alb15, Ban14a, BS16, Boo11, Can19a, Coz17, Goo10b, Hol15, Mic11, Mye12, Pos14, Rad14, San15a, Wyl11, Zai16, Ban14b, Bri10, CR14, End18b, Gep12, KP14, LMS13, Pet14, Pha13, RC14, RC11, TW10, Wil11c, Bow09, CC11a, GMWR13, Sch08b]. Twentieth-Century [Ban14a, Boo11, Coz17, Goo10b, Pha13, Bow09, CC11a, Sch08b]. Two [Jam10a, Kus18, Mil11d, Sim17b, Fis14, Mor17, Ort09]. Tyas [Joh11]. Tylor [Lar13]. Tyndall [Jac15]. Tyne [Gas10, Mar12b]. type [HA19a]. types [Van11a]. Tyranny [FO11, Kid11b].

U [Gou13b]. U.S [Alb15]. U.S. [RC14]. UCL [Fin10b, Hay18]. Uekötter [Web13]. Uhren [BW13, vL13a]. UK [Sto14, Spi11, Spi12b, Wal12a]. Ulf [Man16]. Ulinka [Reg17]. Ultimate [Tys17]. unborn [Ell17]. Uncertainty [Cow10, Goo08, Spi15]. Uncommon [MHO13, Sal15]. Unconventional [Gue19]. Understand [Rad13, Rus11b]. Understanding [Bad19, Loi16, Mar12a, Rie16, Ste11b, dRLE09, Llo12, Mil08, Eva13, Hug12]. understandings [OS13]. Uneasy [Tat17a]. UNESCO [GR18]. Unfinished [Dun12a, Sta10]. Unicopli [Cle18]. unification [Sta12a]. Uniformity [GS08, YM10]. Unipress [Hig16]. United [Ban15, Jon10b, Kel17, Smi11b, Vol16, Wai18, Wil16a, Bro09, Eps14, Moo10, Nov12]. Uniting [Rad13, Rus11b]. Universal [Cle17, Dun12a, Lov10, Lee15a, Wee19]. universalism [Lee15a]. Universe [Mar15, Win11, Edg09, GS13, Row10]. universitaires [Esp10]. Universitat [Car16a]. Universitetsforlag [Hig11]. University [Adl16, Aga17a, Aga17b, Alb15, Alb16, Ale11, And12a, And12b, Arb17, Arn16, Ayl17, Ayl18, Ayl19, Bad16, Bal14b, Ban15, Ban12, Bar13b, Bar15b, Bar17, Bar15c, Bar08, Bec12, Bec18a, Bec18c, Bel18b, BZ11a, Ber12, Ber15, Bha11, Big10, Biv10, Bix19, Bjö11, Bjö14, Bla17, Boa18, Bod19, Boo11, Boo12, Boo14, Boo17, Bou15, Bow16, Bow17, Bro11a, Bro16, Buc11, Bui19, Bur11, Cal13, Can15, Can16, Can19b, Can19a, Cap15, Car11, Car12, Cha10, Chi12, Cle14, Cle15, Cle16, Coe10, Con15, Coo14, Coz15, Coz16, Cul13, Cun10b, Cuo12, Cuo13, Cur18, Das18, Daw10, Daw12, Daw14, Daw15a, Daw15b, Daw16b, Dho15, Die15, Dih16, Dih17, Dij12, Dra17, Dry10, Dun12a, Dun16a, Dun16b, ES16, Ebe17, Ebe18, Egm13, Eic12a, Eic12b, Eva13, Eve18, Fal18]. University [Far10a, Far11, Far16, Far19, Fin10c, Fin17, Fra17, Ful10a, Gib14, Gif16, Gil16b, Goo12a, Goo18, Goo10b, Gou13b, Gou17, Gou12c, Gou14, Gre14b, Hah16, Hai16, Hal10, Hal14b, Hal17, Ham10, Han16, Har10a, Hay19b, Hem19, Hen14, Hen15a, Hen15b, Hen13a, Hen17, Hen13b, Hen19, Hig10, HA16, Hol13a, Hol14a, Hol15, Hol16b, Hol10, HA19b, Hoq10, Huf16, Hug10, Hun18, Hus15, Ins19, Jam10a, Jam11b, Jam12, Jam17, Jar16, Jar17, Joh15, Joh13b, Joh14, Joh16, Jon10a, Kao12, KK14, Kel17, Kip16, Kir13, Kir15, Kne10, Kni15, Kni16, Kot16, Ksi13, Ksi16, Kum16, Kus15, Lan11a, Lan12, Lan19, Lan15b, LN14, Lau15, Law17, Lee15b, Leh12a, Les13b, Les15, LD18, Lou16b, Lou17, Lou19, Mac13, Mac10, Mac11a, Mae12, Mai12a]. University [Man17, Man16, Mar16, Mar17a, Mar18a, Mar15, Mar12c, Mar18b, May17, McC19, Mca16, Met11, Mey11, Mil11a, Mil11b, Mil13, Mil15, Mil16, Mil17, Mil11c, Mil11d, Mis11, Mon10, Moo17, Mor15b, Mor13, Mos16b, Mul13, MY18, Mur12, Nal12, Nal15, Nas16, Nav14a, Nav15a, Nav15b, Nav16, O’C17, Opi14a, Osl10, Par11, Par12, Pea18, Pea12, Pea19, Pet10a, Pha10, Pic10a, Pic10b, Pin12, Pis16, Pis17, Pos15a, Pri12, Prk17, Pro11, Pyl10, Rad13, Rad14, Rad17, Rap12, Ree14, Ree16a, Ree16b, Ree17, Ree18, Ree19a, Reg17, Rei15, Ren19, Ren10, Ren15, Rie16, Rih12, Rob14, Rob15, Rob10b, Roo12b, Roo15, Rot16, Row10, Rus11a, Sac16a, Sac16b, Sac19, San12, San14, San15a, San15b, San16, Sau13, SS11, Sch15a, Sch19, Sea10, SS12, SS13a, SS15]. University [SS16, SS17, Ser12, Ser14a, Ser15, She13, Shi16, Sho18, Sim17a, Sin15, Sle10, Sle11a, Sle12, Smi10c, Smi11d, Smi12, Smi14, Smi16a, Smi18b, Smi19, Sno10, Som15, Sta11c, Sta11d, Sta12b, Sta14, Sta15, Sta16a, Ste10, Ste11b, Ste14b, Ste16, SM19, Str11, Stu10, Stu11, SJ11, Tam14a, Tam16b, Tat16, Tat18, Tau11, Tay14, Tee11, Tes11, TS14, Tho16b, Tot10, Tun19, Tur14, Tur15, Tur16b, Tys17, Und12, Val14, Val11, Van14, Ver16, Vol16, Vos11, Wad11, Wai18, Wal16a, Wal18, Wal19a, Wal15, Wal11b, War13, War19, Web10, Web14, Web15, Web17, Web18a, Web18b, Wes12b, Whi15, Whi16, Wil12, Wil14, Wil16b, Win15, Win11, Wyl11, Yal17, vN19, vL13b, vL16, Hut08, Yi15, HJ13, Ver16]. Unknown [Hol15, Gli12, KP14]. Unmaking [Goo12a, BJW11]. Unnatural [Fin10c, Jay09]. Unnaturalness [Tys17]. Unofficial [Zai16]. Unpublished [Edw14, Sto12, Fis10, FF11, Lew12]. Unrolling [Mos14]. Unseen [GS13, Mar15]. Unsuitable [End14]. Until [Bro10a, Goo12b]. Untimely [May14, Ren15]. Upcycled [Mar18a]. upon [Gas10, Mar12b]. Uppsala [Gil16a]. Uranium [Hug12, Sar14, Ada16, Kri12]. Urban [Hig14a, L^{+}10, Mis13, Bud11]. Ursumde [Chr10]. US$41.95 [Rob15]. use [Dil11, Dil12, Dil13]. uses [Lam11, Sha12]. using [Pet18]. Usselman [Jam17]. Uta [Fin10b]. Ute [Win15]. Utopian [End18a]. Uwe [Gli11, Gli12].

V [Goo10a, Hig16, Mai12c, San14]. V. [OS13]. Vaccination [Law17, Few10]. vaccine [BW14]. València [Car16a]. Valentine [Hen13a, Elm13]. Valier [Daw17]. Valleriani [Rap11]. Values [Ber17, Lan19]. Vandendriessche [Tho16a]. Vandenhoeck [Web13]. Vanessa [And12b, Joh16]. Vanhaesebrouck [Bui19]. Variety [Goo12b, Bro10a]. Vaupel [Hen11a]. Vegetable [Kni16, Gib15]. vein [Mak13]. Velikovsky [Tur14, Gor12]. Venereal [Bow18]. Venezia [Cle18]. Venomous [Hal17]. Venus [Hig13, HA12, Mar17b]. Vera [Hai16, Mar14a]. Verlag [Dun12b, Egm12, Gli11, Mar14a, Mic11, Ser14b, Sta11a, Wal11a, vL13a, vL15]. vernacular [Bil18]. Vernadskii [OS13]. Vernon [Dan19]. Veronika [Lou19]. Véronique [Mar12b]. version [Fis10]. Versita [Ree15]. Verso [Kid11a]. versus [Rab17, Ver12]. Vertebrate [Sun13]. Vertesi [Lo15, Smi16a]. Very [Aga13b, Kip16, Mor13, Ric14, Cou17, Jam10b]. Vesuvius [Cle14, Coc13]. veterinary [Hei18]. vexations [Mer12, Mer13]. vi [Bar13a, Bou15, Col13, Eva13, Gri15, Stu11, Tos11]. Vibrations [Gou13a, Ken12]. Vicky [Gil16b]. Víctor [Car16a, Hol15]. Victoria [Alb15]. Victorian [Alb11a, Bow10, Con15, Daw14, Daw15a, Dra17, Hal14a, Hig14a, Hun18, Kni15, Mil17, OP16, Pea19, Ric12, Wal18, Wal19a, Bar14, Car16b, Cou17, Cow13, DeW13, Dew14, Edd17, Fin09, FW15, Hes11, Lam11, Lig09, LZ14, Mac15, Mis13, Nay10, Opi14b, Sec14]. Victorians [Gil16b, Hoc16a, Cla09, Ito14, Sle10]. Vidal [Pin12]. Videnskabernes [Hun19]. Video [Fin10b]. Videogame [Lea16]. Vielmas [Kee12]. Vienna [Sva18]. View [Jim19, BG11, McA13]. vii [BT14, Biv10, Buc11, Dar11, Fin10c, Goo12a, Ken13, Man13b, Mar15, Mur12, Pet10a, San15a, SS12, Tam16a]. viii [Adl16, Ber12, Big18, Car18, Coz15, Cuo12, Far10a, Ful10a, Ham15, Hol13a, Hul10, Jam11b, Kir13, Kno14, Kum16, Mai12a, Man17, Mar12a, MY18, Nav14a, Pea12, Pea19, Rei15, Roo12b, Rud13, Rus11a, Sal14, Sch19, Sha10, She13, TS14, Val14, War13, Wit11a]. Vincenzo [Gat17]. Violani [Kus18]. Virginia [Sle12]. virtues [Pes12]. Virtuosi [Roo15, Yeo14]. Visible [Ble12, SS13a, SS13a]. Vision [Nic17a, Row10, vL16, Edg09, Hal14d]. Visioneers [Rob14, McC12]. Visions [Lan14, Sec14, Wis12, Kni15]. visit [End18b]. visiting [Nie14]. Visual [Kus15, MY18, Rud13, AD12, Ble12, Hun13, Ver10, D’A18, Nas16]. Visualizing [Bel18b]. Vitalism [Kno14, Pac12, Pet14]. vitalisme [Gis19]. Vitruvian [Ben19a]. Viviani [Gat17]. Vivisection [Hal17]. Vocabulary [Win15, FSS^{+}14]. Voice [Lou16a]. Vol [SK12, Bur13, Daw16b, Eve13, Ham16, Pay16]. Volker [Sim14, Kao12, Sia18]. Volpe [Boa10]. vols [Bak15a, Daw15b, Hen11b, Jam13, Siv10]. Voltaire [Dun12a, Hei15, Sha10, dC14]. Volume [Chr10, Fin10a, YM10, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14c, Ano14a, Ano14d, Ano14b, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano18b, Ano18a, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Cuo12]. Volumes [Kus18]. Vos [Smi11d]. Vox [Cur18]. Voyage [Keh12, Rad13, And18, CV09, Ste11a, Ric18]. Voyages [Sim17b]. Voyaging [War19]. Vrin [Hed13].

W [Aga17a, And12a, Boo17, Cha14, Cha10, Daw16c, Jam13, Lou16b, Mis11, Smi10b, Ste11b, Stu11, vL13b]. W. [Rob14, Swe15]. W.J [Sim16]. Wachelder [Ful18]. Wagner [Tys17]. Wahrheit [Uek10, Web13]. Wahrnehmung [Kle13, vL15]. Wake [Hay18]. Wales [Bro11a, Hun18, Tee11, GT09]. Walker [Doh14, Edd08, Mic11, Rob10a, Sto14, SJ11]. Walking [Wot12, Sal14]. Wallace [Les13a, Les14, Les15, Ben13b, Fla19, SDS13, WC13]. Waller [Mox16b]. Wallis [Ban12]. Walls [Coe10, Mar14a]. Wallstein [Hen11a, Web11, Web14]. Walsh [Nav15a]. Walter [Coe10, Mar14a, Kem19]. Walters [Cuo12]. Walther [Wal11a]. Wandering [JA15]. Wang [Ser14b]. want [Hic19, Wat14]. War [Cal13, Har16, Hol13a, Jon16a, Kel17, Lau15, Lin17, Mat13, McR14, Pha10, Ree16b, Rei15, Rie16, Rob17, Sol13, Sør17, Tur16b, Whi16, Wol13, Wri16, Bri15, JS08, OK14, Sin08, SC12, Ste14a, Ste16, Tur12a, Vic11, Aga15, Asp11, Bix19, Bri15, CC14, EKD^{+}13, Hec11, Nic19, Pha10, Sar13, Sin08, TR14]. Wardhaugh [Ale11, Bar13a]. Warfare [Can19b, Man16, Sch15c]. Warner [Bec18b]. Warren [Prk18]. Warrington [Fin10b]. Wars [Mac11a, Tur14, Gor12, Joh09, JA15]. wartime [Mou17]. Warwick [Ham10]. Was [BW13, Ksi16, Ric13a, vL13a, Rud14, Lan15b]. Washington [Sta16a, Wal12c]. watches [BW13]. Watching [Coc13, Cle14]. Water [Boa15, Opi14a, Hol13b]. Waterwheel [Dry10, Ale08]. Watkins [Dun15b, GMML14]. Watt [Far15, Ins19, Mar12a, Mil11a, Hun10, Mil08, Rus14b, Tom12a, Tom13a]. Waves [Nic17b]. Way [Kus18]. Ways [GH13, Sim14]. Wazeck [Nav15b]. weapon [Mag12]. Weapons [Lau15, Smi11b, Sto14, Tur16b, Bri15, McR14, Moo10, Wal12a]. Weart [Hal14b]. Weather [Har08a, Rob10b, Leh07, Mah18, Tur10]. Weathering [Wil15b]. Web [Roo11, Sta11a, Fan10, Fon12]. Weber [Ren10, Rad09, Rin04]. Webster [Cun10b]. Weindling [Fot18]. Wellcome [Fin10b]. Wellmann [Web11]. Wen [D’A16, Hud14]. Wen-Tsun [D’A16, Hud14]. Werdens [Web11]. Werner [Hen11b, Rec10]. Werrett [Buc11, Ins19]. Wertheimer [Pin13, Spi12a]. West [Dys14, Mul13, Mur14, RVG12, Mue11]. Western [Dun16a, Siv10, Sør17, Sel08, Dev16]. Westport [Ade10]. Westwick [Smi14]. Wexler [Ric16]. Whaling [Lüt19]. Whig [Mor10a, Bor09]. Whitaker [Hen15b]. whitemen [And08, Ham10]. Whiteside [GG10]. Whitmarsh [Tot10]. Whitmer [Web17]. Who [McG15, Ren19, Ros10, Stu11]. Whose [GS08, YM10]. Wicked [Hun13, Kus15]. Wigelsworth [Bak11]. Wijnands [Kus18]. Wilbur [Hig13]. Wiles [Bar08]. Wiley [Daw16a, Ham14b, Sch17a]. Wiley-Blackwell [Ham14b]. Wilhelm [HSW09, Mül11, SJ11]. Wilkins [Tot10]. Wille [Ayl18, Rad14, Rad14, San17]. William [Cuo12, Daw12, Gli12, Gou13b, Hen15a, Jam17, Mor18, Win11, Bor08, BDR13, Die15, Edw14, Fyf12, Gre14a, Hal14c, Hun18, Kip16, Kir15, Lew12, Liv15, Mol14, Pos15a, Rus14a, Š^{+}11, Sta15, Ste10, Wat14, WM11]. Willy [Smi19]. Wilmington [Dan19]. Wils [Tho16a]. Wilson [Cuo12, Mye12]. wilting [Mac16]. Winchell [Liv15]. Window [Row10, Edg09]. wings [Val10]. Winter [Bal14b]. Winterbottom [Ham16, Pay16]. Wireless [And09, Hug10]. Wires [Bec18b]. Wirken [Hen11b, Rec10]. wirkenden [GG12, RRL12]. Wirtén [Far16]. Wisnioski [Lan14]. Wissenschaft [BTW09, Hen11b, Rec10, Wal11a]. Wissenschaft-Wissenschaftler [Wal11a]. Wissenschaftler [BTW09, Wal11a]. Wissenschaftlern [GG12, RRL12]. wissenschaftliche [Hen11b, Rec10]. Wissensgeschichte [Uek10, Web13]. Witch [Reg17]. Withers [And12a, Lou16b, Mis11, Sim16]. Withey [Doh16]. Within [Man17, Smi10b, Har08c]. without [Pis17, San15b]. Wittje [Dun12b, Jam18]. Wodehouse [Ber18]. Wolfe [Gis19, Hol13a]. Wolff [Hen11a]. Wolffram [Som11]. Wollaston [Jam17]. Woloshyn [MY18]. Woman [Ren19]. Women [AM11, KB09, Sac16b, Tat17b, Ray15, SV17, Bha11, Mar12b]. Wonder [Sim17a]. Woodbridge [Hoc16a]. Woodmansee [Goo12a]. Woolgar [Lo15]. Worboys [Pea19]. Wordes [Tim12, LG11]. Words [Sma13]. Wordsworth [GP18, Jim19]. Work [D’A16, Loi16, Mae12, Rec10, Und12, Bro12, Far18, FMW14, Hud14, Mes10, RRL12, Hen15b]. working [Nas10, Win19b]. workmen [ZR15]. workplace [SM12]. Works [Sau13, Sto12, Boo15a, BB12, FF11]. Workshop [Sim13, BFGG11, Boo12, Nie10a]. World [Aga14, Asp11, Bec16, Bix19, Coo11, Far15, Gil12, Gou16, Gou17, Ham16, Kel17, Lau15, Lea16, Mar18b, Mil11c, Mor18, Opi14a, Pay16, Pha10, Pis17, Pos14, Rob10b, Ros17, Row14, San15b, Sin08, Sør17, Tot10, Tys17, WM11, Ale14, BM11, Dew10, Egm10, Fer11b, Fox12b, GWW09, Hol13b, Kla09, KCM13, Leh07, Leh19, LMS13, McA13, Poo10, Por09, Pri15, Rab17, Ros18, Rus14b, Bar15a, Far12, Fix16, McR14, Par12, Tes11, AK16]. Worldmakers [Shi16]. Worldmaking [Rih12, Leh12b]. Worlds [Joh11, Sma13, KMB18, Mar09]. Worse [Buk14, LOvT12]. Wotton [Sal14]. Wouldn’t [Can19b]. Wreckage [McC19]. Wrestling [HNS11, Bec12]. Wright [Goo17]. write [Hec17]. Writing [Ban14b, Chi08, Les13b, Lou16b, Ver16, And18, Dho15, Fer12, KWB15, Hig16, Til10]. Writings [Eve13, Tat18, SK12]. Wu [D’A16, Hud14]. Wuensch [vL13a, BW13]. wussten [BW13, vL13a]. Wyhe [Rad13]. Wynne [Wal12c].

X [Wal19a, And12a, And12b, Bak15a, Bal14b, Bec12, Bel17, Bil17, Biv10, Bjö14, Boo14, Bud11, Bur11, Cuo12, Dob14, Doh14, Doh16, Fan12, Fra17, Gou12a, Hen14, Hen15b, Hoc16a, Keh12, Lan11a, Les13b, Mar16, Mil11a, Mye12, Nas16, Osl10, Pri12, Ree10, Ric16, Sle12, Smi10c, Söd10, Sta14, Ste10, Tur16a, Tur14, Tur15, Tur16b, Tur12b, Whi16, Wol11, Wri16, McG19]. X-rays [McG19]. xi [Alb11a, Arb17, Bar16, Ber17, Biv10, Boo11, Car11, Cha10, Cle14, Cul11, Daw15c, Eve18, Far11, Far19, Gra11, Gre14b, Hig14a, Hoc14, Hoc16a, Hol16a, Jam10a, Joh14, Jon10b, Mac13, Mor18, Nav15a, Nor12, Ree19a, Rot16, SS15, Sha11, Sim17a, Sta15, Ste14b, Van14, Ver16, Wag14]. xii [Ale11, Ben19a, Boo17, Can15, Car18, Cha10, D’A16, Dij12, Dub10, Dun16b, Eic12a, Eic12b, Fer11a, Gar14, Gib14, Gou13a, Ham14a, Ham14b, Har16, Hen10, Hig10, Jam11a, Kee12, Kid11b, Ksi13, Lan12, Lee15b, Les15, Lou15, Lou17, Mac11b, Met11, Mil12, Mil15, Oos12, Pea18, Raj15, Ree16b, Ren15, Ric12, Rih12, Rob14, Sau13, SS11, Sch15b, SS13a, Sle11a, Smi14, Söd10, Sta11c, Sum11, Tim12, Tur10, Wag15, Web10, Wes15, Wig11, vL15]. xiii [And12d, Bak11, Bec16, BZ11a, Bil15a, Bix19, Boa12, Dun12a, Feo18, Fin10a, Fin11, Gou12c, Hal17, Hen13a, Hen13b, Hol14a, Hol15, Hol10, Hom16, Hoq10, Kni15, LN14, Law17, Les13a, Mac10, Mos16a, Pyl10, Rad14, Rob17, Roo12a, Sal15, San13, Sea10, Sle10, Smi19, Smo10, Sno10, Tat17a, Tys17, Wil11a, Wil11b]. XIV [Mil11b, Sho18, Alb11a, Alb15, Ban12, Bow16, Buk14, Chi12, Coo11, Cun10b, Daw10, Dun16a, Edw14, Fin10a, For10, Goo10a, Gou13b, Gou16, Gou12b, Har10a, Jam13, Mae12, Mul13, Nal12, Nic19, Opi14a, Pha10, Pin12, Qui12, Ree16a, Rob10b, Rod15a, Rod15b, Ros16, Sim14, Smi16b, Tam14a, Tee11, Tes11, Und13, Val11, Wil14, Wyl11, YM10, vN19, Dew09]. xix [Car12, Cle17, Ebe17, Fal13a, Far12, Goo10b, KK14, Ren10, Siv10, Sta11d]. XIXe [BFGG11, Sim13]. XIXe-XXe [Sim13]. xliv [Daw16b]. xlvi [Bak14]. xlviii [Mor13]. xv [And12c, Bar15b, Bui19, Coe10, Cuo13, Dih16, GP18, Jac15, Jam12, Jar10, Joh11, Pro11, Rap11, Røs15b, Sin15, Sto14, Tot10, Wad11, Wil12, Wol14, vL13b]. xvi [Ban15, Bel18b, Bow17, Cap15, Cur18, Daw12, Daw16a, Dun12a, Dun17b, Dys14, Eic15, Gas10, Gil12, Hei15, Hus15, Jar17, Kip16, Loi16, Mar17a, Mil11b, Mor13, Nal15, OP16, Par12, Pin13, Row10, Smi11b, Tar15, Tau11, Win11, Wit11b]. XVIe [Chr10, JS08]. xvii [Aga13a, Aga14, Ban14a, Bro11a, Bro16, Daw15a, Dry10, Egm13, Goo11a, Lan14, Mis11, Rap12, Ree17, Ric14, Roo15, San15b, Sch17a, Ser14a, Und12, Vol08]. XVII-th [Vol08]. XVIIe [Boa10, Vol08]. xviii [Ase17, Bak15b, Far10b, Gou17, Joh13a, Rad14, Ree14, Row14, Sør17, Win13, WM11]. xviiii [Mat13]. xx [Ade10, Bow10, Bud13a, Car16a, Chr10, Fon12, Lyk16, Mar18b, Pum12, Sar14, Ser12, Smi12, Sta19c, Tay14, Web10]. XXe [BFGG11, Sim13]. xxi [Bjö11, Boa15, Dun11, Hen11b, McO10, Pos14, Pri13c, Rob10a, Sma13, dC14]. xxii [Hay18, Kir15, Kum18, Rad13, Stu10]. xxiii [Art11, Big10, Bur13, GG12, Moo17, Smi11d, Tat16]. xxiv [Ber11a, Bur13, Chr10, Daw15b, Hig11, Hig13, Ser15, Smi11c, Sta11c, SJ11, Wal18]. xxix [LD18, Rob15]. xxv [Hug10, Mar12b]. xxvi [Ade10, Dun15b, Elt12, Fre10, Gil15a, Mai12c, Mar12c, Moo13, Rad13, Ren19, Ric10, Sho14, Tat18]. xxvii [Hoq10, Jon10a, Van11c, Wal12c]. xxviii [Lau15]. xxx [Bak15a, Bha11]. xxxi [Man16, Ros17]. xxxii [Kid11a, Sta16a]. xxxiv [Hig15, Leh12a, Rad13, Sto12]. xxxvii [Dun14].

Yahgan [Yan13]. Yalcinkaya [Tam16b]. Yale [Bod19, Cun10b, Die15, Far10a, Far11, Fin10c, Hol10, SS11, Sch17b, Ser14a, Sle10, Sno10, Wal11b]. Yaqzan [Wil14, BZ11b]. Year [Hal10, KKNN08]. Years [Aga14, KCM13, Sum14b]. Yellow [Coo14]. Yi [San16]. York [Ada10, Ade10, Aga14, And12b, Ase17, Ber18, Big18, Boa15, Bod18, Boo12, Bow18, Buk14, Cal12, D’A16, D’A18, Das18, Duf18, Dun16a, Dun17b, Ens12, Far12, Feo18, Fix16, Ful18, Goo10a, Gou17, Har16, Hol16c, Hus15, Joh13a, Kaa18, Ken19, Kid11a, Kum18, Lin17, Mac11b, Mar17a, Mil16, Moo13, Nal15, Nav16, Nes10, Oos12, Pea18, Pis16, Pos15a, Pri13c, Pum12, Ree17, Ree10, Ren15, Ric14, Rob17, Rob10b, Sea10, Siv10, Smi10b, Smi18a, Smo10, Som11, Sum11, Tat16, Tra16, Tys17, Wal17, Wig11, Wol14, Rot16]. you’re [Law15b]. Youth [Wol14, Rec10, Ter13]. Yves [Sim13, Tho19].

Z [Ksi13, Mur12, Sch15a]. Zachary [Bec18a]. Zadig [Man13a, Coh11]. Zakariya [Sac19]. Zaken [Wil14]. Zambesi [Dri10, Mac11b]. Zanetti [Ben19a]. Zebras [Sle12, Rit10].

Zeit [Hen11a]. Zeitmessung [BW13, vL13a]. Zelda [Hig16]. Zhores [Mai11]. Zimring [Mar18a]. Zipser [Ree15]. Zittel [Sho14]. Zodiac [Mur12, BJ10]. Zon [Daw15a]. Zoo [Hoc16a, vN19, Ito14]. zoology [Erl11, Erl13, Qui16, Sun13]. Zur [Sma13]. Zurich [Egm12]. Zwijnenberg [Wal16b].

References

Anonymous:2012:MIP

[A^{+}12] Anonymous et al., editors. Medical imaging and philosophy: challenges, reflections and actions: conference proceedings, volume 3 of KulturAnamnesen; Wissenschaftsforschung. Steiner, Stuttgart, Germany, 2012. ISBN 3-515-10046-6 (paperback). LCCN ????

Abraham:2009:SAC

[Abr09] Itty Abraham, editor. South Asian cultures of the bomb: atomic publics and the state in India and Pakistan. Indiana University Press, Bloomington, IN, USA, 2009. ISBN 0-253-35253-3 (cloth), 0-253-22032-7 (paperback). 222 pp. LCCN UA840 .S645 2009. URL http://www.

loc.gov/catdir/toc/
ecip0824/2008031607.html.

Aubin:2010:HEO

[ABS10] David Aubin, Charlotte Bigg, and Heinz Otto Sibum, editors. The Heavens on Earth: observatories and astronomy in Nineteenth-Century science and culture. Science and cultural theory. Duke University Press, Durham, NC, USA, 2010. ISBN 0-8223-9250-X (PDF), 0-8223-4628-1, 0-8223-4640-0. xii + 384 pp. LCCN QB32 .H43 2010.

URL http://site.ebrary.com/
lib/yale/Doc?id=10382983.

Abt:2011:AEL

[Abt11] Jeffrey Abt. American Egyptologist: the life of James Henry Breasted and the creation of his Oriental Institute. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-00110-5 (hardcover). xviii + 510 pp. LCCN D15.B67 A26 2011.

Anderson:2012:EEV

[AD12] Nancy (Nancy A.) Anderson and Michael R. Dietrich, editors. The educated eye: visual culture and pedagogy in the life sciences. Dartmouth College Press, Hanover, NH, USA, 2012. ISBN 1-61168-043-3, 1-61168-044-1 (paperback), 1-61168-212-6 (e-book). viii + 318 pp. LCCN QH315 .E326 2012.

Adam:2010:BRC

[Ada10] Alison Adam. Book review: Claire G. Jones, Femininity, Mathematics and Science, 1880– 1914. Basingstoke and New York: Palgrave Macmillan, 2009. Pp. ix + 264. ISBN 978-0-230-55521-1. £55.00 (hardback). British Journal for the History of Science, 43(3):494–496, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Adamson:2016:LDR

[Ada16] Matthew Adamson. Les liaisons dangereuses: resource surveillance, uranium diplomacy and secret French–American collaboration in 1950s Morocco. British Journal for the History of Science, 49(1):79–105, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Adelman:2009:CSN

[Ade09] Juliana Adelman. Communities of science in Nineteenth-Century Ireland. Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2009. ISBN 1-85196-653-6 (hardcover), 1-85196-654-4 (e-book). xi + 221 pp. LCCN Q175.52.I73 ADE 2009.

Adelman:2010:BRP

[Ade10] Juliana Adelman. Book reviews: Peter M. J. Hess and Paul L. Allen, Catholicism and Science. Westport, CT: Greenwood Press, 2008. Pp. xxvi + 241. ISBN 978-0-313-33190-9. $65.00 (hardback). Don O’Leary, Roman Catholicism and Modern Science: A History. London and New York: Continuum, 2006. Pp. xx + 356. ISBN 0-8264-1868-6. £18.99 (hardback). British Journal for the History of Science, 43(2):295– 296, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Adlem:2016:BRM

[Adl16] Jennifer Adlem. Book reviews:

Margaret E. Derry, Masterminding Nature: The Breeding of Animals, 1750–2010. Toronto: University of Toronto Press, 2015. Pp. viii + 310. ISBN 978-1-4426-2652-2. £18.99 (paperback). British Journal for the History of Science, 49(1):126–127, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Arnold-Forster:2016:PPL

[AF16] Agnes Arnold-Forster. A pathology of progress? locating the historiography of cancer. British Journal for the History of Science, 49(4):627–634, December 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/product/
1B685127054533F238801064AA49F123.

Al-Gailani:2017:DAC

[AG17] Salim Al-Gailani. ‘Drawing aside the curtain’: natural childbirth on screen in 1950s Britain. British Journal for the History of Science, 50(3):473–493, September 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/product/
EFF6B350947315C5CAD69E58788AD515.

Agar:2013:BRA

[Aga13a] Jon Agar. Book reviews: Alexander C. T. Geppert (ed.), Imagining Outer Space: European Astroculture in the Twentieth Century. Basingstoke: Palgrave Macmillan, 2012. Pp. xvii + 393. ISBN 978-0-230-23172-6. £70.00 (hardback). British Journal for the History of Science, 46(2):352– 354, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Agassi:2013:VIM

[Aga13b] Joseph Agassi. The Very Idea of Modern Science: Francis Bacon and Robert Boyle, volume 298 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346. xvii + 315 pp. LCCN Q175 .A318 2013; Q174

.B67 v. 298. URL https://link.
springer.com/book/10.1007/
978-94-007-5351-8.

Agar:2014:BRJ

[Aga14] Jon Agar. Book reviews: John Krige, Angelina Long Callahan and Ashok Maharaj, NASA in the World: Fifty Years of International Collaboration in Space. New York: Palgrave Macmillan, 2013. Pp. xvii + 353. ISBN 978-1-137-34092-4. £22.00 (paperback). British Journal for the History of Science, 47(4):757–759, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Agar:2015:BRN

[Aga15] Jon Agar. Book reviews: Naomi Oreskes and John Krige (eds.), Science and Technology in the Global Cold War. Cambridge, MA: MIT Press, 2014. Pp. 464. ISBN 978-0-262-52653-1. £25.95 (paperback). British Journal for the History of Science, 48(4): 715–716, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Agar:2017:BRD

[Aga17a] Jon Agar. Book review: David Kaiser and W. Patrick McCray (eds.), Groovy Science: Knowledge, Innovation and American Counterculture. Chicago: The University of Chicago Press, 2016. Pp. 426. ISBN 978-0-226-37291-4. £17.50/$25.00 (paperback). British Journal for the History of Science, 50(4):743– 744, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
1063186ED54111F1B87440965871898E.

Agar:2017:BRJ

[Aga17b] Jon  Agar.  Book

review: Jonathan Coopersmith, Faxed: The Rise and Fall of the Fax Machine. Baltimore: Johns Hopkins University Press, 2015. Pp. 320. ISBN 978-1-4214-1591-8. $54.95 (hardcover). British Journal for the History of Science, 50(2):370–371, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
4181790EFD58528E159D5A1CB83FCF3A.

Alfonso-Goldfarb:2013:GGG

[AGF13] Ana Maria Alfonso-Goldfarb and Marcia H. M. Ferraz. Gur, Ghur, Guhr or Bur? The quest for a metalliferous prime matter in early modern times. British Journal for the History of Science, 46(1):23–37, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anderson:2012:CJB

[AJ12] R. G. W. Anderson and Jean Jones, editors. The Correspondence of Joseph Black. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 0-7546-0131-5. xiv + 1564 pp. LCCN ????

Antonovics:2016:TLD

[AK16] Janis Antonovics and Jacobus Kritzinger. A translation of the Linnaean dissertation The Invisible World. British Journal for the History of Science, 49(3): 353–382, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
0B68BB5F50788F57518C006F10AE7618.

Ackermann:2014:SID

[AKM14] Silke Ackermann, Richard L.

(Richard Lynn) Kremer, and Mara Miniati, editors. Scientific instruments on display, volume 46(4) of History of science and medicine library; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-26439-6 (hardcover), 90-04-26440-X (e-book). ISSN 1872-0684. xxxiv + 231 pp. LCCN Q185 .S39 2014.

Akopyan:2019:BRG

[Ako19] Ovanes Akopyan. Book review: Gerrit Jasper Schenk (ed.), Historical Disaster Experiences: Towards a Comparative and Transcultural History of Disasters across Asia and Europe. Cham: Springer, 2017. Pp. ix + 436. ISBN 978-3-3194-9162-2. $139.00 (paperback). British Journal for the History of Science, 52(2):374– 375, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/gerrit-jasper-
                                        schenk-ed-historical-
                                        disaster-experiences-
                                        towards-a-comparative-and-
                                        transcultural-history-of-
                                        disasters-across-asia-and-
                                        europe-cham-springer-2017-
                                        pp-ix-436-isbn-9783319491622-
                                        13900-paperback/439576B2A98F03FA231F89293AAD61AA.

Alberti:2009:NCO

[Alb09] Samuel J. M. M. Alberti. Nature and culture: objects, disciplines and the Manchester Museum. Manchester University Press, Manchester, UK, 2009. ISBN 0-7190-8114-9 (hardcover). xi + 239 pp. LCCN ????

Alberti:2011:BRD

[Alb11a] Samuel J. M. M. Alberti. Book reviews: Diarmid A. Finnegan, Natural History Societies and Civic Culture in Victorian Scotland. London: Pickering and Chatto, 2009. Pp. xi + 254. ISBN 978-1-85196-658-5. £60.00 (hardback). Simon Naylor, Regionalizing Science: Placing Knowledges in Victorian England. London: Pickering and Chatto, 2010. Pp. xiv + 245. ISBN 978-1-85196-636-3. £60.00 (hardback). British Journal for the History of Science, 44(2):294–296, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Alberti:2011:MCM

[Alb11b] Samuel J. M. M. Alberti. Morbid curiosities: medical museums in Nineteenth-Century Britain. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-958458-3. xiii + 238 pp. LCCN R871 .A43 2011.

Alberti:2015:BRK

[Alb15] Samuel J. M. M. Alberti. Book reviews: Karen A. Rader and Victoria E.M. Cain, Life on Display: Revolutionizing U.S. Museums of Science and Natural History in the Twentieth Century. Chicago and London: The University of Chicago Press, 2014. Pp. xiv + 467. ISBN 978-0-226-07966-0. $45.00/£31.50 (hardback). British Journal for the History of Science, 48(4):719– 721, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Alberti:2016:BRS

[Alb16] Samuel J. M. M. Alberti. Book review: Samuel J. Redman, Bone Rooms: From Scientific Racism to Human Prehistory in Museums. Cambridge, MA: Harvard University Press, 2016. Pp. 373. ISBN 978-0-674-66041-0. £22.95 (hardback). British Journal for the History of Science, 49(4): 500–501, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        16FEB94FF7C541F1C7FA0883C92E740F.

Alexander:2008:MEW

[Ale08] Jennifer Karns Alexander. The mantra of efficiency: from waterwheel to social control. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8693-7 (hardcover). xvii + 233 pp. LCCN HC79.T4 A443

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0726/2007018866-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0726/
                                        2007018866-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0717/2007018866.html.

Alexander:2010:DDH

[Ale10] Amir Alexander. Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-04661-7. 307 pp. LCCN QA10.7 .A44 2010.

Alexander:2011:BRB

[Ale11]  Amir Alexander.  Book re-

view: Benjamin Wardhaugh, How to Read Historical Mathematics. Princeton and Oxford:

Princeton University Press, 2010. Pp. xii + 116. ISBN 978-0-691-14014-8. $22.95 (hardback). British Journal for the History of Science, 44(3):456–458, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Alexander:2014:IHD

[Ale14] Amir R. Alexander. Infinitesimal: how a dangerous mathematical theory shaped the modern world. Oneworld, London, UK, 2014. ISBN 1-78074-532-X (hardcover). 352 pp. LCCN ????

Allen:2012:SDJ

[All12] Julia Allen. Swimming with Dr.

Johnson and Mrs. Thrale: sport, health and exercise in Eighteenth-Century England. Lutterworth Press, Cambridge, UK, 2012. ISBN 0-7188-9276-3, 0-7188-4099-2 (e-book). viii + 297 pp. LCCN GV605 .A45 2012.

URL http://site.ebrary.com/
lib/yale/Doc?id=10634115.

Andreolle:2011:WSC

[AM11] Donna Spalding Andréolle and Véronique Molinari, editors. Women and science: 17th century to present: pioneers, activists and protagonists. Cambridge Scholars, Newcastle upon Tyne, UK, 2011. ISBN 1-4438-2918-8. xxv + 272 + 11 pp. LCCN Q130 .W6524 2011.

Abou-Nemeh:2018:BRA

[AN18] Catherine Abou-Nemeh. Book review: Ann Blair and Anja-Silvia Goeing (eds.), For the Sake of Learning: Essays in Honor of Anthony Grafton. Leiden and Boston, MA: Brill, 2016. Pp. lxxvii + 1082. ISBN 978-90-04-26330-7. EUR248.00 (hardcover). British Journal for the History of Science, 51(4): 717–719, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ann-blair-
and-anjasilvia-goeing-eds-
for-the-sake-of-learning-
essays-in-honor-of-anthony-
grafton-leiden-and-boston-
ma-brill-2016-pp-lxxvii-
1082-isbn-9789004263307-
24800-hardcover/
56C28E45862E9DD4D3B3D8CAC118D4FF.

Anderson:2008:CLS

[And08] Warwick Anderson. The collectors of lost souls: turning Kuru scientists into whitemen. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-9040-3 (hardcover). 318 pp. LCCN RC394.K8 A53

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0829/2008007840-b.html;
http://www.loc.gov/catdir/
enhancements/fy0829/
2008007840-d.html;
http://www.loc.gov/catdir/
toc/fy0904/2008007840.html.

AnduagaEgaana:2009:WEG

[And09] Aitor Anduaga Egaäna. Wireless and empire: geopolitics, radio industry, and ionosphere in the British Empire, 1918–1939. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956272-5. xxv + 386 pp. LCCN QC973.4.I6 A526

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0909/2008049036-b.html;
http://www.loc.gov/catdir/
enhancements/fy0909/
2008049036-d.html.

Andersen:2012:BRD

[And12a] Casper Andersen. Book review: David N. Livingstone and Charles W. J. Withers (eds.), Geographies of Nineteenth-Century Science. London and Chicago: The University of Chicago Press, 2011. Pp. x + 526. ISBN 978-0-226-48726-7. £35.50 (hardback). British Journal for the History of Science, 45(1):136–138, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [LW11].

Anderson:2012:BRV

[And12b] Julie Anderson. Book review:

Vanessa Heggie, A History of British Sports Medicine. Manchester and New York: Manchester University Press, 2011. Pp. x + 222. ISBN 978-0-7190-8261-0. £60.00 (hardback). British Journal for the History of Science, 45 (1):141–142, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Heg11].

Anderson:2012:BRE

[And12c] Nancy Anderson. Book reviews:

Eden Medina, Cybernetic Revolutionaries: Technology and Politics in Allende’s Chile. Cambridge, MA and London: MIT Press, 2011. Pp. xv + 326. ISBN 978-0-262-01649-0. £22.95 (hardback). British Journal for the History of Science, 45(4):704– 706, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anderson:2012:BRC

[And12d] Robert G. W. Anderson. Book reviews: Colin A. Russell and John A. Hudson, Early Railway Chemistry and Its Legacy. Cambridge: Royal Society of Chemistry, 2012. Pp. xiii + 193. ISBN 978-1-84973-326-7. £29.99 (paperback). British Journal for the History of Science, 45(4):688–690, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anderson:2013:LDL

[And13a] Ted R. Anderson. The life of David Lack: father of evolutionary ecology. Oxford University Press, 2013. ISBN 0-19-992264-0 (hardcover). x + 246 pp. LCCN QL31.L27 A53 2013.

Anderson:2013:AMG

[And13b] Thomas J. Anderson. Aepyornis as moa: giant birds and global connections in nineteenthcentury science. British Journal for the History of Science, 46(4): 675–693, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anderson:2018:RWS

[And18] Katharine Anderson. Reading and writing the scientific voyage: FitzRoy, Darwin and John Clunies Ross. British Journal for the History of Science, 51(3): 369–394, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/reading-
and-writing-the-scientific-
voyage-fitzroy-darwin-and-
john-clunies-ross/
B9312E2A974B546724AF70EBF355412C.

Anonymous:2010:BVIa

[Ano10a] Anonymous. BJH volume 43 issue 1 cover and back matter. British Journal for the History of Science, 43(1):b1–b7, March 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVIb

[Ano10b] Anonymous. BJH volume 43 issue 1 cover and front matter. British Journal for the History of Science, 43(1):f1–f2, March 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVIc

[Ano10c] Anonymous. BJH volume 43 issue 2 cover and back matter. British Journal for the History of Science, 43(2):b1–b9, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVId

[Ano10d] Anonymous. BJH volume 43 issue 2 cover and front matter. British Journal for the History of Science, 43(2):f1–f2, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVIe

[Ano10e] Anonymous. BJH volume 43 issue 3 cover and back matter. British Journal for the History of Science, 43(3):b1–b7, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVIf

[Ano10f] Anonymous. BJH volume 43 issue 3 cover and front matter. British Journal for the History of Science, 43(3):f1–f2, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVIh

[Ano10g] Anonymous. BJH volume 43 issue 4 cover and back matter. British Journal for the History of Science, 43(4):b1–b8, December 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BVIg

[Ano10h] Anonymous. BJH volume 43 issue 4 cover and front matter. British Journal for the History of Science, 43(4):f1–f2, December 2010.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BRa

[Ano10i] Anonymous. Books received.

British Journal for the History of Science, 43(1):145–147, March 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BRb

[Ano10j] Anonymous. Books received.

British Journal for the History of Science, 43(2):323–325, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BRc

[Ano10k] Anonymous. Books received.

British Journal for the History of Science, 43(3):509–512, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BRd

[Ano10l] Anonymous. Books received.

British Journal for the History of Science, 43(4):625–627, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:BRe

[Ano10m] Anonymous. Books reviewed.

British Journal for the History of Science, 43(4):631–634, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2010:IA

[Ano10n] Anonymous. Index of authors. British Journal for the History of Science, 43(4):629, December 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIb

[Ano11a] Anonymous. BJH volume 44 issue 1 cover and back matter. British Journal for the History of Science, 44(1):b1–b3, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIa

[Ano11b] Anonymous. BJH volume 44 issue 1 cover and front matter. British Journal for the History of Science, 44(1):f1–f2, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVId

[Ano11c] Anonymous. BJH volume 44 issue 2 cover and back matter. British Journal for the History of Science, 44(2):b1–b4, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIc

[Ano11d] Anonymous. BJH volume 44 issue 2 cover and front matter. British Journal for the History of Science, 44(2):f1–f2, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIf

[Ano11e] Anonymous. BJH volume 44 issue 3 cover and back matter. British Journal for the History of Science, 44(3):b1–b7, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIe

[Ano11f] Anonymous. BJH volume 44 issue 3 cover and front matter. British Journal for the History of Science, 44(3):f1–f2, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIh

[Ano11g] Anonymous. BJH volume 44 issue 4 cover and back matter. British Journal for the History of Science, 44(4):b1–b4, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BVIg

[Ano11h] Anonymous. BJH volume 44 issue 4 cover and front matter. British Journal for the History of Science, 44(4):f1–f2, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BRa

[Ano11i] Anonymous. Books received.

British Journal for the History of Science, 44(1):157–159, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BRb

[Ano11j] Anonymous. Books received. British Journal for the History of Science, 44(2):313–316, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BRc

[Ano11k] Anonymous. Books received. British Journal for the History of Science, 44(3):475–478, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BRd

[Ano11l] Anonymous. Books received. British Journal for the History of Science, 44(4):619–621, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:BRe

[Ano11m] Anonymous. Books reviewed. British Journal for the History of Science, 44(4):625–628, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2011:IA

[Ano11n] Anonymous. Index of authors. British Journal for the History of Science, 44(4):623, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIb

[Ano12a] Anonymous. BJH volume 45 issue 1 cover and back matter. British Journal for the History of Science, 45(1):b1–b3, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIa

[Ano12b] Anonymous. BJH volume 45 issue 1 cover and front matter. British Journal for the History of Science, 45(1):f1–f2, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVId

[Ano12c] Anonymous. BJH volume 45 issue 2 cover and back matter. British Journal for the History of Science, 45(2):??, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIc

[Ano12d] Anonymous. BJH volume 45 issue 2 cover and front matter. British Journal for the History of Science, 45(2):??, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIf

[Ano12e] Anonymous. BJH volume 45 issue 3 cover and back matter. British Journal for the History of Science, 45(3):b1–b2, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIe

[Ano12f] Anonymous. BJH volume 45 issue 3 cover and front matter. British Journal for the History of Science, 45(3):f1–f2, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIh

[Ano12g] Anonymous. BJH volume 45 issue 4 cover and back matter. British Journal for the History of Science, 45(4):??, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BVIg

[Ano12h] Anonymous. BJH volume 45 issue 4 cover and front matter. British Journal for the History of Science, 45(4):??, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BRa

[Ano12i] Anonymous. Books received. British Journal for the History of Science, 45(1):149–151, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BRb

[Ano12j] Anonymous. Books received. British Journal for the History of Science, 45(2):315–317, June 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BRc

[Ano12k] Anonymous. Books received. British Journal for the History of Science, 45(3):477–478, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BRd

[Ano12l] Anonymous. Books received.

British Journal for the History of Science, 45(4):711–713, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:BRe

[Ano12m] Anonymous. Books reviewed.

British Journal for the History of Science, 45(4):717–719, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2012:IA

[Ano12n] Anonymous. Index of authors.

British Journal for the History of Science, 45(4):715, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIb

[Ano13a] Anonymous. BJH volume 46 issue 1 cover and back matter. British Journal for the History of Science, 46(1):??, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIa

[Ano13b] Anonymous. BJH volume 46 issue 1 cover and front matter. British Journal for the History of Science, 46(1):??, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVId

[Ano13c] Anonymous. BJH volume 46 issue 2 cover and back matter. British Journal for the History of Science, 46(2):??, June 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIc

[Ano13d] Anonymous. BJH volume 46 issue 2 cover and front matter. British Journal for the History of Science, 46(2):??, June 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIf

[Ano13e] Anonymous. BJH volume 46 issue 3 cover and back matter. British Journal for the History of Science, 46(3):??, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIe

[Ano13f] Anonymous. BJH volume 46 issue 3 cover and front matter. British Journal for the History of Science, 46(3):??, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIh

[Ano13g] Anonymous. BJH volume 46 issue 4 cover and back matter. British Journal for the History of Science, 46(4):??, December 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BVIg

[Ano13h] Anonymous. BJH volume 46 issue 4 cover and front matter. British Journal for the History of Science, 46(4):??, December 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BRa

[Ano13i] Anonymous. Books received.

British Journal for the History of Science, 46(1):175–177, March 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BRb

[Ano13j] Anonymous. Books received.

British Journal for the History of Science, 46(2):357–358, June 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BRc

[Ano13k] Anonymous. Books received.

British Journal for the History of Science, 46(3):547–549, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BRd

[Ano13l] Anonymous. Books received.

British Journal for the History of Science, 46(4):739–742, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:BRe

[Ano13m] Anonymous. Books reviewed.

British Journal for the History of Science, 46(4):745–747, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2013:IA

[Ano13n] Anonymous. Index of authors. British Journal for the History of Science, 46(4):743, December 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIb

[Ano14a] Anonymous. BJH volume 47 issue 1 cover and back matter. British Journal for the History of Science, 47(1):??, March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVId

[Ano14b] Anonymous. BJH volume 47 issue 1 cover and back matter. British Journal for the History of Science, 47(1), March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIa

[Ano14c] Anonymous. BJH volume 47 issue 1 cover and front matter. British Journal for the History of Science, 47(1):??, March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIc

[Ano14d] Anonymous. BJH volume 47 issue 1 cover and front matter. British Journal for the History of Science, 47(1), March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIf

[Ano14e] Anonymous. BJH volume 47 issue 2 cover and back matter. British Journal for the History of Science, 47(2):??, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIe

[Ano14f] Anonymous. BJH volume 47 issue 2 cover and front matter. British Journal for the History of Science, 47(2):??, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIh

[Ano14g] Anonymous. BJH volume 47 issue 3 cover and back matter. British Journal for the History of Science, 47(3):??, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIg

[Ano14h] Anonymous. BJH volume 47 issue 3 cover and front matter. British Journal for the History of Science, 47(3):??, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIj

[Ano14i] Anonymous. BJH volume 47 issue 4 cover and back matter. British Journal for the History of Science, 47(4):??, December 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BVIi

[Ano14j] Anonymous. BJH volume 47 issue 4 cover and front matter. British Journal for the History of Science, 47(4):??, December 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BRa

[Ano14k] Anonymous. Books received. British Journal for the History of Science, 47(1):195–197, March 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BRb

[Ano14l] Anonymous. Books received. British Journal for the History of Science, 47(2):393–396, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BRc

[Ano14m] Anonymous. Books received. British Journal for the History of Science, 47(3):585–586, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BRd

[Ano14n] Anonymous. Books received. British Journal for the History of Science, 47(4):763–766, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:BRe

[Ano14o] Anonymous. Books reviewed.

British Journal for the History of Science, 47(4):769–771, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2014:IA

[Ano14p] Anonymous. Index of authors.

British Journal for the History of Science, 47(4):767–768, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIb

[Ano15a] Anonymous. BJH volume 48 issue 1 cover and back matter. British Journal for the History of Science, 48(1):??, March 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIa

[Ano15b] Anonymous. BJH volume 48 issue 1 cover and front matter. British Journal for the History of Science, 48(1):??, March 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVId

[Ano15c] Anonymous. BJH volume 48 issue 2 cover and back matter. British Journal for the History of Science, 48(2):??, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIc

[Ano15d] Anonymous. BJH volume 48 issue 2 cover and front matter. British Journal for the History of Science, 48(2):??, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIf

[Ano15e] Anonymous. BJH volume 48 issue 3 cover and back matter. British Journal for the History of Science, 48(3):??, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIe

[Ano15f] Anonymous. BJH volume 48 issue 3 cover and front matter. British Journal for the History of Science, 48(3):??, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIh

[Ano15g] Anonymous. BJH volume 48 issue 4 cover and back matter. British Journal for the History of Science, 48(4):??, December 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BVIg

[Ano15h] Anonymous. BJH volume 48 issue 4 cover and front matter. British Journal for the History of Science, 48(4):??, December 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:BR

[Ano15i] Anonymous. Books reviewed.

British Journal for the History of Science, 48(4):729–731, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2015:IA

[Ano15j] Anonymous. Index of authors.

British Journal for the History of Science, 48(4):727, December 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2016:BVIb

[Ano16a] Anonymous. BJH volume 49 issue 1 cover and back matter. British Journal for the History of Science, 49(1):??, March 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2016:BVIa

[Ano16b] Anonymous. BJH volume 49 issue 1 cover and front matter. British Journal for the History of Science, 49(1):??, March 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2016:BVId

[Ano16c] Anonymous. BJH volume 49 issue 2 cover and back matter. British Journal for the History of Science, 49(2):??, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2016:BVIc

[Ano16d] Anonymous. BJH volume 49 issue 2 cover and front matter. British Journal for the History of Science, 49(2):??, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2016:BVIf

[Ano16e] Anonymous. BJH volume 49 issue 3 cover and back matter. British Journal for the History of Science, 49(4):b1– b2, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
0CDC82F1299972FAC2CEC66E97391557.

Anonymous:2016:BVIe

[Ano16f] Anonymous. BJH volume 49 issue 3 cover and front matter. British Journal for the History of Science, 49(4):f1–f2, December 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
72B2C7932DAE6868E9F66B26846B4CF2.

Anonymous:2016:BVIh

[Ano16g] Anonymous. BJH volume 49 issue 4 cover and back matter. British Journal for the History of Science, 49(4):b1– b4, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
C3087F46C1DB7EF9A16DB34281AF3143.

Anonymous:2016:BVIg

[Ano16h] Anonymous. BJH volume 49 issue 4 cover and front matter. British Journal for the History of Science, 49(4):f1–f2, December 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
CE46141DFFDDB73B34B2BDDDB5BBF515.

Anonymous:2017:BVIb

[Ano17a] Anonymous. BJH volume 50 issue 1 cover and back matter. British Journal for the History of Science, 50(1):b1–b4, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E3533236EDE7C033C24C149B55445933.

Anonymous:2017:BVIa

[Ano17b] Anonymous. BJH volume 50 issue 1 cover and front matter. British Journal for the History of Science, 50(1):f1–f2, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
617B45BB57F7553FAB2632415A9BE9B8.

Anonymous:2017:BVId

[Ano17c] Anonymous. BJH volume 50 issue 2 cover and back matter. British Journal for the History of Science, 50(2):b1–b3, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
22B2E3A369294937BE09DEE63458C65A.

Anonymous:2017:BVIc

[Ano17d] Anonymous. BJH volume 50 issue 2 cover and front matter. British Journal for the History of Science, 50(2):f1–f2, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        B014A4ECB83C69705360A830EE841C86.

Anonymous:2017:BVIf

[Ano17e] Anonymous. BJH volume 50 issue 4 cover and back matter. British Journal for the History of Science, 50(4):b1– b3, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        CB7E31DF5BB3F619FBA723C43D85D87A.

Anonymous:2017:BVIe

[Ano17f] Anonymous. BJH volume 50 issue 4 cover and front matter. British Journal for the History of Science, 50(4):f1–f2, December 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        F5969325D0F6825E986E680550DDD396.

Anonymous:2017:PEF

[Ano17g] Anonymous. Past editors’ favourite papers published during their time in office. British Journal for the History of Science, 50(2):173–179, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.

org/core/product/
C1D0853C57ECEED1D72D992AAA4C45BB.

Anonymous:2018:BVIb

[Ano18a] Anonymous. BJH volume 51 issue 4 cover and back matter. British Journal for the History of Science, 51(4):b1– b3, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
51-issue-4-cover-and-back-
matter/
DC903F217038D010B712D9C3E5676644.

Anonymous:2018:BVIa

[Ano18b] Anonymous. BJH volume 51 issue 4 cover and front matter. British Journal for the History of Science, 51(4):f1–f2, December 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-51-
issue-4-cover-and-front-
matter/
A330474AAB766F4C277B7D85BDBD8EB2.

Anonymous:2019:BVIb

[Ano19a] Anonymous. BJH volume 52 issue 1 cover and back matter. British Journal for the History of Science, 52(1):b1–b3, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

                                        journal-for-the-history-of-
                                        science/article/bjh-volume-
                                        52-issue-1-cover-and-back-
                                        matter/
                                        9E51BA60C427693E45FD5B18F9EA21D6.

Anonymous:2019:BVIa

[Ano19b] Anonymous. BJH volume 52 issue 1 cover and front matter. British Journal for the History of Science, 52(1):f1–f2, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bjh-volume-
                                        52-issue-1-cover-and-front-
                                        matter/
                                        DCD1FEEE29FAB0AC16CA56D6324A1738.

Anonymous:2019:BVId

[Ano19c] Anonymous. BJH volume 52 issue 2 cover and back matter. British Journal for the History of Science, 52(2):b1–b2, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bjh-volume-
                                        52-issue-2-cover-and-back-
                                        matter/
                                        6F74F3137F9B6D9ADBE2FA1C4B8F3E2C.

Anonymous:2019:BVIc

[Ano19d] Anonymous. BJH volume 52 issue 2 cover and front matter. British Journal for the History of Science, 52(2):f1–f2, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.

org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
52-issue-2-cover-and-front-
matter/
EDB54B8A0ECB4A5F14A57EEE3B41AE7A.

Anonymous:2019:BVIf

[Ano19e] Anonymous. BJH volume 52 issue 3 cover and back matter. British Journal for the History of Science, 52(3):b1– b3, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
52-issue-3-cover-and-back-
matter/
0275896C2A1D09BB3BD48E3A7D9C213E.

Anonymous:2019:BVIe

[Ano19f] Anonymous. BJH volume 52 issue 3 cover and front matter. British Journal for the History of Science, 52(3):f1–f2, September 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-52-
issue-3-cover-and-front-
matter/
6F25DA0D20C52AB9E31EFD2A272A0BD1.

Anonymous:2019:F

[Ano19g] Anonymous. Forum. British Journal for the History of Science, 52(4):??, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/what-oral-
                                        historians-and-historians-
                                        of-science-can-learn-from-
                                        each-other/
                                        A2A5ADC7166C7ED18DE21332BC515F66.

Anderson:2007:SMP

[ANPR07] Julie Anderson, Francis Neary, John V. Pickstone, and James Raftery. Surgeons, manufacturers and patients: a transatlantic history of total hip replacement. Science, technology, and medicine in modern history. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2007. ISBN 0-230-55314-1 (hardcover). xiv + 222 pp. LCCN RD549 .A53 2007.

Anstey:2011:JLN

[Ans11] Peter R. Anstey. John Locke and natural philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-958977-1 (hardcover). x + 252 pp. LCCN B1284 .A57 2011.

Aratus:2010:P

[APoC10] Solensis Aratus, Aaron Poochigian, and Eudoxus of Cnidus. Phaenomena. Johns Hopkins new translations from antiquity. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9465-4 (hardcover), 0-8018-9466-2 (paperback). xxxi + 72 pp. LCCN QB802 .A713 2010. Translated and with an introduction and notes and by Aaron Poochigian.

Appel:2012:ATS

[App12] Andrew W. Appel, editor. Alan Turing’s systems of logic: the Princeton thesis. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15574-7 (hardcover). xv + 142 pp. LCCN QA9.2 .T86 2012. Reproduction of Turing’s thesis [Tur38].

Arbel:2017:BRA

[Arb17] Tal Arbel. Book review: Aaron Panofsky, Misbehaving Science: Controversy and the Development of Behavior Genetics. Chicago: The University of Chicago Press, 2014. Pp. xi + 321. ISBN 978-0226-05845-

0. £19.50 (paperback). British Journal for the History of Science, 50(2):376–377, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
20BFC89A3DA801AC8C29C6747704B723.

Arnold:2016:BRP

[Arn16] David Arnold. Book review: Paul R. Josephson, Fish Sticks, Sports Bras, and Aluminum Cans: The Politics of Everyday Technologies. Baltimore: Johns Hopkins University Press, 2015. Pp. ix + 212. ISBN 978-1-4214-1783-7. $29.95 (paperback). British Journal for the History of Science, 49 (4):521–523, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        6ED826E74617C5C993D214514294C6EC.

Arthur:2011:BRN

[Art11] Richard T. W. Arthur. Book review: Niccolò Guicciardini, Isaac Newton on Mathematical Certainty and Method. Cambridge, MA and London: MIT Press, 2009. Pp. xxiii + 422. ISBN 978-0-262-01317-8. $55.00 (hardback). British Journal for the History of Science, 44(1):122– 124, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ash:2012:NSK

[AS12] Mitchell G. Ash and Jan Surman, editors. The nationalization of scientific knowledge in the Habsburg Empire, 1848–1918. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, 2012. ISBN 0-230-28987-8. xi + 258 pp. LCCN Q127.E81 N28 2012.

Aselmeyer:2017:BRM

[Ase17] Norman Aselmeyer. Book review: Maria Paula Diogo and Dirk van Laak, Europeans Globalizing: Mapping, Exploiting, Exchanging. London and New York: Palgrave Macmillan, 2016. Pp. xviii + 352. ISBN 978-0-230-27963-6. £60.00 (hardcover). British Journal for the History of Science, 50 (4):737–738, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.

org/core/product/
61D2BB9DF95A4A0D61CCF86FB7A160E5.

Aspengren:2011:SKC

[Asp11] Henrik C. Aspengren. Sociological knowledge and colonial power in Bombay around the First World War. British Journal for the History of Science, 44 (4):533–548, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Aylward:2017:BRL

[Ayl17]  Alex  Aylward.

Book review: Léna Soler, Emiliano Trizio and Andrew Pickering (eds.), Science as It Could Have Been: Discussing the Contingency/Inevitability Problem. Pittsburgh: The University of Pittsburgh Press, 2015. Pp. 462. ISBN 978-0-8229-4445-4. $61.95 (hardback). British Journal for the History of Science, 50(2):379–381, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
14EBF54BAC1627D8819AB08A05EAB5ED.

Aylward:2018:BRH

[Ayl18]  Alex  Ayl-

ward. Book review: Hans–Jörg Rheinberger and Staffan Müller– Wille, The Gene: From Genetics to Postgenomics. Translated by Adam Bostanci. Chicago and London: The University of Chicago Press, 2017. Pp. 147. ISBN 978-0-2265-1000-2. $25.00 (paperback). British Journal for the History of Science, 51(3):

534–535, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/hansjorg-
rheinberger-and-staffan-
mullerwille-the-gene-from-
genetics-to-postgenomics-
translated-by-adam-
bostanci-chicago-and-
london-the-university-of-
chicago-press-2017-pp-147-isbn-
9780226510002-2500-paperback/
50E87B11422F94E190978CA01A5B7D2A.

Aylward:2019:BRE

[Ayl19] Alex Aylward. Book review: Efram Sera-Shriar (ed.), Historicizing Humans: Deep Time, Evolution, and Race in Nineteenth-Century British Sciences. Pittsburgh, PA: University of Pittsburgh Press, 2018. Pp. 326. ISBN 978-0-8229-4529-1. $45.00 (hardcover). British Journal for the History of Science, 52(2):369– 370, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/efram-serashriar-
ed-historicizing-humans-
deep-time-evolution-and-
race-in-nineteenthcentury-
british-sciences-
pittsburgh-pa-university-
of-pittsburgh-press-2018-
pp-326-isbn-9780822945291-4500-
hardcover/1046FFB032EF24D242CF3DC8CD67550F.

Ayres:2008:APD

[Ayr08] Peter G. Ayres. The aliveness of plants: the Darwins at the dawn of plant science. Pickering and Chatto, London, UK, 2008. ISBN 1-85196-970-5 (hardcover). xiii + 227 pp. LCCN QK26.A97; QK26.A97 2008.

Ayres:2012:SEL

[Ayr12] P. G. (Peter G.) Ayres. Shaping ecology: the life of Arthur Tansley. Wiley-Blackwell, Chichester, West Sussex, UK, 2012. ISBN 0-470-67156-4 (hardcover), 0-470-67154-8 (paperback). xii + 213 pp. LCCN QH31.T37 A87 2012; QH31.T37 A87X 2012 (LC).

Azzolini:2012:DSA

[Azz12] Monica Azzolini. The duke and the stars: astrology and politics in Renaissance Milan. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06663-4. 370 pp. LCCN BF1729.P6 A99 2012.

Bleichmar:2009:SSP

[B^{+}09] Daniela Bleichmar et al., editors. Science in the Spanish and Portuguese empires, 1500–1800. Stanford University Press, Stanford, CA, USA, 2009. ISBN 0-8047-5358-X (hardcover). xxii + 427 pp. LCCN Q127.L38 S36 2009; M10.E03051. URL http:

//www.loc.gov/catdir/toc/
ecip0828/2008041298.htm.

Badino:2016:BRJ

[Bad16] Massimiliano Badino. Book review: J.L. Heilbron, Physics: A Short History from Quintessence to Quarks. Oxford: Oxford University Press, 2015. Pp. 224. ISBN 978-0-19-874685-0. £10.99 (hardback). British Journal for the History of Science, 49(3): 474–476, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
599AFDE65EE5CE2D0BC42158084475A7.

Badino:2019:BRE

[Bad19] Massimiliano Badino. Book review: Eric Johnson, Anxiety and the Equation: Understanding Boltzmann’s Entropy. Cambridge, MA: MIT Press, 2018. Pp. ix + 179. ISBN 978-0-2620-3861-4. $22.95/£17.99 (cloth). British Journal for the History of Science, 52(4):721–722, December 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/eric-johnson-
anxiety-and-the-equation-
understanding-boltzmanns-
entropy-cambridge-ma-mit-
press-2018-pp-ix-179-isbn-
9780262038614-22951799-
cloth/
A8E39E6348FEC79D2FCA4F1EEBCE9228.

Baker:2011:BRJ

[Bak11] Alexi Baker. Book review: Jeffrey R. Wigelsworth, Selling Science in the Age of Newton: Advertising and the Commoditization of Knowledge. Farnham: Ashgate, 2010. Pp. xiii + 203. ISBN 978-1-4094-0075-2. £65.00 (hardback). British Journal for the History of Science, 44(4):595–597, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Baker:2014:BRA

[Bak14] Alexi Baker. Book review: Alison D. Morrison-Low, Sven Dupré, Stephen Johnston and Giorgio Strano (eds.), From Earth-Bound to Satellite: Telescopes, Skills and Networks. Leiden: Brill, 2012. Pp. xlvi + 265. ISBN 978-90-04-21150-6. EUR99.00 (hardback). British Journal for the History of Science, 47(1):181– 182, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Baker:2015:BRS

[Bak15a] Alexi Baker. Book reviews:

Scholars in Action: The Practice of Knowledge and the Figure of the Savant in the 18th Century. 2 vols. Leiden: Brill, 2013. Pp. xxx + 501 and x + 425. ISBN 978-90-04-24390-3. EUR229.00 (hardback). André Holenstein, Hubert Steinke and Martin Stuber (eds.). British Journal for the History of Science, 48(1):178– 180, March 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Baker:2015:BRJ

[Bak15b] Alexi Baker. Book reviews:

James Sumner, Brewing Science, Technology and Print, 1700– 1880. London: Pickering and Chatto, 2013. Pp. xviii + 295. ISBN 978-1-84893-423-8. £60.00 (hardback). British Journal for the History of Science, 48(4): 693–694, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ballstadt:2008:DNP

[Bal08] Kurt P. A. (Kurt Paul Alexander) Ballstadt. Diderot: natural philosopher, volume 2008:09 of SVEC. Voltaire Foundation, Oxford, UK, 2008. ISBN 0-7294-0948-1 (paperback). viii + 246 pp. LCCN PQ2105.A2 B482 2008 no.9; PQ1979 .B35 2008; B482 2008.

Baldwin:2014:KRP

[Bal14a] Melinda Baldwin. ‘Keeping in the race’: physics, publication speed and national publishing strategies in Nature, 1895–1939. British Journal for the History of Science, 47(2):257–279, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://journals.
                                        cambridge.org/abstract_
                                        S0007087413000381. Published

online, but not yet assigned to a journal volume.

Balmer:2014:BRA

[Bal14b] Andrew S. Balmer. Book reviews: Alison Winter, Memory: Fragments of a Modern History. Chicago and London: The University of Chicago Press, 2012. Pp. x + 319. ISBN 978-0-226-90258-6. £21.00 (hardback).

British Journal for the History of Science, 47(4):750–751, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Baldwin:2015:MNH

[Bal15] Melinda Clare Baldwin. Making Nature: the history of a scientific journal. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-26145-X (hardcover), 0-226-26159-X (e-book). 309 pp. LCCN Q1.N23 B35 2015.

Baldassarri:2019:MLP

[Bal19] Fabrizio Baldassarri. The mechanical life of plants: Descartes on botany. British Journal for the History of Science, 52(1):41– 63, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/mechanical-life-of-
plants-descartes-on-botany/
44FD912FBCA79F91176BD509E07496D7.

Baneke:2011:STD

[Ban11] David Baneke. Synthetic technocracy: Dutch scientific intellectuals in science, society and culture, 1880–1950. British Journal for the History of Science, 44 (1):89–113, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Banham:2012:BRC

[Ban12] Debby Banham. Book review:

Calvin B. Kendall and Faith Wallis (trans.), Bede: On the Nature of Things and On Times. Translated Texts for Historians 56. Liverpool: Liverpool University Press, 2010. Pp. xiv + 222. ISBN 978-1-8463-1496-4. £16.99 (paperback). British Journal for the History of Science, 45 (1):125–126, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [KW10].

Banerjee:2014:BRJ

[Ban14a] Somaditya Banerjee. Book reviews: Jahnavi Phalkey, Atomic State: Big Science in Twentieth-Century India. Ranikhet: Permanent Black, 2013. Pp. xvii + 335. ISBN 81-7824-376-8. INR 795.00 (hardback). British Journal for the History of Science, 47 (4):754–756, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bangham:2014:WPS

[Ban14b] Jenny Bangham. Writing, printing, speaking: Rhesus bloodgroup genetics and nomenclatures in the mid-twentieth century. British Journal for the History of Science, 47(2):335–361, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bangham:2015:BRJ

[Ban15] Jenny Bangham. Book reviews: Joseph November, Biomedical Computing: Digitizing Life in the United States. Baltimore: Johns Hopkins University Press, 2012.

Pp. xvi + 344. ISBN 978-1-4214-0468-4. £31.00 (hardback). British Journal for the History of Science, 48(3):536–537, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bartlett:2008:NSM

[Bar08]  Robert Bartlett.  The natu-

ral and the supernatural in the Middle Ages: the Wiles lecture given at the Queen’s University of Belfast, 2006. The Wiles lectures. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87832-2 (hardcover), 0-521-70255-0 (paperback). x + 170 pp. LCCN BF1411 .B27

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0808/2007044426-b.html;
http://www.loc.gov/catdir/
enhancements/fy0808/
2007044426-d.html;
http://www.loc.gov/catdir/
enhancements/fy0808/
2007044426-t.html.

Barany:2013:BRB

[Bar13a] Michael J. Barany. Book reviews:

Benjamin Wardhaugh (ed.), The History of the History of Mathematics: Case Studies for the Seventeenth, Eighteenth and Nineteenth Centuries. Oxford: Peter Lang, 2012. Pp. vi + 187. ISBN 978-3-0343-0708-6. £32.00 (paperback). British Journal for the History of Science, 46(2):344– 345, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barany:2013:BRH

[Bar13b] Michael J. Barany. Book reviews: Hawking Incorporated: Stephen Hawking and the Anthropology of the Knowing Subject. Chicago and London: The University of Chicago Press, 2012. Pp. ix + 266. ISBN 978-0-226-52228-9. £18.50 (paperback), by Hélène Mialet. British Journal for the History of Science, 46(3): 544–546, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barany:2014:SNE

[Bar14]  Michael J. Barany.  Savage

numbers and the evolution of civilization in Victorian prehistory. British Journal for the History of Science, 47(2):239–255, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barany:2015:BRA

[Bar15a] Michael J. Barany. Book reviews: Amir Alexander, Infinitesimal: How a Dangerous Mathematical Theory Shaped the Modern World. London: Oneworld, 2014. Pp. 352. ISBN 978-1-78074-532-9. £20.00 (paperback). British Journal for the History of Science, 48(2):362–364, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barany:2015:BRI

[Bar15b] Michael J. Barany. Book reviews: Ian Hacking, Why Is There Philosophy of Mathematics at All?, Cambridge: Cambridge University Press, 2014. Pp. xv + 290. ISBN 978-1-107-65815-8. £17.99 (paperback). British Journal for the History of Science, 48(4): 686–687, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barrett:2015:BRC

[Bar15c] Katy Barrett. Book reviews:

Coral: Something Rich and Strange. Liverpool: Liverpool University Press, 2013. Pp. 128. ISBN 978-1-84631-959-4. £19.99 (paperback). Marion Endt-Jones (ed.). British Journal for the History of Science, 48(1):175–176, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barany:2016:BRC

[Bar16] Michael J. Barany. Book reviews:

Christopher Hollings, Mathematics across the Iron Curtain: A History of the Algebraic Theory of Semigroups. Providence, RI: American Mathematical Society, 2014. Pp. xi + 441. ISBN 978-1-4704-1493-1. £79.95 (hardback). British Journal for the History of Science, 49(1):140– 141, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barberis:2017:BRS

[Bar17] Daniela S. Barberis. Book review: Sabine Arnaud, On Hysteria: The Invention of a Medical Category between 1670 & 1820. Chicago: The University of Chicago Press, 2015. Pp. 376. ISBN 978-0-226-27554-

3. £38.50 (hardback). British Journal for the History of Science, 50(1):148–149, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        D595C1A1A32EA289A076F5083B7FCD91.

Bashford:2014:GPH

[Bas14] Alison Bashford. Global population: history, geopolitics, and life on earth. Columbia studies in international and global history. Columbia University Press, New York, NY, USA, 2014. ISBN 0-231-14766-X. xii + 466 pp. LCCN HB849.44.

Bates:2010:ARK

[Bat10] A. W. (Alan W.) Bates. The anatomy of Robert Knox: murder, mad science and medical regulation in Nineteenth-Century Edinburgh. Sussex Academic, Brighton, UK, 2010. ISBN 1-84519-381-4 (hardcover). ix + 228 pp. LCCN QM16.K6 2010.

Baxfield:2013:MSP

[Bax13] C. R. C. Baxfield. ‘To mend the scheme of Providence’: Benjamin Franklin’s electrical heterodoxy. British Journal for the History of Science, 46(2):179–197, June 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Everett:2012:EIQ

[BB12] Jeffrey Alan Barrett and Peter Byrne, editors. The Everett interpretation of quantum mechanics: collected works 1955–1980 with commentary. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14507-5. xii + 389 pp. LCCN QC174.12 .E96 2012.

Bertrams:2013:SHM

[BCH^{+}13] Kenneth Bertrams, Nicolas Coupain, Ernst Homburg, G. Kurgan-van Hentenryk, and Philippe Mioche. Solvay: history of a multinational family firm. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-02480-3 (hardcover). xvii + 630 pp. LCCN HD9656.B44 S653 2013.

Beretta:2012:SSN

[BCP12] Marco Beretta, Francesco Citti, and Lucia Pasetti, editors. Seneca e le Scienze Naturali. (Italian) [Seneca and natural science]. Leo S. Olschki Editore, Florence, Italy, 2012. ISBN 88-222-6198-5. vi + 273 pp. LCCN ???? EUR 29.00.

Brusius:2013:WHF

[BDR13] Mirjam Brusius, Katrina Dean, and Chitra Ramalingam, editors. William Henry Fox Talbot: beyond photography, volume 23 of Studies in British art. The Yale Center for British Art, The Paul Mellon Centre for Studies in British Art, New Haven, CT, 2013. ISBN 0-300-17934-0 (hardcover). viii + 308 pp. LCCN TR140.T3 W55 2013.

Bredekamp:2015:TBT

[BDS15] Horst Bredekamp, Vera Dünkel, and Birgit Schneider, editors. Technische Bild. The technical image: a history of styles in scientific imagery. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-25884-X (hardcover), 0-226-25898-X (e-book). LCCN T11.8 .T4413 2015.

Beatty:2011:NDL

[Bea11] Heather Beatty. Nervous disease in late eighteenth-century Britain: the reality of a fashionable disorder, volume 6 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-309-6 (Pickering and Chatto), 1-283-29204-1 (MyiLibrary), 1-84893-308-8 (ebook). vii + 241 pp. LCCN RC506 .B43 2011.

Beck:2012:BRP

[Bec12] David Beck. Book review: Peter Harrison, Ronald L. Numbers and Michael H. Shank (eds.), Wrestling with Nature: From Omens to Science. Chicago and London: Chicago University Press, 2011. Pp. x + 416. ISBN 978-0-226-31783-0. £22.50 (paperback). British Journal for the History of Science, 45(2):282– 283, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Beck:2015:KNE

[Bec15] David Beck. Knowing nature in early modern Europe, volume 5 of Warwick series in the humanities. Pickering and Chatto, London, UK, 2015. ISBN 1-84893-518-8, 1-78144-496-X (WebPDF), 1-78144-497-8 (ePub). xii + 227 pp. LCCN BD161 .K5635 2015.

Beck:2016:BRC

[Bec16] Emily E. Beck. Book review:

Cristina Bellorini, The World of Plants in Renaissance Tuscany: Medicine and Botany. Farnham: Ashgate, 2016. Pp. xiii + 261. ISBN 978-1-4724-6622-8. £95.00 (hardback). British Journal for the History of Science, 49(4): 641–643, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A0EB0A5FDB8734A99EEC8D4902CAFBEC.

Beck:2018:BRZ

[Bec18a] Emily E. Beck. Book review:

Zachary A. Matus, Franciscans and the Elixir of Life: Religion and Science in the Later Middle Ages. Philadelphia: University of Pennsylvania Press, 2017. Pp. 203. ISBN 978-0-8122-4921-

7. £52.00 (cloth). British Journal for the History of Science, 51 (4):703–705, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/zachary-a-
matus-franciscans-and-the-
elixir-of-life-religion-
and-science-in-the-later-
middle-ages-philadelphia-

                                        university-of-pennsylvania-
                                        press-2017-pp-203-isbn-
                                        9780812249217-5200-cloth/
                                        AA065E121BA1806A4B7E643104629180.

Beckers:2018:BRD

[Bec18b] Danny Beckers. Book review: Daniel Warner, Live Wires: A History of Electronic Music. London: Reaktion Books, 2017. Pp.

208. ISBN 978-1-78023-824-1. £16.00 (cloth cover). British Journal for the History of Science, 51(4):716–717, December 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/daniel-warner-live-
                                        wires-a-history-of-
                                        electronic-music-london-
                                        reaktion-books-2017-pp-208-
                                        isbn-9781780238241-1600-
                                        cloth-cover/
                                        8137C43E98FC9D0FF0CEB293C49F4A7B.

Beckers:2018:BRR

[Bec18c] Danny Beckers. Book review: Ronald S. Calinger, Leonhard Euler: Mathematical Genius in the Enlightenment. Princeton, NJ and Oxford: Princeton University Press, 2016. Pp. 669. ISBN 978-0-691-11927-4. $55.00/ EUR40.95 (cloth). British Journal for the History of Science, 51(1):155–156, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-

science/article/ronald-s-
calinger-leonhard-euler-
mathematical-genius-in-the-
enlightenment-princeton-nj-
and-oxford-princeton-
university-press-2016-pp-
669-isbn-9780691119274-
55004095-cloth/6F4EB8125512F1EA40BC35DA16CB6542.

Beeley:2019:PMM

[Bee19]  Philip Beeley.  Practical

mathematicians and mathematical practice in later seventeenthcentury London. British Journal for the History of Science, 52(2):225–248, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/practical-
mathematicians-and-
mathematical-practice-in-
later-seventeenthcentury-
london/
CB8DABA63754F007A51AAE9491D4840A.

Belknap:2012:BRC

[Bel12] Geoffrey Belknap. Book review:

Christopher Pinney, Photography and Anthropology. London: Reaktion Books, 2011. Pp. 174. ISBN 978-1-86189-804-3. £17.95 (paperback). British Journal for the History of Science, 45(3): 469–470, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Belknap:2017:BRS

[Bel17] Geoffrey Belknap. Book review: Sarah K. Gillespie, The Early American Daguerreotype:

Cross-currents in Art and Technology. Cambridge, MA: MIT Press, 2016, Pp. x + 213. ISBN 978 0 2620 3410 4. £23.95 (cloth). British Journal for the History of Science, 50(3): 551–552, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
5CF366B99FB541485059C0665E3F76D3.

Belknap:2018:INH

[Bel18a]  Geoffrey  Belk-

nap. Illustrating natural history: images, periodicals, and the making of nineteenth-century scientific communities. British Journal for the History of Science, 51(3): 395–422, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
illustrating-natural-
history-images-periodicals-
and-the-making-of-
nineteenthcentury-
scientific-communities/
8B1DA7E7CA3CE9FBDA3311F5BF5F514F.

Bellis:2018:BRD

[Bel18b] Richard T. Bellis. Book review: Domenico Bertoloni Meli, Visualizing Disease: The Art and History of Pathological Illustrations. Chicago and London: The University of Chicago Press, 2017. Pp. xvi + 294. ISBN 978-0-2261-1029-5. $55.00 (cloth). British Journal for the History of Science, 51(3): 522–524, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/domenico-
bertoloni-meli-visualizing-
disease-the-art-and-
history-of-pathological-
illustrations-chicago-and-
london-the-university-of-
chicago-press-2017-pp-xvi-
294-isbn-9780226110295-
5500-cloth/FD9C5AE4FDC504C12E659E2A27E3747A.

Bentley:2011:LTH

[Ben11] Michael Bentley. The life and thought of Herbert Butterfield: history, science, and God. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-00397-0 (hardcover). xv + 381 pp. LCCN

????  URL http://assets.
cambridge.org/97811070/
03972/cover/9781107003972.
jpg;   http://www.loc.gov/
catdir/enhancements/fy1106/
2011001054-b.html;
http://www.loc.gov/catdir/
enhancements/fy1106/
2011001054-d.html;
http://www.loc.gov/catdir/
enhancements/fy1106/2011001054-
t.html.

Benjamin:2013:PSB

[Ben13a] Ruha Benjamin. People’s science:

bodies and rights on the stem cell frontier. Stanford University Press, Stanford, CA, USA, 2013. ISBN 0-8047-8296-2 (hardcover), 0-8047-8297-0 (paperback), 0-8047-8673-9 (e-book). xiii + 249 pp. LCCN QH588.S83 B46 2013.

Benton:2013:ARW

[Ben13b] Ted Benton. Alfred Russel Wallace: Explorer, Evolutionist, Public Intellectual: A Thinker for Our Own Times? Siri Scientific Press, Manchester, UK, 2013. ISBN 0-9574530-2-7. 224 pp. LCCN ????

Bennett:2019:BRC

[Ben19a] Jim Bennett. Book review: Cristiano Zanetti, Janello Torriani and the Spanish Empire: A Vitruvian Artisan at the Dawn of the Scientific Revolution. Leiden, Boston, Paderborn and Singapore: Brill, 2018. Pp. xii + 450. ISBN 978-9-0043-2089-5. EUR95.00/$110.00 (hardcover). British Journal for the History of Science, 52(3):525–526, September 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/cristiano-zanetti-
                                        janello-torriani-and-the-
                                        spanish-empire-a-vitruvian-
                                        artisan-at-the-dawn-of-the-
                                        scientific-revolution-
                                        leiden-boston-paderborn-
                                        and-singapore-brill-2018-
                                        pp-xii-450-isbn-9789004320895-
                                        950011000-hardcover/566405EB801E00A133CC37385DA52843.

Bennett:2019:MBI

[Ben19b] Jim Bennett. Mathematicians on board: introducing lunar distances to life at sea. British Journal for the History of Science, 52(1):65–83, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
mathematicians-on-board-
introducing-lunar-
distances-to-life-at-sea/
81D23ECFEEB71E604CBC8952BC3E5619.

Berkowitz:2011:BRA

[Ber11a] Carin Berkowitz. Book review: Andrew Cunningham, The Anatomist Anatomis’d: An Experimental Discipline in Enlightenment Europe. Farnham: Ashgate, 2010. Pp. xxiv + 443. ISBN 978-0-7546-6338-6. £65.00 (hardback). British Journal for the History of Science, 44(2):291– 293, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

BertoloniMeli:2011:MED

[Ber11b] Domenico Bertoloni Meli. Mechanism, experiment, disease: Marcello Malpighi and Seventeenth-Century anatomy. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 0-8018-9903-6 (hardcover), 0-8018-9904-4 (paperback). xii + 439 pp. LCCN QM16.M33 B47 2011.

Berry:2012:BRM

[Ber12] Dominic Berry. Book reviews:

Margaret E. Derry, Art and Science in Breeding: Creating Better Chickens. Toronto, Buffalo and London: University of Toronto Press, 2012. Pp. viii + 281. ISBN 978-1-4426-4395-6. $65.00 (hardback). British Journal for the History of Science, 45(4):695– 697, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Berkowitz:2013:SDM

[Ber13] Carin Berkowitz. Systems of display: the making of anatomical knowledge in Enlightenment Britain. British Journal for the History of Science, 46(3): 359–387, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Berry:2015:BRO

[Ber15] Dominic Berry. Book reviews: Oren Harman and Michael R. Dietrich (eds.), Outsider Scientists: Routes to Innovation in Biology. Chicago and London: The University of Chicago Press, 2013. Pp. ix + 374. ISBN 978-0-226-07840-3. £24.50 (paperback). British Journal for the History of Science, 48(2):378–380, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Berry:2017:BRM

[Ber17] Dominic J. Berry. Book review: Maurizio Meloni, Political Biology: Science and Social Values in Human Heredity from Eugenics to Epigenetics. Basingstoke: Palgrave Macmillan, 2016. Pp. xi + 284. ISBN 978-1-137-37771-5. £68.00 (hardback). British Journal for the History of Science, 50(1):164–165, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
BB0FDA63DA0770B09BFB159F6C1BC4D9.

Berry:2018:BRA

[Ber18] Dominic J. Berry. Book review: Annie Tindley and Andrew Wodehouse, Design, Technology and Communication in the British Empire, 1830–1914. London and New York: Palgrave Macmillan, 2016. Pp. 131. ISBN 978-1-1375-9797-7. £37.99 (hardcover). British Journal for the History of Science, 51(3): 527–529, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/annie-
tindley-and-andrew-
wodehouse-design-
technology-and-
communication-in-the-
british-empire-18301914-
london-and-new-york-
palgrave-macmillan-2016-pp-
131-isbn-9781137597977-3799-
hardcover/C474AD7B6C687DE7B382DFD31A5622AD.

Buchwald:2013:NOC

[BF13] Jed Z. Buchwald and Mordechai Feingold. Newton and the origin of civilization. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-15478-3 (hardcover). ix + 528 pp. LCCN QC16.N7 B93 2013.

Buchwald:2014:OHH

[BF14] Jed Z. Buchwald and Robert Fox, editors. The Oxford Handbook of the History of Physics. Oxford Handbooks in Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-969625-X. ix + 945 pp. LCCN Q124.6-127.2.

Bouvier:2011:AAL

[BFGG11] Yves Bouvier, Robert Fox, Pascal Griset, and Anna Guagnini, editors. De l’atelier au laboratoire: recherche et innovation dans l’industie électrique, XIXe–XXe siècles. (French) [From workshop to laboratory: research and innovation in electric industry, 19– 20th centuries], volume 1 of Histoire de l’énergie = History of energy. P. Lang, Bruxelles, Belgium, 2011. ISBN 90-5201-656-9 (paperback). LCCN TK15 .D43 2010.

Boantza:2011:AEP

[BG11] Victor D. Boantza and Ofer Gal. The ‘absolute existence’ of phlogiston: the losing party’s point of view. British Journal for the History of Science, 44(3): 317–342, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Biess:2014:SEA

[BG14] Frank Biess and Daniel M. Gross, editors. Science and emotions after 1945: a transatlantic perspective. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-12634-X (hardcover), 0-226-12648-X (paperback), 0-226-12651-X (e-book). v + 432 pp. LCCN BF531 .S38 2014.

Bennett:2019:LCN

[BH19] Jim Bennett and Rebekah Higgitt. London 1600–1800: communities of natural knowledge and artificial practice. British Journal for the History of Science, 52(2):183–196, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/london-
16001800-communities-of-
natural-knowledge-and-
artificial-practice/
CFD764E398662F8006B631914227E5DE.

Bhattacharya:2011:BRN

[Bha11] Nandini Bhattacharya. Book review: Neelam Kumar (ed.), Women and Science in India: A Reader. New Delhi: Oxford University Press, 2009. Pp. xxx + 351. ISBN 978-0-19-569705-6. £21.99 (hardback). British Journal for the History of Science, 44 (2):305–306, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Biagioli:2012:CCS

[Bia12] Mario Biagioli. From ciphers to confidentiality: secrecy, openness and priority in science. British Journal for the History of Science, 45(2):213–233, June 2012.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bigg:2010:BRR

[Big10] Charlotte Bigg. Book review: Roger Hutchins, British University Observatories 1772–1939. Aldershot and Burlington: Ashgate, 2008. Pp. xxiii + 533. ISBN 978-0-7546-3250-4. £65.00 (hardback). British Journal for the History of Science, 43(4):612, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bigotti:2018:BRC

[Big18]  Fabrizio Bigotti.  Book re-

view: Cristiano Casalini, Aristotle in Coimbra: The Cursus Conimbricensis and the Education at the College of Arts. Translated by Luana Salvarani. Abingdon and New York: Routledge, 2017. Pp. viii + 195. ISBN 978-1-472-46410-1. £110.00 (hardcover). British Journal for the History of Science, 51 (2):329–330, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/cristiano-
casalini-aristotle-in-
coimbra-the-cursus-
conimbricensis-and-the-
education-at-the-college-
of-arts-translated-by-
luana-salvarani-abingdon-
and-new-york-routledge-

2017-pp-viii-195-isbn-9781472464101-
11000-hardcover/D312E0006221FFCFAF78D3734CAEBF7F.

Bil:2015:BRM

[Bil15a] Geoff Bil. Book reviews: Martin Priestman, The Poetry of Erasmus Darwin: Enlightened Spaces, Romantic Times. Farnham: Ashgate, 2013. Pp. xiii + 310. ISBN 978-1-4724-1954-5. £70.00 (hardback). British Journal for the History of Science, 48(3):516–517, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bilgili:2015:ORD

[Bil15b] Alper Bilgili. An Ottoman response to Darwinism: Ismail Fennî on Islam and evolution. British Journal for the History of Science, 48(4):565–582, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bil:2017:BRT

[Bil17] Geoff Bil. Book review: Tom Kennett, The Lord Treasurer of Botany: Sir James Edward Smith and the Linnaean Collections. London: The Linnean Society, 2016. Pp. x + 388. ISBN 978-0-9935510-0-0. £25.00 (paperback). British Journal for the History of Science, 50(3): 549–551, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
9EADD0BD58DD1F02EAC7DA613DC59EC5.

Bil:2018:IVP

[Bil18] Geoff Bil. Imperial vernacular: phytonymy, philology and disciplinarity in the Indo-Pacific, 1800–1900. British Journal for the History of Science, 51(4): 635–658, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/imperial-
vernacular-phytonymy-
philology-and-
disciplinarity-in-the-
indopacific-18001900/
5880ABE58DD69B50F546CFB53F6AA2CF.

Bingham:2012:MBP

[Bin12] Adrian Bingham. ‘the monster’? The British popular press and nuclear culture, 1945–early 1960s. British Journal for the History of Science, 45(4):609–624, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bivins:2010:BRC

[Biv10]  Roberta  Bivins.

Book reviews: Christopher Hamlin, Cholera: The Biography. Oxford University Press, 2009. Pp x + 344. ISBN 978-0-19-954624-

4. £12.99 (hardback). Mark Jackson, Asthma: The Biography. Oxford University Press, 2009. Pp xi + 249. ISBN 978-0-19-923795-1. £12.99 (hardback). Andrew Scull, Hysteria: The Biography. Oxford University Press, 2009. Pp vii + 223. ISBN 978-0-19-956096-7.

£12.99 (hardback). Robert Tattersall, Diabetes: The Biography. Oxford University Press, 2009. Pp x + 229. ISBN 978-0-19-954136-2. £12.99 (hardback). British Journal for the History of Science, 43(3):476–478, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bix:2019:BRP

[Bix19] Amy Sue Bix. Book review: Patricia Fara, A Lab of One’s Own: Science and Suffrage in the First World War. Oxford: Oxford University Press, 2018. Pp. xiii + 319. ISBN 978-0-19-879-498-

1. £18.99 (hardcover). British Journal for the History of Science, 52(1):171–173, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/patricia-
fara-a-lab-of-ones-own-
science-and-suffrage-in-
the-first-world-war-oxford-
oxford-university-press-
2018-pp-xiii-319-isbn-
9780198794981-1899-
hardcover/
AFA8FF547C4AB9C859ACB6B491CCCEF8.

Buchwald:2010:ZPH

[BJ10] Jed Z. Buchwald and Diane Greco Josefowicz. The zodiac of Paris: how an improbable controversy over an ancient Egyptian artifact provoked a modern debate between religion and science. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14576-8 (hardcover). vi + 428 + 8 pp. LCCN Q127.E8 B83 2010.

Bjorkman:2011:BRA

[Bjö11] Maria Björkman. Book review: Alison Bashford and Philippa Levine (eds.), The Oxford Handbook of the History of Eugenics. Oxford: Oxford University Press, 2010. Pp. xxi + 586. ISBN 978-0-19-537314-1. £85.00 (hardback). British Journal for the History of Science, 44(3):472–473, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bjorkman:2014:BRP

[Bjö14] Maria Björkman. Book reviews: Peter Mansfield, The Long Road to Stockholm: The Story of MRI. An Autobiography. Oxford: Oxford University Press, 2013. Pp. x + 241. ISBN 978-0-19-966454-2. £25.00 (hardback). British Journal for the History of Science, 47(2):391–392, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Biagioli:2011:MUI

[BJW11] Mario Biagioli, Peter Jaszi, and Martha Woodmansee, editors. Making and unmaking intellectual property: creative production in legal and cultural perspective. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-90708-2 (hardcover), 0-226-90709-0 (paperback). vi + 466 pp. LCCN KF2979 .M329 2011.

Bashford:2010:OHH

[BL10] Alison Bashford and Philippa Levine, editors. The Oxford handbook of the history of eugenics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-537314-6 (hardcover). xx + 586 pp. LCCN HQ751 .O97 2010.

Blancke:2014:CE

[Bla14] Stefaan Blancke, editor. Creationism in Europe. Medicine, science, and religion in historical context. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1562-3 (hardcover), 1-4214-1563-1 (ebook). xviii + 276 pp. LCCN BS651 .C75 2014.

Blake:2017:BRC

[Bla17] Liza Blake. Book review: Claire Preston, The Poetics of Scientific Investigation in Seventeenth-Century England. Oxford: Oxford University Press, 2015. Pp.

320. ISBN 978-0-19-870480-5. £60.00 (hardback). British Journal for the History of Science, 50(2):351–352, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
82A05C0983E4DF2490E8B6FA36D89CED.

Bleichmar:2012:VEB

[Ble12] Daniela Bleichmar. Visible empire: botanical expeditions and visual culture in the Hispanic Enlightenment. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-05853-0 (hardcover). xii + 286 pp. LCCN QK21.S7 B54 2012.

Bender:2010:CD

[BM10] John B. Bender and Michael Marrinan. The culture of diagram. Stanford University Press, Stanford, CA, USA, 2010. ISBN 0-8047-4504-8 (hardcover), 0-8047-4505-6 (paperback). xvii + 265 pp. LCCN BF367 .B46 2010; BLI m10/.15788.

Bleichmar:2011:CAC

[BM11] Daniela Bleichmar and Peter C. Mancall, editors. Collecting across cultures: material exchanges in the early modern Atlantic world. Early modern Americas. University of Pennsylvania Press, Philadelphia, 2011. ISBN 0-8122-4305-6 (hardcover). xvi + 361 + 8 pp. LCCN GN406 .C65 2011.

Badino:2013:RPH

[BN13] Massimiliano Badino and Jaume Navarro, editors. Research and pedagogy: a history of quantum physics through its textbooks, volume 2 of Max Planck research library for the history and development of knowledge. Studies. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5871-X. 302 pp. LCCN QC173.98. URL

                                        http://www.edition-open-
                                        access.de/studies/2/.

Boantza:2010:BRT

[Boa10] Victor D. Boantza. Book review: Tony Volpe, Science et théologie dans les débats savants de la seconde moitié du XVIIe siècle: La Genèse dans les Philosophical Transactions et le Journal des savants(1665–1710). Preface by Louis Châtellier. Bibliothèque de l’Ecole des Hautes Etudes Sciences Religieuses, 133. Turnhout: Brepols Publishers, 2008. Pp. 467. ISBN 978-2-503-52584-6. EUR65.00 (paperback). British Journal for the History of Science, 43(2):296–297, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boantza:2012:BRJ

[Boa12] Victor D. Boantza. Book review:

John G. McEvoy, The Historiography of the Chemical Revolution: Patterns of Interpretation in the History of Science. London: Pickering & Chatto, 2010. Pp. xiii + 328. ISBN 978-1-84893-030-8. £60.00/$99.00 (hardback). British Journal for the History of Science, 45(3):464–465, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boantza:2015:BRW

[Boa15] Victor Boantza. Book reviews: Is Water H_{2}O? Evidence, Realism and Pluralism. Heidelberg, London and New York: Springer, 2012. Pp. xxi + 316. ISBN 978-94-007-3931-4. £126.00 (hardback). Hasok Chang. British Journal for the History of Science, 48(1):181–183, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boantza:2018:BRJ

[Boa18] Victor D. Boantza. Book review: Jan Golinski, The Experimental Self: Humphry Davy and the Making of a Man of Science. Chicago: The University of Chicago Press, 2016. Pp. 259. ISBN 978-0-226-35136-0. $30.00 (hardback). British Journal for the History of Science, 51(2):307–309, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/jan-
                                        golinski-the-experimental-
                                        self-humphry-davy-and-the-
                                        making-of-a-man-of-science-
                                        chicago-the-university-of-
                                        chicago-press-2016-pp-259-
                                        isbn-9780226351360-3000-
                                        hardback/
                                        2A9EF7BF6BA7998D4CB4AD454BA3D270.

Boddice:2014:PEM

[Bod14] Rob Boddice, editor. Pain and emotion in modern history. Palgrave studies in the history of emotions. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, 2014. ISBN 1-137-37242-7 (hardcover). xi + 284 pp. LCCN RB127 .P3321225 2014.

                                        URL http://www.netread.com/
                                        jcusers2/bk1388/420/
                                        9781137372420/image/
                                        lgcover.9781137372420.jpg.

Boddice:2018:BRS

[Bod18] Rob Boddice. Book review: Susan Broomhall (ed.), Spaces for Feeling: Emotions and Sociabilities in Britain, 1650–1850. Abingdon and New York: Routledge, 2015. Pp. 241. ISBN 978-1-138-82817-

9. £31.99 (paperback). British Journal for the History of Science, 51(1):161–162, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/susan-
broomhall-ed-spaces-for-
feeling-emotions-and-
sociabilities-in-britain-
16501850-abingdon-and-new-
york-routledge-2015-pp-241-
isbn-9781138828179-3199-
paperback/FC988FD6F1F319DAF6E048A0F5431D3C.

Boddice:2019:BRS

[Bod19] Rob Boddice. Book reviews:

Susan Lanzoni, Empathy: A History. New Haven, CT and London: Yale University Press, 2018. Pp. ix + 392. ISBN 978-0-3002-2268-5. $30.00 (hardcover).

— Cathy Gere, Pain, Pleasure, and the Greater Good: From the Panopticon to the Skinner Box and Beyond. Chicago and London: The University of Chicago Press, 2017. Pp. 282. ISBN 978-0-2265-0185-7. $30.00 (cloth). British Journal for the History of Science, 52(3): 534–535, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/susan-
                                        lanzoni-empathy-a-history-
                                        new-haven-ct-and-london-
                                        yale-university-press-2018-
                                        pp-ix-392-isbn-
                                        9780300222685-3000-
                                        hardcover-cathy-gere-pain-
                                        pleasure-and-the-greater-good-
                                        from-the-panopticon-to-the-skinner-
                                        box-and-beyond-chicago-and-london-
                                        the-university-of-chicago-press-
                                        2017-pp-282-isbn-9780226501857-
                                        3000-cloth/F24E62909ECDDBA038B7CCF65D452387.

Boner:2011:CCE

[Bon11] Patrick J. Boner, editor. Change and continuity in early modern cosmology, volume 27 of Archimedes: new studies in the history and philosophy of science and technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-0036-9 (hardcover), 94-007-0037-7 (e-book). ISSN 1385-0180. LCCN QB981 .C5325 2011.

Boon:2011:BRP

[Boo11] Timothy Boon. Book review: Peter J. Bowler, Science for All: The Popularization of Science in Early Twentieth-Century Britain. Chicago and London: University of Chicago Press, 2009. Pp. xi + 339. ISBN 978-0-226-06863-3. £31.00 (hardback). British Journal for the History of Science, 44 (2):304–305, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boon:2012:BRL

[Boo12] Timothy Boon. Book review:

Louis Niebur, Special Sound: The Creation and Legacy of the BBC Radiophonic Workshop. New York and London: Oxford University Press, 2010. Pp. 259. ISBN 978-0-19-536841-3. £17.99 (paperback); 978-0-19-536840-6. £60.00 (hardback). British Journal for the History of Science, 45(2):307–309, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boon:2014:BRM

[Boo14] Tim Boon. Book review: Marcel Chotkowski LaFollette, Science on American Television: A History. Chicago and London: The University of Chicago Press, 2013, Pp. x + 306. ISBN 978-0-226-92199-0. £29.00 (hardback). British Journal for the History of Science, 47(1):189– 191, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boon:2015:SFR

[Boo15a] Tim Boon. Sounding the field:

recent works in sound studies. British Journal for the History of Science, 48(3):493–502, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boon:2015:TSP

[Boo15b] Timothy Boon. The televising of science is a process of television’: establishing Horizon, 1962– 1967. British Journal for the History of Science, 48(1):87–121, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boon:2017:BRT

[Boo17] Tim Boon. Book reviews: Thomas W. Patteson, Instruments for New Music: Sound, Technology, and Modernism. Oakland: University of California Press, 2016. Pp. xii + 236. ISBN 978-0-520-28802-7 £32.95 (paperback). Andrew J. Nelson, The Sound of Innovation: Stanford and the Computer Music Revolution. Cambridge, MA: MIT Press, 2015. Pp. 248. ISBN 978-0-262-02876-9. £29.95 (hardback). British Journal for the History of Science, 50(3):560–562, September 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
ACC85D0EA992654064DEB0180100BDCF.

Bordogna:2008:WJB

[Bor08] Francesca Bordogna. William James at the boundaries: philosophy, science, and the geography of knowledge. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-06652-5 (cloth). x + 382 pp. LCCN B945.J24 B47

2008. URL http://www.loc.
gov/catdir/enhancements/

fy0827/2008014908-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2008014908-d.html;
http://www.loc.gov/catdir/
toc/fy0903/2008014908.html.

Bord:2009:SWM

[Bor09] Joe Bord. Science and Whig manners: science and political style in Britain, c. 1790– 1850. Studies in modern history. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2009. ISBN 0-230-57484-X. ix + 213 pp. LCCN Q127.G4 B67

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0916/2008037191-b.html;
http://www.loc.gov/catdir/
enhancements/fy0916/
2008037191-d.html;
http://www.loc.gov/catdir/
enhancements/fy0916/
2008037191-t.html.

Bourke:2015:BRF

[Bou15] Joanna Bourke. Book reviews:

Frank Biess and Daniel M. Gross (eds.), Science and Emotions after 1945: A Transatlantic Perspective. Chicago and London: The University of Chicago Press, 2014. Pp. vi + 432. ISBN 978-0-226-12648-7. £28.00 (paperback). British Journal for the History of Science, 48(4):708–709, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bowler:2009:SAP

[Bow09] Peter J. Bowler. Science for all: the popularization of science in early Twentieth-Century Britain. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-06863-3 (cloth). x + 339 pp. LCCN Q225.2.G7 B69 2009.

Bowler:2010:BRB

[Bow10]  Peter J. Bowler.  Book re-

view: Bernard Lightman, Evolutionary Naturalism in Victorian Britain: The ‘Darwinians’ and Their Critics. Aldershot: Ashgate, 2009. Pp. xx + 326. ISBN 978-0-7546-5987-7. £70.00 (hardback). British Journal for the History of Science, 43(1):129– 130, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bowler:2011:CP

[Bow11] Peter J. Bowler. Complexity in practice. British Journal for the History of Science, 44(2): 275–280, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bowen:2016:BRJ

[Bow16] Elliott Bowen. Book review: John C. Burnham, Health Care in America: A History. Baltimore: Johns Hopkins University Press, 2015. Pp. xiv + 596. ISBN: 978-1-421-41608-3. $34.95 (paperback). British Journal for the History of Science, 49(4): 504–505, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.

org/core/product/
0DC8F833BC8D734300936DB950D7486D.

Bowler:2017:BRR

[Bow17]  Peter J. Bowler.  Book re-

view: Robert J. Richards and Michael Ruse, Debating Darwin. Chicago: The University of Chicago Press, 2016. Pp. xvi + 299. ISBN 978-0-226-38442-9. £21.00/$30.00 (cloth). British Journal for the History of Science, 50(1):158–159, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
D63276263930A35D63A948B18D01C5DB.

Bowen:2018:BRA

[Bow18] Elliott Bowen. Book review: Anne R. Hanley, Medicine, Knowledge and Venereal Diseases in England, 1886–1916. London and New York: Palgrave Macmillan, 2017. Pp. 318. ISBN 978-3-319-32454-

8. £66.99 (hardcover). British Journal for the History of Science, 51(1):167–168, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/anne-r-
hanley-medicine-knowledge-
and-venereal-diseases-in-
england-18861916-london-
and-new-york-palgrave-
macmillan-2017-pp-318-isbn-
9783319324548-6699-
hardcover/BC44CF8B6559FEBB5A15B3BF0D63382D.

Bradley:2012:CDH

[Bra12] Margaret Bradley. Charles Dupin (1784–1873) and his influence on France: the contributions of a mathematician, educator, engineer, and statesman. Cambria Press, Amherst, NY, USA, 2012. ISBN 1-60497-751-5 (hardcover). xx + 368 pp. LCCN DC255.D86 B73 2012.

Brain:2015:PMP

[Bra15] Robert Michael Brain. The pulse of modernism: physiological aesthetics in Fin-de-Siècle Europe. In Vivo. University of Washington Press, Seattle, WA, USA, 2015. ISBN 0-295-99320-0 (hardcover), 0-295-99321-9 (paperback). xxxii + 348 pp. LCCN BH301.P45 B73 2015.

Brinkman:2010:SJD

[Bri10] Paul D. Brinkman. The second Jurassic dinosaur rush: museums and paleontology in America at the turn of the twentieth century. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-07472-2 (hardcover). xiv + 345 pp. LCCN QE705.U6 B75

                                        2010. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1101/2009046342-b.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1101/
                                        2009046342-d.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1101/
                                        2009046342-t.htm.

Bridger:2015:SWE

[Bri15] Sarah Bridger. Scientists at war: the ethics of Cold War weapons research. Harvard University Press, Cambridge, MA, USA, 2015. ISBN 0-674-73682-6. x + 350 pp. LCCN Q125 .B728 2015.

Brown:2009:HMD

[Bro09] Julie K. Brown. Health and medicine on display: international expositions in the United States, 1876–1904. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-02657-0 (hardcover). xi + 326 pp. LCCN RA438.U5 B76 2009.

Brown:2010:UDS

[Bro10a] B. Ricardo Brown. Until Darwin, science, human variety and the origins of race. Pickering and Chatto, London, UK, 2010. ISBN 1-84893-100-X (hardcover). ix + 199 pp. LCCN GN281 .B765 2010.

Browne:2010:FCA

[Bro10b] Janet Browne. A field comes of age. British Journal for the History of Science, 43(3):469–473, September 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brock:2011:BRR

[Bro11a] W. H. Brock. Book review:

Roy MacLeod, Archibald Liversidge, FRS: Imperial Science under the Southern Cross. Sydney: The Royal Society of New South Wales and Sydney University Press, 2009. Pp. xvii + 637, illustrated. ISBN 978-1-920898-80-9. AU$59.95 (paperback). British Journal for the History of Science, 44(4):610–611, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brown:2011:PMM

[Bro11b] Michael Brown. Performing medicine: medical culture and identity in provincial England, c. 1760–1850. Manchester University Press, Manchester, UK, 2011. ISBN 0-7190-7797-4 (hardback). viii + 254 pp. LCCN R487 .B76 2011.

Brown:2012:KNC

[Bro12] Andrew Brown. Keeper of the nuclear conscience: the life and work of Joseph Rotblat. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-958658-6 (hardcover). xviii + 347 pp. LCCN QC774.R67 B768 2012.

Brooke:2016:BRS

[Bro16] John Hedley Brooke. Book reviews: Stefaan Blancke, Hans Henrik Hjermitslev and Peter C. Kjærgaard (eds.), Creationism in Europe. Baltimore: Johns Hopkins University Press, 2014. Pp. xvii + 276. ISBN 978-1-4214-1562-8. £26.00 (hardback). British Journal for the History of Science, 49(1):131–132, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brush:2015:MCS

[BS15] Stephen G. Brush and Ariel Segal. Making 20th Century science: how theories became knowledge. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-997815-8. xvii + 531 pp. LCCN Q174.8 .B78 2015.

Bertomeu-Sanchez:2016:BRA

[BS16] José Ramón Bertomeu-Sánchez.

Book review: Alison Adam, A History of Forensic Science: British Beginnings in the Twentieth Century. Abingdon: Routledge, 2016. Pp. 236. ISBN 978-0-415-85642-3. £90.00 (hardback). British Journal for the History of Science, 49(4):509– 511, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
EF9D16AAE9D9AA2FD7941676912616F1.

Bennett:2013:CEP

[BT13] J. A. (James A.) Bennett and Sofia Talas, editors. Cabinets of experimental philosophy in eighteenth-century Europe, volume 40(3) of History of science and medicine library, ; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-25296-7 (hardcover), 90-04-25297-5 (e-book). ISSN 1872-0684. xxiii + 253 pp. LCCN Q185 .C27 2013.

Bek-Thomsen:2014:BRG

[BT14] Jakob Bek-Thomsen. Book review: Germano Maifreda, From Oikonomia to Political Economy: Constructing Economic Knowledge from the Renaissance to the Scientific Revolution. Farnham: Ashgate, 2012. Pp. vii + 304. ISBN 978-1-4094-3301-9. £70.00 (hardback). British Journal for the History of Science, 47(1):179– 181, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Berg:2009:FSM

[BTW09] Matthias Berg, Jens Thiel, and Peter Th. Walther, editors. Mit Feder und Schwert: Militär und Wissenschaft, Wissenschaftler und Krieg, volume Band 7 of Geschichte (Franz Steiner Verlag); Wissenschaft, Politik und Gesellschaft. Franz Steiner Verlag, Stuttgart, Germany, 2009. ISBN 3-515-09606-X. 380 pp. LCCN U27 .M58 2009.

Buchanan:2011:BRS

[Buc11] Brenda J. Buchanan. Book review: Simon Werrett, Fireworks: Pyrotechnic Arts and Sciences in European History. Chicago and London: University of Chicago Press, 2010. Pp. vii + 359. ISBN 978-0-226-89377-8. £29.00 (hardback). British Journal for the History of Science, 44(1):124– 125, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bud:2008:PTT

[Bud08] Robert Bud. Penicillin: triumph and tragedy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-954161-2. ix + 330 pp.

                                        LCCN ???? URL http://www.

gbv.de/dms/bowker/toc/
9780199541614.

Bud:2011:BRM

[Bud11]  Robert Bud.  Book review:

Miriam R. Levin, Sophie Forgan, Martina Hessler, Robert H. Kargon and Maurice Low, Urban Modernity: Cultural Innovation in the Second Industrial Revolution. Cambridge, MA and London: MIT Press, 2010. Pp. x + 272. ISBN 978-0-262-01398-7. £22.95 (hardback). British Journal for the History of Science, 44 (2):301–302, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bud:2013:BRC

[Bud13a] Robert Bud.

Book reviews: Charles Dupin (1784–1873) and His Influence on France: The Contributions of a Mathematician, Educator, Engineer, and Statesman. Amherst, NY: Cambria Press, 2012. Pp. xx + 368. ISBN 978-1-60497-751-6. £71.99 (hardback), by Margaret Bradley. British Journal for the History of Science, 46(3): 529–530, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bud:2013:LDM

[Bud13b] Robert Bud. Life, DNA and the model. British Journal for the History of Science, 46(2):311– 334, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bui:2019:BRT

[Bui19] Hanh Bui. Book review: Tomas Macsotay, Cornelis van der Haven and Karel Vanhaesebrouck (eds.), The Hurt(ful) Body: Performing and Beholding Pain, 1600– 1800. Manchester: Manchester University Press, 2017. Pp. xv + 311. ISBN: 978-1-7849-9516-4. £80.00 (hardcover). British Journal for the History of Science, 52 (4):711–712, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/tomas-
macsotay-cornelis-van-der-
haven-and-karel-
vanhaesebrouck-eds-the-
hurtful-body-performing-
and-beholding-pain-
16001800-manchester-
manchester-university-
press-2017-pp-xv-311-isbn-9781784995164-
8000-hardcover/35084F28876169C3432FA59E17B6DCDF.

Buklijas:2014:BRA

[Buk14] Tatjana Buklijas. Book reviews: Annette Lykknes, Donald L. Opitz and Brigitte van Tiggelen (eds.), For Better or for Worse? Collaborative Couples in the Sciences. Heidelberg, New York and London: Birkhäuser, 2012. Pp. xiv + 319. ISBN 978-3-0348-0285-7. EUR 23.53 (hardback). British Journal for the History of Science, 47(4):749–750, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bulstrode:2016:IAD

[Bul16] Jenny Bulstrode. The industrial archaeology of deep time. British Journal for the History of Science, 49(1):1–25, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bunn:2012:TMS

[Bun12] G. C. (Geoffrey C.) Bunn. The truth machine: a social history of the lie detector. Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0530-X (hardcover), 1-4214-0651-9 (ebook). viii + 246 pp. LCCN HV8078 .B86 2012.

Burnett:2011:BRR

[Bur11] Charles Burnett. Book review:

Ruth Glasner, Averroes’ Physics: A Turning Point in Medieval Natural Philosophy. Oxford: Oxford University Press, 2009. Pp. x + 229. ISBN 978-0-19-956773-7. £37.50 (hardback). British Journal for the History of Science, 44 (1):121–122, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Burkett:2012:CND

[Bur12] Jodi Burkett. The campaign for nuclear disarmament and changing attitudes towards the Earth in the nuclear age. British Journal for the History of Science, 45 (4):625–639, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Burnett:2013:BRM

[Bur13] Charles Burnett. Book reviews: Muzaffar Iqbal (ed.), New Perspectives on the History of Islamic Science. Islam and Science: Historic and Contemporary Perspectives, vol. 3. Farnham: Ashgate, 2012. Pp. xxiv + 546. ISBN 978-0-75462-2914-6. £140.00 (hardback). Muzaffar Iqbal (ed.), Studies in the Making of Islamic Science: Knowledge in Motion. Islam and Science: Historic and Contemporary Perspectives, vol. 4. Farnham: Ashgate, 2012. Pp. xxiii + 552. ISBN 978-0-75462-2916-0. £140.00 (hardback). British Journal for the History of Science, 46(2):341–342, June 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Burns:2016:SRG

[Bur16] William E. Burns. The scientific revolution in global perspective. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-998933-8 (paperback). xv + 198 pp. LCCN Q125 .B9497 2016.

Busch:2010:SCC

[Bus10] John Laurence Busch. STEAM COFFIN: Captain Moses Rogers and the steamship Savannah break the barrier. Hodos Historia, New Canaan, CT, USA, 2010. ISBN 1-893616-00-2. 726 pages pp. LCCN VK140.R65 B87

2010. URL http://www.loc.
gov/catdir/enhancements/

fy1106/2009942158-b.html;
http://www.loc.gov/catdir/
enhancements/fy1106/
2009942158-d.html.

Bussotti:2018:BRM

[Bus18] Paolo Bussotti. Book review:

Michel Blay, Critique de l’histoire des sciences. Paris: CNRS Editions, 2017. Pp. 302. ISBN 978-2-271-09184-0. EUR22.00 (paperback). British Journal for the History of Science, 51(1):153– 155, 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/michel-blay-
critique-de-lhistoire-des-
sciences-paris-cnrs-
editions-2017-pp-302-isbn-
9782271091840-2200-
paperback/
B309638B65FA55BC05B393A9A734842E.

Buxton:2018:HDT

[Bux18] Hilary Buxton. Health by design:

teaching cleanliness and assembling hygiene at the nineteenthcentury sanitation museum. British Journal for the History of Science, 51(3):457–485, September 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/health-by-design-
teaching-cleanliness-and-
assembling-hygiene-at-the-
nineteenthcentury-

                                        sanitation-museum/
                                        406101987A597D794EDBB3CAFE916DD2.

Borchardt:2013:AZR

[BW13] Ludwig Borchardt and Daniela Wuensch. Die altägyptische Zeitmessung: Reprint der Ausgabe von 1920 mit einer Einleitung von Daniela Wuensch: Was die alten Ägypter über Uhren und Zeitmessung wussten. (German) [Ancient Egyptian time measurement: Reprint of the Edition of 1920 with an introduction by Daniela Wuensch: About watches and the ancient Egyptians’ knowledge of timekeeping]. Termessos, Göttingen, Germany, 2013. ISBN 3-938016-14-0. 192 + 25 + 20 + 2 pp. LCCN ????

Bresalier:2014:SLO

[BW14] Michael Bresalier and Michael Worboys. ‘Saving the lives of our dogs’: the development of canine distemper vaccine in interwar Britain. British Journal for the History of Science, 47(2):305– 334, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Byrne:2013:GRN

[Byr13] Angela Byrne. Geographies of the romantic north: science, antiquarianism, and travel, 1790– 1830. Palgrave studies in cultural and intellectual history. Palgrave Macmillan, New York, NY, USA, 2013. ISBN 1-137-31131-2. xiv + 265 pp. LCCN G630.B7 B87 2013.

Ben-Zaken:2011:BRT

[BZ11a] Avner Ben-Zaken. Book review:

Toby E. Huff, Intellectual Curiosity and the Scientific Revolution: A Global Perspective. Cambridge: Cambridge University Press, 2011. Pp. xiii + 354. ISBN 978-0-521-17052-9. £60.00 (hardback). British Journal for the History of Science, 44(4): 585–587, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ben-Zaken:2011:RHI

[BZ11b] Avner Ben-Zaken. Reading Hayy Ibn-Yaqzan: a cross-cultural history of autodidacticism. History of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 0-8018-9739-4 (hardcover), 0-8018-9972-9. xiii + 191 pp. LCCN B753.I54 B46X

2011. URL http://muse.jhu.
edu/books/9780801899720/.

Campbell:2012:QAA

[C^{+}12] Steven C. Campbell et al., editors. 100 questions and answers about kidney cancer. Jones and Bartlett Learning, Sudbury, MA, USA, updated. edition, 2012. ISBN 1-4496-2192-9. ???? pp. LCCN RC280.K5 A35 2012.

Caleb:2012:BRC

[Cal12] Amanda Mordavsky Caleb. Book review: Charlotte Sleigh, Literature and Science. Outlining Literature Series. Basingstoke and New York: Palgrave Macmillan, 2011. Pp. 232. ISBN 978-0-230-21817-8. £16.99/$27.00 (paperback). British Journal for the History of Science, 45(1): 145–146, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Sle11b].

Calver:2013:BRS

[Cal13] Neil Calver. Book reviews: Simone Turchetti, The Pontecorvo Affair: A Cold War Defection and Nuclear Physics. Chicago and London: The University of Chicago Press, 2012. Pp. iv + 292. ISBN 978-0-226-81664-7. £29.00 (hardback). British Journal for the History of Science, 46 (1):173–174, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Camilleri:2009:HIQ

[Cam09] Kristian Camilleri. Heisenberg and the interpretation of quantum mechanics: the physicist as philosopher. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88484-5 (hardcover). xii + 199 pp. LCCN QC16.H395 .C34 2009. URL http://www.

loc.gov/catdir/toc/
ecip0825/2008034113.html.

Campe:2012:GPL

[Cam12] Rüdiger Campe. The game of probability: literature and calculation from Pascal to Kleist. Cultural memory in the present. Stanford University Press, Stanford, CA, USA, 2012. ISBN 0-8047-6864-1 (hardcover), 0-8047-6865-X (paperback), 0-8047-8466-3 (ebook). viii + 486 pp. LCCN PN744 .C3613 2012. Translated by Ellwood H. Wiggins, Jr.

Cantor:2011:RGE

[Can11] Geoffrey N. Cantor. Religion and the Great Exhibition of 1851. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-959667-0 (hardcover). xi + 226 pp. LCCN T690.B1 C367 2011.

Cantor:2015:BRD

[Can15] Geoffrey Cantor. Book reviews:

Daniel Pick, The Pursuit of the Nazi Mind: Hitler, Hess and the Analysts. Oxford: Oxford University Press, 2014. Pp. xii + 357. ISBN 978-0-19-967851-8. £10.99 (paperback). British Journal for the History of Science, 48(3): 528–529, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cantor:2016:BRP

[Can16] Geoffrey Cantor. Book review:

Peter Harrison, The Territories of Science and Religion. Chicago and London: The University of Chicago Press, 2015. Pp. 320. ISBN 978-0-226-18448-7. $30.00 (hardback). British Journal for the History of Science, 49(4): 527–528, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
544B607FF401B8F8014008D46D925AC3.

Cantor:2019:BRJa

[Can19a] Geoffrey Cantor. Book review: Jaume Navarro (ed.), Ether and Modernity: The Recalcitrance of an Epistemic Object in the Early Twentieth Century. Oxford: Oxford University Press, 2018. Pp. 272. ISBN 978-0-1987-9725-8. £65.00 (hardcover). British Journal for the History of Science, 52(1):173– 174, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/jaume-navarro-ed-
ether-and-modernity-the-
recalcitrance-of-an-
epistemic-object-in-the-
early-twentieth-century-
oxford-oxford-university-
press-2018-pp-272-isbn-
9780198797258-6500-hardcover/
D7D6686070F16EFA522632EA48EFF748.

Cantor:2019:BRJ

[Can19b] Geoffrey Cantor. Book review: Jeff Hardin, Ronald L. Numbers and Ronald A. Binzley (eds.), The Warfare between Science and Religion: The Idea That Wouldn’t Die. Baltimore: Johns Hopkins University Press, 2018. Pp. 355. ISBN 978-1-4214-2618-

1. $39.95 (paperback). British Journal for the History of Science, 52(2):375–376, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

journal-for-the-history-of-
science/article/jeff-
hardin-ronald-l-numbers-
and-ronald-a-binzley-eds-
the-warfare-between-
science-and-religion-the-
idea-that-wouldnt-die-
baltimore-johns-hopkins-
university-press-2018-pp-
355-isbn-9781421426181-3995-
paperback/FD96842D9EF8CA9EB2CB792E757C387B.

Capelotti:2013:SCF

[Cap13] P. J. (Peter Joseph) Capelotti.

Shipwreck at Cape Flora: the expeditions of Benjamin Leigh Smith, England’s forgotten Arctic explorer, volume 16 of Northern lights series. University of Calgary Press, Calgary, AB, Canada, 2013. ISBN 1-55238-705-4 (paperback), 1-55238-713-5 (ebook), 1-55238-712-7 (PDF), 1-299-31685-9, 1-55238-714-3. ISSN 1701-0004. xxix + 269 pp. LCCN G875.S65 C36 2013.

Capocci:2014:BRR

[Cap14] Mauro Capocci. Book reviews:

Renato G. Mazzolini and Hans-Jörg Rheinberger (eds.), Differing Routes to Stem Cell Research: Germany and Italy. Bologna and Berlin: Il Mulino/Duncker and Humblot, 2012. Pp. 271. ISBN 978-3-428-13849-4. EUR 22.00 (paperback). British Journal for the History of Science, 47(4): 759–760, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Capocci:2015:BRA

[Cap15]  Mauro Capocci.  Book re-

views: Angela N.H. Creager, Life Atomic: A History of Radioisotopes in Science and Medicine. Chicago: The University of Chicago Press, 2013. Pp. xvi + 489. ISBN 978-0-226-01780-8. £31.50 (hardback). British Journal for the History of Science, 48 (4):712–713, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Carter:2009:MFG

[Car09] Christopher Ray Carter. Magnetic fever: global imperialism and empiricism in the nineteenth century, volume 99 (part 4) of Transactions of the American Philosophical Society held at Philadelphia for promoting useful knowledge. American Philosophical Society, Philadelphia, PA, USA, 2009. ISBN 1-60618-994-8. xxvi + 168 pp. LCCN QC807.6.G7 C37 2009.

Carson:2010:HAA

[Car10] Cathryn Carson. Heisenberg in the atomic age: science and the public sphere. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 2010. ISBN 0-521-82170-3 (hardcover). xvi + 541 + 8 pp. LCCN QC16.H45 C37 2010.

Carroy:2011:BRJ

[Car11] Jacqueline Carroy. Book review: Jan Goldstein, Hysteria Complicated by Ecstasy: The Case of Nanette Leroux. Princeton and Oxford: Princeton University Press, 2010. Pp. xi + 246.

ISBN 978-0-691-01186-8. £20.95 (hardback). British Journal for the History of Science, 44(1): 131–132, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Carruthers:2012:BRJ

[Car12] William Carruthers. Book reviews: Jeffrey Abt, American Egyptologist: The Life of James Henry Breasted and the Creation of His Oriental Institute. Chicago and London: The University of Chicago Press, 2011. Pp. xix + 510. ISBN 978-0-226-00110-4. £29.00 (hardback). British Journal for the History of Science, 45 (4):691–692, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Carolino:2016:BRM

[Car16a] Luís Miguel Carolino. Book reviews: Miguel A. Granada, Adam Mosley and Nicholas Jardine, Christoph Rothmann’s Discourse on the Comet of 1585. An Edition and Translation with Accompanying Essays. Leiden and Boston: Brill, 2014. Pp. xx + 379. ISBN 978-90-04-26023-4. £115.00 (hardback). Víctor Navarro Brotóns, Disciplinas, Saberes y Prácticas: Filosofia Natural, Matemáticas y Astronomía en la Sociedad Española de la Época Moderna. València: Universitat de València, 2014. Pp. 496. ISBN 978-84-370-9446-5. £26.75 (paperback). British Journal for the History of Science, 49(2):292–293, June 2016.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Carpenter:2016:BRS

[Car16b] Oliver Carpenter. Book review: Simone Fari, Victorian Telegraphy before Nationalization. Basingstoke: Palgrave Macmillan, 2015. Pp. 248. ISBN 978-1-137-40651-4. £63.00 (hardback). British Journal for the History of Science, 49 (4):496–497, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
982688CB4C67016C12A52DA58A2C709A.

Carolino:2018:BRP

[Car18] Luís Miguel Carolino. Book reviews: Pietro Daniel Omodeo (ed.), in collaboration with Karin Friedrich, Duncan Liddel (1561– 1613): Networks of Polymathy and the Northern European Renaissance. Leiden and Boston: Brill, 2016. Pp. xii + 322. ISBN 978-9-0043-1065-0, EUR135.00, $175.00 (hardback). — Kuni Sakamoto, Julius Caesar Scaliger, Renaissance Reformer of Aristotelianism: A Study of His Exotericae Exercitationes. Leiden and Boston: Brill, 2016. Pp. viii + 213. ISBN 978-9-0043-1009-4, EUR104.00, $135.00 (paperback). British Journal for the History of Science, 51(3): 515–518, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

org/core/journals/british-
journal-for-the-history-of-
science/article/pietro-
daniel-omodeo-ed-in-
collaboration-with-karin-
friedrich-duncan-liddel-
15611613-networks-of-
polymathy-and-the-northern-
european-renaissance-
leiden-and-boston-brill-2016-
pp-xii-322-isbn-9789004310650-
13500-17500-hardback-kuni-sakamoto-
julius-caesar-scaliger-renaissance-
reformer-of-aristotelianism-
a-study-of-his-exotericae-exercitationes-
leiden-and-boston-brill-2016-
pp-viii-213-isbn-9789004310094-
10400-13500-paperback/B6D4C3B41F143AF1FA95655BD1462C23.

Cassidy:2011:SHP

[Cas11] David C. Cassidy. A short history of physics in the American century. New histories of science, technology, and medicine. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-04936-5. 211 pp. LCCN QC9.U5 C36 2011.

Cozzoli:2011:MBT

[CC11a] Daniele Cozzoli and Mauro Capocci. Making biomedicine in Twentieth-Century Italy: Domenico Marotta (1886– 1974) and the Italian Higher Institute of Health. British Journal for the History of Science, 44 (4):549–574, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Croucher:2011:MJT

[CC11b] John S. Croucher and Rosalind F.

Croucher. Mrs Janet Taylor’s ‘Mariner’s Calculator’: assessment and reassessment. British Journal for the History of Science, 44(4):493–507, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cohen-Cole:2014:OMC

[CC14] Jamie Nace Cohen-Cole. The open mind: Cold War politics and the sciences of human nature. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-09216-X (hardcover), 0-226-09233-X (e-book). 397 pp. LCCN BF39.3 .C645 2014.

Chabas:2012:SEA

[CG12] José Chabás and Bernard R. Goldstein. A Survey of European Astronomical Tables in the Late Middle Ages. Time, Astronomy, and Calendars. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-23058-0 (hardcover), 90-04-23059-9 (e-book). xix + 250 pp. LCCN QB23 2012.

Catala-Gorgues:2013:CLB

[CGC13a] Jesús Catalá-Gorgues and Ana Carneiro. Corrigendum: Like birds of a feather: the cultural origins of Iberian geological cooperation and the European Geological Map of 1896. British Journal for the History of Science, 46(1): 71, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Catala-Gorgues:2013:LBF

[CGC13b] Jesús Catalá-Gorgues and Ana Carneiro. Like birds of a feather: the cultural origins of Iberian geological cooperation and the European geological map of 1896. British Journal for the History of Science, 46(1):39– 70, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Chalmers:2009:SAP

[Cha09] A. F. (Alan Francis) Chalmers.

The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms, volume 279 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346. xii + 287 pp. LCCN QD461

.C375 2009. URL https://link.
springer.com/book/10.1007/
978-90-481-2362-9.

Charnley:2010:BRP

[Cha10]  Berris Charnley.  Book re-

views: Philip J. Pauly, Fruits and Plains: The Horticultural Transformation of America. Cambridge, MA and London: Harvard University Press, 2008. Pp. xi + 336, ISBN 978-0-674-02663-6. £29.95 (hardback). Alan L. Olmstead and Paul W. Rhode, Creating Abundance: Biological Innovation and American Agricultural Development. Cambridge: Cambridge University Press, 2009. Pp. xii + 467. ISBN 978-0-521-67387-7. £15.99 (paperback). British Journal for the History of Science, 43(2):308–309, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Chakrabarti:2012:BBI

[Cha12] Pratik Chakrabarti. Bacteriology in British India: laboratory medicine and the tropics, volume 22 of Rochester studies in medical history. University of Rochester Press, Rochester, NY, 2012. ISBN 1-58046-408-4 (hardcover). ISSN 1526-2715. x + 304 pp. LCCN QR46 .C43 2012.

Chaney:2014:BRR

[Cha14] Sarah Chaney. Book review: Robert G. W. Kirk and Neil Pemberton, Leech. London: Reaktion Books, 2013. Pp. 208. ISBN 978-1-78023-033-7. £9.99 (paperback). British Journal for the History of Science, 47(1):173– 175, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Chemla:2012:HMP

[Che12] Karine Chemla, editor. The history of mathematical proof in ancient traditions. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-01221-X. xv + 596 pp. LCCN QA22 .H588

                                        2012. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1305/2012472526-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1305/
                                        2012472526-d.html;
                                        http://www.loc.gov/catdir/

enhancements/fy1305/
2012472526-t.html.

Chimisso:2008:WHM

[Chi08] Cristina Chimisso. Writing the history of the mind: philosophy and science in France, 1900 to 1960s. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hants, England and Farnham, England, 2008. ISBN 0-7546-5705-1. vi + 209 pp. LCCN BD418.3 .C46 2008. URL

http://www.loc.gov/catdir/
toc/ecip083/2007044797.html.

Chiang:2012:BRT

[Chi12] Howard Chiang. Book reviews:

Thomas S. Mullaney, Coming to Terms with the Nation: Ethnic Classification in Modern China. Berkeley, Los Angeles and London: University of California Press, 2010. Pp. xiv + 232. ISBN 978-0-520-26278-2. £41.95 (hardback). British Journal for the History of Science, 45(4):702– 703, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Christianson:2010:BRN

[Chr10] J. R. Christianson. Book reviews:

Nicholas Jardine and Alain-Philippe Segonds, La Guerre des Astronomes: La Querelle au sujet de l’origine du système géohéliocentrique à la fin du XVIe siècle. Volume 1: Introduction. Paris: Société d’Edition Les Belles Lettres, 2008. Pp. xxiv + 290. ISBN 978-2-251-34513-0. EUR40.00 (paperback). Nicholas Jardine and Alain-Philippe Segonds, La Guerre des Astronomes: La Querelle au sujet de l’origine du système géo-héliocentrique à la fin du XVIe siècle. Volume 2/1: LeContra Ursumde Jean Kepler, Introduction et textes préparatoires and Volume 2/2: LeContra Ursumde Jean Kepler, Édition critique, traduction et notes. Paris: Société d’Edition Les Belles Lettres, 2008. Pp. xx + 222 and 223-634. ISBN 978-2-251-34512-3. EUR85.00 (paperback). British Journal for the History of Science, 43(3):483–485, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Christensen:2013:HCO

[Chr13] Dan Ch. Christensen. Hans Christian Ørsted: reading nature’s mind. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966926-0 (hardcover), 0-19-164711-X (e-book). xix + 743 pp. LCCN QC16.O66 C56 2013.

Cantor:2010:ODC

[CJ10] Geoffrey Cantor and Frank James. Obituary: David Charles Gooding (21 November 1947–13 December 2009). British Journal for the History of Science, 43(3): 459–467, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Clark:2009:BVI

[Cla09] J. F. M. (John F. M.) Clark. Bugs and the Victorians: insects, science, and experts in Nineteenth-Century England. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-15091-1. xiii + 322 pp. LCCN QL482.G7 C58 2009.

Claessens:2014:SCN

[Cla14] Guy Claessens. A sixteenthcentury Neoplatonic synthesis: Francesco Piccolomini’s theory of mathematics and imagination in the Academicae contemplationes. British Journal for the History of Science, 47(3):421–431, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Clericuzio:2014:BRS

[Cle14] Antonio Clericuzio. Book reviews: Sean Cocco, Watching Vesuvius: A History of Science and Culture in Early Modern Italy. Chicago and London: The University of Chicago Press, 2013. Pp. xi + 332. ISBN 978-0-226-92371-0. £29.00 (hardback). British Journal for the History of Science, 47(3):572–573, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Clericuzio:2015:BRK

[Cle15] Antonio Clericuzio. Book reviews: Klaas van Berkel, Isaac Beeckman on Matter and Motion: Mechanical Philosophy in the Making. Baltimore: Johns Hopkins University Press, 2013. Pp. ix + 265. ISBN 978-142140936-8. £21.00 (paperback). British Journal for the History of Science, 48(2):359–360, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Clericuzio:2016:BRO

[Cle16] Antonio Clericuzio. Book review: Ohad Nachtomy and Justin E.H. Smith (eds.), The Life Sciences in Early Modern Philosophy. Oxford: Oxford University Press, 2014. Pp. 272. ISBN 978-0-19-998731-3. £51.00 (hardback). British Journal for the History of Science, 49(3):482– 484, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        cambridge.org/core/product/
                                        6AE305B8A0090FA756B1E845D0D21C5B.

Clericuzio:2017:BRG

[Cle17] Antonio Clericuzio. Book review: Georgiana D. Hedesan, An Alchemical Quest for Universal Knowledge: ‘The Christian Philosophy’ of Jan Baptist Van Helmont (1579–1644). London: Routledge, 2016. Pp. 244 + xix. ISBN 978-1472469168. £95.00 (hardback). British Journal for the History of Science, 50(2):348–349, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        A809E846B6A7E8776F029595499C15A7.

Clericuzio:2018:BRS

[Cle18] Antonio Clericuzio. Book review: Sabrina Minuzzi, Sul filo dei segreti: Farmacopea, libri e pratiche terapeutiche a Venezia in età moderna. Milan: Edizioni Unicopli, 2016. Pp. 349. ISBN 978-8-8400-1869-0. EUR25.00 (paperback). British Journal for the History of Science, 51(3):520– 521, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/sabrina-
minuzzi-sul-filo-dei-
segreti-farmacopea-libri-e-
pratiche-terapeutiche-a-
venezia-in-eta-moderna-
milan-edizioni-unicopli-
2016-pp-349-isbn-
9788840018690-2500-
paperback/CED2E0E35730739320768949FA942C2F.

Cocchi:2008:FRS

[CMC08] Igino Cocchi, Giuseppe Meneghini, and Pietro Corsi. Fossils and reputations: a scientific correspondence: Pisa, Paris, London, 1853–1857. Pisa University Press, Pisa, Italy, 2008. ISBN 88-8492-564-9. 411 pp. LCCN QE705 COC 2008.

Cocco:2013:WVH

[Coc13] Sean Cocco. Watching Vesuvius:

a history of science and culture in early modern Italy. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92371-1 (hardcover), 0-226-92373-8 (ebook). xi + 322 pp. LCCN QE523.V5 C63 2013.

Coen:2010:BRL

[Coe10] Deborah R. Coen. Book reviews: Laura Dassow Walls, The Passage to Cosmos: Alexander von Humboldt and the Shaping of America. Chicago and London: University of Chicago Press, 2009. Pp. xv + 404. ISBN 978-0-226-87182-0. £24.00 (hardback). H. Walter Lack, Alexander von Humboldt and the Botanical Exploration of the Americas. London: Prestel, 2009. Pp. 278. ISBN 978-3-7913-4142-2. £125.00 (hardback). British Journal for the History of Science, 43(2):302– 303, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Coen:2013:EOD

[Coe13] Deborah R. Coen. The earthquake observers: disaster science from Lisbon to Richter. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-11181-4 (paperback), 0-226-11183-0 (ebook). 348 pp. LCCN QE539 .C64 2013.

Cohen:2011:MZT

[Coh11] Claudine Cohen. La méthode de Zadig: la trace, le fossile, la preuve. Science ouverte. Éditions du Seuil, Paris, France, 2011. ISBN 2-02-040298-X. 341 + 9 + 16 pp. LCCN QE704 .C64 2011.

Cohen:2015:RMS

[Coh15] H. Floris Cohen. The rise of modern science explained: a comparative history. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-12006-3 (hardcover), 1-107-54560-9 (paperback). ???? pp. LCCN Q125.2 .C6413 2015.

Cohen:2018:SBL

[Coh18] H. Floris Cohen. Stock and bulk in the latest Newton scholarship. British Journal for the History of Science, 51(4): 687–701, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/stock-and-
bulk-in-the-latest-newton-
scholarship/
8E0280DA180DD0F63ACA1E2190CF34BB.

Colonna:2013:BRS

[Col13]  Federica  Turriziani

Colonna. Book reviews: Seneca e le Scienze Naturali. Florence: Leo S. Olschki Editore, 2012. Pp vi + 273. ISBN-13 978-88-222-6198-5. EUR 29.00 (paperback), by Marco Beretta, Francesco Citti and Lucia Pasetti (eds.). British Journal for the History of Science, 46(3):519–520, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Collins:2014:WAS

[Col14] Harry M. Collins. Are we all scientific experts now? Polity Press, Cambridge, UK, 2014. ISBN 0-7456-8203-0 (hardcover), 0-7456-8204-9 (paperback), 0-7456-8272-3 (e-book). vi + 144 pp. LCCN Q175.5.

Colonna:2016:HED

[Col16]  Federica  Turriziani

Colonna. Heredity, evolution and development in their (epistemic) environment at the turn of the nineteenth century. British Journal for the History of Science, 49 (1):107–113, March 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Comfort:2012:SHP

[Com12] Nathaniel C. Comfort. The science of human perfection: how genes became the heart of American medicine. Yale University Press, New Haven, CT, USA, 2012. ISBN 0-300-16991-4 (hardcover), 0-300-18887-0 (e-book). xvii + 316 pp. LCCN RB155 .C66 2012.

Conlin:2015:BRP

[Con15] Jonathan Conlin. Book reviews: Piers J. Hale, Political Descent: Malthus, Mutualism, and the Politics of Evolution in Victorian England. Chicago and London: The University of Chicago Press, 2014. Pp. 442. ISBN 978-0-226-10849-0. $45.00/£31.50 (hardback). British Journal for the History of Science, 48(4):702– 704, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cook:2011:BRF

[Coo11] Harold J. Cook. Book review: Florike Egmond, The World of Carolus Clusius: Natural History in the Making 1550–1610. London: Pickering and Chatto, 2010.

Pp. xiv + 292. ISBN 978-1-84893-008-7. £60.00 (hardback). British Journal for the History of Science, 44(2):286–287, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cook:2012:JJR

[Coo12] Alexandra Cook. Jean-Jacques Rousseau and botany: the salutary science, volume 2012/12 of SVEC. Voltaire Foundation, Oxford, UK, 2012. ISBN 0-7294-1055-2 (paperback). ISSN 0435-2866. xxi + 436 pp. LCCN ????

Cook:2014:BRM

[Coo14] Alexandra Cook. Book reviews:

Michael Keevak, Becoming Yellow: A Short History of Racial Thinking. Princeton: Princeton University Press, 2011. Pp. ix + 219. ISBN 978-0-691-14301-5. £24.95 (hardback). British Journal for the History of Science, 47(2):378–379, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cordle:2012:PPB

[Cor12] Daniel Cordle. Protect/protest:

British nuclear fiction of the 1980s. British Journal for the History of Science, 45(4):653– 669, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Courtney:2017:VDL

[Cou17] Stephen Courtney. ‘A very diadem of light’: exhibitions in Victorian London, the Parliamentary light and the shaping of the Trinity House lighthouses. British Journal for the History of Science, 50(2):249–265, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
43C39FCF31D91AB312642C0F43BFA348.

Cowan:2008:HHC

[Cow08] Ruth Schwartz Cowan. Heredity and hope: the case for genetic screening. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02424-

9. 292 pp. LCCN RB155.65 .C69 2008. URL http://www.

loc.gov/catdir/toc/fy0805/
2007039201.html.

Cowan:2010:BRG

[Cow10] Ruth Schwartz Cowan. Book review: Graeme Gooday, Domesticating Electricity: Technology, Uncertainty and Gender, 1880– 1914. London: Pickering and Chatto, 2008. Pp. ix + 292. ISBN 978-1-85196-975-3. £60.00 (hardback). British Journal for the History of Science, 43(3): 498–500, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cowles:2013:VEA

[Cow13] Henry M. Cowles. A Victorian extinction: Alfred Newton and the evolution of animal protection. British Journal for the History of Science, 46(4):695–714, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cowie:2014:EAN

[Cow14] Helen (Helen Louise) Cowie. Exhibiting animals in nineteenthcentury Britain: empathy, education, entertainment. Palgrave Macmillan, Basingstoke, Hampshire, England, 2014. ISBN 1-137-38443-3. ix + 256 pp. LCCN QL76.5.G7 C68 2014.

Cozzoli:2015:BRA

[Coz15] Daniele Cozzoli. Book reviews:

Abena Dove Osseo-Asare, Bitter Roots: The Search for Healing Plants in Africa. Chicago and London: The University of Chicago Press, 2014. Pp. viii + 300. ISBN 978-0-226-08602-6. £24.50 (paperback). British Journal for the History of Science, 48(2):381–382, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cozzoli:2016:BRS

[Coz16] Daniele Cozzoli. Book reviews:

Scott H. Podolsky, The Antibiotic Era: Reform, Resistance, and the Pursuit of a Rational Therapeutics. Baltimore: Johns Hopkins University Press, 2014. Pp. 328. ISBN 978-1-4214-1593-2. $34.95 (hardback). British Journal for the History of Science, 49(2):317– 318, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cozzoli:2017:BRA

[Coz17] Daniele Cozzoli. Book review:

Amparo Gómez, Antonio Fco. Canales and Brian Balmer (eds.), Science Policies and Twentieth-Century Dictatorships: Spain, Italy and Argentina. Farnham: Ashgate, 2015. Pp. 244. ISBN 978-1-4724-2232-3. £63.00 (hardback). British Journal for the History of Science, 50(2): 371–373, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
88DD519CF81D635149D2AC3239561724.

Cantor:2014:SSA

[CR14] David Cantor and Edmund Ramsden, editors. Stress, shock, and adaptation in the twentieth century. Rochester studies in medical history. University of Rochester Press, Rochester, NY, USA, 2014. ISBN 1-58046-476-9 (hardcover). ISSN 1526-2715. vi + 367 pp. LCCN RC49 .S77 2014.

Cain:2009:DDI

[CRB09] Joe Cain, Michael Ruse, and Frederick Burkhardt, editors. Descended from Darwin: insights into the history of evolutionary studies, 1900–1970, volume 99, part 1 of Transactions of the American Philosophical Society Held at Philadelphia for Promoting Useful Knowledge. American Philosophical Society, Philadelphia, PA, USA, 2009. ISBN 1-60618-991-3. LCCN QH361 .D47 2009.

Creese:2010:MAG

[Cre10] David E. (David Evan) Creese. The monochord in ancient Greek harmonic science. Cambridge classical studies. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-84324-3 (hardcover). xvi + 409 pp. LCCN ML3809 .C87 2010. URL http:

//assets.cambridge.org/
97805218/43249/cover/
9780521843249.jp.

Creager:2013:LAH

[Cre13] Angela N. H. Creager. Life atomic: a history of radioisotopes in science and medicine. Synthesis: a series in the history of chemistry, broadly construed. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-01780-X (hardcover), 0-226-01794-X (e-book). xvi + 489 pp. LCCN QC798.A1 C743 2013.

Crook:2012:GES

[Cro12] D. P. (David Paul) Crook.

Grafton Elliot Smith, Egyptology and the diffusion of culture: a biographical perspective. Sussex Academic Press, Brighton, UK, 2012. ISBN 1-84519-481-0 (hardcover). vii + 160 pp. LCCN GN21.S63 C76 2012.

Csiszar:2017:HLB

[Csi17] Alex Csiszar. How lives became lists and scientific papers became data: cataloguing authorship during the nineteenth century. British Journal for the History of Science, 50(1):23–60, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
2DA1C4FE6547421E1515A1925D804EB4.

Cullen:2011:BRJ

[Cul11] Clara Cullen. Book review: Juliana Adelman, Communities of Science in Nineteenth-Century Ireland. London: Pickering and Chatto, 2009. Pp. xi + 221. ISBN 978-1-85196-653-0. £60.00 (hardback). British Journal for the History of Science, 44(1): 137–138, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cullen:2013:BRI

[Cul13] Clara Cullen. Book reviews: Irish Catholicism and Science: From ‘Godless Colleges’ to the ‘Celtic Tiger’. Cork: Cork University Press, 2012. Pp. 343. ISBN 979-185918-497-4. EUR 39.00, £35.00 (hardback), by Don O’Leary. British Journal for the History of Science, 46(3):532–534, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cunningham:2010:AAE

[Cun10a] Andrew Cunningham. The anatomist anatomis’d: an experimental discipline in Enlightenment Europe. History of medicine in context. Ashgate, Aldershot, Hants, England and Farnham, England, 2010. ISBN 0-7546-6338-8 (hardback). xxii + 443 pp. LCCN QM11 .C863 2010; QS11.GA1 2010.

Cunningham:2010:BRC

[Cun10b] Andrew Cunningham. Book review: Charles Webster, Paracelsus: Medicine, Magic and Mission at the End of Time. New Haven and London: Yale University Press, 2008. Pp. xiv + 326. ISBN 978-0-300-13911-2. £30.00 (hardback). British Journal for the History of Science, 43(2):292– 295, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cuomo:2012:BRR

[Cuo12] Serafina Cuomo. Book review:

Reviel Netz, William Noel, Natalie Tchernetska and Nigel Wilson (eds.), The Archimedes Palimpsest. 2 volume set. Cambridge: Cambridge University Press for the Walters Art Museum, 2011. Pp. viii + 342; x + 344. ISBN 978-1-107-01457-2. £150/$215.00 (hardback). British Journal for the History of Science, 45(4): 679–681, December 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A86C8E1B5C97A5CB5170049E3E8F4D14.

Cuomo:2013:BRH

[Cuo13]  Serafina Cuomo.  Book re-

views: The History of Mathematical Proof in Ancient Traditions. Cambridge: Cambridge University Press, 2012. Pp. xv + 596. ISBN-13 978-1-107-01221-

9. £100.00 (hardback), by Karine Chemla (ed.). British Journal for the History of Science, 46(3): 517–519, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Curry:2014:GBG

[Cur14] Helen Anne Curry. From garden biotech to garage biotech: amateur experimental biology in historical perspective. British Journal for the History of Science, 47(3):539–565, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Curto:2018:BRL

[Cur18] Gemma Curto. Book review: Lisa Vox, Existential Threats: American Apocalyptic Beliefs in the Technological Era. Philadelphia: University of Pennsylvania Press, 2017. Pp. xvi + 266. ISBN 978-0-8122-4919-4. $55.00 (cloth). British Journal for the History of Science, 51(2): 319–321, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/lisa-vox-
                                        existential-threats-
                                        american-apocalyptic-
                                        beliefs-in-the-
                                        technological-era-
                                        philadelphia-university-of-
                                        pennsylvania-press-2017-pp-
                                        xvi-266-isbn-9780812249194-
                                        5500-cloth/0455B920CA8DD7425F3737DB56BDF5EC.

Chancellor:2009:CDN

[CV09] Gordon R. Chancellor and John Van Wyhe, editors. Charles Darwin’s notebooks from the voyage of the Beagle. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-51757-5 (cased). xxxiii + 615 + 8 pp. LCCN QH31.D2 C528 2009.

Coopmans:2014:RSP

[CVLW14] Catelijne Coopmans, Janet Vertesi, Michael Lynch, and Steve Woolgar, editors. Representation in scientific practice revisited. Inside technology. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-52538-0 (paperback), 0-262-31915-2 (ebook). 380 (est.) pp. LCCN Q180.55.M4 R455 2014.

DAmbrosio:2016:BRJ

[D’A16] Ubiratan D’Ambrosio. Book reviews: Jiří Hudeček, Reviving Ancient Chinese Mathematics: Mathematics, History and Politics in the Work of Wu Wen-Tsun. New York: Routledge, 2014. Pp. xii + 210. ISBN 978-0-4157-0296-6. £90.00 (hardback). British Journal for the History of Science, 49(1):143– 144, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

DAmbrosio:2018:BRI

[D’A18] Ubiratan D’Ambrosio. Book review: Ingrid Alexander-Skipnes (ed.), Visual Culture and Mathematics in the Early Modern Period. Abingdon and New York: Routledge, 2017. Pp. ix + 204. ISBN 978-1-138-67938-2. £110.00 (hardcover). British Journal for the History of Science, 51(1):158– 160, 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/ingrid-
                                        alexanderskipnes-ed-visual-
                                        culture-and-mathematics-in-
                                        the-early-modern-period-
                                        abingdon-and-new-york-
                                        routledge-2017-pp-ix-204-
                                        isbn-9781138679382-11000-
                                        hardcover/2964C64A54A6D43BEE84D6AD29AB39C9.

Dallal:2010:ISC

[Dal10] Ahmad S. Dallal. Islam, science, and the challenge of history. The Terry lectures. Yale University Press, New Haven, CT, USA, 2010. ISBN 0-300-15911-0 (hardcover). xii + 239 pp. LCCN BP190.5.S3 D35 2010.

Danielson:2014:PLC

[Dan14] Dennis Richard Danielson. Paradise Lost and the cosmological revolution. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-03360-8 (hardcover). xxiii + 220 pp. LCCN PR3592.S3 D36 2014.

Danne:2019:BRP

[Dan19]  Nicholas Danne.  Book re-

view: Pere Grapí, Inspiring Air: A History of Air-Related Science. Wilmington: Vernon Press, 2019. Pp. ix + 352. ISBN 1-62273-738-5. £44.00 (paperback). British Journal for the History of Science, 52(4): 717–719, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/pere-grapi-
inspiring-air-a-history-of-
airrelated-science-
wilmington-vernon-press-
2019-pp-ix-352-isbn-
1622737385-4400-paperback/
F8AFC182EE6867F777998C87652C9B09.

Darrigol:2011:BRS

[Dar11] Olivier Darrigol. Book review:

Suman Seth, Crafting the Quantum: Arnold Sommerfeld and the Practice of Theory, 1890– 1926. Cambridge, MA and London: MIT Press, 2010. Pp. vii + 378. ISBN 978-0-262-01373-4. £23.95 (hardback). British Journal for the History of Science, 44 (1):143–144, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Darrigol:2012:HOG

[Dar12] Olivier Darrigol. A history of optics from Greek antiquity to the nineteenth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-964437-3 (hardcover). xii + 327 pp. LCCN QC355.3 .D37 2012.

Darrigol:2014:PNR

[Dar14] Olivier Darrigol. Physics and necessity: rationalist pursuits from the Cartesian past to the quantum present. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-871288-X. xiv + 400 pp. LCCN QC6 .D32

2014. URL http://www.loc.
gov/catdir/enhancements/
fy1604/2013953857-b.html;
http://www.loc.gov/catdir/
enhancements/fy1604/
2013953857-d.html;
http://www.loc.gov/catdir/
enhancements/fy1604/
2013953857-t.html.

Das:2018:BRJ

[Das18] Anup Kumar Das. Book review: J. Benjamin Hurlbut, Experiments in Democracy: Human Embryo Research and the Politics of Bioethics. New York: Columbia University Press, 2017. Pp. 359. ISBN 978-0-231-17954-

6. $65.00/£54.00 (cloth). British Journal for the History of Science, 51(2):323–324, June 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/j-benjamin-hurlbut-
experiments-in-democracy-
human-embryo-research-and-
the-politics-of-bioethics-
new-york-columbia-
university-press-2017-pp-
359-isbn-9780231179546-
65005400-cloth/9DECA52616CE433A13B0B794BE41D6C6.

Davies:2019:RRE

[Dav19]  Stephen M. Davies.  Roth-

schild reversed: explaining the exceptionalism of biomedical research, 1971–1981. British Journal for the History of Science, 52(1):143–163, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/rothschild-
reversed-explaining-the-
exceptionalism-of-
biomedical-research-
19711981/
6528F49F7AA0F526040F7FE5B854D6B2.

Dawson:2010:BRJ

[Daw10] Gowan Dawson. Book review: John Holmes, Darwin’s Bards: British and American Poetry in the Age of Evolution. Edinburgh: Edinburgh University Press, 2009. Pp. xiv + 288. ISBN 978-0-7486-3940-3. £60.00 (hardback). British Journal for the History of Science, 43(2):305– 306, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2012:BRF

[Daw12] Gowan Dawson. Book reviews:

Fyfe, Steam-Powered Knowledge: William Chambers and the Business of Publishing, 1820–1860. Chicago and London: The University of Chicago Press, 2012. Pp. xvi + 313. ISBN 978-0-226-27651-9. £32.50 (hardback). British Journal for the History of Science, 45(4):687–688, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2014:BRA

[Daw14] Gowan Dawson. Book reviews:

Anne Dewitt, Moral Authority, Men of Science, and the Victorian Novel. Cambridge: Cambridge University Press, 2013. Pp. ix + 273. ISBN 978-1-107-03617-8. £60.00 (hardback). British Journal for the History of Science, 47(4):735–737, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2015:BRB

[Daw15a] Gowan Dawson. Book reviews: Bernard Lightman and Bennett Zon (eds.), Evolution and Victorian Culture. Cambridge: Cambridge University Press, 2014. Pp. xvii + 320. ISBN 978-1-107-02842-5. £60.00 (hardback). British Journal for the History of Science, 48(2):370–371, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2015:BRC

[Daw15b] Gowan Dawson. Book reviews: Catherine Marshall, Bernard Lightman and Richard England (eds.), The Papers of the Metaphysical Society, 1869–1880: A Critical Edition. 3 vols. Oxford: Oxford University Press, 2015. Pp. xxiv + 1,288. ISBN 978-0-19-964303-5. £320.00 (hardback). British Journal for the History of Science, 48(4): 705–706, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2015:BRE

[Daw15c] Gowan Dawson. Book reviews:

Efram Sera-Shriar, The Making of British Anthropology, 1813– 1871. London: Pickering and Chatto, 2013. Pp. xi + 255. ISBN 978-1-84893-394-1. £60.00 (hardback). British Journal for the History of Science, 48(3): 517–518, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2016:BRB

[Daw16a] Gowan Dawson. Book review:

Bernard Lightman (ed.), A Companion to the History of Science. Oxford: Wiley Blackwell, 2016. Pp. xvi + 601. ISBN 978-1-118-62077-9. £120.00 (hardback). British Journal for the History of Science, 49(4): 635–637, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
3A328F89946B7707C38B6AFEB1C394A5.

Dawson:2016:BRF

[Daw16b] Gowan Dawson. Book review:

Frederick Burkhardt, James A. Secord and the editors of the Darwin Correspondence Project (eds.), The Correspondence of Charles Darwin, vol. 23: 1875. Cambridge: Cambridge University Press, 2015. Pp. xliv + 799. ISBN 978-1-107-13436-2. £94.99 (hardback). British Journal for the History of Science, 49(4): 497–498, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/product/
                                        3BC5566069E3A63456122CF4DFA7B1F7.

Dawson:2016:BRM

[Daw16c] Gowan Dawson. Book reviews: Mark Francis and Michael W. Taylor (eds.), Herbert Spencer: Legacies. Abingdon: Routledge, 2015. Pp. 284. ISBN 978-1-84465-587-8. £65.00 (hardback). British Journal for the History of Science, 49(1):132–133, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawes:2017:BRH

[Daw17] Laura Dawes. Book review: Helen Valier, A History of Prostate Cancer: Cancer, Men and Medicine. Basingstoke: Palgrave Macmillan, 2016. Pp. 241. ISBN 978-1-4039-8803-4. £66.99 (hardback). British Journal for the History of Science, 50 (2):365–366, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        93E542A94D5E5A258AEA7B89C0579C9C.

daCosta:2009:SMK

[dC09] Palmira Fontes da Costa. The singular and the making of knowledge at the Royal Society of London in the Eighteenth Century. Cambridge Scholars Pub., Newcastle upon Tyne, UK, 2009. ISBN 1-4438-0357-X. xvi + 214 pp. LCCN Q41.L85 C67 2009.

daCosta:2014:BRA

[dC14] Palmira Fontes da Costa. Book reviews: Alexandra Cook, Jean-Jacques Rousseau and Botany: The Salutary Science. Oxford: Voltaire Foundation, 2012. Pp. xxi + 436. ISBN 978-0-7294-1055-7. £70.00 (paperback). British Journal for the History of Science, 47(4): 730–731, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Delisle:2009:PDN

[Del09] Richard G. Delisle. Les philosophies du néo-darwinisme: conceptions divergentes sur l’homme et le sens de l’évolution. (French) [Neo-Darwinism philosophies: divergent ideas on man and the meaning of evolution ]. Presses universitaires de France, Paris, France, 2009. ISBN 2-13-056626-X. x + 411 pp. LCCN ????

Delbourgo:2010:GLD

[Del10] James Delbourgo. Gardens of life and death. British Journal for the History of Science, 43(1):113– 118, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Derry:2012:ASB

[Der12] Margaret Elsinor Derry. Art and science in breeding: creating better chickens. University of Toronto Press, Toronto, ON, Canada, 2012. ISBN 1-4426-4395-1. viii + 281 pp. LCCN SF488.C2 D47 2012.

Derry:2015:MNB

[Der15] Margaret E. (Margaret Elsinor) Derry. Masterminding nature: the breeding of animals, 1750– 2010. University of Toronto Press, Toronto, ON, Canada, 2015. ISBN 1-4426-2652-6 (paperback), 1-4426-4904-6 (hardcover). viii + 310 pp. LCCN SF105 .D47 2015.

Desmond:2009:DSC

[Des09] Adrian J. Desmond. Darwin’s sacred cause: race, slavery and the quest for human origins. Allen Lane, London, UK, 2009. ISBN 1-84614-035-8. xxi + 485 + 16 pp. LCCN GN281.4 DES 2009; M09.E02866.

Desmarais:2012:JBH

[Des12]  Ralph  Des-

marais. Jacob Bronowski: a humanist intellectual for an atomic age, 1946–1956. British Journal for the History of Science, 45(4): 573–589, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

DeVun:2009:PAE

[DeV09] Leah DeVun. Prophecy, alchemy, and the end of time: John of Rupescissa in the late Middle Ages. Columbia University Press, New York, NY, USA, 2009. ISBN 0-231-14538-1 (cloth), 0-231-51934-6 (e-book). xiii + 255 pp. LCCN BR115.A57 D48

                                        2009. URL http://www.loc.
                                        gov/catdir/toc/ecip0823/
                                        2008030108.html.

Deveson:2016:PPP

[Dev16] Edward Deveson. Parasites, politics and public science: the promotion of biological control in Western Australia, 1900–1910. British Journal for the History of Science, 49(2):231–258, June 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dew:2009:OLX

[Dew09] Nicholas Dew. Orientalism in Louis XIV’s France. Oxford historical monographs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-923484-1 (hardcover). xv + 301 pp. LCCN DS61.9.F8 DEW 2009.

Dew:2010:STA

[Dew10] Nicholas Dew. Scientific travel in the Atlantic world: the French expedition to Gorée and the Antilles, 1681–1683. British Journal for the History of Science, 43(1): 1–17, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

DeWitt:2013:MAM

[DeW13] Anne DeWitt. Moral authority, men of science, and the Victorian novel. Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03617-8 (hardcover), 1-107-24796-9 (e-book), 1-139-56638-5 (e-book). ix + 273 pp. LCCN ????

Dewis:2014:LGP

[Dew14] Sarah Dewis. The Loudons and the gardening press: a Victorian cultural industry. The Nineteenth Century Series. Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4094-6922-0 (hardcover), 1-4094-6923-9 (e-book PDF), 1-4094-6924-7 (e-book ePUB). xvi + 278 pp. LCCN PR4891.L65 Z62 2014.

Dhondt:2015:UJU

[Dho15] Pieter Dhondt, editor. University jubilees and university history writing: a challenging relationship, volume Volume 13 of Scientific and Learned Cultures and Their Institutions. Brill, Leiden, The Netherlands, 2015. ISBN 90-04-21696-0, 90-04-26507-4 (ebook). ISSN 2352-1325. xi + 255 pp. LCCN LA621 .U55 2015.

Dick:2013:DCA

[Dic13] Steven J. Dick. Discovery and classification in astronomy: controversy and consensus. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03361-6 (hardcover). xvi + 458 pp. LCCN QB61 .D53 2013.

Diekmann:2015:BRM

[Die15] Nele Diekmann. Book reviews: Mirjam Brusius, Katrina Dean and Chitra Ramalingam (eds.), William Henry Fox Talbot: Beyond Photography. New Haven: Yale University Press, 2013. Pp. ix + 308 + 109 colour illus. ISBN 978-0-300-17934-7. £50.00 (cloth). British Journal for the History of Science, 48(3): 519–520, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dihal:2016:BRO

[Dih16] Kanta Dihal. Book reviews:

Olivier Darrigol, Physics and Necessity: Rationalist Pursuits from the Cartesian Past to the Quantum Present. Oxford: Oxford University Press, 2014. Pp. xv + 400. ISBN 978-0-19-871288-6. £39.00 (hardback). British Journal for the History of Science, 49(2):321–322, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dihal:2017:BRH

[Dih17] Kanta Dihal. Book review: Hans Christian von Baeyer, QBism: The Future of Quantum Physics. Cambridge, MA: Harvard University Press, 2016. Pp. 257. ISBN 978-0-674-50464-6. £18.95 (hardback). British Journal for the History of Science, 50 (2):378–379, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
8FB2AD2ACE468F8B3580ED68F7DADEA9.

Dijksterhuis:2012:BRO

[Dij12] Fokko Jan Dijksterhuis. Book reviews: Oliver Darrigol, A History of Optics: From Greek Antiquity to the Nineteenth Century. Oxford: Oxford University Press, 2012. Pp. xii + 327. ISBN 978-0-19-964437-7. £35.00 (hardback).

British Journal for the History of Science, 45(4):678–679, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dilley:2011:CDU

[Dil11] Stephen Dilley. Charles Darwin’s use of theology in the Origin of Species. British Journal for the History of Science, 44(4):??, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dilley:2012:CDU

[Dil12] Stephen Dilley. Charles Darwin’s use of theology in the Origin of Species. British Journal for the History of Science, 45(1):29– 56, March 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). See erratum [Dil13].

Dilley:2013:CDU

[Dil13] Stephen Dilley. Charles Darwin’s use of theology in the Origin of Species — erratum. British Journal for the History of Science, 46(1):149, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Dil12].

DiMoia:2013:RBB

[DiM13] John Paul DiMoia. Reconstructing bodies: biomedicine, health, and nation-building in South Korea since 1945. Studies of the Weatherhead East Asian Institute, Columbia University. Stanford University Press, Stanford, CA, USA, 2013. ISBN 0-8047-8411-6 (hardcover). x + 280 pp. LCCN R627 .D56 2013.

Dobre:2014:BRG

[Dob14] Mihnea Dobre. Book reviews:

Gideon Manning (ed.), Matter and Form in Early Modern Science and Philosophy. Leiden and Boston, MA: Brill, 2012. Pp. x + 248. ISBN 978-90-04-21870-3. EUR105.00 (hardback). British Journal for the History of Science, 47(2):375–376, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Doherty:2014:BRA

[Doh14] Meghan C. Doherty. Book reviews: Alison Walker, Arthur MacGregor and Michael Hunter (eds.), From Books to Bezoars: Sir Hans Sloane and His Collections. London: The British Library, 2012. Pp. x + 310. ISBN 978-0-7123-5880-4. £45.00 (hardback). British Journal for the History of Science, 47(4): 728–730, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Doherty:2016:BRA

[Doh16]  Caitlín Doherty.  Book re-

view: Alun Withey, Technology, Self-Fashioning and Politeness in Eighteenth-Century Britain: Refined Bodies. Basingstoke: Palgrave Macmillan, 2016. Pp. x + 140. ISBN 978-1-137-46747-8. £45.00 (hardback). British Journal for the History of Science, 49(3):490–491, September 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        cambridge.org/core/product/
                                        927E1ED2ADB59EBE80F8A2810BAA9063.

Doherty:2017:TTC

[Doh17] Caitlín Róisín Doherty. ‘Transporting thought’: cultures of balloon flight in Britain, 1784–1785. British Journal for the History of Science, 50(2):229–247, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        C7856622D686CDFD20EF2EBCF48D26D1.

Doig:2013:EEM

[Doi13] Kathleen Hardesty Doig. From Encylopédie to Encyclopédie méthodique: revision and expansion, volume 11 of SVEC. Voltaire Foundation, Oxford, UK, 2013. ISBN 0-7294-1077-3 (paperback). xiv + 314 pp. LCCN AE25.E6 D65 2013.

Donnelly:2014:BSP

[Don14] Kevin Donnelly. On the boredom of science: positional astronomy in the nineteenth century. British Journal for the History of Science, 47(3):479–503, September 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ducheyne:2017:PVM

[DP17] Steffen Ducheyne and Pieter Present. Pieter van Musschenbroek on laws of nature. British Journal for the History of Science, 50(4): 637–656, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
C6E077799F75295A9B0263DC49D793BB.

Dixon:2009:DI

[DR09] Michael F. Dixon and Gregory Radick. Darwin in Ilkley. History Press, Stroud, Gloucestershire, UK, 2009. ISBN 0-7524-5283-5. 126 pp. LCCN QH31.D2 D58 2009.

Draucker:2017:BRB

[Dra17] Shannon Draucker. Book review:

Benjamin Morgan, The Outward Mind: Materialist Aesthetics in Victorian Science and Literature. Chicago and London: The University of Chicago Press, 2017. Pp. 373. ISBN 978-0-226-46220-2. $35.00 (paperback). British Journal for the History of Science, 50 (4):734–735, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
9BD4879C8E9124BDA0E43F5E186C8072.

Dritsas:2010:ZDL

[Dri10]  Lawrence Dritsas.  Zambesi:

David Livingstone and expeditionary science in Africa, volume 1 of Tauris historical geography series. I. B. Tauris, London, UK, 2010. ISBN 1-84511-705-0 (hardcover). xii + 242 pp. LCCN DT1190.Z36 D75 2010.

deRegt:2009:SUP

[dRLE09] Henk  W.  de  Regt,

Sabina Leonelli, and Kai Eigner, editors. Scientific Understanding: Philosophical Perspectives. University of Pittsburgh Press, Pittsburgh, PA, USA, 2009. ISBN 0-8229-4378-6 (hardcover). ix + 352 pp. LCCN Q175 .S42415 2009.

Dry:2010:BRJ

[Dry10] Sarah Dry. Book review: Jennifer Karns Alexander, The Mantra of Efficiency: From Waterwheel to Social Control. Baltimore: The Johns Hopkins University Press, 2008. Pp. xvii + 233. ISBN 978-0-8018-8693-5. £33.50 (hardback). British Journal for the History of Science, 43(3):505–506, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Daston:2008:NLL

[DS08] Lorraine Daston and Michael Stolleis, editors. Natural law and laws of nature in early modern Europe: jurisprudence, theology, moral and natural philosophy. Ashgate, Aldershot, Hants, England and Farnham, England, 2008. ISBN 0-7546-5761-2. xii + 338 pp. LCCN K415 .N36

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip067/
                                        2006002164.html.

Dubow:2010:BRS

[Dub10]  Saul Dubow.  Book review:

Siegfried Huigen, Knowledge and Colonialism: Eighteenth-Century Travellers in South Africa. Leiden: Brill, 2009. Pp. xii + 273. ISBN 978-90-04-17743-7. EUR99.00 (hardback). British Journal for the History of Science, 43(3):490–491, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Duda:2014:PLC

[Dud14] Roman Duda, editor. Pearls from a lost city: the Lvov school of mathematics, volume 40 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2014. ISBN 1-4704-1076-1. LCCN QA14.U382 L85313 2014.

Duffy:2018:BRM

[Duf18] Timothy Duffy. Book review:

Meredith K. Ray, Margherita Sarrocchi’s Letters to Galileo: Astronomy, Astrology, and Poetics in Seventeenth-Century Italy. London and New York: Palgrave Macmillan, 2016. Pp. 98. ISBN 978-1-137-59769-4. £45.00 (hardcover). British Journal for the History of Science, 51(1):160–161, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/meredith-k-
ray-margherita-sarrocchis-
letters-to-galileo-
astronomy-astrology-and-
poetics-in-
seventeenthcentury-italy-
london-and-new-york-
palgrave-macmillan-2016-pp-98-

                                        isbn-9781137597694-4500-hardcover/
                                        16FA9D8722E5811DB716212F8009A8DD.

Dunn:2011:BRP

[Dun11] Richard Dunn. Book review: Peter J. T. Morris (ed.), Science for the Nation: Perspectives on the History of the Science Museum. Basingstoke: Palgrave Macmillan, 2010. Pp. xxi + 350. ISBN 978-0-230-23009-5. £65.00 (hardback). British Journal for the History of Science, 44 (1):152–153, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2012:BRJ

[Dun12a] Rachel L. Dunn. Book reviews: Joanna Stalnaker, The Unfinished Enlightenment: Description in the Age of the Encyclopedia. Ithaca and London: Cornell University Press, 2010. Pp. xvi + 240. ISBN 978-0-8014-4864-5. £27.95 (hardback). Jeff Loveland, An Alternative Encyclopedia? Dennis de Coetlogon’s Universal History (1745). Oxford: Voltaire Foundation, 2010. Pp. xiii + 256. ISBN 978-0-7294-0992-6. £60.00 (paperback). British Journal for the History of Science, 45(2):294– 296, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2012:BRP

[Dun12b] Richard Dunn. Book review: Peter Heering and Roland Wittje (eds.), Learning by Doing: Experiments and Instruments in the History of Science Teaching. Stuttgart: Franz Steiner Verlag, 2011. Pp. 362. ISBN 978-3-515-09842-7. EUR49.00 (paperback). British Journal for the History of Science, 45(2):310– 312, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2014:BRJ

[Dun14] Richard Dunn. Book reviews:

Jim Bennett and Sofia Talas (eds.), Cabinets of Experimental Philosophy in Eighteenth-Century Europe. Leiden: Brill, 2013. Pp. xxxvii + 253. ISBN 978-90-04-25296-7. $147.00 (hardback). British Journal for the History of Science, 47 (4):732–734, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2015:BRA

[Dun15a] Richard Dunn. Book reviews:

The Art and History of Globes. London: The British Library, 2014. Pp. 224. ISBN 978-07-123-5868-2. £30.00 (hardback). British Journal for the History of Science, 48(2):357–358, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2015:BRB

[Dun15b] Richard Dunn. Book reviews:

Brian Gee (ed. Anita McConnell and A.D. Morrison-Low), Francis Watkins and the Dollond Telescope Patent Controversy. Farnham: Ashgate, 2014. Pp. xxvi + 392. ISBN 978-1-4094-6643-7. £85.00 (hardback). British Journal for the History of Science, 48 (4):696–697, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2016:BRB

[Dun16a] Richard Dunn. Book review: Brian P. Copenhaver, Magic in Western Culture: From Antiquity to the Enlightenment. New York: Cambridge University Press, 2015. Pp. xiv + 612. ISBN 978-1-107-07052-3. £80.00 (hardback). British Journal for the History of Science, 49(3): 476–477, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
2B4F63365DA2FE92F85125476DAB387E.

Dunn:2016:BRD

[Dun16b] Richard Dunn. Book review: Darin Hayton, The Crown and the Cosmos: Astrology and the Politics of Maximilian I. Pittsburgh, PA: University of Pittsburgh Press, 2015. Pp. xii + 312. ISBN 978-0-8229-4443-0. $45.00 (cloth). British Journal for the History of Science, 49(4): 640–641, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
073BF2707817CDC88580BD94AB3C8547.

Dunn:2017:BRA

[Dun17a] Richard Dunn. Book review: Alison D. Morrison-Low, Sara J. Schechner and Paolo Brenni (eds.), How Scientific Instruments Have Changed Hands. Leiden and Boston: Brill, 2016. Pp. 239. ISBN 978-90-04-32492-3. EUR125.00/$150.00 (paperback). British Journal for the History of Science, 50(1):149–150, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
161B6152ADF5B6D511A9BE0AFAFD11DD.

Dunn:2017:BRM

[Dun17b] Richard Dunn. Book review:

Michael Hunter, The Image of Restoration Science: The Frontispiece to Thomas Sprat’s History of the Royal Society (1667). Abingdon and New York: Routledge, 2017. Pp. xvi + 150. ISBN 978-1-4724-7872-6. £115.00 (hardcover). British Journal for the History of Science, 50(4): 729–730, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
259D30D9802928EF345A572A1C196C56.

Dyde:2015:GCC

[Dyd15] Sean Dyde. George Combe and common sense. British Journal for the History of Science, 48(2): 233–259, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dysert:2014:BRA

[Dys14] Anna Dysert. Book reviews: Anne Van Arsdall and Timothy Graham (eds.), Herbs and Healers from the Ancient Mediterranean through the Medieval West: Essays in Honor of John M. Riddle. Farnham: Ashgate, 2012. Pp. xvi + 377. ISBN 978-1-4094-0038-7. £80.00 (hardback). British Journal for the History of Science, 47(2):369–371, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eagleton:2010:MMS

[Eag10]  Catherine Eagleton.  Monks,

Manuscripts and Sundials: The Navicula in Medieval England. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-27665-3. xi + 292 pp. LCCN ???? EUR 99.00.

Eberhart:2017:BRR

[Ebe17] Marlene L. Eberhart. Book review: Rebecca Cypess, Curious & Modern Inventions: Instrumental Music as Discovery in Galileo’s Italy. Chicago: The University of Chicago Press, 2016. Pp. xix + 307. ISBN 978-0-226-31944-5. £38.50 (cloth). British Journal for the History of Science, 50(2):349–351, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        A5C6BF7A505C7E13701DDAC9CFFE086B.

Eberhart:2018:BRJ

[Ebe18] Marlene L. Eberhart. Book review: James Q. Davies and Ellen Lockhart (eds.), Sound Knowledge: Music and Science in London, 1789–1851. Chicago and London: The University of Chicago Press, 2016. Pp. 257. ISBN 978-0-2264-0207-9. $55.00 (cloth). British Journal for the History of Science, 51(3): 525–527, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/james-q-
davies-and-ellen-lockhart-
eds-sound-knowledge-music-
and-science-in-london-
17891851-chicago-and-
london-the-university-of-
chicago-press-2016-pp-257-
isbn-9780226402079-5500-
cloth/E3B81CC213DBAF2EF54F7E6D12145811.

Echeverria:2011:HFT

[Ech11] Virginia Iommi Echeverría. Hydrostatics on the fray: Tartaglia, Cardano and the recovering of sunken ships. British Journal for the History of Science, 44(4): 479–491, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Echterholter:2012:SGP

[Ech12] Anna Echterhölter. Schattengefechte: genealogische Praktiken in Nachrufen auf Naturwissenschaftler (1710–1860). (German) [Shadow battles: genealogical practices in obituaries of scientists]. Wissenschaftsgeschichte. Wallstein, Göttingen, Germany, 2012. ISBN 3-8353-1071-2 (paperback). 365 pp. LCCN Q125 .E3625 2012.

Eddy:2008:LMJ

[Edd08] Matthew Eddy, editor. The language of mineralogy: John Walker, chemistry and the Edinburgh Medical School, 1750–1800. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hants, England and Farnham, England, 2008. ISBN 0-7546-6332-9 (hardcover). xx + 309 pp. LCCN QH31.W16 E33

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip0819/
                                        2008023651.html.

Eddy:2017:PCL

[Edd17] Matthew Daniel Eddy. The politics of cognition: liberalism and the evolutionary origins of Victorian education. British Journal for the History of Science, 50(4): 677–699, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        1FC5A97C0CD0D8CE70491DC9DB671E45.

Edgerton:2009:MWT

[Edg09] Samuel Y. Edgerton. The mirror, the window, and the telescope: how Renaissance linear perspective changed our vision of the universe. Cornell University Press, Ithaca, NY, USA, 2009. ISBN 0-8014-4758-5 (hardcover), 0-8014-7480-9 (paperback). xvi + 199 pp. LCCN NC748 .E34

                                        2009. URL http://www.loc.
                                        gov/catdir/toc/ecip0824/
                                        2008031951.html.

Edwards:2014:BRR

[Edw14] Michael Edwards. Book reviews: Rhodri Lewis, William Petty on the Order of Nature: An Unpublished Manuscript Treatise. Tempe: Arizona Centre for Medieval and Renaissance Studies, 2012. Pp. xiv + 176. ISBN 978-0-86698-447-8. £40.00 (hardback). British Journal for the History of Science, 47(4):726–727, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Effros:2017:BGP

[Eff17] Bonnie Effros. Berber genealogy and the politics of prehistoric archaeology and craniology in French Algeria (1860s–1880s). British Journal for the History of Science, 50(1):61–81, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
9849242C6132CB92AC26B6478390A51F.

Egmond:2010:WCC

[Egm10] Florike Egmond. The world of Carolus Clusius: natural history in the making, 1550–1610, volume 6. Pickering and Chatto, London, UK, 2010. ISBN 1-84893-008-9 (hardcover), (paperback), 1-84893-009-7 (paperback). xiv + 292 pp.

Egmond:2012:BRL

[Egm12] Florike Egmond. Book review:

Lissa Roberts (ed.), Centres and Cycles of Accumulation in and around the Netherlands during the Early Modern Period. Zurich and Berlin: LIT Verlag, 2011. Pp. ii + 290. ISBN 978-3-643-90095-1. EUR34.95 (paperback). British Journal for the History of Science, 45(3):459–460, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Egmond:2013:BRS

[Egm13] Florike Egmond. Book reviews: Sachiko Kusukawa, Picturing the Book of Nature: Image, Text and Argument in Sixteenth-Century Human Anatomy and Medical Botany. Chicago and London: The University of Chicago Press, 2012. Pp. xvii + 331. ISBN 978-0-226-46529-6. $45.00/£29.00 (hardback). British Journal for the History of Science, 46(2):342– 344, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eichberg:2012:BRD

[Eic12a] Stephanie Eichberg. Book review: Domenico Bertoloni Meli, Mechanism, Experiment, Disease: Marcello Malpighi and Seventeenth-Century Anatomy. Baltimore: Johns Hopkins University Press, 2011. Pp. xii + 439. ISBN 978-0-8018-9904-1. £23.50 (paperback). British Journal for the History of Science, 45(2):290– 291, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eichberg:2012:BRJ

[Eic12b] Stephanie Eichberg. Book review: Justin E. H. Smith, Divine Machines: Leibniz and the Sciences of Life. Princeton and Oxford: Princeton University Press, 2011. Pp. xii + 380. ISBN 978-0-691-1478-7. £30.95 (hardback). British Journal for the History of Science, 45(1):131–132, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Smi11a].

Eichberg:2015:BRA

[Eic15] Stephanie Eichberg. Book reviews: Angelique Richardson (ed.), After Darwin: Animals, Emotions, and the Mind. Amsterdam: Rodopi, 2013. Pp. xvi + 369. ISBN 978-90-420-3747-2. EUR85.00 (hardback). British Journal for the History of Science, 48(3):523–525, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Endt-Jones:2013:CSR

[EJ13]  Marion  Endt-Jones,

editor. Coral: something rich and strange. Liverpool University Press, Liverpool, UK, 2013. ISBN 1-84631-959-5 (paperback). 128 pp. LCCN SF458.C64 C67 2013.

Erickson:2013:HRA

[EKD^{+}13] Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca M. (Rebecca Maura) Lemov, Thomas Sturm, and Michael D. Gordin. How reason almost lost its mind: the strange career of Cold War rationality. University of Chicago Press, Chicago, IL, USA, 2013.

ISBN 0-226-04663-X (hardcover), 0-226-04677-X (e-book). 259 pp. LCCN D843 .E69 2013.

Ekholm:2010:TRM

[Ekh10] Karin J. Ekholm. Tartaglia’s ragioni: a maestro d’abaco’s mixed approach to the bombardier’s problem. British Journal for the History of Science, 43(2):181– 207, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Elliott:2010:EMS

[Ell10] Paul A. Elliott. Enlightenment, Modernity and Science: Geographies of Scientific Culture and Improvement in Georgian England. I. B. Tauris, London, UK and New York, NY, USA, 2010. ISBN 1-84885-366-1 (hardcover). xii + 363 pp. LCCN ????

Ellis:2017:CUM

[Ell17] Patrick Ellis. A cinema for the unborn: moving pictures, mental pictures and Electra Sparks’s New Thought film theory. British Journal for the History of Science, 50(3):411–428, September 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        6FF001B9660638B3FE6D814984E44C95.

Elmer:2013:MCV

[Elm13] Peter Elmer. The miraculous conformist: Valentine Greatrakes, the body politic, and the politics of healing in Restoration Britain. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966396-3 (hardcover). xiii + 279 pp. LCCN DA447.G74 E56 2013.

Elton:2012:BRD

[Elt12] Julia Elton. Book review: A. D.

Morrison-Low, Northern Lights: The Age of Scottish Lighthouses. Edinburgh: National Museums Scotland in conjunction with The Royal Scottish Society of Arts, 2010. Pp. xxvi + 262. ISBN 978-1-905267-47-7. £17.99 (paperback). British Journal for the History of Science, 45(2):300– 301, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Endersby:2014:UL

[End14] Jim Endersby. Unsuitable for ladies? British Journal for the History of Science, 47(2):363– 367, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Endersby:2016:DOS

[End16] Jim Endersby. Deceived by orchids: sex, science, fiction and Darwin. British Journal for the History of Science, 49(2):205– 229, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Endersby:2018:UB

[End18a] Jim Endersby. Utopian biologies. British Journal for the History of Science, 51(1):147– 152, 2018. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/utopian-biologies/
E18FDFDD6A336B659DBFDDB7CCAF8444.

Endersby:2018:VBG

[End18b] Jim Endersby. A visit to Biotopia: genre, genetics and gardening in the early twentieth century. British Journal for the History of Science, 51(3): 423–455, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/visit-to-
biotopia-genre-genetics-
and-gardening-in-the-early-
twentieth-century/
BFB996938A7185EA7F50C1AE80C6D3A2.

Enebakk:2014:HMO

[Ene14] Vidar Enebakk. Hansteen’s magnetometer and the origin of the magnetic crusade. British Journal for the History of Science, 47 (4):587–608, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ensmenger:2012:BRM

[Ens12] Nathan Ensmenger. Book review: Mark Priestley, A Science of Operations: Machines, Logic, and the Invention of Programming. London, Heidelberg and New York: Springer, 2011. Pp. ix + 341. ISBN 978-1-84882-554-3.

£90.00 (hardback). British Journal for the History of Science, 45 (1):144–145, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Pri10].

Epstein:2014:TIM

[Eps14] Katherine C. Epstein. Torpedo:

Inventing the military–industrial complex in the United States and Great Britain. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72526-3. 305 pp. LCCN V850 .E67 2014.

Erlingsson:2011:IIE

[Erl11] Steindór J. Erlingsson. Institutions and innovation: experimental zoology and the creation of the British Journal of Experimental Biology and the Society for Experimental Biology. British Journal for the History of Science, 44(4):??, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Erlingsson:2013:IIE

[Erl13] Steindór J. Erlingsson. Institutions and innovation: experimental zoology and the creation of the British Journal of Experimental Biology and the Society for Experimental Biology. British Journal for the History of Science, 46(1):73–95, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Easterby-Smith:2016:BRM

[ES16] Sarah Easterby-Smith. Book reviews: Mary Terrall, Catching Nature in the Act: Réaumur and the Practice of Natural History in the Eighteenth Century. Chicago and London: The University of Chicago Press, 2014. Pp. 275. ISBN 978-0-2260-8860-0. £28.00 (hardback). British Journal for the History of Science, 49(1):125– 126, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Esposito:2010:BRG

[Esp10] Maurizio Esposito. Book review: G. Richard Delisle, Les Philosophies du néo-darwinisme. Preface by Jean Gayon. Paris: Presses universitaires France, 2009. Pp. 456. ISBN 978-2-13-056626-7. Canadian $56.70 (paperback). British Journal for the History of Science, 43(2):315– 317, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Esposito:2013:RB

[Esp13] Maurizio Esposito. Romantic biology: 1890–1945, volume 1 of History and philosophy of biology. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-420-3 (hardcover), 1-78144-046-8 (e-book). viii + 257 pp. LCCN QH305 .E87 2013.

Evans:2013:BRG

[Eva13] Jessica Evans. Book reviews: G. E. R. Lloyd, Being, Humanity, and Understanding. Oxford: Oxford University Press, 2012. Pp. vi + 136. ISBN 978-0-19-965472-7. £25.00 (hardback).

British Journal for the History of Science, 46(4):715–716, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Everest:2013:BRO

[Eve13] James Everest. Book reviews:

The Oxford Francis Bacon, vol. 1: Early Writings, 1584–1596. Oxford: Clarendon Press, 2012. Pp. lxi + 1066. ISBN 978-0-19-818313-6. £210.00 (hardback). Alan Stewart with Harriet Knight (eds.). British Journal for the History of Science, 46(3):522– 523, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Everest:2018:BRM

[Eve18] James Everest. Book review:

A. Mark Smith, From Sight to Light: The Passage from Ancient to Modern Optics. Chicago and London: The University of Chicago Press, 2017. Pp xi + 457. ISBN 978-0-226-52857-1. $36.00 (paperback). British Journal for the History of Science, 51(4): 705–706, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mark-smith-
from-sight-to-light-the-
passage-from-ancient-to-
modern-optics-chicago-and-
london-the-university-of-
chicago-press-2017-pp-xi-
457-isbn-9780226528571-

                                        3600-paperback/
                                        DB3DBBE8E549A7318DA777AF423C0E22.

Falk:2013:BRS

[Fal13a] Seb Falk. Book reviews: A Survey of European Astronomical Tables in the Late Middle Ages. Leiden: Brill, 2012. Pp. xix + 250. ISBN 978-90-04-23058-3. EUR 107.00 (hardback), by José Chabás and Bernard R. Goldstein. British Journal for the History of Science, 46(3):520–522, September 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fallwell:2013:MGM

[Fal13b] Lynne Anne Fallwell. Modern German midwifery, 1885– 1960, volume 13 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-428-9 (hardcover), 1-78144-053-0 (e-book). xiii + 263 pp. LCCN RG950 .F348 2013.

Falconer:2018:BRF

[Fal18] Isobel Falconer. Book review: Friedrich Steinle, Exploratory Experiments: Ampère, Faraday and the Origins of Electrodynamics. Translated by Alex Levine. Pittsburgh, PA: University of Pittsburgh Press, 2016. Pp. 494. ISBN 978-0-8229-4450-8. $65.00 (hardcover). British Journal for the History of Science, 51 (2):328–329, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/friedrich-
steinle-exploratory-
experiments-ampere-faraday-
and-the-origins-of-
electrodynamics-translated-
by-alex-levine-pittsburgh-
pa-university-of-
pittsburgh-press-2016-pp-
494-isbn-9780822944508-6500-
hardcover/BBAB8CD68295F601D4C2034287101896.

Fangerau:2010:SSW

[Fan10] Heiner Fangerau. Spinning the scientific web: Jacques Loeb (1859– 1924) und sein Programm einer internationalen biomedizinischen Grundlagenforschung. Akademie-Verlag, Berlin, Germany, 2010. ISBN 3-05-004528-0. 280 pp. LCCN QH31.L73 F36 2010.

Fanger:2012:BRS

[Fan12]  Claire Fanger.  Book re-

views: Stuart McWilliams, Magical Thinking: History, Possibility and the Idea of the Occult. London: Continuum, 2012. Pp. x + 187. ISBN 978-1-4411-1697-0. £65.00 (hardback). British Journal for the History of Science, 45 (4):681–682, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2010:BRC

[Far10a] Patricia Fara. Book review:

Celina Fox, The Arts of Industry in the Age of Enlightenment. New Haven and London: Yale University Press, 2010. Pp. viii + 576. ISBN 978-0-300-16042-0. £40.00 (hardback). British Journal for the History of Science, 43(3): 488–489, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2010:BRF

[Far10b] Patricia Fara. Book review: Faidra Papanelopoulu, Agustí Nieto-Galan and Enrique Perdiguero (eds.), Popularizing Science and Technology in the European Periphery, 1820–2000. Aldershot: Ashgate, 2009. Pp. xviii + 284. ISBN 978-0-7546-6269-3. £60.00 (hardback). British Journal for the History of Science, 43(2):311– 312, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2011:BRP

[Far11] Patricia Fara. Book review: P. M. Harman, The Culture of Nature in Britain 1680–1860. New Haven and London: Yale University Press, 2009. Pp. xi + 393. ISBN 978-0-300-15197-8. £45.00 (hardback). British Journal for the History of Science, 44(1): 126–127, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2012:BRF

[Far12] Patricia Fara. Book reviews: Ferreiro, Measure of the Earth: The Enlightenment Expedition that Reshaped the World. New York: Basic Books, 2011. Pp. xix + 353. ISBN 978-0-465-01723-2. $28.00 (hardback). British Journal for the History of Science, 45 (4):685–686, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2013:BRG

[Far13] Patricia Fara. Book reviews: Gregory Lynall, Swift and Science: The Satire, Politics and Theology of Natural Knowledge, 1690– 1730. London: Palgrave Macmillan, 2012. Pp. ix + 209. ISBN 978-0-230-34364-1. £50.00 (hardback). British Journal for the History of Science, 46(1):165– 166, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2015:BRB

[Far15] Patricia Fara. Book reviews:

Ben Russell, James Watt: Making the World Anew. London: Reaktion Books, 2014. Pp. 280. ISBN 978-1-78023-375-8. £17.95 (hardback). British Journal for the History of Science, 48(4): 694–696, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2016:BRE

[Far16] Patricia Fara. Book reviews:

Eva Hemmungs Wirtén, Making Marie Curie: Intellectual Property and Celebrity Culture in an Age of Information. Chicago and London: The University of Chicago Press, 2015. Pp. 223. ISBN 978-0-2262-3584-4. £24.50, $35.00 (hardback). British Journal for the History of Science, 49 (1):134–135, March 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2018:PWP

[Far18] Patricia Fara. The past as a work in progress. British Journal for the History of Science, 51(1):1–15, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/past-as-a-
work-in-progress/
5E893FE0C8B59237A15CCBFACE28D8C8.

Fara:2019:BRT

[Far19] Patricia Fara. Book review: Tita Chico, The Experimental Imagination: Literary Knowledge and Science in the British Enlightenment. Stanford, CA: Stanford University Press, 2018. Pp. xi + 242. ISBN 978-1-5036-0544-

2. $60.00 (hardcover). British Journal for the History of Science, 52(2):366–367, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/tita-chico-
the-experimental-
imagination-literary-
knowledge-and-science-in-
the-british-enlightenment-
stanford-ca-stanford-
university-press-2018-pp-
xi-242-isbn-9781503605442-
6000-hardcover/B278C0D7236627FA3B652BD14CDB85D6.

Feingold:2019:RHF

[Fei19] Mordechai Feingold. Response to H. Floris Cohen’s essay review on Newtonian scholarship. British Journal for the History of Science, 52(2):353–357, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/response-
to-h-floris-cohens-essay-
review-on-newtonian-
scholarship/
33D9C6F1023F94F2C42E8627B5347199.

Feola:2018:BRS

[Feo18] Vittoria Feola. Book review:

Sarah R. Kyle, Medicine and Humanism in Late Medieval Italy: The Carrara Herbal in Padua. Abingdon and New York: Routledge, 2017. Pp. xiii + 243. ISBN 978-1-4724-4652-

7. £110.00 (hardcover). British Journal for the History of Science, 51(1):157–158, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/sarah-r-
kyle-medicine-and-humanism-
in-late-medieval-italy-the-
carrara-herbal-in-padua-
abingdon-and-new-york-
routledge-2017-pp-xiii-243-
isbn-9781472446527-11000-
hardcover/5282EA15D789B1A8A617591DDB4B27AA.

Feola:2019:AAH

[Feo19] Vittoria Feola. Agnes Arber, historian of botany and Darwinian sceptic. British Journal for the History of Science, 52(3): 515–523, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/agnes-
                                        arber-historian-of-botany-
                                        and-darwinian-sceptic/
                                        54B0E722EAD1C3B81863B2D3957DCA55.

Ferreiro:2011:BRJ

[Fer11a] Larrie D. Ferreiro. Book review: John Laurence Busch, Steam Coffin: Captain Moses Rogers and the Steamship Savannah Break the Barrier. New Canaan, CT: Hodos Historia LLC, 2010. Pp. xii + 726. ISBN 978-1-893616-00-4. $35.00 (hardback). British Journal for the History of Science, 44 (4):605–606, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ferreiro:2011:MEE

[Fer11b] Larrie D. Ferreiro. Measure of the Earth: the Enlightenment expedition that reshaped our world. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02345-2 (e-book), 0-465-01723-1 (hardcover). xix + 353 pp. LCCN QB296.E9 F47 2011.

Ferguson:2012:DSE

[Fer12]  Christine Ferguson.  Deter-

mined spirits: eugenics, heredity and racial regeneration in Anglo–American spiritualist writing, 1848–1930. Edinburgh critical studies in Victorian culture. Edinburgh University Press, Edinburgh, Scotland, 2012. ISBN 0-7486-3965-9, 0-7486-5066-0 (Web PDF), 0-7486-5067-9 (epub), 0-7486-5068-7 (e-book). x + 230 pp. LCCN BF1275.L58 F47 2012.

Ferber:2013:BHP

[Fer13] Sarah Ferber. Bioethics in historical perspective. Palgrave Macmillan, Basingstoke, Hampshire, UK, 2013. ISBN 1-4039-8723-8 (hardcover), 1-4039-8724-6 (paperback). xiv + 233 pp. LCCN QH332 .F46 2013.

Few:2010:CSK

[Few10] Martha Few. Circulating smallpox knowledge: Guatemalan doctors, Maya Indians and designing Spain’s smallpox vaccination expedition, 1780–1803. British Journal for the History of Science, 43(4):519–537, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Feyerabend:2010:AM

[Fey10]  Paul  Feyerabend.

Against method. Verso, London, UK, updated edition, 2010. ISBN 1-84467-443-6 (hardcover), 1-84467-442-8 (paperback). xxxii + 296 pp. LCCN Q175 .F42 2010. With a new introduction by Ian Hacking.

Fletcher:2011:SLW

[FF11] John Edward Fletcher and Elizabeth Fletcher, editors. A study of the life and works of Athanasius Kircher, “Germanus incredibilis”: with a selection of his unpublished correspondence and an annotated translation of his autobiography, volume 12 of Aries book series. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-20712-0 (hardcover). ISSN 1871-1405. ???? pp. LCCN CT1098.K46 F53 2011.

Field:2010:OAR

[Fie10] J. V. Field. Obituaries: Alfred Rupert Hall (26 July 1920– 5 February 2009) and Marie Boas Hall (18 October 1919–23 February 2009). British Journal for the History of Science, 43(1):99– 103, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Finnegan:2009:NHS

[Fin09] Diarmid A. Finnegan. Natural history societies and civic culture in Victorian Scotland, volume 9 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2009. ISBN 1-85196-658-7 (hardcover). xi + 254 pp. LCCN 501; M09.E10358.

Findlen:2010:BRP

[Fin10a] Paula Findlen. Book reviews: Pamela O. Long, David McGee and Alan M. Stahl (eds.), The Book of Michael of Rhodes: A Fifteenth-Century Maritime Manuscript. Volume 1: Fascimile. Cambridge, MA and London: The MIT Press, 2009. Pp. xiv + 519. ISBN 978-0-262-13503-0. $65.00 (hardback). Pamela O. Long, David McGee and Alan M. Stahl (eds.). The Book of Michael of Rhodes: A Fifteenth-Century Maritime Manuscript. Volume 2: Transcription and Translation. Transcription by Franco Rossi. Cambridge, MA and London: The MIT Press, 2009. Pp. lii + 679. ISBN 978-0-262-19590-4. $75.00 (hardback). Pamela O. Long, David McGee and Alan M. Stahl (eds.). The Book of Michael of Rhodes: A Fifteenth-Century Maritime Manuscript. Volume 3: Studies. Cambridge, MA and London: The MIT Press, 2009. Pp. xiii + 370. ISBN 978-0-262-12308-2. $45.00 (hardback). British Journal for the History of Science, 43(3):481–482, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Finn:2010:BRG

[Fin10b] Michael A. Finn. Book review:

Geoffrey Burnstock, Richard Frackowiak, Uta Frith, Richard Gregory, Terry Jones, Sir Peter Mansfield, Salvador Moncada, Alan North, Roger Ordidge, Sir Michael Rutter, Ann Silver and Elizabeth Warrington, Today’s Neuroscience, Tomorrow’s History: A Video Archive Project, Interviews by Richard Thomas. London: UCL and Wellcome Trust, 2009. 12 DVDs. No price given. British Journal for the History of Science, 43(4): 622–623, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Finn:2010:BRM

[Fin10c] Michael A. Finn. Book review: Mike Jay, The Atmosphere of Heaven: The Unnatural Experiments of Dr Beddoes and His Sons of Genius. New Haven and London: Yale University Press, 2009. Pp. vii + 294. ISBN 978-0-300-12439-2. £20.00 (hardback). British Journal for the History of Science, 43(1):125– 126, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Finn:2011:BRL

[Fin11] Michael A. Finn. Book review: Laura Salisbury and Andrew Shail (eds.), Neurology and Modernity: A Cultural History of Nervous Systems 1800–1950. Basingstoke: Palgrave Macmillan, 2010. Pp. xiii + 298. ISBN 978-0-230-23313-3. £55.00 (hardback). British Journal for the History of Science, 44(2):300– 301, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Finnegan:2017:BRP

[Fin17] Diarmid Finnegan. Book review: Patrick Manning and Daniel Rood (eds.), Global Scientific Practice in an Age of Revolutions, 1750–1850. Pittsburgh: University of Pittsburgh Press, 2016. Pp. 416. ISBN 978-0822944546. £40.00 (hardback). British Journal for the History of Science, 50(2):354–356, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
3FD45019C28C60C5BC8156F45A6F6473.

Fisher:2010:CRC

[Fis10] John Fisher. Conjectures and reputations: The composition and reception of James Bradley’s paper on the aberration of light with some reference to a third unpublished version. British Journal for the History of Science, 43(1):19–48, March 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fisch:2014:BTL

[Fis14] Menachem Fisch. Babbage’s two lives. British Journal for the History of Science, 47(1):95–118, March 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fix:2016:BRA

[Fix16] Adam Fix. Book review: Amir Alexander, Infinitesimal: How a Dangerous Mathematical Theory Shaped the Modern World. New York: Farrar, Straus, and Giroux, 2015. Pp. 352. ISBN 978-0-374-53499-8. $16.00 (paperback). British Journal for the History of Science, 49(3): 480–481, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
6BC08BE48951FD4AD5ED85197AE8DFA4.

Fix:2019:BRD

[Fix19] Adam Fix. Book review: Davide Crippa, The Impossibility of Squaring the Circle in the 17th Century: A Debate among Gregory, Huygens and Leibniz. Basel: Birkhäuser, 2019. Pp. ix + 184. ISBN 978-3-0300-1637-1. $54.99 (paperback). British Journal for the History of Science, 52(4): 709–711, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/davide-
crippa-the-impossibility-
of-squaring-the-circle-in-
the-17th-century-a-debate-
among-gregory-huygens-and-
leibniz-basel-birkhauser-
2019-pp-ix-184-isbn-
9783030016371-5499-
paperback/80C1C6A015B7DFFF64D2E85B6047D8DC.

Faraday:2011:CHC

[FJ11] Michael Faraday and Frank A. J. L. James, editors. The chemical history of a candle. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, sesquicentenary edition, 2011. ISBN 0-19-969491-5 (hardcover). xlviii + 152 pp. LCCN QD39 .F2 2011.

Fleming:2014:TAB

[FJB14]  James  Rodger

Fleming, Ann Johnson, and Peter Brimblecombe, editors. Toxic airs: body, place, planet in historical perspective. University of Pittsburgh Press, Pittsburgh, PA, USA, 2014. ISBN 0-8229-6290-X, 1-306-55511-6, 0-8229-7952-7 (e-book). xiv + 284 pp. LCCN TD883 .T66

2014. URL http://muse.jhu.
edu/books/9780822979524;
http://www.jstor.org/
stable/10.2307/j.ctt5vkgsj.

Flannery:2019:BRP

[Fla19] Michael A. Flannery. Book review: Patrick Armstrong, Alfred Russel Wallace. London: Reaktion Books, 2019. Pp. 175, ISBN 978-1-7891-4085-9, £11.99 (paperback). British Journal for the History of Science, 52(3): 526–528, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/patrick-
armstrong-alfred-russel-
wallace-london-reaktion-
books-2019-pp-175-isbn-
9781789140859-1199-
paperback/
EA8D244CBC65B0AC292487CA752DFD10.

Flood:2014:JCM

[FMW14] Raymond Flood, Mark McCartney, and Andrew Whitaker, editors. James Clerk Maxwell: perspectives on his life and work. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-966437-4. x + 364 + 4 pp. LCCN QC16.M4

                                        J36 2014. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1410/2013942195-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1410/
                                        2013942195-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1410/
                                        2013942195-t.html.

Feyerabend:2011:TS

[FO11] Paul Feyerabend and Eric Oberheim, editors. Tyranny of science. Polity Press, Cambridge, UK, 2011. ISBN 0-7456-5189-5 (hardcover), 0-7456-5190-9 (paperback). xii + 153 pp. LCCN Q175 .F49 2011.

FontesDaCosta:2012:BRA

[Fon12] Palmira Fontes Da Costa. Book review: Anna Marie Roos, Web of Nature: Martin Lister (1639– 1712), the First Arachnologist. Leiden and Boston: Brill, 2011. Pp. xx + 478. ISBN 978-90-04-20703-5. EUR129.00 (hardback). British Journal for the History of Science, 45(2):293– 294, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Forgan:2010:BRD

[For10] Sophie Forgan. Book review: David Knight, The Making of Modern Science: Science, Technology, Medicine, and Modernity, 1789–1914. Cambridge and Malden, MA: Polity, 2009. Pp. xiv + 370. ISBN 978-0-7456-3676-4. £17.99 (paperback). British Journal for the History of Science, 43(2):301–302, June 2010.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fotinos:2018:BRP

[Fot18] Nicoletta I. Fotinos. Book review: Paul Weindling (ed.), From Clinic to Concentration Camp: Reassessing Nazi Medical and Racial Research, 1933–1945. London: Routledge, 2017. Pp. 376. ISBN 978-1-4724-8461-1. £105.00 (hardcover). British Journal for the History of Science, 51(3): 532–533, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/paul-
weindling-ed-from-clinic-
to-concentration-camp-
reassessing-nazi-medical-
and-racial-research-
19331945-london-routledge-
2017-pp-376-isbn-
9781472484611-10500-
hardcover/8AF8FBDB09B73F8300359940741F8BA2.

Fox:2009:AIA

[Fox09] Celina Fox. The arts of industry in the Age of Enlightenment. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-16042-9. vii + 576 pp. LCCN N72.T4 F69 2009.

Fox:2012:SSS

[Fox12a] Robert Fox. The savant and the state: science and cultural politics in nineteenth-century France. The Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0878-3 (e-book), 1-4214-0878-3, 1-4214-0522-9 (hardcover). 424 pp. LCCN Q127.F8 F694

2012. URL http://muse.jhu.
edu/books/9781421408781/.

Fox:2012:THH

[Fox12b] Robert Fox, editor. Thomas Harriot and his world: mathematics, exploration, and natural philosophy in early modern England. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 0-7546-6960-2 (hardcover). xviii + 255 pp. LCCN Q143.H36 T46 2012.

Frampton:2016:HSI

[Fra16] Sally Frampton. Honour and subsistence: invention, credit and surgery in the nineteenth century. British Journal for the History of Science, 49(4): 561–576, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
BE52ACD29C853E6F7BC325900C53561E.

Frampton:2017:BRS

[Fra17]  Sally Frampton.  Book re-

view: Susan E. Cayleff, Nature’s Path: A History of Naturopathic Healing in America. Baltimore: Johns Hopkins University Press, 2016. Pp x + 397. ISBN 978-1-4214-1903-9. $39.95 (hardback). British Journal for the History of Science, 50 (2):362–363, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
CF8FA7BA24A34A41858ED87188B8B51E.

Freire:2010:BRC

[Fre10]  Lucas G. Freire.  Book re-

view: Christopher Carter, Magnetic Fever: Global Imperialism and Empiricism in the Nineteenth Century. Philadelphia: American Philosophical Society, 2009. Pp. xxvi + 168. ISBN 978-1-60618-994-8. $35.00 (paperback). British Journal for the History of Science, 43(4):614–616, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fritsche:2011:BPM

[Fri11] Johannes Fritsche. The biological precedents for medieval impetus theory and its Aristotelian character. British Journal for the History of Science, 44(1):1–27, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Scheer:2014:ELC

[FSS^{+}14] Ute Frevert, Monique Scheer, Anne Schmidt, Pascal Eitler, Bettina Hitzer, Nina Verheyen, Benno Gammerl, Christian Bailey, and Margrit Pernau. Emotional lexicons: continuity and change in the vocabulary of feeling 1700–2000. Emotions in history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-965573-1. x + 288 pp. LCCN BF582 .E44

                                        2014. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1503/2013940854-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1503/
                                        2013940854-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1503/
                                        2013940854-t.html.

Francis:2015:HSL

[FT15] Mark Francis and Michael Taylor, editors. Herbert Spencer: legacies. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2015. ISBN 1-84465-587-3, 1-84465-626-8. viii + 275 pp. LCCN B1657 .H474 2015.

Fuller:2010:GRL

[Ful10a] Steve Fuller. Book review: G. E. R. Lloyd, Disciplines in the Making. Oxford: Oxford University Press, 2009. Pp. viii + 215. ISBN 978-0-19-956787-4. £25.00 (hardback). British Journal for the History of Science, 43(4): 607–609, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fuller:2010:S

[Ful10b] Steve Fuller. Science. The art of living. Acumen, Durham, UK, 2010. ISBN 1-84465-204-1 (paperback). v + 170 pp. LCCN ????

Fulford:2018:BRT

[Ful18]  Tim Fulford.  Book review:

Trevor Levere, Larry Stewart and Hugh Torrens, with Joseph Wachelder, The Enlightenment of Thomas Beddoes: Science, Medicine, and Reform. Abingdon and New York: Routledge, 2017. Pp. 263. ISBN 978-1-4724-8829-

9. £110.00 (hardcover). British Journal for the History of Science, 51(1):162–164, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/trevor-
levere-larry-stewart-and-
hugh-torrens-with-joseph-
wachelder-the-
enlightenment-of-thomas-
beddoes-science-medicine-
and-reform-abingdon-and-
new-york-routledge-2017-pp-
263-isbn-9781472488299-11000-
hardcover/76E1432C4CD1E6A18B375329609F79DD.

Finnegan:2015:CSC

[FW15]  Diarmid  A.

Finnegan and Jonathan Jeffrey Wright. Catholics, science and civic culture in Victorian Belfast. British Journal for the History of Science, 48(2):261–287, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fyfe:2012:SPK

[Fyf12]  Aileen Fyfe.  Steam-powered

knowledge: William Chambers and the business of publishing, 1820–1860. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-27651-1 (hardcover). xvi + 313 pages pp. LCCN Z325.C44 F94 2012.

Gabriel:2016:PPT

[Gab16] Joseph M. Gabriel. Pharmaceutical patenting and the transformation of American medical ethics. British Journal for the History of Science, 49(4): 577–600, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        F716579F8528CE4E2DA2AB42AC42BB52.

Gainty:2017:ICS

[Gai17] Caitjan Gainty. ‘Items for criticism (not in sequence)’: Joseph DeLee, Pare Lorentz and The Fight for Life (1940). British Journal for the History of Science, 50(3):429–449, September 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        CB4C265FBB59158047E8CC60C885CC6D.

Galison:2012:BSF

[Gal12]  Peter Galison.  Blacked-out

spaces: Freud, censorship and the re-territorialization of mind. British Journal for the History of Science, 45(2):235–266, June 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gambarotto:2017:LON

[Gam17] Andrea Gambarotto. Lorenz Oken (1779–1851): Naturphilosophie and the reform of natural history. British Journal for the History of Science, 50 (2):329–340, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
AA5EBBBE4ED2FDBBA7E23C4466D854C7.

See erratum [Gam18].

Gambarotto:2018:LON

[Gam18] Andrea Gambarotto. Lorenz Oken (1779–1851): Naturphilosophie and the reform of natural history — Erratum. British Journal for the History of Science, 51 (3):511, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/lorenz-
oken-17791851-
naturphilosophie-and-the-
reform-of-natural-history-
erratum/
8DC2776284E191EB99A53928C9E184A8.

See [Gam17].

Garber:2009:LBS

[Gar09a] Daniel Garber. Leibniz: body, substance, monad. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956664-X. xxi + 428 pp. LCCN B2599.M8 .G37 2009.

Garber:2009:LHT

[Gar09b] Daniel Garber. Leibniz: body, substance, monad. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956664-X. xxi + 428 pp. LCCN B2599.M7 .G37 2009.

Garbutt:2011:BRA

[Gar11] Gwyndaf Garbutt. Book review: Andrew Cunningham and Sachiko Kusukawa (eds.), Natural Philosophy Epitomised: Books 8–11 of Gregor Reisch’s Philosophical Pearl (1503). Farnham: Ashgate, 2010. Pp. lxxiv + 346. ISBN 978-0-7546-0612-3. £65.00 (hardback). British Journal for the History of Science, 44(3): 459–461, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Garrisi:2014:BRJ

[Gar14]  Diana  Garrisi.

Book reviews: Julia Allen, Swimming with Dr Johnson and Mrs Thrale: Sport, Health and Exercise in Eighteenth-Century England. Cambridge: The Lutterworth Press, 2012. Pp. xii + 297. ISBN 978-0-7188-9276-0. £20.00 (paperback). British Journal for the History of Science, 47(2):376– 378, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gascoigne:2010:BRP

[Gas10]  John Gascoigne.  Book re-

view: Palmira Fontes Da Costa, The Singular and the Making of Knowledge at the Royal Society of London in the Eighteenth Century. Newcastle-upon-Tyne: Cambridge Scholars Publishing, 2009. Pp. xvi + 214. ISBN 978-1-4438-0357-1. £39.99 (hardback). British Journal for the History of Science, 43(1):124– 125, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gattei:2017:GLC

[Gat17] Stefano Gattei. Galileo’s legacy:

a critical edition and translation of the manuscript of Vincenzo Viviani’s Grati Animi Monumenta. British Journal for the History of Science, 50(2): 181–228, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A6853F5702F4ECDCFD250303D84045DC.

GomezRodriguez:2009:CF

[GC09] Amparo Gómez Rodríguez and Antonio Francisco Canales Serrano, editors. Ciencia y fascismo, volume 101 of Laertes. Laertes, Madrid, Spain, 2009. ISBN 84-7584-657-2. 358 pp. LCCN Q127.S7 C4954 2009.

Grell:2010:CME

[GCA10] Ole Peter Grell, Andrew Cunningham, and Jon Arrizabalaga, editors. Centres of medical excellence?: medical travel and education in Europe, 1500–1789. The history of medicine in context. Ashgate, Aldershot, Hants, England and Farnham, England, 2010. ISBN 0-7546-6699-9 (hardcover). xiii + 335 pp. LCCN R771 CEN 2010.

Geppert:2012:IOS

[Gep12] Alexander C. T. Geppert, editor. Imagining outer space: European astroculture in the twentieth century. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2012. ISBN 0-230-23172-1 (hardcover). xvii + 393 pp. LCCN TL789.6.E9 I43 2012.

Gere:2009:KPM

[Ger09] Cathy Gere. Knossos and the prophets of modernism. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-28953-2 (cloth). x + 277 pp. LCCN DF261.K55 G47 2009.

Gradmann:2009:LDR

[GF09] Christoph Gradmann and Elborg Forster. Laboratory disease: Robert Koch’s medical bacteriology. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9313-5 (hardcover). 318 pp. LCCN QR46 .G6713 2009. Translation of: Krankheit im Labor: Robert Koch und die medizinische Bakteriologie / Christoph Gradmann. c2005.

Grattan-Guinness:2010:RES

[GG10] Ivor Grattan-Guinness. The role of an editor: some remarks on Whiteside’s edition of Newton’s mathematical papers. British Journal for the History of Science, 43(1):105–112, March 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grattan-Guinness:2012:BRK

[GG12] I. Grattan-Guinness. Book review: Karin Reich and Elena Roussanova (eds.), Carl Friedrich

Gauss  und  Russland:

Sein Briefwechsel mit in Russland wirkenden Wissenschaftlern. Berlin and Boston: De Gruyter, 2012. Pp. xxiii + 905. ISBN 978-3-11-025306-1. EUR149.95 (hardback). British Journal for the History of Science, 45(3):467– 468, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gaudilliere:2013:WRD

[GH13] Jean-Paul Gaudillière and Volker Hess, editors. Ways of regulating drugs in the 19th and 20th centuries. Science, technology and medicine in modern history. Palgrave MacMillan, Houndmills, Basingstoke, Hampshire, 2013. ISBN 0-230-30196-7. xiv + 327 pp. LCCN RS189 .W39 2013.

Gibbs:2014:BRN

[Gib14]  Fred Gibbs.  Book review:

Nancy G. Siraisi, Communities of Learned Experience: Epistolary Medicine in the Renaissance. Baltimore: Johns Hopkins University Press, 2013. Pp. xii + 163. ISBN 978-1-4214-0749-4. £23.50 (hardback). British Journal for the History of Science, 47(1):178– 179, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gibson:2015:AVM

[Gib15] Susannah Gibson. Animal, vegetable, mineral?: how eighteenthcentury science disrupted the natural order. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-870513-1 (hardcover). xv + 215 pp. LCCN QH83 .G53 2015.

Giffard:2016:BRE

[Gif16] Hermione Giffard. Book review: Erik M. Conway, Exploration and Engineering: The Jet Propulsion Laboratory and the Quest for Mars. Baltimore: Johns Hopkins University Press, 2015. Pp. 416. ISBN 978-1-4214-1604-5. $34.95 (hardback). British Journal for the History of Science, 49(4): 517–518, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        33814289BDD3BFFC35E8870339C00061.

Gigante:2009:LOF

[Gig09] Denise Gigante. Life: organic form and Romanticism. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-13685-

4. xiii + 302 + 16 pp. LCCN

PR575.L54 G54 2009.

Gillispie:2011:BRM

[Gil11] Charles C. Gillispie. Book review: Michael R. Lynn, The Sublime Invention: Ballooning in Europe, 1783–1820. London: Pickering and Chatto, 2010. Pp. ix + 240. ISBN 978-1-84893-016-2. £60.00 (hardback). British Journal for the History of Science, 44(1): 130–131, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gillies:2012:BRE

[Gil12] Donald Gillies. Book review: Erling Norrby, Nobel Prizes and Life Sciences. London and Singapore: World Scientific Press, 2010. Pp. xvi + 317. ISBN 978-981-4299-37-4. £25.00 (paperback). British Journal for the History of Science, 45(1):142–143, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Nor10].

Gillan:2015:BRS

[Gil15a] Caroline Gillan. Book reviews:

Sue Ann Prince (ed.), Of Elephants & Roses: French Natural History, 1790–1830. Philadelphia: American Philosophical Society, 2013. Pp. xxvi + 268. ISBN 978-0-87169-267-2. $50.00 (paperback). British Journal for the History of Science, 48(3): 514–515, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gillott:2015:SBA

[Gil15b] David Gillott. Samuel Butler against the professionals. Rethinking Lamarckism 1860–1900. Legenda, Oxford, UK, 2015. ISBN 1-909662-25-9 (hardcover). 208 pp. LCCN ????

Gillan:2016:BRM

[Gil16a] Caroline Gillan. Book review:

Marie-Christine Skuncke, Carl Peter Thunberg: Botanist and Physician — Career-Building across the Oceans in the Eighteenth Century. Uppsala: Swedish Collegium for Advanced Study, 2014. Pp. 376. ISBN 978-91-981948-0-7. Kr 275.00 (hardback). British Journal for the History of Science, 49(3):484– 485, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
CFD900E2C7B111FDF956F7A077DC3DAB.

Gillin:2016:BRV

[Gil16b] Edward J. Gillin. Book review: Vicky Albritton and Fredrik Albritton Jonsson, Green Victorians: The Simple Life in John Ruskin’s Lake District. Chicago: The University of Chicago Press, 2016. Pp. 209. ISBN 978-0-226-33998-6. £28.00/ $40.00 (cloth). British Journal for the History of Science, 49(4): 649–651, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
D7958C07BBE8231CF97A579839537603.

Gissis:2019:BRC

[Gis19]  Snait Gissis.  Book review:

Charles Wolfe, La philosophie de la biologie avant la biologie: Une histoire du vitalisme. Paris: Classiques Garnier, 2019. Pp. 514. ISBN 978-2-4060-8072-5. EUR93.00 (hardback). British Journal for the History of Science, 52(4):719– 721, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/charles-
wolfe-la-philosophie-de-la-

biologie-avant-la-biologie-
une-histoire-du-vitalisme-
paris-classiques-garnier-
2019-pp-514-isbn-
9782406080725-9300-
hardback/0841BB6D42D126D161252AB335FDE4A1.

Gerbino:2009:CRAa

[GJ09a] Anthony Gerbino and Stephen Johnston. Compass and rule: architecture as mathematical practice in England, 1500–1750. Museum of the History of Science, Oxford, UK, 2009. ISBN 0-300-15093-8 (hardcover). 208 pp. LCCN NA1995 .G47 2009.

Gerbino:2009:CRAb

[GJ09b] Anthony Gerbino and Stephen Johnston. Compass and rule: architecture as mathematical practice in England, 1500–1750. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-15093-8. 208 pp. LCCN NA1995 .G47 2009. Published to accompany the exhibition at the Museum of the History of Science, Oxford from 16 June to 6 September 2009 and the Yale Center for British Art, New Haven, from 18 February to 30 May 2010.

Glasner:2009:APT

[Gla09] Ruth Glasner. Averroës’ physics:

a turning point in medieval natural philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956773-5 (hardcover). ix + 229 pp. LCCN QC7 .G65 2009.

Gliboff:2008:HGB

[Gli08] Sander Gliboff. H. G. Bronn, Ernst Haeckel, and the origins of German Darwinism: a study in translation and transformation. Transformations. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-07293-9 (hardcover). xii + 259 pp. LCCN QH366.2 GLI 2008.

Gliboff:2011:BRM

[Gli11] Sander Gliboff. Book review: Michal Simunek, Uwe Hoßfeld, Olaf Breidbach and Miklós Müller (eds.), Mendelism in Bohemia and Moravia, 1900–1930. Stuttgart: Franz Steiner Verlag, 2010. Pp. 276. ISBN 978-3-515-09602-7. EUR48.00 (paperback). British Journal for the History of Science, 44(4):611–612, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gliboff:2012:BRM

[Gli12] Sander Gliboff. Book reviews: Michal Simunek, Uwe Hoßfeld, Florian Thümmler and Olaf Breidbach (eds.), The Mendelian Dioskuri: Correspondence of Armin with Erich von Tschermak-Seysenegg, 1898–1951. Studies in the History of Sciences and Humanities 27. Prague: Institute of Contemporary History of the Academy of Sciences, Prague, and Department of Genetics/‘Mendelianum’ of the Moravian Museum, Brno, 2011. Pp. 259. ISBN 978-80-87378-67-0. Price unknown (paperback). Michal Simunek, Uwe Hoßfeld, Florian Thümmler, and Jirí Sekerák (eds.), The Letters on G. J. Mendel: Correspondence of William Bateson, Hugo Iltis, and Erich von Tschermak-Seysenegg with Alois and Ferdinand Schindler, 1902– 1935. Studies in the History of Sciences and Humanities 28. Prague: Institute of Contemporary History of the Academy of Sciences, Prague, and Department of Genetics/‘Mendelianum’ of the Moravian Museum, Brno, 2011. Pp. 131. ISBN 978-80-87378-73-1. Price unknown (paperback). British Journal for the History of Science, 45(2):303– 305, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gee:2014:FWD

[GMML14] Brian Gee, Anita McConnell, and A. D. Morrison-Low, editors. Francis Watkins and the Dollond Telescope Patent Controversy. Science, technology, and culture. Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4094-6643-4 (hardcover), 1-4094-6644-2 (e-book), 1-4094-6645-0 (epub). ISSN 1700-1945. xxvi + 392 pp. LCCN QB88 .G44 2014.

Gausemeier:2013:HHT

[GMWR13] Bernd Gausemeier, Staffan Müller-Wille, and Edmund Ramsden, editors. Human heredity in the Twentieth Century, volume 15 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-426-2 (hardcover), 1-78144-051-4 (e-book). 256 pp. LCCN QH.

Goldstein:2010:HCE

[Gol10] Jan Goldstein. Hysteria complicated by ecstasy: the case of Nanette Leroux. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-01186-9 (hardcover). xi + 246 pp. LCCN RC532 2010.

Gooday:2008:DET

[Goo08] Graeme Gooday. Domesticating electricity: technology, uncertainty and gender, 1880–1914, volume 7 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2008. ISBN 1-85196-975-6. viii + 292 pp. LCCN HD9685.A2 G66 2008.

Gooday:2010:BRJ

[Goo10a] Graeme J. N. Gooday. Book review: Julie Anderson Francis Neary and John V. Pickstone, Surgeons, Manufacturers and Patients: A Transatlantic History of Total Hip Replacement. Basingstoke and New York: Palgrave Macmillan, 2007. Pp. xiv + 222. ISBN 0-230-55314-1. £45.00 (hardback). British Journal for the History of Science, 43(1):139– 141, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goodrum:2010:BRS

[Goo10b] Matthew R. Goodrum. Book review: Sigrid Schmalzer, The People’s Peking Man: Popular Science and Human Identity in Twentieth-Century China. Chicago and London: University of Chicago Press, 2008. Pp. xix + 346. ISBN 978-0-226-73860-4. $26.00 (paperback). British Journal for the History of Science, 43 (1):132–134, March 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gooday:2011:BRB

[Goo11a] Graeme Gooday. Book review:

Basil Mahon, Oliver Heaviside: Maverick Mastermind of Electricity. London: Institution of Engineering and Technology Press, 2009. Pp. xvii + 183. ISBN 978-0-86341-965-2. £15.00/ $30.00 (paperback). British Journal for the History of Science, 44(1): 138–140, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goodey:2011:HII

[Goo11b] C. F. Goodey. A history of intelligence and ‘intellectual disability’: the shaping of psychology in early modern Europe. Ashgate, Aldershot, Hants, England and Farnham, England, 2011. ISBN 1-4094-2021-3 (hardcover), 1-4094-2022-1 (e-book). 381 pp. LCCN BF441 .G656 2011.

Gooday:2012:BRM

[Goo12a] Graeme Gooday. Book review: Mario Biagioli, Peter Jaszi and Martha Woodmansee (eds.), Making and Unmaking Intellectual Property: Creative Production in Legal and Cultural Perspective. Chicago and London: University of Chicago Press, 2011. Pp. vii + 466. ISBN 978-0-226-90709-3. £26.00 (paperback). British Journal for the History of Science, 45(2):312– 313, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goodrum:2012:BRB

[Goo12b] Matthew R. Goodrum. Book review: B. Ricardo Brown, Until Darwin: Science, Human Variety and the Origin of Race. London: Pickering & Chatto, 2010. Pp. ix + 199. ISBN 978-1-84893-100-8. £60.00 (hardback). British Journal for the History of Science, 45 (1):135–136, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Bro10a].

Goodrum:2016:BHP

[Goo16] Matthew R. Goodrum. The beginnings of human palaeontology: prehistory, craniometry and the ‘fossil human races’. British Journal for the History of Science, 49(3):387–409, September 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
2AE212CF529CCE6E89113B800D1C8BB3.

Goodman:2017:BRD

[Goo17] Matthew Goodman. Book review: Diarmid A. Finnegan and Jonathan Jeffrey Wright (eds.), Spaces of Global Knowledge: Exhibition, Encounter and Exchange in an Age of Empire. Farnham: Ashgate, 2015. Pp. 300. ISBN 978-1-47244436-

3. £63.00 (hardback). British Journal for the History of Science, 50(1):151–153, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
4E61F390B34C3A191BB64FD85168163C.

Gooday:2018:BRA

[Goo18] Graeme Gooday. Book review:

Andreas Marklund and Mogens Rüdiger (eds.), Historicizing Infrastructure. Aalborg: Aalborg University Press, 2017. Pp. 235. ISBN 978-87-7112-594-8. DKr 298.00 (hardcover). British Journal for the History of Science, 51(2):324–326, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/andreas-
marklund-and-mogens-
rudiger-eds-historicizing-
infrastructure-aalborg-
aalborg-university-press-
2017-pp-235-isbn-
9788771125948-dkr-29800-
hardcover/
0A7C296804D3CF7174F596B697F24C7E.

Gordin:2012:PWI

[Gor12] Michael D. Gordin. The pseudoscience wars: Immanuel Velikovsky and the birth of the modern fringe. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-30442-6 (hardcover), 0-226-30443-4 (e-book). x + 291 pp. LCCN Q172.5.P77 G674 2012.

Goulding:2010:DHR

[Gou10] Robert Goulding. Defending Hypatia: Ramus, Savile, and the Renaissance rediscovery of mathematical history, volume 25 of Archimedes, 1385-0180. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3541-9 (hardcover), 90-481-3542-7 (e-book). xx + 201 pp. LCCN QA24 .G68 2010.

Gouk:2012:BRM

[Gou12a] Penelope Gouk. Book review: Maria Semi, Music as a Science of Mankind in Eighteenth-Century Britain. Farnham: Ashgate, 2012. Pp. x + 185. ISBN 978-1-4094-2868-8. £55.00 (hardback). British Journal for the History of Science, 45(3):462– 463, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouyon:2012:BRD

[Gou12b] Jean-Baptiste Gouyon. Book review: David A. Kirby, Lab Coats in Hollywood: Science, Scientists, and Cinema. Cambridge, MA and London: MIT Press, 2011. Pp. xiv + 265. ISBN 978-0-262-01478-6. £20.95 (hardback). British Journal for the History of Science, 45(1):146–148, March 2012.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Kir11].

Gouyon:2012:BRM

[Gou12c] Jean-Baptiste Gouyon. Book review: Michael H. Graham, Joan Parker and Paul K. Dayton (eds.), The Essential Naturalist: Timeless Readings in Natural History. Chicago and London: The University of Chicago Press, 2011. Pp. xiii + 534. ISBN 978-0226-30570-7. $39.00 (paperback). British Journal for the History of Science, 45(3):451– 453, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouk:2013:BRB

[Gou13a] Penelope Gouk. Book reviews: Bad Vibrations: The History of the Idea of Music as a Cause of Disease. Farnham: Ashgate, 2012. Pp. xii + 213. ISBN 978-1-4094-2642-4. £65.00 (hardback), by James Kennaway. British Journal for the History of Science, 46(3):525–526, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouk:2013:BRC

[Gou13b] Penelope Gouk. Book reviews:

C. U. M. Smith, Eugenio Frixione, Stanley Finger and William Clower, The Animal Spirit Doctrine and the Origins of Neurophysiology. Oxford: Oxford University Press, 2012. Pp. xiv + 277. ISBN 978-0-19-976649-9.

£75.00 (hardback). British Journal for the History of Science, 46 (1):161–162, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouyon:2014:BRK

[Gou14] Jean-Baptiste Gouyon. Book review: Kristian H. Nielsen, Michael Harbsmeier and Christopher J. Ries (eds.), Scientists and Scholars in the Field: Studies in the History of Fieldwork and Expeditions. Aarhus: Aarhus University Press, 2012. Pp. 476. ISBN 978-87-7124-014-6. £50.00 (hardback). British Journal for the History of Science, 47(1):188– 189, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouk:2016:BRJ

[Gou16] Penelope Gouk. Book reviews: Jacomien Prins, Echoes of an Invisible World: Marsilio Ficino and Francesco Patrizi on Cosmic Order and Music Theory. Leiden: Brill, 2015. Pp. xiv + 461. ISBN 978-9-0042-7437-2. EUR49.00 (hardback). British Journal for the History of Science, 49(1):119–120, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouk:2017:BRA

[Gou17]  Penelope Gouk.  Book re-

view: Andrew Hicks, Composing the World Harmony in the Medieval Platonic Cosmos. New York: Oxford University Press, 2017. Pp. xviii + 321. ISBN 978-0-19-065820-5. £29.99 (hardback). British Journal for the History of Science, 50(3):545– 547, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
81C3AAC7F072627C26E6F086EB6926BC.

Gamez-Perez:2018:BRE

[GP18]  Carlos Gámez-Pérez.  Book

review: Emily B. Stanback, The Wordsworth–Coleridge Circle and the Aesthetics of Disability. London: Palgrave Macmillan, 2016. Pp. xv + 337. ISBN 978-1-137-51139-3. £90.00 (hardcover). British Journal for the History of Science, 51(4):708–709, December 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/emily-b-stanback-
the-wordsworthcoleridge-
circle-and-the-aesthetics-
of-disability-london-
palgrave-macmillan-2016-pp-
xv-337-isbn-9781137511393-
9000-hardcover/
8C396348B8CA3D1F0806A6DEE05227BC.

Gil-Riano:2018:RAR

[GR18] Sebastián Gil-Riaño. Relocating anti-racist science: the 1950 UNESCO Statement on Race and economic development in the global South. British Journal for the History of Science, 51 (2):281–303, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/relocating-
                                        antiracist-science-the-
                                        1950-unesco-statement-on-
                                        race-and-economic-
                                        development-in-the-global-
                                        south/
                                        83DC153C208EFC078BAEAEE1130BCA67.

Grant:2011:BRP

[Gra11] Florence Grant. Book review: Peter Heering, Oliver Hochadel and David J. Rhees (eds.), Playing with Fire: Histories of the Lightning Rod. Philadelphia: American Philosophical Society, 2009. Pp. xi + 290. ISBN 978-1-60618-995-5. $35.00 (paperback). British Journal for the History of Science, 44(1):127–128, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gray:2013:HPS

[Gra13] Jeremy Gray. Henri Poincaré: a scientific biography. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-15271-3 (hardback). xiii + 592 pp. LCCN Q143.P7 G73 2013.

Grant:2015:MEM

[Gra15] Florence Grant. Mechanical experiments as moral exercise in the education of George III. British Journal for the History of Science, 48(2):195–212, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gregory:2014:WHA

[Gre14a] Andrew Gregory. William Harvey, Aristotle and astrology. British Journal for the History of Science, 47(2):199–215, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grell:2014:BRD

[Gre14b] Ole Peter Grell. Book reviews: Daniel Stolzenberg, Egyptian Oedipus: Athanasius Kircher and the Secrets of Antiquity. Chicago and London: The University of Chicago Press, 2013. Pp. xi + 307. ISBN 978-0-226-92414-4. £32.50 (hardback). British Journal for the History of Science, 47(3):573–574, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grinnell:2015:BRH

[Gri15] Frederick Grinnell. Book reviews: Harry Collins, Are We All Scientific Experts Now? Cambridge: Polity Press, 2014. Pp. vi + 144. ISBN 978-0-7456-8204-4. £9.99 (paperback). British Journal for the History of Science, 48(3):540–541, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gross:2009:HHM

[Gro09] Charles G. Gross. A hole in the head: more tales in the history of neuroscience. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-01338-X. x + 356 pp. LCCN WL11.1 2009; RC338 .G76 2010.

Grossman:2017:JDO

[Gro17] Mark I. Grossman. John Dalton and the origin of the atomic theory: reassessing the influence of Bryan Higgins. British Journal for the History of Science, 50(4): 657–676, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
17EA9670B7CD71428E71C4556B4D887D.

Gooday:2008:WSS

[GS08] Graeme Gooday and James Sumner, editors. By Whose Standards? Standardization, Stability and Uniformity in the History of Information and Electrical Technologies. History of Technology. Continuum, London, UK, 2008. ISBN 0-8264-3875-X. xiv + 171 pp. LCCN ????

Gradmann:2010:EST

[GS10] Christoph Gradmann and Jonathan Simon, editors. Evaluating and standardizing therapeutic agents, 1890–1950. Science, technology and medicine in modern history. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-20281-0 (hardback). xiv + 266 pp. LCCN RS189 .E88 2010.

Garcia-Sancho:2012:BCH

[GS12] Miguel García-Sancho. Biology, computing, and the history of molecular sequencing: from proteins to DNA, 1945– 2000. Science, technology and medicine in modern history. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2012. ISBN 0-230-25032-7 (hardcover). xiii + 242 pp. LCCN ???? £55.00.

Graham-Smith:2013:UCUa

[GS13] Francis Graham-Smith. Unseen Cosmos: the Universe in Radio. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966058-1 (hardcover). vii + 238 + 8 pp. LCCN QB477 .G733 2013; QB477 .G74 2013.

Gomez:2018:E

[GS18] Pablo F. Gómez and Sujit Sivasundaram. Epilogue. British Journal for the History of Science, 51 (4):679–686, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/epilogue/
A538C718E58AB1604A6EB8928CA4EC9B.

Glennie:2009:SDH

[GT09] Paul Glennie and N. J. Thrift.

Shaping the day: a history of timekeeping in England and Wales 1300–1800. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-927820-2. xiv + 456 pp. LCCN QB213 .G54

2009. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0907/2008045775-b.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0907/
                                        2008045775-d.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0907/
                                        2008045775-t.htm.

Guessoum:2011:IQQ

[Gue11] Nidhal Guessoum. Islam’s quantum question: reconciling Muslim tradition and modern science. I. B. Tauris, London, UK, 2011. ISBN 1-84885-517-6 (hardcover), 1-84885-518-4 (paperback). 403 pp. LCCN P190.5.S3 G84 2011.

Guerrini:2019:RUP

[Gue19]  Anita Guerrini.  Retrospec-

tives: Unconventional paths. British Journal for the History of Science, 52(4):696–706, December 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/retrospectives-
                                        unconventional-paths/
                                        4E4BB3534242D750888399490CF7CC3A.

Guicciardini:2009:INM

[Gui09] Niccolò Guicciardini. Isaac Newton on mathematical certainty and method. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-01317-7 (hardcover). xxiii + 422 pp. LCCN QA29.N4 G85 2009.

Gulizia:2018:BRJ

[Gul18] Stefano Gulizia. Book review:

Joyce van Leeuwen, The Aristotelian Mechanics: Text and Diagrams. Berlin: Springer, 2016. Pp. 258. ISBN 978-3-3192-5925-

3. $119.00 (hardcover). British Journal for the History of Science, 51(3):513–515, September 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/joyce-van-leeuwen-
the-aristotelian-mechanics-
text-and-diagrams-berlin-
springer-2016-pp-258-isbn-
9783319259253-11900-
hardcover/
460980C8C6FFADA937FBFFED173398A9.

Gill:2009:GWK

[GWW09] Christopher Gill, Tim Whitmarsh, and John Wilkins, editors. Galen and the world of knowledge, Greek culture in the Roman world. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-76751-2 (hardcover). LCCN B577.G24 G345 2009. Based on a conference held at the University of Exeter in 2005.

Hess:2008:CS

[HA08] Peter M. J. Hess and Paul L.

Allen. Catholicism and science. Greenwood guides to science and religion. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2008. ISBN 0-313-33190-1. xxvi + 241 pp. LCCN BX1795.S35 H47 2008. URL

                                        http://www.loc.gov/catdir/
                                        toc/ecip081/2007039200.html.

Horrocks:2012:VSS

[HA12] Jeremiah Horrocks and Wilbur Applebaum. Venus seen on the sun: the first observations of a transit of Venus, volume 29 / 18 of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-22193-X (hardcover), 90-04-22194-8 (e-book). xxiv + 82 pp. LCCN QB509.5 .H6713 2012.

Hill-Andrews:2016:BRL

[HA16] Oliver Hill-Andrews. Book review: Loren Graham, Lysenko’s Ghost: Epigenetics and Russia. Cambridge, MA: Harvard University Press, 2016. Pp. 200. ISBN 978-0-674-08905-1. £18.95 (hardback). British Journal for the History of Science, 49(4): 511–512, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        47C91EA18AA95602D0F8DE90A55B6F7C.

Hill-Andrews:2019:NHT

[HA19a] Oliver Hill-Andrews. ‘A new and hopeful type of social organism’: Julian Huxley, J. G. Crowther and Lancelot Hogben on Roosevelt’s New Deal. British Journal for the History of Science, 52 (4):645–671, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/new-and-
hopeful-type-of-social-
organism-julian-huxley-jg-
crowther-and-lancelot-
hogben-on-roosevelts-new-
deal/
10C0143F70BCC9446C4CDCE00E1AC594.

Honeybun-Arnolda:2019:BRS

[HA19b] Elliot Honeybun-Arnolda. Book review: Susanna Gibson, The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford: Oxford University Press, 2019. Pp. ix + 363. ISBN 978-0-1988-3337-6. £25.00 (hardback). British Journal for the History of Science, 52(2):367– 369, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/susanna-gibson-the-
spirit-of-inquiry-how-one-
extraordinary-society-
shaped-modern-science-
oxford-oxford-university-
press-2019-pp-ix-363-isbn-
9780198833376-2500-
hardback/58DDAC56AF1634ECEE4D1B403FDE634C.

Hacking:2014:WTP

[Hac14] Ian Hacking. Why is there philosophy of mathematics at all? Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-05017-0, 1-107-65815-2. xv + 290 pp. LCCN QA8.4 .H29

                                        2014. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1408/2014395332-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1408/
                                        2014395332-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1408/
                                        2014395332-t.html.

Hahn:2016:BRL

[Hah16] Andre Michael Hahn. Book review: Lynn Gamwell, Mathematics and Art: A Cultural History. Princeton, NJ: Princeton University Press, 2016. Pp. 576. ISBN 978-0-691-16528-8. £34.95 (hardback). British Journal for the History of Science, 49(4): 524–525, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        3FFC21B7CDFE87BE44DB929C087D072F.

Haines:2016:BRH

[Hai16]  Elizabeth Haines.  Book re-

views: Horst Bredekamp, Vera Dünkel and Birgit Schneider (eds.), The Technical Image: A History of Styles in Scientific Imagery. Chicago: The University of Chicago Press, 2015. Pp. 208. ISBN 978-0-2262-5884-3. £35.00 (hardback). British Journal for the History of Science, 49(1):146– 147, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Halford:2010:BRH

[Hal10] Jacob Halford. Book review: Helge S. Kragh, Peter C. Kjaergaard, Henry Nielsen and Kristian Hvidfelt Nielsen, Science in Denmark: A Thousand Year History. Aarhus: Aarhus University Press, 2008. Pp. 607. ISBN 978-87-7934-317-7. £37.95 (hardback). British Journal for the History of Science, 43(1):131– 132, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hale:2014:PDM

[Hal14a] Piers J. Hale. Political descent:

Malthus, mutualism, and the politics of evolution in Victorian England. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-10849-X (hardcover), 0-226-10852-X (e-book). 442 pp. LCCN JC336 .H254 2014.

Hall:2014:BRS

[Hal14b] Alexander Hall. Book review:

Spencer R. Weart, The Rise of Nuclear Fear. Cambridge, MA and London: Harvard University Press, 2012. Pp. ix + 367. ISBN 978-0-674-05233-8. £16.95 (paperback). British Journal for the History of Science, 47(1):193– 194, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hall:2014:MMC

[Hal14c] Kersten T. Hall. The man in the monkeynut coat: William Astbury and the forgotten road to the double-helix. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-870459-3 (hardcover), 0-19-100988-1 (e-book). ix + 242 pp. LCCN QH506 .H35 2014.

Halpern:2014:BDH

[Hal14d] Orit Halpern. Beautiful data: a history of vision and reason since 1945. Experimental futures. Duke University Press, Durham, NC, USA, 2014. ISBN 0-8223-5730-5 (hardcover), 0-8223-5744-5 (paperback). ix + 342 pp. LCCN Q310 .H35 2014.

Hall:2017:BRP

[Hal17] James R. Hall. Book review: Peter Hobbins, Venomous Encounters: Snakes, Vivisection and Scientific Medicine in Colonial Australia. Manchester: Manchester University Press, 2017. Pp. xiii + 202. ISBN 978-1-5261-0144-0. £70.00 (hardback). British Journal for the History of Science, 50(3):552–554, September 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        237E6C83A71C19EB6A1622C2D8366055.

Hamlin:2009:CB

[Ham09] Christopher Hamlin. Cholera: the biography. Biographies of disease. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954624-X (hardcover). x + 344 pp. LCCN RA644.C3 H28 2009.

Hamlin:2010:BRW

[Ham10] Christopher Hamlin. Book review: Warwick Anderson, The Collectors of Lost Souls: Turning Kuru Scientists into Whitemen. Baltimore: Johns Hopkins University Press, 2008. Pp. 318. ISBN 978-0-8018-9040-6. £13.00 (hardback). British Journal for the History of Science, 43(1):141– 142, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hamilton:2014:BRC

[Ham14a] Catriona Gilmour Hamilton.

Book reviews: Carsten Timmermann and Elizabeth Toon (eds.), Cancer Patients, Cancer Pathways: Historical and Sociological Perspectives. Basingstoke: Palgrave Macmillan, 2012. Pp. xii + 270. ISBN 978-1-137-27207-2. £55.00 (hardback). British Journal for the History of Science, 47(2):389–391, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hammond:2014:BRP

[Ham14b] Andy Hammond. Book reviews: Peter Ayres, Shaping Ecology: The Life of Arthur Tansley. Chichester: Wiley-Blackwell, 2012. Pp. xii + 213. ISBN 978-0-470-67154-2. £19.99 (paperback). British Journal for the History of Science, 47(2):383– 384, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hammond:2015:BRM

[Ham15] Andy Hammond. Book reviews:

Maurizio Esposito, Romantic Biology, 1890–1945. London: Pickering & Chatto, 2013. Pp. viii + 257. ISBN 978-1-8489-3420-7 [not 978-1-84893-430-6]. £60.00 (hardback). British Journal for the History of Science, 48(2): 377–378, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hamza:2016:BRA

[Ham16] Shireen Hamza. Book review: Anna Winterbottom and Facil Tesfaye (eds.), Histories of Medicine and Healing in the Indian Ocean World, vol. 1: The Medieval and Early Modern Period. Basingstoke: Palgrave MacMillan, 2016. Pp. 220. ISBN 978-1-137-56760-5. £63.00 (hardback). British Journal for the History of Science, 49(4): 643–644, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        7EC692BCE931810C2A9B014D670ACC7A.

Hanson:2015:TSH

[Han15] Thor Hanson. The triumph of seeds: how grains, nuts, kernels, pulses, and pips, conquered the plant kingdom and shaped human history. Basic Books, a member of the Perseus Books Group, New York, NY, USA, 2015. ISBN 0-465-05599-0 (hardcover), 0-465-04872-2 (e-book). xxv + 277 pp. LCCN QK661 .H36 2015.

Hanley:2016:BRG

[Han16] Anne Hanley. Book review: Graham Mooney, Intrusive Interventions: Public Health, Domestic Space, and Infectious Disease Surveillance in England, 1840– 1914. Rochester, NY: University of Rochester Press, 2015. Pp. 292. ISBN 978-1-58046-527-4. £80.00 (hardback). British Journal for the History of Science, 49(4): 505–507, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
5948CBDCEF202CB9ED2260BAB96D6609.

Handley:2017:BRE

[Han17] Neil Handley. Book review:

Elizabeth Hallam, Anatomy Museum: Death and the Body Displayed. London: Reaktion, 2016. Pp. 444. ISBN 978-1-86189-375-

8. £35.00 (hardback). British Journal for the History of Science, 50(1):154–156, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
FBAA78D69E6F2BA2F918E5644A3409C1.

Harper:2008:WNG

[Har08a] Kristine Harper. Weather by the numbers: the genesis of modern meteorology. Transformations. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-08378-7. ix + 308 pp. LCCN QC855 .H374 2008. URL http:

//www.loc.gov/catdir/toc/
ecip081/2007039860.html.

Harper:2008:SHM

[Har08b] Peter S. Harper. A short history of medical genetics, volume 57 of Oxford monographs on medical genetics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-518750-4, 1-59463-530-7. xi + 557 pp. LCCN RB155 .H357

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip088/
                                        2008002454.html.

Harrington:2008:CWH

[Har08c] Anne Harrington. The cure within: a history of mind-body medicine. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-06563-4. 336 pp. LCCN RC49 .H333 2008.

Harman:2009:CNB

[Har09] P. M. (Peter Michael) Harman. The culture of nature in Britain, 1680–1860. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-15197-7 (hardcover). xi + 393 pp. LCCN BH301.N3 H36 2009.

Harrington:2010:MMH

[Har10a] Anne Harrington. Book review: Mark S. Micale, Hysterical Men: the Hidden History of Male Nervous Illness. Cambridge, MA and London: Harvard University Press, 2008. Pp. xiv + 366. ISBN 978-0-674-03166-1. £22.95 (hardback). British Journal for the History of Science, 43 (4):619–620, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Mic08].

Harwood:2010:HR

[Har10b] Jonathan Harwood. Heredity reconceived? British Journal for the History of Science, 43(2):277– 282, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Harper:2016:BRS

[Har16] Kristine C. Harper. Book reviews: Simone Turchetti and Peder Roberts (eds.), The Surveillance Imperative: Geosciences during the Cold War and Beyond. New York: Palgrave Macmillan, 2014. Pp. xii + 278. ISBN 978-1-137-43872-0. £60.00 (hardback). British Journal for the History of Science, 49(2):315–317, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hayes:2018:BRM

[Hay18]  Emily Hayes.  Book review:

Michael Boulter, Bloomsbury Scientists: Science and Art in the Wake of Darwin. London: UCL Press, 2017. Pp. xxii + 175. ISBN 978-1-7873-5005-2. £35.00 (paperback). British Journal for the History of Science, 51(3): 529–530, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michael-
boulter-bloomsbury-
scientists-science-and-art-
in-the-wake-of-darwin-
london-ucl-press-2017-pp-
xxii-175-isbn-
9781787350052-3500-

paperback/
133F2EAA519DF853A97C0D56DFBDF71E.

Hayes:2019:FLP

[Hay19a] Emily Hayes. Fashioned in the light of physics: the scope and methods of Halford Mackinder’s geography. British Journal for the History of Science, 52(4): 569–594, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/fashioned-
in-the-light-of-physics-
the-scope-and-methods-of-
halford-mackinders-
geography/
04DA149E9EE462EE626F85FD0B8F8EFE.

Hayward:2019:BRA

[Hay19b] Rhodri Hayward. Book review: Anthony Chaney, Runaway: Gregory Bateson, the Double Bind, and the Rise of Ecological Consciousness. Chapel Hill, NC: University of North Carolina Press, 2017. Pp. 304. ISBN 978-1-4696-3173-8. $32.95 (cloth). British Journal for the History of Science, 52(3): 536–537, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/anthony-
chaneyrunaway-gregory-
bateson-the-double-bind-
and-the-rise-of-ecological-
consciousness-chapel-hill-

nc-university-of-north-
carolina-press-2017-pp-304-
isbn-9781469631738-3295-
cloth/CD3B817EC75111A55B486FA67C6BE94E.

Harman:2013:OSR

[HD13]  Oren  Solomon

Harman and Michael R. Dietrich, editors. Outsider scientists: routes to innovation in biology. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07837-X (hardcover), 0-226-07840-X (paperback), 0-226-07854-X (e-book). 374 pp. LCCN QH26 .O98 2013.

Hecht:2011:EGE

[Hec11] Gabrielle Hecht, editor. Entangled geographies: empire and technopolitics in the global Cold War. Inside technology. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-51578-4 (paperback). viii + 337 pp. LCCN D843 .E58 2011.

Hecht:2017:WWW

[Hec17] David K. Hecht. Why we write (nuclear) history. British Journal for the History of Science, 50(3): 537–543, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
D090C00A5BB07F82F84A5C3A0C31601C.

Hedesan:2013:BRD

[Hed13]  Jo Hedesan.  Book reviews:

Descartes et la chimie. Paris: Librairie Philosophique J. Vrin, 2011. Pp. 256. ISBN 978-2-7116-2330-3. EUR24.00 (paperback), by Bernard Joly. British Journal for the History of Science, 46(3): 523–524, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heggie:2011:HBS

[Heg11] Vanessa Heggie. A history of British sports medicine. Manchester University Press, Manchester, UK, 2011. ISBN 0-7190-8261-7 (hardcover). x + 222 pp. LCCN RC1210 .H39 2011. UK£60.00.

Heggie:2013:EPE

[Heg13] Vanessa Heggie. Experimental physiology, Everest and oxygen: from the ghastly kitchens to the gasping lung. British Journal for the History of Science, 46(1):123– 147, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heintzman:2015:BRK

[Hei15] Kit Heintzman. Book reviews: Kathleen Hardesty Doig, From ‘Encyclopédie’ to ‘Encyclopédie Méthodique’: Revision and Expansion. Oxford: Voltaire Foundation, 2013. Pp. xvi + 314. ISBN 978-0-7294-1077-9. £65.00 (paperback). British Journal for the History of Science, 48(3): 513–514, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heintzman:2018:COD

[Hei18] Kit Heintzman. A cabinet of the ordinary: domesticating veterinary education, 1766–1799. British Journal for the History of Science, 51(2):239–260, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/cabinet-of-
the-ordinary-domesticating-
veterinary-education-
17661799/
6021EB7317F4B67F6EBC150842D267E4.

HemmungsWirten:2015:MMC

[Hem15] Eva Hemmungs Wirtén. Making Marie Curie: intellectual property and celebrity culture in an age of information. Science.culture. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-23584-X (hardcover), 0-226-23598-X (ebook). 223 pp. LCCN QD22.C8 H46 2015.

Hempstead:2019:BRD

[Hem19] Colin Hempstead. Book review:

David Cahan, Helmholtz: A Life in Science. Chicago: The University of Chicago Press, 2018. Pp. 937. ISBN 978-0-226-48114-

2. $55.00 (cloth). British Journal for the History of Science, 52(2):372–373, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/david-
cahan-helmholtz-a-life-in-
science-chicago-the-

                                        university-of-chicago-
                                        press-2018-pp-937-isbn-
                                        9780226481142-5500-cloth/
                                        39D02738112029AE7C48D7F5F0EDFA22.

Henry:2010:BRL

[Hen10] John Henry. Book review: Lorraine Daston and Michael Stolleis (eds.), Natural Law and Laws of Nature in Early Modern Europe: Jurisprudence, Theology, Moral and Natural Philosophy. Aldershot: Ashgate, 2008. Pp. xii + 338. ISBN 978-0-7546-5761-3. £60.00 (hardback). British Journal for the History of Science, 43 (1):123–124, March 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hentschel:2011:BRE

[Hen11a] Klaus Hentschel. Book review: Elisabeth Vaupel and Stefan L. Wolff (eds.), Das Deutsche Museum in der Zeit des Nationalsozialismus: Eine Bestandsaufnahme. Göttingen: Wallstein, 2010. Pp. 710. ISBN 978-3-8353-0596-0. EUR39.90. British Journal for the History of Science, 44 (2):306–307, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hentschel:2011:BRH

[Hen11b] Klaus Hentschel. Book review: Helmut Rechenberg, Werner Heisenberg — Die Sprache der Atome. Leben und Wirken — Eine wissenschaftliche Biographie. Die ‘Fröhliche Wissenschaft’ (Jugend bis Nobelpreis). 2 vols. Heidelberg: Springer, 2010. Pp. xxi + 603. ISBN 978-3-540-69221-8. £168.00 (hardback). British Journal for the History of Science, 44(4):612–614, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Henry:2013:BRP

[Hen13a] John Henry. Book reviews: Peter Elmer, The Miraculous Conformist: Valentine Greatrakes, the Body Politic, and the Politics of Healing in Restoration Britain. Oxford: Oxford University Press, 2013. Pp. xiii + 279. ISBN 978-0-19-966396-5. £65.00 (hardback). British Journal for the History of Science, 46(4):718–719, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hentschel:2013:BRJ

[Hen13b] Klaus Hentschel. Book reviews:

Jeremy Gray, Henri Poincaré: A Scientific Biography. Princeton and Oxford: Princeton University Press, 2013. Pp. xiii + 592. ISBN 978-0-691-15271-4. £24.95 (hardback). British Journal for the History of Science, 46(4): 725–726, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hendrickson:2014:BRT

[Hen14] Kenneth E. Hendrickson. Book reviews: Ted R. Anderson, The Life of David Lack: Father of Evolutionary Ecology. Oxford: Oxford University Press, 2013. Pp. x + 246. ISBN 978-0-19-992264-2. £37.50 (hardback).

British Journal for the History of Science, 47(4):753–754, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hendrickson:2015:BRK

[Hen15a] Kenneth E. Hendrickson. Book reviews: Kersten T. Hall, The Man in the Monkeynut Coat: William Astbury and the Forgotten Road to the Double Helix. Oxford: Oxford University Press, 2014. Pp. ix + 242. ISBN 978-0-19-870459-1. £18.99 (hardback). British Journal for the History of Science, 48(2):380–381, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hendrickson:2015:BRR

[Hen15b] Kenneth E. Hendrickson. Book reviews: Raymond Flood, Mark McCartney and Andrew Whitaker (eds.), James Clerk Maxwell: Perspectives on His Life and Work. Oxford: Oxford University Press, 2014, Pp. x + 364. ISBN 978-0-19-966437-5. £39.99. British Journal for the History of Science, 48(3): 520–521, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Henry:2017:BRJ

[Hen17] John Henry. Book review: Jessica Riskin, The Restless Clock: A History of the Centuries-Long Argument over What Makes Living Things Tick. Chicago: The University of Chicago Press, 2016. Pp. 544. ISBN 978-0-226-30292-8. $40.00 (hardback). British Journal for the History of Science, 50(2):345–346, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A01FC9C9399B7C59CCBB670ABAE91223.

Hentschel:2019:BRJ

[Hen19] Klaus Hentschel. Book review:

Joseph D. Martin, Solid State Insurrection: How the Science of Substance Made American Physics Matter. Pittsburgh, PA: University of Pittsburgh Press, 2018. Pp. ix + 280. ISBN 978-0-8229-4538-3. $49.95 (cloth). British Journal for the History of Science, 52(2):379–381, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/joseph-d-
martin-solid-state-
insurrection-how-the-
science-of-substance-made-
american-physics-matter-
pittsburgh-pa-university-
of-pittsburgh-press-2018-
pp-ix-280-isbn-9780822945383-
4995-cloth/B9B8ECDCCE94F5D2EE7D9B245603E923.

Hesketh:2009:AAE

[Hes09] Ian Hesketh. Of apes and ancestors: evolution, Christianity, and the Oxford debate. University of Toronto Press, Toronto, ON, Canada, 2009. ISBN 0-8020-9284-5. viii + 144 pp. LCCN BL263 .H475 2009; BL263 HES 2009.

Hesketh:2011:SHV

[Hes11] Ian Hesketh. The science of history in Victorian Britain: making the past speak, volume 12 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-126-3 (hardcover), 1-84893-127-1 (e-book). xi + 229 pp. LCCN ????

Heukelom:2015:HAP

[Heu15] Floris Heukelom. A history of the Allais paradox. British Journal for the History of Science, 48 (1):147–169, March 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heering:2009:PFH

[HHR09] Peter Heering, Oliver Hochadel, and David J. Rhees, editors. Playing with fire: histories of the lightning rod, volume 99 (part 5) of Transactions of the American Philosophical Society held at Philadelphia for promoting useful knowledge. American Philosophical Society, Philadelphia, PA, USA, 2009. ISBN 1-60618-995-6. xi + 290 pp. LCCN TH9061 .P58 2009.

Hickman:2019:WPG

[Hic19] Clare Hickman. ‘The want of a proper Gardiner’: late Georgian Scottish botanic gardeners as intermediaries of medical and scientific knowledge. British Journal for the History of Science, 52 (4):543–567, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/want-of-a-
proper-gardiner-late-
georgian-scottish-botanic-
gardeners-as-
intermediaries-of-medical-
and-scientific-knowledge/
C28DBFED916D9DA8D25739A00316BBF3.

Higgitt:2010:DAC

[Hig10]  Rebekah Higgitt.  Book re-

view: David Aubin, Charlotte Bigg and H. Otto Sibum (eds.), The Heavens on Earth: Observatories and Astronomy in Nineteenth-Century Science and Culture. Durham, NC and London: Duke University Press, 2010. Pp. xii + 384. ISBN 978-0-8223-4640-1. £16.99 (paperback). British Journal for the History of Science, 43(4):613–614, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Higgitt:2011:BRK

[Hig11] Rebekah Higgitt. Book review:

Kurt Møller Pedersen and Peter de Clercq (eds.), An Observer of Observatories: The Journal of Thomas Bugge’s Tour of Germany, Holland and England in 1777. Aarhus: Aarhus Universitetsforlag, 2010. Pp. xxiv + 220. ISBN 978-87-7934-311-5. £25.00 (hardback). British Journal for the History of Science, 44(4): 597–598, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Higgitt:2013:BRJ

[Hig13] Rebekah Higgitt. Book reviews: Jeremiah Horrocks and Wilbur Applebaum (trans.), Venus Seen on the Sun: The First Observation of a Transit of Venus by Jeremiah Horrocks. Leiden: Brill, 2012. Pp. xxiv + 82. ISBN 978-90-04-22193-2. EUR99.00 (hardback). British Journal for the History of Science, 46(1):164– 165, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Higgitt:2014:BRL

[Hig14a] Rebekah Higgitt. Book reviews: Louise Miskell, Meeting Places: Scientific Congresses and Urban Identity in Victorian Britain. Farnham: Ashgate, 2013. Pp. xi + 192. ISBN 978-1-4094-5237-9. £60.00 (hardback). British Journal for the History of Science, 47 (4):737–738, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Higgitt:2014:BNO

[Hig14b] Rebekah Higgitt. A British national observatory: the building of the New Physical Observatory at Greenwich, 1889–1898. British Journal for the History of Science, 47(4):609–635, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Higgitt:2015:BRS

[Hig15] Rebekah Higgitt. Book reviews:

Silke Ackermann, Richard L. Kremer and Mara Miniati (eds.), Scientific Instruments on Display. Leiden: Brill, 2014. Pp. xxxiv + 231. ISBN 978-90-04-26439-7. £88.00 (hardback). British Journal for the History of Science, 48 (4):697–699, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Higgitt:2016:BRP

[Hig16]  Rebekah  Hig-

gitt. Book review: Paola Govoni and Zelda Alice Franceschi (eds.), Writing about Lives in Science: (Auto)Biography, Gender, and Genre. Goettingen: V & R Unipress, 2015. Pp. 287. ISBN 978-3-8471-0263-2. EUR44.99 (hardback). British Journal for the History of Science, 49(4):529– 530, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
0D1BD4D05EC1A863FF23460AE2152391.

Higgitt:2019:GNL

[Hig19] Rebekah Higgitt. ‘Greenwich near London’: the Royal Observatory and its London networks in the seventeenth and eighteenth centuries. British Journal for the History of Science, 52(2):297–322, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

                                        journal-for-the-history-of-
                                        science/article/greenwich-
                                        near-london-the-royal-
                                        observatory-and-its-london-
                                        networks-in-the-
                                        seventeenth-and-eighteenth-
                                        centuries/
                                        112F81E1DA2A6EFB6BC458BEEF9030C7.

Hilgartner:2012:SFK

[Hil12] Stephen Hilgartner. Selective flows of knowledge in technoscientific interaction: information control in genome research. British Journal for the History of Science, 45(2):267–280, June 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hirai:2011:MHN

[Hir11] Hiro Hirai. Medical Humanism and Natural Philosophy: Renaissance Debates on Matter, Life, and the Soul, volume 26(17) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-21871-8 (hardcover). xiii + 227 pp. LCCN B778 .H57 2011.

Heffernan:2013:RTD

[HJ13] Michael Heffernan and Heike Jöns. Research travel and disciplinary identities in the University of Cambridge, 1885–1955. British Journal for the History of Science, 46(2):255–286, June 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Horstmanshoff:2012:BST

[HKZ12] H. F. J. Horstmanshoff, Helen King, and Claus Zittel, editors. Blood, sweat, and tears: the changing concepts of physiology from antiquity into early modern Europe, volume 25 of Intersections: interdisciplinary studies in early modern culture. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-22918-3 (hardcover), 90-04-22920-5 (e-book). xxvi + 772 pp. LCCN R135 .B56 2012.

Hogg:2012:IBN

[HL12] Jonathan Hogg and Christoph Laucht. Introduction: British nuclear culture. British Journal for the History of Science, 45(4): 479–493, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hajdenko-Marshall:2015:PMS

[HMLE15] Catherine Hajdenko-Marshall, Bernard V. Lightman, and Richard England, editors. The papers of the Metaphysical Society, 1869–1880: a critical edition. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-964303-2 (set), 0-19-964304-0 (vol. 1), 0-19-964305-9 (vol. 2), 0-19-964306-7 (vol. 3). ???? (3 volumes) pp. LCCN B1614 .P37 2015.

Harrison:2011:WNO

[HNS11] Peter Harrison, Ronald L. Numbers, and Michael H. Shank, editors. Wrestling with nature: from omens to science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-31781-1 (hardcover), 0-226-31783-8 (paperback), 0-226-31803-6 (ebook). x + 416 pp. LCCN Q125 .W86 2011.

Hochadel:2014:BRM

[Hoc14] Oliver Hochadel. Book reviews: Mitchell G. Ash and Jan Surman (eds.), The Nationalization of Scientific Knowledge in the Habsburg Empire, 1848–1918. Basingstoke: Palgrave Macmillan, 2012. Pp. xi + 258. ISBN 978-0-230-28987-1. £50.00 (hardback). British Journal for the History of Science, 47(4):742–743, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hochadel:2016:BRH

[Hoc16a] Oliver Hochadel. Book reviews: Helen Cowie, Exhibiting Animals in Nineteenth-Century Britain: Empathy, Education, Entertainment. Basingstoke: Palgrave Macmillan, 2014. Pp. x + 256. ISBN 978-1-137-38443-0. £60.00 (hardback). Takashi Ito, London Zoo and the Victorians, 1828–1859. Woodbridge: Boydell, 2014. Pp. xi + 204. ISBN 978-0-86193-321-1. £50.00 (hardback). British Journal for the History of Science, 49(2):298– 300, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hochadel:2016:SMM

[Hoc16b] Oliver Hochadel. Spain’s magic mountain: narrating prehistory at Atapuerca. British Journal for the History of Science, 49(3):453–472, September 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
BC591A018EA73259D8E28043896D118E.

Hodacs:2011:LOT

[Hod11] Hanna Hodacs. Linnaeans outdoors: the transformative role of studying nature ‘on the move’ and outside. British Journal for the History of Science, 44(2): 183–209, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hogg:2012:FFT

[Hog12] Jonathan Hogg. ‘the family that feared tomorrow’: British nuclear culture and individual experience in the late 1950s. British Journal for the History of Science, 45 (4):535–549, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Holmes:2009:DBB

[Hol09] John Holmes. Darwin’s bards:

British and American poetry in the age of evolution. Edinburgh University Press, Edinburgh, Scotland, 2009. ISBN 0-7486-3940-3 (hardcover). xiv + 288 pp. LCCN PR585.S33 H65 2009.

Holmes:2010:BRD

[Hol10] John Holmes. Book review:

Denise Gigante, Life: Organic Form and Romanticism. New Haven and London: Yale University Press, 2009. Pp. xiii + 302. ISBN 978-0-300-13685-2. £27.95 (hardback). British Journal for the History of Science, 43(2):303– 305, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hollings:2013:BRA

[Hol13a] Christopher Hollings. Book reviews: Audra J. Wolfe, Competing with the Soviets: Science, Technology, and the State in Cold War America. Baltimore: Johns Hopkins University Press, 2013. Pp viii + 166. ISBN 978-1-4214-0771-5. £10.50 (paperback). British Journal for the History of Science, 46(4):729–731, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Holway:2013:FEA

[Hol13b] Tatiana M. Holway. The flower of empire: an Amazonian water lily, the quest to make it bloom, and the world it created. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-537389-8 (hardcover), 0-19-970604-2 (e-book). xii + 306 + 16 pp. LCCN QK495.N97 H65 2013.

Hollings:2014:BRS

[Hol14a] Christopher Hollings. Book reviews: Scott L. Montgomery, Does Science Need a Global Language? Chicago and London: The University of Chicago Press, 2013. Pp. xiii + 226. ISBN 978-0-226-53503-6. £16.00 (hardback).

British Journal for the History of Science, 47(4):760–762, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hollings:2014:MAI

[Hol14b] Christopher Hollings. Mathematics across the Iron Curtain: a history of the algebraic theory of semigroups, volume 41 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2014. ISBN 1-4704-1493-7. xi + 441 pp. LCCN QA182 .H65 2014.

Hollings:2015:BRV

[Hol15] Christopher Hollings. Book reviews: Victor J. Katz and Karen Hunger Parshall, Taming the Unknown: A History of Algebra from Antiquity to the Early Twentieth Century. Princeton: Princeton University Press, 2014. Pp. xiii + 485. ISBN 978-0-691-14905-9. £34.95 (hardback). British Journal for the History of Science, 48(4):687–689, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hollings:2016:BRR

[Hol16a] Christopher Hollings. Book reviews: Roman Duda, Pearls from a Lost City: The Lvov School of Mathematics. Providence, RI: American Mathematical Society, 2014. Pp. xi + 231. ISBN 978-1-4704-1076-6. £28.50 (hardback). British Journal for the History of Science, 49(1):141– 143, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hollings:2016:BRS

[Hol16b] Christopher D. Hollings. Book review: Slava Gerovitch, Soviet Space Mythologies: Public Images, Private Memories, and the Making of a Cultural Identity. Pittsburgh: University of Pittsburgh Press, 2015. Pp. 256. ISBN 978-0-8229-6363-9. $27.95 (paperback). British Journal for the History of Science, 49(4): 516–517, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        C26905E4D5480FC77D63E2310FD406E5.

Holmes:2016:BRT

[Hol16c] Matthew Holmes. Book reviews: Thor Hansen, The Triumph of Seeds: How Grains, Nuts, Kernels, Pulses, & Pips Conquered the Plant Kingdom and Shaped Human History. New York: Basic Books, 2015. Pp. 288. ISBN 978-0-4650-5599-9. $26.99/£17.99 (hardback). British Journal for the History of Science, 49(1):147– 148, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Homei:2016:BRL

[Hom16] Aya Homei. Book reviews: Lynne Fallwell, Modern German Midwifery, 1885–1960. London: Pickering & Chatto, 2013. Pp. xiii + 263. ISBN 978-1-84893-428-3.

£60.00 (hardback). British Journal for the History of Science, 49(2):308–309, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hoquet:2010:BRJ

[Hoq10] Thierry Hoquet. Book reviews:

Jonathan Hodge and Gregory Radick (eds.), The Cambridge Companion to Darwin, Second Edition. Cambridge: Cambridge University Press, 2009. Pp. xiii + 548. ISBN 978-0-521-71184-5. £19.99 (paperback). Michael Ruse and Robert J. Richards (eds.), The Cambridge Companion to the ‘Origin of Species’. Cambridge: Cambridge University Press, 2009. Pp. xxvii + 395. ISBN 978-0-521-87079-5. £45.00 (hardback). British Journal for the History of Science, 43(1):127– 128, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hodge:2009:CCD

[HR09] M. J. S. (Michael Jonathan Sessions) Hodge and Gregory Radick, editors. The Cambridge companion to Darwin. Cambridge companions to philosophy. Cambridge University Press, Cambridge, UK, second edition, 2009. ISBN 0-521-88475-6, (paperback), 0-521-71184-3 (paperback). xiii + 548 pp. LCCN QH31.D2 C185

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0903/2008038888-b.html;
http://www.loc.gov/catdir/

                                        enhancements/fy0903/
                                        2008038888-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0903/
                                        2008038888-t.html.

Holenstein:2013:SAP

[HSSR13] André Holenstein, Hubert Steinke, Martin Stuber, and Philippe Rogger, editors. Scholars in action: the practice of knowledge and the figure of the savant in the 18th Century, volume 34 / 9 of History of science and medicine library (34); Scientific and learned cultures and their institutions (9). Brill, Leiden, The Netherlands, 2013. ISBN 90-04-24390-9 (hardcover), 90-04-24391-7 (e-book). LCCN AZ604 .S36 2013.

Heim:2009:KWS

[HSW09] Susanne  Heim,  Car-

ola Sachse, and Mark Walker, editors. The Kaiser Wilhelm Society under National Socialism. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-87906-X (hardcover). xxiv + 477 pp. LCCN Q49.K14174 H45

                                        2009. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0914/2008054455-b.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0914/
                                        2008054455-d.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0914/
                                        2008054455-t.htm.

Hudecek:2014:RAC

[Hud14] Jiří Hudeček. Reviving ancient Chinese mathematics: mathematics, history, and politics in the work of Wu Wen-Tsun. Needham Research Institute series. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2014. ISBN 0-415-70296-8 (hardcover), 0-203-79509-1 (ebook). xii + 210 pp. LCCN QA29.W82 H83 2014.

Huff:2011:ICS

[Huf11] Toby E. Huff. Intellectual curiosity and the scientific revolution: a global perspective. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-00082-3 (hardcover), 0-521-17052-4 (paperback). xiii + 354 pp. LCCN Q127.E8 H84

2011. URL http://assets.
cambridge.org/97805211/
70529/cover/9780521170529.
jpg;   http://www.loc.gov/
catdir/enhancements/fy1010/
2010021876-b.html;
http://www.loc.gov/catdir/
enhancements/fy1010/
2010021876-d.html;
http://www.loc.gov/catdir/
enhancements/fy1010/
2010021876-t.html.

Huff:2016:BRH

[Huf16] Toby E. Huff. Book reviews: H.

Floris Cohen, The Rise of Modern Science Explained: A Comparative History. Cambridge: Cambridge University Press, 2015. Pp. 301. ISBN 978-1-107-54560-1. £18.99 (paperback). British Journal for the History of Science, 49(2):289–291, June 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hughes:2010:BRA

[Hug10] Jeff Hughes. Book review: Aitor Anduaga, Wireless and Empire: Geopolitics, Radio Industry and Ionosphere in the British Empire, 1918–1939. Oxford: Oxford University Press, 2009. Pp. xxv + 386. ISBN 978-0-19-956272-5. £39.95 (hardback). British Journal for the History of Science, 43(2):312–314, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hughes:2011:GBB

[Hug11] Sally Smith Hughes. Genentech: the beginnings of biotech. Synthesis. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-35918-2 (hardcover). xv + 213 pp. LCCN HD9999.B444 H85 2011.

Hughes:2012:WBN

[Hug12] Jeff Hughes. What is British nuclear culture? Understanding uranium 235. British Journal for the History of Science, 45(4): 495–518, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Huigen:2009:KCE

[Hui09] Siegfried Huigen. Knowledge and colonialism: Eighteenth-Century travellers in South Africa, volume 18 of Atlantic world. Brill, Leiden, The Netherlands, 2009. ISBN 90-04-17743-4 (hardcover). ISSN 1570-0542. xii + 273 + 39 pp. LCCN DT2020 .H85 2009.

Hull:2010:BRJ

[Hul10] Andrew Hull. Book review: L.

S. Jacyna, Medicine and Modernism: A Biography of Sir Henry Head. London, Pickering and Chatto, 2008. Pp. viii + 353. ISBN 978-1-85196-907-4. £60.00 (hardback). British Journal for the History of Science, 43(1):136– 138, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hunter:2009:BBG

[Hun09] Michael Cyril William Hunter.

Boyle: between God and science. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-12381-7. xii + 366 pp. LCCN Q143.B69 H86 2009.

Hunt:2010:PPL

[Hun10] Bruce J. Hunt. Pursuing power and light: technology and physics from James Watt to Albert Einstein. Johns Hopkins introductory studies in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9358-5 (hardcover), 0-8018-9359-3 (paperback). 182 pp. LCCN T173.8 .H92 2010.

Hunter:2013:WIV

[Hun13] Matthew C. Hunter. Wicked intelligence: visual art and the science of experiment in Restoration London. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-01729-X (hardcover), 0-226-01732-X (e-book). xv + 329 + 10 pp. LCCN Q222 .H86 2013.

Hunt:2018:BRI

[Hun18]  Bruce J. Hunt.  Book re-

view: Iwan Rhys Morus, William Robert Grove: Victorian Gentleman of Science. Cardiff: University of Wales Press, 2017. Pp. 172. ISBN 978-1-78683-004-

3. £16.99 (paperback). British Journal for the History of Science, 51(1):164–165, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/iwan-rhys-
morus-william-robert-grove-
victorian-gentleman-of-
science-cardiff-university-
of-wales-press-2017-pp-172-
isbn-9781786830043-1699-
paperback/
04CF659CB3FB9FC576F42870B51FBB57.

Hunt:2019:BRH

[Hun19] Bruce J. Hunt. Book review: Helge Kragh, Ludvig Lorenz: A Nineteenth-Century Theoretical Physicist. Copenhagen: Det Kongelige Danske Videnskabernes Selskab, 2018. Pp. 280. ISBN 978-8-7730-4417-9. DKK 240,00 (paperback). British Journal for the History of Science, 52(2):370–372, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/helge-kragh-ludvig-
lorenz-a-nineteenthcentury-

theoretical-physicist-
copenhagen-det-kongelige-
danske-videnskabernes-
selskab-2018-pp-280-isbn-
9788773044179-dkk-24000-paperback/
DE959C41B1F48D19217A2D8687FAD75C.

Hussain:2015:BRM

[Hus15] Rawaa Mahmoud Hussain. Book reviews: Michel Janssen and Christoph Lehner (eds.), The Cambridge Companion to Einstein. New York: Cambridge University Press, 2014. Pp. xvi + 562. ISBN 978-0-521-82834-5. £65.00 (hardback). British Journal for the History of Science, 48 (2):374–375, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hutchins:2008:BUO

[Hut08] Roger Hutchins. British university observatories, 1772–1939. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hants, England and Farnham, England, 2008. ISBN 0-7546-3250-4 (hardcover). xxiii + 533 pp. LCCN QB82.G7 H88 2008.

Heering:2011:LDE

[HW11] Peter Heering and Roland Wittje, editors. Learning by doing: experiments and instruments in the history of science teaching. Geschichte (Franz Steiner Verlag). Steiner, Stuttgart, Germany, 2011. ISBN 3-515-09842-9 (paperback). 362 pp. LCCN ????

Hunter:2012:BBS

[HWM12] Michael Hunter, Alison Walker, and Arthur MacGregor. From books to bezoars: Sir Hans Sloane and his collections. British Library, London, UK, 2012. ISBN 0-7123-5880-3 (hardcover). x + 310 pp. LCCN ????

Illy:2012:PEE

[Ill12] József Illy. The practical Einstein: experiments, patents, inventions. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0457-5 (hardcover), 1-4214-0533-4 (e-book). xi + 202 pp. LCCN QC16.E5 I45

2012. URL http://muse.jhu.
edu/books/9781421405339/.

Inkster:2008:HTV

[Ink08] Ian Inkster, editor. History of technology, volume 28. Continuum, London, UK, 2008. ISBN 0-8264-3875-X (e-book), 1-4411-4121-9 (e-book). xiv + 171 pp. LCCN T15 .H57 2008.

Insley:2019:BRD

[Ins19] Jane Insley. Book reviews: David Philip Miller, The Life and legend of James Watt: Collaboration, Natural Philosophy, and the Improvement of the Steam Engine. Pittsburgh: University of Pittsburgh Press, 2019. Pp. ix + 420. ISBN 978-0-8229-4558-1. $50.00 (cloth). - Simon Werrett, Thrifty Science: Making the Most of Materials in the History of Experiment. Chicago and London: The University of Chicago Press, 2019. Pp. 315. ISBN 978-0-2266-1025-2. $45.00 (cloth). British Journal for the History of Science, 52(4):

716–717, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/david-
philip-miller-the-life-and-
legend-of-james-watt-
collaboration-natural-
philosophy-and-the-
improvement-of-the-steam-
engine-pittsburgh-
university-of-pittsburgh-
press-2019-pp-ix-420-isbn-9780822945581-
5000-cloth-simon-werrett-thrifty-
science-making-the-most-of-materials-
in-the-history-of-experiment-
chicago-and-london-the-university-
of-chicago-press-2019-pp-315-
isbn-9780226610252-4500-cloth/
9A6C8F0B0EF4D876D82BFB5ABC2276A4.

Ito:2014:LZV

[Ito14] Takashi Ito. London Zoo and the Victorians, 1828–1859. Studies in history new series. The Royal Historical Society, London, UK, 2014. ISBN 0-86193-321-4. ISSN 0269-2244. xi + 204 pp. LCCN QL76.5.G72 L6648 2014.

Jones:2015:WAI

[JA15] Ross L. Jones and Warwick Anderson. Wandering anatomists and itinerant anthropologists: the antipodean sciences of race in Britain between the wars. British Journal for the History of Science, 48(1):1– 16, March 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jacyna:2008:MMB

[Jac08] L. S. Jacyna. Medicine and modernism: a biography of Sir Henry Head, volume 6 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2008. ISBN 1-85196-907-1 (hardcover). 353 pp. LCCN RC339.52.H43 JAC 2008; WZ112.5.N4 JAC 2008.

Jackson:2009:AB

[Jac09] Mark Jackson. Asthma: the biography. Biographies of disease. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-923795-6 (hardcover). xi + 249 pp. LCCN RC591 .J315 2009.

Jackson:2013:ASS

[Jac13] Mark Jackson. The age of stress: science and the search for stability. Oxford historical monographs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-958862-7 (hardcover). ix + 311 pages pp. LCCN ???? £35.00.

Jacob:2014:FKE

[Jac14] Margaret C. Jacob. The first knowledge economy: human capital and the European economy, 1750–1850. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-61983-1 (paperback), 1-107-04401-4 (hardcover). ix + 257 pp. LCCN HC240 .J335 2014.

Jackson:2015:BRA

[Jac15] Roland Jackson. Book reviews: The Age of Scientific Naturalism:

Tyndall and His Contemporaries. London: Pickering & Chatto, 2014. Pp. xv + 256. ISBN 978-1-84893-463-1. £60.00 (hardback). Bernard Lightman and Michael S. Reidy (eds.). British Journal for the History of Science, 48(1):183– 184, March 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

James:2010:BRG

[Jam10a] Frank A. J. L. James. Book review: Guy Ortolano, The Two Cultures Controversy: Science, Literature and Cultural Politics in Postwar Britain. Cambridge: Cambridge University Press, 2009. Pp. xi + 295. ISBN 978-0-521-89204-9. £55.00 (hardback). British Journal for the History of Science, 43(1):143– 144, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

James:2010:MFV

[Jam10b] Frank A. J. L. James. Michael Faraday: a very short introduction, volume 253 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-957431-6 (paperback). xvi + 162 pp. LCCN QC16.F2 J36 2010.

James:2011:BRA

[Jam11a] Frank A. J. L. James. Book review: Andrew Lambert, Franklin: Tragic Hero of Polar Navigation. London: Faber and Faber, 2009. Pp. xii + 428. ISBN 978-0-571-23160-7. £20.00 (hardback).

British Journal for the History of Science, 44(1):135–137, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

James:2011:BRI

[Jam11b] Frank A. J. L. James. Book review: Ian Hesketh, Of Apes and Ancestors: Evolution, Christianity and the Oxford Debate. Toronto: University of Toronto Press, 2009. Pp. viii + 144. ISBN 978-0-8020-9284-7. CDN$29.95 (hardback). British Journal for the History of Science, 44(2):298–300, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

James:2012:BRM

[Jam12] Frank A. J. L. James. Book reviews: Michael Bentley, The Life and Thought of Herbert Butterfield: History, Science and God. Cambridge: Cambridge University Press, 2011. Pp. xv + 381. ISBN 978-1-107-00397-2. £50.00 (hardback). British Journal for the History of Science, 45(4): 700–701, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

James:2013:BRC

[Jam13] Frank A. J. L. James. Book reviews: The Correspondence of Joseph Black. 2 vols. Farnham: Ashgate, 2012. Pp. xiv + 1564. ISBN 978-0-7546-0131-0. £300.00 (hardback), by R. G. W. Anderson and Jean Jones (eds.). British Journal for the History of Science, 46(3):526–527, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

James:2017:BRM

[Jam17] Frank A. J. L. James. Book review: Melvyn C. Usselman, Pure Intelligence: The Life of William Hyde Wollaston. Chicago: The University of Chicago Press, 2015. Pp. 424. ISBN 978-0-226-24573-7. £24.50 (hardback). British Journal for the History of Science, 50(2):358–359, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
930ABDA9039B8DE6E175CA4FA5D5A44F.

Jamieson:2018:BRR

[Jam18] Annie Jamieson. Book review:

Roland Wittje, The Age of Electroacoustics: Transforming Science and Sound. Cambridge, MA and London: MIT Press, 2016. Pp. 297. ISBN 978-0-262-03526-

2. $40.00/£32.95. British Journal for the History of Science, 51(2):326–328, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/roland-
wittje-the-age-of-
electroacoustics-
transforming-science-and-
sound-cambridge-ma-and-
london-mit-press-2016-pp-
297-isbn-9780262035262-

                                        40003295/
                                        02D170B8B6499FD593EC1B3E477A8559.

Jardine:2010:MJR

[Jar10] Boris Jardine. Book review: Marc J. Ratcliff, The Quest for the Invisible: Microscopy in the Enlightenment. Aldershot: Ashgate, 2009. Pp. xv + 315. ISBN 978-0-7546-6150-4. £65.00 (hardback). British Journal for the History of Science, 43(4):610–611, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jardine:2015:SO

[Jar15] Boris Jardine. The shock of the odd. British Journal for the History of Science, 48(2): 353–356, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jardine:2016:BRE

[Jar16] Boris Jardine. Book reviews: Edward Juler, Grown but Not Made: British Modernist Sculpture and the New Biology. Manchester: Manchester University Press, 2015. Pp. 239. ISBN 978-0-7190-9032-4. £75.00 (hardback). British Journal for the History of Science, 49(2):311– 312, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jardine:2017:BRE

[Jar17] Boris Jardine. Book review: Erik L. Peterson, The Life Organic: The Theoretical Biology Club and the Roots of Epigenetics. Pittsburgh: University of Pittsburgh Press, 2016. Pp. xvi + 334. ISBN 978-0-8229-4466-9. $45.00 (hardback). British Journal for the History of Science, 50 (4):740–741, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
6B8654205B90C34C727BD40485A873E8.

Jay:2009:AHU

[Jay09] Mike Jay. The atmosphere of heaven: the unnatural experiments of Dr. Beddoes and his sons of genius. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-12439-2 (cloth). vii + 294 + 16 pp. LCCN R489.B4 J39 2009.

Jenkins:2015:PHP

[Jen15] Bill Jenkins. Phrenology, heredity and progress in George Combe’s Constitution of Man. British Journal for the History of Science, 48(3): 455–473, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jimenez:2019:BRC

[Jim19] Bárbara Jiménez. Book review:

Charles Morris Lansley, Charles Darwin’s Debt to the Romantics: How Alexander von Humboldt, Goethe and Wordsworth Helped Shape Darwin’s View of Nature. Oxford: Peter Lang, 2018. Pp. 273. ISBN 978-1-78707-138-4. £60.00/$90.95 (hardback). British Journal for the History of Science, 52(1):170–171, 2019.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/charles-morris-
lansley-charles-darwins-
debt-to-the-romantics-how-
alexander-von-humboldt-
goethe-and-wordsworth-
helped-shape-darwins-view-
of-nature-oxford-peter-
lang-2018-pp-273-isbn-9781787071384-
60009095-hardback/F2615BD631A3C323653C2E3A342CD6AA.

Jin:2019:TTO

[Jin19] Xiaoxing Jin. Translation and transmutation: the Origin of Species in China. British Journal for the History of Science, 52(1):117–141, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
translation-and-
transmutation-the-origin-
of-species-in-china/
3F5AED9F9D60EF2D7265927BB2B6AB3A.

Jolivette:2014:BAN

[JJ14] Catherine Jolivette and Carol Jacobi, editors. British art in the nuclear age. British art: histories and interpretations since 1700. Ashgate, Burlington, VT, USA, 2014. ISBN 1-4724-1276-1 (hardcover). xv + 275 pp. LCCN N72.S6 B75 2014.

Janssen:2014:CCE

[JL14] Michel Janssen and Christopher Lehner, editors. The Cambridge Companion to Einstein. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). xvi + 562 pp. LCCN QC16.E5 C36 2014.

Johns:2009:PIP

[Joh09] Adrian Johns. Piracy: the intellectual property wars from Gutenberg to Gates. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-40118-9 (cloth). 626 pp. LCCN K1401 .J64 2009.

Johnston:2011:BRA

[Joh11] Stephen Johnston. Book review: Alexander Marr (ed.), The Worlds of Oronce Fine: Mathematics, Instruments and Print in Renaissance France. Donington: Shaun Tyas, 2009. Pp. xv + 224. ISBN 978-1-900289-96-2. £40.00 (hardback). British Journal for the History of Science, 44 (3):461–462, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johnson:2012:SAN

[Joh12] Paige Johnson. Safeguarding the atom: the nuclear enthusiasm of Muriel Howorth. British Journal for the History of Science, 45 (4):551–571, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

John:2013:BRB

[Joh13a] Steffan John. Book reviews: British Sociology’s Lost Biological Roots: A History of Futures Past. Basingstoke and New York: Palgrave Macmillan, 2012. Pp. xviii + 240. ISBN 978-0-230-35616-0. £55.00 (hardback), by Chris Renwick. British Journal for the History of Science, 46(3):534–535, September 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johnston:2013:BRT

[Joh13b] Sean Johnston. Book reviews: The Truth Machine: A Social History of the Lie Detector. Baltimore: Johns Hopkins University Press, 2012. Pp. ix + 246. ISBN 978-1-4214-0530-8. £18.00 (hardback), by Geoffrey C. Bunn. British Journal for the History of Science, 46(3):540–541, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johnston:2014:BRJ

[Joh14] Sean Johnston. Book reviews: József Illy, The Practical Einstein: Experiments, Patents, Inventions. Baltimore: Johns Hopkins University Press, 2012. Pp. xi + 202. ISBN 978-1-4214-0457-8. £31.00 (hardback). British Journal for the History of Science, 47(2):382–383, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johansen:2015:BRM

[Joh15] Thomas Palmelund Johansen.

Book reviews: Margaret C. Jacob, The First Knowledge Economy: Human Capital and the European Economy, 1750–1850. Cambridge: Cambridge University Press, 2014. Pp. ix + 257. ISBN 978-1-107-61983-8. £19.99 (paperback). British Journal for the History of Science, 48(2): 367–368, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johnston:2016:BRV

[Joh16] Scott Johnston. Book reviews:

Vanessa Ogle, The Global Transformation of Time, 1870–1950. Cambridge, MA: Harvard University Press, 2015. Pp. 288. ISBN 978-0-674-28614-6. $39.95/£29.95 (hardback). British Journal for the History of Science, 49(2):306– 307, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Joly:2011:DC

[Jol11] Bernard Joly. Descartes et la chimie. (French) [Descartes and chemistry]. Mathesis, 1765-8055. Vrin, Paris, France, 2011. ISBN 2-7116-2330-0 (paperback). 256 pp. LCCN QD6 .J65 2011.

Jones:2009:FMS

[Jon09] Claire G. (Claire Gwen) Jones.

Femininity, mathematics, and science, 1880–1914. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2009. ISBN 0-230-55521-7 (hardcover). ix + 264 pp. LCCN QA27.5 .J66 2009.

Jones:2010:BRE

[Jon10a] Alexander Jones. Book review: Eleanor Robson, Mathematics in Ancient Iraq: A Social History. Princeton and Oxford: Princeton University Press, 2008. Pp. xxvii + 441. ISBN 978-0-691-09182-2. £35.00 (hardback). British Journal for the History of Science, 43(2):286–288, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jones:2010:BRJ

[Jon10b] Claire Jones. Book review: Julie K. Brown, Health and Medicine on Display: International Expositions in the United States, 1876– 1904. Cambridge, MA and London: The MIT Press, 2009. Pp. xi + 326. ISBN 978-0-262-02657-4. £33.95 (hardback). British Journal for the History of Science, 43(3):493–494, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jones:2010:DSP

[Jon10c] Susan D. Jones. Death in a small package: a short history of anthrax. Johns Hopkins biographies of disease. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9696-7. xviii + 329 pp. LCCN QR201.A6 J66 2010.

Jones:2013:MTC

[Jon13] Claire Jones. The medical trade catalogue in Britain, 1870–1914, volume 22 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-443-2 (hardcover), 1-78144-067-0 (ebook). xii + 264 pp. LCCN R487 .J66 2013.

Jones:2014:ESB

[Jon14a] Allan Jones. Elite science and the BBC: a 1950s contest of ownership. British Journal for the History of Science, 47(4):701–723, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jones:2014:SRE

[Jon14b] Colin Jones. The smile revolution in eighteenth century Paris. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-871581-1. xiv + 231 pp. LCCN DC729 .J66

2014. URL http://www.loc.
gov/catdir/enhancements/
fy1506/2013953963-b.html;
http://www.loc.gov/catdir/
enhancements/fy1506/
2013953963-d.html;
http://www.loc.gov/catdir/
enhancements/fy1506/
2013953963-t.html.

Jones:2016:JGC

[Jon16a] Allan Jones. J. G. Crowther’s war: Institutional strife at the BBC and British Council. British Journal for the History of Science, 49(2):259–278, June 2016.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jones:2016:BMT

[Jon16b] Claire L. Jones. A barrier to medical treatment? British medical practitioners, medical appliances and the patent controversy, 1870–1920. British Journal for the History of Science, 49(4): 601–625, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
C9C599FCC36D2554523AD0754F9F8D0D.

Jardine:2008:GAQ

[JS08]  Nicholas  Jardine

and A. Ph Segonds. La guerre des astronomes: la querelle au sujet de l’origine du système géo-héliocentrique à la fin du XVIe siècle. (French) [The astronomer’s war: the quarrel about the origin of the geo-heliocentric system at the end of the Sixteenth Century], volume 9.–10 of Science et humanisme. Belles lettres, Paris, France, 2008. ISBN 2-251-34513-2 (v. 1), 2-251-34512-4 (v. 2/1-2). v. 1–2 in 3 pp. LCCN QB603.M6 J37 2008.

Josephson:2017:PMP

[JS17] Paul Josephson and Aleksandr Sorokin. Physics moves to the provinces: the Siberian physics community and Soviet power, 1917–1940. British Journal for the History of Science, 50(2): 297–327, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
C32BF69CEB9091E1E1D6AE6950971709.

Kaartinen:2013:BCE

[Kaa13] Marjo Kaartinen. Breast cancer in the eighteenth century. Studies for the International Society for Cultural History. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-364-9 (hardcover), 1-84893-365-7 (e-book). xii + 181 pp. LCCN RC280.B8 K32 2013.

Kaalund:2018:BRH

[Kaa18] Nanna Katrine Lüders Kaalund.

Book review: Heather Ellis, Masculinity and Science in Britain, 1831–1918. London and New York: Palgrave Macmillan, 2017. Pp. 240. ISBN 978-1-137-31173-3. £66.99 (hardcover). British Journal for the History of Science, 51 (4):710–711, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/heather-
ellis-masculinity-and-
science-in-britain-
18311918-london-and-new-
york-palgrave-macmillan-
2017-pp-240-isbn-
9781137311733-6699-
hardcover/
645E4BA5C5DB529353990F4903405A78.

Kaoukji:2012:BRV

[Kao12] N. Kaoukji. Book review: Volker Remmert, Picturing the Scientific Revolution. Philadelphia: St Joseph’s University Press, 2011. Pp. 295. ISBN 978-0-916101-67-1. $65.00 (hardback). British Journal for the History of Science, 45(2):288–290, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kassler:2014:STR

[Kas14]  Jamie Croy Kassler.  Seek-

ing truth: Roger North’s notes on Newton and correspondence with Samuel Clarke c. 1704–1713. Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4094-4921-1 (hardcover), 1-4094-4922-X (e-book), 1-4724-0646-X (epub). xii + 374 pp. LCCN DA437.N87 K38 2014.

Kumar:2009:WSI

[KB09] Neelam Kumar and Antoinette M. Burton, editors. Women and science in India: a reader. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-569705-7. xxx + 351 pp. LCCN Q130 .W6525

                                        2009. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0907/2009341053-b.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0907/
                                        2009341053-d.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0907/
                                        2009341053-t.htm.

Krige:2013:NWF

[KCM13] John Krige, Angelina Long Callahan, and Ashok Maharaj. NASA in the world: fifty years of international collaboration in space.

Palgrave studies in the history of science and technology. Palgrave Macmillan, New York, NY, 2013. ISBN 1-137-34091-6 (hardcover), 1-137-34092-4 (paperback). xvii + 353 pp. LCCN TL521.312 .K75 2013.

Keene:2012:BRL

[Kee12] Melanie Keene. Book review:

Laurence Talairach-Vielmas (ed.), Science in the Nursery: The Popularisation of Science in Britain and France, 1761–1901. Newcastle: Cambridge Scholars Publishing, 2011. Pp. xii + 309. ISBN 978-1-4438-2680-8. £44.99 (hardback). British Journal for the History of Science, 45(2):302– 303, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Keene:2017:BRD

[Kee17] Melanie Keene. Book review:

Donald L. Opitz, Staffan Bergwik and Brigitte van Tiggelen (eds.), Domesticity in the Making of Modern Science. Basingstoke: Palgrave Macmillan, 2016. Pp. 320. ISBN 978-1-137-49272-2. £63.00 (hardback). British Journal for the History of Science, 50(3): 548–549, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E3D95982BF46C540709F74D3FC7D7285.

Kehoe:2012:BRE

[Keh12] S. Karly Kehoe. Book review:

Emily Steel, He Is No Loss:

Robert McCormick and the Voyage of HMS Beagle. BSHS Monograph 14. Norwich: British Society for the History of Science, 2011. Pp. x + 63. ISBN 978-0-906450-18-5. £10.00 (paperback). British Journal for the History of Science, 45(2):301– 302, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Keller:2010:MSB

[Kel10] Evelyn Fox Keller. The mirage of a space between nature and nurture. Duke University Press, Durham, NC, USA, 2010. ISBN 0-8223-4714-8 (hardcover), 0-8223-4731-8 (paperback). 107 pp. LCCN Wel PAJ Kel.

Keller:2012:MTS

[Kel12] Vera Keller. Mining Tacitus: secrets of empire, nature and art in the reason of state. British Journal for the History of Science, 45(2):189–212, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kelsey:2017:BRH

[Kel17]  Tom Kelsey.  Book

review: Hermione Giffard, Making Jet Engines in World War II: Britain, Germany, and the United States. Chicago: The University of Chicago Press, 2016. Pp. 349. ISBN 978-0-226-38859-5. $45.00 (cloth). British Journal for the History of Science, 50(3): 558–560, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
57017C64C585AE12BCDF76EC678FECDD.

Kempley:2019:JTW

[Kem19] Eilís Kempley. Julian Trevelyan, Walter Maclay and Eric Guttmann: drawing the boundary between psychiatry and art at the Maudsley Hospital. British Journal for the History of Science, 52(4): 617–643, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/julian-
trevelyan-walter-maclay-
and-eric-guttmann-drawing-
the-boundary-between-
psychiatry-and-art-at-the-
maudsley-hospital/
F782CF7756D221D3D102D8954654F7E7.

Kennaway:2012:BVH

[Ken12] James Gordon Kennaway. Bad vibrations: the history of the idea of music as cause of disease. History of medicine in context. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-2642-4 (hardcover), 1-4094-2643-2 (e-book). xii + 213 pp. LCCN ML3830 .K43 2012.

Kennaway:2013:BRH

[Ken13] James Kennaway. Book reviews:

Heather R. Beatty, Nervous Disease in Late Eighteenth-Century Britain: The Reality of a Fashionable Disorder. London: Pickering and Chatto, 2012. Pp. vii + 241. ISBN 978-1-84893-308-8. £60.00 (hardback). British Journal for the History of Science, 46 (4):719–720, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kennaway:2019:BRL

[Ken19] James Kennaway. Book review: Lynda Payne, The Best Surgeon in England: Percivall Pott, 1713–88. New York: Peter Lang Publishing, 2017. Pp. 236. ISBN 978-1-4331-2319-1. £64.00 (hardcover). British Journal for the History of Science, 52(4): 714–716, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/lynda-
                                        payne-the-best-surgeon-in-
                                        england-percivall-pott-
                                        171388-new-york-peter-lang-
                                        publishing-2017-pp-236-
                                        isbn-9781433123191-6400-
                                        hardcover/
                                        24F50DE156C6E3603A359F4598CCE11F.

Kershaw:2014:TSE

[Ker14] Michael Kershaw. ‘A thorn in the side of European geodesy’: measuring Paris–Greenwich longitude by electric telegraph. British Journal for the History of Science, 47 (4):637–660, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kichigina:2009:ILE

[Kic09] Galina Kichigina. The imperial laboratory: experimental physiology and clinical medicine in post-Crimean Russia, volume 87 of Clio medica: The Wellcome series in the history of medicine. Rodopi, Amsterdam, The Netherlands, 2009. ISBN 90-420-2658-8. 374 pp. LCCN W1 CL871 2009.

Kidd:2011:BRF

[Kid11a] Ian James Kidd. Book review: Feyerabend Paul, Against Method. London and New York: Verso, 2010. Pp. xxxii + 296. ISBN 1-84467-442-8. £10.24 (paperback). British Journal for the History of Science, 44(2): 311–312, June 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
41DCCA046A134C5DCD445269D5CB8BEC.

Kidd:2011:BRP

[Kid11b] Ian James Kidd. Book review:

Paul Feyerabend, The Tyranny of Science, ed. Eric Oberheim. London: Polity, 2011. Pp. xii + 153. ISBN 978-0-7456-5190-3. £12.95 (paperback). British Journal for the History of Science, 44(4): 576–577, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

King:2013:OSB

[Kin13] Helen King. The one-sex body on trial: the classical and early modern evidence. The history of medicine in context. Ashgate, Farnham, Surrey, 2013. ISBN 1-4094-6335-4 (hardcover). xii + 274 pp. LCCN QP278 .K56 2013.

Kiprijanov:2016:BRW

[Kip16] Konstantin S. Kiprijanov. Book review: William H. Brock, The History of Chemistry: A Very Short Introduction. Oxford: Oxford University Press, 2016. Pp. xvi + 151. ISBN 978-0-19-871648-

8. £7.99 (paperback). British Journal for the History of Science, 49(4):637–638, December 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        68D2637CEDF871D1D2058F9EDF75DD58.

Kirby:2011:LCH

[Kir11] David A. (David Anthony) Kirby. Lab coats in Hollywood: science, scientists, and cinema. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-01478-5 (hardcover). xiv + 265 pp. LCCN PN1995.9.S265 K57 2011. UK£20.95.

Kirchhelle:2013:BRR

[Kir13]  Claas Kirchhelle.  Book re-

views: Risk: Negotiating Safety in American Society. Baltimore: Johns Hopkins University Press, 2013. Pp. viii + 329. ISBN 978-1-4214-0790-6. £28.50 (hardback), by Arwen P. Mohun. British Journal for the History of Science, 46(3):538–540, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kirwan:2015:BRC

[Kir15] Adrian James Kirwan. Book reviews: Charles Mollan (ed.), William Parsons, 3rd Earl of Rosse: Astronomy and the Castle in Nineteenth-Century Ireland. Manchester: Manchester University Press, 2014. Pp. xxii + 368. ISBN 978-0-7190-9144-5. £70.00 (hardback). British Journal for the History of Science, 48 (4):699–700, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kirby:2017:RCS

[Kir17] David A. Kirby. Regulating cinematic stories about reproduction: pregnancy, childbirth, abortion and movie censorship in the US, 1930–1958. British Journal for the History of Science, 50(3): 451–472, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
2A1AA5620E20D7E46CE14837999896D2.

Katrine:2014:BRD

[KK14]  Nanna  Ka-

trine and Lüders Kaalund. Book reviews: Dan Ch. Christensen, Hans Christian Ørsted: Reading Nature’s Mind. Oxford: Oxford University Press, 2013. Pp. xix + 743. ISBN 978-0-19-966926-4. £39.99 (hardback). British Journal for the History of Science, 47(2):379–380, June 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kragh:2008:SDT

[KKNN08] Helge S. Kragh, Peter C. Kjaergaard, Henry Nielsen, and Kristian Hvidfelt Nielsen, editors. Science in Denmark: a thousandyear history. Aarhus University Press, Århus, Denmark, 2008. ISBN 87-7934-317-1. 607 pp. LCCN Q127.D4 S344 2008. DKR 498,00.

Klaver:2009:SEA

[Kla09] J. M. I. Klaver. Scientific expeditions to the Arab world (1761– 1881), volume 4 of Studies in the Arcadian Library. Arcadian Library in association with Oxford University Press, London, UK, 2009. ISBN 0-19-956889-8 (hardcover). 255 pp. LCCN Q115 .K54 2009.

Klemm:2013:BWK

[Kle13] Tanja Klemm. Bildphysiologie: Wahrnehmung und Körper in Mittelalter und Renaissance. (German) [Image Physiology: perception and body in the Middle Ages and the Renaissance]. Akademie-Verlag, Berlin, Germany, 2013. ISBN 3-05-006478-1 (hardcover). xii + 324 pp. LCCN B828.45 .K64 2013.

Kroupa:2018:SII

[KMB18] Sebestian Kroupa, Stephanie J. Mawson, and Dorit Brixius. Science and islands in Indo-Pacific worlds. British Journal for the History of Science, 51(4): 541–558, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/science-
and-islands-in-indopacific-
worlds/
1C6F5911677AA44563AF25DF1D5EF1DF.

Knell:2010:PCF

[Kne10] Simon Knell. Book review: Pietro Corsi, Fossils and Reputations. A Scientific Correspondence: Pisa, Paris, London, 1853–1857. Pisa: Edizioni Plus — Pisa University Press, 2008. Pp. 411. ISBN 978-88-8492-564-0. EUR25.00 (paperback). British Journal for the History of Science, 43(4):616– 617, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Knight:2009:MMS

[Kni09] David M. Knight. The making of modern science: science, technology, medicine and modernity: 1789–1914. Polity, Cambridge, UK, 2009. ISBN 0-7456-3675-6 (hardcover), 0-7456-3676-4 (paperback). xii + 370 pp. LCCN Q125 .K566 2009.

Knight:2015:BRJ

[Kni15] David Knight. Book reviews:

James A. Secord, Visions of Science: Books and Readers at the Dawn of the Victorian Age. Oxford: Oxford University Press, 2014. Pp. xiii + 306. ISBN 978-0-19-967526-5. £18.99 (hardback). British Journal for the History of Science, 48(2):368–370, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Knight:2016:BRS

[Kni16] David Knight. Book reviews: Susannah Gibson, Animal, Vegetable, Mineral? How Eighteenth-Century Science Disrupted the Natural Order. Oxford: Oxford University Press, 2015. Pp. 240. ISBN 978-0-19-870513-0. £16.99 (hardback). British Journal for the History of Science, 49(2):293– 294, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Knoeff:2014:BRC

[Kno14]  Rina Knoeff.  Book review:

Catherine Packham, Eighteenth-Century Vitalism: Bodies, Culture, Politics. Basingstoke: Palgrave Macmillan, 2012. Pp. viii + 250. ISBN 978-0-230-27618-5. £50.00 (hardback). British Journal for the History of Science, 47 (1):182–183, March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Koberer:2016:BRR

[Köb16] Wolfgang Köberer. Book review: Richard Dunn and Rebekah Higgitt (eds.), Navigational Enterprises in Europe and Its Empires, 1730–1850. Basingstoke: Palgrave Macmillan, 2016. Pp. 272. ISBN 978-1-137-52063-0. £63.00 (hardback). British Journal for the History of Science, 49(3): 492–494, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
2B6F22DD38DDB74CCF7107AC3EB82FA1.

Kohn:2008:DGE

[Koh08] David Kohn. Darwin’s garden:

an evolutionary adventure. New York Botanical Garden, Bronx, NY, USA, 2008. ISBN 0-89327-970-6. 60 pp. LCCN ????

Kohlstedt:2010:TCS

[Koh10] Sally Gregory Kohlstedt. Teaching children science: hands-on nature study in North America, 1890–1930. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-44990-4 (cloth). xv + 363 pp. LCCN LB1585.3 .K64 2010.

Kotevska:2016:BRS

[Kot16] Laura Kotevska. Book review:

Snezana Lawrence and Mark Mc-Cartney (eds.), Mathematics and Their Gods: Interactions between Mathematics and Religious Beliefs. Oxford: Oxford University Press, 2015. Pp. 304. ISBN 978-0-19-870305-1. £24.99 (hardback). British Journal for the History of Science, 49(4): 526–527, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
46B58C2F9FFF7BB06684F9D98B8FA5B6.

Kutzinski:2012:AHA

[KOW12] Vera M. Kutzinski, Ette Ottmar, and Laura Dassow Walls, editors. Alexander von Humboldt and the Americas, volume 3 of Potsdamer inter- und transkulturelle Texte (Pointe). Verlag Walter Frey, Berlin, Germany, 2012. ISBN 3-938944-63-3. 295 pp. LCCN Q143.H9 A524 2012.

Kirk:2013:L

[KP13] Robert G. W. Kirk and Neil Pemberton. Leech. Animal. Reaktion Books, London, UK, 2013. ISBN 1-78023-033-8. 208 pp. LCCN ????

Katz:2014:TUH

[KP14] Victor J. Katz and Karen Hunger Parshall. Taming the unknown: a history of algebra from antiquity to the early twentieth century. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-14905-4 (hardcover). xiii + 485 pp. LCCN QA151 .K38 2014.

Krige:2012:HKT

[Kri12] John Krige. Hybrid knowledge: the transnational co-production of the gas centrifuge for uranium enrichment in the 1960s. British Journal for the History of Science, 45(3):337–357, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ksiazkiewicz:2013:BRJ

[Ksi13] Allison Ksiazkiewicz. Book reviews: Jed Z. Buchwald and Mordechai Feingold, Newton and the Origin of Civilization. Princeton and Oxford: Princeton University Press, 2013. Pp. xii + 528. ISBN 978-0-691-15478-7. £34.95 (hardback). British Journal for the History of Science, 46(2):345– 347, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ksiazkiewicz:2015:BRG

[Ksi15] Allison Ksiazkiewicz. Book reviews: Geographies of the Romantic North: Science, Antiquarianism, and Travel, 1790–1830. Basingstoke: Palgrave Macmillan, 2013. Pp. ixv + 265. ISBN 978-1-137-31131-3. £55.00 (hardback). Angela Byrne. British Journal for the History of Science, 48(1):180–181, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ksiazkiewicz:2016:BRM

[Ksi16] Allison Ksiazkiewicz. Book reviews: Martin J. S. Rudwick, Earth’s Deep History: How It Was Discovered and Why It Matters. Chicago and London: The University of Chicago Press, 2014. Pp. ix + 360. ISBN 978-0-226-20393-5. $30.00/£21.00 (hardback). British Journal for the History of Science, 49(1):116– 117, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kilburn-Toppin:2019:PGT

[KT19]  Jasmine Kilburn-Toppin.  ‘A

place of great trust to be supplied by men of skill and integrity’: assayers and knowledge cultures in late sixteenth- and seventeenthcentury London. British Journal for the History of Science, 52(2):197–223, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/place-of-
great-trust-to-be-supplied-
by-men-of-skill-and-
integrity-assayers-and-
knowledge-cultures-in-late-
sixteenth-and-
seventeenthcentury-london/
602964BC32CA4CC1069A5A6DE1BD5639.

Kumar:2016:BRD

[Kum16] Deepak Kumar. Book review: David Arnold, Toxic Histories: Poison and Pollution in Modern India. Cambridge: Cambridge University Press, 2016. Pp. viii + 241. ISBN 978-1-107-12697-8. £34.99 (hardback). British Journal for the History of Science, 49 (4):646–648, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
378DD639FDC53998F00EF63928886981.

Kumar:2018:BRP

[Kum18] Deepak Kumar. Book review: Pankaj Jain, Science and Socioreligious Revolution in India: Moving the Mountains. Abingdon and New York: Routledge, 2017. Pp. xxii + 98. ISBN 978-1-138-02359-8. £48.99 (hardcover). British Journal for the History of Science, 51(2):305–307, June 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/

journals/british-journal-
for-the-history-of-science/
article/pankaj-jain-
science-and-socioreligious-
revolution-in-india-moving-
the-mountains-abingdon-and-
new-york-routledge-2017-pp-
xxii-98-isbn-9781138023598-
4899-hardcover/
3B84EB491F11E67DEE7465590A83328F.

Kusukawa:2011:PBN

[Kus11]  Sachiko Kusukawa.  Pictur-

ing the book of nature: image, text, and argument in Sixteenth-Century human anatomy and medical botany. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-46529-2 (hardcover). xvii + 331 pp. LCCN QH46.5 .K87 2011.

Kusukawa:2015:BRM

[Kus15] Sachiko Kusukawa. Book reviews: Matthew C. Hunter, Wicked Intelligence: Visual Art and the Science of Experiment in Restoration London. Chicago and London: The University of Chicago Press, 2013. ISBN 978-0-226-01729-7. £38.50 (hardback). British Journal for the History of Science, 48(2):361–362, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kusukawa:2018:BRH

[Kus18] Sachiko Kusukawa. Book review: Henrietta McBurney, Ian Rolfe, Caterina Napoleone and Paula Findlen, with Carlo Violani, Onno Wijnands, Arturo Morales-Muñiz, Eufrasia Roselló-Izquierdo, Arthur McGregor and Kathie Way, Birds, Other Animals and Natural Curiosities (Volumes One and Two). London: Royal Collection Trust (in association with Harvey Miller Publishers), 2017. Pp. 927. ISBN 978-1-9094-0060-3. EUR170.00 (hardcover). British Journal for the History of Science, 51(3):518– 520, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/henrietta-
mcburney-ian-rolfe-
caterina-napoleone-and-
paula-findlen-with-carlo-
violani-onno-wijnands-
arturo-moralesmuniz-
eufrasia-roselloizquierdo-
arthur-mcgregor-and-kathie-
way-birds-other-animals-and-
natural-curiosities-volumes-
one-and-two-london-royal-collection-
trust-in-association-with-harvey-
miller-publishers-2017-pp-927-
isbn-9781909400603-17000-hardcover/
DD926FA182A89EF80575B42ED7022B36.

Kendall:2010:BNT

[KW10] Calvin B. Kendall and Faith Wallis. Bede: On the nature of things and On times, volume 56 of Translated texts for historians. Liverpool University Press, Liverpool, UK, 2010. ISBN 1-84631-495-X (hardcover), 1-84631-496-8 (paperback). xiv + 222 pp. LCCN ????

Keighren:2015:TPE

[KWB15] Innes M. Keighren, Charles W. J.

Withers, and Bill Bell. Travels into print: exploration, writing, and publishing with John Murray, 1773–1859. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-42953-9 (hardcover), 0-226-23357-X (e-book). xiii + 364 + 16 pp. LCCN Z286.T8 K45 2015.

Knoeff:2015:FAC

[KZ15] Rina Knoeff and Robert Zwijnenberg, editors. The fate of anatomical collections. The history of medicine in context. Ashgate, Farnham, Surrey, 2015. ISBN 1-4094-6815-1 (hardcover), 1-4094-6816-X (e-book), 1-4094-6817-8 (epub). xx + 305 pp. LCCN QM21 .F35 2015.

Levin:2010:UMC

[L^{+}10] Miriam R. Levin et al., editors.

Urban modernity: cultural innovation in the Second Industrial Revolution. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01398-3 (hardcover). x + 272 pp. LCCN HT361 .U7173 2010.

Lack:2009:AHB

[Lac09] H. Walter Lack. Alexander von Humboldt and the botanical exploration of the Americas. Prestel, Munich, Germany, 2009. ISBN 3-7913-4142-1 (hardcover). 278 pp. LCCN QH11.L29 L33 2009. Translated from German by Stephen Telfer.

LaFollette:2013:SAT

[LaF13] Marcel C. (Marcel Chotkowski) LaFollette. Science on American television: a history. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92199-9 (hardcover), 0-226-92201-4 (ebook). x + 306 pp. LCCN PN1992.8.S33 L34 2013.

Lambert:2009:FTH

[Lam09] Andrew D. Lambert. Franklin: tragic hero of polar navigation. Faber and Faber, London, UK, 2009. ISBN 0-571-23160-8 (hardcover). xii + 428 + 8 pp. LCCN G660 .L26 2009.

Lambert:2011:UAJ

[Lam11] Kevin Lambert. The uses of analogy: James Clerk Maxwell’s ‘On Faraday’s lines of force’ and early Victorian analogical argument. British Journal for the History of Science, 44(1):61–88, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lamont:2013:EBH

[Lam13] Peter Lamont. Extraordinary beliefs: a historical approach to a psychological problem. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-01933-8 (hardcover), 1-107-68802-7 (paperback). xi + 321 pp. LCCN BF1031 .L255 2013.

Lancaster:2011:BRS

[Lan11a] James A. T. Lancaster. Book review: Stephen Gaukroger, The Collapse of Mechanism and the Rise of Sensibility: Science and the Shaping of Modernity, 1680– 1760. Oxford: Oxford University Press, 2010. Pp. x + 505. ISBN 978-0-19-959493-1. £35.00 (hardback). British Journal for the History of Science, 44(4): 593–595, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lane:2011:GMSa

[Lan11b] K. Maria D. Lane. Geographies of Mars: seeing and knowing the Red Planet. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-47078-4 (hardcover). xiii + 266 pp. LCCN QB641 .L366 2011.

Lancaster:2012:BRP

[Lan12] James A. T. Lancaster. Book review: Peter R. Anstey, John Locke and Natural Philosophy. Oxford: Oxford University Press, 2011. Pp. xii + 252. ISBN 978-0-19-958977-7. £35.00 (hardback). British Journal for the History of Science, 45(1):129–130, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Ans11].

Langrish:2014:BRM

[Lan14] John Z. Langrish. Book reviews:

Matthew Wisnioski, Engineers for Change: Competing Visions of Technology in 1960s America. Cambridge, MA and London: MIT Press, 2012. Pp. xvii + 286. ISBN 978-0-262-01826-5. £24.95 (hardback). British Journal for the History of Science, 47(2):387– 388, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lancaster:2015:BRC

[Lan15a] Cheryl Lancaster. Book reviews: Carsten Timmermann, A History of Lung Cancer: The Recalcitrant Disease. Basingstoke: Palgrave Macmillan, 2014. Pp. 256. ISBN 978-1-4039-8802-7. £53.00 (hardback). British Journal for the History of Science, 48(3): 534–536, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Langrish:2015:BRR

[Lan15b] John Z. Langrish. Book reviews: Robert J. Richards, Was Hitler a Darwinian? Disputed Questions in the History of Evolutionary Theory. Chicago and London: The University of Chicago Press, 2013. Pp. 269. ISBN 978-0-226-05893-1. $27.50 (paperback). British Journal for the History of Science, 48(4):704–705, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lane:2019:BRL

[Lan19] Sydney Lane. Book review: Lisa M. Osbeck, Values in Psychological Science: Re-imagining Epistemic Priorities at a New Frontier. Cambridge: Cambridge University Press, 2018. Pp. ix + 145. ISBN 978-1-1071-3490-4. £80.00 (hardback). British Journal for the History of Science, 52(3): 537–539, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/lisa-m-
osbeck-values-in-
psychological-science-
reimagining-epistemic-
priorities-at-a-new-
frontier-cambridge-
cambridge-university-press-
2018-pp-ix-145-isbn-
9781107134904-8000-hardback/
6AB7218D4BF70B9C6E27EA28D616616B.

Larsen:2013:BTR

[Lar13] Timothy Larsen. E. B. Tylor, religion and anthropology. British Journal for the History of Science, 46(3):467–485, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Larsen:2014:SGA

[Lar14] Timothy Larsen. The slain God:

anthropologists and the Christian faith. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-965787-4, 0-19-178557-1 (e-book). 256 pp. LCCN BR115.C8 L37

2014. URL http://www.loc.
gov/catdir/enhancements/
fy1503/2014934932-b.html;
http://www.loc.gov/catdir/
enhancements/fy1503/
2014934932-d.html;
http://www.loc.gov/catdir/
enhancements/fy1503/
2014934932-t.html.

Laucht:2012:APA

[Lau12] Christoph Laucht. Atoms for the people: the Atomic Scientists’ Association, the British state and nuclear education in the Atom Train exhibition, 1947– 1948. British Journal for the History of Science, 45(4):591– 608, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Laucht:2015:BRK

[Lau15] Christoph Laucht. Book reviews: Kenneth D. McRae, Nuclear Dawn: F.E. Simon and the Race for Atomic Weapons in World War II. Oxford: Oxford University Press, 2014. Pp. xxviii + 284. ISBN 978-0-19-968718-3. £35.00 (hardback). British Journal for the History of Science, 48(3):532–533, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lawrence:2015:ADE

[Law15a] Natalie Lawrence. Assembling the dodo in early modern natural history. British Journal for the History of Science, 48(3): 387–408, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lawson:2015:BEG

[Law15b] Ian Lawson. Bears in Eden, or, this is not the garden you’re looking for: Margaret Cavendish, Robert Hooke and the limits of natural philosophy. British Journal for the History of Science, 48 (4):583–605, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lawrence:2017:BRC

[Law17] Farrah Mary Lawrence. Book review: Christine Holmberg, Stuart Blume and Paul Greenough (eds.), The Politics of Vaccination: A Global History. Manchester: Manchester University Press, 2017. Pp. xiii + 343. ISBN 978-1-5261-1088-6. £75.00 (hardcover). British Journal for the History of Science, 50(4): 741–743, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A38642900985F719A177A7AD757E4B56.

Lynch:2008:TMC

[LCMJ08] Michael Lynch, Simon A. Cole, Ruth McNally, and Kathleen Jordan. Truth machine: the contentious history of DNA fingerprinting. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-49806-9 (cloth). xxii + 391 pp. LCCN RA1057.55 .L96 2008. URL http://www.

loc.gov/catdir/toc/ecip088/
2008001902.html.

Lidwell-Durnin:2018:BRJ

[LD18] John Lidwell-Durnin. Book review: Janet Browne (ed.), The Quotable Darwin. Princeton, NJ and Oxford: Princeton University Press, 2018. Pp. xxix + 348. ISBN 978-0-691-16935-4. $24.95 (hardcover). British Journal for the History of Science, 51(4):

711–713, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/janet-
browne-ed-the-quotable-
darwin-princeton-nj-and-
oxford-princeton-
university-press-2018-pp-
xxix-348-isbn-
9780691169354-2495-
hardcover/
C1CC39EBC72E7D949229301A6B61E3A2.

Lean:2016:BRC

[Lea16] Thomas Lean. Book review: Casey O’Donnell, Developer’s Dilemma: The Secret World of Videogame Creators. Cambridge, MA: The MIT Press, 2014. Pp. 320. ISBN 978-0-262-02819-6. £22.95 (hardback). British Journal for the History of Science, 49 (4):523–524, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
7C61B464C6005779CCF8B74737C630F3.

Lee:2015:BUR

[Lee15a] Jung Lee. Between universalism and regionalism: universal systematics from imperial Japan. British Journal for the History of Science, 48(4):661–684, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lee:2015:BRJ

[Lee15b] Victoria Lee. Book reviews: John P. DiMoia, Reconstructing Bodies: Biomedicine, Health, and Nation-Building in South Korea since 1945. Stanford, CA: Stanford University Press, 2013. Pp. xii + 280. ISBN 978-0-8047-8411-5. $45.00 (hardback). British Journal for the History of Science, 48(4):716–717, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Leggett:2013:RRP

[Leg13] Don Leggett. Replication, replacing and naval science in comparative context, c. 1868–1904. British Journal for the History of Science, 46(1):1–21, March 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lehoux:2007:AWC

[Leh07]  Daryn Lehoux.  Astronomy,

weather, and calendars in the ancient world: parapegmata and related texts in classical and Near Eastern societies. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-85181-5 (hardcover). xiv + 566 pp. LCCN QC999 .L44

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0703/2006038547-d.html;
http://www.loc.gov/catdir/
enhancements/fy0703/
2006038547-t.html;
http://www.loc.gov/catdir/
enhancements/fy0729/
2006038547-b.html.

Lehoux:2012:BRA

[Leh12a] Daryn Lehoux. Book review: Aaron Poochigian (trans.), Aratus: Phaenomena. Baltimore: Johns Hopkins University Press, 2010. Pp. xxxiv + 72. ISBN 978-0-8018-9466-4. $50.00 (hardback), $25.00 (paperback). British Journal for the History of Science, 45 (1):123–125, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [APoC10].

Lehoux:2012:WDR

[Leh12b] Daryn Lehoux. What did the Romans know?: an inquiry into science and worldmaking. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-47114-4 (hardcover). xii + 275 pp. LCCN Q124.95 .L44 2012.

Lehoux:2019:WDA

[Leh19] Daryn Lehoux. Why does Aristotle think bees are divine? Proportion, triplicity and order in the natural world. British Journal for the History of Science, 52(3): 383–403, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/why-does-
                                        aristotle-think-bees-are-
                                        divine-proportion-
                                        triplicity-and-order-in-
                                        the-natural-world/
                                        64B9CB767F0DA3646F58E7DB630777D1.

Lester:2013:BRCb

[Les13a] Ahren Lester. Book reviews:

Charles H. Smith and Megan Derr (eds.), Alfred Russel Wallace’s 1886–1887 Travel Diary: The North American Lecture Tour. Manchester: SIRI Scientific Press, 2013. Pp. xiii + 258. ISBN 978-0-9567795-8-8. £21.00 (paperback). British Journal for the History of Science, 46(4): 722–724, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lester:2013:BRCa

[Les13b] Ahren Lester. Book reviews:

Christine Ferguson, Determined Spirits: Eugenics, Heredity and Racial Regeneration in Anglo-American Spiritualist Writing, 1848–1930. Edinburgh: Edinburgh University Press, 2012. Pp. x + 230. ISBN 978-0-7486-3965-6. £70.00 (hardback). British Journal for the History of Science, 46(2):351–352, June 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lester:2014:BRT

[Les14] Ahren Lester. Book reviews: Ted Benton, Alfred Russel Wallace: Explorer, Evolutionist, Public Intellectual — a Thinker for Our Own Times? Manchester: Siri Scientific Press, 2013. Pp. 223. ISBN 978-0-9574530-2-9. £21.00 (paperback). British Journal for the History of Science, 47(4): 740–742, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lester:2015:BRJ

[Les15] Ahren Lester. Book reviews: James T. Costa (ed.), On the Organic Law of Change: A Facsimile Edition and Annotated Transcription of Alfred Russel Wallace’s Species Notebook of 1855– 1859. Cambridge, MA: Harvard University Press, 2013. Pp. xii + 573. ISBN 978-0-674-72488-4. £36.95 (hardback). British Journal for the History of Science, 48(3):521–523, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Levitt:2009:SEO

[Lev09] Theresa Levitt. The shadow of enlightenment: optical and political transparency in France 1789– 1848. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954470-0 (hardcover). 192 pp. LCCN TA1522 .L48 2009. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0907/
                                        2008048422-b.htm;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0907/
                                        2008048422-d.htm;
                                        http://www.loc.gov/catdir/
                                        toc/fy0904/2008048422.htm.

Levitin:2016:NSP

[Lev16] Dmitri Levitin. Newton and scholastic philosophy. British Journal for the History of Science, 49(1):53–77, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lewis:2012:WPO

[Lew12] Rhodri Lewis. William Petty on the order of nature: an unpublished manuscript treatise, volume 399 of Medieval and Renaissance texts and studies. ACMRS (Arizona Center for Medieval and Renaissance Studies), Tempe, AZ, USA, 2012. ISBN 0-86698-447-X. xiii + 176 pp. LCCN BT695.5 .P4335 2012.

Lock:2011:MWN

[LG11] Humfrey Lock and Peter Grund.

Misticall wordes and names infinite: an edition and study of Humfrey Lock’s Treatise on Alchemy, volume 367 of Medieval and Renaissance texts and studies. ACMRS (Arizona Center for Medieval and Renaissance Studies), Tempe, AZ, USA, 2011. ISBN 0-86698-415-1. xii + 350 pp. LCCN QD25 .L63 2011.

Lightman:2009:ENV

[Lig09] Bernard V. Lightman. Evolutionary naturalism in Victorian Britain: the ‘Darwinians’ and their critics, volume 919 of Variorum collected studies series. Ashgate/Variorum, Aldershot, Hants, England and Farnham, England, 2009. ISBN 0-7546-5987-9 (hardcover). various pp. LCCN Q127.G4 L539 2009.

Link:2017:BRP

[Lin17] Fabian Link. Book review: Paul Betts and Stephen A. Smith (eds.), Science, Religion and Communism in Cold War Europe. New York: Palgrave Macmillan, 2016. Pp. 307. ISBN 978-1-137-54638-8. £60.00 (hardback). British Journal for the History of Science, 50(2):373–375, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
24D38208C911FC26E6CFBA5E836E0055.

Lipkowitz:2014:SNH

[Lip14]  Elise S. Lipkowitz.  Seized

natural-history collections and the redefinition of scientific cosmopolitanism in the era of the French revolution. British Journal for the History of Science, 47(1):15–41, March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Livingstone:2015:FRA

[Liv15] David N. Livingstone. Finding revelation in anthropology: Alexander Winchell, William Robertson Smith and the heretical imperative. British Journal for the History of Science, 48(3):435–454, September 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lloyd:2009:DMC

[Llo09]  G.  E.  R.

(Geoffrey Ernest Richard) Lloyd. Disciplines in the making: crosscultural perspectives on elites, learning, and innovation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956787-5 (hardcover). viii + 215 pp. LCCN AZ105 .L495 2009.

Lloyd:2012:BHU

[Llo12]  G.  E.  R.  (Geof-

frey Ernest Richard) Lloyd. Being, humanity, and understanding: studies in ancient and modern societies. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-965472-7 (hardcover). 136 pp. LCCN BD331 .L66 2012.

Lightman:2013:CKB

[LMS13] Bernard V. Lightman, Gordon McOuat, and Larry Stewart, editors. The circulation of knowledge between Britain, India, and China: the early-modern world to the twentieth century, volume 36 (or 3??) of History of science and medicine library; Knowledge infrastructure and knowledge economy. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-24441-7 (hardback), 90-04-25141-3 (ebook). xxi + 339 pp. LCCN Q126.8 .C56 2013.

Long:2009:BMR

[LMSR09] Pamela O. Long, David McGee, Alan M. Stahl, and Franco Rossi, editors. The book of Michael of Rhodes: a Fifteenth-Century maritime manuscript. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-13503-5 (hardback: vol. 1), 0-262-19590-9 (hardback: vol. 2), 0-262-12308-8 (hardback: vol. 3). ???? pp. LCCN ????

Lanuza-Navarro:2014:BRM

[LN14] Tayra M. C. Lanuza-Navarro. Book review: Monica Azzolini, The Duke and the Stars: Astrology and Politics in Renaissance Milan. Cambridge, MA and London: Harvard University Press, 2013. Pp. xiii + 370. ISBN 978-0-674-06663-2. £36.95 (hardback). British Journal for the History of Science, 47(1):176– 178, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lo:2015:BRC

[Lo15]  Melissa  Lo.

Book reviews: Catelijne Coopmans, Janet Vertesi, Michael Lynch and Steve Woolgar (eds.), Representation in Scientific Practice Revisited. Cambridge, MA: MIT Press, 2014. Pp. ix + 366. ISBN 978-0-262-52538-1. £24.95 (paperback). British Journal for the History of Science, 48(4): 722–724, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Loison:2016:BRR

[Loi16] Coline Loison. Book review: Robert Gilmore McKinnell, The Understanding, Prevention and Control of Human Cancer: The Historic Work and Lives of Elizabeth Cavert Miller and James A. Miller. Leiden: Brill, 2016. Pp. xvi + 196. ISBN 978-9-004-28679-5. EUR45.00 (paperback). British Journal for the History of Science, 49(4):508– 509, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
15510F6201FEE129CE3E61013AAD198C.

Lorimer:2013:SRR

[Lor13] Douglas A. Lorimer. Science, race relations and resistance: Britain, 1870–1914. Studies in Imperialism. Manchester University Press, Manchester, UK, 2013. ISBN 0-7190-3357-8. xi + 344 pp. LCCN DA125.A1 L67 2013.

Loukopoulou:2015:BRC

[Lou15] Eleni Loukopoulou. Book reviews: Claire L. Jones, The Medical Trade Catalogue in Britain, 1870–1914. London: Pickering & Chatto, 2013. Pp. xii + 264. ISBN 978-1-84893-443-6. £60.00 (hardback). British Journal for the History of Science, 48(3): 525–526, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Loukopoulou:2016:BRF

[Lou16a] Eleni Loukopoulou. Book review: Frank Brandsma, Carolyne Larrington and Corinne Saunders (eds.), Emotions in Medieval Arthurian Literature: Body, Mind, Voice. Cambridge: D.S. Brewer, 2015. Pp. 221. ISBN 978-1-84384-421-1. £60.00 (hardback). British Journal for the History of Science, 49(4): 638–640, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

org/core/product/
0D70D8C970B8CB80B92F6A0860B1CDE6.

Loukopoulou:2016:BRI

[Lou16b] Eleni Loukopoulou. Book reviews: Innes M. Keighren, Charles W. J. Withers and Bill Bell, Travels into Print: Exploration, Writing, and Publishing with John Murray, 1773–1859. Chicago and London: The University of Chicago Press, 2015. Pp. 392. ISBN 978-0-226-42953-3. $45.00 (hardback). British Journal for the History of Science, 49(2):295–296, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Loukopoulou:2017:BRF

[Lou17] Eleni Loukopoulou. Book review: Fay B. Alberti, This Mortal Coil: The Human Body in History and Culture. Oxford: Oxford University Press, 2016. Pp. xii + 289. ISBN 978-0-19-959903-5. £20.00 (hardback). British Journal for the History of Science, 50(2):343–344, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
2360C86D88B807DC9B5FB5417126C2BF.

Loukopoulou:2019:BRV

[Lou19] Eleni Loukopoulou. Book review: Veronika Fuechtner, Douglas E. Haynes and Ryan M. Jones (eds.), A Global History of Sexual Science 1880–1960. Oakland, CA: University of California Press, 2018. Pp. 477. ISBN 978-0-520-29339-7. $34.95 (paperback).

British Journal for the History of Science, 52(1):175–176, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/veronika-fuechtner-
douglas-e-haynes-and-ryan-
m-jones-eds-a-global-
history-of-sexual-science-
18801960-oakland-ca-
university-of-california-
press-2018-pp-477-isbn-
9780520293397-3495-paperback/
896480400E17F6163D493AB11A774D66.

Loveland:2010:AED

[Lov10] Jeff Loveland. An alternative encyclopedia?: Dennis de Coetlogon’s “Universal history of arts and sciences” (1745), volume 2010:03 of SVEC. Voltaire Foundation, Oxford, UK, 2010. ISBN 0-7294-0992-9 (paperback). ISSN 0435-2866. xiii + 256 pp. LCCN AE5 .L88 2010.

URL http://www.voltaire.ox.
ac.uk/www\_vf/svec/svec.ssi.

Lykknes:2012:BWC

[LOvT12] Annette Lykknes, Donald Opitz, and Brigitte van Tiggelen, editors. For better or for worse?: collaborative couples in the sciences, volume 44 of Science networks. Historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2012. ISBN 3-0348-0285-4 (hardcover), 3-0348-0286-2 (e-book). xiv + 319 pp. LCCN Q125 .F5935

                                        2012. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1304/2012940246-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1304/
                                        2012940246-t.html.

Lowy:2011:SCB

[Löw11] Ilana Löwy. ‘sexual chemistry’ before the Pill: science, industry and chemical contraceptives, 1920–1960. British Journal for the History of Science, 44(2): 245–274, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Leong:2011:SKM

[LR11] Elaine Yuen Tien Leong and Alisha Michelle Rankin, editors. Secrets and knowledge in medicine and science, 1500–1800. The history of medicine in context. Ashgate, Aldershot, Hants, England and Farnham, England, 2011. ISBN 0-7546-6854-1 (hardcover), 0-7546-9501-8 (e-book). ix + 247 pp. LCCN R484 .S43 2011.

Lundgren:2013:PPF

[Lun13] Frans Lundgren. The politics of participation: Francis Galton’s Anthropometric Laboratory and the making of civic selves. British Journal for the History of Science, 46(3):445–466, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Luttge:2019:WIN

[Lüt19] Felix Lüttge. Whaling intelligence: news, facts and US– American exploration in the Pacific. British Journal for the History of Science, 52(3):425– 445, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/whaling-
intelligence-news-facts-
and-usamerican-exploration-
in-the-pacific/
0E5F5FC5CFDBEC50B29E8119E25D3F12.

Livingstone:2011:GNC

[LW11] David N. Livingstone and Charles W. J. Withers, editors. Geographies of Nineteenth-Century science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-48726-1 (hardcover). x + 526 pp. LCCN Q127.G4 G46 2011. UK£35.50.

Lykknes:2016:BRM

[Lyk16] Annette Lykknes. Book review:

Mansoor Niaz, Chemistry Education and Contributions from History and Philosophy of Science. Heidelberg: Springer International Publishing, 2016. Pp. xx + 250. ISBN 978-3-319-26246-8. £74.50 (hardback). British Journal for the History of Science, 49 (4):530–531, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E256CD33F1D8E60885BED8BF318C156D.

Lynn:2010:SIB

[Lyn10] Michael R. Lynn. The sublime invention: ballooning in Europe, 1783–1820, volume 17 of The Enlightenment world. Pickering and Chatto, London, UK, 2010. ISBN 1-84893-016-X (hardcover), 1-84893-017-8 (paperback). ix + 240 pp. LCCN Q175.52.E85 2010; GV762.5.E85 L966 2010.

Lynall:2012:SSS

[Lyn12] Gregory Lynall. Swift and science: the satire, politics, and theology of natural knowledge, 1690– 1730. Palgrave Macmillan, Basingstoke, UK, 2012. ISBN 0-230-34364-3 (hardcover). xi + 209 pp. LCCN PR3728.S35 L96 2012.

Lightman:2014:EVC

[LZ14] Bernard V. Lightman and Bennett Zon, editors. Evolution and Victorian culture. Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02842-6 (hardcover). xviii + 320 pp. LCCN HM626 .E945 2014.

MacLeod:2009:ALF

[Mac09] Roy M. MacLeod. Archibald Liversidge, FRS: imperial science under the Southern Cross. The Royal Society of New South Wales, Sydney, NSW, Australia, 2009. ISBN 1-920898-80-8 (paperback). xvii + 637 pp. LCCN ????

Macleod:2010:BRS

[Mac10] Christine Macleod. Book review: Samuel J. M. M. Alberti, Nature and Culture: Objects, Disciplines and the Manchester Museum. Manchester: Manchester University Press, 2009. Pp. xiii + 239. ISBN 978-0-7190-8114-9. £60.00 (hardback). British Journal for the History of Science, 43 (4):620–622, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Alb09].

Macleod:2011:BRA

[Mac11a] Christine Macleod. Book review: Adrian Johns, Piracy: The Intellectual Property Wars from Gutenberg to Gates. Chicago and London: University of Chicago Press, 2009. Pp. 626. ISBN 978-0-226-40118-8. $35.00 (hardback). British Journal for the History of Science, 44(1):153–155, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

MacLeod:2011:BRL

[Mac11b] Roy MacLeod.

Book review: Lawrence Dritsas, Zambesi: David Livingstone and Expeditionary Science in Africa. New York: I. B. Tauris, 2010. Pp. xii + 242. ISBN 978-1-84511-705-4. £54.40 (hardback). British Journal for the History of Science, 44(2):296–297, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

MacLeod:2013:BRG

[Mac13] Roy MacLeod. Book reviews: Geoffrey Cantor, Religion and the Great Exhibition of 1851. Oxford: Oxford University Press, 2011. Pp. xi + 226. ISBN: 978-0-19-959667-6. £70.00 (hardback). British Journal for the History of Science, 46(2):347– 348, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

MacDonald:2015:MKO

[Mac15] Lee T. MacDonald. Making Kew Observatory: the Royal Society, the British Association and the politics of early Victorian science. British Journal for the History of Science, 48(3):409–433, September 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mackintosh:2016:AOG

[Mac16] Alan Mackintosh. Authority and ownership: the growth and wilting of medicine patenting in Georgian England. British Journal for the History of Science, 49 (4):541–559, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E8F521DD40D01E9168CBD27F77B18E34.

Madison:2016:MBH

[Mad16] Paige Madison. The most brutal of human skulls: measuring and knowing the first Neanderthal. British Journal for the History of Science, 49(3): 411–432, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
7EED9A8CAE36E7AC7B6639EF14DD7FC6.

Maerker:2012:BRR

[Mae12] Anna Maerker. Book review:

Rebecca Messbarger, The Lady Anatomist: The Life and Work of Anna Morandi Manzolini. Chicago and London: The University of Chicago Press, 2010. Pp. xiv + 234. ISBN 978-0-226-52081-0. £35.00 (hardback). British Journal for the History of Science, 45(1):132–133, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Mes10].

Maguire:2012:NCW

[Mag12] Richard Maguire. ‘never a credible weapon’: nuclear cultures in British government during the era of the H-bomb. British Journal for the History of Science, 45 (4):519–533, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mahon:2009:OHM

[Mah09] Basil Mahon. Oliver Heaviside:

maverick mastermind of electricity, volume 36 of IET history of technology series. Institution of Engineering and Technology, London, UK, 2009. ISBN 0-86341-965-8. xvii + 183 + 16 pp. LCCN TK140.H43 M34 2009; QC515.H4 M346 2009; QC515 M34 2009. 60.00 USD.

URL http://lib.myilibrary.
com/detail.asp?ID=226335.

Mahoney:2011:HC

[Mah11] Michael S. (Michael Sean) Mahoney, editor. Histories of computing. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-05568-3. vi + 250 pp. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh.

Mahony:2018:GSC

[Mah18] Martin Mahony. The ‘genie of the storm’: cyclonic reasoning and the spaces of weather observation in the southern Indian Ocean, 1851–1925. British Journal for the History of Science, 51 (4):607–633, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/genie-of-
the-storm-cyclonic-
reasoning-and-the-spaces-
of-weather-observation-in-
the-southern-indian-ocean-
18511925/
E6A6B57055E1E51928083F18FDF87D40.

Main:2011:BRP

[Mai11] Steven J. Main. Book review: Paul R. Josephson, Lenin’s Laureate: Zhores Alferov’s Life in Communist Science. Cambridge, MA and London: MIT Press, 2010. Pp. ii + 307. ISBN 978-0-262-01458-8. £22.95 (hardback). British Journal for the History of Science, 44(4):614–615, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Maienschein:2012:BRA

[Mai12a] Jane Maienschein. Book reviews: Angus McLaren, Reproduction by Design: Sex, Robots, Trees, and Test-Tube Babies. Chicago and London: The University of Chicago Press, 2012. Pp. viii + 235. ISBN 978-0-226-56069-4. $55.00 (hardback). British Journal for the History of Science, 45 (4):697–699, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Maifreda:2012:OPE

[Mai12b] Germano Maifreda. From oikonomia to political economy: constructing economic knowledge from the Renaissance to the scientific revolution. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-3301-3 (hardcover), 1-4094-3302-1 (e-book - PDF). vii + 304 pp. LCCN HB81 .M36 2012.

Main:2012:BRA

[Mai12c] Steven J. Main. Book review:

Anthony Heywood, Engineer of Revolutionary Russia: Iurii V. Lomonosov (1876–1952) and the Railways. Farnham: Ashgate, 2011. Pp. xxvi + 400. ISBN 978-0-7546-553908. £75.00 (hardback). British Journal for the History of Science, 45(2):305– 306, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Makrides:2013:DVS

[Mak13] Vasilios N. Makrides. In a different vein? Scientific development in the Greek Orthodox East. British Journal for the History of Science, 46(2):335–340, June 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Manning:2012:MFE

[Man12] Gideon Manning, editor. Matter and form in early modern science and philosophy, volume 28 of History of science and medicine library, Scientific and learned cultures and their institutions; v. 6. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-21870-X (hardcover), 90-04-22114-X (e-book). ISSN 1872-0684. x + 248 pp. LCCN Q175.32.H95 M38 2012.

Manias:2013:BRC

[Man13a] Chris Manias. Book reviews: Claudine Cohen, La méthode du Zadig: La trace, le fossile, la preuve. Paris: Editions de Seuil, 2011. Pp. 342. ISBN 978-2-02-040298-9. EUR23.00 (paperback). British Journal for the History of Science, 46(1):166– 168, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Manias:2013:BRP

[Man13b] Chris Manias. Book reviews: Paul Crook, Grafton Elliot Smith, Egyptology & the Diffusion of Culture: A Biographical Perspective. Brighton: Sussex Academic Press, 2012. Pp. vii + 160. ISBN 978-1-84519-481-9. £19.95 (paperback). British Journal for the History of Science, 46(4):726–728, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mansfield:2013:LRS

[Man13c] P. Mansfield. The long road to Stockholm: the story of magnetic resonance imaging (MRI): an autobiography. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966454-4a. x + 241 pp. LCCN QC16.M346 A3 2013; Q143.M2265

A3 2013. URL http://www.loc.
gov/catdir/enhancements/
fy1403/2012946411-b.html;
http://www.loc.gov/catdir/
enhancements/fy1403/
2012946411-d.html;
http://www.loc.gov/catdir/
enhancements/fy1403/
2012946411-t.html.

Manning:2014:DBA

[Man14] Gideon Manning. Descartes and the Bologna affair. British Journal for the History of Science, 47(1):1–13, March 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Manias:2015:SBH

[Man15] Chris Manias. Sinanthropus in Britain: human origins and international science, 1920–1939. British Journal for the History of Science, 48(2):289–319, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mankoo:2016:BRU

[Man16] Alex Mankoo. Book reviews:

Ulf Schmidt, Secret Science:

A Century of Poison Warfare and Human Experiments. Oxford: Oxford University Press, 2015. Pp. xxxi + 637. ISBN 978-0-1992-9979-9. £25.00 (hardback). British Journal for the History of Science, 49(1):136– 138, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Manias:2017:BRM

[Man17] Chris Manias. Book review: Marianne Sommer, History Within: The Science, Culture, and Politics of Bones, Organisms, and Molecules. Chicago: The University of Chicago Press, 2016. Pp. viii + 544. ISBN 978-0-226-3473205. $50.00 (cloth). British Journal for the History of Science, 50(2):359–360, June 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
69BFA5FD101C304C8DA2FA3C3DE4A3CD.

Marr:2009:WOF

[Mar09] Alexander Marr, editor. The worlds of Oronce Fine: mathematics, instruments, and print in Renaissance France. Shaun Tyas, Donington, UK, 2009. ISBN 1-900289-96-2. xv + 224 pp. LCCN QA11.A1 W67 2009.

Marr:2011:BRG

[Mar11]  Alexander Marr.  Between

Raphael and Galileo: Mutio Oddi and the mathematical culture of late Renaissance Italy. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-50628-2 (hardcover). xiii + 359 pp. LCCN QA27.I8 M35 2011.

Marsden:2012:BRD

[Mar12a] Ben Marsden. Book review:

David Philip Miller, James Watt, Chemist: Understanding the Origins of the Steam Age. Science and Culture in the Nineteenth Century. London: Pickering & Chatto, 2009. Pp. viii + 241. ISBN 978-1-85196-974-6. £60.00 (hardback). British Journal for the History of Science, 45(2):298– 300, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Martin:2012:BRD

[Mar12b] Alison E. Martin. Book review: Donna Spalding Andréolle and Véronique Molinari (eds.), Women and Science, 17th Century to Present: Pioneers, Activists and Protagonists. Newcastle upon Tyne: Cambridge Scholars, 2011. Pp. xxv + 272. ISBN 978-1-4438-2918-2. £44.99 (hardback). British Journal for the History of Science, 45(2):309– 310, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Martin:2012:BRK

[Mar12c] Alison E. Martin. Book reviews: Kutzinski and Ette (eds.), Alexander von Humboldt. Political Essay on the Island of Cuba: A Critical Edition. Chicago and London: The University of Chicago Press, 2011. Pp. xxvi + 519. ISBN 978-0-226-46568-5. £45.00 (hardback). British Journal for the History of Science, 45 (4):686–687, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Martin:2014:BRV

[Mar14a] Alison E. Martin. Book review: Vera M. Kutzinski, Ottmar Ette and Laura Dassow Walls (eds.), Alexander von Humboldt and the Americas. Berlin: Verlag Walter Frey, 2012. Pp. 295. ISBN 978-3-938944-63-9. EUR19.80 (paperback). British Journal for the History of Science, 47(1):185– 187, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Martin:2014:SAR

[Mar14b] Craig Martin. Subverting Aristotle: religion, history, and philosophy in early modern science. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1316-7 (hardcover), 1-4214-1317-5 (e-book). 262 pp. LCCN B485 .M267 2014.

Marsh:2015:BRF

[Mar15] Oliver Marsh. Book reviews: Francis Graham-Smith, Unseen Cosmos: The Universe in Radio. Oxford: Oxford University Press, 2013. Pp. vii + 256. ISBN 978-0-19-966058-2. £20.00 (hardback). British Journal for the History of Science, 48(4):721–722, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Marschall:2016:BRS

[Mar16] Wythe Marschall. Book review:

Sigrid Schmalzer, Red Revolution, Green Revolution: Scientific Farming in Socialist China. Chicago: The University of Chicago Press, 2016. Pp. x + 320. ISBN 978-0-226-33015-0. £31.50 (hardback). British Journal for the History of Science, 49(4): 512–514, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
35EAABA5608DC0B24FE478703B10D45F.

Marschall:2017:BRB

[Mar17a] Wythe Marschall. Book review:

B. Estella Leopold, Stories from the Leopold Shack: Sand County Revisited. New York: Oxford University Press, 2016. Pp. xvi + 325. ISBN 978-0-19-046322-9. $27.95 (hardback). British Journal for the History of Science, 50(2):363–365, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
21221CAF32CE9B3AA63EF613FF3D3958.

Martin:2017:BRJ

[Mar17b] Rebecca Martin. Book review:

Joanna Ebenstein, The Anatomical Venus. London: Thames & Hudson, 2016. Pp. 224. ISBN 978-0-500-25218-5. £19.95 (quarter bound/PLC). British Journal for the History of Science, 50(2):352–354, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
798F69016983977130369627B5A2ACFC.

Marschall:2018:BRC

[Mar18a] Wythe Marschall. Book review: Carl A. Zimring, Aluminum Upcycled: Sustainable Design in Historical Perspective. Baltimore: Johns Hopkins University Press, 2017. Pp. 199. ISBN 978-1-4214-2186-5. $39.95 (hardcover). British Journal for the History of Science, 51 (2):317–319, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/carl-a-
zimring-aluminum-upcycled-
sustainable-design-in-
historical-perspective-
baltimore-johns-hopkins-
university-press-2017-pp-
199-isbn-9781421421865-
3995-hardcover/
138445A68E8E7BBADC1250D2D4534A28.

Martin:2018:BRR

[Mar18b] Rebecca Martin. Book review: Rana A. Hogarth, Medicalizing Blackness: Making Racial Difference in the Atlantic World, 1780–1840. Chapel Hill: University of North Carolina Press, 2017. Pp. xx + 268. ISBN 978-1-4696-3287-2. $27.95 (paperback). British Journal for the History of Science, 51(2): 321–322, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/rana-a-
hogarth-medicalizing-
blackness-making-racial-
difference-in-the-atlantic-
world-17801840-chapel-hill-
university-of-north-
carolina-press-2017-pp-xx-
268-isbn-9781469632872-2795-
paperback/D175A16C1E1970FEAB086F7B63044861.

Marsh:2019:LCS

[Mar19] Oliver Marsh. Life cycle of a star:

Carl Sagan and the circulation of reputation. British Journal for the History of Science, 52(3): 467–486, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/life-cycle-
of-a-star-carl-sagan-and-
the-circulation-of-
reputation/
43F3EF8A1496EB2E6BE5DA58FABA8D9C.

Mata:2013:BRC

[Mat13] Tiago Mata. Book reviews:

Cold War Social Science: Knowledge Production, Liberal Democracy, and Human Nature. Basingstoke: Palgrave Macmillan, 2012. Pp. xviiii + 270. ISBN 978-0-230-34050-3. £55.00 (hardback), by Mark Solovey and Hamilton Cravens (eds.). British Journal for the History of Science, 46(3):542–543, September 2013. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-

001X (electronic).

Mayer:2010:RIP

[May10] Thomas F. Mayer. The Roman Inquisition’s precept to Galileo (1616). British Journal for the History of Science, 43(3):327– 351, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mayhew:2014:MLL

[May14] Robert J. Mayhew. Malthus: The Life and Legacies of an Untimely Prophet. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72871-8 (hardcover), 0-674-41940-5 (ebook). 284 (est.) pp. LCCN ????

                                        URL http://www.jstor.org/
                                        stable/10.2307/j.ctt6wprn5.

Maydom:2017:BRM

[May17] Katrina Maydom. Book review: Martha Few, For All of Humanity: Mesoamerican and Colonial Medicine in Enlightenment Guatemala. Tucson: University of Arizona Press, 2015. Pp. 304. ISBN 978-0-8165-3187-5. $34.95 (paperback). British Journal for the History of Science, 50 (2):356–358, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        64AB620309AD1D807F5BA834922FDD6A.

McAleer:2013:SWE

[McA13] John McAleer. ‘stargazers at the world’s end’: telescopes, observatories and ‘views’ of empire in the Nineteenth-Century British Empire. British Journal for the History of Science, 46(3):389–413, September 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mcalpine:2016:BRC

[Mca16] Katherine Mcalpine. Book review: Caroline Grigson, Menagerie: The History of Exotic Animals in England, 1100–1837. Oxford: Oxford University Press, 2016. Pp. 368. ISBN 978-0-19-871470-5. £25.00 (hardback). British Journal for the History of Science, 49(3): 485–487, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
996CEECD6C5763E8D53833329E2426A9.

McCray:2012:VHG

[McC12] Patrick (W. Patrick) McCray.

The visioneers: how a group of elite scientists pursued space colonies, nanotechnologies, and a limitless future. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-13983-0 (hardcover). xii + 351 pp. LCCN Q125 .M417 2012.

McCahey:2019:BRD

[McC19] Daniella McCahey. Book review: David Alff, The Wreckage of Intentions: Projects in British Culture, 1660–1730. Alembics. Philadelphia: University of Pennsylvania Press, 2018. Pp. 239. ISBN 978-0-8122-4959-0. £54.00 (cloth). British Journal for the History of Science, 52(2):365– 366, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/david-alff-the-
wreckage-of-intentions-
projects-in-british-
culture-16601730-alembics-
philadelphia-university-of-
pennsylvania-press-2018-pp-
239-isbn-9780812249590-
5400-cloth/1E1A516A288363AE11BCDB8127D80CD5.

McEvoy:2010:HCR

[McE10] John G. McEvoy. The historiography of the chemical revolution: patterns of interpretation in the history of science. Pickering and Chatto, London, UK, 2010. ISBN 1-84893-030-5 (hardcover), 1-84893-031-3 (e-book). xiii + 328 pp. LCCN ????

McGrath:2013:IAO

[McG13] Mary Quinlan McGrath. Influences: art, optics, and astrology in the Italian Renaissance. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92284-7 (hardcover), 0-226-92285-5 (e-book). xi + 284 + 12 pp. LCCN N6915 .M42 2013.

McGuire:2015:BRR

[McG15] Coreen McGuire. Book review: Ruth Oldenziel and Mikael Hård, Consumers, Tinkerers, Rebels: The People Who Shaped Europe. Basingstoke: Palgrave, 2013.

Pp. 440. IBSN 978-0-230-30801-5. £65.00 (hardback). British Journal for the History of Science, 48 (4):711–712, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McGuire:2019:XRD

[McG19] Coreen McGuire. ‘X-rays don’t tell lies’: the Medical Research Council and the measurement of respiratory disability, 1936– 1945. British Journal for the History of Science, 52(3):447– 465, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/xrays-dont-
tell-lies-the-medical-
research-council-and-the-
measurement-of-respiratory-
disability-19361945/
0B366A9BF0B67F17410E44BCCA907FED.

McLaren:2012:RDS

[McL12] Angus McLaren. Reproduction by design: sex, robots, trees, and test-tube babies in interwar Britain. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-56069-4 (hardcover). viii + 235 pp. LCCN PR478.F87 M37 2012.

McMahon:2018:HTR

[McM18] Richard McMahon. The history of transdisciplinary race classification: methods, politics and institutions, 1840s–1940s. British Journal for the History of Science, 51(1):41–67, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/history-of-
                                        transdisciplinary-race-
                                        classification-methods-
                                        politics-and-institutions-
                                        1840s1940s/
                                        F252BE85A4C64596A15DB32C68B0D72B.

McOuat:2010:BRA

[McO10] Gordon McOuat. Book review: Adrian Desmond and James Moore, Darwin’s Sacred Cause: Race, Slavery and the Quest for Human Origins. London: Allen Lane, 2009. Pp. xxi + 485. ISBN 978-1-84614-035-8. £25.00 (hardback). British Journal for the History of Science, 43 (1):119–121, March 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McRae:2014:NDF

[McR14] Kenneth D. (Kenneth Douglas) McRae. Nuclear dawn: F. E. Simon and the race for atomic weapons in World War II. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-968718-8 (hardcover), 0-19-151084-X (e-book). xxviii + 284 pp. LCCN QC774.S5 M37 2014.

McWilliams:2011:MTH

[McW11] Stuart McWilliams. Magical thinking: history, possibility and the idea of the Occult. Continuum, London, UK, 2011. ISBN 1-4411-1697-4 (hardcover), 1-4411-9295-6 (PDF e-book), 1-4411-8429-5 (ePub e-book). 187 pp. LCCN BF1411 .M385 2011.

Medina:2011:CRT

[Med11] Eden Medina. Cybernetic revolutionaries: technology and politics in Allende’s Chile. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-01649-4 (hardcover). xiv + 326 pp. LCCN JL2631 .M43 2011.

Merchant:2012:FBV

[Mer12] Carolyn Merchant. Francis Bacon and the ‘vexations of art’: experimentation as intervention. British Journal for the History of Science, ??(??):1–49, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). Published online, but not yet assigned to a journal volume.

Merchant:2013:FBV

[Mer13] Carolyn Merchant. Francis Bacon and the ‘vexations of art’: experimentation as intervention. British Journal for the History of Science, 46(4):551–599, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Merchant:2019:WOH

[Mer19] Paul Merchant. What oral historians and historians of science can learn from each other. British Journal for the History of Science, 52 (4):673–688, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/what-oral-
historians-and-historians-
of-science-can-learn-from-
each-other/
A2A5ADC7166C7ED18DE21332BC515F66.

Messbarger:2010:LAL

[Mes10] Rebecca Marie Messbarger. The lady anatomist: the life and work of Anna Morandi Manzolini. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-52081-1 (hardcover). xiii + 234 pp. LCCN QM16.M35 M47 2010; NLS HB3.211.1.137.

Methuen:2011:BRK

[Met11] Charlotte Methuen. Book review: Kathleen M. Crowther, Adam and Eve in the Protestant Reformation. Cambridge: Cambridge University Press, 2010. Pp. xii + 293. ISBN 978-0-52119-236-

1. £50.00 / $85.00 (hardback). British Journal for the History of Science, 44(4):582–584, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Meyers:2011:BRH

[Mey11] Todd Meyers. Book review: Hans-Jörg Rheinberger, On Historicizing Epistemology: An Essay. Stanford: Stanford University Press, 2010. Pp. ix + 114. ISBN 978-0-8047-6289-2. $18.95 (paperback). British Journal for the History of Science, 44(1): 120–121, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Marsden:2013:UCE

[MHO13] Ben Marsden, Hazel Hutchison, and Ralph O’Connor, editors. Uncommon contexts: encounters between science and literature, 1800–1914, volume 23 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-362-2 (hardcover). xiii + 239 pp. LCCN PR468.S34 U53 2013.

Mialet:2012:HIS

[Mia12] Hélène Mialet. Hawking Incorporated: Stephen Hawking and the anthropology of the knowing subject. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-52226-1 (hardcover), 0-226-52228-8 (paperback). ix + 266 pp. LCCN QC16.H33 M53 2012.

Micale:2008:HMH

[Mic08] Mark S. Micale. Hysterical men:

the hidden history of male nervous illness. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-03166-0 (hardcover). xiv + 366 pp. LCCN RC532 .M533 2008.

Mick:2011:BRH

[Mic11] Christoph Mick. Book review:

Helmuth Trischler and Mark Walker (eds.), Physics and Politics: Research Support in Twentieth Century Germany in International Perspective. Stuttgart: Franz Steiner Verlag, 2010. Pp. 285. ISBN 978-3-515-09601-0. EUR44.00 (paperback). British Journal for the History of Science, 44(3):470– 472, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2008:JWC

[Mil08] David Philip Miller. James Watt, chemist: understanding the origins of the steam age, volume 8 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2008. ISBN 1-85196-974-8, 1-85196-685-4 (ebook). viii + 241 pp. LCCN TJ140.W38 M55 2009.

Miller:2011:BRB

[Mil11a] David Philip Miller. Book review: Bruce J. Hunt, Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Baltimore: Johns Hopkins University Press, 2010. Pp. x + 182. ISBN 978-0-8018-9359-9. £10.50 (paperback). British Journal for the History of Science, 44(4):609–610, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2011:BRN

[Mil11b] Peter N. Miller. Book review: Nicholas Dew, Orientalism in Louis XIV’s France. Oxford: Oxford University Press, 2009. Pp. xvi + 301. ISBN 978-0-19-923484-4. £55.00 (hardback). British Journal for the History of Science, 44(2):290–291, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mills:2011:BRJ

[Mil11c] Simon Mills. Book review: J.

M. I. Klaver, Scientific Expeditions to the Arab World 1761– 1881. London: Oxford University Press, 2009. Pp. 255. ISBN 978-0-19-956889-5. £95.00 (hardback). British Journal for the History of Science, 44(1):128–130, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mills:2011:BRM

[Mil11d] Simon Mills. Book review:

Maurits H. van den Boogert, Aleppo Observed: Ottoman Syria through the Eyes of Two Scottish Doctors, Alexander and Patrick Russell. Oxford: Oxford University Press, 2010. Pp. 254. ISBN 978-0-19-958856-5. £120.00 (hardback). British Journal for the History of Science, 44(4): 598–600, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2012:BRL

[Mil12] David Philip Miller. Book review: Leslie Tomory, Progressive Enlightenment: The Origins of the Gaslight Industry, 1780– 1820. Cambridge, MA and London: MIT Press, 2012. Pp. xii + 348. ISBN 978-0-262-01675-9. £19.95 (hardback). British Journal for the History of Science, 45 (3):466–467, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2013:BRA

[Mil13] Ian Miller. Book reviews: The Age of Stress: Science and the Search for Stability. Oxford: Oxford University Press, 2013. Pp. iv + 311. ISBN: 978-0-19-9588626. £35.00 (hardback), by Mark Jackson. British Journal for the History of Science, 46(3): 537–538, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2015:BRP

[Mil15]  Ian Miller.  Book reviews:

Pain and Emotion in Modern History. Basingstoke: Palgrave Macmillan, 2014. Pp. xii + 284 ISBN: 978-1-137-37242-0. £65.00 (hardback). Joanna Bourke, The Story of Pain: From Prayer to Painkillers. Oxford: Oxford University Press, 2014. Pp. xii + 396. ISBN: 978-0-19-968942-2. £20.00 (hardback). Rob Boddice (ed.). British Journal for the History of Science, 48(1):191– 193, March 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2016:BRM

[Mil16]  Ian Miller.  Book reviews:

Matthew Smith, Another Person’s Poison: A History of Food Allergy. New York: Columbia University Press, 2015. Pp. 312. ISBN 978-0-2311-6484-9. £19.95 (hardback). British Journal for the History of Science, 49(1):148– 150, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2017:BRR

[Mil17] Ian Miller. Book review: Rob Boddice, The Science of Sympathy: Morality, Evolution, and Victorian Civilization. Springfield: University of Illinois Press, 2016. Pp. iv + 179. ISBN 978-0-252-08205-4. EUR28.00 (paperback). British Journal for the History of Science, 50(3):554– 555, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
91E9925594A8DAEECAA097DE0470F457.

Mimura:2015:AOP

[Mim15] Taro Mimura. The Arabic original of (ps.) Masha’allah’s Liber de orbe: its date and authorship. British Journal for the History of Science, 48(2):321–352, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miskell:2011:BRC

[Mis11] Louise Miskell. Book review:

Charles W. J. Withers, Geography and Science in Britain 1831– 1939: A Study of the British Association for the Advancement of Science. Manchester: Manchester University Press, 2010. Pp. xvii + 278. ISBN 978-0-7190-7976-4. £60.00 (hardback). British Journal for the History of Science, 44 (2):297–298, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miskell:2013:MPS

[Mis13] Louise Miskell, editor. Meeting places: scientific congresses and urban identity in Victorian Britain. Ashgate, Aldershot, Hants, England and Farnham, England, 2013. ISBN 1-4094-5237-9, 1-4094-5238-7 (e-book). LCCN HS67 .M57 2013.

Mitchell:2012:ADE

[Mit12]  Piers D. Mitchell,  edi-

tor. Anatomical dissection in Enlightenment England and beyond: autopsy, pathology, and display. The history of medicine in context. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-1886-3 (hardcover), 1-4094-1887-1 (e-book). viii + 186 pp. LCCN QM33.4 .M58 2012.

URL http://site.ebrary.com/
lib/yale/Doc?id=10556675.

Mitchell:2018:CDP

[Mit18]  Daniel Jon Mitchell.  From

corps to discipline, part one: Charles d’Almeida, Pierre Bertin and French experimental physics, 1840–1880. British Journal for the History of Science, 51(3): 333–368, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/from-corps-
to-discipline-part-one-
charles-dalmeida-pierre-
bertin-and-french-
experimental-physics-

18401880/
4A56BC28EE744F180324F17D457BEAF8.

Mody:2010:TOO

[ML10a] Cyrus C. M. Mody and Michael Lynch. Test objects and other epistemic things: a history of a nanoscale object. British Journal for the History of Science, 43(3): 423–458, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morrison-Low:2010:NLA

[ML10b] A. D. Morrison-Low. Northern lights: the age of Scottish lighthouses. National Museums Scotland, in conjunction with The Royal Scottish Society of Arts, Edinburgh, Scotland, 2010. ISBN 1-905267-47-9 (paperback), 1-905267-57-6 (hardcover). xxv + 262 pp. LCCN VK1061 .M67 2010.

Morrison-Low:2012:EBS

[ML^{+}12] A. D. Morrison-Low et al., editors. From Earth-bound to satellite: telescopes, skills, and networks, volume 23(2) of History of science and medicine library; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-21150-0. xxix + 265 pp. LCCN QB88 .F86 2012.

Mody:2017:BRJ

[Mod17] Cyrus C. M. Mody. Book review:

John Krige, Sharing Knowledge, Shaping Europe: US Technological Collaboration and Nonproliferation. Cambridge, MA: MIT Press, 2016. Pp. 227. ISBN 978-0-262-03477-7. £24.95 (cloth). British Journal for the History of Science, 50(1):165–167, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
4E30963D2275FB93D7928570C738ACB7.

Mohun:2013:RNS

[Moh13] Arwen Mohun. Risk: negotiating safety in American society. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0790-6 (hardcover), 1-4214-0825-2 (e-book). viii + 329 pp. LCCN HD4605 .M63 2013.

Mollan:2014:WPE

[Mol14] R. C. (R. Charles) Mollan, editor. William Parsons, 3rd Earl of Rosse: astronomy and the castle in nineteenth-century Ireland, volume 4 of Royal Dublin society–Science and Irish culture series. Manchester University Press, Manchester, UK, 2014. ISBN 0-7190-9144-6 (hardcover). xxii + 368 pp. LCCN QB36.R77 W55 2014.

Montgomery:2010:BRA

[Mon10] Georgina M. Montgomery. Book review: Amanda Rees, The Infanticide Controversy: Primatology and the Art of Field Science. Chicago: University of Chicago Press, 2009. Pp. ix + 288. ISBN 978-0-226-70711-2. £40.00 (hardback). British Journal for the History of Science, 43(3):502– 503, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Monk:2012:ICL

[Mon12] Ray Monk. Inside the centre:

the life of J. Robert Oppenheimer. Jonathan Cape, London, UK, 2012. ISBN 0-224-06262-X (hardcover). xiv + 818 + 32 pp. LCCN QC16.O62 M66 2012.

Montgomery:2013:DSN

[Mon13] Scott L. Montgomery. Does science need a global language?: English and the future of research. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-53503-7 (hardcover), 0-226-01004-X (e-book). xiii + 226 pp. LCCN Q226 .M658 2013. With a foreword by David Crystal.

Moore:2010:NIN

[Moo10] Richard Moore. Nuclear illusion, nuclear reality: Britain, the United States and nuclear weapons, 1958–64. Nuclear weapons and international security since 1945. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-23067-9 (hardcover). xv + 332 + 8 pp. LCCN UA647 .M66 2010.

Moosa:2013:BRN

[Moo13] Ebrahim Moosa. Book reviews:

Nidhal Guessoum, Islam’s Quantum Question: Reconciling Muslim Tradition and Modern Science. London and New York: I.

B. Tauris, 2011. Pp. xxvi + 403. ISBN 978-1-84885-518-2. £16.99 (paperback). British Journal for the History of Science, 46(4): 736–738, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Moore:2017:BRE

[Moo17] Martin D. Moore. Book review: Erwin H. Ackerknecht, A Short History of Medicine. Baltimore: Johns Hopkins University Press, 2016, pp. xxiii + 247. ISBN 978-1-4214-1954-

1. $29.95 (paperback). British Journal for the History of Science, 50(1):143–145, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
D519D8FECE2ED4196CD8BFC84EF2093B.

Morrell:2010:BRJ

[Mor10a] Jack Morrell. Book review: Joe Bord, Science and Whig Manners: Science and Political Style in Britain, c. 1790–1850. Basingstoke: Palgrave Macmillan, 2009. Pp. ix + 213. ISBN 978-0-230-57484-7. £50.00 (hardback). British Journal for the History of Science, 43(1):121– 122, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morris:2010:SNP

[Mor10b] Peter John Turnbull Morris, editor. Science for the nation: perspectives on the history of the Science Museum. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-23009-1 (hardcover). xxi + 350 + 16 pp. LCCN Q105.G7 S36 2010.

Morus:2013:BRF

[Mor13] Iwan Rhys Morus. Book reviews:

Frank A. J. L. James, Michael Faraday: A Very Short Introduction. Oxford: Oxford University Press, 2010. Pp. xvi + 162. ISBN 978-0-19-957431-5. £7.99 (paperback). Frank A. J. L. James (ed.), Michael Faraday. The Chemical History of a Candle. Oxford: Oxford University Press, 2011. Pp. xlviii + 152. ISBN 978-0-19-969491-4. £14.99 (hardback). British Journal for the History of Science, 46(1):168– 169, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morar:2015:RMT

[Mor15a] Florin-Stefan Morar. Reinventing machines: the transmission history of the Leibniz calculator. British Journal for the History of Science, 48(1):123–146, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Moruno:2015:BRC

[Mor15b] Dolores Martín Moruno. Book reviews: Colin Jones, The Smile Revolution in Eighteenth Century Paris. Oxford: Oxford University Press, 2014. Pp. 256. ISBN 978-0-19-871581-8. £22.99 (hardback). British Journal for the History of Science, 48(3):511–512, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mori:2017:MSS

[Mor17] Giuliano Mori. Mathematical subtleties and scientific knowledge: Francis Bacon and mathematics, at the crossing of two traditions. British Journal for the History of Science, 50(1):1–21, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/product/
0BCD4D7C513321CBC8E364BD0D6FCFF6.

Morris:2018:BRH

[Mor18] Peter J. T. Morris. Book review: Hannah Gay and William P. Griffith, The Chemistry Department at Imperial College: A History, 1845–2000. London: World Scientific Publishing, 2017. Pp. xi + 569 + illus. ISBN 978-1-78326-973-0. £56.00 (hardcover). British Journal for the History of Science, 51(2): 309–311, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/hannah-gay-
and-william-p-griffith-the-
chemistry-department-at-
imperial-college-a-history-
18452000-london-world-
scientific-publishing-2017-
pp-xi-569-illus-isbn-
9781783269730-5600-hardcover/
A5D5F3D8ECA2AC0009CC89066026AA8C.

Moshenska:2014:UEM

[Mos14] Gabriel Moshenska. Unrolling Egyptian mummies in nineteenth-century Britain. British Journal for the History of Science, 47(3): 451–477, September 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mosley:2016:BRP

[Mos16a] Adam Mosley. Book review:

Pietro Daniel Omodeo, Copernicus in the Cultural Debates of the Renaissance: Reception, Legacy, Transformation. Leiden: Brill, 2014. Pp. xiii + 433. ISBN 978-90-04-25178-6. EUR150.00/$205.00 (hardback). British Journal for the History of Science, 49(3): 477–478, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
C4EBD77A96C62A3E02C96F88D6D57A92.

Mosley:2016:BRC

[Mos16b] Adam Mosley. Book reviews:

Craig Martin, Subverting Aristotle: Religion, History, and Philosophy in Early Modern Science. Baltimore: Johns Hopkins University Press, 2014. Pp. 262. ISBN 978-1-4214-1316-7. £35.50 (hardback). British Journal for the History of Science, 49(1):117– 119, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mougey:2017:NCH

[Mou17] Thomas Mougey. Needham at the crossroads: history, politics and international science in wartime China (1942–1946). British Journal for the History of Science, 50(1):83–109, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        BBFAAD904327C2D5E4BBEE87D8844C87.

Moxham:2016:BRM

[Mox16a] Noah Moxham. Book review: Michael Hunter, Boyle Studies: Aspects of the Life and Thought of Robert Boyle (1627–91). Farnham: Ashgate, 2015. Pp. 266. ISBN 978-1-4724-2810-3. £63.00 (hardcover). British Journal for the History of Science, 49(3): 481–482, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        cambridge.org/core/product/
                                        709E7B8E91153B43B459DAB2088A2BB0.

Moxham:2016:ELE

[Mox16b] Noah Moxham. An experimental ‘life’ for an experimental life: Richard Waller’s biography of Robert Hooke (1705). British Journal for the History of Science, 49(1):27–51, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Moxham:2019:NKI

[Mox19] Noah Moxham. Natural Knowledge, Inc.: the Royal Society as a metropolitan corporation. British Journal for the History of Science, 52(2):249–271, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/natural-
knowledge-inc-the-royal-
society-as-a-metropolitan-
corporation/
96AAB89201CCAFDA1F48C3FDE55DD4F3.

Mazzolini:2012:DRS

[MR12] Renato G. (Renato Giuseppe)

Mazzolini and Hans-Jörg Rheinberger, editors. Differing routes to stem cell research: Germany and Italy, volume 26 of Annali dell’Istituto storico italo-germanico in Trento. Contributi. Societa editrice il Mulino, Bologna, Italy, 2012. ISBN 3-428-13849-X (Duncker and Humblot: paperback), 88-15-23878-6 (Societa editrice il Mulino: paperback). LCCN QH588.S83 D57 2012.

Morris-Reich:2013:ASM

[MR13] Amos Morris-Reich. Anthropology, standardization and measurement: Rudolf Martin and anthropometric photography. British Journal for the History of Science, 46(3):487–516, September 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mueggler:2011:PRA

[Mue11]  Erik Mueggler.  The paper

road: archive and experience in the botanical exploration of West China and Tibet. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-26902-0 (hardcover), 0-520-26903-9 (paperback). xiv + 361 pp. LCCN QK355 .M83 2011.

Mullaney:2010:CTN

[Mul10] Thomas S. (Thomas Shawn) Mullaney. Coming to terms with the nation: ethnic classification in modern China, volume 18 of Asia: local studies/global themes. University of California Press, Berkeley, CA, USA, 2010. ISBN 0-520-26278-6 (hardcover). xxiv + 232 pp. LCCN DS730 .M85 2011.

Muller:2011:JWH

[Mül11] Falk Müller. Johann Wilhelm Hittorf and the material culture of Nineteenth-Century gas discharge research. British Journal for the History of Science, 44(2): 211–244, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mullaney:2013:BRP

[Mul13] Thomas S. Mullaney. Book reviews: The Paper Road: Archive and Experience in the Botanical Exploration of West China and Tibet. Berkeley: University of California Press, 2011. Pp. xiv + 361. ISBN 978-0-520-26903-3. £20.95 (paperback), by Erik Mueggler. British Journal for the History of Science, 46(3): 530–532, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Murphy:2010:LSE

[Mur10] Jane H. Murphy. Locating the sciences in Eighteenth-Century Egypt. British Journal for the History of Science, 43(4): 557–571, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Murphy:2012:BRJ

[Mur12] Jane Murphy. Book review: Jed Z. Buchwald and Diane Greco Josefowicz, The Zodiac of Paris: How an Improbable Controversy over an Ancient Egyptian Artifact Provoked a Modern Debate between Religion and Science. Princeton and Oxford: Princeton University Press, 2010. Pp. vii + 428. ISBN 978-0-691-14576-1. £25.95 (hardback). British Journal for the History of Science, 45(2):296–298, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Murphy:2014:MFA

[Mur14] Kathleen S. Murphy. To make Florida answer to its name: John Ellis, Bernard Romans and the Atlantic science of British West Florida. British Journal for the History of Science, 47(1):43–65, March 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Muller-Wille:2012:CHH

[MWR12] Staffan Müller-Wille and Hans-Jörg Rheinberger. A cultural history of heredity. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-54570-9 (hardcover). xiii + 323 pp. LCCN QH438.5 .M8513 2012.

Murphy-Young:2018:BRT

[MY18] Sarah Murphy-Young. Book review: Tania Anne Woloshyn, Soaking Up the Rays: Light Therapy and Visual Culture in Britain, c.1890– 1940. Manchester: Manchester University Press, 2017. Pp. viii + 273. ISBN 978-1-7849-9512-6. £45.00 (hardcover). British Journal for the History of Science, 51(2):316–317, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/tania-anne-
woloshyn-soaking-up-the-
rays-light-therapy-and-
visual-culture-in-britain-
c18901940-manchester-
manchester-university-
press-2017-pp-viii-273-
isbn-9781784995126-4500-hardcover/
2A8DC0D8C686192DAF2E9086C6024737.

Myelnikov:2012:BRD

[Mye12] Dmitriy Myelnikov. Book review: Duncan Wilson, Tissue Culture in Science and Society: The Public Life of a Biological Technique in Twentieth Century Britain. Basingstoke: Palgrave Macmillan, 2011. Pp. x + 182. ISBN 978-0-230-28427-2. £50.00 (hardback). British Journal for the History of Science, 45 (3):475–476, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Myelnikov:2017:CCE

[Mye17] Dmitriy Myelnikov. Cuts and the cutting edge: British science funding and the making of animal biotechnology in 1980s Edinburgh. British Journal for the History of Science, 50(4): 701–728, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
38AA4178089C6F1B4A33C5B2D44AD6F1.

Netz:2011:AP

[N^{+}11] Reviel Netz et al., editors. The Archimedes palimpsest. The Archimedes palimpsest publications. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-01684-3 (hardcover: set), 1-107-01457-3 (hardcover: vol. 1), 1-107-01437-9 (hardcover: vol. 2). ???? pp. LCCN QA31 .A832 2011. URL http:

//assets.cambridge.org/
97811070/14572/cover/
9781107014572.jpg. Published

for the Walters Art Museum.

Nall:2012:BRK

[Nal12] Joshua Nall. Book reviews: K.

Maria D. Lane, Geographies of Mars: Seeing and Knowing the Red Planet. Chicago and London: The University of Chicago Press, 2011. Pp. xiv + 266. ISBN 978-0-226-47078-8. $45.00 (hardback). British Journal for the History of Science, 45(4):692–694, 2012.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nall:2015:BRD

[Nal15] Joshua Nall. Book reviews: Discovery and Classification in Astronomy: Controversy and Consensus. New York: Cambridge University Press, 2013. Pp. xvi + 458. ISBN 978-1-107-03361-0. £30.00/$45.00 (hardback). Steven J. Dick. British Journal for the History of Science, 48(1):189–191, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nasim:2010:OWI

[Nas10] Omar W. Nasim. Observation, working images and procedure: the ‘Great Spiral’ in Lord Rosse’s astronomical record books and beyond. British Journal for the History of Science, 43(3): 353–389, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nasim:2016:BRK

[Nas16] Omar W. Nasim. Book reviews: Klaus Hentschel, Visual Cultures in Science and Technology: A Comparative History. Oxford: Oxford University Press, 2014. Pp. x + 496. ISBN 978-0-19-871781-4. £60.00 (hardback). British Journal for the History of Science, 49(2):288– 289, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Navarro:2010:EDC

[Nav10] Jaume Navarro. Electron diffraction chez Thomson: early responses to quantum physics in Britain. British Journal for the History of Science, 43(2):245– 275, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Navarro:2014:BRS

[Nav14a] Jaume Navarro. Book reviews: Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Cambridge, MA and London: Harvard University Press, 2012. Pp. viii + 579. ISBN 978-0-674-06587-1. £25.95 (hardback). British Journal for the History of Science, 47(3):579–580, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

NavarroBrotons:2014:DSPa

[Nav14b] Víctor Navarro Brotons. Disciplinas, saberes y prácticas: filosofía natural, matemáticas y astronomía eb la sociedad espaänola de la época moderna. (Spanish) [Disciplines, knowledge and practices: natural philosophy, mathematics, and astronomy in modern Spanish society]. Publicacions de la Universitat de València, València, Spain, 2014. ISBN 84-370-9446-1, 84-370-9611-1 (Llibree). 496 pp. LCCN ???? URL http://uv.

odilotk.es/?id=00004324.

Navarro:2015:BRL

[Nav15a] Jaume Navarro. Book reviews: Lynda Walsh, Scientists as Prophets: A Rhetorical Genealogy. Oxford: Oxford University Press, 2013. Pp. xi + 264. ISBN 978-0-19-985711-1. £22.50 (paperback). British Journal for the History of Science, 48(3): 529–530, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Navarro:2015:BRM

[Nav15b] Jaume Navarro. Book reviews: Milena Wazeck, Einstein’s Opponents: The Public Controversy about the Theory of Relativity in the 1920s. Translated by Geoffrey S. Koby. Cambridge: Cambridge University Press, 2014. ISBN 978-1-107-01744-3. £65.00/$99.00 (hardback). British Journal for the History of Science, 48(2): 375–377, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Navarro:2016:BRS

[Nav16] Jaume Navarro. Book reviews: Stephen G. Brush, Making 20th Century Science: How Theories Became Knowledge. New York: Oxford University Press, 2015. Pp. 552. ISBN 978-0-19-997815-1. £25.99 (hardback). British Journal for the History of Science, 49(2):320–321, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Naylor:2010:RSP

[Nay10] Simon Naylor. Regionalizing science: placing knowledges in Victorian England. Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2010. ISBN 1-85196-636-6 (hardcover). xiv + 245 pp. LCCN Q127.G4 NAY 2010.

Naylor:2014:PSF

[Nay14] Ron Naylor. Paolo Sarpi and the first Copernican tidal theory. British Journal for the History of Science, 47(4):661–675, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Neswald:2010:BRG

[Nes10] Elizabeth Neswald. Book review: Galina Kichigina, The Imperial Laboratory: Experimental Physiology and Clinical Medicine in Post-Crimean Russia. Amsterdam and New York: Rodopi. Pp. ii + 374. ISBN 978-90-420-2658-2. £72.20 (hardback). British Journal for the History of Science, 43(3):491–493, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Neuss:2018:BMH

[Neu18] Michael J. Neuss. Blood money:

Harvey’s De motu cordis (1628) as an exercise in accounting. British Journal for the History of Science, 51(2):181–203, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/blood-
money-harveys-de-motu-
cordis-1628-as-an-exercise-
in-accounting/
5E160F467F6A6A5A195CF882C263AE6E.

Nielsen:2012:SSF

[NHR12] Kristian Hvidtfelt Nielsen, Michael Harbsmeier, and Christopher Jacob Ries, editors. Scientists and scholars in the field: studies in the history of fieldwork and expeditions. Aarhus University Press, Aarhus, Danmark, 2012. ISBN 87-7124-014-4. 476 pp. LCCN Q181 .S385 2012.

Nichols:2017:BRC

[Nic17a] Tiffany Nichols. Book review: Chris Bernhardt, Turing’s Vision: The Birth of Computer Science. Cambridge, MA: MIT Press, 2016. Pp. 189. ISBN 978-0-262-03454-

8. £19.95 (cloth). British Journal for the History of Science, 50(2):366–368, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
6FBE184804281FFF83D386E3D7902756.

Nichols:2017:BRH

[Nic17b] Tiffany Nichols. Book review: Harry Collins, Gravity’s Kiss: The Detection of Gravitational Waves. Cambridge, MA and London: MIT Press, 2017. Pp. 416. ISBN 978-0-262-03618-4. £24.95 (hardcover). British Journal for the History of Science, 50(4): 745–746, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
62920C16E2C5D6F7A4C654C5FEE1DE52.

Nichols:2019:BRR

[Nic19] Tiffany Nichols. Book review:

Roberto Lalli, Building the General Relativity and Gravitation Community during the Cold War. Cham: Springer, 2017. Pp. xiv + 168. ISBN 978-3-319-54653-

7. £49.99 (softcover). British Journal for the History of Science, 52(1):178–180, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/roberto-
lalli-building-the-general-
relativity-and-gravitation-
community-during-the-cold-
war-cham-springer-2017-pp-
xiv-168-isbn-9783319546537-
4999-softcover/
8924361F207D0CFF7B18BA353238DE34.

Niebur:2010:SSC

[Nie10a] Louis Niebur, editor. Special sound: the creation and legacy of the BBC Radiophonic Workshop. The Oxford music/media series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-536840-1, 0-19-536841-X. 259 pp. LCCN ML32.G7 N54 2010.

Nielsen:2010:PPD

[Nie10b] Kristian Hvidtfelt Nielsen. Postcolonial partnerships: deep sea research, media coverage and (inter)national narratives on the Galathea Deep Sea Expedition from 1950 to 1952. British Journal for the History of Science, 43(1):75–98, March 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nielsen:2014:WTM

[Nie14] Kristian H. Nielsen. ‘What things mean in our daily lives’: a history of museum curating and visiting in the Science Museum’s Children’s Gallery from c. 1929 to 1969. British Journal for the History of Science, 47(3):505–538, September 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Noakes:2014:IRE

[Noa14] Richard Noakes. Industrial research at the Eastern Telegraph Company, 1872–1929. British Journal for the History of Science, 47(1):119–146, March 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Norrby:2010:NPL

[Nor10] Erling Norrby. Nobel Prizes and life sciences. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2010. ISBN 981-4299-36-7 (hard cover), 981-4299-37-5 (paperback). xv + 317 pp. LCCN QH315 .N667 2010.

Normandin:2012:BRB

[Nor12] Sebastian Normandin. Book reviews: Brian Regal, Searching for Sasquatch: Crackpots, Eggheads, and Cryptozoology. Basingstoke: Palgrave Macmillan, 2011. Pp. xi + 249. ISBN 978-0-230-11147-9. £55.00 (hardback). British Journal for the History of Science, 45 (4):699–700, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

November:2012:BCD

[Nov12]  Joseph  Adam  Novem-

ber. Biomedical computing: digitizing life in the United States, volume 130 series 2 of Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0468-0 (hardcover). xvi + 344 pp. LCCN R858 .N68 2012.

Osseo-Asare:2014:BRS

[OA14] Abena Dove Agyepoma Osseo-Asare. Bitter roots: the search for healing plants in Africa. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-08552-X (hardcover), 0-226-08602-X (paperback), 0-226-08616-X (e-book). vii + 300 pp. LCCN RS181 .O87 2014.

OBrien:2013:SSH

[O’B13] James F. O’Brien. The scientific Sherlock Holmes: cracking the case with science and forensics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-979496-0 (hardcover). xx + 175 pp. LCCN HV8073 .O36 2012.

OConnor:2017:BRG

[O’C17] Ralph O’Connor. Book review: Gowan Dawson, Show Me the Bone: Reconstructing Prehistoric Monsters in Nineteenth-Century Britain and America. Chicago: The University of Chicago Press, 2016. Pp. 476. ISBN 978-0-226-33273-4. $50.00 (cloth). British Journal for the History of Science, 50(1):156–158, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
F9B0A9B5B62D28F3613D7DBEC287D193.

Olszynko-Gryn:2016:FLE

[OG16]  Jesse Olszynko-Gryn.  Film

lessons: early cinema for historians of science. British Journal for the History of Science, 49(2):279– 286, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Olszynko-Gryn:2017:TBL

[OG17] Jesse Olszynko-Gryn. Thin blue lines: product placement and the drama of pregnancy testing in British cinema and television. British Journal for the History of Science, 50(3):495– 520, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
D47CCF928477D0E635F199BFD1E1923F.

Olszynko-Gryn:2017:MRL

[OGE17] Jesse Olszynko-Gryn and Patrick Ellis. ‘A machine for recreating life’: an introduction to reproduction on film. British Journal for the History of Science, 50(3): 383–409, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
65181AD7E67C074DE863BFE01BE3AA44.

Ogle:2015:GTT

[Ogl15] Vanessa Ogle. The global transformation of time: 1870–1950. Harvard University Press, Cambridge, MA, USA, 2015. ISBN 0-674-28614-6 (hardcover). 279 pp. LCCN QB223 .O35 2015.

Oreskes:2014:STG

[OK14] Naomi Oreskes and John Krige, editors. Science and technology in the global cold war. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02795-X (hardcover), 0-262-52653-0 (paperback). ix + 456 pp. LCCN T14.5 .S3768 2014.

OLeary:2006:RCM

[O’L06] Don O’Leary. Roman Catholicism and modern science: a history. Continuum, New York, NY, USA, 2006. ISBN 0-8264-1868-6 (hardcover). xx + 356 pp. LCCN BX1795.S35 O44

2006. URL http://www.loc.
gov/catdir/toc/ecip0617/
2006023343.html.

OLeary:2012:ICS

[O’L12] Don O’Leary. Irish Catholicism and Science: From ‘Godless Colleges’ to the ‘Celtic Tiger’. Cork University Press, Cork, Ireland, 2012. 343 pp.

Olohan:2014:GKL

[Olo14] Maeve Olohan. Gate-keeping and localizing in scientific translation publishing: the case of Richard Taylor and scientific memoirs. British Journal for the History of Science, 47(3):433–450, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Oosterhoff:2012:BRP

[Oos12] Richard Oosterhoff. Book review: Patrick J. Boner (ed.), Change and Continuity in Early Modern Cosmology. Heidelberg, London and New York: Springer, 2011. Pp. xii + 181. ISBN 978-94-007-0036-9. £90.00 (hardback). British Journal for the History of Science, 45(3):457–459, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Oghina-Pavie:2016:BRS

[OP16] Cristiana Oghina-Pavie. Book reviews: Sarah Dewis, The Loudons and the Gardening Press: A Victorian Cultural Industry. Farnham: Ashgate, 2014. Pp. xvi + 278. ISBN 978-1-4094-6922-3. £65.00 (hardback). British Journal for the History of Science, 49(2):296–298, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Opitz:2014:BRT

[Opi14a] Donald L. Opitz. Book reviews:

Tatiana Holway, The Flower of Empire: An Amazonian Water Lily, the Quest to Make It Bloom, and the World It Created. Oxford: Oxford University Press, 2013. Pp. xiv + 306. ISBN 978-0-19-537389-9. £18.99 (hardback). British Journal for the History of Science, 47(4):738–740, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Opitz:2014:SHP

[Opi14b] Donald L. Opitz. ‘The sceptre of her pow’r’: nymphs, nobility, and nomenclature in early Victorian science. British Journal for the History of Science, 47(1):67– 94, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Oswald:2011:JRC

[OPR11] P. H. Oswald, C. D. (Chris)

Preston, and J. Ray, editors. John Ray’s Cambridge catalogue (1660). Ray Society, London, UK, 2011. ISBN 0-903874-43-1. ix + 612 pp. LCCN ????

Olmstead:2008:CAH

[OR08] Alan L. Olmstead and Paul Webb Rhode. Creating abundance: biological innovation and American agricultural development. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85711-2 (hardcover), 0-521-67387-9 (paperback). xii + 467 + 8 pp. LCCN M08.E15094.

MichaelofRhodes:2009:BMR

[oRLM^{+}09] Michael of Rhodes, Pamela O. Long, David McGee, Alan M. Stahl, and Franco Rossi. The book of Michael of Rhodes: a Fifteenth-Century maritime manuscript. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-13503-5 (vol. 1: hardback), 0-262-19590-9 (vol. 2: hardback), 0-262-12308-8 (vol. 3: hardback). xiv + 519 (vol. 1), lii + 679 (vol. 2), xii + 370 (vol. 3) pp. LCCN V46 .M56 2009. URL http://

www.loc.gov/catdir/toc/
ecip0812/2008008611.html.

Ortolano:2009:TCC

[Ort09] Guy Ortolano. The two cultures controversy: science, literature and cultural politics in postwar Britain. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89204-X (hardcover). xi + 295 pp. LCCN DA566.4 .O78 2009. URL http://www.

loc.gov/catdir/toc/
ecip0826/2008036974.html.

Oldfield:2013:VVD

[OS13] Jonathan D. Oldfield and Denis J. B. Shaw. V. I. Vernadskii and the development of biogeochemical understandings of the biosphere, c. 1880s–1968. British Journal for the History of Science, 46(2):287–310, June 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Osler:2010:BRR

[Osl10] Margaret J. Osler. Book review: Robert Bartlett, The Natural and the Supernatural in the Middle Ages. Cambridge: Cambridge University Press, 2008. Pp. x + 170. ISBN 978-0-521-70255-3. £17.99. British Journal for the History of Science, 43(2):291– 292, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Packham:2012:ECV

[Pac12] Catherine Packham. Eighteenthcentury vitalism: bodies, culture, politics. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, 2012. ISBN 0-230-27618-0 (hardback). viii + 251 pp. LCCN PR447 .P33 2012.

URL http://www.netread.com/
jcusers2/bk1388/185/
9780230276185/image/
lgcover.9780230276185.jpg.

Parshall:2011:BRA

[Par11] Karen Hunger Parshall. Book review: Amir Alexander, Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Cambridge, MA and London: Harvard University Press, 2010. Pp. 307. ISBN 978-0-674-04661-0. £21.95 (hardback). British Journal for the History of Science, 44 (1):134–135, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Parsons:2012:BRD

[Par12] Christopher M. Parsons. Book review: Daniela Bleichmar and Peter C. Mancall (eds.), Collecting across Cultures: Material Exchanges in the Early Modern Atlantic World. Philadelphia: University of Pennsylvania Press, 2011. Pp. xvi + 361. ISBN 978-0-8122-4305-5. £32.50 (hardback). British Journal for the History of Science, 45(1): 128–129, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Par12].

Parker:2018:PIP

[Par18] Katherine Parker. Pepys Island as a Pacific stepping stone: the struggle to capture islands on early modern maps. British Journal for the History of Science, 51 (4):659–677, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/pepys-
island-as-a-pacific-
stepping-stone-the-
struggle-to-capture-
islands-on-early-modern-
maps/
132B1CF739CB5BCD048581872EE7A728.

Pattison:2009:DS

[Pat09] Andrew Pattison. The Darwins of Shrewsbury. History Press, Stroud, UK, 2009. ISBN 0-7524-4867-6 (paperback). 127 pp. LCCN CT788.D37 P38 2009; QH31.

Pauly:2007:FPH

[Pau07] Philip J. Pauly. Fruits and plains:

the horticultural transformation of America. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02663-2. xi + 336 pp. LCCN SB319 .P38

2007. URL http://www.loc.
gov/catdir/toc/ecip0721/
2007025274.html.

Payne:2016:BRA

[Pay16]  Gabriale Payne.  Book re-

view: Anna Winterbottom and Facil Tesfaye (eds.), Histories of Medicine and Healing in the Indian Ocean World, vol. 2: The Modern Period. Basingstoke: Palgrave Macmillan, 2016. Pp. 296. ISBN 978-1-137-56761-1. £63.00 (hardback). British Journal for the History of Science, 49(4): 645–646, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
5243A3A6DDE92F16EB37EC2AB50D8F17.

Paul:2013:PPS

[PB13] Diane B. Paul and Jeffrey P. (Jeffrey Paul) Brosco. The PKU paradox: a short history of a genetic disease. Johns Hopkins biographies of disease. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-1131-8 (paperback), 1-4214-1132-6 (e-book). xxiv + 289 pp. LCCN RC632.H87 P38 2013.

Pedersen:2010:OOJ

[PdCB10] Kurt Møller Pedersen, Peter de Clercq, and Thomas Bugge. An Observer of Observatories: the Journal of Thomas Bugge’s Tour of Germany, Holland and England in 1777. Aarhus University Press, Aarhus, Denmark, 2010. ISBN 87-7934-311-2. xxiii + 220 pp. LCCN ????

Parker:2011:ENT

[PDG11] Joan Parker, Paul K. Dayton, and Michael H. Graham, editors. The essential naturalist: timeless readings in natural history. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-30569-4, 0-226-30570-8. xiii + 534 pp. LCCN ????

Pearce:2012:BRR

[Pea12] Jacob V. Pearce. Book review: Robert H. Sanders, The Dark Matter Problem: A Historical Perspective. Cambridge: Cambridge University Press, 2010. Pp. viii + 205. ISBN 978-0-521-11301-4. £35.00 (hardback). British Journal for the History of Science, 45(2):306–307, June 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Peace:2018:BRJ

[Pea18] David Peace. Book review: J. David Archibald, Origins of Darwin’s Evolution: Solving the Species Puzzle through Time and Place. New York and Chichester: Columbia University Press, 2017. Pp. xii + 192. ISBN 978-0-231-17684-2. £54.95 (hardcover). British Journal for the History of Science, 51(2):314–316, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/j-david-
archibald-origins-of-
darwins-evolution-solving-
the-species-puzzle-through-
time-and-place-new-york-
and-chichester-columbia-
university-press-2017-pp-
xii-192-isbn-9780231176842-5495-
hardcover/2D302D6C4126973E98F17CC394FFF88F.

Pearson:2019:BRM

[Pea19] Chris Pearson. Book review:

Michael Worboys, Julie-Marie Strange and Neil Pemberton, The Invention of the Modern Dog: Breed and Blood in Victorian Britain. Baltimore: Johns Hopkins University Press, 2018. Pp. viii + 282. ISBN 978-1-4214-2658-7. $39.95/£29.50 (hardcover). British Journal for the History of Science, 52(3): 528–529, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michael-
worboys-juliemarie-strange-
and-neil-pemberton-the-
invention-of-the-modern-
dog-breed-and-blood-in-
victorian-britain-

                                        baltimore-johns-hopkins-
                                        university-press-2018-pp-viii-
                                        282-isbn-9781421426587-39952950-
                                        hardcover/A4D8594868ACBE9F21A4BBB2EA2886EE.

Perovic:2012:CRF

[Per12] Sanja Perovic. The calendar in revolutionary France: perceptions of time in literature, culture, politics. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-02595-8. xiv + 276 pp. LCCN CE77 .P47

                                        2012. URL http://assets.
                                        cambridge.org/97811070/
                                        25950/cover/9781107025950.
                                        jpg;   http://www.loc.gov/
                                        catdir/enhancements/fy1210/
                                        2012013663-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1210/
                                        2012013663-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1210/
                                        2012013663-t.html.

Pestre:2012:DTS

[Pes12] Dominique Pestre. Debates in transnational and science studies: a defence and illustration of the virtues of intellectual tolerance. British Journal for the History of Science, 45(3):425–442, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Petrunic:2010:BRE

[Pet10a] Josipa G. Petrunić. Book review: Eleanor Robson and Jacqueline Stedall (eds.), The Oxford Handbook of the History of Mathematics. Oxford: Oxford University Press, 2009. Pp. vii + 918.

ISBN 978-0-19-921312-2. £85.00 (hardback). British Journal for the History of Science, 43(2):289– 291, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pettit:2010:PRI

[Pet10b] Michael Pettit. The problem of raccoon intelligence in behaviourist America. British Journal for the History of Science, 43(3):391–421, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pettit:2013:SDP

[Pet13] Michael (Michael John) Pettit.

The science of deception: psychology and commerce in America. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92374-6 (hardcover), 0-226-92375-4 (e-book). ix + 311 pp. LCCN BF637.D42 P48 2013.

Peterson:2014:CVE

[Pet14] Erik Peterson. The conquest of vitalism or the eclipse of organicism? The 1930s Cambridge organizer project and the social network of mid-twentieth-century biology. British Journal for the History of Science, 47(2):281–304, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Petrou:2018:ASD

[Pet18] Marissa H. Petrou. Apes, skulls and drums: using images to make ethnographic knowledge in imperial Germany. British Journal for the History of Science, 51(1): 69–98, 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/apes-skulls-and-
drums-using-images-to-make-
ethnographic-knowledge-in-
imperial-germany/
8E2156F18100DB6AF64478215190E554.

Phalkey:2010:BRJ

[Pha10]  Jahnavi Phalkey.  Book re-

views: Jagdish N. Sinha, Science, War and Imperialism: India in the Second World War. Leiden and Boston: Brill, 2008. Pp. xiv + 278. ISBN 978-90-04-16645-5. EUR79.00 (paperback). Itty Abraham (ed.), South Asian Cultures of the Bomb: Atomic Publics and the State in India and Pakistan. Bloomington: Indiana University Press, 2009. Pp. ix + 222. ISBN 978-0-253-22032-5. $24.95 (paperback). British Journal for the History of Science, 43(2):285–286, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Phalkey:2013:ASB

[Pha13] Jahnavi Phalkey. Atomic state: big science in twentieth-century India. The Indian century. Permanent Black, Ranikhet, India, 2013. ISBN 81-7824-376-8. xvii + 335 pp. LCCN QC773.3.I4 P43 2013.

Phillips:2012:AND

[Phi12] Denise Phillips. Acolytes of nature: defining natural science in Germany, 1770–1850. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-66737-5 (hardcover), 0-226-66739-1 (ebook). viii + 356 pp. LCCN

Q127.G3. URL http://public.
eblib.com/choice/
publicfullrecord.aspx?p=
914967;   http://site.
ebrary.com/id/10559564.

Pickstone:2010:BRRa

[Pic10a] John V. Pickstone. Book review:

Robert Bud, Penicillin: Triumph and Tragedy. Oxford: Oxford University Press, 2008. Pp ix + 330. ISBN 978-0-19-954161-4. £16.99 (paperback). British Journal for the History of Science, 43(1):138–139, March 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pickstone:2010:BRRb

[Pic10b] John V. Pickstone. Book reviews:

Ruth Schwartz Cowan, Heredity and Hope: The Case for Genetic Screening. Cambridge, MA and London: Harvard University Press, 2008. Pp. 292. ISBN 978-0-674-02424-3. £20.95 (hardback). S. Peter Harper, A Short History of Medical Genetics. Oxford: Oxford University Press, 2008. Pp. ix + 557. ISBN 978-0-19-518750-2. £22.99 (hardback). British Journal for the History of Science, 43(2):317–319, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pickstone:2012:OO

[Pic12] John V. Pickstone. Observations on observations. British Journal for the History of Science, 45(4): 671–675, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pick:2014:PNM

[Pic14] Daniel Pick. The pursuit of the Nazi mind: Hitler, Hess, and the analysts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-967851-0 (paperback). xii + 357 + 16 pp. LCCN DD256.5 .P495 2014.

Pinney:2011:PA

[Pin11] Christopher Pinney. Photography and anthropology. Exposures. Reaktion, London, UK, 2011. ISBN 1-86189-804-5 (paperback). 174 pp. LCCN ???? £17.95.

Pind:2012:BRF

[Pin12] Jörgen L. Pind. Book review: Fernando Vidal, The Sciences of the Soul: The Early Modern Origins of Psychology. Chicago and London: Chicago University Press, 2011. Pp. xiv + 413. ISBN 978-0-226-85586-8. £35.50 (hardback). British Journal for the History of Science, 45(2):283– 285, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pind:2013:BRM

[Pin13] Jörgen L. Pind. Book reviews: Max Wertheimer: On Perceived Motion and Figural Organization. Cambridge, MA and London: MIT Press, 2012. Pp. xvi + 296. ISBN 978-0-262-01746-6. £27.95 (hardback), by Lothar Spillmann (ed.). British Journal for the History of Science, 46(3):536–537, September 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pisano:2016:BRS

[Pis16] Raffaele Pisano. Book review:

Stephen Gaukroger, The Natural and the Human: Science and the Shaping of Modernity 1739– 1841. New York: Oxford University Press, 2016. Pp. 416. ISBN 978-0-19-875763-4. £30.00 (hardback). British Journal for the History of Science, 49(3): 488–490, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
3B2311025F8017BD1699A6EDBC51BC38.

Pisano:2017:BRR

[Pis17]  Raffaele Pisano.  Book re-

view: Robert Fox, Science without Frontiers: Cosmopolitanism and National Interests in the World of Learning, 1870–1940. Corvallis: Oregon State University Press, 2016, Pp. 168. ISBN 978-0-87071-867-0. $22.95 (paperback). British Journal for the History of Science, 50(4): 735–736, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/product/
                                        1EB05E4BD5D172F28EA33949387D384B.

Papanelopoulou:2009:PST

[PNGP09] Faidra Papanelopoulou, Agustí Nieto-Galan, and Enrique Perdriguero, editors. Popularizing science and technology in the European periphery, 1800–2000. Science, technology and culture, 1700–1945. Ashgate, Aldershot, Hants, England and Farnham, England, 2009. ISBN 0-7546-6269-1 (hardcover). xviii + 284 pp. LCCN Q225.2.E85 PAP 2008.

Podgorny:2013:FDP

[Pod13] Irina Podgorny. Fossil dealers, the practices of comparative anatomy and British diplomacy in Latin America, 1820– 1840. British Journal for the History of Science, 46(4):647–674, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Podolsky:2015:AER

[Pod15] Scott H. Podolsky. The antibiotic era: reform, resistance, and the pursuit of a rational therapeutics. The Johns Hopkins University Press, Baltimore, MD, USA, 2015. ISBN 1-4214-1593-3 (hardcover), 1-4214-1594-1 (ebook). xii + 309 pp. LCCN RM267 .P64 2015.

Poole:2010:WMS

[Poo10] William Poole. The world makers: scientists of the Restoration and the search for the origins of the earth. The past in the present.

Peter Lang, Oxford, UK, 2010. ISBN 1-906165-08-4. xvii + 234 pp. LCCN Sci.478; NC/POO; QE11 POO 2010.

Portuondo:2009:SSS

[Por09] María M. Portuondo. Secret science: Spanish cosmography and the new world. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-67534-3 (hardcover). xii + 335 + 8 pp. LCCN QB29 .P67 2009. URL http://

www.loc.gov/catdir/toc/
ecip0825/2008034282.htm.

Poskett:2014:BRB

[Pos14]  James Poskett.  Book

reviews: Bernard Lightman, Gordon McOuat and Larry Stewart (eds.), The Circulation of Knowledge between Britain, India and China: The Early-Modern World to the Twentieth Century. Leiden: Brill, 2013. pp. xxi + 339. ISBN: 978-90-04-24441-2. $146.00 (hardback). British Journal for the History of Science, 47(3): 567–569, September 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Poskett:2015:BRW

[Pos15a] James Poskett. Book reviews:

William E. Burns, The Scientific Revolution in Global Perspective. New York: Oxford University Press, 2015. Pp. 216. ISBN 978-0-19-998933-1. £16.99 (paperback). British Journal for the History of Science, 48(4):689–690, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Poskett:2015:SSM

[Pos15b] James Poskett. Sounding in silence: men, machines and the changing environment of naval discipline, 1796–1815. British Journal for the History of Science, 48(2):213–232, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Powers:2012:ICH

[Pow12] John C. Powers. Inventing chemistry: Herman Boerhaave and the reform of the chemical arts. Synthesis. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-67760-5 (hardcover). viii + 260 pages pp. LCCN QD15 .P69 2012.

Priestley:2010:SOM

[Pri10] Mark Priestley. A science of operations: machines, logic and the invention of programming. History of computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-554-4 (hardcover), 1-84882-555-2 (e-book). ix + 341 pp. LCCN QA76.6 .P737 2010.

Prince:2011:ERE

[Pri11] Sue Ann Prince, editor. Of elephants and roses: encounters with French natural history, 1790–1830: an exhibition guide. American Philosophical Society Museum, Philadelphia, PA, USA, 2011. ISBN 0-87169-267-8 (paperback). xxvi + 268 pp. LCCN QH147 .O33 2011.

Priestley:2012:BRM

[Pri12] Mark Priestley. Book reviews:

Michael Sean Mahoney, Histories of Computing. Cambridge, MA and London: Harvard University Press, 2011. Pp. x + 250. ISBN 978-0-674-05568-1. £36.95 (hardback). British Journal for the History of Science, 45(4): 703–704, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Priestman:2013:PED

[Pri13a] Martin Priestman. The poetry of Erasmus Darwin: enlightened spaces, romantic times. Ashgate, Farnham, Surrey, 2013. ISBN 1-4724-1954-5 (hardcover), 1-4724-1955-3 (e-book PDF), 1-4724-1956-1 (e-book ePUB). xiii + 310 pp. LCCN PR3396 .P75 2013.

Principe:2013:SA

[Pri13b] Lawrence Principe. The secrets of alchemy. Synthesis. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-68295-1 (hardcover), 0-226-92378-9 (e-book), 0-226-10379-X. v + 281 + 8 pp. LCCN QD13 .P75 2013.

Pringle:2013:BRG

[Pri13c]  Yolana Pringle.  Book re-

views: Graham Richards, ‘Race’, Racism and Psychology: Towards a Reflexive History. London and New York: Routledge, 2012. Pp. xxi + 511. ISBN 978-0-415-56142-6. £25.95 (paperback).

British Journal for the History of Science, 46(4):735–736, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Prins:2015:EIW

[Pri15] Jacomien Prins. Echoes of an invisible world: Marsilio Ficino and Francesco Patrizi on cosmic order and music theory, volume 234 of Brill studies in intellectual history. Brill, Leiden, The Netherlands, 2015. ISBN 90-04-27437-5 (hardcover), 90-04-28176-2 (e-book). xii + 461 pp. LCCN ML290.2 .P75 2015.

Primbault:2019:BRF

[Pri19a] Simon Dumas Primbault. Book review: Flavia Marcacci, Cieli in contraddizione: Giovanni Battista Riccioli e il terzo sistema del mondo. Perugia: Aguaplano, 2018. Pp. 257. ISBN 978-8-8858-0311-4. EUR20.00 (paperback). British Journal for the History of Science, 52(4):707– 709, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/flavia-
marcacci-cieli-in-
contraddizione-giovanni-
battista-riccioli-e-il-
terzo-sistema-del-mondo-
perugia-aguaplano-2018-pp-
257-isbn-9788885803114-
2000-paperback/
BB0FDA749227B00C331653F90FD2011E.

Prins:2019:BRK

[Pri19b]  Jacomien Prins.  Book re-

view: Karsten Mackensen, Musik und die Ordnung der Dinge im ausgehenden Mittelalter und in der frühen Neuzeit. Frankfurt am Main: Peter Lang, 2017. Pp. 353. ISBN 978-3-631-71992-

3. EUR74.95/£51.95 (hardcover). British Journal for the History of Science, 52(2):361–363, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/karsten-mackensen-
musik-und-die-ordnung-der-
dinge-im-ausgehenden-
mittelalter-und-in-der-
fruhen-neuzeit-frankfurt-
am-main-peter-lang-2017-pp-
353-isbn-9783631719923-
74955195-hardcover/A05D56EBF5B40900D61EE3DB33C54DB0.

Prkachin:2017:BRK

[Prk17] Yvan Prkachin. Book review:

Katja Guenther, Localization and Its Discontents: A Genealogy of Psychoanalysis & the Neurodisciplines. Chicago: The University of Chicago Press, 2016. Pp. 310. ISBN 978-0-226-28820-

8. £24.50 (cloth). British Journal for the History of Science, 50(1):162–164, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
CDE53F2843FB7049A4B89E43B8A0A2C7.

Prkachin:2018:BRT

[Prk18] Yvan Prkachin. Book review: Tara H. Abraham, Rebel Genius: Warren S. McCulloch’s Transdisciplinary Life in Science. Cambridge, MA and London: MIT Press, 2016. Pp. 305. ISBN 978-0-262-03509-5. $40.00 (hardcover). British Journal for the History of Science, 51 (2):313–314, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/tara-h-
                                        abraham-rebel-genius-
                                        warren-s-mccullochs-
                                        transdisciplinary-life-in-
                                        science-cambridge-ma-and-
                                        london-mit-press-2016-pp-
                                        305-isbn-9780262035095-
                                        4000-hardcover/
                                        B0FF66E09AD99A0B07EAFF18012DD2D4.

Proctor:2011:BRS

[Pro11] Katie Proctor. Book review: Sally Gregory Kohlstedt, Teaching Children Science: Hands-On Nature Study in North America 1890–1930. Chicago and London: Chicago University Press, 2010. Pp. xv + 363. ISBN 978-0-226-44990-6 £29.00 (hardback). British Journal for the History of Science, 44(2):302–304, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pumfrey:2011:YAO

[Pum11] Stephen Pumfrey. ‘your astronomers and ours differ exceedingly’: the controversy over the ‘new star’ of 1572 in the light of a newly discovered text by Thomas Digges. British Journal for the History of Science, 44 (1):29–60, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pumfrey:2012:BRR

[Pum12] Stephen Pumfrey. Book review: Robert Goulding, Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History. Heidelberg, London and New York: Springer, 2010. Pp. xx + 201. ISBN 978-90-481-3541-7. £90.00 (hardback). British Journal for the History of Science, 45(2):286– 288, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pyenson:2012:ARN

[Pye12] Lewis Pyenson. Athena’s retinue: Nineteenth-Century scientists embedded in the army. British Journal for the History of Science, 45(3):377–400, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pyle:2010:BRD

[Pyl10] Andrew Pyle. Book review: Daniel Garber, Leibniz: Body, Substance, Monad. Oxford: Oxford University Press, 2009. Pp. xiii + 428. ISBN 978-0-19-956664-8. £35.00 (hardback). British Journal for the History of Science, 43(3): 487–488, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Quinlan-McGrath:2013:IAO

[QM13] Mary Quinlan-McGrath. Influences: art, optics, and astrology in the Italian Renaissance. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92284-7 (hardcover), 0-226-92285-5 (e-book). xi + 284 pp. LCCN N6915 .Q56 2013.

Quirke:2012:BRC

[Qui12]  Viviane Quirke.  Book re-

view: Christoph Gradmann and Jonathan Simon (eds.), Evaluating and Standardizing Therapeutic Agents 1890–1950. Basingstoke: Palgrave Macmillan, 2010. Pp. xiv + 266. ISBN 978-0-230-20281-8. £60.00 (hardback). British Journal for the History of Science, 45(3):470–472, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Quick:2016:CSM

[Qui16] Tom Quick. A capital Scot: microscopes and museums in Robert E. Grant’s zoology (1815– 1840). British Journal for the History of Science, 49(2):173– 204, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Qureshi:2011:PPE

[Qur11] Sadiah Qureshi. Peoples on parade: exhibitions, empire, and anthropology in Nineteenth-Century Britain. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-70096-8 (hardcover). 382 pp. LCCN GN36.G72.

Rabinovitch:2017:LAD

[Rab17] Oded Rabinovitch. A learned artisan debates the system of the world: Le Clerc versus Mallemant de Messange. British Journal for the History of Science, 50 (4):603–636, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
696DA0182DCAA1F4850C28181B66E77B.

Radkau:2009:MWB

[Rad09] Joachim Radkau. Max Weber:

a biography. Polity, Cambridge, UK, 2009. ISBN 0-7456-4147-4. xix + 683 pp. LCCN HM479.W42 R3313 2009.

Radick:2013:BRG

[Rad13] Gregory Radick. Book reviews:

Gordon Chancellor and John van Wyhe (eds.), Charles Darwin’s Notebooks from the Voyage of the ‘Beagle’. Foreword by Richard Darwin Keynes. Cambridge: Cambridge University Press, 2009. Pp. xxxiv + 615. ISBN 978-0-521-51757-7. £104.00 (hardback). John van Wyhe (ed.), Charles Darwin’s Shorter Publications 1829–1883. Cambridge: Cambridge University Press, 2009. Foreword by Janet Browne and Jim Secord. Pp. xxvi + 529. ISBN 978-0-521-88809-7.

£97.00 (hardback). Edmund Russell, Evolutionary History: Uniting History and Biology to Understand Life on Earth. Studies in Environment and History. Cambridge: Cambridge University Press, 2011. Pp. xxii + 216. ISBN 978-0-521-74509-3. £16.99 (paperback). British Journal for the History of Science, 46(2):349– 351, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Radick:2014:BRS

[Rad14]  Gregory Radick.  Book re-

views: Staffan Müller–Wille and Hans-Jörg Rheinberger, A Cultural History of Heredity. Chicago and London: The University of Chicago Press, 2012. Originally published in German in 2009. Pp. xiii + 323. ISBN 978-0-226-54570-7. $50.00 (hardback). Bernd Gausemeier, Staffan Müller–Wille and Edmund Ramsden (eds.), Human Heredity in the Twentieth Century. Studies for the Society for the Social History of Medicine, 15. London: Pickering & Chatto, 2013. Pp. xviii + 302. ISBN 978-1-84893-426-9. £60.00 (hardback). British Journal for the History of Science, 47 (4):747–748, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Radick:2016:PAE

[Rad16] Gregory Radick. Presidential address: Experimenting with the scientific past. British Journal for the History of Science, 49(2):153– 172, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Radick:2017:BRR

[Rad17] Gregory Radick. Book review:

Robert J. Richards and Lorraine Daston (eds.), Kuhn’s Structure of Scientific Revolutions at Fifty: Reflections on a Science Classic. Chicago: The University of Chicago Press, 2016. Pp. 208. ISBN 978-0-226-31720-5. £17.50 (paperback). British Journal for the History of Science, 50(3): 562–563, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
133381163B69A277166B19DBCE86DB49.

Raj:2010:ICL

[Raj10] Kapil Raj. Introduction: circulation and locality in early modern science. British Journal for the History of Science, 43(4): 513–517, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rajaram:2015:BRB

[Raj15] Megha Rajaram. Book reviews:

Breast Cancer in the Eighteenth Century. London: Pickering & Chatto, 2013. Pp. xii + 181. ISBN 978-1-84893-364-4. £60.00 (hardback). Marjo Kaartinen. British Journal for the History of Science, 48(1):177–178, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Raphael:2011:BRM

[Rap11] Renée J. Raphael. Book review: Matteo Valleriani, Galileo Engineer. Dordrecht: Springer, 2010. Pp. xv + 320. ISBN 978-90-481-8644-0. £90.00 (hardback). British Journal for the History of Science, 44(4):589–590, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Raphael:2012:BRA

[Rap12]  Renée Raphael.  Book re-

view: Alexander Marr, Between Raphael and Galileo: Mutio Oddi and the Mathematical Culture of Late Renaissance Italy. Chicago and London: The University of Chicago Press, 2011. Pp. xvii + 359. ISBN 978-0-226-50628-9. £29.00 (hardback). British Journal for the History of Science, 45 (1):126–128, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Mar11].

Rasmussen:2014:GJL

[Ras14] Nicolas Rasmussen. Gene jockeys: life science and the rise of biotech enterprise. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1340-X (hardcover), 1-4214-1341-8 (electronic). viii + 249 pp. LCCN RS380 .R37 2014.

Ratcliff:2009:QIM

[Rat09] Marc Ratcliff. The quest for the invisible: microscopy in the Enlightenment. Ashgate, Aldershot, Hants, England and Farnham, England, 2009. ISBN 0-7546-6150-4 (hardcover). xii + 315 pp. LCCN QH204 .R38 2009. URL

http://www.loc.gov/catdir/
toc/ecip081/2007039630.html.

Ratcliff:2016:TMC

[Rat16] Jessica Ratcliff. Travancore’s magnetic crusade: geomagnetism and the geography of scientific production in a princely state. British Journal for the History of Science, 49(3):325– 352, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
6ECB938D50CCCBC3885F3FA4601203CA.

Raven:2011:WNE

[Rav11] Diederick Raven. What needs to be explained about modern science? British Journal for the History of Science, 44(3): 449–454, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ray:2015:DAW

[Ray15] Meredith K. Ray. Daughters of alchemy: women and scientific culture in early modern Italy. Harvard University Press, Cambridge, MA, USA, 2015. ISBN 0-674-50423-2. 291 pp. LCCN Q130 .R37 2015.

Ruiz-Castell:2011:PCP

[RC11]  Pedro Ruiz-Castell.  Priority

claims and public disputes in astronomy: E. M. Antoniadi, J. Comas i Solà and the search for authority and social prestige in the early twentieth century. British Journal for the History of Science, 44(4):509–531, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rey-Courtel:2013:MHQ

[RC13] Anne-Lise Rey-Courtel. Méthode et histoire: quelle histoire font les historiens des sciences et des techniques?. (French) [Method and history: what history are historians of science and technology making?], volume 6 of Histoire et philosophie des sciences. Classiques Garnier, Paris, France, 2013. ISBN 2-8124-1419-7 (paperback), 2-8124-1420-0. ISSN 2117-3508. 513 pp. LCCN Q126.9 .M45 2014.

Rader:2014:LDR

[RC14] Karen A. (Karen Ann) Rader and Victoria E. M. Cain. Life on display: revolutionizing U.S. museums of science and natural history in the twentieth century. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-07966-X (hardcover). xiv + 467 pp. LCCN QH70.U6 R33 2014.

Reisch:2010:NPE

[RCK10] Gregor Reisch, Andrew Cunningham, and Sachiko Kusukawa. Natural philosophy epitomised: a translation of books 8–11 of Gregor Reisch’s Philosophical pearl (1503). Ashgate, Aldershot, Hants, England and Farnham, England, 2010. ISBN 0-7546-0612-0 (hardcover). lxxiv + 346 pp. LCCN QC17 .R4513 2010.

Rechenberg:2010:WHS

[Rec10] Helmut Rechenberg. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN QC16.H35 R434 2010. URL http://

ebooks.ciando.com/book/
index.cfm/bok_id/46829;
http://www.ciando.com/img/
books/width167/3540692223_
k.jpg;   http://www.ciando.
com/pictures/bib/
3540692223bib_t_1.jpg.

Rectenwald:2013:SCN

[Rec13] Michael Rectenwald. Secularism and the cultures of Nineteenth-Century scientific naturalism. British Journal for the History of Science, 46(2):231–254, June 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rees:2009:ICP

[Ree09] Amanda Rees. The infanticide controversy: primatology and the art of field science. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-70711-3 (cloth). 288 pp. LCCN QL762.5 .R44 2009.

Rees:2010:BRG

[Ree10] Amanda Rees. Book review: Gavin Schaffer, Racial Science and British Society, 1930–62. London and New York: Palgrave Macmillan, 2008. Pp. x + 234. ISBN 978-0-230-00892-2. £45.00 (hardback). British Journal for the History of Science, 43(3): 501–502, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Reed:2014:BRK

[Ree14] Peter Reed. Book reviews: Kenneth Bertrams, Nicholas Coupain and Ernst Homburg, Solvay: History of a Multinational Family Firm. Cambridge: Cambridge University Press, 2013. Pp. xviii + 630. ISBN 978-1-107-02480-9. £60.00 (hardback). British Journal for the History of Science, 47 (4):743–745, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Reeds:2015:BRB

[Ree15] Karen Reeds. Book reviews: Barbara Zipser (ed.), Simon of Genoa’s Medical Lexicon. London: Versita, 2013, Pp. 164. ISBN 978-83-7656-022-9. $69.95 (hardback). British Journal for the History of Science, 48(3): 508–509, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Reed:2016:BRJ

[Ree16a] Peter Reed. Book reviews: James Rodger Fleming and Ann Johnson (eds.), Toxic Airs: Body, Place, Planet in Historical Perspective. Pittsburgh: University of Pittsburgh Press, 2014. Pp. xiv + 284. ISBN 978-0-8229-6290-8. $28.95 (paperback). British Journal for the History of Science, 49 (1):150–151, March 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Reed:2016:BRK

[Ree16b] Peter Reed. Book reviews:

Kathryn Steen, The American Synthetic Organic Chemicals Industry: War and Politics, 1910– 1930. Chapel Hill: University of North Carolina Press, 2014. Pp. xii + 403. ISBN 978-1-4696-1290-

4. £32.50 (paperback). British Journal for the History of Science, 49(2):312–314, June 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rees:2016:I

[Ree16c] Amanda Rees. Introduction.

British Journal for the History of Science, 49(3):383–386, September 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

cambridge.org/core/product/
821D7900FE8C77330F04BE4C32C29BB9.

Rees:2016:SSB

[Ree16d] Amanda Rees. Stories of stones and bones: disciplinarity, narrative and practice in British popular prehistory, 1911–1935. British Journal for the History of Science, 49(3):433–451, September 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

cambridge.org/core/product/
F670EFF0011919D058E52BD83014A217.

Reed:2017:BRR

[Ree17] Peter Reed. Book review: Raymond G. Stokes and Ralf Banken, Building on Air: The International Industrial Gases Industry, 1886–2006. New York: Cambridge University Press, 2016. Pp. xvii + 465. ISBN 978-1-107-03312-2. £64.99/$99.99 (hardback). British Journal for the History of Science, 50(3):555– 556, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E679A9247A66B38D617F39ED5108C252.

Reed:2018:BRH

[Ree18] Peter Reed. Book review: Hallam Stevens, Biotechnology and Society: An Introduction. Chicago: The University of Chicago Press, 2016. Pp. 397. ISBN 978-0-226-04601-3. £21.00/$30.00 (paperback). British Journal for the History of Science, 51(1):172– 173, 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/hallam-stevens-

biotechnology-and-society-
an-introduction-chicago-
the-university-of-chicago-
press-2016-pp-397-isbn-
9780226046013-21003000-
paperback/
D809328B2CF28A58D92CBC1C61A6262D.

Reed:2019:BRC

[Ree19a] Peter Reed. Book review: Carin Berkowitz and Bernard Lightman (eds.), Science Museums in Transition: Cultures of Display in Nineteenth-Century Britain and America. Pittsburgh, PA: University of Pittsburgh Press, 2017. Pp. xi + 375. ISBN 978-0-8229-4475-

1. $45.00 (hardcover). British Journal for the History of Science, 52(1):168–170, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/carin-
berkowitz-and-bernard-
lightman-eds-science-
museums-in-transition-
cultures-of-display-in-
nineteenthcentury-britain-
and-america-pittsburgh-pa-
university-of-pittsburgh-press-
2017-pp-xi-375-isbn-9780822944751-
4500-hardcover/E3D1E404C6BCB74289BED7BA1EF7D145.

Rees:2019:E

[Ree19b] Amanda Rees. Editorial. British Journal for the History of Science, 52(4):541, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/editorial/
                                        E39BEE5444A2372EF6888DC6C6B01792.

Regier:2017:BRU

[Reg17]  Jonathan Regier.  Book re-

view: Ulinka Rublack, The Astronomer and the Witch: Johannes Kepler’s Fight for His Mother. Oxford: Oxford University Press, 2015. Pp. 400. ISBN 978-0-19-873677-6. £20.00 (hardback). British Journal for the History of Science, 50 (2):346–348, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        91E12F3E81B579F0F040D937D613C94D.

Reisner:2015:BRP

[Rei15]  Clemens Reisner.  Book re-

views: Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca Lemov, Thomas Sturm and Michael D. Gordin, How Reason Almost Lost Its Mind: The Strange Career of Cold War Rationality. Chicago and London: The University of Chicago Press, 2013. Pp. viii + 259. ISBN 978-0-226-04663-1. £24.50 (hardback). British Journal for the History of Science, 48(3):539–540, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Remmert:2011:PSR

[Rem11] Volker R. Remmert. Picturing the scientific revolution, volume 4 of Early Modern Catholicism and the Visual Arts Series. Saint Josephs University Press, Philadelphia, PA, USA, 2011. ISBN 0-916101-67-3. 295 pp. LCCN Q222 .R4513 2011.

Renwick:2010:BRJ

[Ren10]  Chris  Renwick.

Book reviews: Joachim Radkau, Max Weber: A Biography. Trans. Patrick Camiller. Cambridge and Malden, MA: Polity, 2009. Pp. xix + 683. ISBN 978-0-7456-4147-8. £25.00 (hardback). Fritz Ringer, Max Weber: An Intellectual Biography. Chicago and London: University of Chicago Press, 2004. Pp. ix + 307. ISBN 978-0-226-72005-0. $21.00 (paperback). British Journal for the History of Science, 43(3):496–498, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Renwick:2011:PES

[Ren11] Chris Renwick. From political economy to sociology: Francis Galton and the social-scientific origins of eugenics. British Journal for the History of Science, 44 (3):343–369, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Renwick:2012:BSL

[Ren12] Chris Renwick. British sociology’s lost biological roots: a history of futures past. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2012.

ISBN 0-230-35616-8 (hardcover). xvii + 240 pp. LCCN ???? Foreword by Steve Fuller.

Renwick:2015:BRA

[Ren15] Chris Renwick. Book reviews: Alison Bashford, Global Population: History, Geopolitics, and Life on Earth. New York: Columbia University Press, 2014. Pp. xii + 466. ISBN 978-0-231-14766-8. £34.50 (hardback); 978-0-231-51952-6. £34.50 (e-book). Robert J. Mayhew, Malthus: The Life and Legacies of an Untimely Prophet. Cambridge, MA and London: The Belknap Press, 2014. Pp. 284. ISBN 978-0-674-72871-4. £20.00 (hardback). British Journal for the History of Science, 48(2):365–367, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rentetzi:2019:BRG

[Ren19] Maria Rentetzi. Book review: Gayle Greene, The Woman Who Knew Too Much: Alice Stewart and the Secrets of Radiation. Ann Arbor: University of Michigan Press, 2017. 2nd edn. Pp. xxvi + 321. ISBN 978-0-472-05356-8. $21.95 (paperback). British Journal for the History of Science, 52(1):176–178, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/gayle-
greene-the-woman-who-knew-

too-much-alice-stewart-and-
the-secrets-of-radiation-
ann-arbor-university-of-
michigan-press-2017-2nd-
edn-pp-xxvi-321-isbn-
9780472053568-2195-
paperback/0593D57B582893E59930B38614C88EC3.

Russell:2012:ERC

[RH12]  Colin  Archibald  Rus-

sell and John A. Hudson. Early railway chemistry and its legacy. Royal Society of Chemistry, Cambridge, UK, 2012. ISBN 1-84973-326-0 (paperback). xiii + 192 pp. LCCN ????

Rheinberger:2010:HEE

[Rhe10] Hans-Jörg Rheinberger. On historicizing epistemology: an essay. Cultural memory in the present. Stanford University Press, Stanford, CA, USA, 2010. ISBN 0-8047-6288-0 (cloth), 0-8047-6289-9 (paperback). 114 pp. LCCN Q175.32.K45 R48513 2010.

Richards:2008:TSL

[Ric08] Robert J. (Robert John)

Richards. The tragic sense of life: Ernst Haeckel and the struggle over evolutionary thought. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-71214-1 (hardcover). xx + 551 + 8 pp. LCCN QH31.H2 R53

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2007039155-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007039155-d.html;
http://www.loc.gov/catdir/
toc/fy0804/2007039155.html.

Richmond:2010:BRJ

[Ric10]  Jesse Richmond.  Book re-

view: Joe Cain and Michael Ruse (eds.), Descended from Darwin: Insights into the History of Evolutionary Studies, 1900–1970. Philadelphia: American Philosophical Society, 2009. Pp. xxvi + 360. ISBN 978-1-60618-991-7. $35.00 (paperback). British Journal for the History of Science, 43(3):500–501, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Richards:2011:RRP

[Ric11a] Graham Richards. Race, racism and psychology: towards a reflexive history. Psychology Press, New York, NY, USA, second edition, 2011. ISBN 0-415-56141-8 (hardcover), 0-415-56142-6 (paperback). xxi + 511 pp. LCCN BF76.45 .R53 2011.

Richards:2011:BRA

[Ric11b] Michael Richards. Book review: Amparo Gómez Rodríguez and Antonio Fco. Canales Serrano (eds.), Ciencia y fascismo: La ciencia española de posguerra. Barcelona: Laertes, 2009. Pp. 358. ISBN 978-84-7584-657-6. EUR19.50 (paperback). British Journal for the History of Science, 44(1):148–149, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Richard:2012:BRI

[Ric12] Nathalie Richard. Book review: Ian Hesketh, The Science of History in Victorian Britain: Making the Past Speak. London: Pickering & Chatto, 2011. Pp. xii + 229. ISBN 978-1-84893-126-8. £60.00 (hardback). British Journal for the History of Science, 45 (1):138–139, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Hes11].

Richards:2013:WHD

[Ric13a] Robert J. (Robert John)

Richards. Was Hitler a Darwinian?: disputed questions in the history of evolutionary theory. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-05876-X (hardcover), 0-226-05893-X (paperback), 0-226-05909-X (e-book). 269 pp. LCCN QH361 .R534 2013.

Richardson:2013:ADA

[Ric13b] Angelique Richardson, editor. After Darwin: animals, emotions, and the mind, volume 93 of Clio medica: perspectives in medical humanities. Rodopi, Amsterdam, The Netherlands, 2013. ISBN 90-420-3747-4 (hardcover), 94-012-0998-7 (e-book). xvi + 369 pp. LCCN QP401 .A42 2013.

Richmond:2014:BRJ

[Ric14] Sheldon Richmond. Book reviews: Joseph Agassi, The Very Idea of Modern Science: Francis Bacon and Robert Boyle. Heidelberg, New York and London: Springer, 2013. Pp. xvii + 315. ISBN 978-94-007-5350-1. £90.00 (hardback). British Journal for the History of Science, 47(3): 570–572, September 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Richter:2016:BRJ

[Ric16] Linda Richter. Book review: James R. Fleming, Inventing Atmospheric Science: Bjerknes, Rossby, Wexler, and the Foundations of Modern Meteorology. Cambridge, MA: MIT Press, 2016. Pp. x + 296. ISBN 978-0-262-03394-

7. £22.95 (cloth). British Journal for the History of Science, 49 (4):652–654, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
4A279C139CC64CDFE6C7F08019351749.

Richardson:2018:BRM

[Ric18] Brian Richardson. Book review: Mel Gooding, David Mabberley and Joe Studholme, Joseph Banks’ Florilegium: Botanical Treasures from Cook’s First Voyage. London: Thames & Hudson, 2017. Pp. 320. ISBN 978-0-500-51936-3. £65.00 (hardcover). British Journal for the History of Science, 51(4):706–708, December 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/mel-gooding-david-
mabberley-and-joe-
studholme-joseph-banks-
florilegium-botanical-

treasures-from-cooks-first-
voyage-london-thames-
hudson-2017-pp-320-isbn-
9780500519363-6500-hardcover/
666796754A33705C624462CD615CD8AB.

Rieppel:2016:BRD

[Rie16] Lukas Rieppel. Book review: David J. Meltzer, The Great Paleolithic War: How Science Forged an Understanding of America’s Ice Ages Past. Chicago: The University of Chicago Press, 2015. Pp. 680. ISBN 978-0-226-29322-6. $55.00 (hardback). British Journal for the History of Science, 49 (4):501–503, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
ADA5443E64E9442A7590C18A124A50C6.

Rihll:2012:BRD

[Rih12] T. E. Rihll. Book review: Daryn Lehoux, What Did the Romans Know? An Inquiry into Science and Worldmaking. Chicago and London: The University of Chicago Press, 2012. Pp. xii + 275. ISBN 978-0-226-47114-3. £29.00 (hardback). British Journal for the History of Science, 45 (3):453–454, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ringer:2004:MWI

[Rin04] Fritz K. Ringer. Max Weber: an intellectual biography. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-72004-7 (hardcover), 0-226-72005-5 (paperback). 307 pp. LCCN HM479.W42 R56 2004.

Ritvo:2010:NCH

[Rit10] Harriet Ritvo. Noble cows and hybrid zebras: essays on animals and history. University of Virginia Press, Charlottesville, VA, USA, 2010. ISBN 0-8139-3060-X (hardcover). x + 239 pp. LCCN QL85 .R58 2010.

Robson:2008:MAI

[Rob08] Eleanor Robson. Mathematics in ancient Iraq: a social history. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-09182-X (cloth). xxvii + 441 pp. LCCN QA22 .R629

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/fy0904/
                                        2007041758.html.

Roberts:2010:BRM

[Rob10a] Lissa Roberts. Book review: Matthew D. Eddy, The Language of Mineralogy: John Walker, Chemistry and the Edinburgh Medical School, 1750–1800. Aldershot: Ashgate, 2008. Pp. xxi + 309. ISBN 978-0-7546-6332-4. £60.00 (hardback). British Journal for the History of Science, 43(2):299– 301, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Robson:2010:BRD

[Rob10b] Eleanor Robson. Book review: Daryn Lehoux, Astronomy, Weather, and Calendars in the Ancient World: Parapegmata and Related Texts in Classical and Ancient Near Eastern Societies. Cambridge and New York: Cambridge University Press, 2007. Pp. xiv + 566. ISBN 978-0-521-85181-7. £65.00 (hardback). British Journal for the History of Science, 43(2):288–289, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roberts:2011:CCA

[Rob11] Lissa Roberts, editor. Centres and cycles of accumulation in and around the Netherlands during the early modern period, volume 2 of Low countries studies on the circulation of natural knowledge. LIT Verlag, Berlin, Germany; Wien, Austria; Zurich, Switzerland, 2011. ISBN 3-643-90095-3. ii + 290 pp. LCCN ????

Roberts:2014:BRW

[Rob14] Peder Roberts. Book reviews: W.

Patrick McCray, The Visioneers: How a Group of Elite Scientists Pursued Space Colonies, Nanotechnologies, and a Limitless Future. Princeton and Oxford: Princeton University Press, 2013. Pp. xii + 351. ISBN 978-0-691-13983-8. £19.95 (hardback). British Journal for the History of Science, 47(3):580–581, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roberts:2015:BRP

[Rob15] Peder Roberts. Book reviews: P.

J. Capelotti, Shipwreck at Cape Flora: The Expeditions of Benjamin Leigh Smith, England’s Forgotten Arctic Explorer. Calgary: University of Calgary Press, 2013. Pp. xxix + 269. ISBN 978-1-55238-705-4. US$41.95 (paperback). British Journal for the History of Science, 48(2):371–372, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roberts:2017:BRR

[Rob17] Peder Roberts. Book review: Ronald E. Doel, Kristine C. Harper and Matthias Heymann (eds.), Exploring Greenland: Cold War Science and Technology on Ice. New York: Palgrave Macmillan, 2016. Pp. xiii + 311. ISBN 978-1-137-59687-1. £63.00 (hardback). British Journal for the History of Science, 50(1):167–168, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
CF9E2AF1214F15A71478855F4146E3A9.

Rodriguez:2015:BRSa

[Rod15a] Ángel R. Rodriguez. Book reviews: Sarah Ferber, Bioethics in Historical Perspective. Basingstoke: Palgrave Macmillan, 2013. Pp. xiv + 233. ISBN 978-1-4039-8724-2. £18.99 (paperback). British Journal for the History of Science, 48(3):537–538, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See corrigendum [Rod15b].

Rodriguez:2015:BRSb

[Rod15b] Ángel R. Rodriguez. Book reviews: Sarah Ferber, Bioethics in Historical Perspective. Basingstoke: Palgrave Macmillan, 2013. Pp. xiv + 233. ISBN 978-1-4039-8724-2. £18.99 (paperback).

— Corrigendum. British Journal for the History of Science, 48(4):725, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Rod15a].

Roe:2017:BRJ

[Roe17]  Alan Roe.  Book review:

Jonathan D. Oldfield and Denis J.B. Shaw, The Development of Russian Environmental Thought: Scientific and Geographical Perspectives on the Natural Environment. Abingdon: Routledge, 2016. Pp. 196. ISBN 978-0-415-58059-5. £95.00 (hardback). British Journal for the History of Science, 50(1):168–170, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
A9DF48F6D511CC1091584DD74E565846.

Roos:2011:WNM

[Roo11] Anna Marie Eleanor Roos. Web of nature: Martin Lister (1639– 1712), the first arachnologist, volume 22 / 16 of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-20703-1 (hardback). ISSN 1872-0684. xx + 478 pp. LCCN QH31.L6913 R66 2011.

Roos:2012:BRH

[Roo12a] Anna Marie Roos. Book review: Hiro Hirai, Medical Humanism and Natural Philosophy: Renaissance Debates on Matter, Life, and the Soul. Leiden: Brill, 2012. Pp. xiii + 227. ISBN 978-90-04-21871-0. EUR99.00 (hardback). British Journal for the History of Science, 45(4):682– 683, December 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
74A336FBE6857C6873DBA467FE027C6B.

Roos:2012:BRJ

[Roo12b] Anna Marie Roos. Book review: John C. Powers, Inventing Chemistry: Herman Boerhaave and the Reform of the Chemical Arts. Chicago and London: The University of Chicago Press, 2012. Pp. viii + 260. ISBN 978-0-226-67760-6. £26.00 (hardback). British Journal for the History of Science, 45(3):460–461, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roos:2012:BRP

[Roo12c] Anna Marie Roos. Book review: P. H. Oswald and C. D. Preston (eds.), John Ray’s Cambridge Catalogue (1660). London: The Ray Society, 2011. Pp. ix + 612. ISBN 978-0903874-43-4. £75.00 (hardback). British Journal for the History of Science, 45(2):291–293, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roos:2015:BRN

[Roo15] Anna Marie Roos. Book reviews: Notebooks, English Virtuosi, and Early Modern Science. Chicago and London: The University of Chicago Press, 2014. Pp. xvii + 398. ISBN 978-0-226-10656-4. $45.00/£31.50 (hardback). British Journal for the History of Science, 48(2):360– 361, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roos:2017:TNT

[Roo17]  Anna Marie Roos.  Taking

Newton on tour: the scientific travels of Martin Folkes, 1733–1735. British Journal for the History of Science, 50(4): 569–601, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
7AF0C34D20AA39F35C8E274D59E1868E.

Rosner:2010:AMB

[Ros10] Lisa Rosner. The anatomy murders: being the true and spectacular history of Edinburgh’s notorious Burke and Hare, and of the man of science who abetted them in the commission of their most heinous crimes. University of Pennsylvania Press, Philadelphia, 2010. ISBN 0-8122-4191-6 (hardcover). viii + 328 pp. LCCN HV6535.G6 R568 2010.

Rostvik:2015:BRL

[Røs15a] Camilla Mørk Røstvik. Book reviews: Legitimizing ESS: Big Science as a Collaboration across Boundaries. Lund: Nordic Academic Press, 2013. Pp. 240. ISBN 978-91-87351-10-5. £30.95 (hardback). Thomas Kaiserfeld and Tom O’Dell (eds.). British Journal for the History of Science, 48 (1):188–189, March 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rostvik:2015:BRC

[Røs15b] Camilla Mørk Røstvik. Book reviews: Catherine Jolivette (ed.), British Art in the Nuclear Age. Farnham: Ashgate, 2014. Pp. xv + 275. ISBN 978-1-4724-1276-8. £70.00 (hardback). British Journal for the History of Science, 48 (4):713–714, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ross:2016:BRT

[Ros16]  Ryan Ross.  Book review:

Tom Crook and Mike Esbester (eds.), Governing Risks in Modern Britain: Danger, Safety and Accidents, c. 1800–2000. Basingstoke: Palgrave Macmillan, 2016. Pp. xiv + 315. ISBN 978-1-137-46744-7. £63.00 (hardback). British Journal for the History of Science, 49(4):654– 655, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
98B59439A1972BF70AFA73EAA8BDD2AE.

Rose:2017:BRJ

[Ros17]  Edwin Rose.  Book review:

James Delbourgo, Collecting the World: The Life and Curiosity of Hans Sloane. London: Allen Lane, 2017. Pp. xxxi + 504. ISBN 978-1-84614-6572. £25.00 (hardback). British Journal for the History of Science, 50(4): 731–732, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
FE02C45B1B566F688D2F1D5A47D6F836.

Rose:2018:SSS

[Ros18] Edwin D. Rose. Specimens, slips and systems: Daniel Solander and the classification of nature at the world’s first public museum, 1753–1768. British Journal for the History of Science, 51(2):205–237, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/specimens-
slips-and-systems-daniel-
solander-and-the-
classification-of-nature-
at-the-worlds-first-public-
museum-17531768/
D9B06CA7B73563A6E217B9EC6F73BBD7.

Rothera:2016:BRJ

[Rot16] Evan C. Rothera. Book review:

John M. Dixon, The Enlightenment of Cadwallader Colden: Empire, Science, and Intellectual Culture in British New York.

Ithaca, NY: Cornell University Press, 2016. Pp. xi + 243. ISBN 978-0-8014-4803-4. $35.00 (hardback). British Journal for the History of Science, 49(4): 648–649, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        928837C0587F6D9FAA54883593AFBB7B.

Rowland:2010:BRS

[Row10] Ingrid D. Rowland. Book review: Samuel Y. Edgerton, The Mirror, the Window, and the Telescope: How Renaissance Linear Perspective Changed Our Vision of the Universe. Ithaca and London: Cornell University Press, 2009. Pp. xvi + 199. ISBN 978-0-8014-7480-4. £16.50 (paperback). British Journal for the History of Science, 43(4):609–610, December 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Edg09].

Rowlinson:2012:SJD

[Row12] J. S. (John Shipley) Rowlinson. Sir James Dewar, 1842– 1923: a ruthless chemist. Science, technology and culture. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-0613-X (hardcover), 1-4094-0614-8 (e-book). ISSN 1700-1945. xviii + 236 pp. LCCN QD22.D345 R69 2012.

Rowlands:2014:BRR

[Row14] Peter Rowlands. Book reviews: Robert Fox (ed.), Thomas Harriot and His World: Mathematics, Exploration and Natural Philosophy in Early Modern England. Farnham: Ashgate, 2012. Pp. xviii + 255. ISBN 978-0-7546-6960.9. £65.00 (hardback). British Journal for the History of Science, 47(3):569–570, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roy:2012:SMN

[Roy12]  Rohan Deb Roy.  Science,

medicine and new imperial histories. British Journal for the History of Science, 45(3):443– 450, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ruse:2009:CCO

[RR09] Michael Ruse and Robert J.

(Robert John) Richards, editors. The Cambridge companion to the “Origin of species”. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-87079-8 (hardcover), 0-521-69129-X (paperback). xxvii + 395 pp. LCCN QH365.O8 R865

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0808/2008000484-b.html;
http://www.loc.gov/catdir/
enhancements/fy0808/
2008000484-d.html;
http://www.loc.gov/catdir/
enhancements/fy0808/
2008000484-t.html.

Reich:2012:CFG

[RRL12] Karin Reich, Elena Roussanova, and Werner Lehfeldt. Carl Friedrich Gauss und Russland: sein Briefwechsel mit in Russland wirkenden Wissenschaftlern. (German) [Carl Friedrich Gauss and Russia: his correspondence with scientists at work in Russia], volume 16 of Abhandlungen der Akademie der Wissenschaften zu Göttingen. Neue Folge. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2012. ISBN 3-11-025306-2, 3-11-025307-0 (ebook). ISSN 0930-4304. xxiii + 905 pp. LCCN QA29.G3 R38 2012.

Robson:2009:OHH

[RS09]  Eleanor  Rob-

son and Jacqueline A. Stedall, editors. Oxford handbook of the history of mathematics. Oxford handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-921312-7 (hardcover). vii + 918 pp. LCCN QA21 .O94

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0907/2008031793-b.html;
http://www.loc.gov/catdir/
enhancements/fy0907/
2008031793-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0823/2008031793.html.

Rietmann:2017:KCM

[RSA^{+}17] Felix Rietmann, Mareike Schildmann, Caroline Arni, Daniel Thomas Cook, Davide Giuriato, Novina Göhlsdorf, and Wangui Muigai. Knowledge of childhood: materiality, text, and the history of science — an interdisciplinary round table discussion. British Journal for the History of Science, 50(1):111–141, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E65847B590B934A42298BFEB10E36CAA.

Rudge:2013:BRN

[Rud13] David Rudge. Book reviews:

Nancy Anderson and Michael R. Dietrich (eds.) The Educated Eye: Visual Culture and Pedagogy in the Life Sciences. Hanover: Dartmouth College Press, 2012. Pp. viii + 318. ISBN 978-1-61168-044-7. $39.95 (paperback). British Journal for the History of Science, 46(4):734–735, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rudwick:2014:EDH

[Rud14] M. J. S. Rudwick. Earth’s deep history: how it was discovered and why it matters. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-20393-X (hardcover), 0-226-20409-X (ebook). ix + 360 pp. LCCN QE11 .R827 2014.

Ruse:2011:BRE

[Rus11a] Michael Ruse. Book review: Evelyn Fox Keller, The Mirage of a Space between Nature and Nurture. Durham and London: Duke University Press, 2010. Pp. viii + 107. ISBN 978-0-8223-4731-

6. £12.99 (paperback). British Journal for the History of Science, 44(4):617–618, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Russell:2011:EHU

[Rus11b] Edmund Russell. Evolutionary history: uniting history and biology to understand life on Earth. Studies in environment and history. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-76211-1 (hardback), 0-521-74509-8 (paperback). xxi + 216 pp. LCCN QH366.2 .R88 2011.

Rushton:2014:WBC

[Rus14a] Alan R. Rushton. William Bateson and the chromosome theory of heredity: a reappraisal. British Journal for the History of Science, 47(1):147–171, March 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Russell:2014:JWM

[Rus14b] Ben Russell. James Watt: making the world anew. Reaktion Books, London, UK, 2014. ISBN 1-78023-375-2. 280 pp. LCCN TA140.W3 R87 2014. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy1512/
                                        2015375141-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1512/
                                        2015375141-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1512/
                                        2015375141-t.html.

Riddle:2012:HHA

[RVG12] John M. Riddle, Anne Van Arsdall, and Timothy Graham, editors. Herbs and healers from the Ancient Mediterranean through the Medieval West: essays in honor of John M. Riddle. Medicine in the medieval Mediterranean. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-0038-7 (hardcover), 1-4094-4723-5 (e-book). xvi + 377 pp. LCCN R135 .H47 2012eb.

URL http://site.ebrary.com/
lib/yale/Doc?id=10568769.

Simeunek:2011:LGJ

[Š^{+}11] Michal Šimêunek et al. The letters on G. J. Mendel: correspondence of William Bateson, Hugo Iltis, and Erich von Tschermak-Seysenegg with Alois and Ferdinand Schindler, 1902–1935, volume 28 of Studies in the History of Sciences and Humanities. Centre for Sciences and Humanities of the Institute of Contemporary History of the Academy of Sciences of the Czech Republic, Prague, Czech Republic, 2011. ISBN 80-87378-73-3. 131 + 8 pp. LCCN QH31.M45 L48 2011.

Sacco:2016:BRH

[Sac16a] Francesco G. Sacco. Book review: Hjalmar Fors, The Limits of Matter: Chemistry, Mining, and Enlightenment. Chicago and London: The University of Chicago Press, 2015. Pp. 241. ISBN 978-0-266-19499-8. £28.00 (hardback). British Journal for the History of Science, 49(3): 487–488, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
B2867E4C0FBC88A1F17B46BB37621204.

Sacco:2016:BRM

[Sac16b] Francesco G. Sacco. Book reviews: Meredith K. Ray, Daughters of Alchemy: Women and Scientific Culture in Early Modern Italy. Cambridge, MA: Harvard University Press, 2015. Pp. 291. ISBN 978-0-6745-0423-

3. $45.00, £33.95 (hardback). British Journal for the History of Science, 49(1):122–123, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sacco:2019:BRN

[Sac19] Francesco Sacco. Book review: Nasser Zakariya, A Final Story: Science, Myth and Beginnings. Chicago and London, The University of Chicago Press, 2017. Pp. 554. ISBN 978-0-2264-7612-

4. $34.00 (cloth). British Journal for the History of Science, 52 (4):722–723, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/nasser-
zakariya-a-final-story-
science-myth-and-
beginnings-chicago-and-
london-the-university-of-
chicago-press-2017-pp-554-
isbn-9780226476124-3400-
cloth/
3E8A652F67BA84DF5DBA62CA4F461C8F.

Salvey:2014:BRR

[Sal14] Courtney Salvey. Book review:

Roger S. Wotton, Walking with Gosse: Natural History, Creation, and Religious Conflicts. Southampton: Clio Publishing, 2012. Pp. viii + 214. ISBN 978-0-9556983-9-2. £11.95 (paperback). British Journal for the History of Science, 47(1):187– 188, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Salvey:2015:BRB

[Sal15] Courtney J. Salvey. Book reviews: Ben Marsden, Hazel Hutchison and Ralph O’Connor (eds.), Uncommon Contexts: Encounters between Science and Literature, 1800–1914. London: Pickering and Chatto, 2013. Pp. xiii + 239. ISBN 978-1-84893-362-0. £60.00 (hardback). British Journal for the History of Science, 48(4): 700–702, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sanders:2010:DMP

[San10a] Robert H. Sanders. The dark matter problem: a historical perspective. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-511-77357-9 (ebook), 0-511-77615-2 (e-book), 0-521-11301-6 (hardcover). viii + 205 pp. LCCN QB791.3 .S25

2010. URL http://assets.
cambridge.org/97805211/
13014/cover/9780521113014.
jpg;   http://www.loc.gov/

                                        catdir/enhancements/fy1006/
                                        2010004532-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1006/
                                        2010004532-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1006/
                                        2010004532-t.html.

Santos:2010:AKC

[San10b] Catarina Madeira Santos. Administrative knowledge in a colonial context: Angola in the Eighteenth Century. British Journal for the History of Science, 43(4): 539–556, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Santos:2011:DMC

[San11] Gildo Magalhães Santos. A debate on magnetic current: the troubled Einstein–Ehrenhaft correspondence. British Journal for the History of Science, 44(3): 371–400, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2012:BRP

[San12] Neeraja Sankaran. Book reviews: Phillip R. Sloan and Brandon Fogel (eds.), Creating a Physical Biology: The Three-Man Paper and Early Molecular Biology. Chicago and London: The University of Chicago Press, 2011. Pp. ix + 319. ISBN 978-0-226-76783-3. £22.50 (paperback). British Journal for the History of Science, 45(4):694–695, 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2013:BRB

[San13] Neeraja Sankaran. Book reviews: Biology, Computing, and the History of Molecular Sequencing: From Proteins to DNA, 1945–2000. Basingstoke: Palgrave Macmillan, 2012. Pp. xiii + 242. ISBN 978-0-230-25032-1. £55.00 (hardback), by Miguel García-Sancho. British Journal for the History of Science, 46(3): 543–544, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2014:BRL

[San14] Neeraja Sankaran. Book reviews: Lawrence M. Principe, The Secrets of Alchemy. Chicago and London: The University of Chicago Press, 2013. Pp. v + 281. ISBN 978-0-226-68295-2. £16.00 (hardback). British Journal for the History of Science, 47(2):372– 374, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2015:BRD

[San15a] Neeraja Sankaran. Book reviews: David Cantor and Edmund Ramsden (eds.), Stress, Shock, and Adaptation in the Twentieth Century. Rochester, NY: University of Rochester Press, 2014. Pp. vii + 367. ISBN: 978-1-58046-476-5. $125.00 (hardback). British Journal for the History of Science, 48(4): 709–710, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2015:ERF

[San15b] Neeraja Sankaran. Essary review: Far from depleted. … Michael Ruse (ed.), The Cambridge Encyclopedia of Darwin and Evolutionary Thought. Cambridge: Cambridge University Press, 2013. Pp. xvii + 568. ISBN 978-0-521-19531-7. £124.99 (hardback). Peter Bowler, Darwin Deleted: Imagining a World without Darwin. Chicago and London: The University of Chicago Press, 2013. Pp. 318. ISBN 978-0-226-06867-1. £21.00 (hardback). British Journal for the History of Science, 48(1):171– 174, March 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2016:BRD

[San16] Neeraja Sankaran. Book reviews: Doogab Yi, The Recombinant University: Genetic Engineering and the Emergence of Stanford Biotechnology. Chicago: The University of Chicago Press, 2015. Pp. 304. ISBN: 978-0-2261-4383-5. $40.00 (cloth). British Journal for the History of Science, 49(2):318–320, June 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sankaran:2017:BRS

[San17] Neeraja Sankaran. Book review: Staffan Müller–Wille and Christina Brandt, Heredity Explored: Between Public Domain and Experimental Science, 1850–1930. Cambridge, MA: MIT Press, 2016. Pp. 472. ISBN: 978-0-262-03443-

2. £36.00 (cloth). British Journal for the History of Science, 50(2):360–362, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
163D4BE08CF46D2EF032B46665ED1F27.

Sarkar:2013:BRG

[Sar13] Jayita Sarkar. Book reviews:

Gabrielle Hecht (ed.), Entangled Geographies: Empire and Technopolitics in the Global Cold War. Cambridge, MA and London: MIT Press, 2011. Pp. ix + 337. ISBN 978-0-262-51578-8. £20.95 (paperback). British Journal for the History of Science, 46(2):354– 355, June 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sarkar:2014:BRG

[Sar14] Jayita Sarkar. Book reviews:

Gabrielle Hecht, Being Nuclear: Africans and the Global Uranium Trade. Cambridge, MA: MIT Press, 2012. Pp. xx + 451. ISBN 978-0-262-01726-1. £20.95 (hardback). British Journal for the History of Science, 47(2):388– 389, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sauer:2013:BRJ

[Sau13] Tilman Sauer. Book reviews:

Jeffrey A. Barrett and Peter Byrne (eds.), The Everett Interpretation of Quantum Mechanics:

Collected Works 1955–1980 with Commentary. Princeton and Oxford: Princeton University Press, 2012, Pp. xii + 389. ISBN 978-0-691-14507-5. £52.00 (hardback). British Journal for the History of Science, 46(4):731–732, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sauer:2015:BRM

[Sau15] Tilman Sauer. Book reviews: Massimiliano Badino and Jaume Navarro (eds.), Research and Pedagogy: A History of Quantum Physics through Its Textbooks. Berlin: Edition Open Access, 2013, Pp. 296. ISBN 978-3-8442-5871-4. EUR37.84 (hardback). British Journal for the History of Science, 48(3):530– 532, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Solovey:2012:CWS

[SC12] Mark Solovey and Hamilton Cravens, editors. Cold war social science: knowledge production, liberal democracy, and human nature. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2012. ISBN 0-230-34050-4 (hardcover). xv + 270 pp. LCCN H62.5.U5 C625 2012.

Schaffer:2008:RSB

[Sch08a] Gavin Schaffer. Racial science and British society, 1930–62. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2008. ISBN 0-230-00892-5. ix + 234 pp. LCCN DA125.A1 S33

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0829/2008016150-d.html;
http://www.loc.gov/catdir/
enhancements/fy0829/
2008016150-t.html.

Schmalzer:2008:PPM

[Sch08b] Sigrid Schmalzer. The people’s Peking man: popular science and human identity in Twentieth-Century China. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-73859-0 (cloth), 0-226-73860-4 (paperback). xix + 346 pp. LCCN GN284.7 .S36

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0829/2007046296-d.html;
http://www.loc.gov/catdir/
enhancements/fy0829/
2007046296-t.html.

Schwartz:2010:DDG

[Sch10] Joel S. Schwartz. Darwin’s disciple: George John Romanes, a life in letters, volume 5 of Lightning rod press series. American Philosophical Society, Philadelphia, PA, USA, 2010. ISBN 1-60618-920-4. xxiv + 806 pp. LCCN QH31.R7 A4 2010.

Schafer:2011:OAD

[Sch11] Daniel Schäfer. Old age and disease in early modern medicine, volume 4 of The body, gender and culture. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-020-8 (hardcover), 1-84893-021-6 (e-book). 287 pp. LCCN ???? Translated by and Patrick Baker.

Schweber:2012:NFM

[Sch12] S. S. (Silvan S.) Schweber. Nuclear forces: the making of the physicist Hans Bethe. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06587-5 (hardcover). viii + 579 + 12 pp. LCCN QC774.B4 S39 2012. £25.95.

SchmidtHorning:2013:CST

[Sch13] Susan Schmidt Horning. Chasing sound: technology, culture, and the art of studio recording from Edison to the LP. Studies in industry and society. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-1022-2 (hardcover), 1-4214-1023-0 (e-book). x + 292 pp. LCCN ML3790 .S346 2013.

Schilt:2015:BRJa

[Sch15a] Cornelis Schilt. Book reviews: Jed Z. Buchwald and Robert Fox (eds.), The Oxford Handbook of the History of Physics. Oxford: Oxford University Press, 2013. Pp. ix + 945. ISBN 978-0-19-969625-3. £95.00 (hardback). British Journal for the History of Science, 48(2):384–385, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Schilt:2015:BRJb

[Sch15b] Cornelis J. Schilt. Book reviews: Jamie C. Kassler, Seeking Truth: Roger North’s Notes on Newton and Correspondence with Samuel Clarke c.1704–1713. Farnham: Ashgate, 2014. Pp. xii + 374. ISBN 978-1-4094-4921-8. £80.00 (hardback). British Journal for the History of Science, 48 (4):691–693, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Schmidt:2015:SSC

[Sch15c] Ulf Schmidt. Secret science: a century of poison warfare and human experiments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-929979-X (hardcover). xxxi + 637 pp. LCCN RA648 .S355 2015.

Schilt:2016:BRD

[Sch16] Cornelis J. Schilt. Book reviews: David Beck (ed.), Knowing Nature in Early Modern Europe. London: Pickering & Chatto, 2015. Pp. 240. ISBN 978-1-8489-3518-1. £60.00 (hardback). British Journal for the History of Science, 49(1):123– 125, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Scheffler:2017:BRG

[Sch17a] Robin Wolfe Scheffler. Book review: Georgina M. Montgomery and Mark A. Largent, A Companion to the History of American Science. Chichester: Wiley Blackwell, 2016. Pp. xvii + 692. ISBN 978-1-4051-5625-7. £120.00 (hardback). British Journal for the History of Science, 50(2):342–343, June 2017.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
78D93B72F3C09B1DF2CAE22558B3278F.

Schilt:2017:BRE

[Sch17b] Cornelis J. Schilt. Book review: Elizabeth Yale, Sociable Knowledge: Natural History and the Nation in Early Modern Britain. Philadelphia: Penn Press, 2016. Pp. 360. ISBN 978-0-8122-4781-

7. £45.50 (hardback). British Journal for the History of Science, 50(1):145–146, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
5BD3CAB7831A8AEFE34F7A89BD7AE5B3.

Schilt:2019:BRA

[Sch19] Cornelis J. Schilt. Book review: Aviva Rothman, The Pursuit of Harmony: Kepler on Cosmos, Confession, and Community. Chicago and London: The University of Chicago Press, 2017. Pp. viii + 355. ISBN 978-0-226-49697-9. $55.00 (cloth cover). British Journal for the History of Science, 52(1):167–168, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/aviva-rothman-the-
pursuit-of-harmony-kepler-
on-cosmos-confession-and-
community-chicago-and-
london-the-university-of-

chicago-press-2017-pp-viii-
355-isbn-9780226496979-
5500-cloth-cover/1457941CE7055F02D086CF2CB5151421.

Scull:2009:HB

[Scu09] Andrew T. Scull. Hysteria: the biography. Biographies of disease. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956096-X (hardcover). vii + 223 pp. LCCN RC532 .S38 2009.

Wallace:2013:ARW

[SDS13] Charles H. (Charles Hyde) Smith, Megan Derr, and Michael Shermer, editors. Alfred Russel Wallace’s 1886–1887 Travel Diary: the North American lecture tour. Siri Scientific Press, Manchester, UK, 2013. ISBN 0-9567795-8-1. xiii + 258 pp. LCCN QH31.W2 S65x 2013.

Seavers:2010:BRL

[Sea10] Stephanie Seavers. Book review:

Leah DeVun, Prophecy, Alchemy, and the End of Time: John of Rupecissa in the Late Middle Ages. New York: Columbia University Press, 2009. Pp. xiii + 255. ISBN 978-0-231-14538-1. £34.50 (hardback). British Journal for the History of Science, 43(3):479–480, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Secord:2014:VSB

[Sec14] James A. Secord. Visions of science: books and readers at the dawn of the Victorian age.

Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-967526-0 (hardcover), 0-19-166275-5 (ebook). xiii + 306 + 8 pp. LCCN PR468.S34 S43 2014.

Seed:2012:SDN

[See12] David Seed. Seven days to noon: containing the atomic threat. British Journal for the History of Science, 45(4):641–652, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Selin:2008:EHS

[Sel08] Helaine Selin, editor. Encyclopaedia of the history of science, technology, and medicine in nonwestern cultures. Springer reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 1-4020-4425-9, 1-4020-4559-X (hardcover), 1-4020-4960-9 (print and electronic). xix + 2416 pp. LCCN Q124.8 .E53 2008. Two volumes.

Semi:2012:MSM

[Sem12] Maria Semi. Music as a science of mankind in Eighteenth-Century Britain. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-2868-0 (hardcover), 1-4094-2869-9 (e-book). vi + 185 pp. LCCN ML3797.2.G7 S45 2012.

Sepkoski:2012:RFR

[Sep12] David Sepkoski. Rereading the fossil record: the growth of paleobiology as an evolutionary discipline. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-74855-3 (hardcover). 432 pp. LCCN QE719.8 .S47 2012.

Serpente:2012:BRJ

[Ser12] Norberto Serpente. Book review: John Bender and Michael Marrinan, The Culture of Diagram. Stanford: Stanford University Press, 2010. Pp. xx + 264. ISBN 978-0-8047-4504-8. $60.00 (hardback). British Journal for the History of Science, 45(1): 133–135, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [BM10].

Serpente:2014:BRN

[Ser14a] Norberto Serpente. Book review: Nathaniel Comfort, The Science of Human Perfection: How Genes Became the Heart of American Medicine. New Haven and London: Yale University Press, 2012. Pp. xvii + 316. ISBN 978-0-300-16991-1. £25.00 (hardback). British Journal for the History of Science, 47(1):191– 192, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Serpente:2014:BRH

[Ser14b] Norberto Serpente. Book reviews: Heiner Fangerau, Rethy Chhem, Irmgard Müller and Shih-Chang Wang (eds.), Medical Imaging and Philosophy: Challenges, Reflections and Actions.

Stuttgart: Franz Steiner Verlag, 2012. Pp. 190. ISBN 978-3-515-10046-5. EUR38.00 (paperback). British Journal for the History of Science, 47(3):581–583, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Serpente:2015:BRD

[Ser15] Norberto Serpente. Book reviews: Diane B. Paul and Jeffrey P. Brosco, The PKU Paradox: A Short History of a Genetic Disease. Baltimore: Johns Hopkins University Press, 2013. Pp. xxiv + 289. ISBN 978-1-4214-1131-6. £16.00 (paperback). British Journal for the History of Science, 48(3):533–534, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Seth:2010:CQA

[Set10] Suman Seth. Crafting the quantum: Arnold Sommerfeld and the practice of theory, 1890–1926. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01373-8 (hardcover). vii + 378 pp. LCCN QC16.S76 2010.

Sloan:2011:CPB

[SF11] Phillip R. Sloan and D. Brandon Fogel, editors. Creating a physical biology: the Three-Man Paper and early molecular biology. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-76782-5 (hardcover), 0-226-76783-3 (paperback). ix + 319 pp. LCCN QH506 .C73 2011.

Smith:2012:ASD

[SFFC12] C. U. M. Smith, Eugenio Frixione, Stanley Finger, and William Clower. Animal spirit doctrine and the origins of neurophysiology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-976649-5 (hardcover). LCCN QP355.2 .A55 2012.

Shank:2010:BRK

[Sha10] J. B. Shank. Book review:

Kurt Ballstadt, Diderot: Natural Philosopher. Oxford: Voltaire Foundation, 2008. Pp. viii + 246. ISBN 978-0-7294-0948-3. £55.00 (paperback). British Journal for the History of Science, 43(2):297– 299, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shank:2011:BRC

[Sha11] Michael H. Shank. Book review:

Catherine Eagleton, Monks, Manuscripts and Sundials: The Navicula in Medieval England. Leiden: Brill, 2010. Pp. xi + 292. ISBN 978-90-04-27665-2. EUR99.00 (hardback). British Journal for the History of Science, 44(4):580–581, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shapin:2012:ITH

[Sha12] Steven Shapin. The Ivory Tower:

the history of a figure of speech and its cultural uses. British Journal for the History of Science, 45(1):1–27, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sheldon:2013:BRD

[She13] Myrna Perez Sheldon. Book reviews: David Sepkoski, Rereading the Fossil Record: The Growth of Paleobiology as an Evolutionary Discipline. Chicago and London: University of Chicago Press, 2012. Pp. viii + 432. ISBN 978-0-226-74855-9. £38.50 (hardback). British Journal for the History of Science, 46(4):732–734, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shiflett:2016:BRA

[Shi16] Stephanie Shiflett. Book review: Ayesha Ramachandran, The Worldmakers: Global Imagining in Early Modern Europe. Chicago: The University of Chicago Press, 2015. Pp. 299. ISBN: 978-0-226-28879-6. $45.00 (hardback). British Journal for the History of Science, 49(3): 478–479, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
0317D40A7064BC7B70D5B72028B374E9.

Shotwell:2014:BRM

[Sho14] R. Allen Shotwell. Book review: Manfred Horstmanshoff, Helen King and Claus Zittel (eds.), Blood, Sweat and Tears:

The Changing Concepts of Physiology from Antiquity into Early Modern Europe. Leiden: Brill, 2012. Pp. xxvi + 772. ISBN 978-90-04-22918-1. EUR217.00 (hardback). British Journal for the History of Science, 47(1):175– 176, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shotwell:2018:BRA

[Sho18] Allen Shotwell. Book review:

Anita Guerrini, The Courtiers’ Anatomists: Animals and Humans in Louis XIV’s Paris. Chicago: The University of Chicago Press, 2015. Pp. xiv + 343. ISBN: 978-0-2262-4766-3. $35.00 (hardback). British Journal for the History of Science, 51(3):524– 525, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/anita-
guerrini-the-courtiers-
anatomists-animals-and-
humans-in-louis-xivs-paris-
chicago-the-university-of-
chicago-press-2015-pp-xiv-
343-isbn-9780226247663-
3500-hardback/5FB06712829EF9F7C944D7734CF731E5.

Shuttleworth:2010:MCC

[Shu10] Sally Shuttleworth. The mind of the child: child development in literature, science, and medicine, 1840–1900. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010.

ISBN 0-19-958256-4. x + 497 pp. LCCN PR468.C5 S58

                                        2010. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1107/2010923442-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1107/
                                        2010923442-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1107/
                                        2010923442-t.html.

Sialaros:2018:BRV

[Sia18]  Michalis Sialaros.  Book re-

view: Volker R. Remmert, Martina R. Schneider and Henrik Kragh Sørensen (eds.), Historiography of Mathematics in the 19th and 20th Centuries. Cham: Birkhäuser and Springer, 2016. Pp. 267. ISBN 978-3-319-39647-

7. £89.50 (hardcover). British Journal for the History of Science, 51(1):165–167, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/volker-r-
                                        remmert-martina-r-
                                        schneider-and-henrik-kragh-
                                        sorensen-eds-
                                        historiography-of-
                                        mathematics-in-the-19th-
                                        and-20th-centuries-cham-
                                        birkhauser-and-springer-
                                        2016-pp-267-isbn-9783319396477-
                                        8950-hardcover/779A65D645BFF20C360DA7B7ADEF73A0.

Simunek:2010:MBM

[Sim10]  Michal Simunek.  Mendelism

in Bohemia and Moravia, 1900– 1930: collection of selected papers, volume 6 of Wissenschaftskultur um 1900. Steiner, Stuttgart, Germany, 2010. ISBN 3-515-09602-7. 276 pp. LCCN ????

Simon:2013:BRY

[Sim13] Josep Simon. Book reviews:

Yves Bouvier, Robert Fox, Pascal Griset and Anna Guagnini (eds.), De l’atelier au laboratoire: Recherche et innovation dans l’industrie électrique XIXe-XXe siècles / From Workshop to Laboratory: Research and Innovation in Electric Industry 19th-20th Centuries. Brussels: Peter Lang, 2011. Pp. 141. ISBN 978-90-5201-656-6. £22.40 (paperback). British Journal for the History of Science, 46(1):169– 170, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Simmons:2014:BRJ

[Sim14] Anna Simmons. Book reviews:

Jean-Paul Gaudillière and Volker Hess (eds.), Ways of Regulating Drugs in the 19th and 20th Centuries. Basingstoke: Palgrave Macmillan, 2013. Pp. xiv + 327. ISBN 978-0-230-30196-2. £60.00 (hardback). British Journal for the History of Science, 47(4): 745–746, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Simpson:2016:BRF

[Sim16] Thomas Simpson. Book review: Fraser MacDonald and Charles W.J. Withers (eds.), Geography, Technology and Instruments of Exploration. Farnham: Ashgate, 2015. Pp. 282. ISBN 978-1-4724-3425-8. £63.00 (hardback). British Journal for the History of Science, 49(4):494– 496, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
510CA9632483453F99765DEE2D0D11F8.

Simmons:2017:BRM

[Sim17a] Anna Simmons. Book review: Matthew J. Crawford, The Andean Wonder Drug: Cinchona Bark and Imperial Science in the Spanish Atlantic, 1630–1800. Pittsburgh, PA: University of Pittsburgh Press, 2016. Pp. xi + 284. ISBN 978-0-8229-4452-

2. $45.00 (hardback). British Journal for the History of Science, 50(1):150–151, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A940F68A6B81380B2172A67CDA404E57.

Simpson:2017:BRC

[Sim17b] Daniel Simpson. Book review: E. Charles Nelson, John Scouler (c. 1804–1871), Scottish Naturalist: A Life, with Two Voyages. Glasgow: The Glasgow Natural History Society, 2014. Pp. 142. ISBN 978-0-9565295-1-

0. £11.00 (paperback). British Journal for the History of Science, 50(1):153–154, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
5E163B84FBA402928584D24EA3FA3552.

Simmons:2019:TKN

[Sim19a] Anna Simmons. Trade, knowledge and networks: the activities of the Society of Apothecaries and its members in London, c. 1670–c. 1800. British Journal for the History of Science, 52(2):273–296, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/trade-
knowledge-and-networks-the-
activities-of-the-society-
of-apothecaries-and-its-
members-in-london-
c1670c1800/
F9D404B9191CCC4C26171FD96D0DF112.

Simon:2019:RHS

[Sim19b] Jonathan Simon. Retrospectives: History of science in France. British Journal for the History of Science, 52(4): 689–695, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
retrospectives-history-of-
science-in-france/
2084539258B7E125926E7E79960D48F2.

Sinha:2008:SWI

[Sin08] Jagdish N. Sinha. Science, war, and imperialism: India in the Second World War. Social sciences in Asia. Brill, Leiden, The Netherlands, 2008. ISBN 90-04-16645-9. xiv + 278 pp. LCCN Q175.52.I4 S56 2008. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0812/2008009918.html.

Singh:2015:BRR

[Sin15] Neil Singh. Book reviews: Ruha Benjamin, People’s Science: Bodies and Rights on the Stem Cell Frontier. Stanford, CA: Stanford University Press, 2013. Pp. xv + 249. ISBN 978-0-8047-8297-5. £17.99 (paperback). British Journal for the History of Science, 48(2):383–384, 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Siraisi:2013:CLE

[Sir13] Nancy G. Siraisi. Communities of learned experience: epistolary medicine in the Renaissance. Singleton Center books in premodern Europe. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0749-3 (hardcover), 1-4214-0784-1 (ebook). 163 pp. LCCN R484 .S57 2013.

Sivasundaram:2010:BRH

[Siv10] Sujit Sivasundaram. Book review: Helaine Selin (ed.), Encyclopaedia of the History of Science, Technology, and Medicine in Non-Western Cultures, Second Edition. 2 vols. Berlin, Heidelberg and New York: Springer, 2008. Pp. xix + 2416. ISBN 978-1-4020-4559-2. £355.00 (hardback).

British Journal for the History of Science, 43(2):284–285, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Szollosi-Janze:2011:BRS

[SJ11] Margit Szöllösi-Janze. Book review: Susanne Heim, Carola Sachse and Mark Walker (eds.), The Kaiser Wilhelm Society under National Socialism. Cambridge: Cambridge University Press, 2009. Pp. xxiv + 477. ISBN 978-0-521-87906-4. £45.00 (hardback). British Journal for the History of Science, 44(1): 151–152, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sastre-Juan:2019:BRR

[SJ19] Jaume Sastre-Juan. Book review: Rachel Plotnick, Power Button: A History of Pleasure, Panic, and the Politics of Pushing. Cambridge, MA: MIT Press, 2018. Pp. ix + 394. ISBN 978-0-2620-3823-2. $40.00/ £30.00 (cloth). British Journal for the History of Science, 52(3): 532–533, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/rachel-
plotnick-power-button-a-
history-of-pleasure-panic-
and-the-politics-of-
pushing-cambridge-ma-mit-
press-2018-pp-ix-394-isbn-

                                        9780262038232-40003000-
                                        cloth/
                                        341C9B6310BBA441FD24AB624FA00EAF.

Stewart:2012:OFB

[SK12] Alan Stewart and Harriet (Harriet Rachel) Knight, editors. The Oxford Francis Bacon. Vol. 1, Early writings, 1584–1596. The Oxford Francis Bacon. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-818313-5 (hardcover). ???? pp. LCCN PR.

Sleigh:2010:BRJ

[Sle10]  Charlotte Sleigh.  Book re-

view: J. F. M. Clark, Bugs and the Victorians. New Haven and London: Yale University Press, 2009. Pp. xiii + 322. ISBN 978-0-300-15091-9. £25.00 (hardback). British Journal for the History of Science, 43(2):310– 311, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sleigh:2011:BRS

[Sle11a] Charlotte Sleigh. Book review: Sally Shuttleworth, The Mind of the Child: Child Development in Literature, Science and Medicine 1840–1900. Oxford: Oxford University Press, 2010. Pp. xii + 497. ISBN 978-0-19-958256-3. £35.00 (hardback). British Journal for the History of Science, 44(4): 606–607, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sleigh:2011:LS

[Sle11b] Charlotte Sleigh. Literature and science. Outlining literature. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2011. ISBN 0-230-21816-4 (hardcover), 0-230-21817-2 (paperback). xi + 220 pp. LCCN PN55 .S59 2011.

Sleigh:2012:BRH

[Sle12] Charlotte Sleigh. Book review:

Harriet Ritvo, Noble Cows and Hybrid Zebras: Essays on Animals and History. Charlottesville and London: University of Virginia Press, 2010. Pp. x + 239. ISBN 978-0-8139-3060-2. $39.50 (hardback). British Journal for the History of Science, 45(3): 468–469, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sellers:2012:TTI

[SM12] Christopher Sellers and Joseph Melling. Towards a transnational industrial-hazard history: charting the circulation of workplace dangers, debates and expertise. British Journal for the History of Science, 45(3):401–424, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stirling-Middleton:2019:BRH

[SM19] Emma Stirling-Middleton. Book review: Hilary A. Smith, Forgotten Disease: Illness Transformed in Chinese Medicine. Stanford, CA:

Stanford University Press, 2017. Pp. 232. ISBN 978-1-5036-0344-

8. $24.95 (paperback). British Journal for the History of Science, 52(1):165–167, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/hilary-a-
smith-forgotten-disease-
illness-transformed-in-
chinese-medicine-stanford-
ca-stanford-university-
press-2017-pp-232-isbn-
9781503603448-2495-
paperback/4AFB0FBC0FE93F520846E2946D58B6CA.

Small:2013:BRS

[Sma13] Margaret Small. Book review: Zur Shalev, Sacred Words and Worlds: Geography, Religion and Scholarship, 1550–1700. Leiden and Boston: Brill, 2012. Pp. xxi + 319. 978-90-04-20935-0. EUR99.00 (hardback). British Journal for the History of Science, 46(1):163–164, March 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2010:AAH

[Smi10a] Melissa Smith. Architects of Armageddon: the Home Office Scientific Advisers’ Branch and civil defence in Britain, 1945– 1968. British Journal for the History of Science, 43(2):149–180, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2010:BRA

[Smi10b] Roger Smith. Book review: Anne Harrington, The Cure Within: A History of Mind-Body Medicine. New York and London: W. W. Norton & Company, 2008. Pp. 336. ISBN 978-0-393-06563-3. £19.99 (hardback). British Journal for the History of Science, 43(2):320–322, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2010:BRC

[Smi10c] Roger Smith. Book review: Cathy Gere, Knossos and the Prophets of Modernism. Chicago and London: University of Chicago Press, 2009. Pp. x + 277. ISBN 978-0-226-28953-3. £19.00 (hardback). British Journal for the History of Science, 43 (1):135–136, March 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2011:DML

[Smi11a] Justin E. H. Smith. Divine machines: Leibniz and the sciences of life. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14178-9 (hardcover). xii + 380 pp. LCCN ????

Smith:2011:BRR

[Smi11b] Melissa Smith. Book review:

Richard Moore, Nuclear Illusion, Nuclear Reality: Britain, the United States and Nuclear Weapons 1958–64. Basingstoke: Palgrave Macmillan, 2010. Pp. xvi + 332. ISBN 978-0-230-23067-5. £65.00 (hardback). British Journal for the History of Science, 44(2):309–311, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2011:BRJ

[Smi11c] Roger Smith. Book review: Joel S. Schwartz, Darwin’s Disciple: George John Romanes, a Life in Letters. Philadelphia: American Philosophical Society, 2010. Pp. xxiv + 806. ISBN 978-1-60618-920-7. $60.00 (paperback). British Journal for the History of Science, 44(4):607–608, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2011:BRD

[Smi11d] Stefan Halikowski Smith. Book review: Daniela Bleichmar, Paula De Vos, Kristin Huffine and Kevin Sheehan (eds.), Science in the Spanish and Portuguese Empires, 1500–1800. Stanford, CA: Stanford University Press, 2009. Pp. xxiii + 427. ISBN 978-0-8047-5358-6. $65.00 (hardback). British Journal for the History of Science, 44(2):285–286, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2012:BRA

[Smi12] Roger Smith. Book review: Alfred I. Tauber, Freud, the Reluctant Philosopher. Princeton and Oxford: Princeton University Press, 2010. Pp. xx + 318. ISBN 978-0-691-14552-5. $24.95 (paperback). British Journal for the History of Science, 45 (1):140–141, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Tau10].

Smith:2014:BRP

[Smi14] Robert W. Smith. Book reviews:

Peter J. Westwick (ed.), Blue Sky Metropolis: The Aerospace Century in Southern California. Berkeley: University of California Press, 2012. Pp. xii + 308. ISBN 978-0-87328-249-9. £30.95 (hardback). British Journal for the History of Science, 47(2):384– 385, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2015:APP

[Smi15] Matthew Smith. Another person’s poison: a history of food allergy. Arts and traditions of the table: perspectives on culinary history. Columbia University Press, New York, NY, USA, 2015. ISBN 0-231-16484-X (hardcover), 0-231-53919-3 (e-book). xii + 290 pp. LCCN RC596 .S39 2015.

Smith:2016:BRJ

[Smi16a] Robert W. Smith. Book reviews: Janet Vertesi, Seeing Like a Rover: How Robots, Teams, and Images Craft Knowledge of Mars. Chicago: The University of Chicago Press, 2015. Pp. 304. ISBN 978-0-226-15596-8. $35.00/£24.50 (hardback). British Journal for the History of Science, 49(2):322– 323, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2016:BRB

[Smi16b] Roger Smith. Book review: Bernard Lightman (ed.), Global Spencerism: The Communication and Appropriation of a British Evolutionist. Leiden: Brill, 2016. Pp. xiv + 310. ISBN 978-9-004-2639-4. EUR115.00 (paperback). British Journal for the History of Science, 49(4):499– 500, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
F890F3F5DDA42206890352E1B2CB7DD8.

Smith:2018:BRB

[Smi18a] Roger Smith. Book review: Brian J. McVeigh, The History of Japanese Psychology: Global Perspectives, 1875–1950. London and New York: Bloomsbury Academic, 2017. Pp. 319. ISBN 978-1-4742-8308-

3. £76.50 (hardback). British Journal for the History of Science, 51(1):169–170, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/brian-j-
mcveigh-the-history-of-
japanese-psychology-global-
perspectives-18751950-
london-and-new-york-
bloomsbury-academic-2017-
pp-319-isbn-9781474283083-
7650-hardback/1E5BE19B580600434EB57D628C1B2FAC.

Smith:2018:BRS

[Smi18b] Roger Smith. Book review:

Stephen T. Casper and Delia Gavrus, The History of the Brain and Mind Sciences. Rochester, NY: University of Rochester Press, 2017. Pp. 310. ISBN 978-1-5800-46595-3. £95.00 (hardcover). British Journal for the History of Science, 51(3): 531–532, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/stephen-t-
casper-and-delia-gavrus-
the-history-of-the-brain-
and-mind-sciences-
rochester-ny-university-of-
rochester-press-2017-pp-
310-isbn-97815800465953-
9500-hardcover/C83812698B6988697492946AE077AE73.

Smith:2019:BRJ

[Smi19] Robert W. Smith. Book review: Jared S. Buss, Willy Ley: Prophet of the Space Age. Gainesville: University of Florida Press, 2017. Pp. xiii + 321. ISBN 978-0-8130-5443-

8. $34.95 (hardcover). British Journal for the History of Science, 52(2):378–379, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jared-s-
buss-willy-ley-prophet-of-
the-space-age-gainesville-

                                        university-of-florida-
                                        press-2017-pp-xiii-321-
                                        isbn-9780813054438-3495-
                                        hardcover/
                                        9671BF08D9DEDCD2D0A25A876E5DFD77.

Smocovitis:2010:BRP

[Smo10] Vassiliki Betty Smocovitis. Book reviews: Peter Ayres, The Aliveness of Plants: The Darwins at the Dawn of Plant Science. London: Pickering & Chatto, 2008. Pp. xiii + 227. ISBN 978-1-85196-970-8. £60.00 (hardback). David Kohn, Darwin’s Garden: An Evolutionary Adventure. New York: New York Botanical Garden, 2008. Pp. 60. ISBN 978-0-89327-970-7. $17.99. British Journal for the History of Science, 43(2):306–308, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Snobelen:2010:BRM

[Sno10] Stephen D. Snobelen. Book review: Michael Hunter, Boyle: Between God and Science. New Haven and London: Yale University Press, 2009. Pp. xiii + 366. ISBN 978-0-300-12381-4. £25.00 (hardback). British Journal for the History of Science, 43(3): 485–486, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Soderqvist:2010:BRS

[Söd10] Thomas Söderqvist. Book reviews: Sherry Turkle (ed.), Evocative Objects: Things We Think With. Cambridge, MA and London: MIT Press, 2007. Pp. ix + 385. ISBN 978-0-262-20168-1. £19.95 (hardback). Sherry Turkle (ed.), Falling for Science: Objects in Mind. Cambridge, MA and London: MIT Press, 2008. Pp. xii + 318. ISBN 978-0-262-20172-8. £19.95 (hardback). Sherry Turkle (ed.), The Inner History of Devices. Cambridge, MA and London: MIT Press, 2008. Pp. x + 208. ISBN 978-0-262-20176-6. £19.95 (hardback). British Journal for the History of Science, 43(3):506–508, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Solovey:2013:SFP

[Sol13] Mark Solovey. Shaky foundations:

the politics–patronage–social science nexus in Cold War America. Studies in modern science, technology, and the environment. Rutgers University Press, New Brunswick, NJ, USA, 2013. ISBN 0-8135-5465-9 (hardcover), 0-8135-5466-7 (e-book). x + 253 pp. LCCN H62.5.U5 S65 2013.

Sommer:2011:BRH

[Som11] Andreas Sommer. Book review: Heather Wolffram, The Stepchildren of Science. Psychical Research and Parapsychology in Germany, c. 1870–1939. Amsterdam and New York: Ropodi, 2009. Pp. 342. ISBN 978-90-420-2728-2. £70.00 (hardback). British Journal for the History of Science, 44(1):144–146, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sommer:2015:BRP

[Som15] Andreas Sommer. Book reviews: Peter Lamont, Extraordinary Beliefs: A Historical Approach to a Psychological Problem. Cambridge: Cambridge University Press, 2013. Pp. 321. ISBN 978-1-107-68802-5. £19.99 (paperback). British Journal for the History of Science, 48(4):707–708, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sorensen:2017:BRD

[Sør17] Henrik Kragh Sørensen. Book review: David Aubin and Catherine Goldstein (eds.), The War of Guns and Mathematics: Mathematical Practice and Communities in France and Its Western Allies around World War I. Providence, RI: American Mathematical Society, 2014. Pp. xviii + 391. ISBN 978-1-4704-1469-6. $126.00 (hardback). British Journal for the History of Science, 50(3): 557–558, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        4271E6A9449BE0FD01A114E56D04DDDC.

Spary:2012:EEF

[Spa12] E. C. (Emma C.) Spary. Eating the Enlightenment: food and the sciences in Paris. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-76886-4 (hardcover), 0-226-76888-0 (e-book). xi + 366 pp. LCCN GT2853.F8 S737 2012.

Spinardi:2011:LSU

[Spi11] Graham Spinardi. The limits to ‘spin-off’: UK defence R&D and the development of gallium arsenide technology. British Journal for the History of Science, 44(4):??, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Spillmann:2012:MWP

[Spi12a] Lothar Spillmann, editor. Max Wertheimer: On perceived motion and figural organization. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01746-6 (hardcover). vi + 296 pp. LCCN BF203 .W4413 2012.

Spinardi:2012:LSU

[Spi12b] Graham Spinardi. The limits to ‘spin-off’: UK defence R&D and the development of gallium arsenide technology. British Journal for the History of Science, 45(1):97–121, March 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Spiegel:2015:JFT

[Spi15] Richard J. Spiegel. John Flamsteed and the turn of the screw: mechanical uncertainty, the skilful astronomer and the burden of seeing correctly at the Royal Observatory, Greenwich. British Journal for the History of Science, 48(1):17–51, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Salisbury:2010:NMC

[SS10]  Laura  Salis-

bury and Andrew Shail, editors. Neurology and modernity: a cultural history of nervous systems, 1800–1950. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-23313-9 (hardback). xiii + 298 pp. LCCN RC338 .N476 2010.

URL http://bvbr.bib-bvb.de:
8991/F?func=service\%26doc\
_library=BVB01\%26doc\_
number=018950403\%26line\_
number=0001\%26func\_code=
DB\_RECORDS\%26service\_
type=MEDI.

Savage-Smith:2011:BRA

[SS11] Emilie Savage-Smith. Book review: Ahmad Dallal, Islam, Science, and the Challenge of History. New Haven and London: Yale University Press, 2010. Pp. xii + 239. ISBN 978-0-300-15911-0. £18.99 (hardback). British Journal for the History of Science, 44(3):455–456, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sera-Shriar:2012:BRS

[SS12] Efram Sera-Shriar. Book reviews: Sadiah Qureshi, Peoples on Parade: Exhibitions, Empire, and Anthropology in Nineteenth-Century Britain. Chicago and London: The University of Chicago Press, 2011. Pp. vii + 382 ISBN 978-0-226-70096-0. £29.00 (hardback). British Journal for the History of Science, 45 (4):690–691, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sera-Shriar:2013:BRV

[SS13a] Efram Sera-Shriar. Book reviews: Visible Empire: Botanical Expeditions and Visible Culture in the Hispanic Enlightenment. Chicago and London: The University of Chicago Press, 2012. Pp. xii + 286. ISBN 978-0-226-05853-5. £33.50 (hardback), by Daniela Bleichmar. British Journal for the History of Science, 46(3):527–529, September 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sera-Shriar:2013:MBA

[SS13b] Efram Sera-Shriar. The Making of British Anthropology, 1813– 1871. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-394-0. xi + 255 pp. LCCN GN308.3.G7 S47 2013.

Sera-Shriar:2015:BRD

[SS15] Efram Sera-Shriar. Book reviews: Douglas A. Lorimer, Science, Race Relations and Resistance: Britain, 1870–1914. Manchester: Manchester University Press, 2013. Pp. xi + 344. ISBN 978-0-7190-3357-5. £80.00 (hardback). British Journal for the History of Science, 48(2): 373–374, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sera-Shriar:2016:BRT

[SS16] Efram Sera-Shriar. Book reviews: Timothy Larsen, The Slain God: Anthropologists and the Christian Faith. Oxford: Oxford University Press, 2014. Pp. 256. ISBN 978-0-1996-5787-2. £25.00 (hardback). British Journal for the History of Science, 49(1):129– 130, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sera-Shriar:2017:BRJ

[SS17] Efram Sera-Shriar. Book review: Jonathan Lamb, Scurvy: The Disease of Discovery. Princeton and Oxford: Princeton University Press, 2017. Pp. 328. ISBN 978-0-691-14782-6. £27.95 (hardback). British Journal for the History of Science, 50(4): 732–734, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        964E56C0EA290E1B718EAA13C3DF8038.

Stalnaker:2010:UED

[Sta10] Joanna Stalnaker. The unfinished Enlightenment: description in the age of the encyclopedia. Cornell University Press, Ithaca, NY, USA, 2010. ISBN 0-8014-6234-7 (e-book), 0-8014-4864-6 (hardcover). xvi + 240 pp. LCCN PQ265 .S72 2010.

Stadler:2011:BRH

[Sta11a] Max Stadler. Book review: Heiner Fangerau, Spinning the Scientific Web: Jacques Loeb (1859–1924) und sein Programm einer internationalen biomedizinischen Grundlagenforschung. Berlin: Akademie-Verlag, 2010. Pp. 280. ISBN 978-3-05-004528-3. EUR59.80 (hardback). British Journal for the History of Science, 44(1):141–143, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stahnisch:2011:PMJ

[Sta11b] Frank W. Stahnisch. Professor Margaret J. Osler (27 November 1942–15 September 2010). British Journal for the History of Science, 44(1):115–118, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stanley:2011:BRK

[Sta11c] Matthew Stanley. Book reviews:

Kristian Camilleri, Heisenberg and the Interpretation of Quantum Mechanics: The Physicist as Philosopher. Cambridge: Cambridge University Press, 2009. Pp. xii + 199. ISBN 978-0-521-88484-6. £45.00 (hardback). Cathryn Carson, Heisenberg in the Atomic Age: Science and the Public Sphere. Cambridge: Cambridge University Press, 2010. Pp. xxiv + 541. ISBN 978-0-521-82170-4. £55.00 (hardback). British Journal for the History of Science, 44 (2):308–309, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stark:2011:BRS

[Sta11d] James F. Stark. Book review: Susan D. Jones, Death in a Small Package: A Short History of Anthrax. Baltimore: Johns Hopkins University Press, 2010. Pp. xix + 329. ISBN 978-0-8018-9696-5. $24.95 (hardback). British Journal for the History of Science, 44 (4):615–617, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stanley:2012:DJC

[Sta12a] Matthew Stanley. By design: James Clerk Maxwell and the evangelical unification of science. British Journal for the History of Science, 45(1):57–73, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stark:2012:BRM

[Sta12b] James F. Stark. Book review: Michael Brown, Performing Medicine: Medical Culture and Identity in Provincial England, c. 1760–1850. Manchester: University of Manchester Press, 2011. Pp. ix + 254. ISBN 978-0-7190-7797-5. £60.00 (hardback). British Journal for the History of Science, 45(3):463–464, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stark:2014:BRP

[Sta14]  James F. Stark.  Book re-

views: Pratik Chakrabarti, Bacteriology in British India: Laboratory Medicine in the Tropics. Rochester, NY: University of Rochester Press. Pp. x + 304. ISBN 978-1-58046-408-6. £60.00 (hardback). British Journal for the History of Science, 47(2):381– 382, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stark:2015:BRW

[Sta15] James F. Stark. Book reviews:

William J. Turkel, Spark from the Deep: How Shocking Experiments with Strongly Electric Fish Powered Scientific Discovery. Baltimore: Johns Hopkins University Press, 2013. Pp. xi + 287. ISBN 978-1-4214-0981-8. £22.50 (hardback). British Journal for the History of Science, 48(2): 364–365, 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stark:2016:BRR

[Sta16a] James F. Stark. Book reviews:

Robert Brain, The Pulse of Modernism: Physiological Aesthetics in Fin-de–Siècle Europe. Seattle: University of Washington Press, 2015. Pp. xxxii + 348. ISBN 978-0-295-99320-1. $50.00 (hardback). British Journal for the History of Science, 49(2):309– 311, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stark:2016:IPP

[Sta16b] James F. Stark. Introduction:

plurality in patenting: medical technology and cultures of protection. British Journal for the History of Science, 49(4): 533–540, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
9FE353637B674BA579D01A8C70D28D98.

Stack:2019:CDS

[Sta19a] David Stack. Charles Darwin and the scientific mind. British Journal for the History of Science, 52(1):85–115, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/charles-
darwin-and-the-scientific-
mind/
2A8C0A8F984753BC98C3D549D48717F1.

Stan:2019:RHF

[Sta19b] Marius Stan. Response to H. Floris Cohen’s essay review on Newtonian scholarship. British Journal for the History of Science, 52(2):359–360, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/response-
to-h-floris-cohens-essay-
review-on-newtonian-
scholarship/
D295539A32A6E423CA07500A854A8A54.

Starkowski:2019:BRG

[Sta19c] Kristen Starkowski. Book review: Graeme Gooday and Karen Sayer, Managing the Experience of Hearing Loss in Britain, 1830– 1930. London: Palgrave Macmillan, 2017. Pp. xx + 126. ISBN 978-1-1374-0687-3. £44.99 (hardcover). British Journal for the History of Science, 52(3):531– 532, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/graeme-
gooday-and-karen-sayer-
managing-the-experience-of-
hearing-loss-in-britain-
18301930-london-palgrave-
macmillan-2017-pp-xx-126-
isbn-9781137406873-4499-
hardcover/
6527D10A1B05ADE0D2CC86DA3DBF5BB1.

Stegenga:2010:BRF

[Ste10] Jacob Stegenga. Book review:

Francesca Bordogna, William James at the Boundaries: Philosophy, Science, and the Geography of Knowledge. Chicago and London: University of Chicago Press. 2008. Pp. x + 382. ISBN 978-0-226-06652-3. £23.00 (hardback). British Journal for the History of Science, 43(1):130– 131, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Steel:2011:NLR

[Ste11a] Emily Steel. He is no loss: Robert McCormick and the voyage of HMS Beagle, volume 14 of BSHS monographs. British Society for the History of Science, London, UK, 2011. ISBN 0-906450-18-7 (paperback). x + 63 pp. LCCN ????

Stegenga:2011:BRH

[Ste11b] Jacob Stegenga. Book review: Henk W. de Regt, Sabina Leonelli and Kai Eigner (eds.), Scientific Understanding: Philosophical Perspectives. Pittsburgh: University of Pittsburgh Press, 2009. Pp. ix + 352. ISBN 978-0-8229-4378-6. $65.00 (hardback). British Journal for the History of Science, 44 (4):578–580, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stevens:2013:LSD

[Ste13] Hallam Stevens. Life out of sequence: a data-driven history of bioinformatics. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-08017-X (hardcover), 0-226-08020-X (paperback), 0-226-08034-X (e-book). 294 pp. LCCN QH324.2 .S726 2013.

Steen:2014:ASO

[Ste14a] Kathryn Steen. The American synthetic organic chemicals industry: war and politics, 1910– 1930. The University of North Carolina Press, Chapel Hill, NC, USA, 2014. ISBN 1-4696-1290-9 (paperback), 1-4696-1291-7 (ebook). xii + 403 pp. LCCN HD9651.5 .S73 2014.

Stewart:2014:BRC

[Ste14b] Larry Stewart. Book reviews: E. C. Spary, Eating the Enlightenment: Food and the Sciences in Paris, 1670–1760. Chicago and London: The University of Chicago Press, 2012. Pp. xi + 366. ISBN 978-0-226-76886-1. £29.00 (hardback). British Journal for the History of Science, 47 (4):731–732, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stewart:2016:BRR

[Ste16]  John Stewart.  Book

review: Roberta Bivins, Contagious Communities: Medicine, Migration and the NHS in Postwar Britain. Oxford: Oxford University Press, 2015. Pp. 448. ISBN 978-0-19-872528-2. £35.00 (hardback). British Journal for the History of Science, 49(4): 507–508, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
CEC33F005B852613EC6ED6F90219E104.

Stolzenberg:2012:BRJ

[Sto12] Daniel Stolzenberg. Book review: John Edward Fletcher, A Study of the Life and Works of Athanasius Kircher, ‘Germanus Incredibilis’. With a Selection of His Unpublished Correspondence and an Annotated Translation of His Autobiography. Leiden and Boston: Brill, 2011. Pp. xxxiv + 607. ISBN 978-90-04-20712-7. EUR184.00 (hardback). British Journal for the History of Science, 45(4):683– 685, December 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/product/
                                        13E460C33BC9BC1BE7C10F8D29FB71E2.

Stolzenberg:2013:EOA

[Sto13] Daniel Stolzenberg. Egyptian Oedipus: Athanasius Kircher and the secrets of antiquity. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92414-9 (hardcover), 0-226-92415-7 (ebook). ???? pp. LCCN D247 .S76 2013.

Stoddart:2014:BRJ

[Sto14] Kristan Stoddart. Book reviews: John R. Walker, Britain and Disarmament: The UK and Nuclear, Biological and Chemical Weapons Arms Control and Programmes 1956–1975. Farnham: Ashgate, 2012. Pp. xv + 305. ISBN 978-1-4094-3580-8. £70.00 (hardback). British Journal for the History of Science, 47(4):756–757, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Strick:2011:BRC

[Str11] James E. Strick. Book review: Christoph Gradmann, Laboratory Disease: Robert Koch’s Medical Bacteriology, translated by Elborg Forster. Baltimore: Johns Hopkins University Press, 2009. Pp. 318. ISBN 978-0-8018-9313-1. £18.00 (hardback). British Journal for the History of Science, 44 (1):140–141, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Strick:2019:MOS

[Str19] James E. Strick. Metaphors and other slippery creatures. British Journal for the History of Science, 52(2):345–352, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/metaphors-
and-other-slippery-
creatures/
AA376EAD2053F8A3F64BBE621F265C49.

Sturdy:2010:BRM

[Stu10] Steve Sturdy. Book review:

Michael Lynch, Simon A. Cole, Ruth McNally and Kathleen Jordan, Truth Machine: The Contentious History of DNA Fingerprinting. Chicago and London: University of Chicago Press, 2009. Pp. xxii + 389. ISBN 978-0-226-49806-5. £22.00 (hardback). British Journal for the History of Science, 43(2):319– 320, June 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sturdy:2011:BRW

[Stu11] Steve Sturdy. Book reviews:

A. W. Bates, The Anatomy of Robert Knox: Murder, Mad Science and Medical Regulation in Nineteenth-Century Edinburgh. Brighton: Sussex Academic Press, 2010. Pp. ix + 228. ISBN 978-1-84519-381-2. £39.95 (hardback). Lisa Rosner, The Anatomy Murders: Being the True and Spectacular History of Edinburgh’s Notorious Burke and Hare and of the Man of Science who Abetted Them in the Commission of Their Most Heinous Crimes. Philadelphia: University of Pennsylvania Press, 2010. Pp. vi + 328. ISBN 978-0-8122-4191-

4. £19.50 (hardback). British Journal for the History of Science, 44(1):133–134, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sturdy:2017:BRR

[Stu17] Steve Sturdy. Book review: Richard Tutton, Genomics and the Reimagining of Personalized Medicine. Farnham: Ashgate, 2014. Pp. 214. ISBN: 978-1-4724-2257-6. £60.00 (hardback). British Journal for the History of Science, 50(2):375–376, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
8E59EFD7631EEAE3DC5007847D58DCE6.

Suman:2011:BRA

[Sum11] Seth Suman. Book review: Alan Chalmers, The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms. Heidelberg, London and New York: Springer, 2009. Pp. xii + 287. ISBN 978-90-481-2361-2. £90.00 (hardback). British Journal for the History of Science, 44(3): 463–465, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sumner:2013:BST

[Sum13] James Sumner. Brewing science, technology and print, 1700–1880, volume 19 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-423-8 (hardcover), 1-78144-048-4 (ebook). xviii + 295 pp. LCCN TP573.A1 S86 2013.

Sumira:2014:AHG

[Sum14a] Sylvia Sumira. The art and history of globes. The British Library, London, UK, 2014. ISBN 0-7123-5868-4. 224 pp. LCCN GA260 .S86 2014.

Sumira:2014:GYE

[Sum14b] Sylvia Sumira. Globes: 400 years of exploration, navigation, and power. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-13900-X, 0-226-13914-X (e-book). 224 (est.) pp. LCCN G3170 .S955

2014. URL http://alltitles.
ebrary.com/Doc?id=10901486.

Sunderland:2013:TNH

[Sun13] Mary E. Sunderland. Teaching natural history at the Museum of Vertebrate Zoology. British Journal for the History of Science, 46(1):97–121, March 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stamhuis:2017:DBG

[SV17] Ida H. Stamhuis and Annette B.

Vogt. Discipline building in Germany: women and genetics at the Berlin Institute for Heredity Research. British Journal for the History of Science, 50 (2):267–295, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
6002D51F9883FA6A613336403BB4BA73.

Svatek:2018:ECP

[Sva18] Petra Svatek. Ethnic cartography and politics in Vienna, 1918–1945. British Journal for the History of Science, 51(1):99– 121, 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/ethnic-cartography-
and-politics-in-vienna-
19181945/
AC995B21EFEAC63D7DF4F3BF40D209F8.

Swenson:2015:MCD

[Swe15] Sarah A. Swenson. ‘Morals can not be drawn from facts but guidance may be’: the early life of W. D. Hamilton’s theory of inclusive fitness. British Journal for the History of Science, 48(4): 543–563, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sysling:2018:SSA

[Sys18] Fenneke Sysling. Science and selfassessment: phrenological charts 1840–1940. British Journal for the History of Science, 51(2): 261–280, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/science-
and-selfassessment-
phrenological-charts-
18401940/
8224EA2B31FB3EA67BEA89AE5B63EAF5.

Szuhan:2018:SLF

[Szu18] Natasha Szuhan. Sex in the laboratory: the Family Planning Association and contraceptive science in Britain, 1929-1959. British Journal for the History of Science, 51(3):487– 510, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/sex-in-the-
laboratory-the-family-
planning-association-and-
contraceptive-science-in-
britain-19291959/
82CB447E28496B7A2D9EF148F3B4679B.

Tampakis:2014:BRS

[Tam14a] Kostas Tampakis. Book reviews:

Sanja Perovic, The Calendar in Revolutionary France: Perceptions of Time in Literature, Culture, Politics. Cambridge: Cambridge University Press, 2012. Pp. xiv + 276. ISBN 978-1-107-02595-0. £55.00 (hardback). British Journal for the History of Science, 47(3):574–575, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tampakis:2014:OFB

[Tam14b] Kostas Tampakis. Onwards facing backwards: the rhetoric of science in nineteenth-century Greece. British Journal for the History of Science, 47(2):217– 237, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tampakis:2016:BRT

[Tam16a] Kostas Tampakis. Book review: Theodore Arabatzis, Jürgen Renn and Ana Simões (eds.), Relocating the History of Science: Essays in Honor of Kostas Gavroglu. Heidelberg: Springer, 2015. Pp. vii + 383. ISBN 978-3-319-14552-5. £93.50 (hardback). British Journal for the History of Science, 49(4): 655–657, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        23CF8CE069A1690A6D7915D764744966.

Tampakis:2016:BRM

[Tam16b] Kostas Tampakis. Book reviews: M. Alper Yalçinkaya, Learned Patriots: Debating Science, State, and Society in the Nineteenth-Century Ottoman Empire. Chicago: The University of Chicago Press, 2015. Pp. 304. ISBN 978-0-226-18420-3. $50.00 (hardback). British Journal for the History of Science, 49(2):304– 306, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tarrant:2012:BRE

[Tar12]  Neil Tarrant.  Book review:

Elaine Leong and Alisha Rankin (eds.), Secrets and Knowledge in Medicine and Science 1500–1800. Farnham: Ashgate, 2011. Pp. ix + 247. ISBN 978-0-7546-6854-1. £60.00 (hardback). British Journal for the History of Science, 45 (3):454–455, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tarrant:2013:GDP

[Tar13] Neil Tarrant. Giambattista Della Porta and the Roman Inquisition: censorship and the definition of Nature’s limits in sixteenthcentury Italy. British Journal for the History of Science, 46(4): 601–625, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tarrant:2015:BRD

[Tar15] Neil Tarrant. Book reviews:

Dario Tessicini and Patrick J. Boner (eds.), Celestial Novelties on the Eve of the Scientific Revolution, 1540–1630. Florence: Leo S. Olschki Editore, 2013. Pp. xvi + 282. ISBN 978-88-222-6254-7. EUR32.00 (paperback). British Journal for the History of Science, 48(4):690–691, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tattersall:2009:DB

[Tat09] Robert Tattersall. Diabetes: the biography. Biographies of disease.

Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954136-1 (hardcover). x + 229 pp. LCCN RC660 .T38 2009.

Tattersall:2016:BRD

[Tat16] Mason Tattersall. Book reviews: Dennis Danielson, Paradise Lost and the Cosmological Revolution. New York: Cambridge University Press, 2014. Pp. xxiii + 220. ISBN 978-1-1070-3360-3. £60.00 (hardback). British Journal for the History of Science, 49(1):120– 122, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tatarchenko:2017:BRB

[Tat17a] Ksenia Tatarchenko. Book review: Benjamin Peters, How Not to Network a Nation: The Uneasy History of the Soviet Internet. Cambridge, MA: MIT Press, 2016. Pp. xiii + 298. ISBN 978-0-262-03418-0. £28.95 (hardback). British Journal for the History of Science, 50(1):170–171, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
DE47EA657BEEEBB20D2E2C011BEDF9C5.

Tatarchenko:2017:BRM

[Tat17b] Ksenia Tatarchenko. Book review: Marie Hicks, Programmed Inequality: How Britain Discarded Women Technologists and Lost Its Edge in Computing. Cambridge, MA: MIT Press, 2017. Pp. 342. ISBN 978-0-262-03554-5. £32.95. British Journal for the History of Science, 50(2):368–369, June 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
2241841D17127DBAC78E32B5C1C78EA4.

Tattersall:2018:BRO

[Tat18]  Mason Tattersall.  Book re-

view: Olivier Darrigol, Atoms, Mechanics, and Probability: Ludwig Boltzmann’s Statistico-Mechanical Writings — An Exegesis. Oxford: Oxford University Press, 2018. Pp. xxvi + 612. ISBN 978-0-19-881617-1. £45.00 (hardcover). British Journal for the History of Science, 51(4): 713–714, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/olivier-
darrigol-atoms-mechanics-
and-probability-ludwig-
boltzmanns-
statisticomechanical-
writings-an-exegesis-
oxford-oxford-university-
press-2018-pp-xxvi-612-isbn-
9780198816171-4500-hardcover/
7366DBBC4394FB4BA499D89847A495AA.

Tauber:2010:FRP

[Tau10] Alfred I. Tauber. Freud, the reluctant philosopher. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14551-2 (hardcover), 0-691-14552-0 (paperback). xx + 318 pp. LCCN BF109.F74 T38 2010.

Taub:2011:BRD

[Tau11] Liba Taub. Book review: David Creese, The Monochord in Ancient Greek Harmonic Science. Cambridge: Cambridge University Press, 2010. Pp. xvi + 409. ISBN 978-0-521-84324-9. £65.00 (hardback). British Journal for the History of Science, 44(2): 282–283, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Taylor:2014:BRJ

[Tay14] Georgette Taylor. Book reviews: James O’Brien, The Scientific Sherlock Holmes: Cracking the Case with Science and Forensics. Oxford: Oxford University Press, 2013. Pp. xx + 175. ISBN 978-0-19-979496-6. £18.99 (hardback). British Journal for the History of Science, 47(3):577–579, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tessicini:2013:CNE

[TB13] Dario Tessicini and Patrick Boner, editors. Celestial novelties on the eve of the scientific revolution: 1540–1630, volume 3 of Biblioteca di Galiliana. L. S. Olschki, Firenze, Italia, 2013. ISBN 88-222-6254-9. xvi + 282 pp. LCCN QB29 .C45 2013. 32.00 EUR.

Tee:2011:BRP

[Tee11] Garry J. Tee. Book review: Paul Glennie and Nigel Thrift, Shaping the Day: A History of Timekeeping in England and Wales 1300– 1800. Oxford: Oxford University Press, 2009. Pp. xiv + 456. ISBN 978-0-19-927820-6. £41.00 (hardback). British Journal for the History of Science, 44(3): 458–459, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Terrall:2010:FIA

[Ter10] Mary Terrall. Following insects around: tools and techniques of Eighteenth-Century natural history. British Journal for the History of Science, 43(4):573–588, December 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Terzian:2013:SEC

[Ter13] Sevan G. Terzian. Science education and citizenship: fairs, clubs and talent searches for American youth, 1918–1958. Historical studies in education. Palgrave Macmillan, New York, NY, USA, 2013. ISBN 1-137-03186-7 (hardcover). xv + 235 pp. LCCN Q182.4 .T47 2013.

Terrall:2014:CNA

[Ter14] Mary Terrall. Catching nature in the act: Réaumur and the practice of natural history in the eighteenth century. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-08860-X (hardcover), 0-226-08874-X (e-book). 275 pp. LCCN QH15 .T47 2014.

Tessicini:2011:BRM

[Tes11] Dario Tessicini. Book review:

María Portuondo, Secret Science:

Spanish Cosmography and the New World. Chicago and London: University of Chicago Press, 2009. Pp. xiv + 335. ISBN 978-0-226-67534-3. £31.00 (hardback). British Journal for the History of Science, 44(3):462–463, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turchetti:2012:IWE

[THB12] Simone Turchetti, Néstor Herran, and Soraya Boudia. Introduction: have we ever been ‘transnational’? Towards a history of science across and beyond borders. British Journal for the History of Science, 45(3):319–336, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Thomas:2016:BRJ

[Tho16a] William Thomas. Book reviews: Joris Vandendriessche, Evert Peeters and Kaat Wils (eds.), Scientists’ Expertise as Performance: Between State and Society, 1860–1960. London: Pickering & Chatto, 2015. Pp. 256. ISBN 978-1-8489-3527-3. £60.00 (hardback). British Journal for the History of Science, 49(2):303– 304, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Thomson:2016:BRE

[Tho16b] Mathew Thomson. Book review: Erik Linstrum, Ruling Minds: Psychology in the British Empire. Cambridge, MA: Harvard University Press, 2016. Pp. 309. ISBN 978-0-674-08866-5. £39.95 (hardback). British Journal for the History of Science, 49(4): 651–652, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
E221C13BBC84CC91C8E759503DFCDED5.

Thomas:2019:BRY

[Tho19] Renny Thomas. Book review:

Yves Gingras, Science and Religion: An Impossible Dialogue. Cambridge: Polity Books, 2017. Pp. 249. ISBN 978-1-5095-1892-

0. £17.99 (paperback). British Journal for the History of Science, 52(2):363–364, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/yves-
gingras-science-and-
religion-an-impossible-
dialogue-cambridge-polity-
books-2017-pp-249-isbn-
9781509518920-1799-
paperback/
1FBC634B12EE2611581A9C5706A10011.

Tiles:2010:BRC

[Til10]  Mary Tiles.  Book review:

Christina Chimisso, Writing the History of the Mind: Philosophy and Science in France, 1900 to 1960s. Aldershot: Ashgate, 2008. Pp. ix + 209. ISBN 978-0-7546-5705-7. £55.00 (hardback). British Journal for the History of Science, 43(1):134– 135, March 2010. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Timmermann:2012:BRP

[Tim12] Anke Timmermann. Book review: Peter J. Grund, ‘Misticall Wordes and Names Infinite’: An Edition and Study of Humfrey Lock’s Treatise on Alchemy. Tempe: Arizona Center for Medieval and Renaissance Studies, 2011. Pp. xii + 350. ISBN 978-0-86698-415-7. $72.00 (hardback). British Journal for the History of Science, 45(3):456–457, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Timmermann:2013:HLC

[Tim13] Carsten Timmermann. A history of lung cancer: the recalcitrant disease. Science, technology and medicine in modern history. Palgrave Macmillan, Houndsmills, Basingstoke, Hampshire, 2013. ISBN 1-4039-8802-1 (hardcover). x + 244 pp. LCCN RC280.L8 T56 2013.

                                        URL http://www.netread.com/
                                        jcusers2/bk1388/027/
                                        9781403988027/image/
                                        lgcover.9781403988027.jpg.

Timmermann:2012:CPC

[TlT12] Carsten Timmermann and Elizabeth (University lecturer) Toon, editors. Cancer patients, cancer pathways: historical and sociological perspectives. Science, technology, and medicine in modern history. Palgrave Macmillan, Houndsmill, Basingstoke, Hampshire, 2012. ISBN 1-137-27207-4 (hardcover). xii + 270 pp. LCCN RC261 .C26 2012; RC261 .C2737

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1403/2012472625-b.html;
http://www.loc.gov/catdir/
enhancements/fy1403/
2012472625-d.html;
http://www.loc.gov/catdir/
enhancements/fy1403/
2012472625-t.html.

Toal:2011:PBA

[Toa11] Ciaran Toal. Preaching at the British Association for the Advancement of Science: sermons, secularization and the rhetoric of conflict in the 1870s. British Journal for the History of Science, 44(4):??, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Toal:2012:PBA

[Toa12] Ciaran Toal. Preaching at the British Association for the Advancement of Science: sermons, secularization and the rhetoric of conflict in the 1870s. British Journal for the History of Science, 45(1):75–95, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tomory:2012:FNI

[Tom12a] Leslie Tomory. Fostering a new industry in the Industrial Revolution: Boulton & Watt and gaslight 1800–1812. British Journal for the History of Science, 46(2):199–229, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tomory:2012:PEO

[Tom12b] Leslie Tomory. Progressive enlightenment: the origins of the gaslight industry, 1780–1820. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01675-3 (hardcover). x + 348 pp. LCCN TP733.G6 T66 2012.

                                        URL http://site.ebrary.com/
                                        id/10534379.

Tomory:2013:FNI

[Tom13a] Leslie Tomory. Fostering a new industry in the Industrial Revolution: Boulton & Watt and gaslight 1800–1812. British Journal for the History of Science, 46(2):199–229, June 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Toms:2013:MHP

[Tom13b] Jonathan Toms. Mental hygiene and psychiatry in modern Britain. Science, technology and medicine in modern history. Palgrave Macmillan, Houndmills, Basingstoke, 2013. ISBN 1-137-32156-3 (hardcover), 1-137-32001-X (e-book). xi + 262 pp. LCCN RA790.7.G7 T66

                                        2013. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1404/2013427780-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1404/
                                        2013427780-d.html;

http://www.loc.gov/catdir/
enhancements/fy1404/
2013427780-t.html.

Tosh:2011:BRS

[Tos11] Nick Tosh. Book review: Steve Fuller, Science. Durham: Acumen, 2010. Pp. vi + 170. ISBN 978-1-84465-204-4. £9.99 (paperback). British Journal for the History of Science, 44(4):577– 578, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Totelin:2010:BRC

[Tot10]  Laurence Totelin.  Book re-

view: Christopher Gill, Tim Whitmarsh and John Wilkins (eds.), Galen and the World of Knowledge. Cambridge: Cambridge University Press, 2009. Pp. xv + 327. ISBN 978-0-521-76751-4. £60.00 (hardback). British Journal for the History of Science, 43(3):478–479, September 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turchetti:2014:SIG

[TR14] Simone Turchetti and Peder Roberts, editors. The surveillance imperative: geosciences during the Cold War and beyond. Palgrave studies in the history of science and technology. Palgrave Macmillan, New York, NY, 2014. ISBN 1-137-43872-X (hardcover). xii + 278 pp. LCCN QH541.15.M64 S87 2014.

Tradii:2016:BRS

[Tra16] Laura Tradii. Book review: Sofie Lachapelle, Conjuring Science: A History of Scientific Entertainment and Stage Magic in Modern France. New York: Palgrave Macmillan, 2015. Pp. 216. ISBN 978-1-137-49768-0. £60.00 (hardback). British Journal for the History of Science, 49(3): 491–492, September 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
cambridge.org/core/product/
177C41FEB5DD263275E88B07784E4C72.

Thomasen:2014:BRR

[TS14] Laura Søvsø Thomasen and Henrik Kragh Sørensen. Book reviews: Rüdiger Campe, The Game of Probability: Literature and Calculation from Pascal to Kleist. Stanford: Stanford University Press, 2012. Pp. viii + 486. ISBN 978-0-8047-6865-8. $35.00 (paperback). British Journal for the History of Science, 47 (4):727–728, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tun:2019:BRP

[Tun19] Molly Tun. Book review: Pablo Calvi, Latin American Adventures in Literary Journalism. Pittsburgh: University of Pittsburgh Press, 2019. Pp. ix + 276. ISBN 978-0-8229-4565-9. $45.00 (hardback). British Journal for the History of Science, 52(4): 723–725, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/pablo-
calvi-latin-american-
adventures-in-literary-
journalism-pittsburgh-
university-of-pittsburgh-
press-2019-pp-ix-276-isbn-
9780822945659-4500-
hardback/
36A5D8D96DF8B029307427FD9999D525.

Turing:1938:SLB

[Tur38] Alan Mathison Turing. Systems of logic based on ordinals: a dissertation. Ph.D. dissertation, Cambridge University, Cambridge, UK, 1938. 161–228

+ 8 pp. URL http://turing.
ecs.soton.ac.uk/browse.php/

B/15. Published by Hodgson & Son, London, UK.

Turkle:2007:EOT

[Tur07] Sherry Turkle, editor. Evocative objects: things we think with. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-20168-2 (hardcover). ix + 385 pp. LCCN BF175.5.T73 E96

2007. URL http://www.loc.
gov/catdir/toc/ecip0619/
2006027966.html.

Turkle:2008:FSO

[Tur08a] Sherry Turkle, editor. Falling for science: objects in mind. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-20172-0 (hardcover). xii + 318 pp. LCCN Q181.3 .F35 2008. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        ecip0718/2007020843.html.

Turkle:2008:IHD

[Tur08b] Sherry Turkle, editor. The inner history of devices. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-20176-3 (hardcover). x + 208 pp. LCCN T14.5 .I5643 2008. URL http://www.

                                        loc.gov/catdir/toc/
                                        ecip0810/2008005530.html.

Turchetti:2010:BRK

[Tur10] Simone Turchetti. Book review: Kristin C. Harper, Weather by the Numbers: The Genesis of Modern Meteorology. Cambridge, MA and London: MIT Press, 2008. Pp. xii + 308. ISBN 978-0-262-08378-2. £25.95 (hardback). British Journal for the History of Science, 43(2):314–315, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turchetti:2012:PAC

[Tur12a] Simone Turchetti. The Pontecorvo affair: a cold war defection and nuclear physics. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-81664-8 (hardcover). 292 pp. LCCN QC774.P66 T8713 2012.

Turner:2012:BRC

[Tur12b] David M. Turner. Book review: C. F. Goodey, A History of Intelligence and ‘Intellectual Disability’: The Shaping of Psychology in Early Modern Europe. Farnham: Ashgate, 2011. Pp. x + 381. ISBN 978-1-4094-2021-7. £35.00 (hardback). British Journal for the History of Science, 45(2):285– 286, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turkel:2013:SDH

[Tur13] William J. Turkel. Spark from the deep: how shocking experiments with strongly electric fish powered scientific discovery. Animals, history, culture. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0981-X (hardcover), 1-4214-0994-1 (ebook). xi + 287 pp. LCCN QL639.1 .T87 2013.

Turchetti:2014:BRM

[Tur14] Simone Turchetti. Book reviews: Michael D. Gordin, The Pseudo-science Wars: Immanuel Velikovsky and the Birth of the Modern Fringe. Chicago and London: The University of Chicago Press, 2012. Pp. x + 291. ISBN 978-0-226-30442-7. £18.50 (hardback). British Journal for the History of Science, 47(2):386– 387, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turchetti:2015:BRC

[Tur15] Simone Turchetti. Book reviews:

Chasing Sound: Technology, Culture and the Art of Studio Recording from Edison to the LP. Baltimore: Johns Hopkins University Press, 2013. Pp. x + 292. ISBN 978-1-4214-1022-7. £29.00 (hardback). Susan Schmidt Horning. British Journal for the History of Science, 48(1):187–188, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turbil:2016:BRD

[Tur16a] Cristiano Turbil. Book reviews: David Gillott, Samuel Butler against the Professionals: Rethinking Lamarckism 1860–1900. London: Legenda, 2015. Pp. x + 198. ISBN 978-1-909662-25-4. £55.00 (hardback). British Journal for the History of Science, 49(2):300–301, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turchetti:2016:BRS

[Tur16b] Simone Turchetti. Book reviews: Sarah Bridger, Scientists at War: The Ethics of Cold War Weapons Research. Cambridge, MA: Harvard University Press, 2015. Pp. x + 350. ISBN 978-0-674-73682-5. £33.95 (hardback). British Journal for the History of Science, 49(2):314–315, June 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Talairach-Vielmas:2011:SNP

[TV11] Laurence Talairach-Vielmas, editor. Science in the nursery: the popularisation of science in Britain and France, 1761–1901. Cambridge Scholars Pub., Newcastle upon Tyne, UK, 2011. ISBN 1-4438-2680-4 (hardcover). xii + 309 pp. LCCN PN55 .S365 2011.

Trischler:2010:PPR

[TW10] Helmuth Trischler and Mark Walker, editors. Physics and politics: research and research support in twentieth century Germany in international perspective, volume Bd. 5 of Beiträge zur Geschichte der Deutschen Forschungsgemeinschaft. Franz Steiner Verlag, Stuttgart, Germany, 2010. ISBN 3-515-09601-

9. 285 pp. LCCN Q180.G4 P49 2010.

Tyson:2017:BRE

[Tys17] Charlie Tyson. Book reviews:

Eric Dietrich, Excellent Beauty: The Naturalness of Religion and the Unnaturalness of the World. New York: Columbia University Press, 2015. Pp. 208. ISBN: 978-0-231-17102-1. £20.50 (hardback). Roger Wagner and Andrew Briggs, The Penultimate Curiosity: How Science Swims in the Slipstream of Ultimate Questions. Oxford: Oxford University Press, 2016. Pp. xiii + 496. ISBN 978-0-19-874795-

6. £25.00 (hardback). British Journal for the History of Science, 50(1):160–162, 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
0923DFD9BA6D8A63B33CD159EBE0572F.

Uekotter:2010:WFW

[Uek10] Frank Uekötter. Die Wahrheit ist auf dem Feld: eine Wissensgeschichte der deutschen Landwirtschaft. (German) [The truth is in the field: a history of knowledge of German agriculture], volume 1 of Umwelt und Gesellschaft. Vandenhoeck and Ruprecht, Gottingen, Germany, third edition, 2010. ISBN 3-525-31705-0 (hardcover). 524 pp. LCCN S465 U35 2010. EUR49.99.

Underwood:2012:BRA

[Und12] Martin Underwood. Book review: Andrew Brown, Keeper of the Nuclear Conscience: The Life and Work of Joseph Rotblat. Oxford: Oxford University Press, 2012. Pp. xvii + 347. ISBN 978-0-19-958658-5. £18.99 (hardback). British Journal for the History of Science, 45(3):473–474, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Underwood:2013:BRR

[Und13] Martin Underwood. Book reviews: Ray Monk, Inside the Centre: The Life of J. Robert Oppenheimer. London: Jonathan Cape. 2012. Pp. xiv + 818. ISBN 978-0-224-06262-6. £30.00 (hardback). British Journal for the History of Science, 46(1):172– 173, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Valente:2010:GWL

[Val10] K. G. Valente. Giving wings to logic: Mary Everest Boole’s propagation and fulfilment of a legacy. British Journal for the History of Science, 43(1):49–74, March 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Valleriani:2011:BRJ

[Val11] Matteo Valleriani. Book review: J. L. Heilbron, Galileo. Oxford: Oxford University Press, 2010. Pp. xiv + 508. ISBN 978-0-19-958352-2. £20.00 (hardback). British Journal for the History of Science, 44(4):590–591, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Valderrama:2014:BRD

[Val14] Lorena B. Valderrama. Book reviews: Deborah R. Coen, The Earthquake Observers: Disaster Science from Lisbon to Richter. Chicago and London: The University of Chicago Press, 2013. Pp. viii + 348. ISBN 978-0-226-11181-0. £22.50 (hardback). British Journal for the History of Science, 47(3):576–577, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Darwin:2009:CDS

[Van09] John Van Wyhe, editor. Charles Darwin’s shorter publications, 1829–1883. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88809-3. xxvi + 529 pp. LCCN QH365.Z9 V36

2009. URL http://www.loc.
gov/catdir/enhancements/
fy1001/2009286120-b.html;
http://www.loc.gov/catdir/
enhancements/fy1001/

                                        2009286120-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1001/
                                        2009286120-t.html.

Rooij:2011:KMD

[Van11a] Arjan Van Rooij. Knowledge, money and data: an integrated account of the evolution of eight types of laboratory. British Journal for the History of Science, 44 (3):427–448, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VanWyhe:2011:BRA

[Van11b] John Van Wyhe. Book reviews: Andrew Pattison, The Darwins of Shrewsbury. Stroud: The History Press, 2009. Pp. 127. ISBN 978-0-7524-4867-1. £14.99 (paperback). Mike Dixon and Gregory Radick, Darwin in Ilkley. Stroud: The History Press, 2009. Pp. 126. ISBN 978-0-7524-5283-8. £12.99 (paperback). British Journal for the History of Science, 44(3): 467–468, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VandenBroecke:2011:BRA

[Van11c] Steven Vanden Broecke. Book review: André Goddu, Copernicus and the Aristotelian Tradition: Education, Reading and Philosophy in Copernicus’s Path to Heliocentrism. Leiden: Brill, 2010. Pp. xxvii + 545. ISBN 978-90-04-18107-6. EUR129.00 (hardback). British Journal for the History of Science, 44(4):587–588, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VandenBroecke:2013:CQP

[Van13] Steven Vanden Broecke. The Copernican question: Prognostication, skepticism, and celestial order. British Journal for the History of Science, 46(1):151– 159, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VandenBroecke:2014:BRM

[Van14] Steven Vanden Broecke. Book reviews: Mary Quinlan-McGrath, Influences: Art, Optics, and Astrology in the Italian Renaissance. Chicago and London: The University of Chicago Press, 2013. Pp. xi + 284. ISBN 978-0-226-92284-3. $35.00 (hardback). British Journal for the History of Science, 47(2):374–375, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vidal:2011:SSE

[VB11]  Fernando  Vidal

and Saskia Brown. The sciences of the soul: the early modern origins of psychology. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-85586-4. ???? pp. LCCN BF101 .V5313 2011.

vanBerkel:2013:IBM

[vB13] Klaas van Berkel. Isaac Beeckman on matter and motion: mechanical philosophy in the making. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0936-4 (paperback), 1-4214-0961-5 (electronic). viii + 265 pp. LCCN Q125.2 .B4713 2013.

Vertesi:2010:IIV

[Ver10]  Janet Vertesi.  Instrumental

images: the visual rhetoric of self-presentation in Hevelius’s machina coelestis. British Journal for the History of Science, 43(2):209–243, June 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vermeir:2012:OVS

[Ver12] Koen Vermeir. Openness versus secrecy? Historical and historiographical remarks. British Journal for the History of Science, 45(2):165–188, June 2012. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Verschueren:2015:BRA

[Ver15a] Pierre Verschueren. Book reviews: Anne-Lise Rey (ed.), Méthode et histoire: Quelle histoire font les historiens des sciences et des techniques? Paris: Classiques Garnier, 2013. Pp. 513. ISBN 987-2-8124-1419-0. £39.55 (paperback). British Journal for the History of Science, 48(3):504–505, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vertesi:2015:SLR

[Ver15b] Janet Vertesi. Seeing Like a Rover: how robots, teams, and images craft knowledge of Mars. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 1-336-19005-1 (e-book), 0-226-15601-X, 0-226-15596-X (hardcover). xi + 318 pp. LCCN TL799.M3 V47 2015eb. URL http://www.

myilibrary.com?id=750291.

Verschueren:2016:BRP

[Ver16] Pierre Verschueren. Book reviews: Pieter Dhondt (ed.), University Jubilees and University Writing: A Challenging Relationship. Leiden: Brill, 2015. Pp. xi + 255. ISBN 978-9-0042-1696-9. £79.00 (hardback). British Journal for the History of Science, 49 (1):151–152, March 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Verschueren:2019:CMH

[Ver19] Pierre Verschueren. Cécile Morette and the Les Houches summer school for theoretical physics; or, how Girl Scouts, the 1944 Caen bombing and a marriage proposal helped rebuild French physics (1951–1972). British Journal for the History of Science, 52(4): 595–616, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/cecile-
morette-and-the-les-
houches-summer-school-for-
theoretical-physics-or-how-
girl-scouts-the-1944-caen-

                                        bombing-and-a-marriage-
                                        proposal-helped-rebuild-
                                        french-physics-19511972/2A382DBF00E68CB293406AD8AFE9EAAE.

Vicedo:2011:SNM

[Vic11] Marga Vicedo. The social nature of the mother’s tie to her child: John Bowlby’s theory of attachment in post-war America. British Journal for the History of Science, 44(3):401–426, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

vonLieven:2013:BRD

[vL13a] Alexandra von Lieven. Book reviews: Daniela Wuensch and Klaus P. Sommer (eds.), Ludwig Borchardt, Die altägyptische Zeitmessung. Reprint mit einer Einleitung von Daniela Wuensch: Was die alten Ägypter über Uhren und Zeitmessung wussten. Göttingen: Termessos Verlag, 2013. Pp. 189. ISBN 978-3938016-14-5. EUR36.80 (hardback). British Journal for the History of Science, 46(4):716– 718, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

vonLunen:2013:BRA

[vL13b] Alexander von Lünen. Book reviews: Andrew W. Appel (ed.), Alan Turing’s System of Logic: The Princeton Thesis. Princeton and Oxford: Princeton University Press, 2012. Pp. xv + 142. ISBN 978-0-691-15574-6. £16.95 (hardback). British Journal for the History of Science, 46(4):

728–729, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

vonLunen:2015:BRT

[vL15] Alexander von Lünen. Book reviews: Tanja Klemm, Bildphysiologie: Wahrnehmung und Körper in Mittelalter und Renaissance. Berlin: Akademie Verlag, 2013. Pp. xii + 324. ISBN 978-3-05-006478-9. EUR99.80 (hardback). British Journal for the History of Science, 48(3):507– 508, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

vonLunen:2016:BRO

[vL16] Alexander von Lünen. Book reviews: Orit Halpern, Beautiful Data: A History of Vision and Reason since 1945. Durham, NC: Duke University Press, 2015. Pp. ix + 342. ISBN 978-0-8223-5744-5. £17.99 (paperback). British Journal for the History of Science, 49(1):144–146, March 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vermeir:2012:SSI

[VM12] Koen Vermeir and Dániel Margócsy. States of secrecy: an introduction. British Journal for the History of Science, 45(2):153– 164, June 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Virdi:2018:PMT

[VM18] Jaipreet Virdi and Coreen McGuire. Phyllis M. Tookey Kerridge and the science of audiometric standardization in Britain. British Journal for the History of Science, 51(1):123–146, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/phyllis-m-
tookey-kerridge-and-the-
science-of-audiometric-
standardization-in-britain/
2A8EDAC6A696AA0821E29DBF4B7C4C4D.

vanNeste:2019:BRB

[vN19] Aaron van Neste. Book review: Ben A. Minteer, Jane Maienschein and James P. Collins (eds.), The Ark and Beyond: The Evolution of Zoo and Aquarium Conservation. Chicago and London: The University of Chicago Press, 2018. Pp. xiv + 454. ISBN 978-0-226-538446-4. $35.00 (paperback). British Journal for the History of Science, 52(1):180– 181, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/ben-a-minteer-jane-
maienschein-and-james-p-
collins-eds-the-ark-and-
beyond-the-evolution-of-
zoo-and-aquarium-
conservation-chicago-and-

london-the-university-of-
chicago-press-2018-pp-xiv-454-
isbn-97802265384464-3500-paperback/
D0AA7FA6D8773D739A58A44672A94691.

Vogel:2010:LCH

[Vog10] Jakob Vogel. Locality and circulation in the Habsburg Empire: disputing the Carlsbad medical salt, 1763–1784. British Journal for the History of Science, 43(4): 589–606, December 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Volpe:2008:STD

[Vol08] Tony Volpe. Science et theologie dans les debats savants de la seconde moitié du XVIIe siecle: la genèse dans les Philosophical transactions et le Journal des savants (1665–1710). (French) [Science and theology in the scientific debates of the second half of the XVII-th Century: the genesis of the Philosophical transactions and the Journal des savants (1665–1710)]. Bibliothèque de l’École des hautes études, section des sciences religieuses. Brepols, Turnhout, Belgium, 2008. ISBN 2-503-52584-9. 467 pp. LCCN BS651 .V65 2008.

Volmar:2016:BRK

[Vol16] Daniel Volmar. Book reviews:

Katherine C. Epstein, Torpedo: Inventing the Military–Industrial Complex in the United States and Great Britain. Cambridge, MA and London: Harvard University Press, 2014. Pp. 305. ISBN 978-0-6747-2526-3. £30.00 (hardback). British Journal for the History of Science, 49(1):135– 136, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VonHumboldt:2011:PEI

[Von11] Alexander Von Humboldt. Political Essay on the Island of Cuba. Alexander von Humboldt in English. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-46568-3 (e-book), 0-226-46567-5 (hardcover). vi + 519 pp. LCCN F1763 .H913 2011. Edited with an introduction by Vera M. Kutzinski and Ottmar Ette. Translated by J. Bradford Anderson, Vera M. Kutzinski, and Anja Becker. With annotations by Tobias Kraft, Anja Becker, and Giorleny D. Altamirano Rayo.

Voskuhl:2011:BRT

[Vos11] Adelheid Voskuhl. Book review: Theresa Levitt, The Shadow of Enlightenment: Optical and Political Transparency in France, 1789–1848. Oxford: Oxford University Press, 2009. Pp. 192. ISBN 978-0-19-954470-7. £39.95 (hardback). British Journal for the History of Science, 44(2): 293–294, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vandendriessche:2015:SEP

[VPW15] Joris Vandendriessche, Evert Peeters, and Kaat Wils, editors. Scientists’ expertise as performance: between state and society, 1860–1960, volume 6 of History and Philosophy of Technoscience. Pickering and Chatto, Basingstoke, UK, 2015. ISBN 1-84893-527-7 (hardback), 1-317-31722-X (ePub ebook), 1-78144-693-8 (ePub ebook), 1-78144-692-X (PDF ebook), 1-317-31723-8 (PDF ebook), 1-317-31721-1 (Mobipocket ebook). xv + 239 pp. LCCN Q175.5 .S3825 2015.

URL http://swbplus.bsz-bw.
de/bsz427739551inh.htm.

vonTschermak-Seysenegg:2011:MDC

[vTSŠ^{+}11] Armin  von  Tschermak-

Seysenegg, Michal Šimêunek, et al. The Mendelian dioskuri: correspondence of Armin with Erich von Tschermak-Seysenegg, 1898–1951, volume 27 of Studies in the History of Sciences and Humanities. Centre for Sciences and Humanities of the Institute of Contemporary History of the Academy of Sciences of the Czech Republic, Prague, Czech Republic, 2011. ISBN 80-87378-67-9. 259 + 12 pp. LCCN QH429.2 .T73 2011; QP26 .T77 2011.

Vaupel:2010:DMI

[VW10] Elisabeth Vaupel and Stefan L.

Wolff, editors. Das Deutsche Museum im Nationalsozialismus: eine Bestandsaufnahme. (German) [The German Museum under National Socialism: an inventory], volume 27 of Deutsches Museum. Abhandlungen und Berichte. Wallstein, Göttingen, Germany, 2010. ISBN 3-8353-0596-4. 710 pp. LCCN T183.G32 M853 2010; AM.

Waddington:2011:BRS

[Wad11] Keir Waddington. Book review: Samuel J. M. M. Alberti, Morbid Curiosities: Medical Museums in Nineteenth-Century Britain. Oxford: Oxford University Press, 2011. Pp. xv + 238. ISBN 978-0-19-958458-1. £55.00 (hardback). British Journal for the History of Science, 44(4):603–604, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Waddington:2019:BRT

[Wad19] Keir Waddington. Book review: Tiago Saraiva, Fascist Pigs: Technoscientific Organisms and the History of Fascism. Cambridge, MA and London: MIT Press, 2016. Pp 326. ISBN 978-0-2620-3503-

2. $40.00 (hardcover). British Journal for the History of Science, 52(2):376–378, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/tiago-
                                        saraiva-fascist-pigs-
                                        technoscientific-organisms-
                                        and-the-history-of-fascism-
                                        cambridge-ma-and-london-
                                        mit-press-2016-pp-326-isbn-
                                        9780262035032-4000-
                                        hardcover/64525A00977797F7A763CAD4A00C08B9.

Wagner:2014:BRP

[Wag14] Darren N. Wagner. Book review: Piers Mitchell (ed.), Anatomical Dissection in Enlightenment England and Beyond: Autopsy, Pathology and Display. Farnham: Ashgate, 2012. Pp. xi + 186. ISBN 978-1-409-41886-3. £60.00 (hardback). British Journal for the History of Science, 47(1):184– 185, March 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wagner:2015:BRH

[Wag15] Darren N. Wagner. Book reviews: Helen King, The One-Sex Body on Trial: The Classical and Early Modern Evidence. Farnham: Ashgate, 2013. Pp. xii + 274. ISBN 978-1-4094-6335-1. £65.00 (hardback). British Journal for the History of Science, 48(3):505–507, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wainman:2018:BRR

[Wai18] Ruth Wainman. Book review:

Rebecca Onion, Innocent Experiments: Childhood and the Culture of Popular Science in the United States. Chapel Hill: University of North Carolina Press, 2016. Pp. 226. ISBN 978-1-4696-2947-6. $29.95 (paperback). British Journal for the History of Science, 51(1):170–171, 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/rebecca-
onion-innocent-experiments-
childhood-and-the-culture-

                                        of-popular-science-in-the-
                                        united-states-chapel-hill-
                                        university-of-north-
                                        carolina-press-2016-pp-226-
                                        isbn-9781469629476-2995-
                                        paperback/811667CB4ED244B85ED50E9B83213135.

Walls:2009:PCA

[Wal09] Laura Dassow Walls. The passage to Cosmos: Alexander von Humboldt and the shaping of America. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-87182-7. xv + 404 pp. LCCN Q143.H9 D388 2009.

Walker:2011:BRM

[Wal11a] Mark Walker. Book review: Matthias Berg, Jens Thiel and Peter Th. Walther (eds.), Mit Feder und Schwert: Militär und Wissenschaft-Wissenschaftler und Krieg. Stuttgart: Franz Steiner Verlag, 2009. Pp. 380. ISBN 978-3-515-09606-5. EUR46.00 (paperback). British Journal for the History of Science, 44(1):149–151, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Walton:2011:BRA

[Wal11b] Steven A. Walton. Book review: Anthony Gerbino and Stephen Johnston, with a contribution by Gordon Higgott, Compass and Rule: Architecture as Mathematical Practice in England, 1500–1750. New Haven and London: Yale University Press, 2009. Pp. 207. ISBN 978-0-300-15093-3. £30.00 (hardback). British Journal for the History of Science, 44 (2):287–289, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Walker:2012:BDU

[Wal12a] John R. Walker. Britain and disarmament: the UK and nuclear, biological and chemical weapons arms control and programmes, 1956–1975. Ashgate, Aldershot, Hants, England and Farnham, England, 2012. ISBN 1-4094-3580-6 (hardcover), 1-4094-3581-4 (e-book). xv + 305 pp. LCCN JZ6009.G7 W35 2012.

Walker:2012:NIT

[Wal12b] Mark Walker. The ‘national’ in international and transnational science. British Journal for the History of Science, 45(3):359– 376, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wallis:2012:BRB

[Wal12c] Max K. Wallis. Book reviews:

Brian Wynne, Rationality and Ritual: Participation and Exclusion in Nuclear Decision-Making. London and Washington, DC: Earthscan, 2011. Pp. xxvii + 228. ISBN 978-1-84971-161-6. £24.99 (paperback). British Journal for the History of Science, 45(4): 708–709, 2012. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Walsh:2013:SPR

[Wal13]  Lynda Walsh.  Scientists as

prophets: a rhetorical genealogy.

Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-985709-1 (hardcover), 0-19-985711-3 (paperback). x + 264 + 1 pp. LCCN Q175.5 .W35 2013.

Walsh:2015:BRI

[Wal15] Sarah Walsh. Book reviews: An Image of God: The Catholic Struggle with Eugenics. London: The University of Chicago Press, 2013. Pp. ix + 226. ISBN 978-0-2260-3898-8. £31.50 (hardback). Sharon M. Leon. British Journal for the History of Science, 48 (1):185–187, March 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wale:2016:BRM

[Wal16a] Matthew Wale. Book reviews: Melinda Baldwin, Making Nature: The History of a Scientific Journal. Chicago: The University of Chicago Press, 2015. Pp. 328. ISBN 978-0-226-26145-4. $45.00 (hardback). British Journal for the History of Science, 49(2):301– 302, June 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wallis:2016:BRR

[Wal16b] Jennifer Wallis. Book reviews: Rina Knoeff and Robert Zwijnenberg (eds.), The Fate of Anatomical Collections. Farnham: Ashgate, 2015. Pp. 336. ISBN 978-1-4094-6815-8. £67.50 (hardback). British Journal for the History of Science, 49(1):127–129, March 2016. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Wallis:2017:BRJ

[Wal17]  Jennifer Wallis.  Book re-

view: Jonathan Sadowsky, Electroconvulsive Therapy in America: The Anatomy of a Medical Controversy. Abingdon and New York: Routledge, 2017. Pp. 172. ISBN 978-1-138-69696-9. £110.00 (hardcover). British Journal for the History of Science, 50(4): 738–740, December 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
B070E339816A8061C701CB2E648210EE.

Wale:2018:BRH

[Wal18]  Matthew Wale.  Book re-

view: Henry A. McGhie, Henry Dresser and Victorian Ornithology: Birds, Books and Business. Manchester: Manchester University Press, 2017. Pp. xxiv + 341. ISBN 978-1-7849-9413-6. £25.00 (hardcover). British Journal for the History of Science, 51(4): 709–710, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/henry-a-
mcghie-henry-dresser-and-
victorian-ornithology-
birds-books-and-business-
manchester-manchester-
university-press-2017-pp-
xxiv-341-isbn-

9781784994136-2500-hardcover/
6A7F4DC9D7B7E2037A0294A287BC1790.

Wale:2019:BRR

[Wal19a] Matthew Wale. Book review: Ruth Barton, The X Club: Power and Authority in Victorian Science. Chicago: The University of Chicago Press, 2018. Pp. 576. ISBN 978-0-2265-5161-6. £41.50/ $60 (hardback). British Journal for the History of Science, 52(3): 529–530, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ruth-
barton-the-x-club-power-
and-authority-in-victorian-
science-chicago-the-
university-of-chicago-
press-2018-pp-576-isbn-
9780226551616-415060-
hardback/
61A6AC69D48ACE7487431738C8686791.

Wale:2019:EEN

[Wal19b] Matthew Wale. Editing entomology: natural-history periodicals and the shaping of scientific communities in nineteenthcentury Britain. British Journal for the History of Science, 52(3): 405–423, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/editing-
entomology-naturalhistory-
periodicals-and-the-

shaping-of-scientific-
communities-in-
nineteenthcentury-britain/
770FC9722EBA5064EF80214E1A1EB364.

Wardhaugh:2010:HRH

[War10]  Benjamin  Wardhaugh.

How to read historical mathematics. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14014-6 (hardcover). xii + 116 pp. LCCN QA21 .W324 2010.

Wardhaugh:2012:HHM

[War12] Benjamin Wardhaugh, editor.

The history of the history of mathematics: case studies for the Seventeenth, Eighteenth, and Nineteenth Centuries. P. Lang, Oxford, UK, 2012. ISBN 3-0343-0708-X (paperback). vi + 187 pp. LCCN QA21 .H59 2012.

Waring:2013:BRR

[War13] Sophie Waring. Book reviews: Robert Fox, The Savant and the State: Science and Cultural Politics in Nineteenth-Century France. Baltimore: Johns Hopkins University Press, 2012. Pp. viii + 394. ISBN 978-1-4214-0522-3. £31.50 (hardback). British Journal for the History of Science, 46(4): 720–722, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Warkentin:2019:BRA

[War19] Germaine Warkentin. Book review: Adriana Craciun and Mary Terrall (eds.), Curious Encounters: Voyaging, Collecting and Making Knowledge in the Long Eighteenth Century. Toronto: University of Toronto Press, 2019. Pp. ix + 244. ISBN 978-1-4875-0367-

3. $75.00 (cloth). British Journal for the History of Science, 52 (4):713–714, December 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/adriana-
craciun-and-mary-terrall-
eds-curious-encounters-
voyaging-collecting-and-
making-knowledge-in-the-
long-eighteenth-century-
toronto-university-of-
toronto-press-2019-pp-ix-244-
isbn-9781487503673-7500-cloth/
96FC0285A19BA164D29B3CFB83073DB8.

Watts:2014:WWN

[Wat14] Iain P. Watts. ‘We want no authors’: William Nicholson and the contested role of the scientific journal in Britain, 1797– 1813. British Journal for the History of Science, 47(3):397–419, September 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wazeck:2014:EOP

[Waz14] Milena Wazeck. Einstein’s opponents: the public controversy about the Theory of Relativity in the 1920s. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-01744-0 (hardcover). xxi + 355 pp. LCCN QC173.585 .W3913 2014. URL http://

assets.cambridge.org/
97811070/17443/
cover/9781107017443.jpg. En-

glish translation by Geoffrey S. Koby.

Wallace:2013:OLC

[WC13]  Alfred  Russel  Wallace

and James T. Costa. On the organic law of change: a facsimile edition and annotated transcription of Alfred Russel Wallace’s Species notebook of 1855–1859. Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-72488-7 (hardcover). xii + 573 pp. LCCN QH375 .W35 2013.

URL http://www.jstor.org/
stable/10.2307/j.ctt6wpn8h.

Weart:2012:RNF

[Wea12] Spencer R. Weart. The rise of nuclear fear. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-05233-1. ix + 367 pp. LCCN QC773 .W44 2012.

Webster:2008:PMM

[Web08] Charles Webster. Paracelsus:

medicine, magic and mission at the end of time. Yale University Press, New Haven, CT, USA, 2008. ISBN 0-300-13911-X. xiv + 326 pp. LCCN R147.P2 W33

2008. URL http://www.loc.
gov/catdir/toc/fy0904/
2008027973.html.

Weber:2010:BRS

[Web10] Thomas P. Weber. Book reviews:

Sander Gliboff, H. G. Bronn, Ernst Haeckel, and the Origins of German Darwinism: A Study in Translation and Transformation. Cambridge, MA and London: MIT Press, 2008. Pp. xii + 259. ISBN 978-0-262-07293-9. £25.95 (hardback). Robert J. Richards, The Tragic Sense of Life: Ernst Haeckel and the Struggle over Evolutionary Thought. Chicago: The University of Chicago Press, 2008. Pp. xx + 551. ISBN 978-0-226-71214-7. £27.00 (hardback). British Journal for the History of Science, 43(4):617–619, December 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Gli08, Ric08].

Weber:2011:BRJ

[Web11] Thomas P. Weber. Book review: Janina Wellmann, Die Form des Werdens. Eine Kulturgeschichte der Embryologie, 1760–1830. Göttingen: Wallstein, 2010. Pp. 429. ISBN 978-3-8353-0594-6. EUR39.00 (paperback). British Journal for the History of Science, 44(4):601–603, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weber:2013:BRF

[Web13] Thomas P. Weber. Book reviews: Frank Uekötter, Die Wahrheit ist auf dem Feld: Eine Wissensgeschichte der deutschen Landwirtschaft. Göttingen: Vandenhoeck & Ruprecht, 2010. Pp. 524. ISBN 978-3-525-31705-1. EUR49.95 (hardback). British Journal for the History of Science, 46(1):170–172, March 2013.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weber:2014:BRA

[Web14] Thomas P. Weber. Book reviews: Anna Echterhölter, Schattengefechte: Genealogische Praktiken in Nachrufen auf Naturwissenschaftler (1710–1860). Göttingen: Wallstein, 2012. Pp. 365. ISBN 978-3-8353-1071-1. EUR 39.90 (paperback). Denise Phillips, Acolytes of Nature: Defining Natural Science in Germany 1770–1850. Chicago and London: The University of Chicago Press, 2012. Pp.

356. ISBN 0-226-66737-9. £29.99 (hardback). British Journal for the History of Science, 47(4): 734–735, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weber:2015:BRN

[Web15] Thomas P. Weber. Book reviews:

Nicolas Rasmussen, Gene Jockeys: Life Science and the Rise of Biotech Enterprise. Baltimore: Johns Hopkins University Press, 2014. Pp. 264. ISBN 978-1-4214-1340-2. $35.00 (hardback). Hallam Stevens, Life out of Sequence: A Data-Driven History of Bioinformatics. Chicago: The University of Chicago Press, 2013. Pp. 304. ISBN 978-0-226-08020-8. $30.00 (paperback). British Journal for the History of Science, 48 (4):717–719, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weber:2017:BRK

[Web17] Thomas P. Weber. Book review: Kelly Joan Whitmer, The Halle Orphanage as Scientific Community: Observation, Eclecticism, and Pietism in the Early Enlightenment. Chicago and London: The University of Chicago Press, 2015. Pp. 200. ISBN 978-0-226-24377-1. £28.00 (hardback). British Journal for the History of Science, 50(1):146–148, 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
36C33BC395FDD914BD1C06E453BF1199.

Weber:2018:BRC

[Web18a] Thomas P. Weber. Book review: Courtney Fullilove, The Profit of the Earth: The Global Seeds of American Agriculture. Chicago: The University of Chicago Press, 2017. Pp. 288. ISBN 978-0-226-45486-3. $40.00 (hardback). British Journal for the History of Science, 51 (2):311–313, June 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/courtney-
fullilove-the-profit-of-
the-earth-the-global-seeds-
of-american-agriculture-
chicago-the-university-of-
chicago-press-2017-pp-288-
isbn-9780226454863-4000-
hardback/
2531994E046F4AFF3105BEDA306ED0F9.

Weber:2018:BRL

[Web18b] Thomas P. Weber. Book review: Laura J. Miller, Building Nature’s Market: The Business and Politics of Natural Food. Chicago: The University of Chicago Press, 2017. Pp. 288. ISBN 978-0-22650-123-9. $105.00 (cloth cover). British Journal for the History of Science, 51(4): 714–716, December 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/laura-j-
miller-building-natures-
market-the-business-and-
politics-of-natural-food-
chicago-the-university-of-
chicago-press-2017-pp-288-
isbn-9780226501239-10500-
cloth-cover/
BA8AB1D5DD710F41C0991149CDBDE0EE.

Weeks:2019:FBD

[Wee19] S. V. Weeks. Francis Bacon’s doctrine of idols: a diagnosis of ‘universal madness’. British Journal for the History of Science, 52(1):1–39, 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/francis-
bacons-doctrine-of-idols-a-
diagnosis-of-universal-
madness/
B6CE0267A4E9E2EE52C25F067B9F4214.

Wellmann:2017:AET

[Wel17] Janina Wellmann. Animating embryos: the in toto representation of life. British Journal for the History of Science, 50(3): 521–535, September 2017. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        8D6B8231E80E7A5004ED53C63986822F.

Werrett:2010:FPA

[Wer10] Simon Werrett. Fireworks: pyrotechnic arts and sciences in European history. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-89377-4 (hardcover). 359 + 16 pp. LCCN TP268 .W47 2010.

Werrett:2013:REM

[Wer13]  Simon  Werrett.

Recycling in early modern science. British Journal for the History of Science, 46(4):627–646, December 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Westwick:2012:BSM

[Wes12a] Peter J. Westwick, editor. Blue sky metropolis: the aerospace century in Southern California, volume 4 of Western histories. Published for Huntington-USC Institute on California and the West by University of California Press and Huntington Library, San Marino, California, Berkeley, CA, USA, 2012. ISBN 0-87328-249-3 (hardcover). xii + 308 pp. LCCN HD9711.5.U63 C23 2012.

Westwick:2012:BRD

[Wes12b] Peter J. Westwick. Book review: David C. Cassidy, A Short History of Physics in the American Century. Cambridge, MA and London: Harvard University Press, 2011. Pp. 211. ISBN 978-0-674-04936-9. £22.95 (hardback). British Journal for the History of Science, 45(3):472–473, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weston:2015:BRJ

[Wes15] Janet Weston. Book reviews:

Jonathan Toms, Mental Hygiene and Psychiatry in Modern Britain. Basingstoke: Palgrave Macmillan, 2013. Pp. xii + 262. ISBN 978-1-137-32156-5. £55.00 (hardback). British Journal for the History of Science, 48(3): 526–527, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

White:2015:BRS

[Whi15] Alice White. Book reviews: The Science of Deception: Psychology and Commerce in America. Chicago and London: The University of Chicago Press, 2013. Pp. 228. ISBN 978-0-226-92374-1. £35.00/$50.00 (hardback). Michael Pettit. British Journal for the History of Science, 48 (1):184–185, March 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

White:2016:BRJ

[Whi16] Alice White. Book reviews: Jamie Cohen-Cole, The Open Mind: Cold War Politics and the Sciences of Human Nature. Chicago: The University of Chicago Press, 2014. Pp. 397. ISBN 978-0-2260-9216-4. £31.50 (hardback). Mark Solovey, Shaky Foundations: The Politics– Patronage–Social Science Nexus in Cold War America. New Brunswick, NJ: Rutgers University Press, 2013. Pp. x + 253. ISBN 978-0-8135-5465-5. £39.95 (hardback). British Journal for the History of Science, 49(1):138– 140, March 2016. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wigelsworth:2010:SSA

[Wig10] Jeffrey R. Wigelsworth. Selling science in the age of Newton: advertising and the commoditization of knowledge. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hants, England and Farnham, England, 2010. ISBN 1-4094-0075-1 (hardcover). xii + 203 pp. LCCN Q180.55.E25 W54 2010.

Wigelsworth:2011:BRP

[Wig11] Jeffrey R. Wigelsworth. Book review: Paul A. Elliott, Enlightenment, Modernity and Science: Geographies of Scientific Culture and Improvement in Georgian England. London and New York: I. B. Tauris, 2010. Pp. xii + 363. ISBN 978-1-84885-366-9. £65.00 (hardback). British Journal for the History of Science, 44(3): 465–467, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilding:2011:BRM

[Wil11a] Nick Wilding. Book review:

Marco Beretta, Antonio Clericuzio and Lawrence M. Principe (eds.), The Accademia del Cimento and Its European Context. Sagamore Beach: Science History Publications, 2009. Pp. xiii + 257. ISBN 978-0-88135-387-7. $49.95 (hardback). British Journal for the History of Science, 44 (4):592–593, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Williams:2011:BRK

[Wil11b] Alison Williams. Book review:

Kirk D. Read, Birthing Bodies in Early Modern France: Stories of Gender and Reproduction. Farnham: Ashgate, 2011. Pp. xiii + 205. ISBN 978-0-7546-6632-5. £55.00 (hardback). British Journal for the History of Science, 44 (4):584–585, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilson:2011:TCS

[Wil11c] Duncan Wilson. Tissue culture in science and society: the public life of a biological technique in twentieth century Britain. Science, technology, and medicine in modern history. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2011. ISBN 0-230-28427-2 (hardback). x + 183 pp. LCCN QH585.2 .W55 2011.

Wilson:2012:BRS

[Wil12] Duncan Wilson. Book reviews: Sally Smith Hughes, Genentech: The Beginnings of Biotech. Chicago and London: The University of Chicago Press, 2011. Pp. xv + 215. ISBN 978-0-226-35918-2. $25.00 (hardback). British Journal for the History of Science, 45(4):706–707, 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilson:2014:BRA

[Wil14] Lydia Wilson. Book reviews: Avner Ben-Zaken, Reading Hayy Ibn Yaqzan: A Cross-cultural History of Autodidacticism. Baltimore: Johns Hopkins University Press, 2011. Pp. xiv + 191. ISBN 978-0-8018-9739-9. £31.00 (hardback). British Journal for the History of Science, 47(2):371– 372, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wille:2015:IFE

[Wil15a] Sheila Wille. The ichneumon fly and the equilibration of British natural economies in the eighteenth century. British Journal for the History of Science, 48(4): 639–660, December 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Williamson:2015:WEM

[Wil15b] Fiona Williamson. Weathering the empire: meteorological research in the early British Straits settlements. British Journal for the History of Science, 48(3): 475–492, September 2015. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wills:2016:BRJ

[Wil16a] John Wills. Book review: Jeff Crane, The Environment in American History: Nature and the Formation of the United States. London: Routledge, 2014. Pp. 440. ISBN 978-0-415-80871-

2. £100.00 (hardback). British Journal for the History of Science, 49(4):503, December 2016. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
BE96D98AF5E9E3A45B8AF5551270037E.

Wilson:2016:BRD

[Wil16b] Daniel C. S. Wilson. Book review: Dan Bouk, How Our Days Became Numbered: Risk and the Rise of the Statistical Individual. Chicago: The University of Chicago Press, 2015. Pp. 304. ISBN 978-0-226-25917-8. £28.00 (hardback). British Journal for the History of Science, 49(4): 520–521, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
500A1B3C51100625E02D454EC9D8A716.

Winterburn:2011:BRM

[Win11] Emily Winterburn. Book review: Michael Hoskin, Discoverers of the Universe: William and Caroline Herschel. Princeton and Oxford: Princeton University Press, 2011. Pp. xvi + 237. £20.95 (hardback). British Journal for the History of Science, 44 (4):600–601, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Winter:2012:MFM

[Win12] Alison Winter. Memory: fragments of a modern history. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-90258-7 (hardcover). x + 319 pp. LCCN BF371 .W545 2012.

Winterburn:2013:BRJ

[Win13] Emily Winterburn. Book reviews: John Rowlinson, Sir James Dewar, 1842–1923: A Ruthless Chemist. Farnham: Ashgate, 2012. Pp. xviii + 236. ISBN 978-1-4094-0613-6. £65.00 (hardback). British Journal for the History of Science, 46(4):724– 725, December 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Winkler:2015:BRU

[Win15] Anita Winkler. Book reviews: Ute Frevertet al., Emotional Lexicons: Continuity and Change in the Vocabulary of Feeling 1700– 2000. Oxford: Oxford University Press, 2014. Pp. 288. ISBN 978-0-19-965573-1. £65.00 (hardback). British Journal for the History of Science, 48(3):509–511, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Winterbottom:2019:ECE

[Win19a] Anna Winterbottom. An experimental community: the East India Company in London, 1600– 1800. British Journal for the History of Science, 52(2):323– 343, 2019. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/an-experimental-
community-the-east-india-
company-in-london-16001800/
1CA4032E72E95A1E4FF1D7F71F8DDB0B.

Wintroub:2019:PIW

[Win19b] Michael Wintroub. The pharmakon of ‘If’: working with Steven Shapin’s A Social History of Truth. British Journal for the History of Science, 52(3): 487–514, September 2019. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/pharmakon-
of-if-working-with-steven-
shapins-a-social-history-
of-truth/
1CC8DDB64218A09A469BE0457DBA12CE.

Wisnioski:2012:ECC

[Wis12] Matthew H. Wisnioski. Engineers for change: competing visions of technology in 1960s America.

Engineering studies series. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01826-8 (hardcover), 0-262-30518-6. xvii + 286 pp. LCCN T14.5 .W5664 2012.

                                        URL http://site.ebrary.com/
                                        lib/yale/Doc?id=10617474.

Withers:2010:GSB

[Wit10] Charles W. J. Withers. Geography and science in Britain, 1831–1939: a study of the British Association for the Advancement of Science. Manchester University Press, Manchester, UK, 2010. ISBN 0-7190-7976-4 (hardcover). xvii + 278 pp. LCCN G141 WIT 2010.

Withey:2011:BRD

[Wit11a] Alun Withey. Book review: Daniel Schäfer, Old Age and Disease in Early Modern Medicine. London: Pickering & Chatto, 2011. Pp. viii + 287. ISBN 978-1-84893-020-9. £60.00 (hardback). British Journal for the History of Science, 44(4):581–582, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Withey:2011:BRO

[Wit11b] Alun Withey. Book review: Ole Peter Grell, Andrew Cunningham and Jon Arrizabalaga (eds.), Centres of Medical Excellence? Medical Travel and Education in Europe, 1500–1789. Farnham: Ashgate, 2010. Pp. xvi + 335. ISBN 978-0-7546-6699-8. £60.00 (hardback). British Journal for the History of Science, 44 (2):283–284, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wragge-Morley:2011:BRW

[WM11] Alexander Wragge-Morley. Book review: William Poole, The World Makers: Scientists of the Restoration and the Search for the Origins of the Earth. Oxford: Peter Lang, 2010. Pp. xviii + 234. ISBN 978-1-906165-08-6. £30.00 (hardback). British Journal for the History of Science, 44(2):289– 290, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Poo10].

Wragge-Morley:2018:BRF

[WM18a] Alexander Wragge-Morley. Book review: Florike Egmond, Eye for Detail: Images of Plants and Animals in Art and Science. London: Reaktion Books, 2017. Pp. 280. ISBN 978-1-7802-3640-7. £35.00 (hardback). British Journal for the History of Science, 51(3): 521–522, September 2018. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/florike-
egmond-eye-for-detail-
images-of-plants-and-
animals-in-art-and-science-
london-reaktion-books-2017-
pp-280-isbn-9781780236407-
3500-hardback/
52257B2952DC0A03CAEF5652B0270793.

Wragge-Morley:2018:RBR

[WM18b] Alexander Wragge-Morley. Robert Boyle and the representation of imperceptible entities. British Journal for the History of Science, 51(1): 17–40, 2018. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/robert-boyle-and-
the-representation-of-
imperceptible-entities/
EAF044425AB23E01295157165B299566.

Wolffram:2009:SSP

[Wol09] Heather Wolffram. The stepchildren of science: psychical research and parapsychology in Germany, c. 1870–1939, volume 88 of The Wellcome series in the history of medicine; Clio medica (Amsterdam, Netherlands). Editions Rodopi B.V., Amsterdam, The Netherlands, 2009. ISBN 90-420-2728-2, 90-420-2729-0. ISSN 0045-7183. 342 pp. LCCN BF1028.5.G3 W65 2009.

Wolf:2011:BRC

[Wol11] Jeff Wolf. Book review: Charles G. Gross, A Hole in the Head: More Tales in the History of Neuroscience. Cambridge, MA and London: MIT Press, 2009. Pp. x + 356. ISBN 978-0-262-01338-3. £25.95 (hardback). British Journal for the History of Science, 44 (1):146–147, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wolfe:2013:CSS

[Wol13] Audra J. Wolfe. Competing with the Soviets: science, technology, and the state in Cold War America. Johns Hopkins introductory series in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0769-8 (hardcover), 1-4214-0771-X (paperback). viii + 166 pp. LCCN Q127.U6 W65 2013.

Wolfe:2014:BRS

[Wol14]  Audra J. Wolfe.  Book re-

views: Sevan G. Terzian, Science Education and Citizenship: Fairs, Clubs, and Talent Searches for American Youth, 1918–1958. New York: Palgrave Macmillan, 2013. Pp. xv + 235. ISBN 978-1-137-03186-0. £55.00 (hardback). British Journal for the History of Science, 47(4):752–753, December 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wotton:2012:WGN

[Wot12] Roger S. Wotton. Walking with Gosse: natural history, creation and religious conflicts. Clio Publishing, Southampton, England, 2012. ISBN 0-9556983-9-1 (paperback). vii + 214 pp. LCCN QH31.G6 W688 2012.

Wright:2016:BRJ

[Wri16] Aaron Sidney Wright. Book review: Jeroen van Dongen (ed.), Cold War Science and the Transatlantic Circulation of Knowledge. Leiden: Brill, 2015. Pp. x + 293. ISBN 978-9-004-26421-2. EUR115.00 (paperback). British Journal for the History of Science, 49(4):514– 515, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
A0F540FF7F71B6746E4E6D85866C6323.

Wylie:2011:BRP

[Wyl11] Caitlin Donahue Wylie. Book review: Paul D. Brinkman, The Second Jurassic Dinosaur Rush: Museums and Paleontology in America at the Turn of the Twentieth Century. Chicago and London: University of Chicago Press, 2010. Pp. xiv + 345. ISBN 978-0-226-07472-6. $49.00 (hardback). British Journal for the History of Science, 44(3):469–470, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wynne:2011:RRP

[Wyn11] Brian Wynne. Rationality and ritual: participation and exclusion in nuclear decision-making. Science in society series. Earthscan, London, UK, 2011. ISBN 1-84971-162-3 (hardcover), 1-84971-161-5 (paperback). xxvii + 228 pp. LCCN HD9698.G72 W96 2011. With a foreword by Gordon MacKerron.

Yalcinkaya:2011:SAR

[Yal11] M. Alper Yalcinkaya. Science as an ally of religion: a Muslim appropriation of ‘the conflict thesis’.

British Journal for the History of Science, 44(2):161–181, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Yalcinkaya:2015:LPD

[Yal15] M. Alper Yalçinkaya. Learned patriots: debating science, state, and society in the nineteenthcentury Ottoman Empire. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-18420-X (hardcover), 0-226-18434-X (e-book). 314 pp. LCCN Q175.52.T9 Y35 2015.

Yalcinkaya:2017:BRM

[Yal17] M. Alper Yalcinkaya. Book review: Miri Shefer-Mossensohn, Science among the Ottomans: The Cultural Creation and Exchange of Knowledge. Austin: University of Texas Press, 2015. Pp. 262. ISBN 978-1-4773-0359-7. $55.00 (hardback). British Journal for the History of Science, 50(3):547–548, September 2017. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
A2E2479585754E18D8952E4071AF59B5.

Yamamoto:2015:MME

[Yam15] Koji Yamamoto. Medicine, metals and empire: the survival of a chymical projector in early eighteenth-century London. British Journal for the History of Science, 48(4):607–637, December 2015. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Yannielli:2013:YKM

[Yan13] Joseph L. Yannielli. A Yahgan for the killing: murder, memory and Charles Darwin. British Journal for the History of Science, 46(3): 415–443, September 2013. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Yeo:2014:NEV

[Yeo14] Richard R. Yeo. Notebooks, English virtuosi, and early modern science. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-10656-X (hardcover), 0-226-10673-X (e-book). xviii + 398 pp. LCCN Q127.G4 Y47 2014.

Yi:2015:RUG

[Yi15] Doogab Yi. The recombinant university: genetic engineering and the emergence of Stanford biotechnology. Synthesis. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-14383-X (hardcover), 0-226-21611-X (e-book). xi + 318 pp. LCCN QH442 .Y54 2015.

Yates:2010:BRG

[YM10] JoAnne Yates and Craig N. Murphy. Book review: Graeme Gooday and James Sumner (eds.), By Whose Standards? Standardization, Stability and Uniformity in the History of Information and Electrical Technologies. History of Technology, Volume 28. Series editor Ian Inkster. London: Continuum, 2008. Pp. xiv + 171. ISBN 978-0-8264-3875-1. £90.00 (hardback). British Journal for the History of Science, 43(3): 503–505, September 2010. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Zaidi:2016:BRJ

[Zai16]  Waqar Zaidi.  Book review:

Jonathan Hogg, British Nuclear Culture: Official and Unofficial Narratives in the Long Twentieth Century. London: Bloomsbury, 2016. Pp. 248. ISBN 978-1-4411-6976-1. £19.99 (paperback). British Journal for the History of Science, 49(4): 518–520, December 2016. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/product/
6468A99E99CE702CFB124EC40199D5B1.

Zallen:2014:PPL

[Zal14] Doris T. Zallen. The power of partnerships: the Liverpool school of butterfly and medical genetics. British Journal for the History of Science, 47(4): 677–699, December 2014. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Zipser:2013:SGM

[Zip13] Barbara Zipser, editor. Simon of Genoa’s medical lexicon. Versita Discipline: History, Archaeology. Versita, Versita Limited, London, England, 2013. ISBN 83-7656-022-0 (hardcover), 83-7656-021-2 (paperback), 83-7656-023-9 (e-book). 164 pp. LCCN R141 .S56 2013. URL http://www.

degruyter.com/books/
9788376560236.

Zuidervaart:2015:MRW

[ZR15] Huib J. Zuidervaart and Marlise Rijks. ‘Most rare workmen’: optical practitioners in early seventeenth-century Delft. British Journal for the History of Science, 48(1):53–85, March 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).